A Complete Bibliography of Publications in SIAM Journal on Mathematical Analysis, 1970–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

31 May 2024 Version 3.18

Title word cross-reference

#11889 [Spe79].

(−1, 1) [LS93]. (−∞, ∞) [Pas74]. (m(t)x′(t))′ + A(t)x(t) = 0 [Ede79]. (p(x)u′(x))′ + g(x)u′(x) + qu(x) = f [Whi79]. (ϕ(y′)) = qf(t, y, y′), 0 < t < 1 [O’R93]. (r(t)ψ(x)x′)′ + a(t)f(x) = 0 [MR79a]. 0 < p(x) ∈ L_{∞}[a, b] [Whi79]. 0 ≤ x ≤ l [CL73]. 2 [AJV94, CW99]. 2m + 1 [Sho70]. 2π [FR91]. 2π/m [FR91]. 3 [KoM98]. 40 [GM94]. 6 [GB72, Gra72]. 6 − j [AW79]. 7 [Ura84]. [−1, 1] [Rah91]. [0, T] [Log83d]. _{0}F_{3} [Car70c]. _{1}F_{2} [FI75]. _{1}ψ_{1} [Kad87]. _{3}F_{2} [Büh87b, Sas91]. _{4}F_{3}(1) [WB95]. A [Ehl73, Aom94, Har80, Ise79]. aK(at), a → 0 [Log79b]. B^{∗} [Rei79]. BC_{1} [Ask82]. β [HT87]. BV (Ω) [AK99]. C^{1} [Coh89]. C^{α} [YL94]. C^{α}(\overline{Ω}) [XA91]. C^{p} [Rea86b]. C_{0} [Yao95]. C_{0}(T) [Wu90]. C_{ℓ} [Mil94]. A [FM99]. D [Har80]. D_{n} × O(2) [ML93]. D_{x}^{r}u(x, t) = D_{t}u(x, t) [Kem82a]. D_{X_{1}}^{k}U(X_{1}, ⋯, X_{r}) = D_{X_{k}}U(X_{1}, ⋯, X_{r})^{∗} [Kem85]. DA [Har80]. δ [Lan83b]. Δ^{2}u = λu [Cof82]. Δ_{p+2}u(x) + F (x)u(x) = 0 [CG71]. Δψ + K^{2}ψ = 0 [Kal75]. Δu + K(|x|)|u|^{p−1}u = 0 [Yan96]. Δu + |u|^{p−1}u − |u|^{q−1}u = 0 [Tro90]. E^{3} [Pie73, Pie74]. e^{iut} [dB76]. e^{x} [Bor85]. E_{2} [Lip72]. ϵẍ = f(t, x, ẋ) [Die86]. exp(−x^{4}) [Nev84]. F [PR71, Sil85]. f(A) [BR82]. F_{1} [Mil73a]. F_{3} [Sas91, Sin80]. F_{4} [Gar90, Iwa88, Car75b]. f ∈ C(Ω) [Whi80]. G [Mil74a, Fie73, Fie83]. g(x) ∈ L_{∞}[a, b] [Whi79]. G_{2} [Gar88, Gus90, Zei87]. H [Kel79a, Leg76, Jua95, Pes95]. H^{∞} [FF92, Gu92]. H^{p} [CR83]. H^{s}(R^{3}) [EV97]. H_{0}^{m} [Jen91]. H_{μ} [Lee74]. H_{p} [Log79a, McC83, BH85]. ∞ [KS86]. ∫_{0}^{∞} e^{−at}J_{μ}(bt)J_{ν}(ct)t^{μ−ν} dt [Ben75].

iU_{t} + Δ_{2}U − (\frac{α}{x_{1}^{2}} + \frac{β}{x_{2}^{2}} ) U = 0 [Boy76]. J [Chi77]. J_{ν}(νx) [Par84]. K [Amo89, CJ99, Rus77]. k = 1 [Kad88b]. K{M_{p}} [Swa72]. L^{1} [FM92, HW84, JW80b, MP77, Nor88b, Peh84b, Han77]. L_{1}^{1}(R) ∩ L_{N}^{1}(R^{+}) [CK87]. L^{2} [Alp93, FF92, Lon87, RW78, KS79a]. l^{2}(Z) [MR91]. L^{∞} [Fis75, Lon83]. L^{p} [DL84, LN98, Mei77, LW95]. L^{p}(R^{n}) [Ho74]. L_{1} [And96, Wan98b]. L_{1}(0, 1) [CPS87]. L_{2} [AK85, LMW96, MM77b, Ven78, AEU94]. L_{2}(R^{s}) [She95]. L_{∞} [BMP92, Hil98, LL90]. L_{p} [Kry96, LM82, Log79b, Log85, May79, McC81, CSW80, Wal97]. L_{p}[0, 1] [PS82]. L_{q}(−1, 1) [BH85]. L_{loc}^{∞} [And91]. λ < 1 [Dol81a]. m [Gut82]. M(λ) [Kra89a, Kra89b]. R^{2} [Qui93]. R^{3} [Qui93]. R^{n} [Dom72]. Z^{d} [JRS94]. C^{3} [McC97]. R [Rot93b]. R^{2} [BW99]. R^{3} [AS79]. R^{4} [Eck92]. R^{k}, k ≥ 3 [And93]. R^{N} [FSW96, NS89, Gru85]. S_{∞/∥+∞}^{∥/∥+∞} [vdP83]. N [AL78, AB93a, AB93b, BH85, Gau90, GM93, MS88, Sac84, San98, Azz80, Azz81, Bog74, Bro75b, CD80, Hen89, Joy97, KMT91, Kow82a, Kow82b, Leu74, Mel85, MM79, Pet84, Sin75, Wyr77, Zio85]. N + 1 [CS99]. ∇^{2}ϕ + λ^{2}ϕ = 0 [Mer76]. n ≥ 2 [Dom72]. SU(2) [Koo93, Koo91]. p [CL91, CK95a, Jan96, LMN87, Man91, Zha86]. P^{2}(F_{2}) [SY89]. P_{n}^{(0,−1)}(x) [WZ94]. Φ′/Φ [Wim85]. pn [Sch90]. p → ∞ [Jan96]. Q [KMM94, AKM87, ASC87, And76, Aom94, Ask82, Bre81, BI83, Car73a, DG92, Eva92, GR83, GR84, Gas85, GR86, Gas89, GR89, Ger94, Gus92, Gus94, Hab88, HM80, Ism86, Kad88a, Kad88b, KM88, Koe94, Koo91, Mil70d, NR85, Rah86a, RV86, Rah86b, RS94a, Sta80b, Sto97, VK91]. q, f ∈ H_{−1}[a, b] [Whi79]. Q_{n}^{−m}(cosh z) [Fre90]. R [Rei86]. r = 2, 3, 4, ⋯ [Kem82a]. R^{n} [Kni72, Yan96]. S [Oja78, LL90]. s > 1 [EV97]. S^{2} [SS92]. σ [Ver72]. ∫_{0}^{1} f(j)A(j, x)/f(x) dj = ∫_{0}^{1} f(x)A(x, j)/f(j) dj [PW82]. SO(4, C) [KM78]. Sp(n) [Gus94]. su(1, 1) [Edw93, MR91]. SU(2) [Koe94]. SU(2) × SU(2) [Koo85]. SU(3)/T (k, l) [Ura84]. SU(n) [HBL76, DG92, Gus94]. T [RA79, Cla83, IK76]. PC^{1} [KS79c]. SL [Hra85]. Sp [Hra85]. × [JS87]. U(n) [Gus87a, Gus87b, Hol80]. u′′(t) = A(t)u(t) [Lev73, Lev73]. u′′ + c(t)f(u)h(u′) = 0 [PJ73]. U_{q}(3) [LB94]. u_{t} + H(u, Du) = 0 [BJL98]. u_{t} = ωu_{xx} [Ise85c]. u_{tt} − Δu − Δu_{t} + f(u) = 0 [ÁP88b]. U_{xx} + U_{yy} + ε^{2}U_{tt} = U_{t} [Lo72]. u ∈ H_{1}[a, b] [Whi79]. v [Ber75b]. ε [Cuy83a]. ε^{2}Δu + ru = f(x, y) [HK90]. ε∇^{2}w = (∂/∂y)w [CL73]. ε∇^{2}w = ∂w/∂y [CL71]. ε → +0 [CL71]. ε → 0^{+} [CL73]. W [k, s; C_{ij}] [She80b]. W^{1,1} [Lar98]. W^{1,2}[−1, 1] [Coh71a, Coh75]. W^{r,2}[−π, π] [Coh71b]. W_{km} [D’A75]. W Z [Zei94]. x′′(t) + p(t)g(x(t)) = 0 [BY72]. x′(t) = −a(t)f(x(t − r(t))) [Had74]. x′(t) = ax(λt) + bx(t) + f(t), 0 < λ < 1 [Lim78]. x′′ + g(t, x) = 0 [Ste80b]. x^{t} log x [Car87]. xD [Tre79, TF84]. xΔ [Tre79]. x ∈ Ω [Whi80]. y′′(x) − λ^{2}p(x)y(x) = 0, y ∈ L_{2}(−∞, +∞) [Wei71]. y′′ + p(x)y = f(x) [Wal76]. y′′ + p_{1}y′ + p_{2}y = 0 [CM80]. y′′ = φ(t)f(y) [Tal81]. y^{(n)} + py = 0. [Jon81]. Z′ [de 78]. Z_{2} ⊕ Z_{2} [Bri88]. |y| ≤ 1 [CL71, CL73].

-acceptability [Ise79]. -accretive [Gut82]. -Algebra [KMM94, Rei79]. -algorithm [Cuy83a]. -analogue [Kad88a]. -Analogues [And76, Car73a, Kad88b]. -Approximation [MP77, Peh84b]. -arrangement [SY89].

-Ary [Zha86]. -Beta [Ask82, DG92, Eva92, Gus92, Gus94]. -binomial [HM80]. -body [MS88]. -Cell [AJV94]. -classes [BH85]. -Clausen [GR89]. -component [KoM98]. -data [DL84]. -Difference [Aom94, Mil70d]. -dimensional [Azz80, Azz81, CD80, Ura84, GM93, AL78]. -Dirichlet [Jan96]. -Equation [Leg76]. -Equations [Kel79a]. -Error [Wal97]. -Estimates [And91]. -extensions [Gas89]. -Fraction [Lan83b]. -Fronts [San98]. -Function [Mil74a]. -functions [Fie73, Fie83]. -Hahn [Sta80b]. -Harmonic [Man91]. -Instability [MM77b]. -integral [Hab88]. -intégrale [Hab88]. -Jacobi [GR86, Ism86, Sto97]. -Junctions [Sch90]. -Laplacian [CK95a]. -Legendre [Koo91, Koe94, VK91]. -Level [Joy97]. -Like [Jua95]. -Linear [RS94a]. -Measure [Pes95]. -norm [LM82]. -orthogonal [Oja78]. -Periodic [FR91]. -perturbations [Lon83]. -Pollaczek [ASC87]. -Quasiconvexity [FM99]. -Racah [GR84]. -remainder [JW80b]. -Scaling [LMW96]. -Selberg [Rah86a]. -separation [Rei86]. -Series [Ger94, AKM87, KM88, Zio85]. -soliton [Sac84]. -Solution [YL94]. -Solutions [LW95, Lon87]. -Spaces [Chi77, May79]. -Sphere [KMT91]. -Stability [MM77b, Ven78]. -Stable [Ehl73]. -symmetric [ML93]. -systems [RA79]. -th [Bog74]. -Theory [Kry96]. -Type [Ver72, Aom94]. -ultraspherical [Bre81, BI83, GR83, Gas85, RV86]. -Variables [Ber75b]. -width [BH85]. -widths [Mel85]. -Wilson [NR85, Rah86b].

10 [Ban79, Car86]. 14 [Log84a]. 15 [GLY88]. 16 [Mad88]. 16th [FLLL89]. 18 [WH88].

2-resonance [vV84]. 21 [Tah91]. 2D [HFL99].

31011a [Ban79]. 34087b [WH88].

58 [Spe79].

8 [Cig79, Spe79]. 80d [Ack80a]. 80j [Obi81]. 81f [Ban79]. 82d [Car86]. 84f [Log84a]. 85j [GLY88]. 86f [Mad88]. 88i [WH88].

9 [Ack80a, Obi81].

Abel

[Atk74, GL90, LW79, Ort80, PS72, Wal79]. Ablation [Ric90]. Abrikosov [WY92]. absolute [Dom88]. Absolutely [AP74, HS86a, Saa74]. Absorption [BNS98, Wol93, CD95, HV87]. Abstract [Ada78, Bra76, Car71a, CN79, Det78, Gri79b, Hir86, Kif80, Lag75, Say71, Str76, Bre82, GS87, Gv85, GH87b, Hea82, Sta78b, TW79]. acceptability [Ise79]. Accessory [Har94a, Har94b]. accretive [Gut82, Mer87, Oht79]. accurate [CL87]. Ackerberg [Olv78]. Acknowlegment [Ano94]. Acoustic [CD96c, Kli97, SU97, Col78, Col82, CK84, CKP89, Coo88, CD95]. acoustique [CD95]. Across [Hab76]. Acting [WY75]. active [Zem79a]. Adapted [MP98]. Addendum [GLY88, KM91a, Noh74, Car86, Spe79]. Addition [BB97, BC90b, Car71c, Dun78, Dur77, Dur79, KB73, Koe94, Koo75, Koo77, Koo91, VK91, Lai82, Rah79, RV86]. Additional [Lev74b, Lew74]. Adiabatic [BSSV98, HS96, Ste77, Was73, Was74, Ste85b]. Adjoint [BH76, Col73, EP97, Gre91, Mil75, Mit75, Nay78, St.84]. Admissibility [Key91]. Advection [EM90, Hil97]. Affine [CL96, CS97b]. after [Coo74]. Against [CO92]. Age [BIT91, HY95, IMP92, KM91c, Wan80, Bra83, Lan85, Mar81, Swi80]. Age-Dependent [HY95, KM91c, Wan80, Lan85, Mar81, Swi80]. Age-Structured [BIT91, IMP92, Bra83]. aggregated [DW82]. Ahmed [EL86b]. Airy [LY70, OT94, PS91b]. Airy-Type [OT94]. Algebra [Edw93, KMM94, KV98, MR91, Rei79, Zei80]. Algebraic [AMRR92, FT94, Kai92, KO96, MW78, MW93, MG70, OK94, Ser75, Vas79b, Vas79a, Vas81, Cha79, Wex80]. Algebras [Mil71b]. Algorithm [Gu92, HM84, MMM97, OFK77, CL87, Cuy83a, Oht79, Cuy84]. Algorithms [MT95, SN71, Thr94, WW88a]. All-Time [HL98]. Allen [EY98, ES97a]. Alloys [AC80, SZZ98]. Almost [Ber76, FSW96, HT71, Mil71a, Muc70, Pog99, Šim85, Dav78a, GH86, Hir85, Per82, Sei87]. almost-periodic [Dav78a]. Along [Her91]. Alternating [Tre73]. Alternative [BH75c, GG77, Kug74]. Ambiguity [AT85]. Among [Pal77]. Ampère [KS91]. Amplitude [St.88, Veg93, Coo88, Sal89]. Anal [Ack80a, Ban79, Car86, Cig79, GLY88, Log84a, Mad88, Obi81, Spe79, Tah91, WH88]. Analogies [DM82]. Analogue [Des98, GM94, Wil72b, Kad88a]. Analogues [And76, Car73a, Ism85, Kad88b, Per86, Per88, Sid85]. Analogy [CS77c]. Analyse [CD95]. Analyses [DGH96]. Analysis [ABH94, AO81, AM95, AJV94, AF96, Ast91, BB90, Bra71a, Bra72, Che92c, CE99, Chr80, CD96b, CZ94, CD96c, Dom72, Dun93, DDMT93, EG90, EF85, FL92, Flo91, FMP98, Fri81b, GS93, Han99, HH88, IM91, JNW94, Lan74, LPSS89, Lou94, Mar83b, Ngu90, NMIF90, Pan96, Ral83, Rus77, Sch90, Smi71, Smi91b, SZ87, SK76, SS95, Urb96, Val94, Vil94, WS80, Yan92, Yue73, Zho96, AA88, BCG89, CD95, Dol79c, DM82, Kha87, SW86, SK87]. Analytic [BG88, BG84, BB97, BC90b, Car74a, Car74b, Car74d, Car79b, Car86, Cha79, Cha78, Col75, FT94, HV97a, Imm90, Koo74, Koo75, KG71, Lan83b, LT92, LR84, MS99, Mil76, Ols74, Pip93, RM78, Sha77, Yu74, Yu77, de 78, BG89, Bra86, Büh87a, Büh88, CV87, CS89, Coh89, Kaz84, Mil79, Mil84]. Analytical [Eng99, IMP92, Lin94, Tem74, IR89, Piq86]. Analyticity [Hay91, LB96, TBD^{+}96, Zha95a, vd86, Str80]. Andrews [ASC87, Ask80]. angle [Cof82]. angular [Tal80]. Anharmonic [JMR96]. Anisotropic [ES97a, Jol92, Urb96]. annihilation [Cam89b]. Annular [LP91]. Antiplane [GN93]. Any [Olv92, Pog99]. Aomoto [Sas91]. apertures [Zio85]. Appearance [CP73]. Appell [Car71b, Car75b, Car76b, Iwa88, Mil73a, Sas91, Sin80]. Applicability [Paz72]. Application [BMP92, Chi74b, Dou98, Dun90a, EM90, Hof99, JNW94, Kha87, Lab99, Mer87, Nii97a, San98, Tem90b, Thi93, BG89, BD86, CS80, GS81, Gun73, Hea88, Kaw79, Lou80, Mad85a, Mad88, Nip86b, Ren84, Tem87, TWW89]. Applications [Ack77, Ada78, AGV82, Ban75, Bar85b, BJL98, BH75b, BVB94, Bud89, Col70, DS81, ES76, Erb77, Gus85, GLP91, Ład94, LM75, Lew79, Mat78, MP77, Mil70b, MNY94, Osl71, Piq89, Ros78, Rus77, SW75, Ubh76, Yao95, Bra86, Chi87, CLP89, Chu84, Coh89, DK83, Dol80, HM80, How85, KS80a, ML84, O’R87, Rab84, Sas91, SS80, Tah90, Tah91, Wal79, Whi80, Zei80, ZD84, dM86, vd86]. applied [Gin85]. Approach [BB77, CS77c, CG71, IT79, KoM98, Leg76, LM96, MLM97b, Mul78c, NW74, Pes95, Pip93, Sch97, Wal75b, BB78, DG88, Geo87, MB91a, MB91b, Sta88, Zei88, Zio85]. Approximant [BR78a]. Approximants [JT75, Luk75, WB95, Cuy84, Wim85, Cuy83a, Cuy83b]. Approximate [FKS99, Gou92, JT94, Wan98b, DK83, Ste73a]. Approximating [CPW98]. Approximation [Ali95, And93, Bea81, BG75, CRSW76, CSW80, Coh71b, Coh71a, De 77a, De 77b, DJL93, EW94, Ewi75, FI90, FGHW91, FT94, GNP95, GL71, GM76, HW81, Hu91b, LJ91, LJC97, LXLC96, Lon74b, LT94, MW91, MP77, Pet98, Pie74, PS98b, QSW93, Ron97, SS75, SN99, Ste70b, Ubh76, YZ94, Bar87, CS80, Coo88, Cyb84, Dol80, Fry78, GH87a, Ise85b, Jer78, Kel89, LS88a, LM82, LS88b, Luc88, Mar82, Peh84b, PS82, RW78, RS78, Sac84, Sho76, TG85]. Approximations [CG94, Ehl73, Het70, Ise85c, Luk70, Olv74, Pao74, Sen96, SV94, AO81, Cur79, HIN85, Ise79, Ise85a, Ise85b, LM83, LS89, LS79, San80, Sha80b, Wie87]. Arbitrarily [BW99, RS86]. Arbitrary [CD93a, Gro76, Gro77, Gru92, Old70, Old72, Olv77a, Tom86, Whi80]. area [Bra81c]. area-preserving [Bra81c]. Argument [Fre92, HL74, Bog74, Büh87b, GR85]. Arguments [GLP91, SS78, AT89, GL84, HW81]. Arising [BBDU93, BHN72, BT91b, DDD96, Hu90, HW92, Id94, Jam98, KoM98, MZ98, Ngu90, Noh74, SZL97, Veg93, XA91, AW78, BG84, DR86, Gar87, GS81, HP83, Hil82, Hiv78, HM89b, Kap79a, Kap79b, Kna86, KS80b, LK89, Lui82a, Lui82b, Lui85, Lui86, MPN89, MRS81, Nau79, NNB88, Piq86, Ter83, Ter88, Vog85, Wex79, van86]. Arithmetic [CS93b, BG84, Sal89]. arithmetic-geometric [Sal89]. ARR [CS77c]. arrangement [SY89]. Arrays [Car70a, RE98]. Artificial [HS89, Hal91, HK86]. Ary [Zha86]. Ascent [RS76]. Askey [Hab88, ASC87, Chi87, Gas73, Gus94, Hab88, Ism86, Kad88a, Koe94, Koo93]. Aspects [DS95, Hol87, Rei79]. assisted [AA83]. Associated [Ask82, Ban73, Bax71, BSX95, Cha76, Cho84, CS93b, Gus92, Kan93, Kim98, Koe94, Let94, McC97, Nev84, PS76, ST90a, Sch76a, Sch76b, Sri90, Ste73b, Val94, Wan98a, Aom87, Blo80, CF85, Ell82, Gar88, ILV89, KM88, Kol79, Las83, Mar89a, Mar86, OP87, SY89, ZW85]. Association [Llo80, Leo82]. assumptions [BF88, BF89b]. Asymmetric [ACG99, HT98]. Asymptotes [Eli72]. Asymptotic [Ada78, Any77, AP92, AB80, AF96, Avr93, BH75b, Ber76, BHL75, Ble77, BG91, Bra71a, Bra72, Bra91, CG85, CG94, CN81, CZ94, CD96c, Dat72, Dun90b, Dun94, EG90, ET74, Erd74, ES84, Fre87b, Fre90, Fre92, FK88, Gal95, GIO90a, GH88, Gor75a, GB72, Gra72, Gre80, Gri81a, Hab76, Had74, HLL72, HL70, HB73, HL74, Har83, Het70, Hol93, How86, How78a, Imm91a, Jon94, JW74, KV91, Kam89, KN97, Kap72, KS85a, Lae75a, Lae75b, Lan74, Leu73, Leu74, Leu75, LM75, LW94, Lim78, Lon71, Lon74a, Lu97, Mar81, Mar70, Mar78, MW87, Mey86, Mis95, Mos83a, Muk72, MR80, Nam78, Nar93, Ngu90, NY97, NP93, O’M71, Olv91b, Olv91a, Ost96, Pao72, PS82, Per82, PS94, San80, SW86, Sel81, Sen96, SS78, SK73, She72a]. Asymptotic [She72b, She86, SS75, SK87, SW79b, Sid85, Šim84, Šim85, Sin75, Smi71, SS73a, SZZ98, Sta79, Ste70b, Ste77, Tal81, TF75, Tem90b, Tre76a, Tre84, Tre88, Val94, VFO87, VK91, Van96, Van98, Vol95, Wan78, Win74, Won76, Won78, Won80, vV84, BD86, CS79, CE85, Dun86, Dun89, Erb76, Esh87, Fie73, Fie83, Fre87a, FW88, Gin85, Gri80b, HS86b, How85, How82, Hru85, HSX87, KT79, Kha87, Lei81, Lei85, Mem89, Moe79, Mur83, Nay84, Nay87, ND79, Nev79, Oja78, Olv93, Sin79, Ski80a, SM85, Son80, Son83, Sta78b, Sta78a, Ste72, Tem79, Tem82, Tem87, Veg88, WW88a, Wei71, Wes78, ZD84]. Asymptotically [KR81, Mil71a, Pie73, Sin76, FS79]. Asymptotics [Ban92, Ben96, BEGv87, CI91, CR88, DGR99, Esh87, GW94, HPR97, Hsu89, Ism86, Ism94, Jos99, JP95, Joy97, Kam92, KV99, MNT86, MZ98, Nev84, Olv70, Pip93, SS85, Wal71a, WW92, Bre91, Ful82, Ful83]. Asynchronous [WG87]. ation [KF90]. Atkinson [PR91, Vol87]. Atmosphere [BB94, Gv85]. Atmospheric [Dou98].

Attain [Gla76]. Attainable [BBF99]. attenuated [Mar84, MQ85]. Attractor [Rob92]. Attractors [Bil90, BGT88, DRF95, FT94, GM91, HG99, Ghi86, Mar89a]. autonome [BH87]. Autonomous [CR79, Gru92, Mar70, Pér93, RB71, Nip86a, Nip86b, Tro81]. autoregressive [Cyb84]. avalanche [Mar85a]. Average [Car75b, Rua98, ND79, Vit79]. Averaged [RC73, Wad80]. Averages [Car71b, Car74c, Chi74b, And88, But80, Car87, Per81, Per82]. Averaging [Kol74a, San83b, Mur83, Vit79]. Axial [CS77b, NT96, Buz86]. Axially [HQW74]. Axialsymmetric [Pen92]. Axisymmetric [KS75, CL89, Fry78, Lef82, Mil79, Tur89, Wei88]. axon [CB89, Cos87].

B [BG96, Rei97]. B-Splines [Rei97]. Back [Den91b, Nii97b]. Background [KV99]. Backscattering [SU97]. Backward [BBDU93, Car71a, Dem96, Ewi75, MM80]. Bailey [GR84]. balanced [ES84]. Balls [HV97b, Xu98]. Banach [AF92, BU97, BHL75, Dat72, FK94, GK84, Gut82, Mar78, Nam78, Oht79, Sho72, Tho77, ki72]. Banana [PFK95]. Bananas [PFK95]. Band [FG92, Mel85]. Band-Limited [FG92, Mel85]. bande [CD95]. Bandlimited [Wal88, Wal92b, Log83b, Log83c, Log83d]. Bands [SS92]. Bandwidth [CP84]. bank [EF85]. Bar [Kim92]. Barenblatt [Hul97]. Bargmann [Sei91]. Barnes’ [KM88, Gus92, RS94a]. Barotropic [Zaj94]. Based [GC95, Vit79]. Bases [ANS96, Alp93, BAR90, CS93b, CD93b, Dau93, Gri93, MP98, Zho96]. Basic [ASV84, Gas81, GM94, Ism81, Jai81, LB94, Vas79b, VJ81, Ask80, GR86, Gus87a, SP79, Ver80, VJ85]. Basis [CC98, DJJ91b, LXLC96, Mei76b, Zha92b, DJJ91a, Kaz84]. Bateman [Ask74b, GR84]. Bau [HT89]. Bboundary [Smi75]. be [Pou89]. Beam [Car71a, FS83, Fit82, Hol87, Lun87]. Beams [Yos85]. begins [ACK83]. Behavior [AP92, Avr93, BY72, BIT91, Car96, Coh75, CW99, CL71, CL73, CD96c, DGG98, Den95, EZ97, Gao93, Had74, HLL72, Han77, How86, Imm91a, JW74, KN97, Kap72, KS85a, LY96, LM75, Lon71, Lon74a, Lu99, Mar70, Nar93, Nii78, NY97, NP93, Ost96, Rao75, Sch97, SS78, SS73a, SS73b, SZZ98, Tak92, Tal81, Wan78, Wei82, Wol93, Zha99a, BS83b, Büh87b, CB89, CN81, Erb76, Gri80b, HW84, How85, HSX87, KT79, Kha87, Lei81, Lei85, Mar81, Mar78, Mem89, Moe79, Nam78, Nor88b, O’D84, Sac85, Sel81, She86, Sil81, Ski80a, Sta78b, Wag89, Yos84, Yos86]. Behaviors [LMN98]. Behaviour [Gal95, GM91, Jan96, KV91, Pap95, VFO87, VK91, CE85]. Belohorec [KW83]. Benjamin [MM77a]. Bergman [Col72, Sei91]. Bernoulli [BT91a, DS87]. Bernstein [BG74, Car78b, Cie87, EL89, LL90, LS93, May79, Mir83, Peh90]. Bertrand [BSZ77]. Bessel [Obi87, Amo89, Car70c, Car80, Dav73, Dun90a, DL91b, Dur79, EL84, EL85, EL86a, Fet77, Fet78, Gas75, GN89, IK76, IM78, Ism81, JM72, LM83, LM77, LMS73, Lor80, LS88c, Lor93, LS94, McC77, Mer75b, Mer77, MR94, Nås78, Obi86, Par84, Tem90b, Wim85]. Best [CRSW76, Coh71a, Gal91, Hen87a, Lon74b, LT94, Mia93, MP77, SS89, Peh84b, PS82]. Beta [Ask82, DG92, Eva92, Gus92, Gus94, RS94b, Gar88, Tem87]. Bethe [Wil73]. Better [CD93b]. Between [Bao98, BFF^{+}90, Imm90, Car70c, Fet77, Fet78, Kin86, MO83, Sym86b]. beyond [JP95]. Bi [HQW74]. Bi-Axially [HQW74]. Biased [Sch90, BCG89]. Biaxially [Wei74]. Biconfluent [LMPS90]. Bidimensional [CD90]. Bieberbach [Gas89]. Bifurcating [San98, TT83, Cic87, Nis82]. Bifurcation [AA88, Ast91, BM92, BT91b, CDT90, FS95, FR91, Gar92, Guc84, Hai92, HJ92, Ko98, KM95, Lab85, Lab89, Lu90a, Lu90b, Mem89, MS95, NT86a, NT86b, Olm77, PFK95, PB85, Pér93, Pos90, RJ86, RS91a, Rob92, Ruf95, Sch87c, Sch98, SM84, She80a, St.96, Stu89, Wol96, AR87, BB86, FLLL89, Fie86, GH84, IR89, Lou80, Mad85a, Mad88, Web79]. Bifurcations [CSv87, Chi92, Den90, Den91a, Kru90, MS88, NT86c, Nii97a, SM84, Buz86, DS84, Fen83a, Fen83b, FR86b, Hea88, Ste85a]. Big [Sto97]. biharmonic [Cof86, DL84, How85]. Bilateral [WS80, MS79]. Bilinear [Ask74b, Rah76c, Rah81b]. Binding [Vol87]. Bingham [Kim90]. Binomial [Car75a, EP97, HM80]. Biochemical [Lu87, Lu90b]. Biodegradation [MX98]. Biography [dBJ82]. Biological [Wei82]. Biorthogonal [CD93a, Han99, KMT91, Sch74, PT78, Rah81a, Sri83]. Biorthogonality [Rah91]. Birth [Mak76, ILV89]. Birth-Death [Mak76]. Bistable [DL92b, IM91, Mis95, NMIF90, Ter83]. blockdiagonalization [Gin78]. blocks [Dol79c]. Blow [Aye83, BF89a, DH99, GV93, Gal95, GW96, Liu95b, Sou98, Suz95, WW93, Wei99, Ali95, Bel87, FL87]. Blow-Up [DH99, Gal95, GW96, Wei99, BF89a, Liu95b, Sou98, Suz95, WW93, Ali95, Bel87, FL87]. Blowup [AC94, BL90, BB92, Bre91, Gal91, Gra98, MY98, PS97, CW89]. Boas [Ism74]. Bodies [Gus98, Wil80b]. Body [BBN93, GK87, HL93, Liu91, LF98, CS85, MS88, Moe84]. Bohl [GS89]. Boltzmann [And96, BD95, Ben96, CPW98, CGP93, Lu99, MM97, Mik88, MP97, NP95]. Bona [MM77a]. Borel [LB94, Sid86]. Bound [BSX95, BR78b, Gal91, LN98, WZ97, Chi83, HKS91, Mil70c]. Boundaries [BV98, HP98, Ren96, Sak93, Day89]. Boundary [Ack78, Ack80b, AGJL92, ABV96, AC80, AC98, AF90, AF96, AL98, BG88, Ban92, Ban73, Bao98, Bar77b, BT91a, BG76, Bax91, Boh90, Bro75b, BV92, Bus78, Car73b, Chi74b, Col94a, CL71, Cos88, DGR99, Day89, Dea71, DL84, Dun93, EP98, Elc75, ET71, FP86, FGHW91, FI96, FR94, GY90, Gil70, Gio93, GK77, Gre91, Hal91, Har72b, Har74, Har79a, HLM92, HP98, HS84b, Hol77, HS95, How75, How78a, How78b, Hsu89, Hu91a, HW92, HP92, Ish96a, Jam98, Joh75, Joh77, Joh80, Kai91, Kap79a, Kel84, KL96, KS96, Kim98, KO95b, Kui76, KM95, KK97, LB74, Lag72b, LW77, Liu91, LMN98, MW75, MW93, McL90, MA88, Mil76, MP98, MNY94, MW74, Mul78a, Nag78, NU95, Nii78, NT96, NP95, O’D84, O’M70, Pao74, Par72a, PS76, Pet84, PP95]. Boundary [Rei73, Rei93, Ren90, SS77, Sch97, SZL97, Smi91b, Spe89, SHM89, Sta71, SS95, Szw92a, Szw92b, Tep74, Tep75, Ter99, Tip74, TY77, Tur84, Wai76, Wal75a, WN81, WW93, Wei99, Wol93, Xie90, Yu74, YL94, Zay82, Zha95b, Zha99b, dM89, Do 93, Ack80a, AGX89, Ale80, AK85, AW78, AK82, BG89, BN78, Ben86, BEGv87, BB79, CFV81, CD80, CR83, Cha72, Cof86, Col82, CS85, Daf82, DeS87, DH84, DK86, EH84, ET72, Gab84, Gus85, HK86, HS89, HLM86, Har78, Hen87a, Hen89, Hin81, HS84a, HP89, How82, Iff87, KKZ84, Kim87, Kna86, Kor88, LS88a, LS89, Lef82, Lev83, Lev88, LS86, LS87, LK89, LP81, Mar82, Mar83b, Mil79, MTH80, Mul81, O’M79, O’R87, Pet81, Rei79, SW86, Sch81, SAW87, Spe79, Sym86b, Tef78]. boundary [Ter83, Ton81, Ton82, Vog85, Wei71, Wei84, Wie87, WW88b, dW80, vP86]. Boundary-Value [EP98, KO95b, Rei93]. Bounded [Avr91, Avr93, BH75a, Ber72, CH92, CR92, Cur79, Cus75, Dek71, FHX96, GL74, Hil98, KK97, Lev73, Lon74a, Prü88, SK76, Swa82, TBD^{+}96, Tro90, Zaj94, Log83a, Log83e, Tro87]. Boundedness [Bak72, Chu76, Han75, HP93, LS71, LZ95, Li85, Mor90, Fit82, Har70b, Har70a, HMP87]. Bounding [KS78]. Bounds [AB93a, AB93b, Ban75, BR78a, Blo80, Bus78, BR71, Car76a, Car97, Det95, EP99, Eli72, Fox72, Fre90, Fre92, Het70, HFL99, Jer99, Lar81, Lev76, LM75, LW94, Lim78, Log88, LS94, Nås78, Olv74, Paş89, Pay70a, RDLH77, SS77, Van96, Wal97, Won76, AMS86, BS83b, Cos84, DK86, EH84, Fre87a, LL79, Sch87a, Spe79, Vit79]. Boussinesq [Cur79, Liu95b]. Box [CSW91]. Branch [Lae75a, Lae75b, Sat77]. Branching [Olm77, SW75, Van90, Zac74]. branchpoint [Cam89b]. Branges [Gas89]. Breakdown [GP99, Nip86a, Nip86b, Lin86]. Breaking [Ast91, Pos90]. Broadwell [Fan97, Pes95]. Brouwer [Kug74]. Bubble [WW98, She83]. bubble-column [She83]. Bubbles [de 95]. Buck [Ism74]. Buckled [KS75]. Buckling [Bar85b, CO92, KLV93, KS79b]. Bulk [FL98]. Bunsen [BH99, Mic92, Mic96]. Burgers’ [Lev88, Sen97a, Sen97b, Dix96, Eng99, Ito94, LMN98, Sle95]. BV} [Lar98, Kim83].

CAD [Kre94]. CAD/CAM [Kre94]. Cahn [EY98, EG96, ES97a, Gra95, Ish96a, WW98]. Calculation [GR75, GR78, GR79, Was74, Sal89]. Calculus [Bus72, Kai92, Lub86, RT94, Shu78, Str77a, Str77b]. CAM [Kre94]. canal [Sac84]. cannot [Pou89]. Canonical [AKM87, Chi77, Gra80, Gra88, Har94a, Imm91b]. Cantilever [HH88]. Cantor [Lan96]. cap [PGRSV89]. Capacity [GR75, Pie83, Sha73, Van96, Lon88]. Capillary [CF96, Kim98, Zaj94]. Caps [KS75]. Carathéodory [Por76]. Cardinal [LS79, Rei97, ST76, dB76]. Cardinal-type [LS79]. Carlitz [Cig79, GW94, Sha76b]. Carried [de 78]. Case [Dou97, ES97a, FM98, LT92, Lew74, McL78, PP93, PP95, Que71, Tep75, Tro90, JSW87a, JSW87b, Lui85, Ral82, Sac84, Zei87]. Cases [BBL97]. catalyst [Veg88]. Catalyzed [HJ92]. Cauchy [Ade74, Ame82, Avr91, Bel81a, Bel81b, BV98, Bra76, Bre82, Car73b, Coh89, Col97, CG71, Col74, CE87, Ell82, GR92, Ish96b, Kni72, MM97, Mar86, Mil84, Pay70a, Say71, Smi71, Ste75, Tsu84, Tsu89, Yu77]. Causal [Dol81b, Fei79b]. Causality [Dol72, Mei76a]. Caustic [Kam92, Kam89]. Caustics [BGG84]. Cavities [Lu90a, LKMT88]. Cavity [LF98]. Celestial [ElB90]. Cell [AJV94, GH87b, WG87]. Center [BB92, Gla76, Tur96]. Center-Unstable [Tur96]. Centers [CD97]. Central [MS95, Wyr77, CR88]. centroid [EL87]. Certain [Ask74a, Bax71, Ber75a, Boe78, BR81, BC90b, Bus77, Car70a, Car72, Car73a, Car78a, CJ76, DP97, Eli72, Esc73, Ewi75, FTZ88, HL74, Hir72, Hir75, Hun73, Kee72, Lee75, LL93, Lin78, Log84c, Old70, Old72, Olm77, Par72a, SN71, Slo75, Smi71, Sri70, Ste70b, Swa72, YH82, ABM87, AH83, BSL87, But79, CDMS81, Dil88, Fie73, KKZ84, Log85, Mon80, PV84, She93, Sim80, Tch96, Tom86]. certains [Tch96]. CFL [GHLS91]. chain [Eng82, EK84, Mor87]. Chains [KZE90, RE98]. Chandrasekhar [Jua95, Leg76]. Change [Key91]. Changing [KV91]. Channel [Lu97]. Channels [KR81]. Chaos [BFF^{+}90, MR98b, FMV98]. Chaotic [HM84, WM87, Yag92, Ber88]. Chaplygin [Kla77]. Chapman [HNT98]. Characteristic [Ban73, BR78b, SS73b, Day89, Zem81]. characteristic-resistance [Zem81]. Characteristics [HS98a, Pey71]. Characterization [ASC72, BVB94, Boj94, CS93b, Eir92, IM98, LW95, Mer75a, Mer76, Peh81, Peh84a, Som80, Som82, vdP83, BC79, FS83]. Characterizations [KKZ86, KLY94, Ver72]. Charges [Pie74]. Charlier [Zen90, Zen90]. Charney [BCT88]. Chebyshev [BC79, GP84, IM87, Lai80, MP77, Nür80, Pas74, ST90a, Som82, TG85]. Chebyshevian [NSSS84]. Chemical [Ale90, Jam98, MS86, GK80, WL82b]. chemically [DJ87]. Chemicals [Ost96]. chemistry [DR86]. Chemostat [HRX98]. Chemostat-Type [HRX98]. Chi [GG77]. Chi-Square [GG77]. Child [BD95, Ben96]. Chiral [AN98a]. Chrestenson [Zha86]. Christoffel [EMN94, Hug75, LS93, NEM94, Pfa86]. Chromatography [Ost96]. Circle [Edw93, Hal84, Spi79]. circular [Mor87]. Clamped [KLV93, Cof82, PP83]. Class [AS87, Alp93, AJV94, AB72, Ban73, BR78a, BS74, BG91, Bra76, BH76, CJ99, Car77a, CA71, CD74, Col70, CS78, Dol79d, Dom72, DH99, Eli72, GL90, Han75, Han76a, How78a, KSK78, Kov95, KS85a, Lag75, Lin91, Lin94, Lin75a, Pan71, Pan72, PS76, SW73, SS78, ST76, She80b, Siv71, SS73a, SH71, Tem90b, Wal71a, Wei82, Wei99, XA91, AL78, AO81, Bel81a, Bel81b, Blo80, BD86, Bra79, BB79, Cig79, DH84, Epp86, Fie83, Hiv78, HP89, Ira89, Log79a, ML93, Nam78, Nau79, Nay84, Nay89, Oja78, Rab84, RS86, Sac85, Som82, Son80, Tah90, Tah91, Tem82, Vic78, Wex79, ZW85, dT78, van86]. classe [Tah90, Tah91]. Classes [BG76, Ber75a, BS76, Coh76, KO95b, WS80, BH85, Fie73, HW84, Kor88]. Classical [ASC72, Bor99, Coh76, ES97b, Gas73, Gas77, Gus92, HL98, KT99, McC97, RT94, Stu96, BG84, FR86a, GR85, Klo85, Ton81, Ton82]. Classification [DS84, SR93]. Clausen [Gas89, GR89]. clines [Pou89, Lui86]. Close [JRRV97, MP97, PM82]. Closed [BHU96]. closure [Ben82a]. cluster [HM89b]. Cnoidal [Dv96]. Co [Let94]. Co-Recursive [Let94]. Coagulation [MLM97a]. Coalescing [Dun90b, Dun94]. Coating [Tro93]. Codes [Llo80]. Codimension [Guc84]. coding [Chi87]. Coefficient [CJM98, Duc95, EI97, FG87, Sym86b, Tre88]. Coefficient-to-Data [Duc95]. Coefficients [AHZ93, Aom94, Ask71, Ask74a, Bao98, Cam77, Chi77, Col75, CC96, Elc74, ET74, ES75, EG91, Imm91a, JRRV97, KL98, KL99, LT92, Lew86, Paş89, Pey71, Rea86a, Ros98, Saf99, Ser75, Šim85, Tre76b, VFO87, Van96, Vas79b, Wyr77, YS93, AW79, Bar79b, Bar81, Can85, CV87, Car79a, Dun74, GN83, HM80, Jer78, MM80, Piq86, Rob83, Tch96, Vas79a, Vas81]. coerciveness [Rab84]. Coffee [FT99]. Cohen [Mar83a]. Coherent [Wil85]. Collection [BV92, KS96]. Collision [ElB90, CLP88]. Collisionless [GS99]. Column [CM98, CM00, She83]. Columns [CO92]. Combinations [Dav73]. Combinatorial [AGV82, Car73a, Car78b, IT79, CR88, GJZ88]. Combustion [Avr93, Bon96, DS95, FT88, HNT98, HP85, Has88, Li93, LY95, Ter90, BNS85, Gar87, HP83, Ter88]. Commutant [FF92]. Commutative [Bra74]. Commutes [Olv92]. Compact [And93, Gut82, Han99, Hir72, Li85, ST90a, Wil80b, Auc89b, Lui82b, MC79]. Compactly [BAR90, CD93b, Dau93, Vol95]. Comparison [AL81, Eli84, Erb77, Gar81, Mul78a, Pao83, Sle95, St.84, Tom83, Tom86, AL78, Ame82, BE83, Cuy83b, DiG81, Kim86, Mur83, Pat79b]. Competing [HV97a, WZ97]. Competition [Ko95a, Ko98]. Competition-Diffusion [Ko95a]. Competitive [Bro83, FG99, Hir90, Smi84, Smi91a, Hir82, Hir85, Smi86b]. Competitive-cooperative [FG99]. Complementing [Yu74]. Complete [AVV90, AVV92, Col75, Far73, HW82, Neu88, QV96, SH71]. Completely [AP74, BS75, Mug75, Mug79, PS72, AZ79, CN81, HW84, Sha77]. Completeness [Pin85, Sac83, Cho76]. Complex [BSSV98, Dav73, Dun90b, Dun94, Fre92, Kai92, KS91, STW88, BKW85, DM82, HIN85, HS79, LS88b, Sch80, Sch84a, Sen97a, Sen97b]. Complicated [DD91]. Component [RJ86, KoM98]. Components [AL77]. comportant [BH87]. Composite [Osl70, Ski88]. compositions [Yom86a]. Comprehensive [FT99]. Compressible [AM95, Che92b, HL98, Hof96, HS96, KK97, Zaj94, CDHL88, CL86, Coo88, Kel88]. compression [Buz86]. Compressive [Bre93]. Computation [Ast91]. Computationally [Bus78]. computations [Ura84]. computer [FS79]. Concave [Her91, Hsu89, Kad72]. Concentrated [GS77]. Concentration [FMP98, WZ97, CP84]. Concept [Hai82, Pal77]. Concerning [Ben75, BV92, Dör91, Elc75, Gas73, Har80, KK92, Lir71a, Pat71, Sib81]. Concurrent [She83]. condensation [Sen97a, Sen97b, AN98b]. Condition [AMS^{+}92, jCcW99, Dat81, HP92, Jam98, Kim98, KK97, LT92, Šim85, Wal75a, Xie90, Yu74, Chi74a, Gri79a, Hin81, Lev88, Peh84b, Rab84, Sin79, Wad80]. Conditional [CS89, Mar80]. conditionedness [MN86]. Conditioning [dM89]. Conditions [AL98, BT91a, BC95, BS74, BY91, Bra95, Bro75b, jCcW99, DGR99, Elc75, ES75, GT94a, Gu92, GLP91, Hal91, HS95, Ish96a, Kai91, Key91, KZ95, LM95, McL90, MP98, Olv78, PJ73, Šim84, Spe89, SHM89, Tre76a, Tur84, Wai76, WW93, Yam90, AW78, AK82, BG89, Bar79b, Bar81, CD80, Cof82, ET72, GN83, HK86, HS89, HS84b, LS89, Lev83, LS86, Tef78, Tre84, Tre88, WW88b, vP86]. Conducting [Van89, Wol96]. Conduction [Dav78b, Gal91, LP94, PW70, DM82, KM86, Nau79]. Conductivities [Bro96]. Conductivity [EI97, FKS99, KSS97, Smi91b, FG87]. Conductor [AR98]. Cone [BRC72]. Cones [AGGO94]. Configurations [MS95, Wil85]. Confluent [Olv91a, Tem90b, Dun89, KS86, Olv93]. Conformal [GR78, Hom79, IL96, KM83]. Conjecture [And80, Ber99, Gar90, ZZ90, ASC87, Kad88a, Zei87]. conjectured [Kad88a, Rah86a]. Conjectures [Gas73, KK92, Mac82, Gas89, HT89, Zei88]. Conjugate [CS93b, Fin74, Hal77, Lip72, Lip73, Tom81, HIN85, Min86]. Connected [Det78, HW92, PS76, Val94, Log83f, Log84a, She88b]. Connection [Aom94, KN72, LMPS90, Leu73, Leu75, MP83, Nau76, Olv77b, Olv77a, PM82, Wyr77, BJL81, BJL88, MO83, SS80, Sch80, Sch84a]. Connections [Den91a, Sat90]. Consequences [Str80, MB91a]. Conservation [CC99, Daf95, DL96, Die95, Fan93, GR92, Gim93, HS98a, IMP90, Jam98, Key91, KKL99, KT97, Nak88, NY97, Sch84b, Sin97, Ter99, TW91, TW97, Wan98b, You99, Daf87, Gar87, How87, JW88, Luc88, MMN85, Wag83]. consisting [Dol79c]. Constant [Bra83, Coh71b, Coo74, For90, GLP91, JRRV97, KM91c, KL98, KL99, Ric89, Šim85, Smy79, GM84, Mar85b, Tre88]. Constant-rate [Bra83]. Constantly [WY75]. constrained [Mad85a, Mad88]. Constraint [Car99a, Shi93]. Constraints [Bry74, Don92, HV97a, Lan78, Pfa86]. Constructing [CD91, PW82, MB91a, MB91b]. Construction [DS99, DGHM96, DGH96, DGH99, Gil70, Gou92, Han99, Her95, Log79c, PS98b, Rah76a, Red71, Swe98, Cof86, LB94, MrS87]. Constructive [Ban74, Elc71, Elc74]. Contact [AMS^{+}92, FL98, Her90]. Containing [CC98, Zac74, Piq89]. contaminant [vd87]. Contaminants [MX98]. Contiguity [Sas91]. Continuation [Car74a, Cha78, Mey76, Yu74, Zha92a, Büh87a, Büh88, Hea88, RS78]. Continued [Ahl93, Fai71, Fie84, Mey86, PH75, BF79]. Continuity [Bou72, Dom88]. Continuous [Art75, CN77, Cha70, Col97, CC99, DK83, Dol70, Dol72, DDMT93, FH90b, Gin79, Han76a, Heg81, Hof96, KV99, Koe94, KM95, LA96, Mei77, Saa74, SV94, Swa72, TW91, Wil80a, Wu90, BS83a, Bos79, BI83, GR83, Gas85, GR86, GK80, HS86a, ILV89, Kaw79, Nür80, RV86, Som80, Som82]. Continuous-Energy [CN77]. Continuous-Time [FH90b]. Continuum [DT81, Ful82, Ful83]. Contour [LY70, Jay88]. contracting [CT88]. Contrast [DDMT93]. Control [MZ98, BGS86, Gri89, van86]. Controlled [PS94]. Convection [EZ97, Gra98, Sai96, Suz95, Rul87]. Convection-Diffusion [EZ97, Gra98, Sai96]. convective [LPSS89]. Converge [Col70]. Convergence [All92, And96, Bea85, Ben96, Ben85, BL74, BLM96, BF79, CG83b, CF98, Esc73, Fie84, GV93, GC95, HJ98, HP93, HT71, IS98, Jam98, JT75, KP92, KT97, LB74, LA96, LM95, MX91, Muc70, Mug79, NV95, Nov88, PH75, PD71, Red84, RTV96, SV83, Šim84, Šim85, ST91, Tre76a, TW97, Wan98b, Chu84, LS88a, LS87, Ngu89, ST90b, TG85, Hir85]. Convergent [Ber75a, DL91b]. converging [Sid86]. Converse [Lu99, Mul87]. Convex [Bea81, BHU96, Ber99, Bry74, BS75, DQ90, DP97, FM92, GNP95, Gus98, Mug75, Mug79, PS72, Wil80b, AZ79, Lon88, MC79, PP83, Sha77, Tah91]. convexes [Tah90, Tah91]. Convexity [Car99a, EL85, GP93, GS80, Hu91b, KI95, Kli97, Sen78, Sin97, LM77]. Convolution [BI75, BL74, BMY77, Car77c, GC95, GW97, Ho74, JSW87a, Kra72, LXLC96, Mad85b, Pan71, Per73, PW70, Sta79, dGdJ75, de 85, Lai80, Lew79, Log79b, Log85, Ste73a]. convolution-type [Ste73a]. Convolutions [ASC76, HL74, KV98, LW94]. Convolved [Car77b]. Cooperative [Hir90, Mie87, Smi91a, Tak92, TKS93, FG99, Hir82, Hir85, Smi84, Smi86b]. Coordinates [Mil74b, Mil74c]. Corner [MW91, O’D84]. Corners [Dau89]. Corrected [CM00]. Correction [Osl73]. Corrector [Bal99, Sen85]. Correlated [Mia93]. Correlations [BFF^{+}90]. Correspondence [RT93]. Corresponding [LMN98, JT79]. Corrigenda [Obi87]. Corrigendum [Fet78, Obi81, Spe79]. Cost [KI95]. Couette [GS86]. Coulomb [Jos99].

Coulomb-like [Jos99]. Countable

[Den91a]. Countably [Zem79a, Zem79b]. counterexample [RR84]. Coupled [Chi95, EW94, GHL96, GGI92, HY95, Jef91, Jua95, RE98, RR95, St.96, WO91, CS86, Cos84, EK84, ET89, JAK88, Mor87, Pao87]. Coupled/Quasi [EW94]. Coupled/Quasi-Static [EW94]. Coupling [Bil90, KZE90, Ros98, WY92, WW99, Zio85]. Crack [AR98]. Cracks [ABV96, BV92, KS96]. Creases [Bra95]. Criteria [Boe95, Chi77, Jon76, Lew74, Rio92, Spi79, Ven78, Cic87]. Criterion [Nak78, Tha93, Bel87, CS79]. Critical [AM94, CP97, Hab76, Lip72, Lip73, SW73, WY92, Yom86a, JSW87a, JSW87b]. Criticality [Bal99]. Cross [IM78, Kot94]. Cross-Product [IM78]. Cross-Sections [Kot94]. Crystalline [IS98, Ryb99]. Cubature [BSX95, Xu94, Xu98]. Cubic [CK92, WK92]. cumulative [Tem82]. Current [Dol77, Ack81]. Currents [Rei93]. Curvature [BS97, CRSW76, CZ91, dPFW99, GJ82]. Curve [KG71]. Curved [AN98a, Zha99b]. Curves [GM96, GPG91, SS89, Smi86a, Bra81c, GJ82, Lon88]. Cusp [Kam92, WK92]. Cutoff [Log88]. Cycle [St.88, Che81, LC88, WG87]. Cycles [GL90, HP93, RS91a, SC86, WK92]. Cyclicly [EW73]. Cylinder [Mer75a, Mil74b, MS84]. cylinders [HK86]. Cylindrical [PP93, PP95, SN99, KS79b, LK89].

D [CW99]. d’absorption [CD95]. d’Alembert [GS87]. Damped [GM91, ÁP88b, GL84, Tsu84]. damping [ÁP88a, GLY84, GLY88]. Daniell [BD78]. Darboux [Bra81a, Bre73, Bre78, Mil73b, You75]. Data [AP93, Ark90, BG88, Bar91, BK95, BV98, Cha70, Chi74b, CL71, Duc95, Eva96a, Hof96, Jos99, LPR92, MP97, MH95, Qui93, BG89, Bra81a, DL84, DW82, Hru85, Lui82a, Lui82b, MN86, Tsu89]. Death [Mak76, ILV89]. Deblurring [Car97]. Decay [AP89, APS93, BT91a, BHL72, Bra79, Bro83, CQ97, DJJ91b, FM98, HP89, Kat97, Kim92, Lew74, LP94, Lu99, Mor90, Qui96, Riv93, RR95, Sig74, dS97, DJJ91a]. Decaying [Dv96, NS89]. d’échelle [Her95]. Decomposable [Heg81]. Decomposition [Bus72, DS99, GM84, Riv93, Gin85]. Decompositions [CL96, DDMT93, HL77, MT95]. Decreasing [CLH92, CF98]. Dedication [AGIN94]. Defined [Dol72, Pal77, But79, Vic78]. Definite [BSX95, BRC72, Jus74, Rea83, Rea86b, AK88, Ha86, Log83b, Log83c, Log83d, Rea84, Sta78a, Ste72]. Definiteness [BY91]. Definition [LT99]. Deflagration [GS93]. Deformation [HH88, LA91, MrS87]. Deformations [IL96, Wei88]. Degeneracies [PK91]. Degeneracy [PFK95, Qi96, TG85]. Degenerate [AHZ93, And91, BBDU93, BK90, DGG98, DS81, EG96, FLLL89, HJ92, KL96, Lab85, Lab89, LB93, NP95, Ren97, Sho75, Suz95, Wai76, WO91, XA91, Xu92, You99, HM89b]. Degree [CSW80, GP84, LS88b, Sch76a, Sch76b]. Delay [Cao95, FW92, KY75, KTV94, KF90, Lou94, O’M71, Yan92, Cho76, Erb76, Gri85, GH87b, MPN89, Mem89, RR84, Ren84, Tud87]. Delay-Differential [Lou94, Cho76, GH87b, Mem89]. Delayed [Win74]. Delays [Boe95, HRX98, Pao87]. Dense [CK84, CK85, WW88b]. Denseness [Hof99]. Densities [HL74]. Density [LY96, Sim90, BH86, Kim87, NNB88, She86]. Dependence [Bro75a, Cha70, CC99, Han76a, Hof96, Ko95a, TW91, Art75, Bel81a, Bel81b, Sei87, Wil80a]. Dependent [AV90, APS93, AMS^{+}92, BCB82, BPL91, CN77, HY95, KM91c, Pey71, Wan80, AT89, BH86, Cos84, Dam87, GH87b, How85, Lan85, LC84, Mar81, Swi80]. Depending [CD96a, HS97, Tak92]. d’équations [Ali95]. Derangement [GIO90a]. Derivation [ADH90, BSZ77, BH75c, GB72, Gra72]. Derivative [EM94, Erd74, HS97, HO94, Osl70, Pal77, CS80, HS79, Log83a, Log83e, TF84]. Derivatives [BLN87, BR71, CD96a, Dav73, Fin74, JK75, LS88c, LS94, Osl71, Osl72, Osl73, Wol71, Amo89, Ban79, Gun73, LS79, Nev79, PP79, Pit80, Sac83, Tre79]. derived [KM86]. Descent [GB72]. Described [MM77b, PT97, GM79]. Describing [Hol93, MR98a, Sil81, CFV81, MR83a, Ren82, She85]. description [Rob89]. Design [CO92, Hol87]. Detected [Yag92]. Determinant [HL77]. Determinants [Gas73, Var88, Wil91]. Determination [AR98, Amo99, BV98, CR91, HL93, KS96, LVLH92, PB85, And88]. Determined [EP98, GM92, YS93, Pag86]. Determining [ABV96, Mia87]. Deterministic [Wan78]. Detonation [GS93, Wag89]. Detonations [LY95]. deux [Ali95]. Developable [Kre94, CL87]. Deviating [SS78, GL84, HW81]. Deviations [BH93]. Device [Ala95, Moc74]. Devices [Zha94]. d’explosion [Ali95]. Diagonal [Nak78, GH86]. Diagonalization [GT98]. Dielectric [Urb96]. dielectrics [Blo80]. Difference [Aom94, AF92, Dat81, FAS87, Gao93, GLP91, Han90, Hun73, Imm91b, KZ95, Mil70d, Mil71b, PR91, Vil92, Wim74, AH88, Bra79, CDMS81, CE89, DL91a, DL92a, Han89, Hen89, HKP87, Pat79b, Pat79a, Pet84, Vic78, Wie87, Zei80]. Differences [Ise85c, AM83, Ise85a, Ise85b]. Different [PP93, PP95, Chi74a, Koo82]. Differentiability [HK90, Bre82]. differentiable [Ha86]. Differential [Any77, AB72, AH72, BH75a, BK95, Ber72, BHL75, BP96, BVB94, BS74, BMP92, Bra71a, Bra71b, Bra72, Bra81b, Bro75b, BH76, BR71, Cam77, Cao95, CPS98, Cha76, Chi77, Chi73, CD91, Chu76, CvL92, Col70, Cos73, Cus75, D’A75, DT96, Dat81, DC85, Dun90a, Dun90b, Dun94, Edw93, Eli72, Erh73, ET71, EW73, ES75, EKLOO92, Fai77, FAS87, FH90a, FMZ93, Fuj92, Gao93, GL71, GR76, GT98, GL74, Gor75a, Gra88, GB72, Gra72, GP93, Gug71, Hal77, HLL72, Har94a, Har94b, Har77, HH72, Her90, Hir90, Hol93, How78a, How82, Hsi73, Hun73, Imm90, Jon76, KO96, KO89, KY75, Kar98, Kee72, KR76, Kry96, KS85a, KLY94, Lag72a, Lag72b, Lag73, LS71, LT92, LMPS90, Leu74, Lew86, Lim78, Lin94, Liu97, Lon73, Lou94, Mag87]. Differential [Mao97, Mar70, MW93, Mat70, McC97, Mil70f, Mil73b, MN75, Mil75, Mit75, MP98, MG70, Mul78b, Nag78, Nau76, O’M70, O’M71, OK94, Olv77b, Olv77a, Olv92, Pao72, Par92, Pog99, Rei70, RB71, Rot93b, RS76, Sap71, Ser75, SS78, SK73, Sha87, Sha91b, Sha76c, ST70, Sho70, Sho75, Šim84, Šim85, Sin75, Smi91a, St.84, Sta71, Tho77, Tre76a, Tro93, Ung74a, Vas79b, Wal71a, Wal76, Wal75a, Wil73, Wil72a, Win74, Wyr77, Wys93, Yok97, ALS87, BJL81, BJL88, Bar79b, BS79a, Blo80, Bog74, BD86, Bra91, Bra79, BHS83, But80, BE83, BR84, BE86, Cam87a, Can85, CH79, Cha79, Cho76, Chu80, CLP88, Dav79, DG89, DR86, Eli84, Eng82, Erb76, ET72, Gau90, GH87a, GS89, GL84, GLY84, GLY88, Gra80, GK84, Gun73, GH87b, HK86, HW82]. differential [HW84, Har70b, Har70a, Har83, Hea82, HW81, HI89, HS86a, Hir83, Hir82, Hir85, Hsu83, JAK88, JW80a, JSW87b, KS86, Lei81, Lou80, MPN89, Mar86, MR83a, Mar78, McC78, MO83, Mem89, MM79, Min86, Mul81, Nam78, ND81, Oja78, Olv93, OP87, Pao80, Per81, PR87, RW78, San83a, SS80, Sch80, Sch84a, Sin80, Smi84, SM85, Spi79, Sta86a, Sta88, Ste80a, Ste85a, Ste84, Tef78, Tom81, Tom83, Tom86, TW79, Tre84, Tre85, Tre88, Tro81, Tud87, Vas79a, Vas81, Web79, Wei84, Wes78, Whi80, Wil71, WL82b, WM87, dW80, van86]. Differential- [Gao93]. Differential-Algebraic [KO96, MW93, OK94]. differential-delay [MPN89]. Differential-Difference [Dat81, FAS87, Bra79]. differintegral [ABM87, Si78]. differintegration [Cam89a]. differintegrations [Cam89b]. Diffraction [NB75, Nay77, NS91, Spe89, SHM89]. Diffractive [BC97]. diffusing [HSX87]. Diffusion [Ala95, BCB82, Ben85, BS85a, BS85b, BLS99, CN77, CJM98, EM90, ET94, EMS98, EZ97, EH90, FHM97, FS96, FRZ95, FH96, Gar92, GGI92, Gra98, HS94, HG99, Hu91a, HY95, Hut86, IM91, Ko95a, Ko98, KoM98, Kel73, Kuw96, Lan74, LS96, Lia94, Mor90, MH95, MNY94, NMIF90, PS97, Qi96, Red85, RB98, Sai96, Sch98, SW91, Sou98, Tak92, TN94, Wol93, Yam90, AO81, AA88, AA83, ARRR81, BH86, CE85, CE87, Cos84, Dam81, DS87, DH84, Feh88, FL87, FH85, Gar81, HMP87, HM89b, Kha87, KA88, Lan85, LPSS89, Mar89a, Mem89, Nis82, NF87, NNB88, Pao87, Sec88, She86, Ter83, van79, van86]. Diffusive [CC91, NT86a, NT86b, NT86c]. Diffusivity [LP94]. Dilated [CS97b, Mad85b]. Dilation [Eir92, LW95]. Dilations [Mas78]. Dimension [ANS96, BR78b, Car95b, GM91, Gru92, Lop89, Tri97, Wei99, Ali95, Bel81a, CGR88, Ghi86, Mar85b, OP87]. Dimensional [Av91, Ala95, AM95, AMS^{+}92, BO91, BBN93, BSX95, BMS83, CvL92, Dau89, GHL96, GM93, GHLS91, Gor95, Gra95, HW89, HT98, HL94, Hir90, Hof96, Hof99, Kal75, KN97, Kim90, Kim95, KI95, Kli97, Kuw96, LA91, Liu91, Maa96, MW75, MW78, MW91, MR98a, Mil70a, Mil70b, NS98, Pes95, Sac74, SR93, Ste91, Sun93, Ves97, WW99, ZZ90, Zha94, AL78, Ali95, Azz80, Azz81, CD80, Coc87, Cur79, Daf82, FSV99, HV87, Kim83, KS80b, Lar88, Mar89a, Mos83b, Nas84, RV87, Sym86a, Ura84, WM82, Ful83].

Dimensionality [GZ92, MM80]. Dimensions [AB93a, AB93b, Daf95, Dou97, EZ97, GL71, MR95, PS91a, Sak93, Yin91, GM83, HP89, Lin86, Smi89, Wag83]. Diminishing [Hor75, vdW75]. diodes [BCG89, Mar85a]. Dirac [BG91, EV97]. Direchlet [Sig76]. direct [Sta88]. Directionally [How85]. Dirichlet [AB93a, Azz80, Azz81, BT91a, BR78b, Car74c, Car87, CC94, EL89, FMT96, GK77, Hsi78, IM98, Ish96a, Jan96, Jer78, Jia88, KK92, KL96, Kla71, KJ82, LS88a, LS89, Mha88, Mos83a, Mos83b, Osh93, Pin88b, Pos90, Ren90, Ros71, Slo75, SN99, Yam90, Zag98, Zha95b]. Dirichlet-to-Neumann [IM98]. Dirichlet/Neumann [BT91a]. disappearance [Die88]. Disc [Wol71]. Disconjugacy [Kim86, Wil71, Wil72a, AH88, Eli84, Tro81]. Disconjugate [Gra88, Har77, Gra80]. Discontinuities [PS91a]. Discontinuity [CK91, Red84]. Discontinuous [CS97a, CL71, Die95, EI97, Elc74, GM93, GR92, Gim93, Hof96, Kel88, Jer78]. Discrete [Ahl93, AEU94, FS89, GHLS91, HP93, MMM97, Mor85, Ros98, SV94, Szw92a, Szw92b, Wal92a, Wil72b, Zin91, AH88, Dri87, HKS91, Rah81a]. Discretization [FD97]. Discretizations [Li97]. Discretized [Lub86]. disease [WL82a]. disfocality [Eli84]. disk [Bar84a]. Disks [FKS99, Lu90a, LKMT88]. disparition [Die88]. Dispersion [HFL99, Kot94, DJ87]. Dispersive [Col94b]. Displacement [AM95, FL92, Fox72, Rei93, Riv93]. Disrupted [CC91]. Dissipation [AL98, Dv96, Lip98, PS97]. Dissipative [Kov95, Lab99, Klo85, Mar89a]. Dissipativity [Thi93]. Dissolution [AHIR93]. Dissolution-Growth [AHIR93]. Distributed [DS95, Fer86, HRX98, Ack81, CG87, Jia88]. Distribution [Coh75, Fai77, GG77, GW94, IK76, Mei76b, Van85, Van01, Wol71, Aro81, Chi80, NNB88, Ric89, Sil85, Tem82, Wei71, Whi80, vd86]. Distributional [Car73b, Car74d, CS77a, DdS74, Liu96, Mey74, Mey77, MP97, MK78, Wal75b, CR83]. Distributions [Dol70, Dol72, Dom72, EM70, FI72, Jan81, Koh70a, Lee75, Nav99, Ort73, Pri75, Sri90, Sri72, Car79b, Car86, Jan84, Mei79, Ort80]. Disturbances [Sun93, WY75]. Divergence [AM94, CF98]. divergent [FS83, Sid86]. dividend [DS82, SDL84]. divisibility [IK79]. division [GH87b]. divisors [GLR79]. Do [Kat97]. Dolezal [Mei72]. Domain [BR78b, DS99, EY98, Häh98, JM96, KK97, LP94, MA88, Urb96, Ale80, Dri87, Gri79a, Hom79, Wex79, Wex80]. Domains [AGJL92, Avr91, Avr93, Ban92, Ber99, CZ91, Col94a, Dau89, FHX96, Hil98, Jan86, Kai91, Leu73, Leu74, Leu75, LP91, Mei76a, PS76, Sak93, SN99, Was70, Azz80, Azz81, Coh89, Cos88, DV87, Lon88, PP83]. Dominance [Nak78]. Dominated [PP93]. d’onde [Ali95]. Double [ADH90, Ber75a, BLM96, Car75b, Dun90b, ES97a, NV95, Sei91, She80a]. Double-Obstacle [ES97a]. Double-Porosity [BLM96]. Doubling [PK91]. Doubly [GS77, Ish96b, Leu74, PS76, BF88, BF89b, NT86a, NT86b, NT86c]. doubly-diffusive [NT86a, NT86b, NT86c]. Down [SS75]. Draining [BP96, Tro93]. drawing [AR87]. Drazin [HS79]. Drift [Ala95, Lia94]. Drift- [Ala95]. Drift-Diffusion [Lia94]. Dual [BAR90, FN75, FK78, Fei78, KF74, Pan72, Pen92, Sri72, Wal75b, Zha92b, ILV89, Wal79]. Duality [Bro75b, Fis75, Tha93, TW89, Leo82]. duals [Per84]. Due [GP99, Gla76, KG71, Lin78]. d’une [CD95]. Dunkl [Rah79]. Dyadic [CL96, Lan96, Pal77]. Dynamic [Che92c, Fan97, GY90, Gor95, NR95, Eng87].

Dynamical [ABH94, Aus77, BBN93, Boe95, Den95, Gru92, HP93, HM84, JPS87, LYC95, Day82, Fuh80, Rob89]. Dynamics [BBL97, BHN72, Col97, DDD96, EM90, EY98, HY95, KoM98, KM91c, LD97, MR98a, Mis95, Noh74, Rua98, SR93, Sen97a, Sen97b, Yag92, ZZ90, Ber88, BH86, HM89b, Mar81, Pao80, Sch87b, Swi80, Wex79]. dynamo [Bar84a]. Dyson [And80, Zei87].

E. [Gre74]. ear [FI90]. Eberlein [Rah78]. Ecology [DL92b, Hsu83]. Edges [Ban92]. Editorial [Ano98]. Effect [BPL91, Liu95a, Yin98, Mar85a]. Effective [JRRV97, Paş89]. Effects [Col94b, FMP98, LL93, LP94]. efficient [MM80]. Eigenfunction [FH90a, Hin81, Nay78, Ful82, Ful83]. Eigenfunctions [AM94, SK73, Pin85, Sac83, She88a, She93]. eigenparameter [Hin81]. Eigenvalue [AB93a, Bal99, Ban75, Bar77b, Bar85a, Bar88, Bar91, BK95, BB77, CS78, Dea71, Edw74, FS77, Fra90, Gre89, Kel73, LB74, Lip98, MW83, McL76, McL78, Par70a, Par70b, She80a, SS92, BB78, CS86, Chu84, DiG81, Gri85, Hin81, Kol79, LL79, Mar85a, Sen85, She88b, Sil85, Val86, WM82]. Eigenvalues [AFT98, AB93b, Bar85b, Bax71, BR78b, BBF99, Car99b, CD91, CK95b, Dri87, Fai77, Fox72, Gre91, Ha86, Hir75, Kar98, KLV93, KS78, LB74, LR84, Log84c, MR80, Pap95, Pin80, Rea83, Rea84, Rot93b, SL78, Wil73, Auc89a, Auc89b, Bar87, Chi83, Har80, Rea85, Rea88, Wei71]. Eigenvectors [Sap71, Sil81]. Elastic [BBN93, DGR99, Fra90, HL93, KLV93, LA91, LSZ92, NU95, PP93, PP95, SS93a, Wol96, Lef82, Lun87, PGRSV89, She85, Sym81, Wei88]. Elastic-Plastic [SS93a]. Elasticity [FH91, CDHL88, DV87, How85]. Elasto [NR95]. Elasto-Plasticity [NR95]. Elastohydrodynamical [CL90]. Elastohydrodynamics [Hu90].

Elastoplasticity [Gor95]. elastostatics [Coh89, Rog88, Str80]. Electrical [Mat78, Zem79b, Kaw79, Zem79a]. Electromagnetic [AN98a, CWZ99, CGP93, CP90, CD90, Tor96, Yin98, CK85, Zio85]. Electromagnetics [KN97]. electromagneto [Rog88]. electromagneto-elastostatics [Rog88]. Electrons [KG71]. electropaint [CF85]. Electropainting [CO88, MS87]. electrophoretic [Avr88]. Electrophotography [HW92]. Electrostatic [BV92, Van89]. Electrostatics [FR86a, TW89]. Elementary [BI75, GRZ83, MW74, Pip93, Saa74, She85, Cos88, MQ85, Rah79, RV87, Vre84]. Elements [Koe94, CT88, Koo85]. Eliminating [Gu92]. Ellipsoid [Sha91a]. Elliptic [AP93, AM94, AVV90, AVV92, BG88, Ban92, BMP92, BR78b, Car77d, Car78c, CG85, CG94, CS77c, CK91, CL91, CD74, CvL92, Col72, Col73, Cus71, DH99, Elc71, Elc74, Fet70, FSW96, GHL96, GL71, Gom86, GP93, Hil98, HS95, Jan86, JS96, KS78, LP76, LL93, LP91, Lin75a, LB93, Pay70a, QV96, QVV98, Rao75, Saf99, Sch73a, SS77, TF75, Tse97, Val94, XA91, Yu74, Yu77, Do 93, AW78, Azz80, Azz81, AK82, BG89, Ban78, Ben86, BB86, BB79, Bud89, CV87, CS86, CG87, FG87, JK86, Kam79, Kui80, Lan80, Mil84, NS89, PR87, Pin88b, Sch87a, Spe79, Swe89, Tch96, Wie87, Yos83, ZW85, dM86, van86]. Elliptic-Hyperbolic [CK91]. Elliptic-Parabolic [CvL92, HS95]. Elliptical [HP98]. elliptiques [Tch96]. elongation [Ren82]. Embedded [KLV93]. Embeddings [AJV94]. Emden [Iff87, LP81]. emerging [SAW87]. Encapsulated [IW98]. Encapsulated-Vortex [IW98]. End [LP94]. Endemic [Thi93]. Energies [FL98]. Energy [BT91a, CN77, CS97b, DG83, FD97, Fuj92, HO94, Jos99, Kim92, Lip98, MP97, Mul78c, PP93, PP95, Shi74, Vil92, MS86, Sta78b]. Eneström [ASV81, Rus78]. Engineering [Jam98]. Enskog [Ark90]. Entering [RB71]. Entire [Joh87, KS91, LP76, Log85]. Entrainment [Chi92, St.88]. Entries [BS96, Har80]. Entropy [Cla83, DGG98, Egg93, GG96, BH85, Sto80]. Entry [GM94]. Environments [CC91]. Enzyme [HJ92]. Enzyme-Catalyzed [HJ92]. Epidemic [BIT91, HY95, IMP92, Wan78, Wan79]. Equ [KF90]. Equal [KJ82]. equality [RV87]. Equally [CR92]. Equation [Ada78, Ahl93, Av91, And96, Ark90, AD86, BG88, BNS98, Ban74, BS95, Bao98, Bar77a, BBDU93, BD95, Ben96, Ben81, BFF^{+}90, BS85b, BCT98, Boy76, Bra71a, Bra72, BB97, Bra91, Bra75, Bre73, Bre78, BS97, BRC72, CS97a, CN77, CPW98, Car71a, Car76a, Cha76, CS77c, CJM98, CGP93, CvL92, CK92, Cus75, D’A75, DGG98, DDD96, DC85, Dek71, Dem96, Den95, Det78, Dix96, Ede79, EY98, EI97, Elc71, Elc72, EG96, ES97a, Eng99, ET94, EZ97, Eva96a, FAS87, FI96, FD97, Gal91, GV93, Gal95, GO71, GM93, GH79, GK77, GW97, GB72, Gra72, Gra98, Gri79b, Gri87, Hab76, Hal77, HK90, Han73, Han74a, Han74b, Han76b, Han77, Har77, Hay91, HNO98, HH72, Hsi73, Ish96b, Ito94, Joh75, Joh77, JKO98, Kai91]. Equation [Kal75, KM78, KV91, KY75, Kif77, Kif80, KL96, Kim98, KV99, Kra74, Kui79, Kuw96, LMPS90, Leg76, Lim78, Liu91, Liu95b, LMN98, Log84c, Lon71, Lon73, Lon74a, Lon77, Lon87, Lu99, Lun90, Lun78, Mac90, MR95, MLM97a, McL90, MM77a, Mer75a, Mer76, Mey74, Mic92, Mic96, Mik88, Mil76, MP97, MY98, NS91, Nor88a, O’M70, Ort94, PS91a, Par75, Par92, PT97, Qi96, Qui96, RS94b, Red85, RZ91, RDLH77, Rot93b, SC86, San98, Sap71, SK73, Shi74, Sho70, Sho72, Sig74, Sig76, Sle95, Slo75, Sta80a, SU97, Str76, TBD^{+}96, Utz70, Ves97, Vil92, Wai76, Wal76, WW98, Wim74, Wyr77, You75, Zha99a, Zha92a, Zha95a, Zin91, dGdJ75, dHW73, Do 93, Akı88, ÁP88b, ÁP88a, AK85, BH87, BF89a, Ben84b, BH86, BF88, BF89b, BW83]. equation [BKW85, Can85, CH79, CH82, Chi74a, CL86, CW89, Chu80, CS89, CK87, CE85, Dam81, DS82, Dol80, DR86, EV97, FIT87, Fit82, Fri81c, FG87, FH85, GT89, Gri78, GH87b, HT87, Hea82, HV87, HI89, Hoc84, Hru85, HR88, HW83, Iwa88, JK86, JSW87b, Keg83, Kha87, KS80a, Kri80, Lar81, Lef82, Lei81, Lev88, LPSS89, Log83f, Log84a, Lun87, MPN89, MR83a, Mem89, Mil79, Min86, Moe79, Nas84, Nor88b, Pet84, PR87, Rea87, Ren82, Sac83, Sen97a, Sen97b, SDL84, She88a, SP79, Sta78b, Sta85, Sta86a, Ste80a, Sym86b, Tad86, Ter83, TW79, Tre88, Tro87, Tsu89, Web79, Wes78, Wex80, WW88b, Yos84, Yos86, BH87]. Equations [AGJL92, AP93, Ala95, AM94, AHZ93, And91, Any77, AP92, Atk74, Aye83, AB72, AH72, BC90a, BH75a, Bar75, BO91, BF98, BK95, BW73, BT91a, BB92, BHL75, BP96, BK90, BO92, BVB94, Bil90, BS74, BMP92, BS76, Bra74, BHN72, Bro78, BC90b, Bry74, BH76, BR71, Cam77, CC91, Cao95, Car96, CP97, Che92b, Che96, Chi77, CL90, CL91, CD74, CD91, Chu76, CMOR86, CJ88, Col94b, CM80, Col70, Col72, Col73, Col74, Col75, CW99, Cos73, DT96, Dat81, Dau89, Dea71, Den91b, DS81, DH99, Dun90a, Dun90b, Dun94, Edw93, Egg93, Eir92, Elc74, ES76, Eli72, Erb77, Erh73, ES75, Ewi75, FI90, Fai77, FMZ93, FSW96, FS96, FSV99, Fuj92, FR91, GL90, GM91, GGI92, GM96, GL71, GR76, GT98, GL74, GM76]. Equations [GM92, Gor75a, GP93, Gug71, GLP91, Hal77, HLL72, HJ98, Han90, Han75, Han76a, Har94a, Har94b, Har77, HO94, HL96, HRX98, Her90, Hil97, Hil98, Hir90, Hir72, Ho74, Hof96, Hog98, Hol77, Hol93, HS95, HY95, Hun73, IW98, Imm90, Imm91b, JS89, JM96, Jon76, JT94, JRRV97, JW74, Jua95, KO96, KM85, Ko95a, Kap72, Kar98, Kee72, Kel79a, KO92, KS79a, KR76, Kni72, Kol74a, KTV94, KZ95, Kry96, Kui85, KM95, KS85a, KS85b, KS91, KK97, KLY94, Ład94, Lae75a, Lae75b, Lag72a, Lag72b, Lag73, Lag75, LS71, LR77, LT92, LW95, LMPS90, LS96, Leu74, Lev76, Lev74a, Lev74b, LM75, LP76, Lew86, LB96, LP91, Lin91, Lin94, Lin95, LB93, Liu97, Lou94, LW79, Lu90a, Mag87, MM97, Mao97, Mar70, MW93, Mat70, McC97, DMV98, MR98a]. Equations [Mil70d, Mil70f, Mil70e, MF71, Mil73b, MN75, Mil75, MM77b, Mis95, MP98, MG70, Mul78c, Nag78, NW74, Nau76, NS91, Nii97b, Noh74, Nus78, O’M71, OK94, Okr91, Ols74, Olv77b, Olv77a, Osh93, Pao72, Par70a, Par70b, Pay70a, Pen92, Per73, PR91, Pey71, Pog99, Pra72, Prü88, Rao75, Rei70, Rei93, Ren91, RB71, RS76, Saf99, Sat77, Sch73a, SW73, SW75, Ser75, SS78, Sha87, Sha73, Sha91b, ST70, Sho75, Šim84, Šim85, Sin75, Sin76, Smi91a, Son70, Sou98, SV94, Sri72, Sta71, Ste75, Stu96, Suz95, Tak92, TF75, Tre76a, Tro93, Van98, Vas79b, Veg93, Vil94, Wal71a, Wal75a, Wal75b, WW93, WZ97, Wil72a, Win74, Wys93, Yam90, Yin91, Yin98, Yok97, Yu74, Yu77, Zac74, Zag98, AKM87, AH88, ABM87, AA88, Ali95, Ame82]. equations [ALS87, Art75, ARRR81, Azz80, Azz81, AK82, BJL81, BJL88, Bar79a, Bar79b, Bel81a, Bel81b, Bel87, BS79a, BB86, Blo80, Bog74, BD86, Bra81a, Bra79, BGT88, Bud89, BHS83, But80, BE83, BR84, Cam87a, Cam87b, CDMS81, Cha79, CL81, CE89, CG87, CLP89, CN79, CN81, CS85, Cos84, CLP88, Cur79, Daf82, DV87, DL91a, DL92a, Dav79, DG89, Die86, Die88, Dol79b, Eli84, Ell82, Eng82, Epp86, Erb76, EW89, ET72, Fie86, FL87, FR86b, Gau90, GH87a, GS81, GL84, GLY84, GLY88, GK84, Gri79a, Gri80b, Gri80a, Gri81a, Gri81b, GM79, Hag82, HK86, Han89, HW82, HW84, Har70b, Har70a, Har83, Hen87a, Hen89, HW81, HI89, Hir83, Hir86, Hir82, Hir85, Hiv78, HS84b, HKP87, HM89b, HJ83, Hsu83, JW80b, KM80, KM81, KM83]. equations [KT79, Kaw79, Kel88, Kim83, KM86, Klo85, KKY84, KS80b, KS86, KA88, Lar88, Lei85, LC84, LM80, Lev83, Lon83, LKMT88, MM80, Mar89b, Mar78, McC78, McC81, MO83, Mer87, Mil84, MM79, MB91a, MB91b, Nam78, NS85, Nip86a, NF87, ND81, NS89, Oja78, Olv93, OP87, Ort80, Pat79b, Pat79a, Per81, Pit80, RW78, Rei86, RS78, Sac85, San83a, Sch87a, SS80, Sch80, Sch84a, SW79a, Sei87, She85, Sho76, Sin79, Smi84, Spi79, Sta88, Ste85a, Ste84, Ste73a, Swa82, Tre84, Tre85, Tro81, Tsu84, Tud87, Tur89, Vas79a, Vas81, Wal79, Wei84, Wei85, Wex79, Wex80, Whi80, Wil71, WL82a, WL82b, WM87, Yos83, dW80, vP86, Die86, Die88]. Equiconvergence [Muc70, Wim74]. Equilateral [Pin80, Pin85]. Equilibria [BM92, CS99, Den90, DP97, GJ82, Kru90, FH85, LC84, ML93, MS88]. Equilibrium [And96, BRC72, JS96, Kel73, Liu95a, MR98a, Ort94, PGRSV89, TW97, BH86, BGS86, NT86c, NT86a]. Equipartition [DG83]. Equivalence [BHL75, KO89, MR79b]. Equivalent [KR76, BKM86]. Equivariant [IW98, Mir94]. Erdélyi [Tem76]. Ergodic [JPS87, Nus90, Wys93]. ergodicity [Bra81c]. Errata [Ack80a, Gro77, Leu75, Log84a, Mad88, Old72, WH88, Wil72a]. Erratum [BF89b, Bra72, Car86, DJJ91a, EMN94, Log83a, Pay74, Ski81, Tah91, Van01]. Error [Car76a, Car97, FKS99, Fre87a, Fre90, Fre92, Het70, KT97, LW94, Luk75, Olv74, Thr94, Van96, Vit79, Wal97, Won76, MS79, Son80]. Errors [JKK96]. Essential [Fox72, Bea85, Stu89]. Essentially [TBD^{+}96]. Estimate [BMP92, BL74, ES76, PR71, Qui96, Rea85]. Estimates [AGJL92, AP89, APS93, And91, Ban78, BF98, BH93, BHN72, CQ97, Daf87, DGG98, DDD96, ET74, FKS99, Far78, Hil97, Hil98, Kar98, KT97, Liu97, Noh74, PP93, PP95, Pit80, QV96, Sch73a, Sch88, Ste77, Ves97, Wei99, BF89a, HP89, Mur83, Rea88, SM85, Ste85b, Nip86b]. Estimating [vdS76]. Estimation [FS77, KSS97, MW78, Ben91, CS80, DW82]. Estimations [LMW96]. états [BH87]. Étude [BH87]. Euclidean [BGG84]. Euler [BT91a, Bra81a, Bre91, Bre73, Bre78, Car74a, FSV99, Li99, Mil73b, Mul75, Tur89, You75]. evaluating [PV84, Piq89, Sim80]. Evaluation [BH75b, Ber75a, Car72, Erd74, FI72, LB74, Luk70, Old70, Old72]. Evaluations [GS82]. Evans [Kap98]. Evaporation [HP92, AN98b]. Even [Bal99, Fia80, GL84]. Eventually [Log83b, Log83c]. Everywhere [HT71, Muc70, Pog99, Hir85, Sid86]. Evolution [Bra74, Car99a, DDD96, DS81, FD97, GPG91, HS98b, KM85, KO92, Shi93, Sho75, Van98, Vis86, Zei94, CF85, CO89, Epp86, Hil82, KA88, Nau79, Ren84, Rob89, Sac85, Sei87, Sho76, Wex80]. Evolutionary [CK95a, Dat72, DL92b]. Evolving [GM96]. Exact [HK86, Sch88, Son80, You99, Wie87]. Examples [BÇ94, BMS83, Dor70]. Exchange [BBDU93, Her91]. Excitation [Kra75]. Excited [Tur96]. Exhibiting [Pos90]. Existence [Ada78, AGX89, AC94, Atk74, Ban74, Bar79a, BCT88, BFGJ97, BBN93, BGS86, BMP92, BH99, Bor99, Bra91, BSL87, BRC72, Car95b, Che96, CL86, CL90, CO88, CE87, Cos87, DGG98, Den91b, DiG81, DL92b, Elc71, Elc74, FS95, FD97, GM93, GGI92, GM96, Gio93, GW96, HL98, Har72b, HZ94, HH72, Hen89, HS96, IW98, Ish96b, Jer78, KH95, Kel72, KO92, Kra74, Lak90, LP91, LKMT88, Lu90a, MLM97a, MM77a, Mor89, MH95, MW74, MX98, NNB88, NR95, O’R93, OFK77, Ols74, Par72b, Pet81, PT99, PD71, Ren90, Ren91, RR95, San83a, SS89, Sch84b, Sho72, Sim90, Sta71, Tip74, Ton81, TW91, Wal75a, WW93, YL94, Zha99a, Zha94, Zha99b, Avr88, BN78, BB79, CD80, CDHL88, Che92a, DH84, DK86, Far78, Fit82, Gab84, Gri78]. existence [Gru85, Han89, Hir83, HMP87, Hru85, Keg83, Klo85, Kor88, Lar88, LK89, Min86, Ton82, Tro87, Wag89, DL91a]. Existence-Uniqueness [Tip74, Pet81]. Existing [Ito94]. exit [Day89]. Expanding [Kol74b]. Expansion [BG91, Car74b, CG85, Coh76, FH90a, Fre87b, Fre90, Fre92, Gor75a, Hab76, HL70, HL74, Lam78, Mac90, MW87, MR80, NB75, Nay78, PS72, Ros78, Ski94, Vol90, Won80, Dun74, Kam89, SW79b, Son80, Son83, Tem79, Tem82, Tem87, Vol87]. Expansions [AZ79, Ask71, Ble77, Car77c, DL91b, GT94a, GG77, HB73, JK75, Joh87, KM91a, KM91b, Kem85, KSK78, Lan83b, LW94, Lo72, LT94, Mey86, MX91, Muk72, OT94, Olv91b, Olv91a, Pil86, Sha76c, She72a, She72b, Ski80b, TF75, Tem90b, TF84, Tre73, Tre76b, Tse97, Van96, Won76, AB80, Dil88, Dun86, Dun89, Fie73, Fie83, Fre87a, FW88, Ful82, Ful83, GK84, Hin81, Kel89, Kem82a, Lan80, Lou84, Mar83a, Mha88, Rap82, Sid85, Ski81, Ski88, Ste84, Ste72, WZ87, Won78, ZW85]. Explicit [AGGO94, BR82, Car72, Cie87, ES76, FTZ88, GR75, Mia93, PJ73, Sch73b, Wyr77, FIT87, Sac83, Wim85]. Exponent [Dun90b, LP81]. Exponential [AL98, DJJ91b, Day82, Ehl73, Joy97, Lag72a, Luk75, Mao97, Olv91b, Sig74, Car79b, Car86, DJJ91a, HIN85, Log85, LS93, Mos83a, Mos83b, WG87, Wes78, Zuc84]. Exponentially [JKK96, Olv91b, Olv91a, Olv93, Tol98]. Exponentially-improved [Olv93]. Exponentials [LJ91]. Exponents [CP97]. expressible [GR85]. expression [Piq86]. Expressions [Mit75, KKZ86]. Extendability [Ber70]. Extended [BGG84, Pas74, PT97, Sta85]. Extendibility [Mei76a]. Extensible [Yos85, Fit82]. Extension [AL81, BHL72, CA71, LA96, LN98, Nii97a, Osl72, Par75, Rus88, Sch76b, Šim88, AL78, ASV81, San80]. Extensions [All79, Car70c, GA98, Hei75, Kla71, Mer75b, Sle72, Ask80, Gas89, KKZ86]. Extent [Olm77]. Exterior [Häh98, Hsi78, Mil76, Hom79, Mil79]. External [CGP93]. Extrapolations [Wel91]. Extrema [WZ94, Whi80]. Extremal [AL77, Bar85b, Bar85a, Bar88, Dör91, GR78, Log83b, Log83c, Log83d, Ack81, WM82]. Extreme [Bax71]. Extremum [PD71, RS91a, GR79].

F. [Vol87]. Faber [Ban73, Joh87]. factor [Dol81a]. Factorable [Bra81b]. Factored [Ste84]. Factorial [DL91b]. Factorizable [FTZ88]. Factorization [GW97, Gra88, HL77]. factorizations [Gra80]. factors [Sid86]. Fading [Lun90, JSW87a, JSW87b]. Falkner [Gab84, HT87]. Families [Col75, KMT91, Maa96, Pie73, Rah81a, RTV96, WB95, Hal84, MB91a, MB91b, Per84, TW89]. Family [BM92, Hor75, Li85, Rah76a, Rah76b, VS79, GH84, WL82b, Yom83, Yom86b]. Far [CKP89, CP90, CH93, CK95b, Col82, CK84, CK85, WW88b]. Far-Field [CP90, CKP89]. Farka [CK77]. Fast [Kat97, Qi96, Bar84b, CL87, Fen83b]. Fast-Diffusion [Qi96]. Feedback [BT91a, Dol81a, Pao83, Pey75, Sel81, van86]. Feedbacks [Van98]. Fejér [CB87, Por76]. Fermi [BG91, Lun78]. ferrofluid [TT83]. Feynman [CS89]. Fiber [BB90, Paş89]. Fiber-Reinforced [Paş89]. Fibers [Che92c]. Field [BFGJ97, CP90, CH93, CK95b, HT71, Pal77, Pér93, Riv93, SS96, Stu96, Wol96, Yin98, Col82, CK84, CK85, CKP89, Fri81c, WW88b]. Fields [AN98a, GP99, KG71, MMM97, ML93]. Fifth [Chu76]. Filament [NT96]. filaments [Ren82]. Filippov [Sha87]. Filled [LF98]. Film [BC95, DDD96]. Filter [Aus77, Log88]. Filters [AEU94, CS93b, Mor85, Wel91]. Final [BB92, Car76a]. financial [SDL84]. Finite [Bai71, Bar91, BK95, BL90, BCT98, Car74d, Coo74, CC96, GW97, HO94, Ise85c, JK75, KS96, LPR92, MW78, Mar89a, PB85, Per73, Chi80, Cur79, Ise85a, Ise85b, Lon83, MR79b, Ral82, Sta78b, Wie87]. Finite-dimensional [Mar89a]. Finite-Order [CC96]. Finite-Time [BL90]. Finitely [BV92, GT98, Mai94, Zha92b]. First [Bal99, Bar77b, Bry74, Car77d, CG85, Egg93, Gra72, Hol77, Lor93, DMV98, Mie87, SW73, SS92, Ski94, Spe89, TKS93, Tol98, Yu77, Ame82, Chi83, Cof86, KM88, MR83c, McC78, Nau79, Sac84, Wei84, WM87]. First-Order [Bal99, Ski94, Tol98, McC78, Nau79]. Fisher [PT97]. Fissured [BS85a, SW91]. FitzHugh [Che92c, Den91b, Flo91, KM86, Nii97b, San98, CB89, SS96]. Five [Rah76b]. Five-Parameter [Rah76b]. Fixed [HP93, Jos99, Kug74, MP77, Par70b, St.96]. Fixed-Point [Kug74, Par70b]. Flame [Mic96, BSL87, Lar88]. Flames [BH99, Gio90b, Mic92]. Flat [Sat90]. Floating [HL93, Liu91, Lu90a, LKMT88, LK89]. Floating-Body [Liu91]. Flooding [TW91, JW88, JW89]. Floquet [Ber76]. Flow [Ack77, AP89, APS93, ADH90, AK99, BBDU93, BC95, BPL91, BLM96, BB94, Bra91, BS97, CQ97, CvL92, Elc82, EMS98, FT99, GT94b, Gio93, Gla76, GS86, HLM92, Her91, Hof96, HR95, HS96, KT99, Lu97, MA88, NT96, RJ86, RS95, Sha76a, SZL97, SS75, Zha99b, Ack81, ACK83, Coo88, Eng87, GK80, She83]. Flows [BP96, Dou98, Eva96b, Gil79, HL98, HFL99, Id94, Kra91, LA91, Li97, LN98, Mir94, SS95, Tro93, Ber88, Ghi86, Gri82, Gus85, HS89, LK89]. Fluid [AC98, Col97, Fan97, Kim90, KT99, LF98, MZ98, MR98a, NP93, Sch97, Sha76a, SLS81, SS75, SW91, SS93b, Zaj94, Ber88, Ghi86, Gri82, Kim87, Lei81]. Fluid-Dynamic [Fan97]. Fluid-Particle [SLS81]. Fluid-Solid [AC98]. Fluid/Structure [MZ98]. Fluids [BPL91, Ish96a, MA88, Ren96, Sim90, Sec88, Ton82, van79]. Flux [BCB82, Die95, GR92, Gim93, LWY98, LC84, SAW87]. Flux-Dependent [BCB82]. Fobenius [Mul75]. focal [Ira89, Pet81, Tom81]. focal-point [Pet81]. Fock [Cho84]. Fokker [JKO98]. Folding [CRSW76]. fonctionelles [Tah91]. fonctionnelles [Tah90]. fonctions [Her95]. Föppl [Mul78c]. Force [CGP93]. Forced [Ruf95, Won88, Yag92, Yos85, FR86b, GH84, KT79, Sch87b]. Forcing [FSW96, St.88, Sta85]. Forcings [FR91]. Forest [Gre74]. Form [AM94, Luk75, Sai96, Wal97, Cam87a]. Formal [Imm90, JT79]. Formalism [Jaf97a, Jaf97b]. Formation [ABH94, EM90, Hol93, MMN85, Nak88, Veg93, Do 93]. Forms [Har94a, Imm91b, Ral83, Kol79, Mad85a, Mad88, Mar80]. Formula [BSZ77, BJL98, BR82, Dur77, Hug75, Koe94, Koo74, Koo75, Koo77, Koo91, Mia93, NC82, Osl72, Sha76b, Sle72, Sri90, VK91, BD78, Ben91, BG96, Büh88, Cie87, Gas85, Gas89, GR89, GN83, GS87, JS87, Kur79, Lai80, Lai82, Mar86, Nay87, Nev79, Peh84a, Sin80, Wim85]. Formulae [BSX95, FTZ88, MP77, Wil79, Xu94, Xu98, Coh89, VJ85]. Formulas [BF98, Car70b, Car77c, Car99b, Coh76, Dur78, Gas81, HQW74, Leu73, Leu75, Olv77b, Olv77a, Pap95, Peh81, Sei91, Vre84, Bar87, Bre81, CH82, Dur79, ES84, FLLL89, GR84, Gas85, GR86, GR89, Kow82a, Kow82b, Lai82, MZ92, NR85, Nay84, RV86, Sta80b, Tre79]. Formulation [Bal99, Joh75, JKO98, Osh93]. Formulations [LVLH92]. Forward [BBDU93, Dem96]. Forward-Backward [BBDU93, Dem96]. Foundation [Mei76b].

Four [Col72, McL76, McL78]. Fourier [BHL72, Car74d, CLH92, CF98, ET74, HT71, Lan80, Lew74, Lip72, Lip73, Mar84, Ort73, Rea86a, Red84, SS84, Sen96, SW79b, Str77b, Wil79, de 78, vdW75]. Fourth [Bro78, CPS98, DGG98, EP97, Eng99, Gre91, Kee72, LS71, Mil75, Par70a, Par70b, Ben86, DiG81, Har70b, Har70a, Mil84, Sch87a]. Fourth-Order [CPS98, DGG98, EP97, Eng99, Gre91]. Fowler [Iff87, LP81]. Fractal [BEHM89, DGHM96, Tri97, Ghi86]. fractals [DL92a]. Fraction [Ahl93, Lan83b]. Fractional [BH75b, DRF95, EM70, Erd74, Lub86, McB75, Osl70, Osl71, Osl72, Osl73, Pra72, Tre79, TF84, Won78]. Fractions [Fai71, Fie84, Mey86, PH75, BF79, Lan80]. Fragmentation [MLM97a, MLM97b]. Frames [CC98, CS93a]. Frankl [AS79]. Fredholm [Av91, Bry74, Egg93, DMV98, Pit80]. Free [Ack78, Ack80b, AC80, AL98, Boh90, Col94a, FP86, FR94, GY90, Gio93, Har94a, Har94b, HP98, Hu91b, Hu91a, HW92, PP95, SV76, Sch97, SS95, Tep74, Tep75, Ter99, Wei99, Xie90, Zaj94, Ack80a, BEGv87, CFV81, DH84, Iff87, Kna86, Kor88, LL79, SAW87, Ter83, TT83, Vog85]. Free-Boundary [GY90, Vog85]. French [Ali95, BH87, Che92a, CD95, Die86, Die88, Hab88, Her95, Miz83, Tah90, Tah91, Tch96, Zen90]. Frequencies [HL93, Lab99]. Frequency [Chi92, CD93b, EK84, GM92, JNW94, Pan96, RT93, Vil92, Wex79, Gri79a, Wex80]. Freud [LL90, LM95, MX91]. Freud-Type [MX91]. Friedrichs’ [Mar85b, KKZ86]. Front [Den91b, FS96, Nii97b]. Fronts [NV95, RTV96, San98]. Fryant [Vem95]. Fuchsian [BC90a, SY89]. Fully [DT96, ES97a, May70]. Function [BG76, BAR90, BHU96, BÇ94, BM73, Car75b, Chi74b, DQ90, Die95, Dut85, EM94, Elc75, Fet78, FG99, For90, FMV98, Fre90, Gau74a, Gau74b, GG77, Gil70, GR92, Jia96, Kad72, Kap98, Lon74b, Luk70, Luk75, Mat78, Mil74a, Olv91a, Olv92, Osl70, She98, Ste70a, Tri97, Vol90, WW92, And88, Bel82, Büh87b, Büh88, Fet77, GN83, HIN85, Iwa88, Log83a, Log83e, Lou84, LS79, NR85, Nay87, Olv93, Par84, PT78, Sal89, Sin80, Sri83, Sta85, Tem87, Wim85]. Functional [AVV90, AVV92, BB97, BC90b, BH76, CL96, DC85, HLL72, KI95, Lim78, Lou94, Mag87, Mao97, Mer75a, Mer76, Par92, Rei75, Rus77, Sat77, Tre85, BHS83, Chu80, GJ82, JSW87b, McC78, Ngu89, RW78, Sin79, Sta85, Sta86a, Ste85a, Tud87, van86]. functional-differential [BHS83, JSW87b, McC78, Sta86a, Ste85a, Tud87, van86]. Functionals [Bar85a, Bar88, Bra95, IL96, Jan96, Jer99, Kra78, Lip98, Sig74, Swa72, Vis90, de 78, BD78, CPS87, Rab84, Tah90, Tah91]. Functions [AHP75, AW98, AVV92, Ask74a, AP74, AT85, BR78a, BEHM89, Ber70, Ber99, BHL72, Ble77, BS76, BR82, BS75, BMY77, CMR81, Car70b, Car77a, Car70c, Car71b, Car74b, Car76b, CS84, CLH92, CK95a, CS97b, Coh76, Col94a, CH92, Dav73, Dol70, DGHM96, Dou98, Dun78, Dun90a, DL91b, Dur78, EL84, EL85, EL89, Far73, FG92, Fin74, FHM97, FG95, Fre92, GNP95, Gas75, Gim93, Gin79, GU77, GK77, Gro76, Gro77, Hei78, HV97a, Het70, Hra85, Hua96, Ism74, IM78, Ism81, JM72, Jaf97a, Jaf97b, Jai81, Jia98, Kan93, KS92, Kat97, KB73, Koh70b, Koo93, Lan83b, LMW96, LLS97, LA96, Ler90, Lew74, Lin74, Lin75a, Lin75b, Lir71a, LY70, Log79c, Log84b, LMS73, LS88c, Lor93, LS94, Man91, McB75, McC77, Mer75a, Mer75b, MP77, Mil70a]. Functions [Mil70b, Mil70f, Mil71b, Mil72, Mil73a, MR94, MK78, Mug75, Mug79, MS84, Mul75, Nås78, Nak91, Neu88, NEM94, Pal77, Pan71, Pan72, Pay70b, Pay74, PS72, Pet98, Pie74, Pil86, PS91b, PS98b, QV99, QSW93, Rah91, RM78, Red71, Ron97, Rua98, Saa74, Sac74, Sch76a, Sch76b, Sch76c, Ser75, Sha91b, Shu78, SS73b, SH71, Sri70, Tem90b, Tse97, Val94, VJ81, Wal88, Wal92b, WZ97, Wil91, Wol71, Zha86, Zho96, Zuc79, ki72, AM83, AZ79, Amo89, AT89, Ban78, Ban79, BG84, BC79, BD86, Bra86, Car80, CG83a, Car79b, CR83, Car86, Chi80, Cig79, Coh89, Dil88, Dun86, Dun89, Dur79, DW82, EL86a, EMN94, Fie73, Fie83, Fre87a, GN89, Gin78, GR79, GM84, GR85, Har78, Her95, How85, Ira89, IK79, Jan84, Jia88, Joh87, JT79, Kaz84]. functions [KS86, LM83, Lan80, LS88b, LM77, Log83b, Log83c, Log83d, Log85, LB94, Lon88, Lor80, LS93, McC83, Mel85, Mer77, Nay84, Nay89, Obi86, Obi87, PP79, Per86, Per88, PV84, Piq86, Piq89, Rah81a, RS85, Rah86b, Ral82, RW78, Rus88, Sas91, SS80, Sha80a, Sha77, Si78, Tem79, Tem82, Ter85, TWW89, Val86, Var88, Vol82, Vol84, Vre84, Wil86, Wim85, Yom83, Yom86a, Yom86b, Zuc84]. Fundamental [CRD88, Rao75, San83b, LL79, Vol84, Vas79a]. Further [Har70a, Osl72, SK76].

G [RA79]. Gain [And96]. Galerkin [AGJL92, AO81, Esh87, Sha80b]. galvanizing [Vog85]. Games [FMV98]. Gamma [Dut85, Fre92, Gau74a, Gau74b, Luk70, Luk75, Fre87a, Tem79, Var88]. Gap [Tha93]. Gaps [GH79]. Gas [BBL97, DS95, HS94, KN98, SR93, ZZ90, Bre91]. Gas/Liquid [HS94]. Gaseous [Lin97]. gases [Hil82, Kap79a, Kap79b]. Gauss [Bel82, BC79]. Gaussian [KS92, WW88a, Xu94]. Gegenbauer [Car71c, Col70, Dea79, Dur79, Rus78]. Gel′fand [Sas91]. Genealogy [RS91a]. General [BB97, BR78b, BC90b, DS82, Hai82, KB73, Lag72b, LA96, O’R93, Per74, Pra72, RDLH77, SDL84, Cam87a, Cha72, ET72, Gut82, HS86b, Ise85a, Ngu89, SS80, Sch80, Sin79]. Generalised [PG93, de 85]. Generalization [BSZ77, Fri81a, GM94, HB73, KM93, Lab99, Lin78, Par94, Pil86, Por76, Rah77]. Generalizations [CK77, For90]. Generalize [Gus94, AW79]. Generalized [AGGO94, AM80, BFGJ97, BO92, Bil74, BS76, Bre78, BR82, BS75, BG74, Car70b, Car73b, Chi73, Cho84, CB87, Col94b, Dol72, EMN94, Fri95, GS80, Hai80, HQW74, IM87, Jer99, Kap98, Koh70b, Köl86, KSK78, Kui85, LSZ92, Liu95b, LW79, McB75, Mei77, Mil72, NW74, NEM94, NV95, NSSS84, Olv91b, Osl71, Pan71, Pan72, Pil86, Que71, Red71, Ros78, SS94, Sri70, Wei74, Zha95a, dS97, Akı88, AZ79, Bry86, Büh88, Cam89a, CE87, Fry78, GT89, Lai80, Lan78, Lou80, LP81, MQ85, Mei79, Ort80, Pin88b, SW79a, Sha77, Som80, Vre84, Wal79]. Generated [BG75, CJ76, FMP98, Sun93, Jay88, Sta85, Sta86a]. Generating [AHP75, BR82, Car77a, Fry91, Hun73, Ism74, Kra78, Sac74, Cig79, NR85, PT78, SW79a, Sri83]. Generation [CV87, Swe98]. Generators [BC77]. Generic [Cos73]. Genericity [Gu92]. Genes [Lin95]. Genetics [BT91b, Lui82a, Lui82b, Lui85, Lui86]. gently [EF85]. Genuinely [Wan98b]. Genus [AJV94]. Geometric [AHP76, AM94, CJ93, Dun93, Eva96a, GS93, Gug71, LM96, NV95, Fen85, Ful82, Ful83, Sal89]. Geometry [Chi77, Nak91]. Geophysical [Col97]. Geostrophic [CW99]. Geronimus’ [Jay88]. Gierer [WW99]. Gilbert [JKV99]. Ginzburg [BCP93, BS97, CJM98, DLY98, GP99, HT98, Jer99, JM96, LD97, RS95, SZZ98, TBD^{+}96]. Given [Gla76]. Glicksberg [GS81]. Global [AV90, Avr88, BBN93, BB86, Bor99, Bra71b, Bra91, Bro96, BRC72, BIT91, CJ99, Car95b, CB89, Che92b, Daf82, DGG98, DRF95, Fen85, Fie86, Fit82, FD97, GM91, GGI92, GT94b, GH86, GH87a, HRX98, Hea88, HMP87, Hol93, Hru85, HS96, JMR96, Kli97, Klo85, Lev74b, LM96, Liu97, Lu90b, Mis95, Mor89, Nis82, NR95, Pos90, Pou83, RR95, RTV96, Sac85, Tha93, Ven78, WW93, Wol93, Wol96, Zha99b, Zha92b, AR87, Che92a, Chu84, Cos87, DL91a, Gin78, Keg83, KS86, Pao80, Tsu84, Veg88]. Global-Stability [LM96]. globale [Che92a]. Globally [Ito94]. Golomb [dBJ82]. Good [Ron97]. Gordon [Bil90, BFF^{+}90, CMOR86, KS85b]. Goursat [Sas91]. Gradient [DP97, Dou97, Eva96b, Ish96a, Li97, Mis95, CW89, Sch87a]. Gradient-Like [Li97, Mis95]. Gradients [FMP98]. Grained [MA88]. Gram [Wil91]. Graph [STW88]. Grating [BC97]. greatest [MNT86]. Green [SV94, Ban78, Har78, Ira89, Lon88]. Green’s [BG76, Chi73, Hei78]. Grey [FMV98]. Grids [Zem82, Zem81]. Ground [FT99, FM98, Wei85]. Grounded [Zem82, Zem81]. Groundwater [CvL92]. Group [Ast91, AT85, BÇ94, KM78, KMM94, Koe94, Koo91, Koo93, Lan96, Sat77, Koo82, LC84]. Groups [BC90a, CR91, Gus92, Hra85, Yok97, Bos79, Sil81]. Growing [Tol98]. Growth [AHIR93, Bra75, CR92, Koh70a, KS85b, LN98, Mak76, Pat79b, Pat79a, QVV98, Str76, Swi76, Wal88, Wal92b, Wil85, BF88, BF89b, Car79b, Car86, GH87b, Kur79, McC83, Wad80, WG87]. Grünbaum [Mer75b]. Guided [BB90]. Gulf [BCT88].

Haar [Zho96]. Hadamard [Bai71, Cha78]. Hahn [Cie87, Det95, Dun78, ILV89, Rah76c, Rah78, Sta80b, Wil70]. Half [CP97, EP99, HKS91, KL98, KL99, Lin75a, Mey76, McC83]. Half-bound [HKS91]. Half-Plane [Mey76, McC83]. Half-Space [CP97, EP99]. halfspace [McC81]. Halfway [BS92]. Hamilton [BO91, BF98, BKW85, CS97a, GV93, KM80, KM81, KM83, Osh93, Zag98]. Hamiltonian [Bar81, BKM86, Dv96, GH87a, GH88, HS84a, Kra89a, Kra89b, Mey84, Rab82, Rei79, Rot93a, Ste85b, vV84]. Hamiltonians [Bea85]. Hammerstein [Dol80]. Hammerstein-type [Dol80].

Hamming [Llo80]. Hankel [AM80, BS96, Bow71, FTZ88, Gas73, GW97, JK75, Koh70a, Koh70b, Lee74, Lee75, Lut78, Muk72, RS74, Sch73b, Ste73b, Won76, de 85, vd86]. Hard [BP90]. Hardening [NR95]. Hardy [GM84, Hei75, Log79a, Wal97]. Harmonic [AN98a, And93, Ber99, BSSV98, Col94a, Dom72, Far73, Gau74a, GK77, HL96, Man91, Mey76, Mil70a, Mil70b, Mug75, NS91, Pay70b, Pay74, Pie74, Rei93, WS80, Wol71, Ban79, Iwa88, PP79]. Harmonics [Fry91, Sac74]. Harmonious [GA98]. Hartree [HNO98]. harvesting [Bra83]. Hausdorff [Wei99]. Having [Dun90a, Dun90b, Dun94, Gri91, Olv77a, BD86, Ell82, Kel89, KS86, Log83a, Log83e]. Hearing [Car95a]. Heat [Ack77, AGGO94, Aye83, BBDU93, Bil74, Bra91, Car76a, Dav78b, Den95, Eva96a, Gal91, GV93, Gal95, Gla76, Hil97, Hsu89, LP94, Lun90, Mey74, PW70, Qui96, RS95, Ves97, Bel87, BF89a, BF88, BF89b, DM82, FL87, HV87, HI89, Nau79, Rei86, Tro87]. Heating [Xu92]. Heavy [HH88]. Heisenberg [CP84]. Hele [ES97b, Gus85]. helicity [Tur89]. Helmholtz’ [Kri80, AK85, KM80, Mil76, Mil79, Wei74, WW88b]. hereditary [Mar81]. Hermite [AGV82, HPR97, JRS94, Lah93, Mar83a, Muc70, Sle72]. Hermitian [Rei70, SL78]. Heteroclinic [Ber88, CDT90, Den91a, Moe84, San98]. Heterogeneous [EM90, Sha91a]. Heterozygote [Lui85]. Heun [KM91a, KM91b, Kui79, LMPS90, Val86]. Hidden [BEHM89]. Hierarchies [Per81]. High [BMY77, Log84b, She83]. High-Pass [Log84b]. Higher [CD74, DP97, Dou97, EZ97, KO92, KS85a, LB93, MR95, NS98, PK91, Ruf95, Yu74, Gun73, How82, Mur83]. Higher-Order [KO92, LB93, PK91]. Hilbert [Log83a, Bar75, Bar77a, Bar79a, BB78, But79, CH79, CH82, Dat81, Dol79a, Dol79d, Eng82, FLLL89, Gla84, GL74, Gri78, HL77, Han74a, Han74b, Han76b, KS80a, Lag73, Lev73, LM75, Log83e, Lon77, Mar80, Mei79, Ort73, Paz72, SS84, Wex80, Won80, Zho89, Zio85]. Hilbertian [Mas78]. Hill [FIT87, GH79, KM95, Ort94]. Hille [St.84]. Hilliard [EG96, Ish96a, WW98]. Hohenberg [Kuw96]. Hölder [AT89]. holds [RV87]. Holes [Kai91]. Holmes [HT89]. Holomorphic [CR83, Vol90, vdP83]. Homeomorphisms [KS79c, MC79, Hal84]. Homoclinic [Che96, CDT90, Den90, Eck92, Gru92, HZ94, Kov95, Rob92, WH87a, WH88, Bra81c, GH84, Gru85, Stu89]. Homogeneous [ANS96, Hai80, Hir72, MM97, Sch76c, Wys93, Yam90, ML93, Pou89, She93, Ura84]. Homogenization [All92, AC98, AHZ93, ADH90, AD86, Bal99, BLM96, DV87, DD91, DQ90, Fra83, MA88, RB98, ka76a, ka76b, ka77, Ngu89]. homogenize [Dam81]. homogenized [MB91a, MB91b]. Hopf [BF98, BJL98, FLLL89, Fie86, Gar92, HJ92, IR89, Lab85, Lab89, PFK95, RJ86, SM84, Ste85a, Ste73b]. Hopf-Type [BF98]. horn [Sym86b]. Hudson [Jan84]. Hull [BHU96]. Hurwitz [Fia80, Kem82b, LR72, YS93]. Hybrid [MZ98]. Hydraulic [DuC97]. Hydrodynamic [Zha94]. hydrostatics [Coh89]. Hyperbolic [AB72, AH72, Car95b, Cha76, CK91, Col73, Coo74, DL96, EG90, EW89, FP86, GHL96, GPG91, Hal91, HY95, KT97, Ład94, Lin74, Lin75b, Mag87, McC97, NY97, Pey71, Qui96, Ren97, Smi71, Sta80a, TW91, Zuc79, vV85, BF89a, BT87b, CL81, Dur79, Esh87, Hru85, HW83, Ira89, JW88, KKY84, LS88a, LS89, Lev83, ML84, Pey75, Ren84, Smi89, Web79, Zuc84]. Hyperbolic-Elliptic [GHL96]. Hyperbolic-parabolic [EW89, BT87b, Esh87]. hyperdifferential [Si78]. Hyperfunctions [ZW82]. Hypergeometric [AVV92, AJV94, AW82, Car70b, CS84, DG92, Gas81, GS82, GR85, Gus90, Hen87b, HBL76, Hra85, Jai81, Kad72, KM91a, KM91b, Kan93, Kui79, Mil70e, Mil72, Mil73b, Olv91a, Pra72, QV99, Sri70, Tem90b, VJ81, Wil80c, Ask80, Bel82, Büh87a, Büh87b, Büh88, CG83a, Dun89, ES84, GR86, Gus87a, Gus87b, Hol80, KS86, LB94, Olv93, Sas91, Sin80, Ver80, VJ85]. Hypergroups [DR74]. Hypersonic [Gio93]. Hypersurfaces [EMS98]. Hypoelastic [Ren90]. Hypoelliptic [BC77]. Hysteresis [HS95, LD97, Vis86].

Ice [HV97b]. Ideal [LF98, Wil79]. identifiability [CG87]. Identification [BV92, FKS99, FG87, HS86b]. Identities [BÇ94, Car78b, DG92, Elc75, For90, Sch76a, Sch76b, Sta86b, WB95]. Identity [BG75, Car75a, Mil94, Mim88]. if [Ols74]. II [Ack80b, Ask71, Ask74a, Azz81, BG89, BJL88, Bar88, Bel81b, BLN87, Bor85, Bre93, BS80, CC91, Car78a, CA71, CS77b, CD96c, DL92a, Dau93, Fen83b, Ful83, GT94a, GM76, Gra88, Gra72, HS94, HP85, Hir85, How78b, HS98b, Ise85b, Jaf97b, Jai81, Koo74, KZ88, Lab89, LPSS89, Lin75b, Log83c, Lui82b, Mel85, Mil70b, Mil79, Mil74c, MB91b, NT86b, Nip86b, Par70b, Peh84a, PP95, Ral82, Rea84, Rea88, Rob92, Sch76b, Sen97b, SHM89, Szw92b, Vas79a, Yom86b, Yos86, ka76b]. II. [WO91]. III [Fri81c, Ise85c, Koo75, Log83d, Lui85, NT86c, ka77]. III. [CD93b]. Ill [Car99a, Fra78, CO89, MN86, Mil70c]. ill-conditionedness [MN86]. Ill-Posed [Car99a, Fra78, Mil70c]. ill-posedness [CO89]. Image [Car97]. Imaginary [Dun90a, Lab99]. Imaging [Bro96]. Imbedding [SW73]. Immersed [EMS98]. Immiscible [BLM96, van79]. Impedance [Bro96, Col82]. Impedance-Imaging [Bro96]. implementing [Piq86]. Implicit [DS81, FD97, Mat78, Cam89a]. imposed [SAW87]. Improperly [Ade73, Ade74, Col73, Ame82]. Improvable [Sch73a]. Improved [Olv91b, Olv91a, Olv93]. Impulse [Lab85, Lab89]. impulsive [Coc87]. Impurity [DLY98]. Inclined [SS75]. Including [Luk75]. Inclusion [CJ76, DLY98]. Incomplete [Luk75, Tem87, Son83, Tem79]. Incompletely [Hal91]. Incompressible [BPL91, Car96, HFL99, KT99, Ren90, Sim90, Che92a, Ghi86, HS89, Kim87, Ton82]. Increasingly [DD91]. Indefinite [CC91, Chi77, MK83, PV84, Piq86, Piq89]. Independence [Wan98a, Zha92b, Lam80]. Independent [Col72, LY96, Mil76, Que71, Mil79, Mil84]. Index [Lip72, Lip73, Nii97a]. Induced [BRC72, Her90]. Induction [Van96]. Inductively [Fry91]. industry [FMP86]. Inequalities [Ack77, AMS86, AW98, AVV90, AVV92, BM73, BI75, Bow71, BS79b, BS80, Car74c, CS77a, CLH92, CS93a, CS97b, CS78, Dör87, Dut85, Fin74, Fin77, Fri81b, Fuj92, Gau74b, GM89, How78a, LM83, LL90, Lip98, Log79a, Lor93, McC77, Mul78b, Pat71, Pay70b, Pay74, PS76, PS91b, QVV98, Rei75, Sig76, Tho77, AH83, BI83, ELS94, Fei79a, Gas89, Gat87, Gun73, Gus85, Har78, HS86b, How82, Lon88, LS93, Mar83a, Mer77, Mon80, Nau79, PP83]. Inequality [Bar77b, Dör91, Gau74a, Hei75, Hu90, Mer75b, Moo73, Swa78, Wal97, AT89, CP84, Elc72, Mar85b, MRS81, Mir83, Par84, RV87, Sto80]. Inertial [Gou92, JT94, MNY94, Tem90a]. infection [WHG88]. infectious [WL82a]. Infectives [Wan79]. inferior [Lui85]. Infinite [BS96, BMS83, CS97a, DJ87, GN89, HW89, HI89, KF90, Lir71b, MW83, McL78, MP97, QV99, Rei73, Van85, Van01, Zem79b, Zem82, CGR88, DL92a, GM83, Gri85, IK79, Kri80, KS80b, Mar82, Mar83b, Old83, RS85, RV87, Ren84, Zem79a, Zem81, Zuc84]. Infinite-Dimensional [HW89, KS80b, RV87]. Infinite-order [HI89]. Infinitely [Den91b, Mai94, JK86].

Infinitesimal [BC77]. Infinity [BHL72, Kat97, Lew74]. Inflow [KK97, Ren96]. Influence [DGR99, LP94]. information [MRS81]. Inhomogeneous [AR98, CWZ99, CP90, FK94, GM76, RS91b, SS77, Sin97, CKP89, FK88, Spe79]. Initial [Ark90, BS74, Car99a, Col73, Dix96, Eva96a, FI96, Gor95, Hof96, LY96, MP97, Moc74, NT96, O’M71, Pao74, PJ73, Ren90, Ren96, SS93a, Zha95a, Ale80, Aro81, CK87, CS85, Daf82, Hru85, Kim87, Lui82a, RW78, SAW87, Sta85, Ton81, Ton82, Tsu89, Lui82b]. Initial-Boundary [FI96, NT96, Pao74, Ren90, Ale80, Daf82, Kim87, Ton81, Ton82]. Initial-Value [Dix96, Ren96]. initially [ACK83]. Inner [Moo73, Bra86, MK83]. input [Coc87]. Instabilities [FR94]. Instability [CD93a, FSV99, Liu95b, LF98, MM77b, TN94, Veg93, de 95, CL89, Hoc84, Lev88, TT83, Wil80a]. Insufficient [Chi74b]. Integer [BAR90]. Integers [HT71]. Integrability [CMOR86, DP97, Dou97, Gri81b, vV84]. Integrable [Kov95, Nak91, BT96, BKM86, GM84]. Integral [AHP76, AGJL92, Av91, Alp93, Ask74b, Ask82, Atk74, Bar77a, BW73, BH75b, Ble77, BS76, Bra76, Bra75, Bro78, Bry74, Bus72, But80, CG85, CRD88, CG71, D’A75, DG92, Egg93, GO71, GK77, Gri79b, HQW74, HB73, Hir72, Joh75, JW74, Jua95, Kam92, Kap72, KS79a, Kla71, Kra74, Lae75a, Lae75b, LJC97, Liu91, LY70, Log84c, Lon77, MW87, McL90, Mer77, DMV98, Mie87, MF71, Nay77, NS91, Nus78, Okr91, Olv91b, Olv91a, Par94, Pen92, Per74, Pra72, RC73, Sch76a, Sch76b, Sch73b, SW73, SW75, Šim84, Šim85, Slo75, Son70, SS73a, SV94, Sri72, Ste75, TKS93, Tre76a, Urb96, Vol82, Vol84, Wal75b, Won70, Yu74, Zei94, dGdJ75, dHW73, dM89, AK88, Akı88, AH83, Art75, BD78, BE83, CS89, Cos88, DS82, Ell82]. integral [Fie86, Fri81c, Gar88, GJZ88, GS81, Gri78, Gri81a, Gri85, GM79, Hab88, Hir75, Hiv78, JW80a, Kad88a, Kad88b, Kam89, KS80a, Log83a, Log83b, Log83e, Log83f, Log84a, Lon83, Lui82a, Lui82b, Lui85, Lui86, MB91a, Nay89, Pit80, Rah86a, SW79a, SDL84, Sha80a, Son80, Sta78b, Ste73a, Tef78, Tre84, Tre88, Val86, Vic78, Wal79, WL82a]. intégrale [Hab88]. Integrals [AVV90, AVV92, Aom94, Bai71, Boe78, BG91, Car77d, Car78c, CG94, Car73b, DP97, Dur78, EM70, Erd74, Esc73, Eva92, Fet70, Fet78, Fuj92, Gas75, Gil75, GC95, Gus92, Gus94, JM72, Jan96, Kan93, Kap98, KJ82, Lan74, Lip72, Lip73, Log84b, MR94, Muk72, OT94, QV96, QVV98, RS94a, RS94b, Rus77, Sen96, Siv71, Ski80b, Ste70b, Tem90b, dK71, Amo89, Aom87, AB80, Ask80, CS89, Fet77, GN89, Lan80, Mim88, PV84, Piq89, Sim80, Ski80a, Ski81, Ski88, Son83, Ste72, Tem82, Tem87, Won78]. Integrands [FM92, KP92]. Integration [CMR81, McB75, Pra72, Ral82, Šim84, Tre76a, Yue73, ki72, Har83, JS87, Old83, Piq86, Tre84, Tre88]. Integro [Bra91, Cus75, Fuj92, Hol93, MN75, SK73, Blo80, Can85, CH79, Dav79, DG89, DR86, Eng82, GK84, HW82, HW84, Hea82, JW80a, Lei81, MR83a, MM79, Pao80, Sta88, Ste80a, TW79]. Integro-Differential [Cus75, Hol93, MN75, SK73, Blo80, Can85, CH79, Dav79, DG89, DR86, Eng82, GK84, HW82, HW84, Hea82, JW80a, Lei81, MR83a, MM79, Pao80, Sta88, Ste80a, TW79]. Integro-Partial [Fuj92]. Integrodifference [Lin95]. Integrodifferential [Ada78, BHN72, Bro75a, Han76b, Han77, Lia94, Lon71, MM77b, Noh74]. Inter [IMP92]. Inter-Intracohort [IMP92]. Interaction [SM84, Smi91b, NT86c]. Interactions [Bar84b, Fan93]. interconnected [GM79]. Interface [BC95, Car94, GGI92, HS94, ka76a, ka76b, ka77, ACK83, Kri80]. Interfaces [CWZ99, CK95a, EY98, TN94, HV87].

Interior [CK91, Gra98, RS91b, Do 93, DeS87, Kel84, KZ88]. Intermediate [FS77, FS95, Bea85]. Intermittently [PS94]. internal [Bri88]. Interpolant [Boj94, dB76]. Interpolants [Pin88a]. Interpolated [Nov88]. Interpolating [BFW86, GJ82]. Interpolation [BEHM89, Bor85, CSW91, DGHM96, Fry78, Gri96, Jen91, JRS94, Lah93, LM95, Mat70, Ray81, SS89, SV83, Wal97, DG88, Geo87, HIN85, Mel85]. Interpretation [Eva96a, GJZ88]. Interpretations [BÇ94, Koo82]. Interreacting [Ost96]. Interreflections [Kra74]. Intertwining [DGH96]. Interval [CMR81, KZ95, Ral83, Rei73, RS91a, Smi75, Van85, Van01, BN78, Hen87a, Log83d, Ral82, Tro81, van84]. Intervals [Kol74b, MW83, GH88, Hoc84, Mar82, Mar83b, Peh90]. Intracohort [IMP92]. Introduction [Rob85]. Invariance [Ste77, Tho77, Was73]. Invariant [Bar83, Bar84a, BSSV98, Bra81c, GZ92, Gor95, Jia96, JKK96, KS92, Kui80, LJC97, Mag87, Mor85, Smi86a, Veg88, Was74, WO91, Ben82a, Lam80, MrS87, Rob89, She86, vd86]. Inverse [Av91, Bar91, BK95, BM73, BK76, BR81, CK92, CH93, CK95b, DT81, DuC97, Duc95, EP97, FKS99, GHLS91, Häh98, Hof99, KSS97, KI95, Kli97, Lin91, McL76, McL78, NU95, RS91b, SU97, SF91, Ung74b, Ves97, Bry86, But79, FIT87, HS79, Kin86, Lan83a, Lan78, Lou80, MO83, PR87, Sym81, Sym86a, Zho89]. Inverses [Jus74, NW74]. Inversion [Lon74b, NC82, PS91a, Per74, PW70, Sym91, Tch96, Wil79, Yao95, Aro81, Mar84]. Invertibility [CSW91, Duc95, FG95, Kaw79, MST96, MQ85, Mul81]. Investigation [WY75, FMP86]. Investigations [Ład94]. Inviscid [HFL99, Ito94]. Involve [Bra74]. Involves [Dea71]. Involving [BP96, BM73, Boe95, Bra71b, CP73, Erd74, Jan96, Old70, Old72, Šim84, Sri70, Tre76a, Won70, Ack81, Ban79, Dol79b, GN89, LMN87, Nay89, PP79, PV84, Tre79, Val86, Wim85, Won78]. ionized [Hil82]. Irreducibility [Par92]. irreducible [Koo85]. Irregular [CM80, FG92, Nau76, BJL81, BJL88, KS86, Sch84a]. Isochronous [CD97]. Isolated [Man91, Sha76a]. isometric [CL87]. Isometry [Kre94]. Isoperimetric [Bar77b, CS78, GM89, Pay70b, Pay74, PS76, PP83, Lon88]. isosceles [Moe84]. Isospectral [CPS98]. Isothermal [DS95, Veg88]. Isotropic [CS77c, NR95]. Iterated [FMV98, Gre74, RS91a, Mur83, Per81, Per82]. Iterates [CG88, Smy79]. Iteration [Kel79a, Kol74a, Nov88]. Iterations [FK94, FK88]. Iterative [FG92, Lun78]. IV [Hir90, Hor75, Lui86].

J [Ack80a, Ban79, Car86, Cig79, GLY88, Log84a, Mad88, Obi81, Spe79, Tah91, WH88]. Jabotinsky [LT81]. Jack [Kan93]. Jackson [Aom94]. Jacobi [EMN94, Log84a, Miz83, Aom87, Ask74b, BO91, BF98, BKW85, Bus77, CS97a, Car74b, Car74c, CI91, Cie87, Dur78, Dur79, Edw93, EL87, GV93, GR86, Gat87, Hor75, Ism86, KM80, KM81, KM83, Koo74, Koo75, Lai82, Let94, Lew86, Lew79, Log83f, LT94, Mar83a, MR91, Miz83, NEM94, Osh93, Pat71, Rah76a, Rah76b, Rah77, RS85, Sto97, Vre84, WW92, WZ94, YH82, Zag98, vdW75]. Jacobian [Lan80, MC79, Ols74, Sal89, Tse97]. Jets [Rei91]. Jordan [Mha88]. Josephson [SC86, Sch87b, St.96]. Jouguet [HNT98]. Joule [Xu92]. Jump [Lar98]. jumping [HS84b]. Jumps [Bra95, Die86, Die88]. junction [Sch87b]. Junctions [Sch90, St.96]. Justification [DW93, Mac90, Sac84].

Kadomtsev [dS97]. Kakeya [ASV81, Rus78]. Karman [Mul78c, MM92, HS84b]. Kassoy [BT87a]. Kato [HM89a]. Katzenelson [PG93]. KdV [Sac83, Sac84]. Kernel [AGGO94, GW97, Han76b, Hil97, Hor75, Hsu89, IM98, Kif77, Mei77, NW74, Rah77, Rah78, dHW73, Ell82, GR83, GR84, GR86, HR88, Mei79, MB91b, NR85, Sta78a]. Kernels [HB73, LR84, Log88, MF71, Rah76a, Rah76b, Rah76c, Rea83, Rea86b, Sri72, AB80, CN81, CB87, Gri80a, Ha86, HW84, Hir86, Log79a, Log79b, Rea84, Rea85, Rea88, Son80, Vic78]. Killing [KM80, KM81, KM83]. Kind [BW73, Bry74, Car77d, Din89, Egg93, DMV98, Pra72, Spe89, SHM89, Dur79, MR83b, Nay87, RS85, Rah86b, Wei84, dW80]. Kinetic [AN98b, Ren91, Xie90, vP86, Kap79a, Kap79b, Lei85]. kinetics [BSL87]. Kink [HR99]. Kirchhoff [MS88]. Klein [Bil90, CMOR86, KS85b]. Knot [Hu91b]. Knots [MP77, CS80]. Kolmogorov [Cus82, PT97]. Konhauser [PT78, Sri83]. Kontorovich [FI72]. Koornwinder [Ask74b]. Korteweg [BO92, BW83, CK87, GT89, HL94, Hay91, Tsu89, Zha92a, Zha95a]. KPP [MR95, Moe79]. Krahn [Ban73]. Krawtchouk [Zen90, Det95, DR74, Koo82, Rah79, Zen90]. Kreın [RA79]. Kreiss [Fri81a]. Kronecker [Sle72]. Kuramoto [Mic92, Mic96, Tad86]. Kwong [CJ93].

L [Gre74]. L. [Cig79]. Lag [KF90]. Lagrange [LM95, Wal97]. Laguerre [BS79b, BS80, Dur77, Dur79, EKLOO92, FW88, GT94a, Gil75, KM93, Koo77, Let94, Mar83a, Mar86, Muc70, Pan72, SS94, Zay82]. Lakes [BLS99]. Lamé [HS77, Sha80a, Vol82, Vol84]. Laminar [Gio90b, HLM92, Lu97]. Lamm [Yos84, Yos86]. Lánczos [CB87]. Landau [JKV99, BCP93, BS97, CJM98, DLY98, GP99, HT98, Jer99, JM96, LD97, RS95, SZZ98, TBD^{+}96, Zha99a]. Langmuir [BD95, Ben96]. Laplace [Miz83, Ski81, Ask74b, BBF99, Car80, DS87, HL70, HL74, Kal75, Lef82, LW94, Lon74b, Miz83, Pri75, SS92, Ski80a, Ski80b, Ski88, Son83, Tem87, Ung74b, Vol90, Wil86, Won70, Yao95]. Laplacian [Tol98, CL91, CK95a, EP99, Elc75, Sch87a, Ura84]. Laplacians [AB93a, AB93b, Var88]. Lardner [Car70c]. Large [Ark90, AH72, BH93, BB94, Car99b, Car96, Dun90a, EG90, HL74, HS77, HH88, Kee77, LA91, Lu97, MM77b, Pap95, RB98, St.88, Tak92, Wal75a, Dun86, Fie73, Fie83, Hru85]. Large-Scale [BB94, MM77b]. Large-Time [Car96, Tak92]. Lateral [DGR99]. Lattice [JRS94, RS94a, Chu84]. Laurent [JT79, Ros78]. Law [BD95, GR92, Nak88, Sin97, Gar87, MMN85, Mar81, Wag83]. Laws [Ast91, CC99, Daf95, DL96, Die95, Fan93, Gim93, HS98a, IMP90, Jam98, Key91, KKL99, KT97, NY97, Sch84b, Ter99, TW91, TW97, Wan98b, You99, Daf87, How87, JW88, Luc88]. Lax [BJL98]. Layer [AF90, AF96, Hab76, Har72b, SS93b, Do 93, DeS87, Kel84, O’D84]. Layered [Sym91]. Layers [AC98, AN98a, Boh90, CWZ99, How87]. leaching [Kna86]. Least [Mil70c, QSW93, Mar85b]. Lebedev [FI72]. Lebesgue [BHL72]. Leffler [Cam89a]. Left [Olm77]. Legendre [Koo91, BD86, Fre90, GJZ88, Koe94, Las83, McC97, MK83, Sch76a, Sch76b, VK91, Wil72b]. Leibniz [Osl72, Osl73]. Lemma [BHL72, BMS83, CLP89, CGR88, GM83, KM88, Sid85, ZD84]. lengths [Hen87a]. Less [Bro96]. Level [FMV98, Joy97, Osh93, Lon88, WHG88]. leveling [Day82]. Levinson [HKS91, MMM97, Pan96]. Levinson-Type [MMM97]. Li [HP83]. Liapunov [Rua98]. Lie [ABM87, Boy76, Edw93, Hra85, KM78, KM85, KV98, MR91, Mil70d, Mil70f, Mil70e, Mil71b, Mil72, Mil73a, Mil74a, Mil74b, Mil74c]. Liénard [Ben81, Ben84b]. Life [MY98]. Lifschitz [JKV99]. Lifshitz [Zha99a]. Lifting [Bra81b, FF92, Swe98]. Like [BS95, Jua95, Li97, Mis95, Xu92, EY98, Fei79a, Jos99]. Likelihood [PR71]. Limit [AS87, BS97, Chr80, Col97, DT81, ES97a, GL90, HNT98, Jan96, Joy97, Lu90a, Mik88, NMIF90, NS98, Nor88a, SC86, STW99, Smy79, St.88, Thr94, WK92, BF79, Che81, CR88, LC88, Spi79, Hir82]. Limit-Cycle [St.88]. limite [CD95]. Limited [FG92, Qui93, Mel85]. Limiting [CD95, Lu99, vd87, Sil85]. Limits [Fan97, FHX96, Gio90b, LT99, Log79b, Mad85b, BNS85, DV87]. Lin [FI90]. Liñan [AV90, HP85]. Line [KL98, Kut84, Sat90]. linéaire [BH87]. linéaires [Ali95]. Linear [ABV96, AP92, BG88, Bak72, Bar83, BW73, BAR90, Ber76, Bra71b, Bra74, Bra76, Bro78, BHS83, Cam77, CR79, Cha76, Chi73, CGP93, Coo74, CC96, Dav73, Dav78b, Dol70, Dol72, DW93, Dun90a, Dun90b, Dun94, Eng99, EW94, FMT96, Fis75, Fra83, FH90b, FD97, Gug71, Han74a, Heg81, Hei78, Imm91b, JPS87, JRRV97, KO96, Kee72, Kim92, Kol74a, KF90, LS96, Lew86, Lun90, Mei76b, Mei77, Mia93, MR98b, Mil75, Mor85, Mul78a, Mul78b, NW74, Nau76, OFK77, OP87, PP93, PP95, PG93, RS94a, Rat74, Riv93, SW73, Ser75, Sha76c, Šim84, Šim85, St.84, Str77a, Suz95, Swa72, TF75, Tre76a, Vas79b, Wan98a, Wil72a, Wyr77, Zha92b, AH88, Bar81, BS79a, Bos79, BD86, Bre82, Bry86, BE83, BR84, Cam87a, Can85, CDMS81, CN79]. linear [DG88, DG89, Dol79c, Fei79b, GS89, HW81, HKP87, Jia88, Kap79b, Kim86, KS86, Lan78, Lar81, Lev83, Lou80, MM79, Mul81, MR79b, O’M79, Pat79b, Pat79a, Pet84, SS80, Sch80, SM85, Ste85b, Str80, Tal80, Tre84, Tro81, Vas79a, Vas81, Wil80a, Wil71, Zei80, vV85]. Linéarisation [Zen90]. Linearization [Bre81, Dol79a, FI96, Las83, Sym86a, Zen90, Dol79b, Gas85, Kus88]. Linearized [Mik88, Par92, Rei91, Kel88, Sac83, Dau89]. linearly [AT89]. Lines [SK76]. Liouville [Bra76, BS75, Cam87b, Car99b, EM94, Fuj92, KKZ84, KKZ86, KZ88, Rap82, RS86, SV94, Wal71b, WZ87, ZW85, dT78]. Lipschitz [AGJL92, ABV96, Cos88, Hen87a, Kaw79, Muk72, SS73b, Wu90]. Lipschitzian [Pou83]. Liquid [HS94, SZL97, Smi91b]. liquids [Eng87, MR83a, Ren82]. Liron [How74]. Little [Koo91, Sto97, VK91, Daf87]. Littlewood [CS93a, CS97b, MT95]. Liver [Hol93]. Local [ANS96, And91, Ben84a, CD74, FG95, Hir83, HT71, Kim95, Lun87, Ren90, Rua98, Zha94, Day89, Ful82, Ful83, MQ85, Ton82, Whi80, DL92a]. Localization [RT93]. Localized [AGJL92, HM89b, RB98, Sun93]. locally [Lon83]. Location [Mil70a, Mil70b]. Logarithmic [Ble77, Car99a, Hab76, AB80, ELS94, Sto80]. Logarithms [Won70, Won78]. Logistic [CC91]. Long [BO92, EZ97, LY96, LB96, Lu99, Wei82, CB89, HS86a]. Long-Time [EZ97, LY96, Wei82, CB89]. Long-Wave [BO92]. Longitudinal [SS93a]. Longtime [GM91]. Loop [Den91a, San98, Sch87c, Mim88, Sel81]. Lorenz [Che96, HZ94, Rob92]. Lotka [Ko98, WHG88]. Low [AB93b]. Lower [BSX95, Bra95, BR78b, FM92, FM99, Fox72, Jer99, LL79, LM75]. lubrication [CL86]. Lucas [BR82]. Lurie [Chu80]. Lyapunov [Ber70, DC85, Erh73, FG99, FHM97, Paz72, RRW88, Sta88].

M. [RA79]. MacDonald

[Gar90, Sta86b, Zei87, Zei88]. Maclaurin [Bre85]. MacMahon [Car77c]. Magnetic [CL90, Wol96, Fri81c]. Magnetohydrodynamic [FS95]. Magnitude [Rea86a]. Mahony [MM77a]. maintaining [Sin79]. Majorization [Cla83]. Manifold [BB92, BW80, MN75, Tur96, Rob89].

Manifolds [DS99, Gou92, JT94, JKK96, Mag87, MNY94, Tem90a, Bar84a, GJ82, MrS87, Ura84]. Many [BV92, Den91b, Gus92, Kai91, Dam87, JK86, RS86]. Map [IM98, RS91a]. Mapping [Yam71, CL87, Hom79]. Mappings [Duc95, GR78, Smi86a, Bra81c, DK83, Kaw79, Oht79]. Maps [CG88, FMV98, Her90, MST96, PFK95, Pou83]. Marchioro [LN98]. Markov [Dör87, Dör91, KR81, LL90, LS93, Mir83, She86]. Markov-Type [Dör91]. Martensitic [SZZ98]. Mass [BBDU93, FM98, PS97]. Masses [Car95a]. Master [Car77c]. Matched [Gor75a, Ski94]. Matching [BSSV98, Mac90]. material [Blo80]. Materials [Dav78b, HL94, KH95, Paş89, Ren90, CDHL88]. Math [Ack80a, Ban79, Car86, Cig79, GLY88, Log84a, Mad88, Obi81, Spe79, Tah91, WH88]. Mathematical [AK99, BB90, CE99, DS95, DW93, FL92, FT99, Fri81c, HR95, Mil71b, CT88, Hol87]. Mathieu [Hoc84, Nay84, Nay87, Nay89]. Matrices [AM80, Bax71, BB77, BS96, Chi73, CD91, Edw93, For90, Jus74, SL78, Auc89a, Chu84, DL92a, MR91, Sil81, Ter85]. Matrix [Ahl93, BB83, Car99b, CS97b, DT81, Fie84, Fri81a, Gin79, GW97, Hog98, Jia98, Kem82b, Koo85, LT81, PR91, Rei70, Sap71, Sle72, AH88, Gin78, GR85, HS79, MC79, MS79, Oja78, Rah81b, Sil85]. Matrix-Dilated [CS97b]. matrix-valued [Gin78]. Maxima [Ber99]. Maximal [Ahl93, BL84, Lop89, Par70a]. Maximization [BG76]. Maximum [Egg93, GG96, PR71, SS77, Spe79, Wan79, Ban79, Lar81, Log83d, PP79, Swe89, Whi80, dM86]. Maxwell [CJ88, CS85, HL96, JKV99, NS91, NP95, Rei93, Ren96, Yin98]. Maxwellian [MP97]. McKendrick [WHG88]. McLeod [CJ93]. Mean [BS97, CZ91, Fin77, GG96, Gau74a, Gau74b, KR81, LM95, McC83, MX91, Mul87, Ral83, SV83, dPFW99, CG83a, Sal89]. Means [Lip72, Luk70, vdP83, Bre85, Yom86a]. Measurable [BU97, Saf99]. Measure [DQ90, Dem96, KP92, Lop89, Pes95, Rah91, WW88a]. Measurement [FKS99, KSS97]. Measurements [ABV96, BV92, KS96]. Measures [AP93, FM99, DN86, Dom88, Fuh80, Lon83, Pag86, Sil81, Tsu89]. Mechanical [AL98]. Mechanics [ElB90, Ngu90, Hea88]. Media [AM95, BS85a, BCT98, CJ88, DuC97, FFW80, Gil79, HS96, RS91b, Kri80, MB91a, MB91b, van79, vd87]. Medium [BNS98, BC95, CQ97, CP90, FL92, FT88, KV91, KLV93, Kim95, NU95, SW91, ACK83, Col78, CKP89, Sym81]. Mehler [Sle72]. Mehta [Zei94]. Meijer [Fie73, Fie83]. Meijers [Mil74a]. Meinhardt [WW99]. Meixner [Zen90, Zen90]. Mel′nikov [Gru85, Kap98, Sch87b, Yag92]. Mellin [Fre87b, Gus92, HP73, KN72, Per73, Sid85]. Melting [HV97b]. Membrane [Bre93, PP93, Chi83, LL79, Lef82, PP83]. Membrane-Dominated [PP93]. membranes [She93]. Memory [BBN93, Dav78b, KH95, Lun90, SZZ98, JSW87a, JSW87b, MMN85, Nau79]. meromorphic [JT79]. Metal [SZL97]. Metals [ABH94]. Method [CK92, FS77, GG96, GK77, GB72, Gre91, HB73, HS98a, Kee77, Kim90, Li99, McL90, Pan96, PG93, RS76, Ser75, Vas79b, Yag92, Yar75, dK71, Dav78a, Esh87, FS79, Gin78, Gin85, Gru85, JAK88, MrS87, MS86, OP87, Per82, Rap82, Sch87b, SW79a, Vas79a, Vas81, Vit79, Zem81]. Methods [Ast91, Ehl73, Fis75, Gil70, GRZ83, Gor75b, KN72, Kol74b, Ler90, NS91, Nov88, Par70b, Pen92, Rei75, Tem90a, Tem74, Feh88, HW81, MM80, Mil70c, Ste72]. Metric [CRSW76, LA96, Wu90, DK83, Nür80, Som80, Som82]. Michael [dBJ82]. microbial [Hsu83]. Microscopes [Die94]. Microstructure [Lar98, SW91]. Mikusi [Bus72]. mild [Tre84, Tre88]. Mildly [Ros71]. Milin [Gas89]. Mindlin [AF90, AF96]. Minimal [Ahl93, GP84, GR75, Stu96, Wel91]. Minimization [Eva96b]. Minimizers [Dou97, HT98, Jan96]. Minimizing [PD71]. Minimum [Fra78, Lam78, Log79a]. Minkowski [Kal75]. Miscible [AM95, FL92, MA88]. Mittag [Cam89a]. Mittag-Leffler [Cam89a]. Mixed [Ban92, Elc75, Fan93, GK77, HS98b, IMP92, Joh75, Joh77, Joh80, McL90, Sho75, You75, YL94, AW78, AK82, Kap79a, LS86]. Mixed-State [HS98b]. Mixture [Che92b, SLS81]. Mobility [EG96]. mode [HS84b]. Model [AM95, ADH90, AF90, AF96, Avr91, BFGJ97, BCB82, BFF^{+}90, BH99, BLM96, BB94, BLS99, BIT91, Col97, DLY98, DL92b, Fan97, FT99, FH99, GP99, HNT98, HJ92, Has88, HT98, HR95, Hul97, Hut86, IMP92, JMR96, Ko98, KN98, KO92, Li93, LB96, Lia94, LY95, LWY98, MX98, NR95, Pes95, Ren91, Smi91b, Swi76, Sym91, Ter90, Thi93, Zha94, Avr88, BCT88, Bre91, CB89, CT88, Cos87, DS87, EF85, Gv85, HR88, HSX87, Lui82a, Lui82b, Lui85, Lui86, Swi80, Ter88, WHG88, WW88a]. Modeling [ABH94, Bra91, FR94, HY95, Ren97, STW99, TW91, Hsu83, JW88, WL82a, WL82b]. Modelling [AS87, Lin95, NP95, SS93a, GH87b, Mar85a]. Models [Avr93, BB83, Bon96, CC91, CE99, ES97b, FW92, Id94, KMM94, MLM97b, PP93, PP95, Wan78, Wan79, Wei82, Yan92, BNS85, HM89a, WG87]. Modes [ACG99, BB90]. Modified [CH93, Fan97, KV99, Nås78, PS72, Ryb99, Tem90b, May79]. Modifier [Fei78]. Modifiers [FK78]. Modulation [MP83]. Modulational [Wei85]. Moduli [Ubh76]. Modulus [EG91, Fet70]. Moment [Kra78, MP77, Sim77, Lan83a]. Moments [Vil92]. Monge [KS91]. Monodromy [BC90a, CR91, Har94b, Yok97]. Monosplines [BL84, MP77]. Monotone [CJ99, CSW80, De 77a, De 77b, Dol79a, HP93, Jia96, Ray81, RTV96, Ubh76, YZ94, Dol79b, LM82, Lui82a]. Monotoneity [Rei70]. Monotonic [AP74, KF90, RM78, HW84]. Monotonicities [LS94]. Monotonicity [Duc95, EL86a, IM78, LM77, LMS73, RE98, Ubh76, HW82]. Monotony [Har78]. Morawetz [AS79]. Morral [Gra95]. Morris [Gar90, Kad88a]. Morse [BMS83, CGR88, GM83, Li97]. Mother [Gus98]. Motion [Gra95, IS98, Kim90, Kim95, LSZ92, NV95, RS95, Smi91b, WY75, Zaj94, Sec88, She85, Vit79]. Motions [GN93, KMM94, Lin97, NP93, STW99]. Motivated [FH91]. move [ACK83]. Moving [GR76, CS85, Gus85]. MR [Ack80a, Ban79, Car86, GLY88, Log84a, Mad88, Obi81, Spe79, WH88]. Muldoon [EL86b]. Multi [BO91, BB78, Dav78a, Gin85]. Multi-Dimensional [BO91]. multi-parameter [BB78]. multi-time [Dav78a]. multi-turning [Gin85]. multicomponent [JW89]. Multidimensional [BH75c, CZ94, ES97b, Eva92, FH96, Lan74, PS98a, Ski80a, Bel81b]. Multidimensions [SF91]. Multiexistence [Cao95]. Multifractal [Jaf97a, Jaf97b]. Multifunctions [BU97]. Multigrid [Tem90a]. Multilateral [Gus87a]. Multilayer [ABH94, Col97]. Multilinear [PT78, Sri83]. Multiparameter [BB77, Bin82, Hai92, AR87]. Multiphase [Vis90]. Multiple [Avr93, Boe78, Car71b, CC94, Coh72, DG92, Gre80, Guc84, Has88, HLM92, Hra85, Kel73, Kla71, KZE90, Lae75a, Lae75b, Lip73, Lu87, MW87, MLM97a, MTH80, Nii97b, Olv77b, Par72b, Pey71, She72a, She72b, Siv71, dK71, Bri88, Car79a, Hag82, Kel72, Sel81, SW79b, Var88, WL82b]. Multiple-Back [Nii97b]. Multiple-Fragmentation [MLM97a]. Multiple-Front [Nii97b]. Multiplicities [Olv77a]. Multiplicity [HLM86, KK92, KO95b, LP91]. Multiplier [GT94a]. Multipliers [Imm90]. Multiply [AA88]. Multipoint [BG76, dM89, Har78]. Multiresolution [CD96b, DGH96, DDMT93, Zho96]. Multiresolutions [AEU94]. Multiscaling [PS98b]. Multispecies [FT99]. Multistratified [CD96c, CD95]. multistratifiée [CD95]. Multivalued [Eva96a]. Multivariable [Ros98, Gus87b]. Multivariate [FG92, HJ98, Han99, LLS97, Sto97, Wal97, Xu93, Xu94, Cuy83b, MS79, Mia87, Sil85]. muscle [CT88]. Mutualism [Hut86]. Mutualist [MH93]. Myelinated [Che92c, CB89, Cos87].

Nagumo [CB89, Che92c, Den91b, Flo91, KM86, Nii97b, San98, SS96, Zin91]. Nagumo-Type [SS96]. nán [HP83]. Nasrallah [Gus94]. Natural [Old70, Old72, TW89]. Nature [Sin75, de 80]. Navier [Che92a, BGT88, Car96, Che92a, Che92b, Dau89, Hof96, JT94, Kel88, Kni72, KK97, Sha73, Shi74]. Naylor [HP73]. Near [Chi92, Chi95, GK87, HL70, Kam92, Kov95, She80a, Ali95, Kam89]. Near- [Kov95]. Nearest [RE98]. Nearest-Neighbor [RE98]. Nearly [IL96]. Necessary [jCcW99, GT94a, GN83, LM95, Sin79, Rab84]. Negative [BRC72, Koh70a, Rot93b, HT87, Log83c, LP81]. Neighbor [RE98]. Neighborhood [AHIR93]. neighboring [Sch80]. Nerve [Che92c, Lab85, Lab89, Cos87, KM86]. Network [BB83, Fer86, HR95, Kaw79]. Networks [Dol77, Dol79a, Dol79d, Mat78, OFK77, Pet98, PG93, Zem79b, Zem79a]. Neumann [AB93b, BT91a, GK77, IM98, MTH80, PS88, PT99, dPFW99]. Neural [Pet98]. Neutral [KF90, Mao97, Pie73, Yan92, Cho76, Chu80, Sta86a]. Neutron [Bal99, CN77, LC84, Pao83]. Nevanlinna [Geo87]. Newton [HM84, Yar75]. Newtonian [NP93, Sha91a, STW88]. Nicholson [Dur78]. Nicholson-Type [Dur78]. Nielsen [Köl86]. Nikol′skiı [LS93]. Nikol′skiĭ [ELS94]. Nikol′skiĭ-type [ELS94]. no [Ack80a, Ban79, BF88, BF89b, Car86, Cig79, GLY88, Log84a, Mad88, Obi81, Spe79, Tah91, WH88]. Nodal [Kut84, She88a, She93]. Node [Che92c, Sch87c, Sch87b]. Nodes [BSX95]. Noise [MZ98]. Non [NP93, Sch73a, BH87, Tah90, Tah91]. non-convex [Tah91]. Non-Newtonian [NP93]. Non-Well-Posed [Sch73a]. nonabelian [Doo89]. Nonabsolutely [Ber75a]. Nonadiabatic [Gio90b]. Nonanalytic [Wal71a]. Nonautonomous [GZ92, Kai91, TKS93, BH87, SM85]. Noncentral [GG77, PR71]. Noncentrality [PR71]. Noncharacteristic [Col74, Ste75, Bel81a, Bel81b]. nonclassical [KA88]. Noncommutative [Fai71, PH75]. Nonconcave [BF98]. nonconducting [Blo80]. Nonconservation [Sai96]. Nonconservative [CH92, LT99]. nonconstant [Can85]. Noncontinuous [Ber70]. Nonconvex [BO91, BF98, FD97, LWY98, Tha93, Vis90, Tah90]. Nonconvolution [Kif77]. noncooperative [Swe89]. noncylindrical [Ale80]. Nonequilibrium [GW96]. Nonexistence [BT87a, Bre93, Che96, Gri82, Kho77, KO92, Lev74a, Lev74b, LP76, NS98, Nür80, Pao80, PT99, Tsu84, Wal75a]. nonfull [Mia87]. nonfull-rank [Mia87]. Nongeneric [Ste85a]. Nonhomogeneous [Any77, Mil70f, Sim90, Tef78, BG89, CH79, LS88a, LS89, She93, Ton82]. Nonhyperbolic [Den90]. nonintegrable [Gri80a]. Nonlinear [AFT98, ACG99, AB72, BH75a, Ban73, Ban75, Ban74, BC97, Bar75, Bar77a, BBDU93, Bax91, BHL75, BVB94, BMP92, BS85b, Bra91, Bra75, BHN72, Bro75a, Bus78, BE83, CC94, CK91, Chi92, CvL92, CS78, Cus71, DT96, DL96, Dea71, Dek71, Dem96, Dol77, Dun93, Elc71, EH84, ET71, FI96, FK94, FR91, FH91, GT98, GL74, Gom86, Gor75a, Gre89, Gre91, Hab76, Har72a, Har74, HO94, HH72, How75, How78a, How78b, Hsi73, HY95, Hua96, Id94, Ish96b, JS89, JMR96, Kel73, KV99, KR76, KS75, KS79b, KM91c, Kui76, KM95, Ład94, Lae75a, Lae75b, LB74, Lag75, LR77, Lan74, Lev76, Lev74a, Lev74b, LP76, LL93, Lia94, Lin94, LY95, Lon71, Lon73, Lon74a, LF98, MW78, MW83, Mat78, May70, Mei76a, Mil71a, MNY94, MW74]. Nonlinear [Nag78, Nak91, Noh74, NP93, Nus78, Nus90, OK94, Oht79, Okr91, Olv92, Ort94, Pao72, Pao74, Par70a, Par70b, Par72a, PT99, PS94, RR95, RB71, Ros71, Rot93b, Sat77, Sei87, Sha73, Sho75, Smi75, Sta80a, Sta71, Van98, Ven78, Wai76, Wal75a, WW93, WZ97, Wan98b, Wol93, Won88, XA91, ÁP88a, ARRR81, BN78, BH87, BS83b, Bar79a, BF89a, BT87b, BF88, BF89b, CF85, CD80, CDHL88, CW89, CN79, CN81, CE85, CE87, Daf82, Dam81, Dam87, DR86, DK86, Epp86, Fri81c, FL87, FK88, FR86b, GJ82, GLY84, GLY88, Gri78, Gri80a, Gri81a, Gri82, GM79, HLM86, HP83, Hen89, HW81, HV87, Hil82, Hir83, Hir86, Hiv78, HP89, HM89b, How82, How81, Jer78, KT79, Kel79b, KKY84, Kui80, Lan85, Lef82, LM80]. nonlinear [Lev83, LS86, Lui82a, Lui82b, Lui85, Lui86, MR83a, Mar85a, Mar89b, MS86, Nam78, OP87, PS88, Pey75, PR87, Rab87, RW78, Rob83, Rog88, Sac85, San83a, Sch87a, Spi79, Ste80a, Ste84, Swi80, TW89, Tsu84, Web79, Wei85, Wie87, WL82a, Zem79a, dT78]. Nonlinear-Diffusion [Wol93]. Nonlinearities [CD96a]. Nonlinearity [HS97, Ład94, Mos83a, Mos83b, Wes78]. Nonlinearly [LA91]. Nonlocal [BS97, Den95, EP98, Ito94, Lin94, Rog88, Shi93, Sou98, Gar87, Nas84]. Nonminimizing [CP97]. Nonmonotone [DeS87, Pou89, Van98]. Nonnegative [CD91, ES75, EG91, Tre73, Tre76b, Dun74, Jan84, Kim87, RA79]. Nonorthogonal [CL93, KM81]. Nonoscillation [ES75, Lon73, Utz70, CE89]. nonoscillations [Sin79]. Nonoscillatory [Har79a, Sin75]. Nonplanar [BH99]. Nonresonant [CSv87]. Nonself [Nay78]. Nonselfadjoint [Hil98, AL78, Pin88b, Tom83]. Nonseparable [BU97, BW99]. Nonsieved [Des98]. Nonsmooth [BS95, Car97, ES97a, LT92, SS95, Tol98]. Nonstandard [Str76]. Nonstationary [CD96b, GT94b, Zaj94]. Nontensor [She95]. nonterminating [GR89, VJ85]. Nontrivial [Kho77, Okr91]. Nonuniform [Wal92b, Zem82]. Nonuniqueness [Dix96, Saf99]. Nonvanishing [KV99]. Nonzero [WL82a, Tur89]. Norm [Bou72, AH83, LM82, Log79a, Log85]. Normal [DLY98, Chu84, Cyb84]. Normalized [Gas73, QV99]. Normed [Kol74a]. Norms [Tri97]. Note [Ano98, Ber75b, Car70a, Car75a, Car77b, Cha76, Dol81b, GU77, Gre80, Har79b, Hog98, Imm90, Lau75, Lev74a, LY70, Log79c, Mil87, MR80, Obi78, PG93, Red71, Sch74, Sha76b, SZL97, Sim77, Ski80a, SL78, SS85, Yam71, CG87, Cig79, EL86b, Jan84, Moe79, Obi81, Sin80, Son83, Sri83]. Notes [And80]. Nucleation [FL98]. null [Lou84]. Number [BL84, BSX95, KS96, Olv77a, RDLH77, Wan79, Log83d, NS85, Wie87]. Numbers [Car73a, Old70, Old72, She83]. Numerical [FI90, IMP92, Lin94, RW78, Sen96, Str77b, Ura84, Wan98b, AMS86, FMP86, HJ83, Kha87, MM80]. Numerical-Valued [Str77b].

O.D.E.s [Cic87]. Object [Tor96]. Obstacle [CL91, ES97a, MM92, NV95, BS83a, Col78]. Obtained [DQ90, Gor75b, Han90]. Obtaining [Ism74, Lun78, Ste72]. Occurring [Boe78, NB75, Nay77]. Ocean [BB94]. Oceanic [Dou98]. Oceanography [Dun93]. odd [Fia80]. Olver [SV94]. O’Malley [Olv78]. One [Ala95, AM95, AMS^{+}92, Avr91, Avr93, Car95b, Col74, DP97, EG91, FKS99, Fan93, Gor95, Gra95, HT98, Hof96, KSS97, Kim95, Kra89a, Lak90, MP77, Ser75, Ste91, Tre73, VS79, Yan92, Zha94, Bel81a, Car79a, Daf82, HV87, HSX87, Kim83, KS86, Lar88, Mar85b, Nas84, OP87, Rab87, Ful83]. One- [Yan92]. One-Dimensional [Ala95, AM95, AMS^{+}92, Gor95, Gra95, HT98, Hof96, Kim95, Ste91, Zha94, Daf82, HV87, Kim83, Lar88, Nas84, Ful83]. One-Parameter [VS79, Rab87]. One-Phase [Fan93]. One-Predator [Lak90]. One-Sided [MP77]. One-Step [Avr91, Avr93]. Ones [MS95, BJL81, BJL88]. Only [CD96a, HS97, BJL81, BJL88]. Onset [Veg93, Bar84a]. opérateurs [Tch96]. Operational [Kai92, Shu78, Str77a, Str77b, Str79, Tre79]. Operator [Amo99, BS96, BBF99, Bus72, CG71, CK95b, Cus75, EP97, FH90a, Fuh80, GHLS91, Gu92, Hua96, Hun73, Kui85, LR77, NW74, Olv92, SW75, SS92, Ung74b, Urb96, ki72, Ame82, Bry86, BF79, Cam89b, Cuy83a, Cuy84, DL84, GLR79, Gri85, KS80a, Kur79, Lan78, Lou80, Lui82a, Lui82b, Lui85, Lui86, Mar86, Sen85, Sta78a]. Operator-Valued [BS96, ki72, Sta78a]. Operators [Alp93, Bak72, BT91a, Ben85, BH75b, Bra74, Bra81b, Bro75b, Bro78, BS75, CPS98, Cho84, CSW91, CS97b, Col72, CH93, Dol70, Dol72, Dol81b, EKLOO92, FTZ88, FK94, Gra88, Hir75, KO89, Kra72, KS78, LJC97, LXLC96, Li85, Mai94, Mei76a, Mei77, Mil71b, Mor85, Rat74, Sen78, Sha76c, Ste75, Ste73b, AH83, Auc89b, Ban78, Bos79, BG75, But79, CV87, CLP88, Cos88, Cyb84, Dol79b, Fei79b, Fuh80, FK88, Gra80, Gut82, HS86a, JSW87a, Mer87, Oja78, Pin88b, Si78, Tch96, Tre79, TF84, Vic78, Zei80, vd86, van86].

Optical [ACG99, AK99, BB90, Mei76b]. Optics [BC97, JMR96, MPN89]. Optimal [BL84, CO92, Dol79b, Han99, Lar98, CS80, Hol87, Mel85, WW88a]. Optimality [Tha93]. Optimization [Ack77, Ack78, Ack80b, Gu92, Tha93, Ack80a]. Optimizing [HV97a]. Orbit [HZ94, Pér93, Nip86a]. Orbits [Che96, CDT90, Eck92, Kov95, Mir94, WH87a, WH87b, Ber88, FS79, Gru85, Stu89, WH88]. Order [AL77, Amo99, AF92, BG88, BC90a, Bal99, Ban92, BP96, Boe95, Bro75b, Bro78, Bus78, CPS98, CD74, Chu76, CC96, DGG98, Dek71, Dun90a, Dun90b, Dun94, EP97, Eli72, Eng99, Erb77, ET71, ES75, Fai77, FGHW91, FH90a, FMZ93, FR91, GM76, GB72, Gra72, Gre91, Gro76, Gro77, Gug71, HS77, Har74, Har77, Hei78, Ise85a, Ise85b, Ise85c, Jef91, Jon76, Kee72, KO92, Koh70a, KS85a, LS71, LW77, Leu73, Leu74, LP76, LB93, Lon73, McL76, McL78, Mil70f, Mil75, MW74, Nau76, Olv77b, Olv77a, Par70a, Par70b, PK91, Pog99, Rea86a, RB71, Rus77, Saf99, Sch76a, Sch76b, Sho70, Sin75, Sin76, Ski94, ST90b, ST91, St.84, Tho77, Tol98, Tro93, Utz70, Van98, Wai76, Wal71a, Won88, Wyr77, Yu74, Yu77, Ame82, Bar79b, Ben86]. order [Bog74, BD86, But80, BR84, BE86, CE89, Coh89, CLP88, DiG81, Erb76, ET72, FR86b, Gau90, GL84, GLY84, GLY88, Gun73, Har70b, Har70a, Hen87a, Hen89, HI89, HS86a, HKP87, How82, Kel79b, Kim86, KS86, LM83, Leu75, LS86, MR83c, McC78, Mil84, MM79, Min86, Mur83, Nau79, O’R87, Pat79b, Pat79a, PS88, Pet84, Pin88b, Sac84, San83a, Sch87a, Sho76, Tal80, Tef78, Tom81, Tom83, Tom86, TW79, Tre85, Wei84, WM87, Yos83]. Order-Preserving [ST90b, ST91]. Ordered [Tho77]. Ordering [Jon81]. Orderings [Zha86]. Orders [JP95, Ron97]. Ordinary [BP96, BR71, CPS98, Fai77, KS85a, Mat70, Nag78, Šim84, Šim85, Sta71, Tre76a, Win74, Wyr77, Bar79b, BS79a, But80, Gus87a, Har83, MO83, Min86, Mul81, ND81, Olv93, Per81, SS80, Sch80, Sch84a, Wei84, Whi80, dW80]. Organic [MX98]. Orientable [AJV94]. Origin [HL70, RB71]. Orr [Hab76]. Orthogonal [ASC72, AMRR92, Ask71, AW82, Bax71, BW99, BSX95, Ber75b, BLN87, Chi80, Coh75, DN86, DGHM96, DGH99, Ell82, ELM^{+}95, FK78, Fei78, FQ84, Gas77, Ger94, GS77, Ism74, KMT91, KF74, KV98, KSK78, Kra78, KLY94, Lan96, Leo82, Log90, Lou84, Maa96, MX91, MK78, Nev84, Ros98, SH71, SK76, Szw92a, Szw92b, Tre73, Tre76b, Val94, Van85, Van01, WN81, Wil80c, Wil91, Xu93, Xu94, Xu98, ZW82, AW79, CI87, DG88, Dom88, Gau89, GN83, Gus87b, Ism85, KM88, Koo85, Kow82b, MNT86, Mha88, ND79, Nev79, Oja78, Peh90, Per84, Sil81, van84]. Orthogonality [Des98, Kow82a, Rei97, Sei91, Jay88, van84]. Orthonormal [ASC76, CS93b, CD93b, DJJ91b, Dau93, Far73, MP98, Pil86, DJJ91a]. Orthonormality [LLS97]. Oscillating [Cao95, HT87]. Oscillation [BE86, CE89, Erb76, FMP98, GL84, GLY84, GLY88, Gre91, GLP91, Jon76, Jon81, KT79, KZ95, KKY84, KW83, Per88, PR91, Wal76, Yos83, But80, DiG81, GS89, Pat79b, Pat79a, Rei79, van84]. Oscillations [Bog74, Chi92, Nar93, Pes95, Won88, Yos85, ALS87, Dam87, Fen83a, Fen83b, Sin79]. Oscillator [BSSV98, JMR96, St.88, Ste77, Was73]. Oscillators [Chi95, KZE90, PS94, RE98, WO91, WH87a, WH87b, Yag92, EK84, ET89, GH84, Mor87, WH88]. Oscillatory [Ale90, BY72, FK78, Fen83a, Fen83b, GS99, HB73, HH72, HKP87, Kee72, KF90, Sch97, SS78, Sin76, Veg93, AB80, Son80]. Oscillatory-Tail [GS99]. Ostrowski [CPS87]. Other [JK75, Olv91a, Sch76b, Sha80a]. Outlets [SN99]. Over-Determined [EP98].

Overdetermined [HP98, Ben86].

Overpotentials [MS87]. Oversampled [CS97b]. Overview [Ful82].

P [Spe79]. P. [Vol87]. Packets [CL93, CD93a, She95]. Padé [Ehl73, JT75, Luk70, Luk75, Cuy83a, Cuy83b, Cuy84, Cyb84, Ise79, TG85, Wim85, WB95]. Padé-approximants [Cuy83a, Cuy83b, Cuy84]. Painlevé [MO83]. Palais [jCcW99]. Paley [CS93a, CS97b, MT95]. panels [KS79b]. Paper [Tem76, Ban79, Cig79, EL86b]. Parabola [SK76]. Parabolic [And91, Ban74, BL90, BB92, BK90, Bro78, CK89, CvL92, Col74, Col75, DGG98, DRF95, Dem96, DH99, Duc95, EI97, ES76, Eva96b, Ewi75, FP86, GM93, Hal91, Hil97, Hil98, Hol77, HS95, HY95, Id94, Ish96b, KL96, Lin91, Lin94, MR95, Mey77, MR98b, Mil74b, Mil74c, MY98, Mor89, Pie83, Ren97, RTV96, RS76, SV76, Sig76, Ste75, Suz95, TBD^{+}96, Wai76, Wal75a, WW93, Wei99, Yin91, Ale80, BT87b, BF88, BF89b, CW89, CLP89, Esh87, EW89, FH85, Hag82, Hea82, HJ83, Kui80, LM80, Lev83, MM80, NS85, RRW88, Ren82]. Parabolic-Hyperbolic [FP86, Ren97]. Paradox [BP90, Olv70]. Parameter [Bra71b, Bra76, Dea71, Dun90a, EM94, Fai77, Gor75a, Han73, Han76a, Han77, Hsi73, Kad72, Ko95a, LB74, MW75, Nag78, PR71, Rah76b, TF75, VS79, BB78, Bre82, Day89, Fie73, Fie83, Hal84, HJ83, JS87, MTH80, Nor88b, Per81, Rab87]. Parameters [BH75a, Bro75a, Gus92, Har94a, Har94b, Tak92, Zac74, CG87, Dun86, HM89a, HS86b, Pfa86]. Parametric [PFK95, SS89]. Parametrically [Tur96]. parasitic [WHG88]. Parity [Bal99]. Parseval [Osl72, Par75, RS74]. Part [Bai71, BS95, Kra75, BG88, CS77b, Che96, CD96c, Dau89, Gra88, GB72, Gra72, HP85, Jaf97a, Jaf97b, Jai81, Lab89, Sac74, She80b, Vas79b, VJ81]. Partial [AB72, AH72, BK95, BS74, BMP92, Cha76, CvL92, Col70, DT96, Fuj92, FH91, GL71, Gor75a, GP93, Hun73, Kry96, Liu97, McC97, MP98, Pog99, RS76, Sha91b, ST70, Sho70, Sho75, Vas79b, HK86, Hsu83, MO83, OP87, Vas79a, Web79, Wes78, WM87, Zei80]. Partially [Col94a, Sun93]. Particle [CPW98, SLS81]. Particular [BL90]. partly [Mar89a]. Parts [Yue73, Fia80]. Pass [Log84b]. Passage [San79]. Past [BC95, Elc82, Zha99b]. Patch [LN98]. Pattern [EM90, Veg93]. Patterns [CP90, NS98, PT97, Col82, CK84, CK85, CKP89, WW88b]. Pauli [CP84]. PDEs [MR98b, ZW85]. peaks [NS85]. Pearcey [Kam89, Par94]. Pearson [RS94b]. Péclet [She83]. penalized [BGT88]. Penetrable [Tor96, Kri80]. Penetrant [Hu91a]. percolation [EF85]. Perfect [MR98a]. Perforations [DD91]. Period [PK91]. Periodic [AHZ93, AB72, AH72, Ber76, BMY77, Cao95, Chi95, Chi77, CDT90, Cus82, FSW96, FS96, FW92, FR86b, FR91, GM96, Gri87, Gri91, How86, Jia96, JS96, JRRV97, KN97, KY75, Kra75, LS96, MM77a, Mil71a, MH95, NS91, Paş89, PT97, Pér93, Rab82, SM84, Smi86b, Smi91a, SH71, Ste80a, Ste80b, Swi80, Tak92, Thr94, Tur84, WH87b, AA88, Bar79b, Bar81, Bri88, BF79, Cic87, Dam87, Dav78a, Far78, FIT87, FS79, Gri89, MB91a, MB91b, Mor87, NT86b, NT86c, NNB88, Per82, Sei87, vV84]. Periodically [Bar83, Mia93, Yag92, GH84]. Periodicity [Nus78]. Periods [Gri91, HP93, Lin75a, Rot93a, Stu96]. Peristaltic [SLS81]. Permanence [CJ99]. Perron [Dat81, Šim88]. Persistence [FS89, HW89, Rua98, Thi93, WO91]. Perturbation [Ade73, Ade74, AF92, CR79, Coh72, Dor70, Feh88, GS93, GLR79, Hsi78, KoM98, Kif80, Kov95, MNY94, Par72b, Par72a, Sch90, Tem74, Tur84, Wan79, vV85, Ben82b, BCG89, Esh87, Fat85, Fen83a, Fen83b, Fen85, Kam79, Kel72, Kel79b, Kel89, LS86, MR83b, MR83c, ML84, PS88, SK87, de 80]. Perturbations [CC98, EG90, Eck92, EP99, Erh73, EW94, Hab76, Hal91, Har72a, Hol77, Hsi73, JS96, KP74, KM95, Lag75, LP94, May70, Tol98, Win74, Yar75, BT87b, BR84, Cha72, EW89, Gut82, HW83, Lon83, Tre85, vd86]. Perturbed [BFF^{+}90, Boh90, FKS99, Fra90, Gar92, Gor95, How75, How78a, How78b, Jef91, JW74, Kel73, MS99, Nii78, O’M70, Šim84, dPFW99, BS79a, Day82, DeS87, Far78, GH86, GH87a, How87, Kel84, Kol79, Nip86a, Nip86b, NF87, O’D84, O’M79, Ren84, SW86, Tre88]. Perturbing [Ber72]. Petviashvili [dS97]. Phase [AS87, ADH90, BFGJ97, BH75c, Bor99, BLM96, CP73, CC99, Fan93, GT94b, GZ95, GW96, Hab76, HB73, Ish96a, Kuw96, MW91, Mir94, Olv74, SS96, SZZ98, dK71, SAW87]. Phase-Field [BFGJ97, SS96]. Phaselocked [RE98]. Phaselocking [ET94]. Phenomena [Imm91a, Kel84, Moe84]. Phenomenon [Imm91b]. Physical [Dun93]. Physics [GY90, Mil71b, Hil82, Piq89]. physiology [MPN89]. Pick [Geo87]. Picone [Gau90]. Piecewise [BG88, DGH96, DGH99, FQ84, GLP91, LJ91, OFK77, PG93, Ray81, Zha99b, BG89, Lin86]. Piecewise-Linear [OFK77]. Piecewise-Polynomial [DGH96]. Pinning [LD97]. Pipe [AP89]. Pipes [Kot94]. Plan [Yu77]. Planar [AR98, CD97, IM98, MS95, Rot93a, TN94, Ter90, Ber88]. Plancherel [Ben91]. Planck [JKO98]. Plane [Dun94, Gio90b, GPG91, KMM94, LB93, Man91, Mey76, Mil76, NY97, Ryb99, ST90a, SS75, Slo75, Yu74, BW83, BSL87, Lei81, Lon88, McC83, PP83, She88b, Wag89]. planetary [Gv85]. Plasma [GY90, Liu95a, BS83b, BGS86]. Plasmas [GS99]. Plastic [Chr80, GN93, Rei91, SS93a]. Plasticity [Car94, NR95]. Plate [AF90, AF96, BP90, Elc82, KLV93, Kim92, MM92, Cof82]. plateaus [EK84]. Plates [Bre93, Chr80, DGR99, SP79, Wei88]. Plemelj [Ort73]. Poincaré [BSZ77, CS79, ZD84]. Point [Car95a, Che92c, CM80, Die95, Dun90b, Dun94, ET71, Gal95, Gre91, HS97, Har74, HJ92, Kra89a, Kug74, MW91, Nii78, O’M70, Ohk91, PM82, Par70b, Par72a, SW73, SZL97, Ski94, Thi93, Tip74, Was70, Was74, WB95, AGX89, BN78, BJL81, BJL88, BD86, Chi80, DiG81, Gin85, Ira89, Lam80, LK89, MR83c, MTH80, Pet81, Sch84a, Sch81, Sib81, de 80]. Point-Dissipativity [Thi93]. Point-Node [Che92c]. Points [AM94, Any77, BS92, CR92, Dor70, GB72, Gra72, HP93, How75, How78b, JS96, Kra89b, KP74, KM95, Lae75a, Lae75b, Leu73, Leu75, Nau76, Olv77b, Olv77a, Sat77, St.96, Wal71b, Bra81c, DeS87, GH87a, GH88, HIN85, HS84a, Kam79, KZ88, Min86, SS80, Sch80, Sha80b, Tom81]. Pointwise [BR71, Cos84, Cho76]. Poised [Gus90, HBL76]. Poisson [Bra81a, Bre73, Bre78, You75, Ban75, Car73b, Elc72, GR83, GR84, GR86, HQW74, Kai91, Pan72, Rah77, Ste91]. Pol [Mac90]. Polarizable [CJ88]. Pole [Dun90b, Dun94]. Poles [RDLH77]. Pollaczek [ASC87, CI87, Ism85, Ism94]. Polya [GM92, Mul87]. Polyconvex [Dou97, IL96]. Polygamma [AW98, Gro76, Gro77]. Polyhedra [Gus98]. Polyhedrons [KS79c]. Polylogarithms [Köl86, AK88]. Polymer [Hu91a, Ren91, TW91, FMP86, JW88, JW89, Ren84]. polymeric [MR83a, Ren82]. polynômes [Zen90]. Polynomial [All79, Bus77, CD97, Coh71a, DJL93, DGH96, DGH99, Hal77, Kem82a, Kem85, Lam78, LT94, Per86, Rat74, Ray81, Rus78, She80b, STW88, Tre76b, Ubh76, Ver72, Wal88, Wal92b, Wyr77, YS93, Bre85, Cam87b, EL87, Fia80, Kaz84, Mha88, PT78, Per84, Per88]. Polynomials [ASC72, ASC76, AMRR92, Ask74b, AW82, AGV82, BFW86, Bax71, BH93, Ber75b, Bil74, BLN87, BR82, BG74, Car72, Car71c, Car74c, CA71, CI91, Coh75, CR92, CJ76, De 77a, Des98, Det95, Dil88, DGH99, DR74, Dun78, Dur77, EG91, ELM^{+}95, EKLOO92, FQ84, Gas73, Gas77, Ger94, GS77, Gil75, GW94, Han90, HS77, HPR97, Hor75, How74, IN90, Ism74, IK76, Ism81, IM87, JNW94, KM91a, KMT91, KM91b, Kan93, KM93, Koe94, KV98, Koo74, Koo75, Koo77, KSK78, Koo91, Koo93, Kra78, KLY94, Let94, LR72, Log90, MX91, Mil87, MK78, Mul75, Nev84, NEM94, Obi78, Pat71, Rah76a, Rah76b, Rah76c, Rah77, Rah78, Ros98, SV76, ST90a, Sch76b, Sch74, SS94, Sle72, SK76, Sto97, Szw92a, Szw92b, Tre73, Val94, Van85, VFO87]. Polynomials [VK91, Van01, WN81, Wil70, Wil72b, Wil80c, WZ94, Xu93, Xu94, Xu98, YH82, ZW82, AMS86, ASC87, Aom87, AW79, Bre81, BS79b, BS80, BI83, CI87, Chi80, Chi87, Cie87, DG88, DN86, Dom88, Dun74, Ell82, EMN94, FR86a, FW88, GR83, GR84, Gas85, GR86, Gat87, Gau89, GN83, GJZ88, GLR79, Gus87b, Ism85, Ism86, ILV89, Ism94, Jay88, KM88, Koo82, Koo85, Kow82a, Kow82b, Kus88, Lai80, Lai82, Las83, Leo82, LS88b, Lew79, Log83f, Log84a, LMN87, MNT86, Mar86, May79, MK83, Mir83, Mon80, NR85, ND79, Nev79, Obi81, Peh90, Rah79, Rah81b, RS85, RV86, RA79, Ric89, Sim80, Sri83, Sta80b, Vre84, Zen90, van84]. Polynomic [Por80]. Population [Bra75, BT91b, CC91, EM90, KoM98, KM91c, Red85, Rua98, Swi76, Wan80, Yan92, BH86, HM89a, HM89b, Lan85, Lui82a, Lui82b, Lui85, Lui86, Mar81, Swi80, WHG88]. populations [Bra83]. Porosity [ADH90, BLM96]. Porous [AM95, BNS98, BC95, BCT98, CQ97, DuC97, FL92, FFW80, FT88, Gil79, HS96, KV91, ACK83, van79, vd87]. Posed [Ade73, Ade74, Car99a, Col73, Fra78, Sch73a, Sho70, Ame82, BW83, Mil70c]. Posedness [FR94, CO89, Tad86]. Positive [Ask71, Ask74a, BI89, CP97, Cus71, Ede79, FS96, FK94, Gas75, Gas77, GU77, Iff87, Jus74, Kem82b, Lae75a, Lae75b, Lah93, Lam78, LC84, LP91, Lor93, LP81, Mug75, Peh81, Rah76a, Rah76b, Rah76c, Rah78, Rah91, Rea83, Rea86b, WZ97, Wol93, Zha95b, Akı88, Bel82, BB86, BB79, CN81, DK86, Ha86, Hir86, Kur79, Log83b, Log83c, Log83d, NS85, Rea84, Sel81, Sta78a, Swa82, Ter85]. positive-definite [Log83b, Log83c, Log83d]. Positivity [AHP75, FI75, GR83, GR86, GRZ83, Han90, IT79, Jan81, MR94, CG83a, CN79, GR84, Mar80]. Possible [Gal91]. Postulational [Mei76b]. Potential [GO71, HQW74, Hof99, PS98a, Sha91a, SS95, WZ97, Zha99b, Nas84, Pag86]. Potentials [Jos99, LZ95, LPR92, Pie74, Sch88, Wei74, ELS94, Fry78, HS86a]. pour [BH87, Che92a, CD95]. Power [Ask74a, BG75, Car77b, DRF95, GH75, Gor75b, Imm91a, KSK78, LT81, Old70, Old72, Sle72, AM83, Car79a]. Powers [LB74, LMN87]. Precessing [LF98]. Precise [HZ94]. Predator [Din89, Hai92, Lak90, MR93, Yam90, Che81, LC88]. Predator-Prey [Din89, Hai92, MR93, Che81, LC88]. predators [HSX87]. prediction [MS79]. Predictor [Mia93, Mia87]. Premixed [AV90, BH99, BSL87]. près [Ali95]. Prescribed [CD91, CZ91, Mil70c, Rul87]. presence [CS85, Coo88, Hea88, IR89, Sec88]. Preservation [AHP76, CC98]. preserve [BR84]. Preserving [BS97, Hu91b, Sen78, ST90b, ST91, Bra81c, CN79]. Presistence [Mag87]. Pressure [Sim90, Sun93]. Prestarlike [CS84]. Prey [Din89, Hai92, Lak90, MR93, Yam90, Che81, HSX87, LC88]. Prey-Predator [Yam90]. price [DS82, SDL84]. price-dividend [DS82, SDL84]. Prime [HP93]. Principal [BS95, Ben81, Rao75]. Principe [CD95]. Principle [Ben84a, Eva96b, Lab99, Yu74, CD95, Pin88b, Swe89, Whi80, dM86]. Principles [Bin82, BY91, Fra78, O’R93, SS77, Auc89a, Auc89b, Ban79, Lar81, Mad85a, Mad88, PP79, Spe79, TW89]. Priori [DDD96, ES76, HFL99, Lev76, Pay70a, PS76, Sig76, DK86, EH84]. prisms [Kin86]. probabilistic [Kad87, Sim80]. probability [NNB88, WG87]. Problem [AHIR93, Ack78, Ack80b, Ade73, Ade74, Av91, AC98, AV90, AC94, AN98b, AK99, Avr91, Aye83, Bal99, BC97, BO91, Bar77b, Bar91, BK95, BT87a, BBN93, Boe78, Boh90, Bor99, BR78b, Bro96, CS99, CP73, Car76a, CC94, Cha70, CL91, CvL92, CZ91, Col94a, Col97, CO88, CG71, Col74, Coo74, CD90, Dea71, DLY98, Dix96, Dör91, DuC97, Dun93, Edw74, EL89, EP97, ET71, FKS99, Fan93, FP86, FN75, FI96, Fra90, FR94, FRZ95, FH96, FH99, GY90, GIO90a, GR92, Gio93, GH75, Gom86, GK77, GZ95, Gre74, GM89, HS97, Häh98, HP85, HLM92, Hof99, Hol77, Hsi78, Hu91a, HW92, HP92, Id94, Ish96b, Joh75, Joh77, Joh80, KSS97, KK92, Kel73, KL96, Kla77, KI95, Kli97, Kni72, KT99, Kut84, Lam78, LMPS90, LSZ92]. Problem [Li93, Lin91, Liu91, MW75, MM97, MS87, McL76, McL78, McL90, MS99, Moc74, Nau76, NB75, NS91, Ngu90, Nii78, Nor88a, NR95, Ohk91, Osh93, Par72b, Pay70a, Por76, Pos90, Rei93, Rei91, Ren90, Ros71, Ryb99, RS91b, SR93, SS93a, Shi93, SZL97, Sig76, Sin97, Ski94, Sle95, Smi71, Smi75, SN99, Ste75, Ste91, Sym91, Szw92a, Szw92b, Tem74, Tep74, Tep75, Ter99, Tor96, Ves97, Wyr77, Xie90, XA91, Xu92, You75, Yu77, YL94, Zag98, ZZ90, vV85, Ack80a, Ack81, Ale80, AW78, AS79, Azz80, Azz81, BS83b, Bel81a, Bel81b, Ben82a, Ben86, BEGv87, BSL87, Bre82, But79, CFV81, CF85, CD80, CS86, Cha72, Chu84, Cof86, Col82, CS85, CE87, Daf82, Dam81, Eng87, FLLL89, FMP86, Fat85, Feh88, FIT87, Gab84, Gri85]. problem [Gus85, HP83, Hil82, Hin81, Hol87, Iff87, JW88, JW89, Kel72, Kel79b, Kim87, Kna86, Lan83a, Lan85, LS88a, Lef82, Lin86, LK89, Mar86, Mar85a, McC81, Mil84, Moe84, Mos83a, Mos83b, NNB88, O’M79, Pao83, PS88, PR87, RA79, Rei79, SS80, Sch80, Sch84a, SK87, Sib81, Smi89, SAW87, Sta78a, Sym81, Sym86a, Ter83, Ton81, Ton82, Tsu84, Tsu89, Val86, Vog85, Wag83, Web79, Wei71, Wei84, Zho89, de 80, dW80]. Problems [Ack77, AC80, AS87, AF92, BG88, Ban92, Ban73, Ban75, Bar85b, Bar85a, Bar88, BG76, Bax91, BP96, BB77, Bra76, Bra75, Bus78, CD96a, Car99a, Car99b, Car94, CK89, Chi74b, Coh72, Coh76, Col73, CS78, Cus71, DRF95, DT81, Dor70, Duc95, EP98, FGHW91, FMT96, Fis75, FFW80, FS77, Fra78, FH91, Gil70, GT98, Gor95, Gre89, Gre91, GW96, Guc84, HG99, Har72b, Har74, Har79a, HL96, HP98, HR99, How75, How78a, How78b, IT79, Jan86, KN97, KO95b, Kui76, LB74, Lag72b, LW77, LVLH92, LM96, MW83, MW93, Mey77, MM92, Mil76, MW74, Mul78a, Nag78, NU95, NT96, NV95, NP95, O’M70, O’M71, Olm77, Pao74, Par70a, Par70b, Par72a, PS76, Pen92, PB85, PT99, PD71, Rei73, Ren96, Ric90, RZ91, Say71, Sch73a]. Problems [SS77, She72a, Sho70, Slo75, SZ87, Spe89, SHM89, Sta71, SS95, SF91, Tip74, TY77, Van89, WW93, Was70, WO91, Wei99, Wol93, Zha95b, dM89, dPFW99, ka76a, ka76b, ka77, AGX89, AO81, AA83, Ame82, AR87, Aro81, BG89, BN78, BJL81, BJL88, Bar84b, Bel81b, BT87b, BB78, BB79, CL89, Dav78a, Day89, DeS87, DG88, DiG81, DL84, DK86, EH84, Esh87, Gin85, GR79, HLM86, Har78, Hen87a, Hen89, HS84a, HP89, How82, Ira89, Jer78, Kam79, Kap79a, KKZ84, Kel89, Kol79, Kor88, KZ88, LS89, Log83b, Log83c, Log83d, LS86, LS87, LP81, Mar82, Mar83b, MR83b, MR83c, MS88, Mil79, Mil70c, MTH80, Mul81, O’D84, O’R87, PP83, PS88, Pet81, Pet84, Rab87, Ren84, Rog88, RS86, Rul87, SW86, Sch81, Spe79, Vis86]. problems [Wie87, WM82, Zio85, dM86, dT78, Bel81a]. Process [BLM96, JS89, Kel73, CF85, MS79, She86, Vog85]. Processes [Dat72, GZ92, Mak76, Mei77, SW75, ILV89, Mei79, Mia87, van86]. Product [Cha78, Dur78, GR84, Hen87b, IM78, Koo74, RV86, She95, Sta80b, AK88, BHS83, Car80, GR89, GJZ88, Kus88, Lai80, Las83, Mar86, MK83, Moo73, Ste84]. Products [BL74, Bra74, CH92, Gil75, Har77, JM72, LT99, LY70, QV99, Bra86, DL92a, PV84, Piq86, Piq89, RS85, Vol84, Zen90]. produits [Zen90]. Professional [dBJ82]. Profile [Eng99, CK87]. Profiles [BB92, FS95, KH95, Mic96, vd87]. Projection [Lai82, NR85, Nür80, Oja78, Som80, Som82]. projections [DK83]. projective [BKW85]. Projectors [BAR90]. prolate [Dun86, Per86, Per88]. Proof [Car71c, CJ93, Eng99, Gar90, Koo74, Koo75, Kug74, Mei72, MW74, Vem95, Gas89, Kad87, Kad88a, Kad88b, Lai82, Mar86, MQ85, Mim88, Rah79, Zei87]. Proofs [Ask74b, Car78c, Sch73b]. propagating [Lei81]. Propagation [BCT98, CS97a, Gao93, Ko95a, Lin95, Can85, Col82, CK84, CK85, Dav79, DM82, Lar88, Sym86a]. Propagators [Mas78]. Proper [Pou83]. properly [Ame82]. Properties [AHP75, AHP76, AMRR92, AFT98, Bak72, Bou72, Coh71a, Cos73, DuC97, Gla84, HK90, How74, JPS87, KS80a, KS85b, LMS73, Lu99, MR79a, PJ73, Pou83, RM78, Rei70, RR95, Ryb99, SS92, Wei74, BG84, BR84, Cam87b, DG89, EL86a, EL87, FK88, Har78, Kol79, KKY84, LM77, ND79, Pat79a, PS82, Per88, Sha77, Sin79, Vit79, Yos83, van84, van86, MB91b]. Property [CC98, Kad72, MST96, Tho77, Per84].

Protocell [FH99]. Provide [Ron97]. Prüfer [Ben84b]. Pseudo [Sig76]. Pseudo-Parabolic [Sig76]. Pseudoinverses [Har79b]. pseudomonotone [AO81]. Pseudoparabolic [BBDU93, BS85b, RS76, ST70, Sig74, RS78]. pulled [PGRSV89]. Pulse [Flo91]. Pure [ABH94]. Purely [Dun90a, Lab99]. Pyramidal [MT95].

QD [Cuy84]. QD-algorithm [Cuy84]. Quadratic [ASV84, Car76b, Edw74, ILV89, Lah93, Rei75, Gas85, GR86, How81, JAK88, Mad85a, Mad88, Mar80, Ver80]. Quadrature [CS93b, LMN87, MP77, Peh81, Sak93, Peh84a, Wim85]. quadratures [BC79]. Quadrilateral [Hug75]. Qualitative [ABH94, Avr91, BS83b, Che92c, DGG98, Lew74, Yan92, GK80]. Quantitative [BHN72, Noh74, Mha88]. Quantum [Koe94, KV98, Koo91, Koo93]. Quarter [Lin75a, NY97, Akı88, BW83]. Quarter-Plane [NY97, BW83]. quarter-space [Akı88]. Quarters [PM82]. Quasi [Bra86, CW99, FMT96, FM92, FD97, Hu90, JT94, JS96, JRRV97, LS96, NS91, SM84, ST90b, Suz95, TF75, Xu94, Yag92, vV85, Ali95]. Quasi-Convex [FM92]. Quasi-Geostrophic [CW99]. Quasi-inner [Bra86]. quasi-linéaires [Ali95]. Quasi-Linear [FMT96, FD97, LS96, Suz95, TF75, vV85]. Quasi-Orthogonal [Xu94]. Quasi-Periodic [JS96, JRRV97, NS91, SM84]. Quasi-Periodically [Yag92]. Quasi-Static [EW94]. Quasi-Stationary [JT94]. Quasi-Variational [Hu90]. Quasiconvexification [Lar98]. Quasiconvexity [FM99]. Quasigeostrophic [BB94]. Quasilinear [AP93, CL91, Gal91, Gal95, Jef91, KA88, Lin91, LB93, Sha91b, WW93, Ali95, ÁP88a, DeS87, EW89, Ren82]. Quasiperiodic [LZ95]. Quenching [CK89, RR84, CL81, LM80, Lev83, Smi89]. Questions [KLV93]. Quotients [QV99].

R [Spe79]. Racah [AW79, GR84, Rah81b]. Radar [AT85]. Radial [CZ91, FM98, Gla76, LXLC96, LP91, QSW93, Yan96, Bra81a]. Radially [BT91b, CLH92]. Radiating [KN98]. Radiation [AMS^{+}92, CS77b, Chi74a]. radiative [Gv85, Mer87]. radius [Vic78]. Radon [Dea79, Fri95, Kat97, Lou84, MN86, MQ85]. Rahman [Gus94]. Raised [Old70, Old72]. Raman [MS92]. Ramanujan [Dut81, GM94, Kad87, Lin78, Mim88, MZ92, RS94a]. Ramanujan-Type [RS94a]. Random [BH93, CG83b, Des98, For90, GG77, HL74, MMM97, Mei77, Lei81, Mia87, Sil81]. Range [Kat97, FS83, HS86a, WM82]. Rank [DP97, KS86, Mia87]. Rank-One [DP97]. Rapid [Koh70a]. Rapidly [Rei91]. Rarefaction [LMN98]. Rate [BL74, CG83b, FM98, TW97, Wan98b, Bra83, Kur79, LM82, LS87]. Rates [BT91a, Mak76, Riv93, Bra79]. rather [BEGv87]. Ratio [Fre92, Bel82, Chi83, Fre87a]. Rational [Ask74a, BR78a, Bor85, GNP95, HIN85, Joh87, Lon74b, Mak76, Mul75, Nås78, Nak91, Rah91, Dil88, Lan80, Rah81a, TG85]. Ratios [AVV90, AB93a, Nås78]. Ray [Qui93, FS83, Mar84]. Rayleigh [Obi81, Ban73, Ben82b, Lir71a, Obi78]. Razumikhin [Mao97]. Razumikhin-Type [Mao97]. reach [DS87]. Reactants [Avr93]. Reacting [Che92b, HSX87]. Reaction [AV90, Avr91, Bon96, Car95b, DS95, EM90, ET94, FHM97, FRZ95, FH96, Gar92, HS94, HG99, HJ92, Hut86, IM91, KoM98, LS96, Mor90, MH95, MNY94, NMIF90, PS97, Sch98, Sou98, Tak92, TN94, AA88, Cos84, DH84, Gar81, HMP87, Kha87, LPSS89, Mar89a, Nis82, NF87, Pao87, Ter83].

Reaction-Diffusion [ET94, FHM97, FRZ95, FH96, Gar92, Hut86, KoM98, LS96, Mor90, MH95, MNY94, NMIF90, PS97, Sch98, Sou98, Tak92, TN94, AA88, Cos84, DH84, Gar81, HMP87, Kha87, LPSS89, Mar89a, Nis82, NF87, Ter83]. Reaction-Diffusion-Advection [EM90]. Reaction-Hyperbolic [Car95b]. Reaction-Variable [IM91]. reactions [MS86, WL82b]. reactive [Bre91, DJ87]. reactive-Euler [Bre91]. Reactor [BHN72, Noh74, Ost96, GK80, Pao80, She83, Wex79]. Reactors [Ale90]. Real [GU77, KS91, Mey86, RB71, Sap71, Ste70a, WZ87, de 78, Auc89a, Bel82, Chu84, LS88b]. Realizability [Bos79]. Realization [AM80]. Rearranged [CF98]. Rearrangements [Dou98]. Reception [CS77b]. Reciprocal [AM83, Fet70, LB74]. reciprocals [LS88b]. Reciprocity [Cha76]. Reconstitution [Rob85]. Reconstruction [FG92, MN86]. Recording [CL90]. Recovering [Jos99]. Recovery [EI97, Log90, LPR92, PS98a]. Rectangle [CL73, Cof86]. Rectangular [Lu90a, LKMT88]. Recurrence [CC96, ELM^{+}95, Lew86, Lir71a, Sri90, DN86, GN83]. Recurrences [Car78a]. Recursion [Jus74, Rea87, Kow82a, Kow82b, Sin80]. Recursions [vdS76]. Recursive [Let94, SN71, OP87]. Recursively [CJ76, Jay88]. Reduced [Joh77, Chi74a]. Reducibility [JRRV97]. reducing [Dol81a]. Reduction [Bra71b, Car70b, Gra72, Imm91b, JS89, MG70, BJL81, BJL88, MM80]. Refinable [Jia98, LLS97, Ron97, She98, Zho96]. Refined [Pan96]. Refinement [GM92, HJ98, Hog98, Vil94]. reflecting [vP86]. Reflection [Lab99, Sym91, Yu74, Gv85, Sym81]. Reflux [Jam98]. reformulation [Old83]. Refraction [Hof99]. Region [HW92, IM98, SK76, Hom79, She88b].

Regional [GV93]. Regions

[HS98b, SV76, Far78, Veg88]. regression [Ric89]. Regular [Bro96, Hal77, BJL81, BJL88, BD86, Kel89, SS80, Sch80, Sch84a]. regularité [Her95]. Regularity [And96, BG88, BG89, BS95, CK95a, FH91, IS98, Jia98, Kim95, LS88a, LMW96, LB93, Luc88, MS99, Rio92, Sak93, Sch81, Vol95, DL91a, DL92a, HV87, Her95, Mar89b, NNB88]. Regularization [BBDU93, EP98, Egg93, Epp86, KO96, Mai94, DMV98, NW74, OK94, Sho76, HS86b]. Regularizations [Sle95]. Regularized [BO92]. Regularizing [Die94, KKZ84]. Reimann [BO91]. Reinforced [Paş89]. Reissner [AF90, AF96]. Related [BS76, Car77c, DLY98, Elc72, EW73, FN75, Hua96, Id94, KM85, KLV93, Koe94, Lan74, Mor85, Mul75, PW70, Son70, Ung74b, Vis90, Ame82, Bre85, Bre81, Car79b, Car86, CS89, DG88, Dil88, Gab84, GR89, How87, Mon80, Ngu89, Str80, Tem82, TF84, Val86, Wim85]. Relating [PP93, PP95, Cie87, GN83]. Relation [CRD88, Fet70, Jus74, LMPS90, Osl72, RS74, Ste91, EL87, Kin86, Sym86b]. Relations [BR81, Car70c, Cha76, CC96, ELM^{+}95, Fet78, Lew86, Ort73, DN86, Fet77, Rea87, Sas91, SW79a, Vol82, Wal79]. Relationship [Imm90]. Relative [Ber99, CS99, Kru90, WZ94, MS88]. Relativistic [And96, HPR97]. Relaxation [FL98, Jam98, Lar98, LWY98, MR98a, Nar93, NY97, STW99, TW97]. Relaxed [AK99, FMT96, Thi93]. Relevance [Imm91a]. Relevant [Tro93]. remainder [JW80b, Lut78]. Remainders [Jon94]. Remark [CM80, GSZ94, JT94, Ban79]. Remarks [Che92a, KP74, Lev74b, MS84, Par72b, Par72a, Rot93a, She88b, Tem76, CGR88]. Remarques [Che92a]. Remez [ELS94]. Remez- [ELS94]. Renewal [SW75, Mar81]. Renormalized [MP97]. Repeated [Amo89, GH84]. representable [Yom83, Yom86b]. Representation [Alp93, Ask74b, Car74d, D’A75, Dol70, KP92, Koh70b, Kra91, LT99, Mei77, Mey74, Mey77, Pan71, Sat77, Sho72, Si78, Yu74, Akı88, BG96, Kap79b, KS86, MB91a, Wil86, Yom86a]. Representations [Ahl93, BT96, Har94b, KMM94, KV98, LY70, Mey76, Ros98, de 78, Koo85, Mer77, Sha80a, Vol84]. Reproducing [Mei79, NW74, Sei91, NR85]. Reproducing-kernel [Mei79]. Reservoir [CE99, Coc87]. Residue [SW75]. resistance [Ack81, Zem81]. Resistive [PG93]. Resistor [OFK77]. Resolution [CD93b, JAK88, MN86]. Resolvent [Ben85, CH82, Han76b, Ros78, KS80a, McC81]. resolvents [GK84, Gri80b, Gri81b, HW82, JW80a, JW80b]. Resonance [Chi92, Chi95, Hal84, Ohk91, Olv78, San79, Bri88, GH84, Rob83, Sib81, de 80, vV84]. Resonances [EP99]. Resonant [Lab99, LVLH92, Pér93]. Respect [EM94, Kad72, Rah91, Zha95a, ki72, Bre82]. Response [Kra75, Coc87]. rest [ET89]. Restricted [Har79b, Mad85a, Mad88, MN86]. Restrictions [Cyb84]. Result [BMP92, Bre93, BV92, CJ93, CO88, Fri95, Jen91, MLM97a, Gri78, Gun73, Ngu89]. Resulting [Ren91]. Results [Dur78, Dut81, GC95, IMP92, Jaf97a, Jam98, KO95b, Let94, Mor90, O’R93, Sri70, BS83a, BE86, CW89, Cos88, Har70a, IR89, Lun87, ML84]. Resurgence [BSSV98]. Retarded [DC85, Sin75, Sin76, ALS87, Bog74]. Reversal [Say71, Bar84a]. Reverse [Sch90, BCG89]. reverse-biased [BCG89]. Reversible [Bon96, Pér93, Van90, ML93]. Reversion [GH75]. Revisited [CF96, MZ92]. Reynolds [CL86]. rheology [Ren84]. Riccati [AH88, Ahl93, BRC72, BE83, Erb77, HKP87, JAK88, Keg83, KS80b, Kui85, LT92, Rei70, RDLH77, Wex80]. Ridge [Pet98]. Riemann [BHL72, Bra76, Fan93, Fuj92, HR99, Ira89, Iwa88, JW88, JW89, LSZ92, Li93, Lin86, Rus77, SR93, Sin97, Sle95, Wag83, ZZ90, Zho89, Zio85, ki72]. Riemannian [BW80]. Riesz [CC98]. rigged [But79]. right [Cof82]. rigid [Blo80]. Rigorous [CC96]. Rikitake [Bar84a]. rim [PGRSV89]. Ring [Col94a, HT71]. Ring-Shaped [Col94a]. rings [Tur89]. Ritz [Sha80b]. Rivers [Die94]. Roads [HR95]. Robertson [Gas89]. Robustness [FSV99]. Rod [Wol96, Buz86]. Rods [Fra90]. Rogers [Gas85, Mil94]. Role [dPFW99]. Roll [Kuw96]. Root [Gar90, Mac82, Gar88, Zei87, Zei88]. root-system [Zei88]. Roots [Bou72, SV83, Obi86, Obi87]. Rotation [KS92]. Rotational [JM96, Dri87]. Rough [CWZ99]. Rule [Osl72, Osl73, Eng82]. Rules [Str79]. Runge [Col82, RS78].

S [IMP92]. Saddle

[Sch87c, Sha80b, Sch87b]. Saddle-Node [Sch87c, Sch87b]. Saint [Kin86]. saline [Kna86]. Salle [RRW88]. Sampling [FG92, GC95, JK75, Liu96, Wal88, Wal92b, Doo89]. sand [EF85]. sand-bank [EF85]. Satisfy [Wim74, Cof82, Wad80]. Satisfying [Elc75, KLY94, Mil70f, AK82]. Saturation [Det78, KN72, May79]. Sauts [Die86, Die88]. Scalar [Die95, Ter99, Wan98b, FH85, GR85, Lin86, Luc88]. Scale [All92, BB94, Gor95, LW95, MM77b, Vil92, DL91a, DL92a, Dav78a]. Scale-Invariant [Gor95]. Scales [ABH94, Bar84b, San80]. Scaling [Ast91, Hua96, LMW96, Mey84, Wan98a, Her95]. Scattering [Av91, Boe78, CWZ99, CRD88, CK92, CH93, CK95b, DG83, Ger94, GK87, GM76, Häh98, HNO98, HL93, HL96, Jos99, JP95, Joy97, KN97, KI95, Lab99, LVLH92, MS92, PS98a, RS91b, Sat90, Tor96, Col78, MO83, Zho89]. Schaefer [Spe79]. Scheme [Swe98]. Schemes [AF92, CD96b, Gri96, HJ98, LA96, Llo80, Rio92, Wal97, YS93, Leo82, OP87]. Schrödinger [BH87, Av91, BH87, Ben82b, Bil90, CS89, CK92, Fat85, FI96, HO94, Kho77, KV99, LMPS90, Nas84, PS91a, Rea87, Rei86, RZ91, Sac83, SW79a, Ste91, Swa82, Tsu84, WZ97, Wei85]. Schur [CPS87, Thr94]. Schur’s [LR72]. Schwartz [Lee75, Ort80]. Schwarz [Hug75, Lab99, Pfa86]. Screen [Pen92]. SDE [Nar93]. Second [AL77, AF92, BG88, BC90a, Ban92, Boe95, Bus78, Dek71, Dun90a, Dun90b, Dun94, Eli72, Erb77, ET71, ET72, Fai77, FGHW91, FH90a, FMZ93, FR91, GM76, Gug71, Har74, Har77, Hei78, Jef91, LW77, Leu73, Leu75, LP76, Lip98, Lon73, LS86, Mil70f, MW74, Nau76, Olv77b, Olv77a, Saf99, Sin76, Spe89, St.84, Swe98, Tho77, Van98, Wai76, Wal71a, Won88, Bar79b, BD86, But80, BR84, BE86, CE89, CLP88, Dur79, FR86b, GLY84, GLY88, HS86a, HKP87, Kel79b, Kim86, KS86, MR83b, Min86, Pat79b, Pat79a, PS88, Pin88b, RS85, Rah86b, San83a, She88b, Sho76, Tal80, Tef78, Tom81, Tom83, TW79, Tre85, Wei84, Yos83, dW80]. Second-Order [AF92, BC90a, Ban92, Boe95, Dun90a, Dun90b, Dun94, FGHW91, FH90a, FMZ93, FR91, Jef91, Olv77b, Olv77a, Saf99, Van98, Bar79b, BD86, CE89, CLP88, GLY88, Kel79b, Kim86, KS86, Pat79b, Tom81, TW79, Wei84]. Secondary [Buz86, Ruf95, She80a]. sectionally [Azz80, Azz81]. Sections [Kot94, Rus88]. Sector [Tem74]. sediment [AW78]. sedimentation [NNB88]. sedimentation-diffusion [NNB88]. Seismology [Sym91]. Selberg [Hab88, Aom87, Ask80, Hab88, Kad88a, Kad88b, Kan93, Mil94, Rah86a]. selection [DK83]. Selections [Wu90, DK83, Nür80, Som80, Som82]. Self [BK76, Col73, EP97, Fan97, Gal95, Gre91, Hul97, Jaf97b, Mil75, Mit75, St.84, Wei99]. Self-Adjoint [Col73, EP97, Gre91, Mil75, Mit75, St.84]. Self-Inverse [BK76]. Self-Similar [Fan97, Gal95, Hul97, Jaf97b, Wei99]. selfadjoint [But79, CE89, Fuh80, KZ88, vd86]. Semi [Zem82, Kin86, Old83, Zem81]. Semi-Infinite [Zem82, Zem81]. semi-integration [Old83]. semi-inverse [Kin86]. semicircle [Gau89]. Semiclassical [JP95]. Semiconductor [Ala95, Lia94, MM97, Moc74, Smi91b, Zha94, BCG89, Mar85a]. Semiconductors [BD95, Ben96, NP95]. Semicontinuity [Bra95, FM92, FM99]. Semidiscrete [LS89]. Semidiscretization [Kim90]. Semidynamical [FS89]. Semiflows [ST90b, ST91]. Semigroup [BH76, FI90, MLM97b, Sch97, Wil80b, Ben82a, Kur79]. Semigroups [LT92, Pao72, Sta86a, Yao95, vdP83, Ben82b, BHS83, CV87, Sta85]. Semilinear [Aye83, BB92, CD96a, CK89, Den95, DH99, Elc74, ES76, EH90, FSW96, FS96, GV93, Kai91, LM75, LP91, MY98, Mor89, Sho72, TBD^{+}96, Bel87, BB79, Bud89, CL81, CLP89, EV97, Hru85, HW83, JK86, Lev88, Mar78, NS85, NS89, O’D84, Swa82, TW79, Tro87, dM86]. Sensitivity [SZ87, Dol79c, Dol81a]. separability [SW79a]. Separable [KS92]. Separated [Elc82]. Separation [Boy76, BKW85, Hai82, Kal75, Mil74b, Mil74c, Mil75, AL78, Avr88, KM80, KM81, KM83, Rei86]. Separatrix [Sch87c]. Separatrix-Loop [Sch87c]. Sequence [FQ84, Rat74, Smy79, VS79]. Sequences [BL74, BK76, BR81, GM92, Gre89, JT79, Mia93, PD71, SV76, Sim77, SH71, VFO87, Sil81]. Sequency [Zha86]. Sequential [Fri81b, DS84]. Series [ASV84, AJV94, Ask74a, BB83, Ber75a, BG75, Car74b, Col70, CF98, DG92, DL91b, FN75, FK78, Fei78, Gas81, Ger94, GS82, Gla71, GH75, GC95, Gor75b, Gre80, GK84, Gus90, Hen87b, HBL76, HT71, Imm91a, KF74, KSK78, LT81, Leu74, Lew86, Lin74, Lin75b, Lip73, Lo72, McC97, Muc70, Olv92, Osl71, Pil86, Red84, Ros98, Saa74, SN71, Sha76c, Tse97, Vol90, Wal71b, WN81, Wim74, Zay82, ZW82, Zuc79, AKM87, Büh87a, Car79a, Cof86, Cyb84, ES84, GR86, Gus87a, Gus87b, Hol80, JT79, KM88, KS86, Lou84, RS85, Sid86, Ver80, VJ85, Wad80, Zio85, Zuc84]. Set [AW82, Far73, Mul75, Osh93, ST90a, Tre73, AW79, Bre85, DK83, GS81, RV87, Whi80]. set-valued [DK83, GS81]. Sets [All79, And93, CPS98, FT94, GG96, GZ92, GR75, Mat70, She80b, Ver72, CK84, CK85, Hir82, Kui80, MC79, PT78, She88a, She93, WW88b]. Several [Boe95, Daf95, Que71, Yin91, Zac74, DW82, GH87a, GH88, Kor88, Peh90, Smi89, Sym86a]. several-dimensional [Sym86a]. shadowing [CLP89]. Shallow [DW93, KS75, Klo85, Ton81]. Shannon [Doo89]. Shape [CM98, CM00, SZ87, SZZ98, GM84]. Shaped [Col94a]. Shapes [DD91]. Sharp [Kar98, LMW96, Log88, QV96]. Sharp-Cutoff [Log88]. sharpness [ASV81, Rea85, Rea88]. Shaw [ES97b, Gus85]. Shear [BBDU93, BPL91, St.88, Eng87]. Shearing [GN93, NP93, Gri82]. sheets [CO89]. Sheffer [All79, BK76, BR81]. Shell [PP93, PP95]. shells [Wei88]. Shift [Hab76, LJC97]. Shift-Invariant [LJC97]. Shifted [CSW91]. Shifting [Bra76]. Shil′nikov [Bud89]. Shock [Eng99, FS95, GSZ94, How87, KH95, KN98]. Shocks [Key91, Nak88]. Shooting [HZ94]. Short [Car78c, Gri91, Hsu89, HS86a, Sin80]. Short-Time [Hsu89]. SIAM [Ack80a, Ban79, Car86, Cig79, GLY88, Log84a, Mad88, Obi81, Spe79, Tah91, WH88]. Sided [MP77, Str77b]. Sieved [Des98, CI87, Ism85, CI87]. Sign [KV91, Sta86b]. signals [Ben82a]. significance [Obi86, Obi87]. Signorini [HP92]. Similar [Fan97, Gal95, Har72b, Hul97, Jaf97b, Wei99]. Similarity [Lu90a, LKMT88]. Simple [BP90, BH99, Cic87, CD96c, DJJ91b, Dun94, Hra85, JP95, Rio92, Ste77, Vem95, Was70, Was73, CD95, DJJ91a]. Simplification [Hug75, Ohk91, GH86, Sib81]. Simplified [Zha94, Nay84, Nay87]. simply [She88b]. Simultaneous [DJL93, FF92, SW79a, Wal79]. Sine [BFF^{+}90]. Sine-Gordon [BFF^{+}90]. Single [ADH90, FW92, Gal95, Nak88, Swi76, Swi80, Wag83]. Single-Species [FW92]. Singular [Ade73, Ade74, AA83, AS87, AF92, Bax91, BBN93, BT87b, Bra71b, Bus78, CO89, Cam77, CR79, CK89, Cha72, Coh72, CM80, Col70, DDD96, Dor70, EG90, Eck92, Esc73, EW94, FGHW91, FH90a, FTZ88, GS93, GT98, Gio90b, Gom86, Hal77, Har79a, Hol77, Hsi73, Hsi78, HW83, Joy97, KoM98, Kov95, KZ88, Kra89a, Kra89b, KP74, Lag73, LW77, MR83b, MW93, MF71, Nau76, NMIF90, NS98, Nor88a, Par72b, Par72a, Sch90, Sch84b, SK73, Sha87, Sha91b, She72a, Tem74, TY77, Wal71b, Wil72a, Yar75, Zha95b, dHW73, vV85, Ale80, AR87, BJL81, BJL88, BNS85, Bra81a, BCG89, Cam87a, DR86, DK86, Ell82, Esh87, EW89, Fat85, Fen83a, Fen83b, Fen85, Fri81c, Hin81, HS84a, HR88, Kam79, KKZ84, KKZ86, Kel72, Kel79b, Kel89, LS86, Mar86]. singular [MR83c, O’M79, PS88, Pit80, SS80, Sch80, Sch84a, SW86, Sch81, SK87, WZ87, Wil71, ZW85, de 80]. Singularities [Ble77, ElB90, FMZ93, GR76, Man91, McC97, Mil70a, Mil70b, Mil76, Mil79, Nav99, PS98a, Qui93, Ruf95, Sha76a, AB80, BJL81, BJL88, MMN85, Sen97a, Sen97b, WZ87, ZW85]. Singularity [Den95, BD86, GM83, KS86, MR83b, Wei84, dW80]. Singularly [BS79a, Boh90, Fra90, Gar92, How75, How78a, How78b, Jef91, Kel73, MS99, Nii78, O’M70, Ren84, dPFW99, DeS87, GH86, Kel84, Nip86a, Nip86b, NF87, O’D84, O’M79, SW86]. singularly-perturbed [O’M79]. SIR} [Wan79]. Sivashinsky [Mic92, Mic96, Tad86]. Size [DL96, KSS97, KM91c, GH87b]. Skan [Gab84, HT87]. Skeleton [RS91a]. Skew [Gu92]. ski [Bus72]. Skin [Liu95a]. Slightly [HL98, CDHL88]. sloping [EF85]. Sloshing [Bar85b, Kut84]. Slow [EY98, Flo91, Gra95, RS95, Bar84b, Car99a, Fen83a]. Slowly [Cao95, Kot94, WH87a, WH87b, vdS76, Rob83, WH88]. Smale [jCcW99, Li97]. Small [Bra71b, Che95, DD91, Gor75a, HM89a, Hsi73, JKK96, Ład94, MP97, Nag78, O’M71, TF75, Coo88, Day89, FL87, HJ83, LM83, MTH80, Per81, She86]. Smallness [Tre76a, Tre88]. Smooth [BW99, CC96, DGH99, GT98, Ron97, Zha99b, Azz80, Azz81, Daf82, Gri82, Rea85, Rea88, Yom83, Yom86b]. Smoothest [Boj94, Pin88a]. Smoothing [Col94b, DG89, Gre74, Lam78, Mai94, RR95]. smoothly [Sym81]. Smoothness [AC94, Bao98, DL96, Mag87, MF71, Lin86, MrS87]. Sobolev [AMRR92, Coh71b, Coh75, Eir92, ELM^{+}95, Ewi75, FG95, Gri93, Jol92, KL98, KL99, Lag72a, Lag72b, MST96, Pen92, Pie83, Rei97, Sho72]. Soft [BP90]. soils [Kna86]. Solid [AC98, DS95, GN93, STW99, Feh88, Hea88]. Solidification [AC80]. Solitary [KO92, LB96, SS93b, dS97]. Solitary-Wave [LB96]. soliton [Sac84]. Solitons [KV99]. soluble [MO83]. solute [DJ87]. Solution [AHIR93, Av91, BG88, BS74, Bor99, BB97, BC90b, Car78a, D’A75, Dea71, GO71, GR92, GW97, HL96, Hol93, How78a, Ito94, KY75, Kel79a, Kol74a, Lon71, Lu97, Mat70, Mey74, O’M71, OFK77, Olm77, Osh93, RDLH77, Sch87a, Shi93, Utz70, Wyr77, YL94, dGdJ75, dHW73, ka76a, ka76b, ka77, ARRR81, BG89, BSL87, Bre82, CE85, DS82, FIT87, How82, How81, Iff87, JW88, Lar81, MM80, Nor88b, Oht79, SDL84, Ste73a, Sym81, Ton81, Ton82, Wes78]. Solutions [Ada78, Ahl93, AL77, AP93, AM94, And91, Any77, AB72, AH72, Ban92, BH75a, Ban75, BFGJ97, BCP93, BCB82, BBN93, Ben81, Ber70, BY72, BO92, BVB94, Bon96, BH99, Bre93, BT91b, BR71, CS97a, CN77, Cao95, CPW98, Car95b, CC94, CP97, Che92b, Chi95, CL90, Chu76, CGP93, Coh72, CK87, CK92, Col75, CW99, CL71, CL73, Coo74, CD96c, Cus71, Cus75, DGG98, DL96, Dek71, Dem96, Den95, DL92b, DH99, Dun90b, Dun94, Ede79, Eir92, Eli72, ES97b, EH90, Fan97, FGHW91, FMZ93, FS96, FW92, FR91, FHX96, Gal91, GHL96, Gar87, Gar92, GGI92, GT94b, GM96, Gil70, GL71, GT98, GL74, GH75, Gor75a, GB72, Gra72, GP93, Gri87, Gri91, Gru92, GW96, Had74, HL98, Hal77, HK90, Har72b, Har77, HLM92, HL94, Hay91]. Solutions [HO94, HH72, Hil98, Hof96, HS96, HW92, Hul97, Hun73, IW98, IM91, Imm90, Ish96b, JM96, KV91, Ko98, KN97, KK92, Kee72, Kem85, Kho77, KO92, KS79a, Kim83, KV99, KF90, KS85a, KS91, Kuw96, Ład94, Lae75a, Lae75b, Lag72a, Lak90, LR77, LW95, Leu73, Leu75, LZ95, Lev73, Lev74b, LP76, Lew86, LB96, Lim78, LP91, Lin94, LB93, Liu95b, Liu97, LMN98, Lo72, Lon74a, Lon87, Lu90a, Lu90b, Lu99, Lun78, MR79a, MM97, MM77a, Mer76, Mey77, Mic92, Mil76, Mil70e, Mil71a, MF71, MP97, MY98, MH95, MW74, Nar93, Nii97a, Nii97b, Nii78, NMIF90, NP95, Okr91, Ols74, Par70a, Par72b, PT97, Pes95, PJ73, Pog99, PT99, Prü88, Rao75, Rei70, RE98, Ren90, Ren91, Ric90, RR95, Ros71, Sai96, Sch84b, Sha87, Sig74, Sin75, Ski94]. Solutions [SZZ98, Ste80b, Stu96, Suz95, TBD^{+}96, Tal81, TF75, Ter90, Tho77, Tol98, TY77, Tro90, Tro93, Vil92, Wal76, Wal75a, Wan98b, WW98, Wol93, Wys93, XA91, Yan96, You99, Zac74, Zha99a, ZZ90, Zha95a, Zha95b, vdS76, Akı88, ABM87, Ali95, Ame82, ÁP88a, Art75, AK82, BCT88, Bel87, BNS85, BB86, Blo80, BD86, Bra79, Bri88, BB79, CO89, Cam87b, Can85, CD80, Cha79, CL81, CB89, CL86, CL89, Cic87, CN81, Cof82, Cof86, CE87, Cos87, Cur79, Daf82, Dam87, DL91a, Dav79, Die86, DK86, EH84, Erb76, Far78, Fit82, FL87, FR86b, GH88, GT89, GK80, Gri81a, Gri82, Gri89, Hag82, Har70b, Har70a, HLM86, HT87, HV87, HM89b, HSX87, Jer78, JK86, JSW87b, Kel72, Kel88, Kem82a, Kha87, Kim87, Kin86, KM86, Klo85]. solutions [Kor88, KS80b, KS86, Lei85, LC84, LM80, Lev83, Lev88, Lin86, LS87, LKMT88, LK89, LP81, Mar82, Mar89b, McC78, Mem89, Mil79, MTH80, Moe79, Mos83a, NT86a, NT86b, NT86c, Nam78, NS85, Nip86b, Nis82, NF87, NS89, Oja78, Olv93, Pao80, Pao83, PW82, Pit80, Rab82, Rea87, RS86, Rul87, Sac83, San83a, Sel81, She85, SP79, Smi86b, Sta78b, Ste80a, Ste84, Swa82, Swi80, Sym86b, Tal80, Ter88, Tro87, Tsu84, Tur89, Vas79a, Vol84, Whi79, Wie87, WL82a, WL82b, WM87, Yos83, Yos84, Yos86, dT78, vV84]. Solvability [Dol79d, JMR96, KT99, MW93, Sch76c, CDMS81, Oht79]. Solvable [KR76, Cam87a]. Solving [Ser75, Vas79b, Vas79a, Vas81]. Some [AP93, AFT98, AP74, Ask80, Bar87, Bax91, BVB94, BMP92, BS76, BR84, Cam87b, Car70b, Car73a, Car77b, Car77c, Car78b, Car79a, Car70c, CW89, Chu76, CGR88, DDD96, Dor70, Dut85, EP98, Elc75, Fei79a, FTZ88, FMZ93, For90, Gau74b, GM93, GG77, Gla84, GC95, GR75, GR78, GP93, Gus92, Gus94, HG99, Har72b, Har79a, HS98a, Hei75, HP98, Hol87, Id94, IT79, Jai81, Jam98, Jol92, Kar98, KTV94, Kui76, LS71, LMW96, LMPS90, Let94, Lev76, Lev74b, Lir71b, LB93, Lon73, Lon88, LMS73, Lor93, Mac82, MR79a, ML84, Mil70e, Mon80, MZ92, MS84, Mur83, Nus78, Nus90, O’R93, Pao74, Par70a, Par70b, Pat71, Pay70b, Pay74, PS76, Pip93, QVV98, Rah76c, Rei79, Ros78, Sch73a, SS77, Sch73b, SS92, Son70, Stu96, Tre79, Val94]. Some [VFO87, VJ81, VJ85, Wan78, Wan79, Wei74, Wil80c, Wim85, WB95, Wys93, Zuc84, dPFW99, ALS87, ARRR81, Ban79, BB86, Cuy83b, DeS87, FI75, GR86, GR89, Gri81a, HM80, Har70b, Har70a, HT89, Kad88b, Kaz84, LM80, Lon83, Nis82, OP87, PP79, Per88, Sch87a, Sid86, Spe79, Wie87]. Sommerfeld [Hab76, Spe89, SHM89]. Source [Die95, Gal91, GM93, Vis86]. Sources [Dol77]. Space [AK99, Bar75, Bar77a, BGG84, Car95b, CP97, Coh75, Col74, Daf95, Dat72, Dat81, EP99, GL74, HL77, Han74a, Han74b, Han76b, Kal75, KM78, Kry96, KL98, KL99, Lag73, Lev73, LM75, Lon77, MM97, Paz72, Pen92, Sho72, Wil80b, Yin91, You75, de 78, Akı88, Bel81a, BB78, CH79, CH82, GK84, Gri78, JS87, JSW87a, JSW87b, KS80a, Lou84, Mar80, Nam78, Pou89, Ter85, Wag83]. Space-Time [KM78]. Spaced [CR92]. Spaces [ANS96, AF92, BU97, BHL75, Chi77, Cho84, Coh71b, DRF95, FK94, Gri93, Hir72, Ho74, Jol92, Kol74a, LA96, Lee74, Lee75, LJC97, NW74, Pie83, Pil86, Rei97, Sei91, Swa72, Tho77, ki72, vdP83, BG89, Bar79a, BKW85, BHS83, But79, Eng82, Gut82, Kaz84, Log85, Mar78, May79, McC83, Mei79, Oht79, Som82, Wex80, vd86]. Span [MY98]. Sparse [Alp93]. Sparsity [Don92]. Spatial [APS93, CQ97, Lu99, MS95, PT97, Qui96, RB98, Lin86]. Spatially [Ale90]. SPDE [KL98]. SPDEs [KL99]. Special [BBL97, BÇ94, IK79, Mil70f, Mil70e, Mil71b, Sha91b, Ter85, AM83, Bra86, PV84, Piq86, Piq89, SS80, Sha80a, Si78]. Specializations [SS94]. Species [FW92, MH93, Swi76, Yan92, Swi80]. Spectra [Amo99, Edw93, Fox72, Bea85, Chi80, HS86a, Ura84]. Spectral [AC98, CD95, CD96c, Dom88, Kap79b, LL93, LPR92, MR91, Yam71, Ben91, GN83, Kol79, Vic78, van86]. spectrale [CD95]. Spectrum [BW80, GH79, KM95, Lop89, MZ98, Sta79, Stu89]. Speed [BCT98, BMY77, CS97a, Can85, Ko95a, Dav79]. Spence [HT89]. Sperb [Spe79]. Sphere [Gla76, Joh75, Joh77, KMT91, SS84]. Spheres [Llo80, Van89, Van96, Xu98].

Spherical [BM92, Dav73, Dun78, Esc73, Fry91, KS75, Koe94, Koo93, Lip72, Lip73, Sac74, SS92, And88, Ben91, CL89, Coc87, GN89, IR89, Mon80, PGRSV89, Vre84]. Spherically [Lin97, Col78]. Spheroidal [KB73, Dun86, Per86, Per88]. Spigler [EL86b]. Spike [Do 93]. Spiral [Sch98]. Splay [Mir94]. Splay-Phase [Mir94]. Spline [AEU94, ANS96, Che95, CSW91, MP77, YZ94, dB76, CS80, GJ82, LM82]. Splines [Bea81, CSW80, De 77b, Hu91b, Lah93, NSSS84, Rei97, ST76, PS82, Som80]. Splits [PFK95]. Splitting [BW80]. Spoiler [Elc82]. Spread [FH90b]. Spreading [Ren97]. Square [BT96, Bou72, EL84, GG77, HK90, DL84, GM84]. Square-integrable [BT96]. squared [Sac83]. Squares [Log90, QSW93, Mil70c]. Stabilities [Her91, LC84]. Stability [Ada78, ABV96, AV90, Aus77, AL98, Bar79b, Bar81, Boe95, Bra91, CS99, CJ99, Chi77, Chu76, Daf95, Dat72, Eng99, Fer86, FHM97, Flo91, Gar92, GS86, GSZ94, Gri89, GM79, HRX98, Hir90, Hol93, Hut86, IM91, KSS97, KY75, Kel73, KTV94, KKL99, Kuw96, Lag72a, LS71, LLS97, LA96, Lev88, LM96, Li97, Li99, LYC95, Lin97, LY95, LWY98, Lou94, Mao97, ML93, Mic96, MM92, MM77b, MM79, Nak78, Nii97a, Nii97b, NF87, NMIF90, NP93, Ort94, Pag86, Pao72, Par92, Pey75, PS94, Qi96, RB71, RTV96, San98, ST90b, SM85, Ter90, Van98, Ven78, Ves97, Wan80, Wan98a, WY75, Win74, Yam90, Zho96, Zin91, CS79, Dav78a, ET89, Gar81, Har70b, Har70a, Hru85, KS79b, Lun87, Mad85a, Mad88, MS86, Mor87, Nip86a, Nip86b]. stability [Pao83, Per82, Sym86a, Veg88, Wei85, Wex79, Wex80]. Stabilization [CJM98, Gil79, BH86, ML84]. Stable [AR98, BM92, BV98, Ehl73, FK94, FH85, Gri91, Han90, HP93, JM96, MN75, NS98, Sym81, Cic87, FS79, JSW87b, Pou89]. Stably [BBDU93]. Stacked [RTV96].

Stäckel [BKM86]. Stäckel-equivalent [BKM86]. Standard [Kre94]. Standing [KM86]. Starlike [CS84, Jia88, Tep75]. Stars [Ise85c, Lin97, Ise85a, Ise85b]. State [BT91b, Gar92, HS98b, Lak90, Lu90b, MR98a, SM84, Sha76a, ET89, GK80, Kel88, MR79b]. States [Ale90, Bro83, FM98, How86, JKV99, Kel73, KS75, LVLH92, Lu87, Mik88, WZ97, Yam90, BH87, HKS91, Hsu83, Wei85]. Static [EW94]. stationarity [DS87]. Stationary [AHIR93, Bak72, BH75c, CPW98, Dau89, HB73, Hol93, IM91, JKV99, JT94, Ko98, KR81, KS85b, MMM97, MW91, MA88, NP95, Olv74, WW98, de 95, dK71, ACK83, BH87, MS79, Mia87]. stationnaires [BH87]. Statistical [Lam80, MR98a, HM80, TWW89]. Steady [Ale90, AP89, BT91b, Gar92, Hsu83, Lak90, Lu87, Lu90b, Mic92, SM84, Sha76a, Sun93, Yam90, Zha99b, GK80, How87, Kel88]. Steady-State [BT91b, Lak90, Lu90b, GK80]. Steepest [GB72]. Stefan [Bor99, CP73, CD80, Cha70, Dam81, Feh88, FRZ95, FH96, FH99, GZ95, GM89, KT99, Ric90, Rul87, Ryb99, SAW87, Xie90, Xu92]. Stefan-Like [Xu92]. Stekloff [AM94, Lip98]. Step [Avr91, Avr93, BSL87, Ter88]. Stieltjes [BD78, BR78a, BB83, Bro75b, IK79, Rap82, Rus77, Sim77, SV94, Wil86, ki72]. Stiff [KT97, Ngu90]. Stimulated [MS92]. stirred [GK80]. Stochastic [DT96, Kry96, Mao97, MR98b, Bar87, BS79a, GM79, JS87, Mei79, MS79, Rah81b, SDL84, Tud87]. stochastically [Day82]. Stokes [Che92a, APS93, BGT88, Car96, Che92a, Che92b, Dau89, GT94b, Hof96, Imm90, Imm91a, Imm91b, JT94, Kel88, Kni72, Kra91, KK97, McC81, Sha73, Shi74, SN99, SS95]. Stommel [BCT88]. Stopped [CG83b]. Strain [How86, NR95]. Strain-Hardening [NR95]. Strange [GS82]. Stratification [CD96c, CD95]. Stratified [BBDU93, Col78, Sym81]. Stream [BCT88]. Stress [AMS^{+}92, AA83]. stress-assisted [AA83]. Stress-Dependent [AMS^{+}92]. Stretching [Rei91, MR83a]. Strict [KI95]. Strictly [Dol81b, Fei79b, Mie87, TKS93]. String [Car95a, BS83a, She85, She88a]. Strings [LSZ92, AR87]. Strip [CD96c, EY98, CD95]. Strip-like [EY98]. Strips [GO71]. Strong [And96, ÁP88a, Ben85, Bri88, Chi77, GP99, KS80b, LY95, Sri90, WW99, LS88a, Swe89]. stronger [Sto80]. Strongly [GM91, Jef91, Pér93, ST90b, ST91, ÁP88b, Cos84, ET89, Fei79b, HR88]. Structural [Hir90, Li97, Li99]. Structure [CZ94, JW80a, Kai92, Ko98, KM95, MW78, MZ98, Yan96, ZZ90, Cof82, Lai80, Nis82, Wag89]. Structured [BIT91, IMP92, Bra83, HM89a, WHG88]. Structures [AC98]. Struve [HR94, Ste70a]. Student [IK76]. Study [AK99, BH87, BF88, BF89b, KS79c, Sen85, Fri81c, GK80, Nas84, Rab87]. Sturm [AL78, AL81, BS75, Cam87b, Car99b, EM94, Gau90, Gre89, KKZ84, KKZ86, KZ88, Rap82, RS86, Wal71b, WZ87, ZW85, dT78]. Sturmian [Rei73]. Subclass [PS72]. Subdifferential [BHU96]. Subdivided [KS79c]. Subdivision [CD96b, Gri96, HJ98, Rio92, YS93]. Subfunctions [CS77a]. Subharmonic [FMZ93, Van90, McC83]. Subject [Liu95a, Tre76a, Tre84]. subjected [vd86]. Sublinear [Zha95b]. Suboptimization [FF92]. Subspaces [JSW87b, Lam80]. Substitution [Gor75b]. Successive [Pao74]. Suction [Lu97]. Sufficient [jCcW99, LM95, Olv78, GN83, Peh84b, Rab84, Sin79]. suggested [Lef82, Sto80]. Sum [Ask74b, BI89, Log90, Kad87, Old83]. Summability [Gor75b, Lip73, SS85, Obi86, Obi87, Rap82, Sid86]. Summary [Dur78, RR84]. Summation [Gas81, Gla71, Gus90, Hol80, Lin74, Lin75b, SN71, Zuc79, Gus87a, VJ85]. Summations [And76, Lin78]. summing [Rap82]. Sums [Bus77, Car77b, Gas77, HL74, Kra72, LB74, Lir71b, Old70, Old72, Rah76c, Rus78, AM83, BG84, CR88, DJ87, LMN87, Rah81b]. Superadditive [TWW89]. Superconductivity [DLY98, GP99, HT98]. superconductor [CFV81]. Superconductors [HS98b]. Supercooling [GZ95]. Supercritical [Tro90]. Superlinear [PR91, RS86]. Superquadratic [Die88]. Superresolution [Don92, GG96]. Supersonic [Zha99b]. Support [BP90, Che95, Han99, Lui82b]. Supported [BAR90, CD93b, Dau93, Vol95, Zha92b]. supremum [Yom83, Yom86b]. Surface [AHIR93, EMS98, HP92, Lip98, SS93b, Tur96, Zaj94, CL87, Coo88, DR86, Gv85, MS86, Pag86, TT83]. Surfaces [AJV94, Kre94]. Surfactants [Ren97]. surquadratiques [Die88]. survey [Rab82]. suspended [AW78]. Sustained [Rob83]. Swallowtail [Kam92]. swelling [FMP86]. Swift [Kuw96]. swirl [Tur89]. Symbol [FTZ88]. symbols [AW79]. Symmetric [ACG99, BT91b, CG94, CD91, HQW74, Hug75, Lin97, Neu88, Sap71, Wei74, Auc89a, ML93, Pey75, She88b, Ter85, Ism85]. Symmetrical [Dur77]. Symmetries [KM85, Mil73b, dS97]. Symmetrization [DR74, KJ82]. Symmetrized [Sle72]. Symmetry [Ast91, BM92, Ben86, CSv87, CG88, CZ91, FM98, FR91, GS86, PS98b, Pos90, Do 93, AKM87, Bri88, CL89, Dri87, FR86b, Hea88, IR89]. Symmetry-Breaking [Ast91]. Synthesis [Por80]. synthetic [Rab87]. System [ABH94, AM80, AL98, BL90, BBL97, BK90, Bil90, BFF^{+}90, BHN72, Bro75a, CK91, Che92c, Chi95, CvL92, Din89, EH90, Fai77, Flo91, FRZ95, GHL96, Gao93, Gar92, Gar90, Gra72, Hai92, HZ94, HL94, Hol93, HS96, Jef91, JKV99, KoM98, KN98, Kem85, Lak90, LYC95, Lin95, Lu87, Lu90b, MW75, McC97, MZ98, Mil71a, MN75, Noh74, NP95, Rah91, Ren97, Sch73a, SS93a, STW99, Shi74, SZZ98, Tak92, TKS93, TW91, TW97, WK92, WW99, Wil73, Che92a, Che81, Gar88, GH86, GH88, HSX87, Hsu83, JW88, KM86, LC88, Pao80, Pao87, Rob89, Rob83, Sas91, Swe89, Zei87, Zei88, dM86, vV84]. system-Dyson [Zei87]. système [Che92a]. Systems [AL77, Aom94, Aus77, Bar83, BCP93, BSX95, Ber76, Ber72, Boe95, Bra71b, Bro83, CJ99, Cam77, CR79, CPW98, Car95b, CL90, Chi73, CD97, Cus75, Cus82, Daf95, Dau89, Dv96, EM90, ET71, EW73, Fan93, FG99, FHM97, FMV98, FS89, FH90b, FH96, GP84, Gra95, Gri91, Gru92, HW89, Hal91, Har94b, Har72a, Har74, Heg81, HS98a, Hei78, HP93, Hir82, Hir85, Hir90, HM84, IM91, Jia96, JPS87, JP95, Joy97, Kho77, Kov95, Kra89a, Kra89b, KT97, Lab99, LL93, LW79, Mac82, MR83c, MW93, May70, Mei76b, Mey84, MP77, Mie87, MM77b, MR93, Mor89, Mor90, MH95, MNY94, Nak91, NMIF90, Ols74, Pas74, PS97, Pog99, Por80, RR95, RTV96, Rot93a, Sai96, Sch98, Sch97, Smi91a, SS96, St.84, TBD^{+}96, TN94, Van90]. Systems [Ven78, Vis90, Wal79, WO91, XA91, Yok97, Yu77, ZZ90, AH88, AL78, Bar81, Ben82a, BG96, BB86, BKM86, BE83, BE86, Cam87a, CS86, CDMS81, CE89, Cho76, Cos84, Day82, DG88, DH84, Dol79c, Fuh80, Ful83, Gar81, GS89, Gri81b, Gru85, GM79, HS84a, HKS91, HMP87, How87, How81, JAK88, JW80a, Kel79b, Kel84, Kim86, Kui80, Mar89a, ML84, MR79b, NT86a, NT86b, NT86c, Nip86a, Nip86b, Nis82, NF87, Nür80, O’D84, Per81, Per82, Pey75, Rab82, RA79, RR84, RRW88, Rei79, SY89, Sel81, Smi84, Smi86b, SM85, Ste85b, Str80, Tal80, Tef78, Tom81, Tom83, Tom86, Vas81, Wei84, WM87, dW80]. Szego [DG88, JNW94, Jus74, Peh90, Rus88].

Tail [GS99, HL74]. Tails [Log88]. Tallest [CM98, CM00]. tank [GK80]. tau [OP87]. Tauberian [AP92, Lut78, RS74, Saa74]. Taylor [GS86, Osl71, Ral83]. Technique [Lun78, PV84, Rob85, How81, Piq86, Piq89]. Techniques [Yar75, AKM87, Kha87]. Tellegen [Dol77]. Temperature [BPL91, Gla76, Yin98, Coc87, LC84, Pao83]. temperature-dependent [LC84]. Temperatures [Hai80]. Tempered [Sri72]. Temple [BG96]. Temporally [GM76]. temps [Ali95]. Tending [Fei78]. Tension [AHIR93, SS93b]. tensors [KM80, KM81, KM83]. Term [And96, For90, FSW96, GM93, Han90, Ito94, Jam98, TW97, BH87, BF88, BF89b, CW89, GLY84, GLY88, Vis86]. terme [BH87]. Terminal [LT92]. Terms [BR78b, Tem90b, Tre73, Wim74, GR85, JAK88, KS86, Son80]. Test [KTV94]. Testing [Dol70]. Tests [Wil72a, Wil71]. th [Bog74, Bro75b, GB72, Gra72, Hen89, Leu74, LMN87, Pet84, Sin75, Wyr77, Gau90]. th-order [Hen89]. Their [AVV90, BLN87, Dav73, Fin74, GM91, Gio90b, GW94, Lee74, Mei76a, MNY94, Zha95a, Ban79, BNS85, BC79, EL87, Ism94, JAK88, Mad85a, Mad88, ML84, PP79, Sin79]. them [BR84]. Theme [Dau93]. Theorem [AL81, Atk74, Ban73, Ban74, Ber76, Car77c, Car71c, Dol77, Dun78, Elc71, Fri81a, Gio93, Gus90, Häh98, Hen87b, Ho74, Hof99, KB73, Kug74, Kut84, KW83, Lev74a, LR72, Li85, Liu96, Mat78, McC97, Mei72, MN75, Mul87, NR95, Nus78, Pan71, Paz72, PR91, PD71, Ren91, Rus78, Saa74, San83b, Sch76c, SW75, Šim88, St.84, Sta71, Tip74, Tur96, Vem95, Yam71, Zha94, ASV81, Aro81, Cam89a, CDHL88, CR88, Col82, Dol80, Doo89, Gab84, Gau90, GS81, Han89, HM89a, HKS91, Jan84, Keg83, Kri80, Lew79, Lut78, May79, Mha88, ND81, Rah79, RRW88, RS78, Rus88, Sib81, Vol87, CK77]. Theorems [Ala95, AP92, Car78c, CD74, CJ76, Det78, Erb77, Fie84, GM93, Jol92, Lon73, Mao97, Mil75, Mul78a, Nus90, OFK77, RS74, Sch73b, Sho72, Str76, Wim74, Wys93, AGX89, BE83, CPS87, DiG81, Eli84, Gar81, GL84, GLY84, GLY88, Gus87a, Hen89, Hir83, Hol80, JW80b, Kim86, Pat79b, Tom83, Tom86, Ise85b]. Theoretic [Ast91, Gil70, Koo82]. Theoretical [BÇ94, Coh71a, FMP86]. Theories [Ren91]. Theory [ADH90, AN98b, AT85, Avr91, BBDU93, BB92, Boy76, Bro75b, BH76, CRD88, CH93, CK95b, DG83, Dav78b, DW93, Erh73, FI90, For90, FI72, Fra90, Ger94, GRZ83, GIO90a, Gio93, GM76, Gug71, GPG91, HQW74, HP85, HLM92, HNO98, Imm91b, Ish96a, Kai92, KM78, Kel73, KN72, Kov95, Kra89a, Kra89b, Kry96, KL98, KL99, Lab99, McB75, Mei76b, MP77, Mil70a, Mil70d, Mil70f, Mil70e, Mil72, Mil73a, Mil74a, Mil74b, Mil74c, Moc74, NB75, Nay77, Par72b, Sat77, Sat90, Šim85, SS96, Vas79b, Was74, AL78, ABM87, AA83, AW78, BG84, Blo80, Bos79, Bud89, Chi87, Cuy83a, Cuy84, DJ87, DeS87, Fen83a, Fen83b, Fen85, Gar87, GS81, GLR79, GM83, Gri79a, HP83, HW81, Hiv78, Ise85a, Ise85b, Kap79a, Kap79b, Kel72, KS80b, Mar82]. theory [MR91, MRS81, Mer87, Ngu89, Rei79, SS80, SDL84, SP79, Ton81, van86]. There [HT98]. Thermal [FR94]. Thermistor [AC94, XA91, YL94]. Thermoelastic [AMS^{+}92, AL98, Kim92, RR95]. Thermoelasticity [EW94, Fra83, Riv93, Wil80a]. Thermostat [Gri87, Gri91, Gri89]. thermostats [GS81]. thermoviscoelasticity [Daf82]. Theta [For90, BG84]. Thin [AN98a, BC95, Bre93, DGR99, DDD96, PP93, PP95, Buz86, DV87, Hom79, Ren82]. Thin-Film [DDD96]. Third [Amo99, BW73, BP96, ES75, Jon76, RB71, SHM89, Tro93, Utz70, Erb76, Hen87a, Nay87, O’R87]. Third-Order [Amo99, BP96, Tro93]. Thomas [Lun78]. Three [Av91, BBN93, Car78c, Coc87, Dau89, GK87, GL71, Han90, Hir90, Kal75, Kim90, KI95, Kli97, LL93, Mil76, PS91a, Sac74, Sun93, Ves97, AGX89, DiG81, GN89, Mil79, Moe84]. Three-Body [GK87, Moe84]. Three-Dimensional [Av91, BBN93, Dau89, Hir90, Kal75, Kim90, KI95, Kli97, Sac74, Sun93, Ves97, Coc87]. three-point [AGX89, DiG81]. Three-Term [Han90]. Threshold [Gao93, Nus78]. Thresholds [GK87]. Tikhonov [Mai94]. Time [AV90, APS93, AN98a, BS95, BL90, BB92, Boe95, CN77, Car96, Coo74, Dam87, DS87, EZ97, Ewi75, FH90b, FD97, Gla76, HL98, HL96, Heg81, Hsu89, Kai91, KM78, LY96, LS96, Lu99, Mei76a, Mor85, NS91, Pey71, RT93, Ray81, Rei93, Say71, Tak92, Ven78, Vil92, Wei82, Zem79b, Ali95, Bar84b, Bel81a, Bel81b, Cam87a, CB89, Cos84, CP84, Cyb84, Dav78a, Day89, FL87, MM80, ML93, Mel85, NT86c, NNB88, Pao87, RR84, San80, Sei87, Zem79a, NT86b]. time- [Mel85]. time-delay [RR84]. Time-Dependent [AV90, APS93, CN77]. Time-Frequency [RT93]. Time-Harmonic [AN98a, HL96, NS91, Rei93]. Time-Invariant [Mor85]. Time-Like [BS95]. Time-Periodic [LS96, Tak92, NNB88]. time-reversible [ML93]. time-scales [San80]. Time-Varying [Mei76a, Zem79b, Zem79a]. Times [Kee77]. Timescales [EG90]. Timing [Kee77, Kol74b]. Titchmarsh [Wal75b]. Toda [Chu84]. Toeplitz [Bax71, BS96, DG88, Gu92, Jus74]. tokamak [BGS86]. Tomographic [Ler90]. Tomography [Qui93]. Toner [HW92]. Topological [MST96, O’R87, Geo87, Kha87]. Tori [Mir94, SM84, WO91]. torsion [PP83]. Total [AHP75, CG83a, Yok97]. Totally [Kem82b]. Trace [BS95, Car99b, Jol92, Pap95, BG89].

Tracking [JKK96]. tract [Lan83a]. Traffic [HR95]. Trajectories [Sin76]. trajectory [Gun73]. Transcendental [Het70, Ste85b, SP79]. transfer [Gv85, Mer87]. Transform [Car74a, Car74d, CLH92, Col94b, Dea79, FI72, Fri95, Gla84, Kat97, KN72, Lon74b, Nav99, Nay77, Pan72, PW70, Pri75, Qui93, RS74, Sch73b, Son70, Ste72, Ste73b, Str77b, Ung74a, Ung74b, Wil79, Won80, Yao95, Car80, FS83, Log83a, Log83e, Lou84, Lut78, MN86, Mar84, Nay89, Val86]. Transformable [Koh70b, Pan71]. Transformation [AHP76, DG92, HR94, Hug75, Koh70a, Lee75, Que71, Ben84b, Gas85, GR86, GS89, Gus87b, How81, Kap79b, Lam80, Miz83, Ver80, Lau75]. Transformations [ASV84, Car76b, Jai81, Lee74, Sal89, Str77a, VS79, VJ81, KKZ84]. Transformed [IN90]. Transforms [BHL72, Ble77, Bow71, BS76, CL96, Dea79, DdS74, Fre87b, HL70, HB73, HP73, JK75, Lew74, Mad85b, Olv91a, Ort73, Per74, RC73, SS84, SS73a, Won76, Yue73, de 78, vdW75, IK79, Log79b, Log85, MQ85, SW79b, Sid85, Son80, Wil86]. Transient [MS92]. Transition [Mac90, Ohk91, CFV81, Sib81, WG87]. Transitional [IMP90]. Transitions [AS87, CC99, Ish96a, JKV99, SZZ98]. Transitive [Rob92]. Translates [BAR90]. Transmission [IMP92, Rei93, RZ91, RS91b, Tor96, WW88b]. Transplantation [Sch73b]. Transport [Bal99, CLP88, SLS81, AW78, Hiv78, KS80b, Lar81, Pao83, Sen85, vd87]. transversality [O’R87]. Transversely [CS77c]. Traveling [ARRR81, BNS85, Bon96, Den91b, Flo91, FHX96, Hag82, Has88, MX98, Nii97a, NMIF90, Ric90, RTV96, Sai96, Ter88, Zin91]. Traveling-Wave [Bon96, Nii97a, Sai96]. Travelling [BFGJ97, DL92b, EH90, Ko95a, MH93, MR93, Nii97b, AA88]. Treatment [GK77, SV94, HJ83]. Triangle [Pin80, Pin85]. Triangular [Car70a]. Triangularization [Gin79, GT98]. Tricomi [GW94]. Tridiagonal [FG99, Smi91a, DG88, Smi84]. Trigonometric [BH93, BI89, Coh71b, FN75, Rat74, ST76, Sri72, Tse97]. Triple [Hen87b]. Trotter [HM89a]. Truncated [BFF^{+}90]. Truncation [Thr94]. Tubes [LA91, CR83]. Tubular [Ale90]. Turán [BI83, MRS81]. Turánians [Lor80]. Turbulence [Hul97]. Turbulent [BBDU93, Id94]. Turcotte [HT89]. Turing [NS98]. Turning [Any77, Dor70, Dun90b, Dun94, GB72, Gra72, How75, How78b, KP74, Leu73, Leu75, Nii78, O’M70, Olv77b, Olv77a, PM82, Ski94, Was70, Was74, BD86, DeS87, Gin85, GH87a, GH88, Kam79, MR83c, de 80]. Turning-Point [PM82, Was70]. Turrittin [Bra71a, Bra72]. Twelve [Tse97]. Twist [ET94]. Twisted [Den91a, San98]. Two [All92, Amo99, BR78a, BSX95, BBL97, BP96, BÇ94, Bor99, BLM96, CDT90, CvL92, DL91a, DL92a, Dau89, Dou97, Dut81, ET71, Fai77, Fet78, Fre92, FSV99, GHL96, Gas73, GO71, GT94b, GL71, GHLS91, GH79, GZ95, GB72, Gra72, Gre91, GW96, Guc84, HS97, Har74, HL94, Hof99, Kal75, KN97, KK92, Kee77, KS96, Kol74b, KSK78, KO95b, Kra89b, Kuw96, Lak90, LA91, LW95, Liu91, Maa96, MW75, MW91, MR98a, Mil70a, Mil70b, Mit75, Nau76, Ost96, Par72a, Pes95, RJ86, RS91a, Sak93, Sch74, SR93, SZL97, Shu78, Sle95, SK76, Str77b, SS93b, Tip74, Van89, Van96, Vil92, WW99, WB95, Yan92, Yok97, You75, ZZ90, AK88, Ali95, BN78, Bar84a, Bar84b, BSL87, Chi83, Coh89, DiG81, Dil88, Fet77, Fre87a]. two [HIN85, Hal84, HS84a, HP89, HSX87, JS87, Koo82, Kus88, LC84, Lin86, LK89, Mil84, MTH80, Mos83b, SS80, Sch80, Sch81, SAW87, Ter88, Wag83, WM82]. Two- [Dau89, Kal75, DiG81]. Two-Component [RJ86]. Two-Dimensional [BSX95, CvL92, GHL96, GHLS91, HL94, Hof99, KN97, Kuw96, LA91, Liu91, Maa96, MW75, MW91, MR98a, Mil70a, Mil70b, Pes95, SR93, WW99, ZZ90, FSV99, Ali95, Mos83b, WM82]. two-disk [Bar84a]. Two-Extremum [RS91a]. two-group [LC84]. Two-Layer [SS93b]. Two-Parameter [Fai77, Hal84, JS87]. Two-Phase [Bor99, BLM96, GT94b, GZ95, GW96, SAW87]. Two-Point [ET71, Gre91, HS97, Har74, Par72a, SZL97, Tip74, WB95, BN78, LK89, MTH80, Sch81]. Two-Prey [Lak90]. Two-Scale [All92, LW95, Vil92, DL91a, DL92a]. Two-Sided [Str77b]. Two-Species [Yan92]. two-step [BSL87, Ter88]. Two-Timing [Kee77, Kol74b]. Two-Variable [BR78a]. Type [AMRR92, Aom94, Ask74b, BF98, BCP93, BB97, BC90b, CJ99, CL91, CS93a, Cla83, Dem96, Dör87, Dör91, Dur78, EG90, EKLOO92, Fan93, HP73, HL94, HO94, HRX98, Ism74, Joh80, Key91, Kho77, Lag72a, Lag72b, Lee74, MMM97, MR95, Mao97, MX91, Muk72, Nak78, OT94, RS94a, Rei73, RR95, Rob92, SW75, Sho75, SS96, Ver72, Wal75b, dGdJ75, vV85, BEGv87, Cho76, Chu80, Die86, Dol80, ELS94, Iff87, Kap79a, Log85, LS86, LS79, Mar83a, MO83, Mha88, MTH80, MK83, Mir83, RRW88, Spi79, Ste73a, de 80, HS98b]. Type-II [HS98b]. Types [Fet78, Jon81, Fet77, GR84].

Ultimate [Ber72]. ultra [Mon80]. ultra-spherical [Mon80]. ultracentrifuge [Yos84, Yos86]. Ultraspherical [LS94, AMS86, Bre81, BS79b, BS80, BI83, Dun74, GR83, Gas85, RV86]. Umbral [RT94]. Unbounded [FSW96, HFL99, Jan86, LT92, Leu73, Leu74, Leu75, Smi75, Was70, Bry86, CV87, Kap79b, Lan78, Tud87]. Uncertainty [Ben84a, Bow71, Per86]. Unconditional [Gri93]. Unconstrained [Auc89a]. undergoing [Coo88]. Unfolding [HJ92, DS84]. Unidirectional [HR95, SS93a]. unification [Koo82]. Unified [SV94, Zei88]. Uniform [And93, Any77, AF92, BT91a, Ber72, BD86, Bro83, BG75, Dat72, Dun86, Dun89, Dun90b, Dun94, Fie73, Fie83, FH90a, FW88, Fre90, Han75, Han77, HW84, Kel89, KI95, Li85, Muk72, Nor88b, Ohk91, OT94, Olv91b, Olv91a, SS75, Tem90b, Sib81, Son83, Tem82, Tem87]. Uniformly [Ski80b, Ski81, Ski88]. Unilateral [SZ87]. Unique [CRSW76, KS96, WL82b, Zha92a, RS78]. Uniquenes [Lev73]. Uniqueness [Ala95, AC94, Bro96, BV92, Che81, CD74, CJ93, Cus71, Dix96, EI97, Fri95, GT89, Häh98, Ho74, Kra74, MLM97a, MM77a, NR95, Sta71, Str76, Tip74, TY77, TW91, Wad80, YL94, AGX89, CL86, ET89, Fen85, Han89, Keg83, Kri80, Lar88, LC88, NNB88, ND81, Pet81, San83a, Ton81, Ton82, Wei88]. Unit [Pie74, Wol71, Büh87b]. Unitary [Son70]. Unity [SV83]. Univalence [Pfa86]. univalent [GR79, Rus88]. Universal [AB93b]. unsettled [RA79]. Unstable [GS99, Tur96, JSW87b]. unsteady [Lar88]. Unusual [Eva96b, BEGv87]. Upper [Gal91, McC83]. Use [Bus72, Mim88, Sim80, Vol84]. used [Gas89, Gri79a]. Useful [Bus78, FI72, Gla84]. Using [Av91, Ast91, DGHM96, Gou92, Yue73, Sil81]. utility [Rob89].

V [Vol87, CI87, Joh80]. Valid

[Jaf97a, Kol74b, Ski80b, Ski81, Ski88]. Validity [BB94, Kee77, TBD^{+}96]. valuation [SDL84]. Value [BG88, Ban92, Ban73, Bar77b, BG76, Bax91, Boh90, Bus78, Car76a, Car99a, Chi74b, Col73, Dea71, Dix96, Dun93, EP98, Fin77, FGHW91, FI96, Gau74b, Gil70, GK77, Gor95, Gre91, Har74, Har79a, HLM92, HP98, Hol77, How75, How78a, How78b, Joh75, Joh77, Joh80, KL96, KO95b, Kui76, LB74, Lag72b, LW77, MW75, MW93, Mil76, Moc74, MW74, Mul78a, Mul87, Nag78, Nii78, NT96, NP95, O’M70, O’M71, Pao74, Par72a, PS76, Ral83, Rei93, Ren90, Ren96, SS77, SS93a, SZL97, Smi75, Sta71, Szw92a, Szw92b, Tip74, TY77, Whi80, YL94, Zha95b, dM89, AGX89, Ale80, AW78, BG89, BN78, Ben86, BB79, Cha72, Cof86, Col82, Daf82, DeS87, DK86, EH84, Gab84, HLM86, Har78, Hen87a, Hen89, HS84a, HP89, How82, Kap79a, KKZ84, Kim87]. value [LS88a, LS89, Lef82, LS87, LK89, LP81, Mar82, Mar83b, Mil79, MTH80, Mul81, O’D84, O’M79, O’R87, Pet81, Pet84, SW86, Sch81, Ton81, Ton82, Wei71, Wei84, Wie87, dW80]. Valued [AP92, BS96, Dou98, Str77b, ki72, DK83, Gin78, GS81, Koo85, Sta78a]. Values [Ban73, Bao98, Car73b, FG92, FTZ88, MNY94, WN81, Zay82, Zha95a, AK85, CG83a, CR83, Sym86b, Yom86a]. Vanishing [Ban92, Sin76]. Variable [BR78a, BEHM89, CJM98, Col74, GG77, IM91, Ser75, Smi91b, Vas79b, CS80, FH85, KM80, KM81, KM83, MM80, Rah81a, Tch96, Vas79a, Vas81]. Variables [Ber75b, Boy76, Col72, Hai82, HL74, Kal75, KSK78, LL93, Mil76, Mil74b, Mil74c, Mit75, Que71, Sch74, Shu78, SK76, Yok97, You75, BKW85, Dil88, DW82, Kow82a, Kow82b, Mil79, Mil84, Tch96]. Variance [Lam78]. Variate [PR71]. Variation [Bai71, CH92, Hor75, Ven78, vdW75, Ste85b, Tal80]. Variational [Auc89b, BB77, Bin82, BY91, Col94a, DLY98, Fis75, FM98, Fri81b, Hu90, JKO98, LVLH92, Ngu90, Rei75, Auc89a, BB78, Gus85, HS86b, Kur79, LS88a, Mad85a, Mad88, Rog88, TW89]. Variations [Dau93, Sta86b, Har80]. Varieties [Mas78]. Varying [Kot94, Mei76a, Ven78, WH87a, WH87b, Zem79b, vdS76, Cam87a, Rob83, WH88, Zem79a]. Vector [AP92, Bra81b, Dou98, Pér93, Riv93, ki72, Jia88, Koo85, ML93, WW88b]. Vector-Valued [AP92, ki72, Koo85]. Vectorial [Zag98]. Vectors [She98, Wan98a]. Velocity [HFL99, Lou94, Sim90, Sym91, SAW87]. Venant [Kin86]. verification [FS79]. Version [Ryb99, Doo89, Eng82, Jay88]. versus [CP84]. Vertical [BLS99, Eli72]. Very [Gus90]. Very-Well-Poised [Gus90]. VI [AA88]. Via [Col94b, Don92, LT94, Nav99, SV94, SS95, ADH90, BB92, Cha78, Dea79, GM83, How78a, Mar86, Old83, RT93, Sal89, Wex80]. Vibrating [CZ94, BS83a, Hol87, SP79]. Vibration [Bar85b]. vibrations [Coo88]. Viewed [Joh80]. Visco [GN93]. Visco-Plastic [GN93]. Viscoelastic [KH95, Kim95, STW99, CT88, Eng87, Gri82]. Viscoelasticity [Ada78, FI90, HR88, Kim83]. Viscoplasticity [Car94]. Viscosity [BPL91, Eng99, SZZ98, Daf87, LS87]. Viscous [Eng99, FS95, GSZ94, KT99, LY95, SS75, Sim90, Sle95, Wan98b, Zaj94, Ghi86, HS89, Kel88, Kim87, Sec88, Ton82, Wag89]. vocal [Lan83a]. Volterra [AP92, Art75, AD86, Bar75, Bar77a, Bar79a, CH82, CN79, CN81, DG89, Fie86, GK84, Gri78, Gri79a, Gri80b, Gri80a, Gri81a, Gri81b, Gri85, GM79, Han73, Han74a, Han74b, Han75, Han76a, HW82, Hea82, Hir86, Hru85, JW74, JW80a, JW80b, Ko98, Kap72, Kif77, KS79a, KS80a, Kif80, Lev76, Lon74a, Lon87, MR83a, Mil71a, MF71, MN75, MM79, Nor88a, Nor88b, Okr91, Prü88, Red85, SV94, Sta78a, Sta80a, Sta88, Ste80a, TW79, Yin91, dHW73]. Volume [BS97]. Volume-Preserving [BS97]. Vortex [IW98, LN98, NT96, Tur89, CO89]. Vortices [CS99, CJM98, LD97, RS95, WY92]. Vorticity [HFL99, LN98]. Vries [BO92, BW83, CK87, GT89, Hay91, Tsu89, Zha92a, Zha95a].

W [Spe79, HL77]. Walk [Des98]. Walks [CG83b]. Wall [Her91, VK91]. Walsh [Red71, Wad80, Yue73]. Water [DW93, Sun93, Klo85, Sac84]. waters [Ton81]. Watson [And76, BS76, DdS74, GR84, GM94, Mar86, Sid85]. Wave [AHZ93, BS95, Bao98, BFGJ97, BO92, Bon96, Det78, DL92b, EH90, Fan93, FS96, GM91, GHLS91, IW98, Joh77, KM78, KO92, Kim98, KB73, Lev74a, Lev74b, LM75, LB96, MP83, Nii97a, Nii97b, NMIF90, Ric90, Sai96, Smi91b, Str76, Ter90, Ali95, ÁP88b, ÁP88a, ARRR81, BNS85, Bra81a, Chi74a, Col82, CK84, CK85, DM82, Hag82, KM86, Mar89b, Mor87, Per86, Per88, She85, Sym86a, Ter88]. Wavefronts [Zin91]. Waveguide [Boe78]. Waveguides [ACG99, Urb96]. Wavelet [CL93, CL96, CS93b, CD93a, DDMT93, MP98, Nav99, She95, Vil94]. Wavelets [AEU94, BW99, BT96, Che95, CS93a, CD93b, DS99, Dau93, DGHM96, DGH96, DGH99, Gou92, Gri93, Han99, Hua96, Kai92, Lan96, Maa96, Swe98, Vil92, Vol95, Wal92a, GM84]. Waves [CWZ99, CP90, Den91b, Dv96, FS95, FHX96, GS93, GSZ94, GS99, Has88, IMP90, Ko95a, KN98, Kim98, LB96, LMN98, MH93, MR93, MX98, RTV96, Sch98, SS93a, SS93b, Sun93, Tor96, Tur96, dS97, AA88, Bar84b, Col78, CKP89, Lei81, Sac84, Wag89, Zio85]. Weak [AP93, BBN93, CN77, GGI92, GT94b, Kim87, KP92, LT99, MP77, Nus90, Ros71, Rul87, Shi93, Van98, Whi79, CD80, Jer78, MR79b, Nür80, Som82]. weak-solutions [CD80]. Weakly [AB72, Kol79, MF71, dHW73, EK84, Mor87, Pit80]. Webster [Sym86b]. Wedge [Zha99b]. Wedges [CF96]. Weierstrass [Lin75a, Que71]. Weight [LL90, Log79c, MK78]. Weighted [AH83, CLH92, Jan81, JW80b, LT94, Rei97, Van85, Van01]. Weights [CC91, Coh71b, LM95, NEM94, EMN94, LS93]. Weil [LB94]. Weinberger [FS77]. Weinstein [Akı88, FS77]. Well [FR94, Gus90, HBL76, Sch73a, Sho70, Tad86]. Well-Poised [HBL76]. Well-Posed [Sho70]. Well-Posedness [FR94, Tad86]. Weyl [CP84, RT93, Rea85, Rea88]. which [CR83, Cof82, Col70, Hir82, Kat97, Pil86, RV87, Wad80]. Whipple [And76, Gus87b]. whirling [AR87]. Whittaker [D’A75, Dun89, Har77]. Whole [Kry96]. Whose [Ask74a, BHL72, Coo74, Dea71, EG91, Hal77, Lew74, Ser75, Wim74, Chi80]. Wick [Ler90]. Widder [Aro81]. width [BH85, Hoc84]. widths [Mel85]. Wiener [Ben91, JS87, MR98b, Pan96, Ste73b]. Wigner [Col94b, Jan81, Jan84, Ler90, Ste91]. Wijngaarden [Lau75]. Wilson [DJJ91a, Chi87, DJJ91b, Gus94, Ism86, Koe94, Koo93, Mil87, NR85, Rah81b, Rah86b]. Wirtinger [Fei79a, Swa78]. Wirtinger-like [Fei79a]. Wishart [Ric89]. Without [BY91, CZ91, Rei93, AL98, NT96, Sin97]. WKB [CC96]. Wronskians [Lor80].

x [FS83, Mar84, Qui93]. x-ray

[FS83, Mar84, Qui93].

Young [Dem96, FM99, KP92]. Yukawa [Bil90].

Zemanian [Pil86]. Zero [Coh75, Col70, EM94, Fei78, FM98, GW94, Ols74, SV76, STW99, Tha93, Van85, Van01, Ver72, Car79a, ES84, Log83b]. zero-balanced [ES84]. Zero-Free [SV76]. zero-one [Car79a]. Zeros [AMRR92, BFW86, BL84, BCP93, Bus77, CJ76, Dav73, EL84, EL85, EG91, HPR97, Het70, IN90, IM78, LS88c, Lor93, LS94, McC78, McC77, MS84, PS91b, ST90a, Sen96, Ste70a, WW92, Xu94, YH82, AM83, AMS86, Cam87b, Chi87, EL86a, EL87, Fia80, FR86a, Gat87, Gau89, Ism94, LM83, LM77, Log83d, Lor80, MNT86, ND79]. Ziglin [CR91]. Zonal [Koo93, Kus88]. Zone [AV90, LK89]. Zones [Hol93, Hal84].

References

Alexiades:1983:SPT

[AA83] Vasilios Alexiades and Elias C.

Aifantis. Singular problems in the theory of stress-assisted diffusion. SIAM Journal on Mathematical Analysis, 14(5):925–933, September 1983. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alexander:1988:BAR

[AA88] James C. Alexander and Giles Auchmuty. Bifurcation analysis of reaction-diffusion equations. VI. Multiply periodic travelling waves. SIAM Journal on Mathematical Analysis, 19(1):100–109, January 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aziz:1972:PSC

[AB72] A. K. Aziz and S. L. Brodsky.

Periodic solutions of a class of weakly nonlinear hyperbolic partial differential equations. SIAM Journal on Mathematical Analysis, 3(2):300–313, May 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Armstrong:1980:AEI

[AB80] Judith A. Armstrong and Norman Bleistein. Asymptotic expansions of integrals with oscillatory kernels and logarithmic singularities. SIAM Journal on Mathematical Analysis, 11 (2):300–307, March 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ashbaugh:1993:MBE

[AB93a] Mark S. Ashbaugh and Rafael D. Benguria. More bounds on eigenvalue ratios for Dirichlet Laplacians in N dimensions. SIAM Journal on Mathematical Analysis, 24(6):1622–1651, November 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ashbaugh:1993:UBL

[AB93b] Mark S. Ashbaugh and Rafael D. Benguria. Universal bounds for the low eigenvalues of Neumann Laplacians in N dimensions. SIAM Journal on Mathematical Analysis, 24(3):557–570, May 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Akuezue:1994:QAD

[ABH94] H. C. Akuezue, R. L. Baker, and M. W. Hirsch. The qualitative analysis of a dynamical system modeling the formation of multilayer scales on pure metals. SIAM Journal on Mathematical Analysis, 25 (4):1167–1175, July 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/19853.

Al-Bassam:1987:LTS

[ABM87] M. A. Al-Bassam and H. L. Manocha. Lie theory of solutions of certain differintegral equations. SIAM Journal on Mathematical Analysis, 18(4):1087– 1100, July 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alessandrini:1996:DLC

[ABV96] Giovanni Alessandrini, Elena Beretta, and Sergio Vessella. Determining linear cracks by boundary measurements: Lipschitz stability. SIAM Journal on Mathematical Analysis, 27 (2):361–375, March 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/26579.

Alexiades:1980:FBP

[AC80] Vasilios Alexiades and John R.

Cannon. Free boundary problems in solidification of alloys. SIAM Journal on Mathematical Analysis, 11(2):254–264, March 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Antontsev:1994:TPE

[AC94] S. N. Antontsev and M. Chipot.

The thermistor problem: Existence, smoothness, uniqueness, blowup. SIAM Journal on Mathematical Analysis, 25(4): 1128–1156, July 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23348.

Allaire:1998:BLH

[AC98] Grégoire Allaire and Carlos Conca. Boundary layers in the homogenization of a spectral problem in fluid-solid structures. SIAM Journal on Mathematical Analysis, 29(2):343–379, March 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/30432.

Arcoya:1999:AMS

[ACG99] David Arcoya, Silvia Cingolani, and José L. Gámez. Asymmetric modes in symmetric nonlinear optical waveguides. SIAM Journal on Mathematical Analysis, 30(6):1391–1400, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/33638.

Acker:1977:HFI

[Ack77] Andrew Acker. Heat flow inequalities with applications to heat flow optimization problems. SIAM Journal on Mathematical Analysis, 8(4):604–618, August 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Acker:1978:FBO

[Ack78] Andrew Acker. A free boundary optimization problem. SIAM Journal on Mathematical Analysis, 9(6):1179–1191, December 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [Ack80a].

Acker:1980:EFB

[Ack80a] Andrew Acker. Errata: “A free boundary optimization problem” [SIAM J. Math. Anal. 9 (1978), no. 6, 1179–1191; MR 80d:49031]. SIAM Journal on Mathematical Analysis, 11(1):209, January 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [Ack78].

Acker:1980:FBO

[Ack80b] Andrew Acker. A free boundary optimization problem. II. SIAM Journal on Mathematical Analysis, 11(1):201–209, January 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Acker:1981:EPI

[Ack81] Andrew Acker. An extremal problem involving current flow through distributed resistance. SIAM Journal on Mathematical Analysis, 12(2):169–172, March 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aronson:1983:HIS

[ACK83] D. G. Aronson, L. A. Caffarelli, and S. Kamin. How an initially stationary interface begins to move in porous medium flow. SIAM Journal on Mathematical Analysis, 14(4):639–658, July 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Attouch:1986:HVE

[AD86] Hédy Attouch and Alain Damlamian. Homogenization for a Volterra equation. SIAM Journal on Mathematical Analysis, 17(6): 1421–1433, November 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Adali:1978:EAS

[Ada78]  Sarp Adali.  Existence and

asymptotic stability of solutions of an abstract integrodifferential equation with applications to viscoelasticity. SIAM Journal on Mathematical Analysis, 9 (1):185–206, February 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Adelson:1973:SPI

[Ade73] L. E. Adelson. Singular perturbation of an improperly posed problem. SIAM Journal on Mathematical Analysis, 4(2):344–366, May 1973. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Adelson:1974:SPI

[Ade74] L. E. Adelson. Singular perturbation of an improperly posed Cauchy problem. SIAM Journal on Mathematical Analysis, 5(3): 417–424, May 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Arbogast:1990:DDP

[ADH90] Todd Arbogast, Jim Douglas, Jr., and Ulrich Hornung. Derivation of the double porosity model of single phase flow via homogenization theory. SIAM Journal on Mathematical Analysis, 21(4): 823–836, July 1990. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aldroubi:1994:DSF

[AEU94] Akram Aldroubi, Murray Eden, and Michael Unser. Discrete spline filters for multiresolutions and wavelets of l_{2}. SIAM Journal on Mathematical Analysis, 25(5): 1412–1432, September 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23408.

Arnold:1990:BLR

[AF90] Douglas N. Arnold and Richard S.

Falk. The boundary layer for the Reissner–Mindlin plate model. SIAM Journal on Mathematical Analysis, 21(2):281–312, March 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ashyralyev:1992:UDS

[AF92] A. Ashyralyev and H. O. Fattorini. On uniform difference schemes for second-order singular perturbation problems in Banach spaces. SIAM Journal on Mathematical Analysis, 23 (1):29–54, January 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Arnold:1996:AAB

[AF96] Douglas N. Arnold and Richard S.

Falk. Asymptotic analysis of the boundary layer for the Reissner–Mindlin plate model. SIAM Journal on Mathematical Analysis, 27(2):486–514, March 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/24527.

Alvino:1998:PSN

[AFT98] Angelo Alvino, Vincenzo Ferone, and Guido Trombetti. On the properties of some nonlinear eigenvalues. SIAM Journal on Mathematical Analysis, 29 (2):437–451, March 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30211.

Angad-Gaur:1994:EHK

[AGGO94] Hendrik W. K. Angad-Gaur, Bernard Gaveau, and Masami Okada. Explicit heat kernel on generalized cones. SIAM Journal on Mathematical Analysis, 25(6): 1562–1576, November 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23062.

Andrews:1994:D

[AGIN94] George Andrews, George Gasper, Mourad Ismail, and Paul Nevai. Dedication. SIAM Journal on Mathematical Analysis, 25(2): vii–ix, March 1994. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Adolfsson:1992:LGE

[AGJL92] V. Adolfsson, M. Goldberg, B. Jawerth, and H. Lennerstad. Localized Galerkin estimates for boundary integral equations on Lipschitz domains. SIAM Journal on Mathematical Analysis, 23 (5):1356–1374, September 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Azor:1982:CAH

[AGV82] Ruth Azor, J. Gillis, and J. D.

Victor. Combinatorial applications of Hermite polynomials. SIAM Journal on Mathematical Analysis, 13(5):879–890, September 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aftabizadeh:1989:EUT

[AGX89] A. R. Aftabizadeh, Chaitan P.

Gupta, and Jian-Ming Xu. Existence and uniqueness theorems for three-point boundary value problems. SIAM Journal on Mathematical Analysis, 20(3): 716–726, May 1989. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aziz:1972:PSH

[AH72] A. K. Aziz and M. G. Horak. Periodic solutions of hyperbolic partial differential equations in the large. SIAM Journal on Mathematical Analysis, 3 (1):176–182, February 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Andersen:1983:WNI

[AH83] K. F. Andersen and H. P. Heinig. Weighted norm inequalities for certain integral operators. SIAM Journal on Mathematical Analysis, 14(4):834–844, July 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ahlbrandt:1988:RMD

[AH88] Calvin D. Ahlbrandt and John W. Hooker. Riccati matrix difference equations and disconjugacy of discrete linear systems. SIAM Journal on Mathematical Analysis, 19(5):1183–1197, September 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Abergel:1993:DGP

[AHIR93] F. Abergel, D. Hilhorst, and F. Issard-Roch. On a dissolutiongrowth problem with surface tension in the neighborhood of a stationary solution. SIAM Journal on Mathematical Analysis, 24 (2):299–316, March 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ahlbrandt:1993:CFR

[Ahl93] Calvin Ahlbrandt. Continued fraction representations of maximal and minimal solutions of a discrete matrix Riccati equation. SIAM Journal on Mathematical Analysis, 24(6):1597– 1621, November 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

A-Hameed:1975:TPP

[AHP75] M. S. A-Hameed and F. Proschan.

Total positivity properties of generating functions. SIAM Journal on Mathematical Analysis, 6(1):81–84, February 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Abdel-Hameed:1976:PGP

[AHP76] M. Abdel-Hameed and F. Proschan.

Preservation of geometric properties under an integral transformation. SIAM Journal on Mathematical Analysis, 7(3):359–365, May 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Amirat:1993:HDW

[AHZ93] Youcef Amirat, Kamel Hamdache, and Abdelhamid Ziani. Homogenization of degenerate wave equations with periodic coefficients. SIAM Journal on Mathematical Analysis, 24(5): 1226–1253, September 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Andrews:1994:HAG

[AJV94] G. E. Andrews, D. M. Jackson, and T. I. Visentin. A hypergeometric analysis of the genus series for a class of 2-cell embeddings in orientable surfaces. SIAM Journal on Mathematical Analysis, 25 (2):243–255, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/22954.

Azzam:1982:SEE

[AK82] A. Azzam and E. Kreyszig. On solutions of elliptic equations satisfying mixed boundary conditions. SIAM Journal on Mathematical Analysis, 13(2):254–262, March 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Angell:1985:HEB

[AK85] T. S. Angell and R. E. Kleinman. The Helmholtz equation with L_{2} boundary values. SIAM Journal on Mathematical Analysis, 16 (2):259–278, March 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Adamchik:1988:DIP

[AK88] V. S. Adamchik and K. S. Kölbig. A definite integral of a product of two polylogarithms. SIAM Journal on Mathematical Analysis, 19 (4):926–938, July 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aubert:1999:MSR

[AK99] G. Aubert and P. Kornprobst. A mathematical study of the relaxed optical flow problem in the space BV (Ω). SIAM Journal on Mathematical Analysis, 30(6):1282–1308, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/34123.

Akin:1988:IRP

[Akı88] Ömer Akın. The integral representation of the positive solutions of the generalized Weinstein equation on a quarterspace. SIAM Journal on Mathematical Analysis, 19(6):1348– 1354, November 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Agarwal:1987:CES

[AKM87] A. K. Agarwal, E. G. Kalnins, and Willard Miller, Jr. Canonical equations and symmetry techniques for q-series. SIAM Journal on Mathematical Analysis, 18(6): 1519–1538, November 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ahmad:1977:CES

[AL77] Shair Ahmad and Alan C. Lazer.

On the components of extremal solutions of second order systems. SIAM Journal on Mathematical Analysis, 8(1):16–23, February 1977. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ahmad:1978:DES

[AL78] Shair Ahmad and A. C. Lazer.

An N-dimensional extension of the Sturm separation and comparison theory to a class of nonselfadjoint systems. SIAM Journal on Mathematical Analysis, 9 (6):1137–1150, December 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ahmad:1981:ESC

[AL81] Shair Ahmad and Alan C. Lazer. On an extension of the Sturm comparison theorem. SIAM Journal on Mathematical Analysis, 12(1):1–9, January 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Avalos:1998:EST

[AL98] George Avalos and Irena Lasiecka. Exponential stability of a thermoelastic system with free boundary conditions without mechanical dissipation. SIAM Journal on Mathematical Analysis, 29 (1):155–182, January 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30082.

See the Editorial Note regarding this paper.

Alabau:1995:NUT

[Ala95] Fatiha Alabau. New uniqueness theorems for the one-dimensional drift- diffusion semiconductor device equations. SIAM Journal on Mathematical Analysis, 26 (3):715–737, May 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/18823.

Alexiades:1980:SPI

[Ale80] Vasilios Alexiades. A singular parabolic initial-boundary value problem in a noncylindrical domain. SIAM Journal on Mathematical Analysis, 11(2):348–357, March 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alexander:1990:SOS

[Ale90] Roger Alexander. Spatially oscillatory steady states of tubular chemical reactors. SIAM Journal on Mathematical Analysis, 21 (1):137–153, January 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alinhac:1995:APD

[Ali95] Serge Alinhac. Approximation près du temps d’explosion des solutions d’équations d’onde quasilinéaires en dimension deux. (French) [Approximation near the blow-up time of solutions of two-dimensional quasilinear wave equations]. SIAM Journal on Mathematical Analysis, 26(3):529–565, May 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24454.

Allaway:1979:ESP

[All79] Wm. R. Allaway. Extensions of Sheffer polynomial sets. SIAM Journal on Mathematical Analysis, 10(1):38–48, January 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Allaire:1992:HTS

[All92] Grégoire Allaire. Homogenization and two-scale convergence. SIAM Journal on Mathematical Analysis, 23(6):1482–1518, November 1992. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alpert:1993:CBS

[Alp93] Bradley K. Alpert. A class of bases in L^{2} for the sparse representation of integral operators. SIAM Journal on Mathematical Analysis, 24(1):246–262, January 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Arino:1987:OSR

[ALS87] O. Arino, G. Ladas, and Y. G. Sficas. On oscillations of some retarded differential equations. SIAM Journal on Mathematical Analysis, 18(1):64–73, January 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Arbib:1980:GHM

[AM80] Michael A. Arbib and Ernest G. Manes. Generalized Hankel matrices and system realization. SIAM Journal on Mathematical Analysis, 11(3):405–424, May 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ahmed:1983:RPS

[AM83] S. Ahmed and Martin E. Muldoon. Reciprocal power sums of differences of zeros of special functions. SIAM Journal on Mathematical Analysis, 14 (2):372–382, March 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alessandrini:1994:EED

[AM94] G. Alessandrini and R. Magnanini. Elliptic equations in divergence form, geometric critical points of solutions, and Stekloff eigenfunctions. SIAM Journal on Mathematical Analysis, 25(5): 1259–1268, September 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24908.

Amirat:1995:AOD

[AM95] Youcef Amirat and Mohand Moussaoui. Analysis of a onedimensional model for compressible miscible displacement in porous media. SIAM Journal on Mathematical Analysis, 26 (3):659–674, May 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23297.

Ames:1982:CSR

[Ame82] Karen A. Ames. On the comparison of solutions of related properly and improperly posed Cauchy problems for first order operator equations. SIAM Journal on Mathematical Analysis, 13 (4):594–606, July 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Amos:1989:RID

[Amo89] Donald E. Amos. Repeated integrals and derivatives of K Bessel functions. SIAM Journal on Mathematical Analysis, 20 (1):169–175, January 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Amour:1999:DTO

[Amo99] L. Amour. Determination of a third-order operator from two of its spectra. SIAM Journal on Mathematical Analysis, 30(5):1010–1028, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/32963.

Alfaro:1992:OPS

[AMRR92] M. Alfaro, F. Marcellan, M. L. Rezola, and A. Ronveaux. On orthogonal polynomials of Sobolev type: Algebraic properties and zeros. SIAM Journal on Mathematical Analysis, 23(3):737–757, May 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ahmed:1986:INB

[AMS86] Shafique Ahmed, Martin E. Muldoon, and Renato Spigler. Inequalities and numerical bounds for zeros of ultraspherical polynomials. SIAM Journal on Mathematical Analysis, 17(4):1000– 1007, July 1986. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Andrews:1992:ODT

[AMS^{+}92] Kevin T. Andrews, Andro Mikelić, Peter Shi, Meir Shillor, and Steve Wright. Onedimensional thermoelastic contact with a stress-dependent radiation condition. SIAM Journal on Mathematical Analysis, 23(6): 1393–1416, November 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ammari:1998:THE

[AN98a] H. Ammari and J. C. Nédélec.

Time-harmonic electromagnetic fields in thin chiral curved layers. SIAM Journal on Mathematical Analysis, 29(2):395–423, March 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/30550.

Arkeryd:1998:CEP

[AN98b] L. Arkeryd and A. Nouri.

A Condensation–Evaporation problem in kinetic theory. SIAM Journal on Mathematical Analysis, 29(1):30–47, January 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/30173.

Andrews:1976:AWW

[And76] George E. Andrews. On qanalogues of the Watson and Whipple summations. SIAM Journal on Mathematical Analysis, 7(3):332–336, May 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Andrews:1980:NDC

[And80] George E. Andrews. Notes on the Dyson conjecture. SIAM Journal on Mathematical Analysis, 11(5):

787–792, September 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Andersson:1988:DFS

[And88] Lars-Erik Andersson. On the determination of a function from spherical averages. SIAM Journal on Mathematical Analysis, 19 (1):214–232, January 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Andreucci:1991:ELS

[And91]  Daniele Andreucci.  L_{loc}^{∞}-

estimates for local solutions of degenerate parabolic equations. SIAM Journal on Mathematical Analysis, 22(1):138–145, January 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Andrievskii:1993:UHA

[And93] Vladimir Andrievskii. Uniform harmonic approximation on compact sets in R^{k}, k ≥ 3. SIAM Journal on Mathematical Analysis, 24(1):216–222, January 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Andreasson:1996:RGT

[And96] Håkan Andréasson. Regularity of the gain term and strong L_{1} convergence to equilibrium for the relativistic Boltzmann equation. SIAM Journal on Mathematical Analysis, 27(5):1386– 1405, September 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Anonymous:1994:A

[Ano94] Anonymous. Acknowlegment.

SIAM Journal on Mathematical Analysis, 25(4):vi, July 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Anonymous:1998:EN

[Ano98]  Anonymous.  Editorial note.

SIAM Journal on Mathematical Analysis, 29(3):v, May 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97406.

Alfeld:1996:DLB

[ANS96] Peter Alfeld, Marian Neamtu, and Larry L. Schumaker. Dimension and local bases of homogeneous spline spaces. SIAM Journal on Mathematical Analysis, 27(5):1482–1501, September 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/27627.

Anyanwu:1977:UAS

[Any77] Donatus U. Anyanwu. Uniform asymptotic solutions of nonhomogeneous differential equations with turning points. SIAM Journal on Mathematical Analysis, 8 (4):710–718, August 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alduncin:1981:AGA

[AO81] G. Alduncin and J. T. Oden.

Analysis of Galerkin approximations of a class of pseudomonotone diffusion problems. SIAM Journal on Mathematical Analysis, 12(6):917–930, November 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aomoto:1987:JPA

[Aom87] K. Aomoto. Jacobi polynomials associated with Selberg integrals. SIAM Journal on Mathematical Analysis, 18(2):545–549, March 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aomoto:1994:CCD

[Aom94] Kazuhiko Aomoto. On connection coefficients for q-difference systems of A-type Jackson integrals. SIAM Journal on Mathematical Analysis, 25(2): 256–273, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/22746.

Askey:1974:SAM

[AP74] Richard Askey and Harry Pollard. Some absolutely monotonic and completely monotonic functions. SIAM Journal on Mathematical Analysis, 5(1):58– 63, February 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ang:1988:SSQ

[ÁP88a] Dáng Dình Áng and A. Pham Ngoc Dinh. Strong solutions of a quasilinear wave equation with nonlinear damping. SIAM Journal on Mathematical Analysis, 19 (2):337–347, March 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ang:1988:SDW

[ÁP88b] Dáng Dình Áng and Alain Pham Ngoc Dinh. On the strongly damped wave equation: u_{tt} − Δu−Δu_{t}+f(u) = 0. SIAM Journal on Mathematical Analysis, 19 (6):1409–1418, November 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ames:1989:DES

[AP89] K. A. Ames and L. E. Payne.

Decay estimates in steady pipe flow. SIAM Journal on Mathematical Analysis, 20(4):789–815, July 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Arendt:1992:VVT

[AP92] Wolfgang Arendt and Jan Prüss.

Vector-valued Tauberian theorems and asymptotic behavior of linear Volterra equations. SIAM Journal on Mathematical Analysis, 23(2):412–448, March 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alaa:1993:WSS

[AP93] Nour Eddine Alaa and Michel Pierre. Weak solutions of some quasilinear elliptic equations with data measures. SIAM Journal on Mathematical Analysis, 24(1):23–35, January 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ames:1993:SDE

[APS93] K. A. Ames, L. E. Payne, and P. W. Schaefer. Spatial decay estimates in time-dependent Stokes flow. SIAM Journal on Mathematical Analysis, 24(6): 1395–1413, November 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Antman:1987:DWS

[AR87] Stuart S. Antman and Michael Reeken. The drawing and whirling of strings: singular global multiparameter bifurcation problems. SIAM Journal on Mathematical Analysis, 18 (2):337–365, March 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alessandrini:1998:SDC

[AR98] Giovanni Alessandrini and Luca Rondi. Stable determination of a crack in a planar inhomogeneous conductor. SIAM Journal on Mathematical Analysis, 30(2):326–340, 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/32550.

Arkeryd:1990:EEL

[Ark90] Leif Arkeryd. On the Enskog equation with large initial data. SIAM Journal on Mathematical Analysis, 21(3):631–646, May 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aronson:1981:WIT

[Aro81] D. G. Aronson. Widder’s inversion theorem and the initial distribution problems. SIAM Journal on Mathematical Analysis, 12 (4):639–651, July 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Atkinson:1981:TWS

[ARRR81] C. Atkinson, G. E. H. Reuter, and C. J. Ridler-Rowe. Traveling wave solution for some nonlinear diffusion equations. SIAM Journal on Mathematical Analysis, 12(6):880–892, November 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Artstein:1975:CDS

[Art75] Zvi Artstein. Continuous dependence of solutions of Volterra integral equations. SIAM Journal on Mathematical Analysis, 6(3): 446–456, May 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aziz:1979:FMP

[AS79] A. K. Aziz and M. Schneider. Frankl–Morawetz problem in R^{3}. SIAM Journal on Mathematical Analysis, 10(5):913–921, September 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alikakos:1987:SLC

[AS87] Nicholas D. Alikakos and K. C. Shaing. On the singular limit for a class of problems modelling phase transitions. SIAM Journal on Mathematical Analysis, 18(5): 1453–1462, September 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Al-Salam:1972:ACC

[ASC72] W. A. Al-Salam and T. S. Chihara. Another characterization of the classical orthogonal polynomials. SIAM Journal on Mathematical Analysis, 3(1):65– 70, February 1972. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Al-Salam:1976:COP

[ASC76] W. A. Al-Salam and T. S. Chihara. Convolutions of orthonormal polynomials. SIAM Journal on Mathematical Analysis, 7 (1):16–28, February 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Al-Salam:1987:PPC

[ASC87] W. A. Al-Salam and T. S. Chihara. q-Pollaczek polynomials and a conjecture of Andrews and Askey. SIAM Journal on Mathematical Analysis, 18(1):228–242, January 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Askey:1971:OEP

[Ask71]  Richard Askey.  Orthogonal

expansions with positive coefficients. II. SIAM Journal on Mathematical Analysis, 2(3): 340–346, August 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Askey:1974:CRF

[Ask74a] Richard Askey. Certain rational functions whose power series have positive coefficients. II. SIAM Journal on Mathematical Analysis, 5(1):53–57, February 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Askey:1974:JPN

[Ask74b] Richard Askey. Jacobi polynomials, I. new proofs of Koornwinder’s Laplace type integral representation and Bateman’s bilinear sum. SIAM Journal on Mathematical Analysis, 5(1): 119–124, February 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Askey:1980:SBH

[Ask80] Richard Askey. Some basic hypergeometric extensions of integrals of Selberg and Andrews. SIAM Journal on Mathematical Analysis, 11(6):938–951, November 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Askey:1982:BIA

[Ask82] Richard Askey. A q-beta integral associated with BC_{1}. SIAM Journal on Mathematical Analysis, 13(6):1008–1010, November 1982. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Aston:1991:ACS

[Ast91] P. J. Aston. Analysis and computation of symmetry-breaking bifurcation and scaling laws using group theoretic methods. SIAM Journal on Mathematical Analysis, 22(1):181–212, January 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Anderson:1981:EEK

[ASV81] N. Anderson, E. B. Saff, and R. S. Varga. An extension of the Eneström–Kakeya theorem and its sharpness. SIAM Journal on Mathematical Analysis, 12 (1):10–22, January 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Al-Salam:1984:QTB

[ASV84] W. A. Al-Salam and A. Verma. On quadratic transformations of basic series. SIAM Journal on Mathematical Analysis, 15 (2):414–421, March 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Auslander:1985:RAF

[AT85] L. Auslander and R. Tolimieri. Radar ambiguity functions and group theory. SIAM Journal on Mathematical Analysis, 16(3): 577–601, May 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Avram:1989:HIF

[AT89] Florin Avram and Murad S.

Taqqu. Hölder’s inequality for functions of linearly dependent arguments. SIAM Journal on Mathematical Analysis, 20(6): 1484–1489, November 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Atkinson:1974:ETA

[Atk74] Kendall E. Atkinson. An existence theorem for Abel integral equations. SIAM Journal on Mathematical Analysis, 5(5): 729–736, October 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Auchmuty:1989:UVP

[Auc89a] Giles Auchmuty. Unconstrained variational principles for eigenvalues of real symmetric matrices. SIAM Journal on Mathematical Analysis, 20(5):1186–1207, September 1989. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Auchmuty:1989:VPE

[Auc89b] Giles Auchmuty. Variational principles for eigenvalues of compact operators. SIAM Journal on Mathematical Analysis, 20(6): 1321–1335, November 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Auslander:1977:FSD

[Aus77] Joseph Auslander. Filter stability in dynamical systems. SIAM Journal on Mathematical Analysis, 8(4):573–579, August 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

AlvarezPereira:1990:GSP

[AV90] Carlos Alvarez Pereira and José M. Vega. Global stability of a premixed reaction zone (time-dependent Liñan’s problem). SIAM Journal on Mathematical Analysis, 21(4):884–904, July 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aktosun:1991:SIS

[Av91] Tuncay Aktosun and Cornelis van der Mee. Solution of the inverse scattering problem for the three-dimensional Schrödinger equation using a Fredholm integral equation. SIAM Journal on Mathematical Analysis, 22(3): 717–731, May 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Avrin:1988:GEM

[Avr88] Joel D. Avrin. Global existence for a model of electrophoretic separation. SIAM Journal on Mathematical Analysis, 19(3): 520–527, May 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Avrin:1991:QTC

[Avr91] Joel D. Avrin. Qualitative theory of the Cauchy problem for a one-step reaction model on bounded domains. SIAM Journal on Mathematical Analysis, 22 (2):379–391, March 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Avrin:1993:ABO

[Avr93] Joel D. Avrin. Asymptotic behavior of one-step combustion models with multiple reactants on bounded domains. SIAM Journal on Mathematical Analysis, 24(2):290–298, March 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Anderson:1990:FIC

[AVV90] G. D. Anderson, M. K. Vamanamurthy, and M. Vuorinen. Functional inequalities for complete elliptic integrals and their ratios. SIAM Journal on Mathematical Analysis, 21(2):536–549, March 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Anderson:1992:FIH

[AVV92] G. D. Anderson, M. K. Vamanamurthy, and M. Vuorinen. Functional inequalities for hypergeometric functions and complete elliptic integrals. SIAM Journal on Mathematical Analysis, 23 (2):512–524, March 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Aronsson:1978:EBV

[AW78] Gunnar Aronsson and Bengt Winzell. On an elliptic boundary value problem with mixed boundary conditions, arising in suspended sediment transport theory. SIAM Journal on Mathematical Analysis, 9(1):142–150, February 1978. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Askey:1979:SOP

[AW79] Richard Askey and James Wilson. A set of orthogonal polynomials that generalize the Racah coefficients or 6 − j symbols. SIAM Journal on Mathematical Analysis, 10(5):1008–1016, September 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Askey:1982:SHO

[AW82] Richard Askey and James Wilson. A set of hypergeometric orthogonal polynomials. SIAM Journal on Mathematical Analysis, 13(4):651–655, July 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Alzer:1998:IPF

[AW98] Horst Alzer and Jim Wells. Inequalities for the polygamma functions. SIAM Journal on Mathematical Analysis, 29(6): 1459–1466, November 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/32507.

Ayeni:1983:BPS

[Aye83] R. O. Ayeni. On the blow up problem for semilinear heat equations. SIAM Journal on Mathematical Analysis, 14(1):138–141, January 1983. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Amir:1979:EGC

[AZ79] Dan Amir and Zvi Ziegler. Expansions of generalized completely convex functions. SIAM Journal on Mathematical Analysis, 10(3):643–654, May 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Azzam:1980:DPE

[Azz80] A. Azzam. On Dirichlet’s problem for elliptic equations in sectionally smooth n-dimensional domains. SIAM Journal on Mathematical Analysis, 11(2): 248–253, March 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Azzam:1981:DPE

[Azz81] A. Azzam. On Dirichlet’s problem for elliptic equations in sectionally smooth n-dimensional domains. II. SIAM Journal on Mathematical Analysis, 12(2): 242, March 1981. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bailey:1971:VHF

[Bai71] Paul B. Bailey. A variation of Hadamard’s finite part integrals. SIAM Journal on Mathematical Analysis, 2(4):500–504, November 1971. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Baker:1972:BPS

[Bak72] James D. Baker. Boundedness properties of stationary linear operators. SIAM Journal on Mathematical Analysis, 3(4): 589–594, November 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bal:1999:FOC

[Bal99] Guillaume Bal. First-order corrector for the homogenization of the criticality eigenvalue problem in the even parity formulation of the neutron transport. SIAM Journal on Mathematical Analysis, 30(6):1208–1240, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33885.

Bandle:1973:RFK

[Ban73] Catherine Bandle. The Rayleigh– Faber–Krahn theorem for the characteristic values associated with a class of nonlinear boundary value problems. SIAM Journal on Mathematical Analysis, 4 (1):8–14, February 1973. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bange:1974:CET

[Ban74] David W. Bange. A constructive existence theorem for a nonlinear parabolic equation. SIAM Journal on Mathematical Analysis, 5 (1):103–110, February 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bandle:1975:BSP

[Ban75] Catherine Bandle. Bounds for the solutions of Poisson problems and applications to nonlinear eigenvalue problems. SIAM Journal on Mathematical Analysis, 6(1):146–152, February 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bandle:1978:EGF

[Ban78] Catherine Bandle. Estimates for the Green’s functions of elliptic operators. SIAM Journal on Mathematical Analysis, 9(6): 1126–1136, December 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bandle:1979:RPP

[Ban79] Catherine Bandle. A remark on a paper: “On some maximum principles involving harmonic functions and their derivatives” [SIAM J. Math. Anal. 10 (1979), no. 1, 96–104; MR 81f:31011a]. SIAM Journal on Mathematical Analysis, 10(6): 1244–1245, November 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [PP79].

Banasiak:1992:ASE

[Ban92] Jacek Banasiak. On asymptotics of solutions of elliptic mixed boundary value problems of second-order in domains with vanishing edges. SIAM Journal on Mathematical Analysis, 23(5): 1117–1124, September 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bao:1998:SBC

[Bao98] Gang Bao. Smoothness between coefficients and boundary values for the wave equation. SIAM Journal on Mathematical Analysis, 29(2):380–394, March 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/29936.

Barbu:1975:NVE

[Bar75] Viorel Barbu. Nonlinear Volterra equations in a Hilbert space. SIAM Journal on Mathematical Analysis, 6(4):728–741, August 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barbu:1977:NVI

[Bar77a] Viorel Barbu. On a nonlinear Volterra integral equation on a Hilbert space. SIAM Journal on Mathematical Analysis, 8 (2):346–355, April 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bareket:1977:IIF

[Bar77b] Miriam Bareket. On an isoperimetric inequality for the first eigenvalue of a boundary value problem. SIAM Journal on Mathematical Analysis, 8(2): 280–287, April 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barbu:1979:ENV

[Bar79a] Viorel Barbu. Existence for nonlinear Volterra equations in Hilbert spaces. SIAM Journal on Mathematical Analysis, 10(3): 552–569, May 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barnes:1979:SCS

[Bar79b] Earl R. Barnes. Stability conditions for second-order ordinary differential equations with periodic coefficients. SIAM Journal on Mathematical Analysis, 10(5): 985–1001, September 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barnes:1981:SCL

[Bar81] Earl R. Barnes. Stability conditions for linear Hamiltonian systems with periodic coefficients. SIAM Journal on Mathematical Analysis, 12(1):60–71, January 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barbey:1983:PIL

[Bar83] Klaus Barbey. Periodically invariant linear systems. SIAM Journal on Mathematical Analysis, 14(1):158–166, January 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barge:1984:IMO

[Bar84a] Marcy Barge. Invariant manifolds and the onset of reversal in the Rikitake two-disk dynamo. SIAM Journal on Mathematical Analysis, 15(3):514–529, May 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barker:1984:IFS

[Bar84b] John W. Barker. Interactions of fast and slow waves in problems with two time scales. SIAM Journal on Mathematical Analysis, 15 (3):500–513, May 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barnes:1985:EPEb

[Bar85a] David C. Barnes. Extremal problems for eigenvalue functionals. SIAM Journal on Mathematical Analysis, 16(6):1284–1294, November 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barnes:1985:EPEa

[Bar85b] David C. Barnes. Extremal problems for eigenvalues with applications to buckling, vibration and sloshing. SIAM Journal on Mathematical Analysis, 16 (2):341–357, March 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barnes:1987:SAF

[Bar87] David C. Barnes. Some approximation formulas for stochastic eigenvalues. SIAM Journal on Mathematical Analysis, 18(4): 933–940, July 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barnes:1988:EPE

[Bar88] David C. Barnes. Extremal problems for eigenvalue functionals. II. SIAM Journal on Mathematical Analysis, 19(5):1151–1161, September 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ben-Artzi:1990:ITC

[BAR90] Asher Ben-Artzi and Amos Ron.

On the integer translates of a compactly supported function: Dual bases and linear projectors. SIAM Journal on Mathematical Analysis, 21(6):1550– 1562, November 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barnes:1991:IEP

[Bar91] David C. Barnes. The inverse eigenvalue problem with finite data. SIAM Journal on Mathematical Analysis, 22(3):732–753, May 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Baxley:1971:EET

[Bax71] John V. Baxley. Extreme eigenvalues of Toeplitz matrices associated with certain orthogonal polynomials. SIAM Journal on Mathematical Analysis, 2 (3):470–482, August 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Baxley:1991:SSN

[Bax91] John V. Baxley. Some singular nonlinear boundary value problems. SIAM Journal on Mathematical Analysis, 22(2):463–479, March 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Binding:1977:VAM

[BB77] Paul Binding and Patrick J. Browne. A variational approach to multiparameter eigenvalue problems for matrices. SIAM Journal on Mathematical Analysis, 8(5):763–777, October 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Binding:1978:VAM

[BB78] Paul Binding and Patrick J. Browne. A variational approach to multi-parameter eigenvalue problems in Hilbert space. SIAM Journal on Mathematical Analysis, 9(6):1054–1067, December 1978. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brown:1979:EPS

[BB79] K. J. Brown and H. Budin. On the existence of positive solutions for a class of semilinear elliptic boundary value problems. SIAM Journal on Mathematical Analysis, 10(5):875–883, September 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Basu:1983:MSS

[BB83] S. Basu and N. K. Bose. Matrix Stieltjes series and network models. SIAM Journal on Mathematical Analysis, 14(2):209–222, March 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Blat:1986:GBP

[BB86] J. Blat and K. J. Brown. Global bifurcation of positive solutions in some systems of elliptic equations. SIAM Journal on Mathematical Analysis, 17(6):1339– 1353, November 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bamberger:1990:MAG

[BB90] A. Bamberger and A. S. Bonnet. Mathematical analysis of the guided modes of an optical fiber. SIAM Journal on Mathematical Analysis, 21(6):1487– 1510, November 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bebernes:1992:FTB

[BB92] J. Bebernes and S. Bricher. Final time blowup profiles for semilinear parabolic equations via center manifold theory. SIAM Journal on Mathematical Analysis, 23 (4):852–869, July 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bourgeois:1994:VQM

[BB94] Alfred J. Bourgeois and J. Thomas Beale. Validity of the quasigeostrophic model for large-scale flow in the atmosphere and ocean. SIAM Journal on Mathematical Analysis, 25(4):1023– 1068, July 1994. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

http://epubs.siam.org/sam-
bin/dbq/article/23498.

Braden:1997:GAS

[BB97] H. W. Braden and V. M. Buchstaber. The general analytic solution of a functional equation of addition type. SIAM Journal on Mathematical Analysis, 28 (4):903–923, July 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29138.

Barenblatt:1993:DPR

[BBDU93] G. I. Barenblatt, M. Bertsch, R. Dal Passo, and M. Ughi. A degenerate pseudoparabolic regularization of a nonlinear forwardbackward heat equation arising in the theory of heat and mass exchange in stably stratified turbulent shear flow. SIAM Journal on Mathematical Analysis, 24(6): 1414–1439, November 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bucur:1999:AEL

[BBF99] Dorin Bucur, Giuseppe Buttazzo, and Isabel Figueiredo. On the attainable eigenvalues of the Laplace operator. SIAM Journal on Mathematical Analysis, 30(3):527–536, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32913.

Bereux:1997:GDS

[BBL97] François Bereux, Eric Bonnetier, and Philippe G. LeFloch. Gas dynamics system: Two special cases. SIAM Journal on Mathematical Analysis, 28(3): 499–515, May 1997. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/28583.

Bellout:1993:EGW

[BBN93] Hamid Bellout, Frederick Bloom, and Jindřich Nečas. Existence of global weak solutions to the dynamical problem for a threedimensional elastic body with singular memory. SIAM Journal on Mathematical Analysis, 24 (1):36–45, January 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Baider:1977:HIG

[BC77] Alberto Baider and Barry Cherkas. Hypoelliptic infinitesimal generators. SIAM Journal on Mathematical Analysis, 8(6): 971–977, November 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Borosh:1979:CFT

[BC79] Itshak Borosh and Charles K.

Chui. On characterization of functions by their Gauss– Chebyshev quadratures. SIAM Journal on Mathematical Analysis, 10(3):532–541, May 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Baider:1990:MGS

[BC90a] A. Baider and R. C. Churchill. On monodromy groups of second-order Fuchsian equations. SIAM Journal on Mathematical Analysis, 21(6):1642–1652, November 1990. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bruschi:1990:GAS

[BC90b] M. Bruschi and F. Calogero. General analytic solution of certain functional equations of addition type. SIAM Journal on Mathematical Analysis, 21 (4):1019–1030, July 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Biedenharn:1994:GTI

[BÇ94] L. C. Biedenharn and A. K. Çiftçi. Group theoretical interpretations of special function identities: Two examples. SIAM Journal on Mathematical Analysis, 25(2):274–287, March 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/23033.

Bayada:1995:ICT

[BC95] Guy Bayada and Michèle Chambat. On interface conditions for a thin film flow past a porous medium. SIAM Journal on Mathematical Analysis, 26(5): 1113–1129, September 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24661.

Bao:1997:NGP

[BC97] Gang Bao and Yunmei Chen.

A nonlinear grating problem in diffractive optics. SIAM Journal on Mathematical Analysis, 28 (2):322–337, March 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/28446.

Bell:1982:SFD

[BCB82] Jonathan Bell, Chris Cosner, and Willy Bertiger. Solutions for a flux-dependent diffusion model. SIAM Journal on Mathematical Analysis, 13(5):758–769, September 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brezzi:1989:SPA

[BCG89] F. Brezzi, A. C. S. Capelo, and L. Gastaldi. A singular perturbation analysis of reverse-biased semiconductor diodes. SIAM Journal on Mathematical Analysis, 20(2):372–387, March 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bauman:1993:ZSG

[BCP93] Patricia Bauman, Neil N. Carlson, and Daniel Phillips. On the zeros of solutions to Ginzburg– Landau type systems. SIAM Journal on Mathematical Analysis, 24(5):1283–1293, September 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barcilon:1988:ESS

[BCT88] V. Barcilon, P. Constantin, and E. S. Titi. Existence of solutions to the Stommel–Charney model of the Gulf Stream. SIAM Journal on Mathematical Analysis, 19 (6):1355–1364, November 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bonafede:1998:FSP

[BCT98] S. Bonafede, G. R. Cirmi, and A. F. Tedeev. Finite speed of propagation for the porous media equation. SIAM Journal on Mathematical Analysis, 29(6): 1381–1398, November 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/29807.

Baker:1978:SIF

[BD78] James D. Baker and James A. Dyer. A Stieltjes integral formula for Daniell functionals. SIAM Journal on Mathematical Analysis, 9(4):743–750, August 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Boyd:1986:UAS

[BD86] W. G. C. Boyd and T. M. Dunster. Uniform asymptotic solutions of a class of second-order linear differential equations having a turning point and a regular singularity, with an application to Legendre functions. SIAM Journal on Mathematical Analysis, 17(2):422–450, March 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

BenAbdallah:1995:CLL

[BD95] Naoufel Ben Abdallah and Pierre Degond. The Child–Langmuir law for the Boltzmann equation of semiconductors. SIAM Journal on Mathematical Analysis, 26 (2):364–398, March 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24656.

Butler:1983:NIR

[BE83] G. J. Butler and L. H. Erbe. Nonlinear integral Riccati systems and comparison theorems for linear differential equations. SIAM Journal on Mathematical Analysis, 14(3):463–473, May 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Butler:1986:ORS

[BE86] G. J. Butler and L. H. Erbe. Oscillation results for second order differential systems. SIAM Journal on Mathematical Analysis, 17 (1):19–29, January 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Beatson:1981:CAS

[Bea81] R. K. Beatson. Convex approximation by splines. SIAM Journal on Mathematical Analysis, 12(4): 549–559, July 1981. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Beattie:1985:CES

[Bea85] Christopher Beattie. Convergence of essential spectra for intermediate Hamiltonians. SIAM Journal on Mathematical Analysis, 16(3):492–499, May 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brauner:1987:ARU

[BEGv87] C.-M. Brauner, W. Eckhaus, M. Garbey, and A. van Harten. Asymptotics of a rather unusual type in a free boundary problem. SIAM Journal on Mathematical Analysis, 18(3):812–841, May 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barnsley:1989:HVF

[BEHM89] M. F. Barnsley, J. Elton, D. Hardin, and P. Massopust. Hidden variable fractal interpolation functions. SIAM Journal on Mathematical Analysis, 20(5): 1218–1242, September 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bell:1981:NCPa

[Bel81a] John B. Bell. The noncharacteristic Cauchy problem for a class of equations with time dependence. I. Problems in one space dimension. SIAM Journal on Mathematical Analysis, 12(5): 759–777, September 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bell:1981:NCPb

[Bel81b] John B. Bell. The noncharacteristic Cauchy problem for a class of equations with time dependence. II. Multidimensional problems. SIAM Journal on Mathematical Analysis, 12(5): 778–797, September 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Belevitch:1982:GHR

[Bel82] Vitold Belevitch. The Gauss hypergeometric ratio as a positive real function. SIAM Journal on Mathematical Analysis, 13(6): 1024–1040, November 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bellout:1987:CBS

[Bel87] Hamid Bellout. A criterion for blow-up of solutions to semilinear heat equations. SIAM Journal on Mathematical Analysis, 18(3): 722–727, May 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Benton:1975:C

[Ben75]  T. C. Benton.  Concern-

ing ∫_{0}^{∞} e^{−at}J_{μ}(bt)J_{ν}(ct)t^{μ−ν} dt.

SIAM Journal on Mathematical Analysis, 6(5):761–765, October 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Benson:1981:PSL

[Ben81]  Donald C. Benson.  Princi-

pal solutions for Liénard’s equation. SIAM Journal on Mathematical Analysis, 12(3):398–412, May 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Benedetto:1982:CPS

[Ben82a] John J. Benedetto. A closure problem for signals in semigroup invariant systems. SIAM Journal on Mathematical Analysis, 13 (2):180–207, March 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Benzinger:1982:RSP

[Ben82b] Harold E. Benzinger. Rayleigh– Schrödinger perturbation of semigroups. SIAM Journal on Mathematical Analysis, 13(4): 515–531, July 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Benedetto:1984:LUP

[Ben84a] John J. Benedetto. A local uncertainty principle. SIAM Journal on Mathematical Analysis, 15(5): 988–995, September 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Benson:1984:PTL

[Ben84b] Donald C. Benson. A Prüfer transformation for Liénard’s equation. SIAM Journal on Mathematical Analysis, 15(4): 656–669, July 1984. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Benzinger:1985:SRC

[Ben85] Harold E. Benzinger. Strong resolvent convergence of diffusion operators. SIAM Journal on Mathematical Analysis, 16(4): 713–724, July 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bennett:1986:SOF

[Ben86] Allan Bennett. Symmetry in an overdetermined fourth order elliptic boundary value problem. SIAM Journal on Mathematical Analysis, 17(6):1354–1358, November 1986. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Benedetto:1991:SWP

[Ben91] John J. Benedetto. The spherical Wiener–Plancherel formula and spectral estimation. SIAM Journal on Mathematical Analysis, 22 (4):1110–1130, July 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

BenAbdallah:1996:CCL

[Ben96] Naoufel Ben Abdallah. Convergence of the Child–Langmuir asymptotics of the Boltzmann equation of semiconductors.

SIAM Journal on Mathematical Analysis, 27(1):92–109, January 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/25964.

Bernfeld:1970:NLF

[Ber70] Stephen R. Bernfeld. Noncontinuous Lyapunov functions and extendability of solutions. SIAM Journal on Mathematical Analysis, 1(4):515–523, November 1970. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bernfeld:1972:PUU

[Ber72] Stephen R. Bernfeld. Perturbing uniform ultimate bounded differential systems. SIAM Journal on Mathematical Analysis, 3(2): 358–370, May 1972. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Berndt:1975:ECC

[Ber75a] Bruce C. Berndt. The evaluation of certain classes of nonabsolutely convergent double series. SIAM Journal on Mathematical Analysis, 6(6):966–977, November 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bertran:1975:NOP

[Ber75b] M. Bertran. Note on orthogonal polynomials in v-variables. SIAM Journal on Mathematical Analysis, 6(2):250–257, April 1975.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Berkey:1976:AFT

[Ber76] Dennis D. Berkey. An asymptotic Floquet theorem for linear almost periodic systems. SIAM Journal on Mathematical Analysis, 7(6):843–847, November 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bertozzi:1988:HOC

[Ber88] Andrea Louise Bertozzi. Heteroclinic orbits and chaotic dynamics in planar fluid flows. SIAM Journal on Mathematical Analysis, 19(6):1271–1294, November 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Berrone:1999:CRM

[Ber99] Lucio R. Berrone. On a conjecture relative to the maxima of harmonic functions on convex domains. SIAM Journal on Mathematical Analysis, 30(6):1185–1207, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/33497.

Busby:1979:CPL

[BF79] Robert C. Busby and Wyman Fair. Convergence of “periodic in the limit” operator continued fractions. SIAM Journal on Mathematical Analysis, 10(3):

512–522, May 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Blanchard:1988:SDN

[BF88] Dominique Blanchard and Gilles A. Francfort. Study of a doubly nonlinear heat equation with no growth assumptions on the parabolic term. SIAM Journal on Mathematical Analysis, 19(5): 1032–1056, September 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [BF89b].

Bellout:1989:BEN

[BF89a] Hamid Bellout and Avner Friedman. Blow-up estimates for a nonlinear hyperbolic heat equation. SIAM Journal on Mathematical Analysis, 20(2):354–366, March 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Blanchard:1989:ESD

[BF89b] D. Blanchard and G. A. Francfort. Erratum: “Study of a doubly nonlinear heat equation with no growth assumptions on the parabolic term”. SIAM Journal on Mathematical Analysis, 20 (3):761–762, May 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [BF88].

Bardi:1998:HTE

[BF98] Martino Bardi and Silvia Faggian. Hopf-type estimates and formulas for nonconvex nonconcave Hamilton–Jacobi equations. SIAM Journal on Mathematical Analysis, 29(5):1067– 1086, September 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30962.

Bishop:1990:CBC

[BFF^{+}90] Alan R. Bishop, Randy Flesch, M. Gregory Forest, David W. McLaughlin, and Edward A. Overman, II. Correlations between chaos in a perturbed Sine-Gordon equation and a truncated model system. SIAM Journal on Mathematical Analysis, 21(6): 1511–1536, November 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bates:1997:ETW

[BFGJ97] P. W. Bates, P. C. Fife, R. A.

Gardner, and C. K. R. T. Jones. The existence of travelling wave solutions of a generalized phasefield model. SIAM Journal on Mathematical Analysis, 28 (1):60–93, January 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/28382.

Barnard:1986:ZIP

[BFW86] Roger W. Barnard, Wayne T.

Ford, and Hsing Y. Wang. On zeros of interpolating polynomials. SIAM Journal on Mathematical Analysis, 17(3):734–744, May 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bustoz:1974:GBP

[BG74] J. Bustoz and C. W. Groetsch. On generalized Bernstein polynomials. SIAM Journal on Mathematical Analysis, 5(2):256–262, April 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bustoz:1975:UAO

[BG75] J. Bustoz and C. W. Groetsch. Uniform approximation operators generated by a power series identity. SIAM Journal on Mathematical Analysis, 6(3):472–479, May 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bates:1976:MGF

[BG76] P. W. Bates and G. B. Gustafson. Maximization of Green’s function over classes of multipoint boundary value problems. SIAM Journal on Mathematical Analysis, 7(6):858–871, November 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Berndt:1984:APA

[BG84] Bruce C. Berndt and Larry A. Goldberg. Analytic properties of arithmetic sums arising in the theory of the classical theta functions. SIAM Journal on Mathematical Analysis, 15(1):143–150, January 1984. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Babuska:1988:RSE

[BG88] I. Babuška and B. Q. Guo. Regularity of the solution of elliptic problems with piecewise analytic data. Part I. boundary value problems for linear elliptic equation of second order. SIAM Journal on Mathematical Analysis, 19 (1):172–203, January 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Babuska:1989:RSE

[BG89] I. Babuška and B. Q. Guo. Regularity of the solution of elliptic problems with piecewise analytic data. II. The trace spaces and application to the boundary value problems with nonhomogeneous boundary conditions. SIAM Journal on Mathematical Analysis, 20(4):763–781, July 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Boersma:1991:AEC

[BG91] J. Boersma and M. L. Glasser.

Asymptotic expansion of a class of Fermi–Dirac integrals. SIAM Journal on Mathematical Analysis, 22(3):810–820, May 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Benzoni-Gavage:1996:RFB

[BG96] Sylvie Benzoni-Gavage. On a representation formula for B. Temple systems. SIAM Journal on Mathematical Analysis, 27(6): 1503–1519, November 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/26883.

Bruce:1984:CEE

[BGG84] J. W. Bruce, P. J. Giblin, and C. G. Gibson. Caustics in extended Euclidean space. SIAM Journal on Mathematical Analysis, 15(1):50–68, January 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Blum:1986:ECP

[BGS86] J. Blum, T. Gallouët, and J. Simon. Existence and control of plasma equilibrium in a tokamak. SIAM Journal on Mathematical Analysis, 17(5):1158– 1177, September 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brefort:1988:APN

[BGT88] B. Brefort, J. M. Ghidaglia, and R. Temam. Attractors for the penalized Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 19(1):1–21, January 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bancroft:1975:BSN

[BH75a] Stephen Bancroft and Thomas G. Hallam. Bounded solutions of nonlinear differential equations with parameters. SIAM Journal on Mathematical Analysis, 6(2):236–241, April 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Berger:1975:AEF

[BH75b] Neil Berger and Richard Handelsman. Asymptotic evaluation of fractional integral operators with applications. SIAM Journal on Mathematical Analysis, 6 (5):766–773, October 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bleistein:1975:MSP

[BH75c] Norman Bleistein and Richard A.

Handelsman. Multidimensional stationary phase an alternative derivation. SIAM Journal on Mathematical Analysis, 6(3): 480–487, May 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Burns:1976:AST

[BH76] J. A. Burns and T. L. Herdman. Adjoint semigroup theory for a class of functional differential equations. SIAM Journal on Mathematical Analysis, 7 (5):729–745, October 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Burchard:1985:WEC

[BH85] H. G. Burchard and K. Höllig.

N-width and entropy of H_{p}- classes in L_{q}(−1, 1). SIAM Journal on Mathematical Analysis, 16 (2):405–421, March 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bertsch:1986:DDD

[BH86] M. Bertsch and D. Hilhorst.

A density dependent diffusion equation in population dynamics: stabilization to equilibrium. SIAM Journal on Mathematical Analysis, 17(4):863–883, July 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bamberger:1987:EES

[BH87] Alain Bamberger and Laurence Halpern. Étude des états stationnaires pour une équation de Schrödinger non linéaire comportant un terme non autonome. (French) [Study of stationary states for a nonlinear Schrödinger equation with a nonautonomous term]. SIAM Journal on Mathematical Analysis, 18(1):97–126, January 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Benke:1993:ELD

[BH93] George Benke and W. J. Hendricks. Estimates for large deviations in random trigonometric polynomials. SIAM Journal on Mathematical Analysis, 24 (4):1067–1085, July 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bonnet:1999:ENS

[BH99] Alexis Bonnet and François Hamel. Existence of nonplanar solutions of a simple model of premixed Bunsen flames. SIAM Journal on Mathematical Analysis, 31(1):80–118, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/31639.

Bleistein:1972:FWF

[BHL72] N. Bleistein, R. A. Handelsman, and J. S. Lew. Functions whose Fourier transforms decay at infinity: An extension of the Riemann–Lebesgue lemma. SIAM Journal on Mathematical Analysis, 3(3):485–495, August 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bernfeld:1975:AEN

[BHL75] S. R. Bernfeld, T. G. Hallam, and V. Lakshmikantham. Asymptotic equivalence of nonlinear differential equations in Banach spaces. SIAM Journal on Mathematical Analysis, 6(1):140–145, February 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bronikowski:1972:QEN

[BHN72] T. A. Bronikowski, J. E. Hall, and J. A. Nohel. Quantitative estimates for a nonlinear system of integrodifferential equations arising in reactor dynamics. SIAM Journal on Mathematical Analysis, 3(4):567–588, November 1972. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Burns:1983:LFD

[BHS83] John A. Burns, Terry L. Herdman, and Harlan W. Stech. Linear functional-differential equations as semigroups on product spaces. SIAM Journal on Mathematical Analysis, 14(1):98–116, January 1983. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Benoist:1996:WSC

[BHU96] J. Benoist and J.-B. Hiriart-Urruty. What is the subdifferential of the closed convex hull of a function? SIAM Journal on Mathematical Analysis, 27(6): 1661–1679, November 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/26593.

Boas:1975:ECI

[BI75] R. P. Boas, Jr. and C. O. Imoru. Elementary convolution inequalities. SIAM Journal on Mathematical Analysis, 6(3):457–471, May 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bustoz:1983:TIU

[BI83] Joaquin Bustoz and Mourad E. H. Ismail. Turán inequalities for ultraspherical and continuous q-ultraspherical polynomials. SIAM Journal on Mathematical Analysis, 14(4):807–818, July 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bustoz:1989:PTS

[BI89] Joaquin Bustoz and Mourad E. H. Ismail. A positive trigonometric sum. SIAM Journal on Mathematical Analysis, 20(1):

176–181, January 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bilodeau:1974:GHP

[Bil74] G. G. Bilodeau. On generalized heat polynomials. SIAM Journal on Mathematical Analysis, 5 (1):43–50, February 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Biler:1990:ASS

[Bil90] Piotr Biler. Attractors for the system of Schrödinger and Klein– Gordon equations with Yukawa coupling. SIAM Journal on Mathematical Analysis, 21(5): 1190–1212, September 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Binding:1982:MVP

[Bin82] Paul Binding. Multiparameter variational principles. SIAM Journal on Mathematical Analysis, 13(5):842–855, September 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Busenberg:1991:GBA

[BIT91] Stavros N. Busenberg, Mimmo Iannelli, and Horst R. Thieme. Global behavior of an agestructured epidemic model. SIAM Journal on Mathematical Analysis, 22(4):1065–1080, July 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Balser:1981:RCP

[BJL81] W. Balser, W. B. Jurkat, and D. A. Lutz. On the reduction of connection problems for differential equations with an irregular singular point to ones with only regular singularities. I. SIAM Journal on Mathematical Analysis, 12(5):691–721, September 1981. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Balser:1988:RCP

[BJL88] W. Balser, W. B. Jurkat, and D. A. Lutz. On the reduction of connection problems for differential equations with an irregular singular point to ones with only regular singularities. II. SIAM Journal on Mathematical Analysis, 19(2):398–443, March 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barron:1998:AHL

[BJL98] E. N. Barron, R. Jensen, and W. Liu. Applications of the Hopf–Lax formula for u_{t} + H(u, Du) = 0. SIAM Journal on Mathematical Analysis, 29 (4):1022–1039, July 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/31996.

Brown:1976:SIS

[BK76] James Ward Brown and Marek Kuczma. Self-inverse Sheffer sequences. SIAM Journal on Mathematical Analysis, 7(5):723–728, October 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bertsch:1990:SDP

[BK90] M. Bertsch and S. Kamin. A system of degenerate parabolic equations. SIAM Journal on Mathematical Analysis, 21(4): 905–916, July 1990. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barnes:1995:IEP

[BK95] David C. Barnes and Roger Knobel. The inverse eigenvalue problem with finite data for partial differential equations. SIAM Journal on Mathematical Analysis, 26(3):616–632, May 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/25313.

Boyer:1986:SEI

[BKM86] C. P. Boyer, E. G. Kalnins, and W. Miller, Jr. Stäckelequivalent integrable Hamiltonian systems. SIAM Journal on Mathematical Analysis, 17(4): 778–797, July 1986. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Boyer:1985:SVH

[BKW85] C. P. Boyer, E. G. Kalnins, and P. Winternitz. Separation of variables for the Hamilton–Jacobi equation on complex projective spaces. SIAM Journal on Mathematical Analysis, 16(1):93–109, January 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bojanic:1974:ERC

[BL74] R. Bojanic and Y. H. Lee. An estimate for the rate of convergence of convolution products of sequences. SIAM Journal on Mathematical Analysis, 5(3):452–462, May 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barrar:1984:OMM

[BL84] R. B. Barrar and H. L. Loeb. Optimal monosplines with a maximal number of zeros. SIAM Journal on Mathematical Analysis, 15 (6):1196–1204, November 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bebernes:1990:FTB

[BL90] J. Bebernes and A. Lacey. Finitetime blowup for a particular parabolic system. SIAM Journal on Mathematical Analysis, 21(6): 1415–1425, November 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bleistein:1977:AEI

[Ble77] Norman Bleistein. Asymptotic expansions of integral transforms of functions with logarithmic singularities. SIAM Journal on Mathematical Analysis, 8(4): 655–672, August 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bourgeat:1996:CHP

[BLM96] Alain Bourgeat, Stephan Luckhaus, and Andro Mikelic. Convergence of the homogenization process for a double-porosity model of immiscible two-phase flow. SIAM Journal on Mathematical Analysis, 27(6):1520– 1543, November 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/27645.

Bonan:1987:OPT

[BLN87] S. Bonan, D. S. Lubinsky, and P. Nevai. Orthogonal polynomials and their derivatives, II. SIAM Journal on Mathematical Analysis, 18(4):1163–1176, July 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bloom:1980:BSC

[Blo80] Frederick Bloom. Bounds for solutions to a class of integrodifferential equations associated with a theory of rigid nonconducting material dielectrics. SIAM Journal on Mathematical Analysis, 11(2):265–291, March 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bresch:1999:VDM

[BLS99] Didier Bresch, Jérome Lemoine, and Jacques Simon. A vertical diffusion model for lakes. SIAM Journal on Mathematical Analysis, 30(3):603–622, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/32294.

Boas:1973:IIF

[BM73] R. P. Boas, Jr. and M. B. Marcus. Inequalities involving a function and its inverse. SIAM Journal on Mathematical Analysis, 4(4): 585–591, November 1973. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barany:1992:FSE

[BM92] E. Barany and I. Melbourne. A family of stable equilibria in bifurcation with spherical symmetry. SIAM Journal on Mathematical Analysis, 23(1):72–80, January 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Boccardo:1992:ESN

[BMP92] L. Boccardo, F. Murat, and J.- P. Puel. L_{∞} estimate for some nonlinear elliptic partial differential equations and application to an existence result. SIAM Journal on Mathematical Analysis, 23 (2):326–333, March 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Buchner:1983:EID

[BMS83] Michael Buchner, Jerrold Marsden, and Stephen Schecter. Examples for the infinite dimensional Morse lemma. SIAM Journal on Mathematical Analysis, 14 (6):1045–1055, November 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See remark [GM83, CGR88].

Bucy:1977:HSC

[BMY77] R. S. Bucy, A. J. Mallinckrodt, and H. Youssef. High speed convolution of periodic functions. SIAM Journal on Mathematical Analysis, 8(4):619–625, August 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bailey:1978:IEN

[BN78] Paul B. Bailey and Michael J.

Norris. On the interval of existence for nonlinear two-point boundary value problems. SIAM Journal on Mathematical Analysis, 9(3):535–557, June 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Berestycki:1985:TWS

[BNS85] Henri Berestycki, Basil Nicolaenko, and Bruno Scheurer. Traveling wave solutions to combustion models and their singular limits. SIAM Journal on Mathematical Analysis, 16(6): 1207–1242, November 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bandle:1998:PME

[BNS98] Catherine Bandle, Tokumori Nanbu, and Ivar Stakgold. Porous medium equation with absorption. SIAM Journal on Mathematical Analysis, 29(5): 1268–1278, September 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/31142.

Bardi:1991:NMD

[BO91] Martino Bardi and Stanley Osher. The nonconvex multidimensional Reimann problem for Hamilton–Jacobi equations. SIAM Journal on Mathematical Analysis, 22(2):344–351, March 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Biagioni:1992:GSK

[BO92] H. A. Biagioni and M. Oberguggenberger. Generalized solutions to the Korteweg–de Vries and the regularized long-wave equations. SIAM Journal on Mathematical Analysis, 23(4): 923–940, July 1992. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Boersma:1978:CMI

[Boe78] J. Boersma. On certain multiple integrals occurring in a waveguide scattering problem. SIAM Journal on Mathematical Analysis, 9(2):377–393, April 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Boese:1995:SCS

[Boe95] F. G. Boese. Stability criteria for second-order dynamical systems involving several time delays. SIAM Journal on Mathematical Analysis, 26(5):1306– 1330, September 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/20084.

Bogar:1974:OOD

[Bog74] G. A. Bogar. Oscillations of n-th order differential equations with retarded argument. SIAM Journal on Mathematical Analysis, 5(3):473–481, May 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bohe:1990:FLS

[Boh90] Adriana Bohé. Free layers in a singularly perturbed boundary value problem. SIAM Journal on Mathematical Analysis, 21(5): 1264–1280, September 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bojanov:1994:CSI

[Boj94] Borislav Bojanov. Characterization of the smoothest interpolant. SIAM Journal on Mathematical Analysis, 25(6):1642– 1655, November 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bonnet:1996:TWS

[Bon96] Alexis Bonnet. Traveling-wave solutions to combustion models for a reversible reaction. SIAM Journal on Mathematical Analysis, 27(5):1270–1285, September 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/26295.

Borwein:1985:RII

[Bor85] Peter B. Borwein. Rational interpolation to e^{x}, II. SIAM Journal on Mathematical Analysis, 16(3): 656–662, May 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Borodin:1999:EGC

[Bor99] M. A. Borodin. Existence of the global classical solution for a two-phase Stefan problem. SIAM Journal on Mathematical Analysis, 30(6):1264–1281, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33253.

Bose:1979:RTC

[Bos79] Ramendra Krishna Bose. Realizability theory of continuous linear operators on groups. SIAM Journal on Mathematical Analysis, 10(4):767–777, July 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bouldin:1972:NCP

[Bou72] Richard Bouldin. The norm continuity properties of square roots. SIAM Journal on Mathematical Analysis, 3(2):206–210, May 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bowie:1971:UIH

[Bow71] Patrick C. Bowie. Uncertainty inequalities for Hankel transforms. SIAM Journal on Mathematical Analysis, 2(4):601–606, November 1971. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Boyer:1976:LTS

[Boy76] Charles P. Boyer. Lie theory and separation of variables for the equation iU_{t} + Δ_{2}U − (\frac{α}{x_{1}^{2}} + \frac{β}{x_{2}^{2}} ) U = 0. SIAM Journal on Mathematical Analysis, 7 (2):230–263, April 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Babuska:1990:PPH

[BP90] I. Babuška and J. Pitkäranta.

The plate paradox for hard and soft simple support. SIAM Journal on Mathematical Analysis, 21 (3):551–576, May 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bernis:1996:TPD

[BP96] F. Bernis and L. A. Peletier. Two problems from draining flows involving third-order ordinary differential equations. SIAM Journal on Mathematical Analysis, 27 (2):515–527, March 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/26084.

Bertsch:1991:ETD

[BPL91] M. Bertsch, L. A. Peletier, and S. M. Verduyn Lunel. The effect of temperature dependent viscosity on shear flow of incompressible fluids. SIAM Journal on Mathematical Analysis, 22 (2):328–343, March 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Butler:1971:PBD

[BR71] Geoffrey Butler and Thomas Rogers. Pointwise bounds on derivatives of solutions to ordinary differential equations. SIAM Journal on Mathematical Analysis, 2(4):521–528, November 1971. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Barnsley:1978:RAB

[BR78a] M. F. Barnsley and P. D. Robinson. Rational approximant bounds for a class of two-variable Stieltjes functions. SIAM Journal on Mathematical Analysis, 9 (2):272–290, April 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Breuer:1978:LBE

[BR78b] Shlomo Breuer and Joseph J. Roseman. A lower bound for the eigenvalues of the elliptic Dirichlet problem for a general domain in terms of its characteristic dimension. SIAM Journal on Mathematical Analysis, 9 (4):774–778, August 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brown:1981:IRC

[BR81] James Ward Brown and Steven M. Roman. Inverse relations for certain Sheffer sequences. SIAM Journal on Mathematical Analysis, 12(2):186–195, March 1981.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bruschi:1982:EFG

[BR82] Massimo Bruschi and Paolo Emilio Ricci. An explicit formula for f(A) and the generating functions of the generalized Lucas polynomials. SIAM Journal on Mathematical Analysis, 13(1): 162–165, January 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Butler:1984:SPS

[BR84] G. J. Butler and V. Sree Hari Rao. Some properties of second order linear differential equations and perturbations that preserve them. SIAM Journal on Mathematical Analysis, 15(5):912–921, September 1984. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Braaksma:1971:AAD

[Bra71a] B. L. J. Braaksma. Asymptotic analysis of a differential equation of turrittin. SIAM Journal on Mathematical Analysis, 2(1):1– 16, February 1971. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [Bra72].

Braaksma:1971:GRL

[Bra71b] B. L. J. Braaksma. Global reduction of linear differential systems involving a small singular parameter. SIAM Journal on Mathematical Analysis, 2(2):149–165, May 1971. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Braaksma:1972:EAA

[Bra72] B. L. J. Braaksma. Erratum: Asymptotic analysis of a differential equation of turrittin. SIAM Journal on Mathematical Analysis, 3(1):175, February 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [Bra71a].

Bragg:1974:LEE

[Bra74] L. R. Bragg. Linear evolution equations that involve products of commutative operators. SIAM Journal on Mathematical Analysis, 5(2):327–335, April 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brauer:1975:NIE

[Bra75] Fred Brauer. On a nonlinear integral equation for population growth problems. SIAM Journal on Mathematical Analysis, 6 (2):312–317, April 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bragg:1976:RLI

[Bra76] L. R. Bragg. The Riemann– Liouville integral and parameter shifting in a class of linear abstract Cauchy problems. SIAM Journal on Mathematical Analysis, 7(1):1–12, February 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brauer:1979:DRS

[Bra79] Fred Brauer. Decay rates for solutions of a class of differentialdifference equations. SIAM Journal on Mathematical Analysis, 10 (4):783–788, July 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bragg:1981:RWE

[Bra81a] L. R. Bragg. The radial wave and Euler–Poisson–Darboux equations with singular data. SIAM Journal on Mathematical Analysis, 12(4):489–501, July 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bragg:1981:VLF

[Bra81b] L. R. Bragg. Vector lifting and factorable differential operators. SIAM Journal on Mathematical Analysis, 12(5):722–730, September 1981. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Braun:1981:ICH

[Bra81c] Martin Braun. Invariant curves, homoclinic points, and ergodicity in area-preserving mappings. SIAM Journal on Mathematical Analysis, 12(4):630–638, July 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brauer:1983:CRH

[Bra83] Fred Brauer. Constant-rate harvesting of age-structured populations. SIAM Journal on Mathematical Analysis, 14(5):947–961, September 1983. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bragg:1986:QIP

[Bra86] L. R. Bragg. Quasi-inner products of analytic functions with applications to special functions. SIAM Journal on Mathematical Analysis, 17(1):220–230, January 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brandon:1991:GEA

[Bra91] Deborah Brandon. Global existence and asymptotic stability for a nonlinear integro differential equation modeling heat flow. SIAM Journal on Mathematical Analysis, 22(1):72–106, January 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Braides:1995:LSC

[Bra95] Andrea Braides. Lower semicontinuity conditions for functionals on jumps and creases. SIAM Journal on Mathematical Analysis, 26(5):1184–1198, September 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/24267.

Bucy:1972:NDE

[BRC72] R. S. Bucy and J. Rodriguez-Canabal. A negative definite equilibrium and its induced cone of global existence for the Riccati equation. SIAM Journal on Mathematical Analysis, 3(4): 644–646, November 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bresters:1973:EEP

[Bre73] D. W. Bresters. On the equation of Euler–Poisson–Darboux. SIAM Journal on Mathematical Analysis, 4(1):31–41, February 1973. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bresters:1978:GEP

[Bre78] D. W. Bresters. On a generalized Euler–Poisson–Darboux equation. SIAM Journal on Mathematical Analysis, 9(5):924–934, October 1978. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bressoud:1981:LRF

[Bre81] D. M. Bressoud. Linearization and related formulas for q-ultraspherical polynomials. SIAM Journal on Mathematical Analysis, 12(2):161–168, March 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brewer:1982:DRP

[Bre82] Dennis W. Brewer. The differentiability with respect to a parameter of the solution of a linear abstract Cauchy problem. SIAM Journal on Mathematical Analysis, 13(4):607–620, July 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brenner:1985:NSP

[Bre85] J. L. Brenner. A new set of polynomial means related to the means of Maclaurin. SIAM Journal on Mathematical Analysis, 16 (6):1309–1316, November 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bressan:1991:BAR

[Bre91] Alberto Bressan. Blowup asymptotics for the reactive-Euler gas model. SIAM Journal on Mathematical Analysis, 22(3):587–601, May 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brewster:1993:TPC

[Bre93] M. E. Brewster. Thin plates and compressive membrane solutions II: A nonexistence result. SIAM Journal on Mathematical Analysis, 24(3):634–647, May 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bridges:1988:SIR

[Bri88] Thomas J. Bridges. Strong internal resonance, Z_{2} ⊕ Z_{2} symmetry, and multiple periodic solutions. SIAM Journal on Mathematical Analysis, 19(5):1015– 1031, September 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bronikowski:1975:DNI

[Bro75a] T. A. Bronikowski. Dependence of a nonlinear integrodifferential system on parameters. SIAM Journal on Mathematical Analysis, 6(1):153–167, February 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brown:1975:DTO

[Bro75b] R. C. Brown. Duality theory for n th order differential operators under Stieltjes boundary conditions. SIAM Journal on Mathematical Analysis, 6(5): 882–900, October 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brown:1978:IOF

[Bro78] Patrick M. Brown. Integral operators for fourth order linear parabolic equations. SIAM Journal on Mathematical Analysis, 9 (1):121–141, February 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brown:1983:DUS

[Bro83] Peter N. Brown. Decay to uniform states in competitive systems. SIAM Journal on Mathematical Analysis, 14(4):659–673, July 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brown:1996:GUI

[Bro96]  Russell M. Brown.  Global

uniqueness in the impedanceimaging problem for less regular conductivities. SIAM Journal on Mathematical Analysis, 27 (4):1049–1056, July 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/27113.

Brynielsson:1974:FIE

[Bry74] Lennart Brynielsson. On Fredholm integral equations of the first kind with convex constraints. SIAM Journal on Mathematical Analysis, 5(6):955–962, November 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bryan:1986:GIU

[Bry86] Robert Neff Bryan. The generalized inverse of an unbounded linear operator. SIAM Journal on Mathematical Analysis, 17(1): 128–131, January 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Billard:1974:SCP

[BS74] L. Billard and Norman C. Severo. Solution of a class of partial differential equations with initial conditions. SIAM Journal on Mathematical Analysis, 5(6): 918–919, November 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Buckholtz:1975:GCC

[BS75] J. D. Buckholtz and J. K. Shaw. Generalized completely convex functions and Sturm– Liouville operators. SIAM Journal on Mathematical Analysis, 6 (5):812–828, October 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Braaksma:1976:SCW

[BS76] B. L. J. Braaksma and A. Schuitman. Some classes of Watson transforms and related integral equations for generalized functions. SIAM Journal on Mathematical Analysis, 7(6):771–798, November 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Blankenship:1979:SPL

[BS79a] G. Blankenship and S. Sachs.

Singularly perturbed linear stochastic ordinary differential equations. SIAM Journal on Mathematical Analysis, 10(2): 306–320, March 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bustoz:1979:IUL

[BS79b] J. Bustoz and N. Savage. Inequalities for ultraspherical and Laguerre polynomials. SIAM Journal on Mathematical Analysis, 10(5):902–912, September 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bustoz:1980:IUL

[BS80] J. Bustoz and N. Savage. Inequalities for ultraspherical and Laguerre polynomials. II. SIAM Journal on Mathematical Analysis, 11(5):876–884, September 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bamberger:1983:NRV

[BS83a] A. Bamberger and M. Schatzman. New results on the vibrating string with a continuous obstacle. SIAM Journal on Mathematical Analysis, 14(3):560–595, May 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bandle:1983:QBB

[BS83b] Catherine Bandle and René P. Sperb. Qualitative behavior and bounds in a nonlinear plasma problem. SIAM Journal on Mathematical Analysis, 14(1): 142–151, January 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bohm:1985:DFM

[BS85a] Michael Böhm and R. E. Showalter. Diffusion in fissured media. SIAM Journal on Mathematical Analysis, 16(3):500–509, May 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bohm:1985:NPD

[BS85b] Michael Böhm and R. E. Showalter. A nonlinear pseudoparabolic diffusion equation. SIAM Journal on Mathematical Analysis, 16(5):980–999, September 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Beyer:1992:HP

[BS92] W. A. Beyer and Blair Swartz. Halfway points. SIAM Journal on Mathematical Analysis, 23(5): 1332–1341, September 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bao:1995:TLT

[BS95] Gang Bao and William W.

Symes. Time-like trace regularity of the wave equation with a nonsmooth principal part. SIAM Journal on Mathematical Analysis, 26(1):129–146, January 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/24426.

Bottcher:1996:ITH

[BS96] Albrecht Böttcher and Bernd Silbermann. Infinite Toeplitz and Hankel matrices with operatorvalued entries. SIAM Journal on Mathematical Analysis, 27(3): 805–822, May 1996. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bronsard:1997:VPM

[BS97] Lia Bronsard and Barbara Stoth.

Volume-preserving mean curvature flow as a limit of a nonlocal Ginzburg–Landau equation. SIAM Journal on Mathematical Analysis, 28(4):769–807, July 1997. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/27927.

Brauner:1987:ESC

[BSL87] C. M. Brauner and Cl. Schmidt-Lainé. Existence of a solution to a certain plane premixed flame problem with two-step kinetics. SIAM Journal on Mathematical Analysis, 18(5):1406– 1415, September 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bonet:1998:AIH

[BSSV98] Carles Bonet, David Sauzin, Tere Seara, and Marta Valencia. Adiabatic invariant of the harmonic oscillator, complex matching and resurgence. SIAM Journal on Mathematical Analysis, 29(6): 1335–1360, November 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32151.

Berens:1995:TDD

[BSX95] H. Berens, H. J. Schmid, and Y. Xu. On two-dimensional definite orthogonal systems and a lower bound for the number of nodes of associated cubature formulae. SIAM Journal on Mathematical Analysis, 26 (2):468–487, March 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/22986.

Baird:1977:PBF

[BSZ77] L. C. Baird, S. Sancaktar, and P. F. Zweifel. The Poincaré– Bertrand formula: a derivation and a generalization. SIAM Journal on Mathematical Analysis, 8(4):580–591, August 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bebernes:1987:NKP

[BT87a] J. Bebernes and W. Troy. Nonexistence for the Kassoy problem. SIAM Journal on Mathematical Analysis, 18(4):1157– 1162, July 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Benaouda:1987:SPN

[BT87b] Ahmed Benaouda and Monique Madaune Tort. Singular perturbations for nonlinear hyperbolic-parabolic problems. SIAM Journal on Mathematical Analysis, 18(1): 137–148, January 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bartolomeo:1991:UED

[BT91a] J. Bartolomeo and R. Triggiani.

Uniform energy decay rates for Euler–Bernoulli equations with feedback operators in the Dirichlet/Neumann boundary conditions. SIAM Journal on Mathematical Analysis, 22(1):46–71, January 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Brown:1991:BRS

[BT91b] K. J. Brown and A. Tertikas.

On the bifurcation of radially symmetric steady-state solutions arising in population genetics. SIAM Journal on Mathematical Analysis, 22(2):400–413, March 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bernier:1996:WSI

[BT96] David Bernier and Keith F. Taylor. Wavelets from squareintegrable representations. SIAM Journal on Mathematical Analysis, 27(2):594–608, March 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/25626.

Barcenas:1997:MMN

[BU97] Diómedes Bárcenas and Wilfredo Urbina. Measurable multifunctions in nonseparable Banach spaces. SIAM Journal on Mathematical Analysis, 28(5):1212– 1226, September 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/29600.

Budd:1989:AST

[Bud89] C. J. Budd. Applications of Shil′nikov’s theory to semilinear elliptic equations. SIAM Journal on Mathematical Analysis, 20(5): 1069–1080, September 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Buhring:1987:ACH

[Büh87a] Wolfgang Bühring. An analytic continuation of the hypergeometric series. SIAM Journal on Mathematical Analysis, 18(3): 884–889, May 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Buhring:1987:BUA

[Büh87b] Wolfgang Bühring. The behavior at unit argument of the hypergeometric function _{3}F_{2}. SIAM Journal on Mathematical Analysis, 18(5):1227–1234, September 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Buhring:1988:ACF

[Büh88] Wolfgang Bühring. An analytic continuation formula for the generalized hypergeometric function. SIAM Journal on Mathematical Analysis, 19(5):1249– 1251, September 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Buschman:1972:DIO

[Bus72] R. G. Buschman. Decomposition of an integral operator by use of mikusi ski calculus. SIAM Journal on Mathematical Analysis, 3 (1):83–85, February 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bustoz:1977:ZCJ

[Bus77] J. Bustoz. On the zeros of certain Jacobi polynomial sums. SIAM Journal on Mathematical Analysis, 8(2):202–205, April 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bushard:1978:CUB

[Bus78] Louis B. Bushard. Computationally useful bounds for singular nonlinear second order boundary value problems. SIAM Journal on Mathematical Analysis, 9 (4):651–659, August 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Butler:1979:IPS

[But79] John B. Butler, Jr. On the inverse problem for selfadjoint operators defined on certain rigged Hilbert spaces. SIAM Journal on Mathematical Analysis, 10 (2):281–291, March 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Butler:1980:IAO

[But80] G. J. Butler. Integral averages and the oscillation of second order ordinary differential equations. SIAM Journal on Mathematical Analysis, 11(1):190–200, January 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Buzano:1986:SBT

[Buz86] Ernesto Buzano. Secondary bifurcations of a thin rod under axial compression. SIAM Journal on Mathematical Analysis, 17 (2):312–321, March 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bryan:1992:URC

[BV92] Kurt Bryan and Michael Vogelius. A uniqueness result concerning the identification of a collection of cracks from finitely many electrostatic boundary measurements. SIAM Journal on Mathematical Analysis, 23(4): 950–958, July 1992. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Beretta:1998:SDB

[BV98] E. Beretta and S. Vessella. Stable determination of boundaries from Cauchy data. SIAM Journal on Mathematical Analysis, 30(1):220–232, 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32573.

Bidaut-Veron:1994:CSS

[BVB94] Marie-Françoise Bidaut-Veron and Mustapha Bouhar. On characterization of solutions of some nonlinear differential equations and applications. SIAM Journal on Mathematical Analysis, 25(3):859–875, May 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23059.

Bart:1973:LIE

[BW73] G. R. Bart and R. L. Warnock.

Linear integral equations of the third kind. SIAM Journal on Mathematical Analysis, 4(4): 609–622, November 1973. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bleecker:1980:SSR

[BW80] David D. Bleecker and Leslie C. Wilson. Splitting the spectrum of a Riemannian manifold. SIAM Journal on Mathematical Analysis, 11(5):813–818, September 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Bona:1983:KVE

[BW83] Jerry Bona and Ragnar Winther. The Korteweg–de Vries equation, posed in a quarter-plane. SIAM Journal on Mathematical Analysis, 14(6):1056–1106, November 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Belogay:1999:ASO

[BW99] Eugene Belogay and Yang Wang. Arbitrarily smooth orthogonal nonseparable wavelets in R^{2}. SIAM Journal on Mathematical Analysis, 30(3):678–697, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/32773.

Bernfeld:1972:BOS

[BY72] Stephen R. Bernfeld and James A. Yorke. The behavior of oscillatory solutions of x′′(t) + p(t)g(x(t)) = 0. SIAM Journal on Mathematical Analysis, 3(4): 654–667, November 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Binding:1991:VPD

[BY91] Paul Binding and Qiang Ye.

Variational principles without definiteness conditions. SIAM Journal on Mathematical Analysis, 22(6):1575–1583, September 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chak:1971:ECP

[CA71] A. M. Chak and A. K. Agarwal. An extension of a class of polynomials. II. SIAM Journal on Mathematical Analysis, 2 (3):352–355, August 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Campbell:1977:LSD

[Cam77] Stephen L. Campbell. Linear systems of differential equations with singular coefficients. SIAM Journal on Mathematical Analysis, 8(6):1057–1066, November 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Campbell:1987:GFS

[Cam87a] Stephen L. Campbell. A general form for solvable linear time varying singular systems of differential equations. SIAM Journal on Mathematical Analysis, 18 (4):1101–1115, July 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Campos:1987:SPZ

[Cam87b] R. G. Campos. Some properties of the zeros of polynomial solutions of Sturm–Liouville equations. SIAM Journal on Mathematical Analysis, 18(6):1664– 1668, November 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Campos:1989:GML

[Cam89a] L. M. B. C. Campos. On a generalized Mittag-Leffler theorem and implicit differintegration. SIAM Journal on Mathematical Analysis, 20(2):454–467, March 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Campos:1989:BOA

[Cam89b] L. M. B. C. Campos. On the branchpoint operator and the annihilation of differintegrations. SIAM Journal on Mathematical Analysis, 20(2):439–453, March 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Canadas:1985:SPS

[Can85] Guy Canadas. Speed of propagation of solutions of a linear integro-differential equation with nonconstant coefficients. SIAM Journal on Mathematical Analysis, 16(1):143–147, January 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cao:1995:MSO

[Cao95] Yulin Cao. Multiexistence of slowly oscillating periodic solutions for differential delay equations. SIAM Journal on Mathematical Analysis, 26(2):436–445, March 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlitz:1970:NCT

[Car70a] L. Carlitz. Note on certain triangular arrays. SIAM Journal on Mathematical Analysis, 1 (3):328–332, August 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlitz:1970:SRF

[Car70b] L. Carlitz. Some reduction formulas for generalized hypergeometric functions. SIAM Journal on Mathematical Analysis, 1(2): 243–245, May 1970. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlson:1970:SEL

[Car70c] B. C. Carlson. Some extensions of Lardner’s relations between 0F_{3} and Bessel functions. SIAM Journal on Mathematical Analysis, 1(2):232–242, May 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carasso:1971:ABB

[Car71a] Alfred Carasso. The abstract backward beam equation. SIAM Journal on Mathematical Analysis, 2(2):193–212, May 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlson:1971:AFM

[Car71b] B. C. Carlson. Appell functions and multiple averages. SIAM Journal on Mathematical Analysis, 2(3):420–430, August 1971. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlson:1971:NPA

[Car71c] B. C. Carlson. New proof of the addition theorem for Gegenbauer polynomials. SIAM Journal on Mathematical Analysis, 2 (3):347–351, August 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlitz:1972:EEC

[Car72] L. Carlitz. Explicit evaluation of certain polynomials. SIAM Journal on Mathematical Analysis, 3(2):352–357, May 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlitz:1973:SAC

[Car73a] L. Carlitz. Some q-analogues of certain combinatorial numbers. SIAM Journal on Mathematical Analysis, 4(3):433–446, May 1973. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carmichael:1973:GCP

[Car73b] Richard D. Carmichael. Generalized Cauchy and Poisson integrals and distributional boundary values. SIAM Journal on Mathematical Analysis, 4(1): 198–219, February 1973. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlson:1974:ACE

[Car74a] B. C. Carlson. Analytic continuation of the Euler transform. SIAM Journal on Mathematical Analysis, 5(2):252–255, April 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlson:1974:EAF

[Car74b] B. C. Carlson. Expansion of analytic functions in Jacobi series. SIAM Journal on Mathematical Analysis, 5(5):797–808, October 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlson:1974:IJP

[Car74c] B. C. Carlson. Inequalities for Jacobi polynomials and Dirichlet averages. SIAM Journal on Mathematical Analysis, 5(4): 586–596, August 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carmichael:1974:ARD

[Car74d] Richard D. Carmichael. Analytic representation of the distributional finite Fourier transform. SIAM Journal on Mathematical Analysis, 5(5):737–761, October 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlitz:1975:NBI

[Car75a] L. Carlitz. Note on a binomial identity. SIAM Journal on Mathematical Analysis, 6(5):

901–905, October 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlson:1975:AFD

[Car75b] B. C. Carlson. Appell’s function F_{4} as a double average. SIAM Journal on Mathematical Analysis, 6(6):960–965, November 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carasso:1976:EBF

[Car76a] Alfred Carasso. Error bounds in the final value problem for the heat equation. SIAM Journal on Mathematical Analysis, 7 (2):195–199, April 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlson:1976:QTA

[Car76b] B. C. Carlson. Quadratic transformations of Appell functions. SIAM Journal on Mathematical Analysis, 7(2):291–304, April 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlitz:1977:CGF

[Car77a] L. Carlitz. A class of generating functions. SIAM Journal on Mathematical Analysis, 8(3): 518–532, May 1977. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlitz:1977:NSC

[Car77b] L. Carlitz. Note on some convolved power sums. SIAM Journal on Mathematical Analysis, 8 (4):701–709, August 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlitz:1977:SEC

[Car77c] L. Carlitz. Some expansions and convolution formulas related to MacMahon’s master theorem. SIAM Journal on Mathematical Analysis, 8(2):320–336, April 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlson:1977:EIF

[Car77d] B. C. Carlson. Elliptic integrals of the first kind. SIAM Journal on Mathematical Analysis, 8 (2):231–242, April 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlitz:1978:SCR

[Car78a] L. Carlitz. Solution of certain recurrences. II. SIAM Journal on Mathematical Analysis, 9(5): 781–786, October 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlitz:1978:SCI

[Car78b] L. Carlitz. Some combinatorial identities of Bernstein. SIAM Journal on Mathematical Analysis, 9(1):65–75, February 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlson:1978:SPT

[Car78c] B. C. Carlson. Short proofs of three theorems on elliptic integrals. SIAM Journal on Mathematical Analysis, 9(3):524–528, June 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlitz:1979:SMP

[Car79a] L. Carlitz. Some multiple power series with zero-one coefficients. SIAM Journal on Mathematical Analysis, 10(5):1022–1040, September 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carmichael:1979:AFR

[Car79b] Richard D. Carmichael. Analytic functions related to the distributions of exponential growth. SIAM Journal on Mathematical Analysis, 10(5):1041–1068, September 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum and addendum [Car86].

Carlson:1980:LTP

[Car80] B. C. Carlson. The Laplace transform of a product of Bessel functions. SIAM Journal on Mathematical Analysis, 11(3): 428–435, May 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carmichael:1986:EAA

[Car86] Richard D. Carmichael. Erratum and addendum: “Analytic functions related to the distributions of exponential growth” [SIAM J. Math. Anal. 10 (1979), no. 5, 1041–1068; MR 82d:46059]. SIAM Journal on Mathematical Analysis, 17(5):1287–1288, September 1986. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [Car79b].

Carlson:1987:DA

[Car87] B. C. Carlson. Dirichlet averages of x^{t} log x. SIAM Journal on Mathematical Analysis, 18 (2):550–565, March 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carstensen:1994:IPV

[Car94] Carsten Carstensen. Interface problems in viscoplasticity and plasticity. SIAM Journal on Mathematical Analysis, 25(6): 1468–1487, November 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23658.

Carlson:1995:HPM

[Car95a] Robert Carlson. Hearing point masses in a string. SIAM Journal on Mathematical Analysis, 26 (3):583–600, May 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24428.

Carr:1995:GES

[Car95b] Danielle D. Carr. Global existence of solutions to reactionhyperbolic systems in one space dimension. SIAM Journal on Mathematical Analysis, 26(2): 399–414, March 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/24805.

Carpio:1996:LTB

[Car96] Ana Carpio. Large-time behavior in incompressible Navier– Stokes equations. SIAM Journal on Mathematical Analysis, 27 (2):449–475, March 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/25678.

Carasso:1997:EBN

[Car97] Alfred S. Carasso. Error bounds in nonsmooth image deblurring. SIAM Journal on Mathematical Analysis, 28(3):656–668, May 1997. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/29021.

Carasso:1999:LCS

[Car99a] Alfred S. Carasso. Logarithmic convexity and the “Slow evolution” constraint in ill-posed initial value problems. SIAM Journal on Mathematical Analysis, 30(3):479–496, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33236.

Carlson:1999:LET

[Car99b] Robert Carlson. Large eigenvalues and trace formulas for matrix Sturm–Liouville problems. SIAM Journal on Mathematical Analysis, 30(5):949–962, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/34041.

Clutton-Brock:1987:GFL

[CB87] M. Clutton-Brock. Generalized Fejér and Lánczos kernels. SIAM Journal on Mathematical Analysis, 18(1):259–272, January 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:1989:GSL

[CB89] Pei-Li Chen and Jonathan Bell.

Global solutions and long-time behavior to a Fitzhugh–Nagumo myelinated axon model. SIAM Journal on Mathematical Analysis, 20(3):567–581, May 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cantrell:1991:DLE

[CC91] Robert Stephen Cantrell and Chris Cosner. Diffusive logistic equations with indefinite weights: Population models in disrupted environments II. SIAM Journal on Mathematical Analysis, 22(4):1043–1064, July 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Castro:1994:MSN

[CC94] Alfonso Castro and Jorge Cossio.

Multiple solutions for a nonlinear Dirichlet problem. SIAM Journal on Mathematical Analysis, 25(6):

1554–1561, November 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23010.

Costin:1996:RWF

[CC96] Ovidiu Costin and Rodica Costin. Rigorous WKB for finite-order linear recurrence relations with smooth coefficients. SIAM Journal on Mathematical Analysis, 27(1):110–134, January 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/24803.

Casazza:1998:FCR

[CC98] Peter G. Casazza and Ole Christensen. Frames containing a Riesz basis and preservation of this property under perturbations. SIAM Journal on Mathematical Analysis, 29(1):266–278, January 1998. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

http://epubs.siam.org/sam-
bin/dbq/article/29425.

Colombo:1999:CDC

[CC99] Rinaldo M. Colombo and Andrea Corli. Continuous dependence in conservation laws with phase transitions. SIAM Journal on Mathematical Analysis, 31(1):34–62, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/33187.

Chow:1974:LUT

[CD74] Shui-Nee Chow and D. R. Dunninger. Local uniqueness theorems for a class of higher order elliptic equations. SIAM Journal on Mathematical Analysis, 5 (2):224–229, April 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cannon:1980:EWS

[CD80] J. R. Cannon and Emmanuele DiBenedetto. On the existence of weak-solutions to an n-dimensional Stefan problem with nonlinear boundary conditions. SIAM Journal on Mathematical Analysis, 11(4):632–645, July 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Crouzeix:1990:BEP

[CD90] Michel Crouzeix and Jean Descloux. A bidimensional electromagnetic problem. SIAM Journal on Mathematical Analysis, 21 (3):577–592, May 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chu:1991:CSN

[CD91] Moody T. Chu and Kenneth R.

Driessel. Constructing symmetric nonnegative matrices with prescribed eigenvalues by differential equations. SIAM Journal on Mathematical Analysis, 22(5): 1372–1387, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cohen:1993:IAB

[CD93a] A. Cohen and I. Daubechies. On the instability of arbitrary biorthogonal wavelet packets. SIAM Journal on Mathematical Analysis, 24(5):1340–1354, September 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cohen:1993:OBC

[CD93b] A. Cohen and Ingrid Daubechies. Orthonormal bases of compactly supported wavelets III. better frequency resolution. SIAM Journal on Mathematical Analysis, 24(2):520–527, March 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Croc:1995:ASB

[CD95] Élisabeth Croc and Yves Dermenjian. Analyse spectrale d’une bande acoustique multistratifiée. I. Principe d’absorption limite pour une stratification simple. (French) [Spectral analysis of a multistratified acoustic strip. I. Limiting absorption principle for a simple stratification]. SIAM Journal on Mathematical Analysis, 26(4): 880–924, July 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24855.

Canada:1996:SPN

[CD96a] A. Cañada and P. Drábek. On semilinear problems with nonlinearities depending only on derivatives. SIAM Journal on Mathematical Analysis, 27(2): 543–557, March 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/26108.

Cohen:1996:NSS

[CD96b] Albert Cohen and Nira Dyn.

Nonstationary subdivision schemes and multiresolution analysis. SIAM Journal on Mathematical Analysis, 27(6):1745–1769, November 1996. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

http://epubs.siam.org/sam-
bin/dbq/article/27429.

Croc:1996:SAM

[CD96c] Elisabeth Croc and Yves Dermenjian. Spectral analysis of a multistratified acoustic strip. Part II: Asymptotic behavior of solutions for a simple stratification. SIAM Journal on Mathematical Analysis, 27(6):1631– 1652, November 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/27123.

Christopher:1997:ICP

[CD97] C. J. Christopher and J. Devlin.

Isochronous centers in planar polynomial systems. SIAM Journal on Mathematical Analysis, 28 (1):162–177, January 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/25924.

Charrier:1988:ETS

[CDHL88] P. Charrier, B. Dacorogna, B. Hanouzet, and P. Laborde. An existence theorem for slightly compressible materials in nonlinear elasticity. SIAM Journal on Mathematical Analysis, 19 (1):70–85, January 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cavaretta:1981:SCS

[CDMS81] A. S. Cavaretta, Jr., W. Dahmen, C. A. Micchelli, and P. W. Smith. On the solvability of certain systems of linear difference equations. SIAM Journal on Mathematical Analysis, 12(6):833–841, November 1981. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chow:1990:BHP

[CDT90] S.-N. Chow, B. Deng, and D. Terman. The bifurcation of homoclinic and periodic orbits from two heteroclinic orbits. SIAM Journal on Mathematical Analysis, 21(1):179–204, January 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cortazar:1985:ABS

[CE85] Carmen Cortazar and Manuel Elgueta. The asymptotic behaviour of the solution of a nonlinear diffusion equation. SIAM Journal on Mathematical Analysis, 16(2):251–258, March 1985.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cortazar:1987:EGS

[CE87] Carmen Cortazar and Manuel Elgueta. Existence of generalized solutions of a nonlinear diffusion Cauchy problem. SIAM Journal on Mathematical Analysis, 18(5): 1385–1393, September 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:1989:ONS

[CE89] Shao Zhu Chen and Lynn H.

Erbe. Oscillation and nonoscillation for systems of selfadjoint second-order difference equations. SIAM Journal on Mathematical Analysis, 20(4):939–949, July 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:1999:MAR

[CE99] Zhangxin Chen and Richard Ewing. Mathematical analysis for reservoir models. SIAM Journal on Mathematical Analysis, 30(2):431–453, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/31915.

Caffarelli:1985:NEP

[CF85] Luis A. Caffarelli and Avner Friedman. A nonlinear evolution problem associated with an electropaint process. SIAM Journal on Mathematical Analysis, 16(5):

955–969, September 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Concus:1996:CWR

[CF96] Paul Concus and Robert Finn. Capillary wedges revisited. SIAM Journal on Mathematical Analysis, 27(1):56–69, January 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/24748.

Cordoba:1998:CDD

[CF98] Antonio Córdoba and Pablo Fernández. Convergence and divergence of decreasing rearranged Fourier series. SIAM Journal on Mathematical Analysis, 29(5):1129–1139, September 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/32070.

Caffarelli:1981:FBP

[CFV81] Luis A. Caffarelli, Avner Friedman, and Augusto Visintin. A free boundary problem describing transition in a superconductor. SIAM Journal on Mathematical Analysis, 12(5):679–690, September 1981. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colton:1971:IOA

[CG71] David Colton and Robert P. Gilbert. An integral operator approach to Cauchy’s problem for Δ_{p+2}u(x)+F (x)u(x) = 0. SIAM Journal on Mathematical Analysis, 2(1):113–132, February 1971. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlson:1983:TPM

[CG83a] B. C. Carlson and John L.

Gustafson. Total positivity of mean values and hypergeometric functions. SIAM Journal on Mathematical Analysis, 14 (2):389–395, March 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chapin:1983:RCS

[CG83b] Jed Chapin and John L. B.

Gamlen. On the rate of convergence of stopped random walks. SIAM Journal on Mathematical Analysis, 14(5):962–973, September 1983. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlson:1985:AEF

[CG85] B. C. Carlson and John L.

Gustafson. Asymptotic expansion of the first elliptic integral. SIAM Journal on Mathematical Analysis, 16(5):1072– 1092, September 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chicone:1987:NID

[CG87] Carmen Chicone and Jürgen Gerlach. A note on the identifiability of distributed parameters in elliptic equations. SIAM Journal on Mathematical Analysis, 18(5):1378–1384, September 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chossat:1988:IMS

[CG88] Pascal Chossat and Martin Golubitsky. Iterates of maps with symmetry. SIAM Journal on Mathematical Analysis, 19(6): 1259–1270, November 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlson:1994:AAS

[CG94] B. C. Carlson and J. L. Gustafson. Asymptotic approximations for symmetric elliptic integrals. SIAM Journal on Mathematical Analysis, 25(2):288–303, March 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/22847.

Chvala:1993:SLB

[CGP93] František Chvála, Tommy Gustafsson, and Rolf Pettersson. On solutions to the linear Boltzmann equation with external electromagnetic force. SIAM Journal on Mathematical Analysis, 24(3): 583–602, May 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Crouzeix:1988:SRA

[CGR88] Michel Crouzeix, Giuseppe Geymonat, and Geneviève Raugel. Some remarks about the Morse lemma in infinite dimension.

SIAM Journal on Mathematical Analysis, 19(2):358–371, March 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [GM83, BMS83].

Carr:1979:NID

[CH79] Ralph W. Carr and Kenneth B.

Hannsgen. A nonhomogeneous integro-differential equation in Hilbert space. SIAM Journal on Mathematical Analysis, 10(5): 961–984, September 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carr:1982:RFV

[CH82] Ralph W. Carr and Kenneth B.

Hannsgen. Resolvent formulas for a Volterra equation in Hilbert space. SIAM Journal on Mathematical Analysis, 13(3):459–483, May 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colombeau:1992:NPB

[CH92] Jean François Colombeau and Arnaud Heibig. Nonconservative products in bounded variation functions. SIAM Journal on Mathematical Analysis, 23(4): 941–949, July 1992. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colton:1993:MFF

[CH93] David Colton and Peter Hähner.

Modified far field operators in inverse scattering theory. SIAM Journal on Mathematical Analysis, 24(2):365–389, March 1993.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chan:1970:CDD

[Cha70] C. Y. Chan. Continuous dependence on the data for a Stefan problem. SIAM Journal on Mathematical Analysis, 1(2): 282–287, May 1970. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chang:1972:SPG

[Cha72] K. W. Chang. Singular perturbations of a general boundary value problem. SIAM Journal on Mathematical Analysis, 3 (3):520–526, August 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chambers:1976:NRR

[Cha76] Ll. G. Chambers. Note on reciprocity relations associated with a linear hyperbolic, partial differential equation. SIAM Journal on Mathematical Analysis, 7(6): 834–836, November 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chambers:1978:ACH

[Cha78] Barbara Fromm Chambers. Analytic continuation via Hadamard’s product. SIAM Journal on Mathematical Analysis, 9(6): 1096–1104, December 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chalkley:1979:ASA

[Cha79] Roger Chalkley. Analytic solutions of algebraic differential equations. SIAM Journal on Mathematical Analysis, 10(4): 778–782, July 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cheng:1981:ULC

[Che81] Kuo Shung Cheng. Uniqueness of a limit cycle for a predatorprey system. SIAM Journal on Mathematical Analysis, 12(4): 541–548, July 1981. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chemin:1992:REG

[Che92a] J. Y. Chemin. Remarques sur l’existence globale pour le système de Navier–Stokes incompressible. (French) [Remarks on global existence for the incompressible Navier–Stokes system]. SIAM Journal on Mathematical Analysis, 23(1):20–28, January 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:1992:GSC

[Che92b] Gui Qiang Chen. Global solutions to the compressible Navier– Stokes equations for a reacting mixture. SIAM Journal on Mathematical Analysis, 23(3): 609–634, May 1992. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:1992:QAM

[Che92c] Pei-Li Chen. Qualitative analysis of myelinated nerve fibers with point-node FitzHugh–Nagumo dynamic system. SIAM Journal on Mathematical Analysis, 23 (1):81–98, January 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:1995:SWS

[Che95] Debao Chen. Spline wavelets of small support. SIAM Journal on Mathematical Analysis, 26 (2):500–517, March 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/24526.

Chen:1996:LEP

[Che96]  Xinfu Chen.  Lorenz equa-

tions. Part I: Existence and nonexistence of homoclinic orbits. SIAM Journal on Mathematical Analysis, 27(4):1057– 1069, July 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/26441.

Chitwood:1973:GGM

[Chi73] Howard Chitwood. Generalized Green’s matrices for linear differential systems. SIAM Journal on Mathematical Analysis, 4 (1):104–110, February 1973. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ching:1974:RWE

[Chi74a] Chin-Hung Ching. The reduced wave equation with a different radiation condition. SIAM Journal on Mathematical Analysis, 5 (5):723–728, October 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ching:1974:AFA

[Chi74b] Chin’Hung Ching. On averages of a function and its application to boundary value problems with insufficient data. SIAM Journal on Mathematical Analysis, 5(6): 941–947, November 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chiou:1977:GIS

[Chi77] Kuo-Liang Chiou. The geometry of indefinite J-spaces and strong stability criteria of canonical differential equations with periodic coefficients. SIAM Journal on Mathematical Analysis, 8 (1):118–126, February 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chihara:1980:OPW

[Chi80]  T. S. Chihara.  Orthogonal

polynomials whose distribution functions have finite point spectra. SIAM Journal on Mathematical Analysis, 11(2):358–364, March 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chiti:1983:BRF

[Chi83] Giuseppe Chiti. A bound for the ratio of the first two eigenvalues of a membrane. SIAM Journal on Mathematical Analysis, 14(6): 1163–1167, November 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chihara:1987:ZAW

[Chi87] Laura Chihara. On the zeros of the Askey–Wilson polynomials, with applications to coding theory. SIAM Journal on Mathematical Analysis, 18(1):191–207, January 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chicone:1992:BNO

[Chi92] Carmen Chicone. Bifurcations of nonlinear oscillations and frequency entrainment near resonance. SIAM Journal on Mathematical Analysis, 23(6):1577– 1608, November 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chicone:1995:PSS

[Chi95] Carmen Chicone. Periodic solutions of a system of coupled oscillators near resonance. SIAM Journal on Mathematical Analysis, 26(5):1257–1283, September 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/24353.

Choudhury:1976:PCD

[Cho76] A. K. Choudhury. On the pointwise completeness of delaydifferential systems of neutral type. SIAM Journal on Mathematical Analysis, 7(6):913–929, November 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cholewinski:1984:GFS

[Cho84] Frank M. Cholewinski. Generalized Fock spaces and associated operators. SIAM Journal on Mathematical Analysis, 15(1): 177–202, January 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Christiansen:1980:LAP

[Chr80] Edmund Christiansen. Limit analysis for plastic plates. SIAM Journal on Mathematical Analysis, 11(3):514–522, May 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chukwu:1976:BSS

[Chu76] E. N. Chukwu. On the boundedness and stability of solutions of some differential equations of the fifth order. SIAM Journal on Mathematical Analysis, 7 (2):176–194, April 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chukwu:1980:NFD

[Chu80]  E. N. Chukwu.  A neutral

functional differential equation of Lurie type. SIAM Journal on Mathematical Analysis, 11(1): 108–114, January 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chu:1984:GCT

[Chu84] Moody T. Chu. On the global convergence of the Toda lattice for real normal matrices and its applications to the eigenvalue problem. SIAM Journal on Mathematical Analysis, 15 (1):98–104, January 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Charris:1987:SOP

[CI87] Jairo A. Charris and Mourad E. H. Ismail. On sieved orthogonal polynomials. V. Sieved Pollaczek polynomials. SIAM Journal on Mathematical Analysis, 18 (4):1177–1218, July 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:1991:AJP

[CI91] Li-Chen Chen and Mourad E. H. Ismail. On asymptotics of Jacobi polynomials. SIAM Journal on Mathematical Analysis, 22(5): 1442–1449, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cicogna:1987:SCS

[Cic87] G. Cicogna. Simple criteria for stable bifurcating periodic solutions of O.D.E.s. SIAM Journal on Mathematical Analysis, 18 (1):34–45, January 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ciesielski:1987:EFR

[Cie87] Z. Ciesielski. Explicit formula relating the Jacobi, Hahn and Bernstein polynomials. SIAM Journal on Mathematical Analysis, 18(6):1573–1575, November 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cigler:1979:NPC

[Cig79] J. Cigler. A note on a paper by L. Carlitz: “A class of generating functions” (SIAM J. Math. Anal. 8 (1977), no. 3, 518–532). SIAM Journal on Mathematical Analysis, 10(1):78–79, January 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Crates:1976:ITZ

[CJ76] Gladys Hayes Crates and John W.

Jayne. Inclusion theorems for the zeros of certain recursively generated polynomials. SIAM Journal on Mathematical Analysis, 7 (2):264–270, April 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cockburn:1988:MEP

[CJ88] Bernardo Cockburn and Patrick Joly. Maxwell equations in polarizable media. SIAM Journal on Mathematical Analysis, 19(6): 1372–1390, November 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Clemons:1993:GPK

[CJ93] C. B. Clemons and C. K. R. T.

Jones. A geometric proof of the Kwong–McLeod uniqueness result. SIAM Journal on Mathematical Analysis, 24(2):436–443, March 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Caifeng:1999:GSP

[CJ99] Tu Caifeng and Jiang Jifa. Global stability and permanence for a class of type K monotone systems. SIAM Journal on Mathematical Analysis, 30(2):360– 378, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/32529.

Chen:1998:SVG

[CJM98] Xu-Yan Chen, Shuichi Jimbo, and Yoshihisa Morita. Stabilization of vortices in the Ginzburg– Landau equation with a variable diffusion coefficient. SIAM Journal on Mathematical Analysis, 29 (4):903–912, July 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/30875.

Craven:1977:GFT

[CK77] B. D. Craven and J. J. Koliha. Generalizations of Farkas’ Theorem. SIAM Journal on Mathematical Analysis, 8(6):983–997, November 1977. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colton:1984:DSF

[CK84] David Colton and Andreas Kirsch. Dense sets and far field patterns in acoustic wave propagation. SIAM Journal on Mathematical Analysis, 15(5): 996–1006, September 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colton:1985:DSF

[CK85] David Colton and Rainer Kress.

Dense sets and far field patterns in electromagnetic wave propagation. SIAM Journal on Mathematical Analysis, 16(5):1049– 1060, September 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cohen:1987:SKV

[CK87] Amy Cohen and Thomas Kappeler. Solutions to the Korteweg– de Vries equation with initial profile in L_{1}^{1}(R) ∩ L_{N}^{1}(R^{+}). SIAM Journal on Mathematical Analysis, 18(4):991–1025, July 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chan:1989:QSS

[CK89] C. Y. Chan and Hans G. Kaper.

Quenching for semilinear singular parabolic problems. SIAM Journal on Mathematical Analysis, 20(3):558–566, May 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:1991:IDN

[CK91] Senhuei E. Chen and R. Bruce Kellogg. An interior discontinuity of a nonlinear elliptichyperbolic system. SIAM Journal on Mathematical Analysis, 22 (3):602–622, May 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cohen:1992:SCS

[CK92] Amy Cohen and Thomas Kappeler. Solutions to the cubic Schrödinger equation by the inverse scattering method. SIAM Journal on Mathematical Analysis, 23(4):900–922, July 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Choe:1995:RIE

[CK95a] Hi Jun Choe and Jongsik Kim. Regularity for the interfaces of evolutionary p-Laplacian functions. SIAM Journal on Mathematical Analysis, 26(4):791–819, July 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/24704.

Colton:1995:EFF

[CK95b] David Colton and Rainer Kress. Eigenvalues of the far field operator and inverse scattering theory. SIAM Journal on Mathematical Analysis, 26(3):601–615, May 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/24946.

Colton:1989:FFP

[CKP89] David Colton, Andreas Kirsch, and Lassi Päivärinta. Far-field patterns for acoustic waves in an inhomogeneous medium. SIAM Journal on Mathematical Analysis, 20(6):1472–1483, November 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cook:1971:BSD

[CL71] L. Pamela Cook and G. S. S.

Ludford. The behavior as ε → +0 of solutions to ε∇^{2}w = ∂w/∂y in |y| ≤ 1 for discontinuous boundary data. SIAM Journal on Mathematical Analysis, 2(4):567–594, November 1971. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cook:1973:BSR

[CL73] L. Pamela Cook and G. S. S.

Ludford. The behavior as ε → 0^{+} of solutions to ε∇^{2}w = (∂/∂y)w on the rectangle 0 ≤ x ≤ l,|y| ≤ 1. SIAM Journal on Mathematical Analysis, 4(1):161–184, February 1973. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chang:1981:QSS

[CL81] Peter H. Chang and Howard A.

Levine. The quenching of solutions of semilinear hyperbolic equations. SIAM Journal on Mathematical Analysis, 12(6): 893–903, November 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chipot:1986:EUS

[CL86] Michel Chipot and Mitchell Luskin. Existence and uniqueness of solutions to the compressible Reynolds lubrication equation. SIAM Journal on Mathematical Analysis, 17(6):1390– 1399, November 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Clements:1987:FAA

[CL87] John C. Clements and L. J. Leon. A fast, accurate algorithm for the isometric mapping of a developable surface. SIAM Journal on Mathematical Analysis, 18(4): 966–971, July 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chossat:1989:IAS

[CL89] Pascal Chossat and Reiner Lauterbach. The instability of axisymmetric solutions in problems with spherical symmetry. SIAM Journal on Mathematical Analysis, 20(1):31–38, January 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chipot:1990:ESE

[CL90] Michel Chipot and Mitchell Luskin. Existence of solutions to the elastohydrodynamical equations for magnetic recording systems. SIAM Journal on Mathematical Analysis, 21(1):1–17, January 1990. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Choe:1991:OPQ

[CL91] Hi Jun Choe and John L. Lewis. On the obstacle problem for quasilinear elliptic equations of p Laplacian type. SIAM Journal on Mathematical Analysis, 22(3): 623–638, May 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chui:1993:NWP

[CL93] Charles K. Chui and Chun Li. Nonorthogonal wavelet packets. SIAM Journal on Mathematical Analysis, 24(3):712–738, May 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chui:1996:DAD

[CL96] Charles K. Chui and Chun Li. Dyadic affine decompositions and functional wavelet transforms. SIAM Journal on Mathematical Analysis, 27(3):865–890, May 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Clausing:1983:TEM

[Cla83] A. Clausing. Type t entropy and majorization. SIAM Journal on Mathematical Analysis, 14 (1):203–208, January 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carton-Lebrun:1992:WFT

[CLH92] C. Carton-Lebrun and H. P.

Heinig. Weighted Fourier transform inequalities for radially decreasing functions. SIAM Journal on Mathematical Analysis, 23 (3):785–798, May 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cosner:1988:TES

[CLP88] Chris Cosner, Suzanne M. Lenhart, and Vladimir Protopopescu. Transport equations with second-order differential collision operators. SIAM Journal on Mathematical Analysis, 19(4):797–813, July 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chow:1989:SLA

[CLP89] Shui-Nee Chow, Xiao-Biao Lin, and Kenneth J. Palmer. A shadowing lemma with applications to semilinear parabolic equations. SIAM Journal on Mathematical Analysis, 20(3):547–557, May 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colombeau:1980:REI

[CM80] J. F. Colombeau and A. Meril. A remark on equations y′′ + p_{1}y′ + p_{2}y = 0 at an irregular singular point. SIAM Journal on Mathematical Analysis, 11(6): 1002–1010, November 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cox:1998:STC

[CM98] Steven J. Cox and C. Maeve McCarthy. The shape of the tallest column. SIAM Journal on Mathematical Analysis, 29(3): 547–554, May 1998. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

                                        http://epubs.siam.org/sam-

bin/dbq/article/31453. See

correction [CM00].

Cox:2000:STC

[CM00] Steven J. Cox and C. Maeve Mc-Carthy. The shape of the tallest column: Corrected. SIAM Journal on Mathematical Analysis, 31(4):940, 2000. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

http://epubs.siam.org/sam-
bin/dbq/article/35495. See

[CM98].

Clarkson:1986:IKG

[CMOR86] Peter A. Clarkson, J. Bryce McLeod, Peter J. Olver, and Alfred Ramani. Integrability of Klein–Gordon equations. SIAM Journal on Mathematical Analysis, 17(4):798–802, July 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Caprani:1981:IIF

[CMR81] Ole Caprani, Kaj Madsen, and L. B. Rall. Integration of interval functions. SIAM Journal on Mathematical Analysis, 12(3): 321–341, May 1981. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cannon:1977:WST

[CN77] John R. Cannon and Paul Nelson. Weak solutions of the time-dependent continuousenergy equation of neutron diffusion. SIAM Journal on Mathematical Analysis, 8(1):155–165, February 1977. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Clement:1979:ALN

[CN79] Ph. Clément and J. A. Nohel. Abstract linear and nonlinear Volterra equations preserving positivity. SIAM Journal on Mathematical Analysis, 10 (2):365–388, March 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Clement:1981:ABS

[CN81] Ph. Clément and J. A. Nohel. Asymptotic behavior of solutions of nonlinear Volterra equations with completely positive kernels. SIAM Journal on Mathematical Analysis, 12(4): 514–535, July 1981. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colli:1988:ERE

[CO88] Pierluigi Colli and Luc Oswald. An existence result for the electropainting problem. SIAM Journal on Mathematical Analysis, 19(6):1314–1323, November 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Caflisch:1989:SSI

[CO89] Russel E. Caflisch and Oscar F. Orellana. Singular solutions and ill-posedness for the evolution of vortex sheets. SIAM Journal on Mathematical Analysis, 20 (2):293–307, March 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cox:1992:ODC

[CO92] Steven J. Cox and Michael L.

Overton. On the optimal design of columns against buckling. SIAM Journal on Mathematical Analysis, 23(2):287–325, March 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cochran:1987:TDT

[Coc87]  James A. Cochran.  Three-

dimensional temperature response to impulsive input outside a spherical reservoir. SIAM Journal on Mathematical Analysis, 18(5):1394–1405, September 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Coffman:1982:SSW

[Cof82] Charles V. Coffman. On the structure of solutions Δ^{2}u = λu which satisfy the clamped plate conditions on a right angle. SIAM Journal on Mathematical Analysis, 13(5):746–757, September 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Coffman:1986:CSS

[Cof86] Charles V. Coffman. On the construction of series solutions to the first biharmonic boundary value problem on a rectangle. SIAM Journal on Mathematical Analysis, 17(2):384–402, March 1986. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Cohen:1971:TPB

[Coh71a] Edgar A. Cohen, Jr. Theoretical properties of best polynomial approximation in W^{1,2}[−1, 1]. SIAM Journal on Mathematical Analysis, 2(2):187–192, May 1971. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cohen:1971:TAS

[Coh71b] Edgar A. Cohen, Jr. Trigonometric approximation in the Sobolev spaces W^{r,2}[−π, π] with constant weights. SIAM Journal on Mathematical Analysis, 2(4):529–535, November 1971. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cohen:1972:MSS

[Coh72] Donald S. Cohen. Multiple solutions of singular perturbation problems. SIAM Journal on Mathematical Analysis, 3(1):72– 82, February 1972. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cohen:1975:ZDB

[Coh75] Edgar A. Cohen, Jr. Zero distribution and behavior of orthogonal polynomials in the Sobolev space W^{1,2}[−1, 1]. SIAM Journal on Mathematical Analysis, 6 (1):105–116, February 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cohen:1976:EPN

[Coh76] M. E. Cohen. On expansion problems: New classes of formulas for the classical functions. SIAM Journal on Mathematical Analysis, 7(5):702–712, October 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cohen:1989:CFF

[Coh89] Jonathan Cohen. Cauchy formulae for functions analytic of order two on C^{1} domains with applications to elastostatics and hydrostatics. SIAM Journal on Mathematical Analysis, 20(1):98–105, January 1989. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colton:1970:ACS

[Col70] David Colton. Applications of a class of singular partial differential equations to Gegenbauer series which converge to zero. SIAM Journal on Mathematical Analysis, 1(1):90–95, February 1970. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colton:1972:BOE

[Col72] David Colton. Bergman operators for elliptic equations in four independent variables. SIAM Journal on Mathematical Analysis, 3(3):401–412, August 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colton:1973:IPI

[Col73] David Colton. Improperly posed initial value problems for selfadjoint hyperbolic and elliptic equations. SIAM Journal on Mathematical Analysis, 4(1):42– 51, February 1973. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colton:1974:NCP

[Col74] David Colton. The noncharacteristic Cauchy problem for parabolic equations in one space variable. SIAM Journal on Mathematical Analysis, 5(2): 263–272, April 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colton:1975:CFS

[Col75] David Colton. Complete families of solutions for parabolic equations with analytic coefficients. SIAM Journal on Mathematical Analysis, 6(6):937–947, November 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colton:1978:SAW

[Col78] David Colton. The scattering of acoustic waves by a spherically stratified medium and an obstacle. SIAM Journal on Mathematical Analysis, 9(5):935–942, October 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colton:1982:RTF

[Col82] David Colton. Runge’s theorem and far field patterns for the impedance boundary value problem in acoustic wave propagation. SIAM Journal on Mathematical Analysis, 13(6):970–977, November 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colesanti:1994:VPH

[Col94a] Andrea Colesanti. A variational problem for harmonic functions in ring-shaped domains with partially free boundary. SIAM Journal on Mathematical Analysis, 25 (4):1122–1127, July 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24654.

Colin:1994:SED

[Col94b] Thierry Colin. Smoothing effects for dispersive equations via a generalized Wigner transform. SIAM Journal on Mathematical Analysis, 25(6):1622– 1641, November 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24322.

Colin:1997:CPC

[Col97] T. Colin. The Cauchy problem and the continuous limit for the multilayer model in geophysical fluid dynamics. SIAM Journal on Mathematical Analysis, 28 (3):516–529, May 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29126.

Cooke:1974:LHP

[Coo74] Kenneth L. Cooke. A linear hyperbolic problem all of whose solutions are constant after finite time. SIAM Journal on Mathematical Analysis, 5(3):482–488, May 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cooper:1988:AAC

[Coo88] Jeffery Cooper. The acoustic approximation for compressible flow in the presence of a surface undergoing small amplitude vibrations. SIAM Journal on Mathematical Analysis, 19 (1):60–69, January 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Costello:1973:GPD

[Cos73] Thomas Costello. Generic properties of differential equations. SIAM Journal on Mathematical Analysis, 4(2):245–249, May 1973. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cosner:1984:PBS

[Cos84] Chris Cosner. Pointwise bounds for strongly coupled time dependent systems of reactiondiffusion equations. SIAM Journal on Mathematical Analysis, 15 (2):350–356, March 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cosner:1987:EGS

[Cos87] Chris Cosner. Existence of global solutions to a model of a myelinated nerve axon. SIAM Journal on Mathematical Analysis, 18(3): 703–710, May 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Costabel:1988:BIO

[Cos88] Martin Costabel. Boundary integral operators on Lipschitz domains: elementary results. SIAM Journal on Mathematical Analysis, 19(3):613–626, May 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cannon:1973:SPI

[CP73] J. R. Cannon and Mario Primicerio. A Stefan problem involving the appearance of a phase. SIAM Journal on Mathematical Analysis, 4(1):141–148, February 1973. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cowling:1984:BVT

[CP84] Michael G. Cowling and John F.

Price. Bandwidth versus time concentration: the Heisenberg– Pauli–Weyl inequality. SIAM Journal on Mathematical Analysis, 15(1):151–165, January 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Colton:1990:FFP

[CP90] David Colton and Lassi Päivärinta.

Far-field patterns for electromagnetic waves in an inhomogeneous medium. SIAM Journal on Mathematical Analysis, 21(6):

1537–1549, November 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cerami:1997:NPS

[CP97] Giovanna Cerami and Donato Passaseo. Nonminimizing positive solutions for equations with critical exponents in the halfspace. SIAM Journal on Mathematical Analysis, 28(4):867–885, July 1997. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/29574.

Chan:1987:SOT

[CPS87] Wai Chan, Frank Proschan, and Jayaram Sethuraman. Schur– Ostrowski theorems for functionals on L_{1}(0, 1). SIAM Journal on Mathematical Analysis, 18 (2):566–578, March 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Caudill:1998:ISF

[CPS98] Lester F. Caudill, Jr., Peter A. Perry, and Albert W. Schueller. Isospectral sets for fourth-order ordinary differential operators. SIAM Journal on Mathematical Analysis, 29(4):935–966, July 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/31119.

Caprino:1998:SPS

[CPW98] S. Caprino, M. Pulvirenti, and W. Wagner. Stationary particle systems approximating stationary solutions to the Boltzmann equation. SIAM Journal on Mathematical Analysis, 29 (4):913–934, July 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30998.

Chadam:1997:SDE

[CQ97] J. Chadam and Y. Qin. Spatial decay estimates for flow in a porous medium. SIAM Journal on Mathematical Analysis, 28 (4):808–830, July 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29056.

Campbell:1979:SPA

[CR79] Stephen L. Campbell and Nicholas J. Rose. Singular perturbation of autonomous linear systems. SIAM Journal on Mathematical Analysis, 10(3): 542–551, May 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carmichael:1983:HFT

[CR83] Richard D. Carmichael and Stephen P. Richters. Holomorphic functions in tubes which have distributional boundary values and which are H^{p} functions. SIAM Journal on Mathematical Analysis, 14(3):596–621, May 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cohen:1988:ACS

[CR88] Paula S. Cohen and Amitai Regev. Asymptotics of combinatorial sums and the central limit theorem. SIAM Journal on Mathematical Analysis, 19(5): 1204–1215, September 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Churchill:1991:DZM

[CR91] Richard C. Churchill and David L. Rod. On the determination of Ziglin monodromy groups. SIAM Journal on Mathematical Analysis, 22(6):1790–1802, September 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Coppersmith:1992:GPB

[CR92] Don Coppersmith and T. J. Rivlin. The growth of polynomials bounded at equally spaced points. SIAM Journal on Mathematical Analysis, 23(4):970–983, July 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cheney:1988:FIR

[CRD88] Margaret Cheney, James H. Rose, and Brian DeFacio. A fundamental integral relation of scattering theory. SIAM Journal on Mathematical Analysis, 19(5): 1090–1102, September 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chui:1976:MCF

[CRSW76] C. K. Chui, E. R. Rozema, P. W. Smith, and J. D. Ward. Metric curvature, folding, and unique best approximation. SIAM Journal on Mathematical Analysis, 7 (3):436–449, May 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carmignani:1977:SDI

[CS77a] Robert Carmignani and Keith Schrader. Subfunctions and distributional inequalities. SIAM Journal on Mathematical Analysis, 8(1):52–68, February 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chee-Seng:1977:ARR

[CS77b] Lim Chee-Seng. Axial radiation reception. Part II. SIAM Journal on Mathematical Analysis, 8 (1):24–51, February 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chee-Seng:1977:TIE

[CS77c] Lim Chee-Seng. A transversely isotropic elliptic equation: An ARR analogy approach. SIAM Journal on Mathematical Analysis, 8(5):843–853, October 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Crooke:1978:IIC

[CS78] P. S. Crooke and R. P. Sperb.

Isoperimetric inequalities in a class of nonlinear eigenvalue problems. SIAM Journal on Mathematical Analysis, 9(4): 671–681, August 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Churchill:1979:PCA

[CS79] R. C. Churchill and J. F. Selgrade. On the Poincaré criterion for asymptotic stability. SIAM Journal on Mathematical Analysis, 10(6):1293–1298, November 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chui:1980:ASA

[CS80] Charles K. Chui and Philip W. Smith. An application of spline approximation with variable knots to optimal estimation of the derivative. SIAM Journal on Mathematical Analysis, 11(4): 724–736, July 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlson:1984:SPH

[CS84] B. C. Carlson and Dorothy B. Shaffer. Starlike and prestarlike hypergeometric functions. SIAM Journal on Mathematical Analysis, 15(4):737–745, July 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cooper:1985:IBP

[CS85] Jeffery Cooper and Walter Strauss. The initial boundary problem for the Maxwell equations in the presence of a moving body. SIAM Journal on Mathematical Analysis, 16(6): 1165–1179, November 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cantrell:1986:EPC

[CS86] Robert Stephen Cantrell and Klaus Schmitt. On the eigenvalue problem for coupled elliptic systems. SIAM Journal on Mathematical Analysis, 17(4):850–862, July 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chung:1989:CAF

[CS89] Dong Myung Chung and David Skoug. Conditional analytic Feynman integrals and a related Schrödinger integral equation. SIAM Journal on Mathematical Analysis, 20(4):950–965, July 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chui:1993:ILP

[CS93a] Charles K. Chui and Xian Liang Shi. Inequalities of Littlewood– Paley type for frames and wavelets. SIAM Journal on Mathematical Analysis, 24(1): 263–277, January 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cohen:1993:ACC

[CS93b] Albert Cohen and Qi Yu Sun.

An arithmetic characterization of the conjugate quadrature filters associated to orthonormal wavelet bases. SIAM Journal on Mathematical Analysis, 24(5): 1355–1360, September 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Camilli:1997:DSH

[CS97a] Fabio Camilli and Antonio Siconolfi. Discontinuous solutions of a Hamilton–Jacobi equation with infinite speed of propagation. SIAM Journal on Mathematical Analysis, 28(6):1420– 1445, November 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29804.

Chui:1997:IMD

[CS97b] Charles K. Chui and Xianliang Shi. Inequalities on matrixdilated Littlewood–Paley energy functions and oversampled affine operators. SIAM Journal on Mathematical Analysis, 28(1): 213–232, January 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/27200.

Cabral:1999:SRE

[CS99] H. E. Cabral and D. S. Schmidt. Stability of relative equilibria in the problem of N + 1 vortices. SIAM Journal on Mathematical Analysis, 31(2):231– 250, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/30212.

Carr:1987:NBS

[CSv87] Jack Carr, Jan A. Sanders, and Stephan A. van Gils. Nonresonant bifurcations with symmetry. SIAM Journal on Mathematical Analysis, 18(3):579–591, May 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chui:1980:DAM

[CSW80] C. K. Chui, P. W. Smith, and J. D. Ward. Degree of L_{p} approximation by monotone splines. SIAM Journal on Mathematical Analysis, 11(3):436–447, May 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chui:1991:ISB

[CSW91] C. K. Chui, J. Stöckler, and J. D.

Ward. Invertibility of shifted box spline interpolation operators. SIAM Journal on Mathematical Analysis, 22(2):543–553, March 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Comincioli:1988:MMC

[CT88] V. Comincioli and A. Torelli. A mathematical model of contracting muscle with viscoelastic elements. SIAM Journal on Mathematical Analysis, 19(3):593–612, May 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Curry:1979:BSF

[Cur79]  James H. Curry.  Bounded

solutions of finite dimensional approximations to the Boussinesq equations. SIAM Journal on Mathematical Analysis, 10 (1):71–77, January 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cushing:1971:UPS

[Cus71] J. M. Cushing. Uniqueness of positive solutions to nonlinear elliptic problems. SIAM Journal on Mathematical Analysis, 2(2): 290–292, May 1971. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cushing:1975:OEB

[Cus75] J. M. Cushing. An operator equation and bounded solutions of integro-differential systems. SIAM Journal on Mathematical Analysis, 6(3):433–445, May 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cushing:1982:PKS

[Cus82] J. M. Cushing. Periodic Kolmogorov systems. SIAM Journal on Mathematical Analysis, 13(5): 811–827, September 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cuyt:1983:APA

[Cuy83a] Annie A. M. Cuyt. The εalgorithm and Padé-approximants in operator theory. SIAM Journal on Mathematical Analysis, 14 (5):1009–1014, September 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cuyt:1983:CSM

[Cuy83b] Annie A. M. Cuyt. A comparison of some multivariate Padéapproximants. SIAM Journal on Mathematical Analysis, 14(1): 195–202, January 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cuyt:1984:QAP

[Cuy84] Annie A. M. Cuyt. The QDalgorithm for Padé-approximants in operator theory. SIAM Journal on Mathematical Analysis, 15 (4):746–752, July 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cannarsa:1987:GAS

[CV87] Piermarco Cannarsa and Vincenzo Vespri. Generation of analytic semigroups by elliptic operators with unbounded coefficients. SIAM Journal on Mathematical Analysis, 18(3):857–872, May 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Clement:1992:NEP

[CvL92] Ph. Clément, C. J. van Duijn, and Shuanhu Li. On a nonlinear elliptic-parabolic partial differential equation system in a two-dimensional groundwater flow problem. SIAM Journal on Mathematical Analysis, 23(4): 836–851, July 1992. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chipot:1989:SBR

[CW89] M. Chipot and F. B. Weissler. Some blowup results for a nonlinear parabolic equation with a gradient term. SIAM Journal on Mathematical Analysis, 20(4): 886–907, July 1989. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Constantin:1999:BSQ

[CW99] Peter Constantin and Jiahong Wu. Behavior of solutions of 2D quasi-geostrophic equations. SIAM Journal on Mathematical Analysis, 30(5):937– 948, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/33733.

Chandler-Wilde:1999:SEW

[CWZ99] Simon N. Chandler-Wilde and Bo Zhang. Scattering of electromagnetic waves by rough interfaces and inhomogeneous layers. SIAM Journal on Mathematical Analysis, 30(3):559– 583, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/32893.

Cybenko:1984:RNO

[Cyb84] George Cybenko. Restrictions of normal operators, Padé approximation and autoregressive time series. SIAM Journal on Mathematical Analysis, 15(4):753–767, July 1984. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Coffman:1991:PMC

[CZ91] Charles V. Coffman and William K.

Ziemer. A prescribed mean curvature problem on domains without radial symmetry. SIAM Journal on Mathematical Analysis, 22(4):982–990, July 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Conca:1994:AAM

[CZ94] Carlos Conca and Enrike Zuazua.

Asymptotic analysis of a multidimensional vibrating structure. SIAM Journal on Mathematical Analysis, 25(3):836–858, May 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/24338.

DArchangelo:1975:IRS

[D’A75] James D’Archangelo. An integral representation for the solution W_{km} of Whittaker’s differential equation. SIAM Journal on Mathematical Analysis, 6(6): 907–912, November 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dafermos:1982:GSS

[Daf82] C. M. Dafermos. Global smooth solutions to the initial-boundary value problem for the equations of one-dimensional nonlinear thermoviscoelasticity. SIAM Journal on Mathematical Analysis, 13(3):397–408, May 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dafermos:1987:ECL

[Daf87] Constantine M. Dafermos. Estimates for conservation laws with little viscosity. SIAM Journal on Mathematical Analysis, 18 (2):409–421, March 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dafermos:1995:SSC

[Daf95] C. M. Dafermos. Stability for systems of conservation laws in several space dimensions. SIAM Journal on Mathematical Analysis, 26(6):1403–1414, November 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/25847.

Damlamian:1981:HHN

[Dam81]  Alain Damlamian.  How to

homogenize a nonlinear diffusion equation: Stefan’s problem. SIAM Journal on Mathematical Analysis, 12(3):306–313, May 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Damon:1987:TDN

[Dam87] James Damon. Time dependent nonlinear oscillations with many periodic solutions. SIAM Journal on Mathematical Analysis, 18(5):

1294–1316, September 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Datko:1972:UAS

[Dat72] Richard Datko. Uniform asymptotic stability of evolutionary processes in a Banach space. SIAM Journal on Mathematical Analysis, 3(3):428–445, August 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Datko:1981:PCD

[Dat81] Richard Datko. The Perron condition for differential-difference equations in a Hilbert space. SIAM Journal on Mathematical Analysis, 12(6):842–852, November 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dauge:1989:SSN

[Dau89]  Monique Dauge.  Station-

ary Stokes and Navier–Stokes systems on two- or threedimensional domains with corners. Part I. Linearized equations. SIAM Journal on Mathematical Analysis, 20(1):74–97, January 1989. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Daubechies:1993:OBC

[Dau93] Ingrid Daubechies. Orthonormal bases of compactly supported wavelets II variations on a theme. SIAM Journal on Mathematical Analysis, 24(2):499–519, March 1993. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Davies:1973:CZL

[Dav73] B. Davies. Complex zeros of linear combinations of spherical Bessel functions and their derivatives. SIAM Journal on Mathematical Analysis, 4(1):128–133, February 1973. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Davis:1978:MTS

[Dav78a] Jon H. Davis. A multi-time scale method in almost-periodic stability problems. SIAM Journal on Mathematical Analysis, 9(6): 979–995, December 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Davis:1978:LTH

[Dav78b] Paul L. Davis. On the linear theory of heat conduction for materials with memory. SIAM Journal on Mathematical Analysis, 9 (1):49–53, February 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Davis:1979:SPS

[Dav79] Paul L. Davis. On the speed of propagation of solutions of integro-differential equations. SIAM Journal on Mathematical Analysis, 10(3):570–576, May 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Day:1982:ELS

[Day82] Martin Day. Exponential leveling for stochastically perturbed dynamical systems. SIAM Journal on Mathematical Analysis, 13(4): 532–540, July 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Day:1989:BLT

[Day89] Martin V. Day. Boundary local time and small parameter exit problems with characteristic boundaries. SIAM Journal on Mathematical Analysis, 20(1): 222–248, January 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

deBoor:1976:CSI

[dB76] Carl de Boor. On the cardinal spline interpolant to e^{iut}. SIAM Journal on Mathematical Analysis, 7(6):930–941, November 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

deBoor:1982:PBM

[dBJ82] Carl de Boor and Joseph W.

Jerome. Professional biography: Michael Golomb. SIAM Journal on Mathematical Analysis, 13(3): vii–xiv, May 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

DeOliveira:1985:LFR

[DC85] J. C. F. De Oliveira and L. A. V.

Carvalho. A Lyapunov functional for a retarded differential equation. SIAM Journal on Mathematical Analysis, 16(6): 1295–1305, November 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Damlamian:1991:HSP

[DD91] Alain Damlamian and Patrizia Donato. Homogenization with small perforations of increasingly complicated shapes. SIAM Journal on Mathematical Analysis, 22 (3):639–652, May 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Davis:1996:SPE

[DDD96] Stephen H. Davis, Emmanuele DiBenedetto, and David J. Diller. Some A priori estimates for a singular evolution equation arising in thin-film dynamics. SIAM Journal on Mathematical Analysis, 27(3):638–660, May 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Duval-Destin:1993:CWD

[DDMT93] M. Duval-Destin, M. A. Muschietti, and B. Torresani. Continuous wavelet decompositions, multiresolution, and contrast analysis. SIAM Journal on Mathematical Analysis, 24(3):739–755, May 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dijksma:1974:DWT

[DdS74] A. Dijksma and H. S. V. de Snoo. Distributional Watson transforms. SIAM Journal on Mathematical Analysis, 5(6):888–892, November 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

DeVore:1977:MAP

[De 77a] Ronald A. De Vore. Monotone approximation by polynomials. SIAM Journal on Mathematical Analysis, 8(5):906–921, October 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

DeVore:1977:MAS

[De 77b] Ronald A. De Vore. Monotone approximation by splines. SIAM Journal on Mathematical Analysis, 8(5):891–905, October 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

deRoever:1978:ARF

[de 78] J. W. de Roever. Analytic representations and Fourier transforms of analytic functionals in Z′ carried by the real space. SIAM Journal on Mathematical Analysis, 9(6):996–1019, December 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

deGroen:1980:NRS

[de 80] P. P. N. de Groen. The nature of resonance in a singular perturbation problem of turning point type. SIAM Journal on Mathematical Analysis, 11(1):1– 22, January 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

deSousaPinto:1985:GHC

[de 85] J. de Sousa Pinto. A generalised Hankel convolution. SIAM Journal on Mathematical Analysis, 16 (6):1335–1346, November 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

deBouard:1995:ISB

[de 95] Anne de Bouard. Instability of stationary bubbles. SIAM Journal on Mathematical Analysis, 26 (3):566–582, May 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/23702.

Deavours:1971:BVP

[Dea71] C. A. Deavours. A boundary value problem whose solution involves equations nonlinear in an eigenvalue parameter. SIAM Journal on Mathematical Analysis, 2(2):168–186, May 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Deans:1979:GTR

[Dea79] Stanley R. Deans. Gegenbauer transforms via the Radon transform. SIAM Journal on Mathematical Analysis, 10(3):577–585, May 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dekleine:1971:BSS

[Dek71] H. Arthur Dekleine. Bounded solutions for a second order nonlinear equation. SIAM Journal on Mathematical Analysis, 2(4): 511–520, November 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Demoulini:1996:YMS

[Dem96] Sophia Demoulini. Young measure solutions for a nonlinear parabolic equation of forwardbackward type. SIAM Journal on Mathematical Analysis, 27 (2):376–403, March 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/26184.

Deng:1990:HBN

[Den90] Bo Deng. Homoclinic bifurcations with nonhyperbolic equilibria. SIAM Journal on Mathematical Analysis, 21(3):693–720, May 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Deng:1991:BCC

[Den91a] Bo Deng. The bifurcations of countable connections from a twisted heteroclinic loop. SIAM Journal on Mathematical Analysis, 22(3):653–679, May 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Deng:1991:EIM

[Den91b] Bo Deng. The existence of infinitely many traveling front and back waves in the FitzHugh– Nagumo equations. SIAM Journal on Mathematical Analysis, 22 (6):1631–1650, September 1991.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Deng:1995:DBS

[Den95] Keng Deng. Dynamical behavior of solutions of a semilinear heat equation with nonlocal singularity. SIAM Journal on Mathematical Analysis, 26 (1):98–111, January 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/22388.

DeSanti:1987:NIL

[DeS87] Albert J. DeSanti. Nonmonotone interior layer theory for some singularly perturbed quasilinear boundary value problems with turning points. SIAM Journal on Mathematical Analysis, 18 (2):321–331, March 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Desesa:1998:OSR

[Des98] Blaise Desesa. Orthogonality of sieved random walk polynomials from a nonsieved analogue. SIAM Journal on Mathematical Analysis, 29(1):251–265, January 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/29843.

Dettman:1978:STC

[Det78] John W. Dettman. Saturation theorems connected with the abstract wave equation. SIAM Journal on Mathematical Analysis, 9(1):54–64, February 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dette:1995:NBH

[Det95] Holger Dette. New bounds for Hahn and Krawtchouk polynomials. SIAM Journal on Mathematical Analysis, 26(6):1647– 1659 (or 1647–1662??), November 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/25566.

Dassios:1983:EES

[DG83] George Dassios and Manoussos Grillakis. Equipartition of energy in scattering theory. SIAM Journal on Mathematical Analysis, 14(5):915–924, September 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Delsarte:1988:TAS

[DG88] P. Delsarte and Y. Genin. The tridiagonal approach to Szegő’s orthogonal polynomials, Toeplitz linear systems, and related interpolation problems. SIAM Journal on Mathematical Analysis, 19 (3):718–735, May 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Desch:1989:SPL

[DG89] W. Desch and R. Grimmer.

Smoothing properties of linear Volterra integro-differential equations. SIAM Journal on Mathematical Analysis, 20(1): 116–132, January 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Denis:1992:BIT

[DG92] R. Y. Denis and R. A. Gustafson. An SU(n) q-beta integral transformation and multiple hypergeometric series identities. SIAM Journal on Mathematical Analysis, 23(2):552–561, March 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

deGroot:1975:SIE

[dGdJ75] E. H. de Groot and E. M. de Jager. On the solution of an integral equation of convolution type. SIAM Journal on Mathematical Analysis, 6(1):208–227, February 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

DalPasso:1998:FOD

[DGG98] Roberta Dal Passo, Harald Garcke, and Günther Grün. On a fourth-order degenerate parabolic equation: Global entropy estimates, existence, and qualitative behavior of solutions. SIAM Journal on Mathematical Analysis, 29(2):321–342, March 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/30617.

Donovan:1996:IMA

[DGH96] George C. Donovan, Jeffrey S.

Geronimo, and Douglas P. Hardin. Intertwining multiresolution analyses and the construction of piecewise-polynomial wavelets. SIAM Journal on Mathematical Analysis, 27(6): 1791–1815, November 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/27616.

Donovan:1999:OPC

[DGH99] G. C. Donovan, J. S. Geronimo, and D. P. Hardin. Orthogonal polynomials and the construction of piecewise polynomial smooth wavelets. SIAM Journal on Mathematical Analysis, 30(5):1029–1056, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/31311.

Donovan:1996:COW

[DGHM96] George C. Donovan, Jeffrey S.

Geronimo, Douglas P. Hardin, and Peter R. Massopust. Construction of orthogonal wavelets using fractal interpolation functions. SIAM Journal on Mathematical Analysis, 27(4):1158– 1192, July 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/25652.

Dauge:1999:ILB

[DGR99] Monique Dauge, Isabelle Gruais, and Andreas Rössle. The influence of lateral boundary conditions on the asymptotics in thin elastic plates. SIAM Journal on Mathematical Analysis, 31(2):305–345, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33302.

Diaz:1984:EFB

[DH84] J. Ildefonso Diaz and Jesús Hernández. On the existence of a free boundary for a class of reaction-diffusion systems. SIAM Journal on Mathematical Analysis, 15(4):670–685, July 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Du:1999:BSC

[DH99] Yihong Du and Qingguang Huang. Blow-up solutions for a class of semilinear elliptic and parabolic equations. SIAM Journal on Mathematical Analysis, 31(1):1–18, 1999. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

                                        http://epubs.siam.org/sam-
                                        bin/dbq/article/35284.

deHoog:1973:SVI

[dHW73] Frank de Hoog and Richard Weiss. On the solution of a Volterra integral equation with a weakly singular kernel. SIAM Journal on Mathematical Analysis, 4(4):561–573, November 1973. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Diener:1986:SSE

[Die86] Francine Diener. Sauts des solutions des équations ϵẍ = f(t, x, ẋ). (French) [Jumps of solutions of equations of the type ϵẍ = f(t, x, ẋ)]. SIAM Journal on Mathematical Analysis, 17(3): 533–559, May 1986. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Diener:1988:ESD

[Die88]  Francine Diener.  Équations

surquadratiques et disparition des sauts. (French) [Superquadratic equations and disappearance of jumps]. SIAM Journal on Mathematical Analysis, 19(5):1127– 1134, September 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Diener:1994:RMR

[Die94] Marc Diener. Regularizing microscopes and rivers. SIAM Journal on Mathematical Analysis, 25 (1):148–173, January 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/21937.

Diehl:1995:SCL

[Die95] Stefan Diehl. On scalar conservation laws with point source and discontinuous flux function. SIAM Journal on Mathematical Analysis, 26(6):1425–1451, November 1995. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

http://epubs.siam.org/sam-
bin/dbq/article/24253.

DiGiallonardo:1981:EOE

[DiG81]  Joseph DiGiallonardo.  Ex-

istence, oscillation and eigenvalue comparison theorems for two- and three-point fourth order problems. SIAM Journal on Mathematical Analysis, 12(6): 866–873, November 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dilcher:1988:PRE

[Dil88] Karl Dilcher. Polynomials related to expansions of certain rational functions in two variables. SIAM Journal on Mathematical Analysis, 19(2):473–483, March 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ding:1989:KPP

[Din89] Sun Hong Ding. On a kind of predator-prey system. SIAM Journal on Mathematical Analysis, 20(6):1426–1435, November 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dix:1996:NUI

[Dix96] Daniel B. Dix. Nonuniqueness and uniqueness in the initialvalue problem for Burgers’ equation. SIAM Journal on Mathematical Analysis, 27(3):708–724, May 1996. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

DeGance:1987:IST

[DJ87] A. E. DeGance and L. E. Johns.

Infinite sums in the theory of dispersion of chemically reactive solute. SIAM Journal on Mathematical Analysis, 18(2):473–485, March 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Daubechies:1991:ESW

[DJJ91a] Ingrid Daubechies, Stéphane Jaffard, and Jean-Lin Journé. Erratum: “A simple Wilson orthonormal basis with exponential decay”. SIAM Journal on Mathematical Analysis, 22(3): 878, May 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [DJJ91b].

Daubechies:1991:SWO

[DJJ91b] Ingrid Daubechies, Stéphane Jaffard, and Jean-Lin Journé. A simple Wilson orthonormal basis with exponential decay. SIAM Journal on Mathematical Analysis, 22(2):554–573, March 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [DJJ91a].

Ditzian:1993:SPA

[DJL93] Z. Ditzian, D. Jiang, and D. Leviatan. Simultaneous polynomial approximation. SIAM Journal on Mathematical Analysis, 24(6): 1652–1661, November 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

deKok:1971:MSP

[dK71] F. de Kok. On the method of stationary phase for multiple integrals. SIAM Journal on Mathematical Analysis, 2(1):76– 104, February 1971. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Deutsch:1983:CSA

[DK83] Frank Deutsch and Petar Kenderov. Continuous selections and approximate selection for setvalued mappings and applications to metric projections. SIAM Journal on Mathematical Analysis, 14(1):185–194, January 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dunninger:1986:PBE

[DK86] D. R. Dunninger and J. C. Kurtz. A priori bounds and existence of positive solutions for singular nonlinear boundary value problems. SIAM Journal on Mathematical Analysis, 17(3):595–609, May 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Diomeda:1984:BPB

[DL84] Lorenza Diomeda and Benedetta Lisena. Boundary problems for the biharmonic operator in a square with L^{p}-data. SIAM Journal on Mathematical Analysis, 15(6):1153–1168, November 1984. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Daubechies:1991:TSD

[DL91a] Ingrid Daubechies and Jeffrey C.

Lagarias. Two-scale difference equations. I. Existence and global regularity of solutions. SIAM Journal on Mathematical Analysis, 22(5):1388–1410, September 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dunster:1991:CFS

[DL91b] T. M. Dunster and D. A. Lutz.

Convergent factorial series expansions for Bessel functions. SIAM Journal on Mathematical Analysis, 22(4):1156–1172, July 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Daubechies:1992:TSD

[DL92a] Ingrid Daubechies and Jeffrey C.

Lagarias. Two-scale difference equations. II. Local regularity, infinite products of matrices and fractals. SIAM Journal on Mathematical Analysis, 23(4):1031– 1079, July 1992. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dockery:1992:ETW

[DL92b] Jack D. Dockery and Roger Lui.

Existence of travelling wave solutions for a bistable evolutionary ecology model. SIAM Journal on Mathematical Analysis, 23(4): 870–888, July 1992. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

DeVore:1996:SSS

[DL96] Ronald A. DeVore and Bradley J. Lucier. On the size and smoothness of solutions to nonlinear hyperbolic conservation laws. SIAM Journal on Mathematical Analysis, 27(3):684–707, May 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ding:1998:VPR

[DLY98] Shijin Ding, Zuhan Liu, and Wanghui Yu. A variational problem related to the Ginzburg– Landau model of superconductivity with normal impurity inclusion. SIAM Journal on Mathematical Analysis, 29(1): 48–68, January 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/30308.

Dunne:1982:ACA

[DM82] Edward G. Dunne and Dale H. Mugler. Analogies from complex analysis and heat conduction for wave propagation. SIAM Journal on Mathematical Analysis, 13(1): 1–15, January 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

deFigueiredo:1986:MPE

[dM86] Djairo G. de Figueiredo and Enzo Mitidieri. A maximum principle for an elliptic system and applications to semilinear problems. SIAM Journal on Mathematical Analysis, 17(4):836–849, July 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

deHoog:1989:CMI

[dM89] F. R. de Hoog and R. M. M.

Mattheij. On the conditioning of multipoint and integral boundary value problems. SIAM Journal on Mathematical Analysis, 20 (1):200–214, January 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Micheli:1998:RFI

[DMV98] Enrico De Micheli, Nicodemo Magnoli, and Giovanni Alberto Viano. On the regularization of Fredholm integral equations of the first kind. SIAM Journal on Mathematical Analysis, 29 (4):855–877, July 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30174.

Dombrowski:1986:OPM

[DN86] Joanne Dombrowski and Paul Nevai. Orthogonal polynomials, measures and recurrence relations. SIAM Journal on Mathematical Analysis, 17(3):752–759, May 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

doNascimento:1993:BIS

[Do 93] Arnaldo S. Do Nascimento.

Boundary and interior spike layer formation for an elliptic equation with symmetry. SIAM Journal on Mathematical Analysis, 24 (2):466–479, March 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dolezal:1970:RLC

[Dol70] Vaclav Dolezal. A representation of linear continuous operators on testing functions and distributions. SIAM Journal on Mathematical Analysis, 1(4):491–506, November 1970. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dolezal:1972:GCL

[Dol72] Vaclav Dolezal. A generalized causality of linear continuous operators defined on distributions. SIAM Journal on Mathematical Analysis, 3(1):170–174, February 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dolezal:1977:NNC

[Dol77] Vaclav Dolezal. Nonlinear networks with current sources and Tellegen’s theorem. SIAM Journal on Mathematical Analysis, 8 (3):473–485, May 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dolezal:1979:LMH

[Dol79a] Vaclav Dolezal. Linearization of monotone Hilbert networks. SIAM Journal on Mathematical Analysis, 10(3):523–531, May 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dolezal:1979:OLE

[Dol79b] Vaclav Dolezal. Optimal linearization of equations involving monotone operators. SIAM Journal on Mathematical Analysis, 10(2):354–364, March 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dolezal:1979:SAS

[Dol79c] Vaclav Dolezal. A sensitivity analysis of systems consisting of linear blocks. SIAM Journal on Mathematical Analysis, 10(6): 1121–1137, November 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dolezal:1979:SCH

[Dol79d] Vaclav Dolezal. Solvability of a class of Hilbert networks. SIAM Journal on Mathematical Analysis, 10(2):404–411, March 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dolezal:1980:ATH

[Dol80] Vaclav Dolezal. An approximation theorem for a Hammersteintype equation and applications. SIAM Journal on Mathematical Analysis, 11(2):392–399, March 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dolezal:1981:FRS

[Dol81a] Vaclav Dolezal. Feedback reducing sensitivity by a factor λ < 1. SIAM Journal on Mathematical Analysis, 12(1):40–48, January 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dolezal:1981:NSC

[Dol81b] Vaclav Dolezal. A note on strictly causal operators. SIAM Journal on Mathematical Analysis, 12(3):469–470, May 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Domar:1972:HAC

[Dom72] Yngve Domar. Harmonic analysis of a class of distributions on R^{n}, n ≥ 2. SIAM Journal on Mathematical Analysis, 3(2): 230–245, May 1972. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dombrowski:1988:SMO

[Dom88]  Joanne Dombrowski.  Spec-

tral measures, orthogonal polynomials, and absolute continuity. SIAM Journal on Mathematical Analysis, 19(4):939–943, July 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Donoho:1992:SSC

[Don92] David L. Donoho. Superresolution via sparsity constraints. SIAM Journal on Mathematical Analysis, 23(5):1309–1331, September 1992. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dooley:1989:NVS

[Doo89] A. H. Dooley. A nonabelian version of the Shannon sampling theorem. SIAM Journal on Mathematical Analysis, 20(3): 624–633, May 1989. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dorr:1970:SES

[Dor70] Fred W. Dorr. Some examples of singular perturbation problems with turning points. SIAM Journal on Mathematical Analysis, 1(2):141–146, May 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dorfler:1987:NIM

[Dör87] P. Dörfler. New inequalities of Markov type. SIAM Journal on Mathematical Analysis, 18 (2):490–494, March 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dorfler:1991:EPC

[Dör91] P. Dörfler. An extremal problem concerning a Markov-type inequality. SIAM Journal on Mathematical Analysis, 22(3): 792–795, May 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dougherty:1997:HGIb

[Dou97] Michael M. Dougherty. Higher gradient integrability of minimizers for a polyconvex case in two dimensions. SIAM Journal on Mathematical Analysis, 28 (3):530–538, May 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29258.

Douglas:1998:RVV

[Dou98] R. J. Douglas. Rearrangements of vector valued functions, with application to atmospheric and oceanic flows. SIAM Journal on Mathematical Analysis, 29 (4):891–902, July 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/28216.

Dougherty:1997:HGIa

[DP97] Michael M. Dougherty and Daniel Phillips. Higher gradient integrability of equilibria for certain rank-one convex integrals. SIAM Journal on Mathematical Analysis, 28(2):270–273, March 1997. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/29338.

delPino:1999:RMC

[dPFW99] Manuel del Pino, Patricio L. Felmer, and Juncheng Wei. On the role of mean curvature in some singularly perturbed Neumann problems. SIAM Journal on Mathematical Analysis, 31(1):63–79, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/33283.

Demengel:1990:CFM

[DQ90] Françoise Demengel and Tang Qi. Convex function of a measure obtained by homogenization. SIAM Journal on Mathematical Analysis, 21(2):409–435, March 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dunkl:1974:KPS

[DR74] Charles F. Dunkl and Donald E.

Ramirez. Krawtchouk polynomials and the symmetrization of hypergroups. SIAM Journal on Mathematical Analysis, 5(3): 351–366, May 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Duchon:1986:NSI

[DR86] Jean Duchon and Raoul Robert.

A nonlinear singular integrodifferential equation arising in surface chemistry. SIAM Journal on Mathematical Analysis, 17(6): 1434–1446, November 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

DeCarvalho:1995:GAP

[DRF95] Alexandre Nolasco De Carvalho and José Gaspar Ruas-Filho. Global attractors for parabolic problems in fractional power spaces. SIAM Journal on Mathematical Analysis, 26(2):415–427, March 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/24642.

Driscoll:1987:EDD

[Dri87] Bonita Hart Driscoll. Eigenvalues on a domain with discrete rotational symmetry. SIAM Journal on Mathematical Analysis, 18 (4):941–953, July 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

DiBenedetto:1981:IDE

[DS81] Emmanuele Di Benedetto and R. E. Showalter. Implicit degenerate evolution equations and applications. SIAM Journal on Mathematical Analysis, 12(5): 731–751, September 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Derzko:1982:GSP

[DS82] N. A. Derzko and S. P. Sethi. General solution of the price-dividend integral equation. SIAM Journal on Mathematical Analysis, 13(1):106–111, January 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dangelmayr:1984:CUS

[DS84] Gerhard Dangelmayr and Ian Stewart. Classification and unfolding of sequential bifurcations. SIAM Journal on Mathematical Analysis, 15(3):423–445, May 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Diaconis:1987:TRS

[DS87] Persi Diaconis and Mehrdad Shahshahani. Time to reach stationarity in the Bernoulli– Laplace diffusion model. SIAM Journal on Mathematical Analysis, 18(1):208–218, January 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Diaz:1995:MAC

[DS95] Jesus Ildefonso Diaz and Ivar Stakgold. Mathematical aspects of the combustion of a solid by a distributed isothermal gas reaction. SIAM Journal on Mathematical Analysis, 26 (2):305–328, March 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24706.

deBouard:1997:SDG

[dS97] Anne de Bouard and Jean-Claude Saut. Symmetries and decay of the generalized Kadomtsev–Petviashvili solitary waves. SIAM Journal on Mathematical Analysis, 28(5):1064– 1085, September 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29766.

Dahmen:1999:WMC

[DS99] Wolfgang Dahmen and Reinhold Schneider. Wavelets on manifolds I: Construction and domain decomposition. SIAM Journal on Mathematical Analysis, 31(1):184–230, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33345.

deMottoni:1978:SCN

[dT78] P. de Mottoni and A. Tesei. On the solutions of a class of nonlinear Sturm–Liouville problems. SIAM Journal on Mathematical Analysis, 9(6):1020–1029, December 1978. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Deift:1981:CLM

[DT81] P. Deift and E. Trubowitz. A continuum limit of matrix inverse problems. SIAM Journal on Mathematical Analysis, 12(6): 799–818, November 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

DaPrato:1996:FNS

[DT96] G. Da Prato and L. Tubaro. Fully nonlinear stochastic partial differential equations. SIAM Journal on Mathematical Analysis, 27 (1):40–55, January 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/25676.

Duchateau:1995:MIC

[Duc95] Paul Duchateau. Monotonicity and invertibility of coefficient-to-data mappings for parabolic inverse problems. SIAM Journal on Mathematical Analysis, 26(6):

1473–1487, November 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/25925.

DuChateau:1997:IPH

[DuC97] Paul DuChateau. An inverse problem for the hydraulic properties of porous media. SIAM Journal on Mathematical Analysis, 28 (3):611–632, May 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/28567.

Dunkl:1974:EUP

[Dun74] Charles F. Dunkl. An expansion in ultraspherical polynomials with nonnegative coefficients. SIAM Journal on Mathematical Analysis, 5(1):51–52, February 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dunkl:1978:ATH

[Dun78] Charles F. Dunkl. An addition theorem for Hahn polynomials: The spherical functions. SIAM Journal on Mathematical Analysis, 9(4):627–637, August 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dunster:1986:UAE

[Dun86] T. M. Dunster. Uniform asymptotic expansions for prolate spheroidal functions with large parameters. SIAM Journal on Mathematical Analysis, 17(6):

1495–1524, November 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dunster:1989:UAE

[Dun89] T. M. Dunster. Uniform asymptotic expansions for Whittaker’s confluent hypergeometric functions. SIAM Journal on Mathematical Analysis, 20(3):744–760, May 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dunster:1990:BFP

[Dun90a] T. M. Dunster. Bessel functions of purely imaginary order, with an application to second-order linear differential equations having a large parameter. SIAM Journal on Mathematical Analysis, 21(4):995–1018, July 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dunster:1990:UAS

[Dun90b] T. M. Dunster. Uniform asymptotic solutions of second-order linear differential equations having a double pole with complex exponent and a coalescing turning point. SIAM Journal on Mathematical Analysis, 21(6): 1594–1618, November 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dunbar:1993:GAN

[Dun93] Steven R. Dunbar. Geometric analysis of a nonlinear boundary value problem from physical oceanography. SIAM Journal on Mathematical Analysis, 24 (2):444–465, March 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dunster:1994:UAS

[Dun94] T. M. Dunster. Uniform asymptotic solutions of second-order linear differential equations having a simple pole and a coalescing turning point in the complex plane. SIAM Journal on Mathematical Analysis, 25 (2):322–353, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/22953.

Durand:1977:SAF

[Dur77] Loyal Durand. A symmetrical addition formula for the Laguerre polynomials. SIAM Journal on Mathematical Analysis, 8 (3):541–546, May 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Durand:1978:PFN

[Dur78] Loyal Durand. Product formulas and Nicholson-type integrals for Jacobi functions. I: Summary of results. SIAM Journal on Mathematical Analysis, 9(1):76– 86, February 1978. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Durand:1979:AFJ

[Dur79] Loyal Durand. Addition formulas for Jacobi, Gegenbauer, Laguerre, and hyperbolic Bessel functions of the second kind.

SIAM Journal on Mathematical Analysis, 10(2):425–437, March 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dutka:1981:TRR

[Dut81] Jacques Dutka. Two results of Ramanujan. SIAM Journal on Mathematical Analysis, 12(3): 471–476, May 1981. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dutka:1985:SGF

[Dut85] Jacques Dutka. On some gamma function inequalities. SIAM Journal on Mathematical Analysis, 16(1):180–185, January 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Damlamian:1987:HLE

[DV87] A. Damlamian and M. Vogelius. Homogenization limits of the equations of elasticity in thin domains. SIAM Journal on Mathematical Analysis, 18(2):435–451, March 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Derks:1996:DHS

[Dv96] G. Derks and E. van Groesen. Dissipation in Hamiltonian systems: Decaying cnoidal waves. SIAM Journal on Mathematical Analysis, 27(5):1424– 1447, September 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/25342.

deHoog:1980:BVP

[dW80] Frank R. de Hoog and Richard Weiss. On the boundary value problem for systems of ordinary differential equations with a singularity of the second kind. SIAM Journal on Mathematical Analysis, 11(1):41–60, January 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dyn:1982:EFS

[DW82] Nira Dyn and Grace Wahba. On the estimation of functions of several variables from aggregated data. SIAM Journal on Mathematical Analysis, 13(1):134–152, January 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Donaldson:1993:LSW

[DW93] James A. Donaldson and Daniel A.

Williams. The linear shallow water theory: A mathematical justification. SIAM Journal on Mathematical Analysis, 24(4): 892–910, July 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Eckhaus:1992:SPH

[Eck92] Wiktor Eckhaus. Singular perturbations of homoclinic orbits in R^{4}. SIAM Journal on Mathematical Analysis, 23(5):1269– 1290, September 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Edelson:1979:PSE

[Ede79]  Allan L. Edelson.  Posi-

tive solutions of the equation (m(t)x′(t))′ + A(t)x(t) = 0. SIAM Journal on Mathematical Analysis, 10(4):789–793, July 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Edward:1974:QEP

[Edw74] Hughes Edward. On a quadratic eigenvalue problem. SIAM Journal on Mathematical Analysis, 5 (2):319–326, April 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Edward:1993:SJM

[Edw93] Julian Edward. Spectra of Jacobi matrices, differential equations on the circle, and the su(1, 1) Lie algebra. SIAM Journal on Mathematical Analysis, 24 (3):824–831, May 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elliott:1985:AMP

[EF85] Charles M. Elliott and Avner Friedman. Analysis of a model of percolation in a gently sloping sand-bank. SIAM Journal on Mathematical Analysis, 16(5): 941–954, September 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Eckhaus:1990:AAL

[EG90] W. Eckhaus and M. Garbey. Asymptotic analysis on large timescales for singular perturbations of hyperbolic type. SIAM Journal on Mathematical Analysis, 21(4):867–883, July 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Evans:1991:PNC

[EG91] Ronald Evans and John Greene.

Polynomials with nonnegative coefficients whose zeros have modulus one. SIAM Journal on Mathematical Analysis, 22 (4):1173–1182, July 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elliott:1996:CHE

[EG96] Charles M. Elliott and Harald Garcke. On the Cahn– Hilliard equation with degenerate mobility. SIAM Journal on Mathematical Analysis, 27 (2):404–423, March 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/26766.

Eggermont:1993:MER

[Egg93] P. P. B. Eggermont. Maximum entropy regularization for Fredholm integral equations of the first kind. SIAM Journal on Mathematical Analysis, 24(6): 1557–1576, November 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Eloe:1984:NBV

[EH84] P. W. Eloe and Johnny Henderson. Nonlinear boundary value problems and a priori bounds on solutions. SIAM Journal on Mathematical Analysis, 15(4): 642–647, July 1984. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Esquinas:1990:TWS

[EH90] J. Esquinas and M. A. Herrero. Travelling wave solutions to a semilinear diffusion system. SIAM Journal on Mathematical Analysis, 21(1):123–136, January 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ehle:1973:SMP

[Ehl73] Byron L. Ehle. A-stable methods and padé approximations to the exponential. SIAM Journal on Mathematical Analysis, 4(4): 671–680, November 1973. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elayyan:1997:URD

[EI97] Alaeddin Elayyan and Victor Isakov. On uniqueness of recovery of the discontinuous conductivity coefficient of a parabolic equation. SIAM Journal on Mathematical Analysis, 28(1): 49–59, January 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/28601.

Eirola:1992:SCS

[Eir92] Timo Eirola. Sobolev characterization of solutions of dilation equations. SIAM Journal on Mathematical Analysis, 23 (4):1015–1030, July 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ermentrout:1984:FPC

[EK84] George Bard Ermentrout and Nancy Kopell. Frequency plateaus in a chain of weakly coupled oscillators. I. SIAM Journal on Mathematical Analysis, 15 (2):215–237, March 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Everitt:1992:DOL

[EKLOO92] W. N. Everitt, A. M. Krall, L. L.

Littlejohn, and V. P. Onyango-Otieno. Differential operators and the Laguerre type polynomials. SIAM Journal on Mathematical Analysis, 23(3):722–736, May 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elbert:1984:SZB

[EL84] Árpád Elbert and Andrea Laforgia. On the square of the zeros of Bessel functions. SIAM Journal on Mathematical Analysis, 15 (1):206–212, January 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elbert:1985:CZB

[EL85] Árpád Elbert and Andrea Laforgia. On the convexity of the zeros of Bessel functions. SIAM Journal on Mathematical Analysis, 16 (3):614–619, May 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elbert:1986:MPZ

[EL86a] Árpád Elbert and Andrea Laforgia. Monotonicity properties of the zeros of Bessel functions. SIAM Journal on Mathematical Analysis, 17(6):1483–1488, November 1986. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elbert:1986:NPA

[EL86b] Árpád Elbert and Andrea Laforgia. A note to the paper of Ahmed, Muldoon and Spigler. SIAM Journal on Mathematical Analysis, 17(4):1008–1009, July 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elbert:1987:NPZ

[EL87] Árpád Elbert and Andrea Laforgia. New properties of the zeros of a Jacobi polynomial in relation to their centroid. SIAM Journal on Mathematical Analysis, 18(6): 1563–1572, November 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elcrat:1989:BFD

[EL89] Alan R. Elcrat and Kirk E. Lancaster. Bernstein functions and the Dirichlet problem. SIAM Journal on Mathematical Analysis, 20(5):1055–1068, September 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

ElBialy:1990:CSC

[ElB90] Mohamed Sami ElBialy. Collision singularities in celestial mechanics. SIAM Journal on Mathematical Analysis, 21(6):1563– 1593, November 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elcrat:1971:CET

[Elc71] Alan R. Elcrat. A constructive existence theorem for a nonlinear elliptic equation. SIAM Journal on Mathematical Analysis, 2 (3):368–374, August 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elcrat:1972:IRP

[Elc72] Alan R. Elcrat. An inequality related to Poisson’s equation. SIAM Journal on Mathematical Analysis, 3(3):544–545, August 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elcrat:1974:CES

[Elc74] Alan R. Elcrat. Constructive existence for semilinear elliptic equations with discontinuous coefficients. SIAM Journal on Mathematical Analysis, 5 (4):663–672, August 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elcrat:1975:SIC

[Elc75] Alan R. Elcrat. Some identities concerning the Laplacian of a function satisfying mixed boundary conditions. SIAM Journal on Mathematical Analysis, 6(3): 496–502, May 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elcrat:1982:SFP

[Elc82] Alan R. Elcrat. Separated flow past a plate with spoiler. SIAM Journal on Mathematical Analysis, 13(4):632–639, July 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Eliason:1972:VAB

[Eli72] S. B. Eliason. Vertical asymptotes and bounds for certain solutions of a class of second order differential equations. SIAM Journal on Mathematical Analysis, 3(3):474–484, August 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elias:1984:CTD

[Eli84] Uri Elias. Comparison theorems for disfocality and disconjugacy of differential equations. SIAM Journal on Mathematical Analysis, 15(5):922–931, September 1984. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elliott:1982:OPA

[Ell82] David Elliott. Orthogonal polynomials associated with singular integral equations having a Cauchy kernel. SIAM Journal on Mathematical Analysis, 13(6): 1041–1052, November 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Evans:1995:RRS

[ELM^{+}95] W. D. Evans, Lance L. Littlejohn, Francisco Marcellan, Clemens Markett, and Andre Ronveaux. On recurrence relations for Sobolev orthogonal polynomials. SIAM Journal on Mathematical Analysis, 26 (2):446–467, March 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/22692.

Erdelyi:1994:RNT

[ELS94] Tamás Erdélyi, Xin Li, and E. B.

Saff. Remez- and Nikol′skiĭtype inequalities for logarithmic potentials. SIAM Journal on Mathematical Analysis, 25 (2):365–383, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23488.

Erdelyi:1970:FID

[EM70] A. Erdélyi and A. C. McBride.

Fractional integrals of distributions. SIAM Journal on Mathematical Analysis, 1(4):547–557, November 1970. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ei:1990:PFH

[EM90] S.-I. Ei and M. Mimura. Pattern formation in heterogeneous reaction-diffusion-advection systems with an application to population dynamics. SIAM Journal on Mathematical Analysis, 21 (2):346–361, March 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elbert:1994:DRP

[EM94] Árpád Elbert and Martin E. Muldoon. On the derivative with respect to a parameter of a zero of a Sturm–Liouville function. SIAM Journal on Mathematical Analysis, 25(2): 354–364, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/22887.

Erdelyi:1994:EGJ

[EMN94] Tamás Erdélyi, Alphonse P. Magnus, and Paul Nevai. Erratum: “Generalized Jacobi weights, Christoffel functions, and Jacobi polynomials”. SIAM Journal on Mathematical Analysis, 25(5):1461, September 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [NEM94].

Escher:1998:SDF

[EMS98] Joachim Escher, Uwe F. Mayer, and Gieri Simonett. The surface diffusion flow for immersed hypersurfaces. SIAM Journal on Mathematical Analysis, 29(6): 1419–1433, November 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32067.

Engler:1982:VCR

[Eng82] Hans Engler. A version of the chain rule and integrodifferential equations in Hilbert spaces. SIAM Journal on Mathematical Analysis, 13(5):801–810, September 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Engler:1987:DSF

[Eng87]  Hans Engler.  On the dy-

namic shear flow problem for viscoelastic liquids. SIAM Journal on Mathematical Analysis, 18(4): 972–990, July 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Engelberg:1999:APL

[Eng99] Shlomo Engelberg. An analytical proof of the linear stability of the viscous shock profile of the Burgers equation with fourth-order viscosity. SIAM Journal on Mathematical Analysis, 30(4):927–936, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/33639.

Elcrat:1997:IPF

[EP97] Alan Elcrat and Vassilis G. Papanicolaou. On the inverse problem of a fourth-order self-adjoint binomial operator. SIAM Journal on Mathematical Analysis, 28 (4):886–896, July 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29093.

Edmunds:1998:NRS

[EP98] D. E. Edmunds and N. I. Popivanov. A nonlocal regularization of some over-determined boundary-value problems I. SIAM Journal on Mathematical Analysis, 29(1):85–105, January 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/30323.

Edward:1999:BRL

[EP99] Julian Edward and David Pravica. Bounds on resonances for the Laplacian on perturbations of half-space. SIAM Journal on Mathematical Analysis, 30(6):1175–1184, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33172.

Epperson:1986:RCN

[Epp86] James F. Epperson. Regularization for a class of nonlinear evolution equations. SIAM Journal on Mathematical Analysis, 17 (1):84–90, January 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Erbe:1976:OAB

[Erb76] L. Erbe. Oscillation and asymptotic behavior of solutions of third order differential delay equations. SIAM Journal on Mathematical Analysis, 7(4): 491–500, August 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Erbe:1977:CTS

[Erb77] L. Erbe. Comparison theorems for second order Riccati equations with applications. SIAM Journal on Mathematical Analysis, 8(6):1032–1037, November 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Erdelyi:1974:AEI

[Erd74] A. Erdélyi. Asymptotic evaluation of integrals involving a fractional derivative. SIAM Journal on Mathematical Analysis, 5 (2):159–171, April 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Erhart:1973:LTP

[Erh73] John V. Erhart. Lyapunov theory and perturbations of differential equations. SIAM Journal on Mathematical Analysis, 4(3): 417–432, May 1973. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Etgen:1975:CNT

[ES75] G. J. Etgen and C. D. Shih.

Conditions for the nonoscillation of third order differential equations with nonnegative coefficients. SIAM Journal on Mathematical Analysis, 6(1):1– 8, February 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elcrat:1976:EPE

[ES76] Alan Elcrat and Vincent G.

Sigillito. An explicit a priori estimate for parabolic equations with applications to semilinear equations. SIAM Journal on Mathematical Analysis, 7 (5):746–753, October 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Evans:1984:AFZ

[ES84] Ronald J. Evans and Dennis Stanton. Asymptotic formulas for zero-balanced hypergeometric series. SIAM Journal on Mathematical Analysis, 15(5): 1010–1020, September 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elliott:1997:LFA

[ES97a] Charles M. Elliott and Reiner Schätzle. The limit of the fully anisotropic double-obstacle Allen–Cahn equation in the nonsmooth case. SIAM Journal on Mathematical Analysis, 28 (2):274–303, March 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/28673.

Escher:1997:CSM

[ES97b] Joachim Escher and Gieri Simonett. Classical solutions of multidimensional Hele–Shaw models. SIAM Journal on Mathematical Analysis, 28(5):1028– 1047, September 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29191.

Escobedo:1973:SCC

[Esc73] Richard Escobedo. Spherical convergence and certain singular integrals. SIAM Journal on Mathematical Analysis, 4(3): 409–416, May 1973. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Esham:1987:AAG

[Esh87] Benjamin F. Esham, Jr. Asymptotics and an asymptotic Galerkin method for hyperbolic-parabolic singular perturbation problems. SIAM Journal on Mathematical Analysis, 18(3):762–776, May 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Etgen:1971:TPB

[ET71] G. J. Etgen and S. C. Tefteller. A two-point boundary problem for nonlinear second order differential systems. SIAM Journal on Mathematical Analysis, 2(1):64– 71, February 1971. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Etgen:1972:SOD

[ET72] G. J. Etgen and S. C. Tefteller.

Second order differential equations with general boundary conditions. SIAM Journal on Mathematical Analysis, 3(3):512–519, August 1972. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Elliott:1974:AEF

[ET74] David Elliott and P. D. Tuan.

Asymptotic estimates of Fourier coefficients. SIAM Journal on Mathematical Analysis, 5(1):1– 10, February 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ermentrout:1989:USR

[ET89] G. Bard Ermentrout and William C. Troy. The uniqueness and stability of the rest state for strongly coupled oscillators. SIAM Journal on Mathematical Analysis, 20 (6):1436–1446, November 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ermentrout:1994:PRD

[ET94] G. Bard Ermentrout and W. C. Troy. Phaselocking in a reactiondiffusion equation with twist. SIAM Journal on Mathematical Analysis, 25(6):1504–1520, November 1994. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

http://epubs.siam.org/sam-
bin/dbq/article/23931.

Escobedo:1997:SDE

[EV97] M. Escobedo and L. Vega. A semilinear Dirac equation in H^{s}(R^{3}) for s > 1. SIAM Journal on Mathematical Analysis, 28 (2):338–362, March 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/28301.

Evans:1992:MBI

[Eva92] Ronald J. Evans. Multidimensional q-beta integrals. SIAM Journal on Mathematical Analysis, 23(3):758–765, May 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Evans:1996:GIH

[Eva96a] Lawrence C. Evans. A geometric interpretation of the heat equation with multivalued initial data. SIAM Journal on Mathematical Analysis, 27 (4):932–958, July 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/27543.

Evans:1996:UMP

[Eva96b] Lawrence C. Evans. An unusual minimization principle for parabolic gradient flows. SIAM Journal on Mathematical Analysis, 27(1):1–4, January 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/26970.

Etgen:1973:CRD

[EW73] G. J. Etgen and N. C. Wong.

Cyclicly related differential systems. SIAM Journal on Mathematical Analysis, 4(1):52–61, February 1973. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Esham:1989:HPS

[EW89] Benjamin F. Esham, Jr. and Richard J. Weinacht. Hyperbolicparabolic singular perturbations for quasilinear equations. SIAM Journal on Mathematical Analysis, 20(6):1344–1365, November 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Esham:1994:SPC

[EW94] B. F. Esham and R. J. Weinacht. Singular perturbations and the coupled/quasi-static approximation in linear thermoelasticity. SIAM Journal on Mathematical Analysis, 25(6):1521–1536, November 1994. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

http://epubs.siam.org/sam-
bin/dbq/article/24370.

Ewing:1975:ACP

[Ewi75] Richard E. Ewing. The approximation of certain parabolic equations backward in time by Sobolev equations. SIAM Journal on Mathematical Analysis, 6 (2):283–294, April 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ei:1998:SDI

[EY98] Shin-Ichiro Ei and Eiji Yanagida. Slow dynamics of interfaces in the Allen–Cahn equation on a strip-like domain. SIAM Journal on Mathematical Analysis, 29 (3):555–595, May 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30720.

Escobedo:1997:LTB

[EZ97] Miguel Escobedo and Enrique Zuazua. Long-time behavior for a convection-diffusion equation in higher dimensions. SIAM Journal on Mathematical Analysis, 28 (3):570–594, May 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/27112.

Fair:1971:NCF

[Fai71] Wyman Fair. Noncommutative continued fractions. SIAM Journal on Mathematical Analysis, 2 (2):226–232, May 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Faierman:1977:DET

[Fai77] M. Faierman. On the distribution of the eigenvalues of a two-parameter system of ordinary differential equations of the second order. SIAM Journal on Mathematical Analysis, 8(5): 854–870, October 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fan:1993:OPR

[Fan93] Haitao Fan. One-phase Riemann problem and wave interactions in systems of conservation laws of mixed type. SIAM Journal on Mathematical Analysis, 24(4): 840–865, July 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fan:1997:SSS

[Fan97] Haitao Fan. Self-similar solutions for a modified Broadwell model and its fluid-dynamic limits. SIAM Journal on Mathematical Analysis, 28(4):831–851, July 1997. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/25641.

Farooqui:1973:CSO

[Far73] A. S. Farooqui. A complete set of orthonormal harmonic functions. SIAM Journal on Mathematical Analysis, 4(2):309–313, May 1973. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Farkas:1978:EER

[Far78] M. Farkas. Estimates on the existence regions of perturbed periodic solutions. SIAM Journal on Mathematical Analysis, 9 (5):876–890, October 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fields:1987:DDE

[FAS87] Jerry L. Fields and Waleed A. Al-Salam. A differential-difference equation. SIAM Journal on Mathematical Analysis, 18(3): 873–879, May 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fattorini:1985:SSP

[Fat85]  H. O. Fattorini.  On the

Schrödinger singular perturbation problem. SIAM Journal on Mathematical Analysis, 16(5): 1000–1019, September 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Friesecke:1997:ITD

[FD97] G. Friesecke and G. Dolzmann.

Implicit time discretization and global existence for a quasilinear evolution equation with nonconvex energy. SIAM Journal on Mathematical Analysis, 28 (2):363–380, March 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/28595.

Fehribach:1988:PMS

[Feh88] Joseph D. Fehribach. Perturbation methods for solid diffusion in a Stefan problem. SIAM Journal on Mathematical Analysis, 19 (1):86–99, January 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Feinerman:1978:DOSb

[Fei78] Robert P. Feinerman. Dual orthogonal series with modifier tending to zero. SIAM Journal on Mathematical Analysis, 9 (4):667–670, August 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Feinberg:1979:SWL

[Fei79a]  Jerry M. Feinberg.  Some

Wirtinger-like inequalities. SIAM Journal on Mathematical Analysis, 10(6):1258–1271, November 1979. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Feintuch:1979:SSS

[Fei79b]  Avraham Feintuch.  Strictly

and strongly strictly causal linear operators. SIAM Journal on Mathematical Analysis, 10(3): 603–613, May 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fenichel:1983:OBSa

[Fen83a] Neil Fenichel. Oscillatory bifurcations in singular perturbation theory. I. Slow oscillations. SIAM Journal on Mathematical Analysis, 14(5):861–867, September 1983. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fenichel:1983:OBSb

[Fen83b] Neil Fenichel. Oscillatory bifurcations in singular perturbation theory. II. Fast oscillations. SIAM Journal on Mathematical Analysis, 14(5):868–874, September 1983. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fenichel:1985:GUG

[Fen85] Neil Fenichel. Global uniqueness in geometric singular perturbation theory. SIAM Journal on Mathematical Analysis, 16(1):1– 6, January 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ferreira:1986:SDN

[Fer86] José M. Ferreira. On the stability of a distributed network. SIAM Journal on Mathematical Analysis, 17(1):38–45, January 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fettis:1970:RMR

[Fet70] Henry E. Fettis. On the reciprocal modulus relation for elliptic integrals. SIAM Journal on Mathematical Analysis, 1(4): 524–526, November 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fettis:1977:NRB

[Fet77] Henry E. Fettis. New relations between two types of Bessel function integrals. SIAM Journal on Mathematical Analysis, 8(6): 978–982, November 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See corrigendum [Fet78].

Fettis:1978:CNR

[Fet78]  Henry E. Fettis.  Corrigen-

dum: New relations between two types of Bessel function integrals. SIAM Journal on Mathematical Analysis, 9(4):779, August 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [Fet77].

Foias:1992:CLS

[FF92] C. Foias and A. E. Frazho. Commutant lifting and simultaneous H^{∞} and L^{2} suboptimization. SIAM Journal on Mathematical Analysis, 23(4):984–994, July 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ford:1980:PMP

[FFW80] Wayne T. Ford, Maria C. Fuente, and Margaret C. Waid. Porous media problems. SIAM Journal on Mathematical Analysis, 11 (2):340–347, March 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Friedman:1987:ICC

[FG87] Avner Friedman and Björn Gustafsson. Identification of the conductivity coefficient in an elliptic equation. SIAM Journal on Mathematical Analysis, 18(3): 777–787, May 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Feichtinger:1992:IRM

[FG92] Hans G. Feichtinger and Karlheinz Gröchenig. Iterative reconstruction of multivariate bandlimited functions from irregular sampling values. SIAM Journal on Mathematical Analysis, 23 (1):244–261, January 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fonseca:1995:LIS

[FG95] I. Fonseca and W. Gangbo. Local invertibility of Sobolev functions. SIAM Journal on Mathematical Analysis, 26(2): 280–304, March 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/25741.

Fiedler:1999:LFT

[FG99] Bernold Fiedler and Tomás Gedeon. A Lyapunov function for tridiagonal competitivecooperative systems. SIAM Journal on Mathematical Analysis, 30(3):469–478, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/31614.

Fink:1991:ASS

[FGHW91] A. M. Fink, Juan A. Gatica, Gaston E. Hernandez, and Paul Waltman. Approximation of solutions of singular second-order boundary value problems. SIAM Journal on Mathematical Analysis, 22(2):440–462, March 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fusco:1985:SES

[FH85] G. Fusco and J. K. Hale. Stable equilibria in a scalar parabolic equation with variable diffusion. SIAM Journal on Mathematical Analysis, 16(6):1152–1164, November 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fitouhi:1990:UEE

[FH90a] A. Fitouhi and M. M. Hamza. A uniform expansion for the eigenfunction of a singular secondorder differential operator. SIAM Journal on Mathematical Analysis, 21(6):1619–1632, November 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Friedman:1990:SCT

[FH90b] Avner Friedman and Michael L. Honig. On the spread of continuous-time linear systems. SIAM Journal on Mathematical Analysis, 21(3):757–770, May 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fusco:1991:PRP

[FH91] Nicola Fusco and J. Hutchinson. Partial regularity in problems motivated by nonlinear elasticity. SIAM Journal on Mathematical Analysis, 22(6):1516–1551, September 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Friedman:1996:SPM

[FH96] Avner Friedman and Bei Hu. A Stefan problem for multidimensional reaction-diffusion systems. SIAM Journal on Mathematical Analysis, 27(5):1212– 1234, September 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/27284.

Friedman:1999:SPP

[FH99] Avner Friedman and Bei Hu. A Stefan problem for a protocell model. SIAM Journal on Mathematical Analysis, 30 (4):912–926, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/33758.

Fitzgibbon:1997:SLF

[FHM97] W. B. Fitzgibbon, S. L. Hollis, and J. J. Morgan. Stability and Lyapunov functions for reactiondiffusion systems. SIAM Journal on Mathematical Analysis, 28 (3):595–610, May 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/27224.

Fusco:1996:TWL

[FHX96] Giorgio Fusco, Jack K. Hale, and Jianping Xun. Traveling waves as limits of solutions on bounded domains. SIAM Journal on Mathematical Analysis, 27(6): 1544–1558, November 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/27536.

Forristall:1972:EDU

[FI72] G. Z. Forristall and J. D. Ingram. Evaluation of distributions useful in Kontorovich–Lebedev transform theory. SIAM Journal on Mathematical Analysis, 3(4): 561–566, November 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fields:1975:PS

[FI75] Jerry L. Fields and Mourad El-Houssieny Ismail. On the positivity of some _{1}F_{2}’s. SIAM Journal on Mathematical Analysis, 6(3): 551–559, May 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fabiano:1990:STN

[FI90] R. H. Fabiano and K. Ito. Semigroup theory and numerical approximation for equations in lin ear viscoelasticity. SIAM Journal on Mathematical Analysis, 21 (2):374–393, March 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fokas:1996:LIB

[FI96] A. S. Fokas and A. R. Its. The linearization of the initialboundary value problem of the nonlinear Schrödinger equation. SIAM Journal on Mathematical Analysis, 27(3):738–764, May 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fialkow:1980:ZOE

[Fia80] Aaron Fialkow. The zeros of the odd and even parts of a Hurwitz polynomial. SIAM Journal on Mathematical Analysis, 11 (1):73–81, January 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fields:1973:UAE

[Fie73] Jerry L. Fields. Uniform asymptotic expansions of certain classes of Meijer G-functions for a large parameter. SIAM Journal on Mathematical Analysis, 4(3):

482–507, May 1973. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fields:1983:UAE

[Fie83] Jerry L. Fields. Uniform asymptotic expansions of a class of Meijer G-functions for a large parameter. SIAM Journal on Mathematical Analysis, 14(6):1204– 1253, November 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Field:1984:CTM

[Fie84] David A. Field. Convergence theorems for matrix continued fractions. SIAM Journal on Mathematical Analysis, 15(6):1220– 1227, November 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fiedler:1986:GHB

[Fie86] Bernold Fiedler. Global Hopf bifurcation for Volterra integral equations. SIAM Journal on Mathematical Analysis, 17(4): 911–932, July 1986. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fink:1974:CIF

[Fin74] A. M. Fink. Conjugate inequalities for functions and their derivatives. SIAM Journal on Mathematical Analysis, 5(3): 399–411, May 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fink:1977:MVI

[Fin77] A. M. Fink. Mean value inequalities. SIAM Journal on Mathematical Analysis, 8(5):785–791, October 1977. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fisher:1975:DML

[Fis75] Stephen D. Fisher. Duality methods for linear variational problems in L^{∞}. SIAM Journal on Mathematical Analysis, 6(5): 840–845, October 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fitzgibbon:1982:GEB

[Fit82] W. E. Fitzgibbon. Global existence and boundedness of solutions to the extensible beam equation. SIAM Journal on Mathematical Analysis, 13(5): 739–745, September 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Finkel:1987:ESI

[FIT87] Allan Finkel, Eli Isaacson, and Eugene Trubowitz. An explicit solution of the inverse periodic problem for Hill’s equation. SIAM Journal on Mathematical Analysis, 18(1):46–53, January 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Feinerman:1978:DOSa

[FK78] Robert P. Feinerman and Robert B. Kelman. Dual orthogonal series with oscillatory modifiers. SIAM Journal on Mathematical Analysis, 9(4):591–594, August 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fujimoto:1988:API

[FK88] Takao Fujimoto and Ulrich Krause. Asymptotic properties for inhomogeneous iterations of nonlinear operators. SIAM Journal on Mathematical Analysis, 19(4):841–853, July 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fujimoto:1994:SII

[FK94] Takao Fujimoto and Ulrich Krause. Stable inhomogeneous iterations of nonlinear positive operators on Banach spaces. SIAM Journal on Mathematical Analysis, 25(4):1195–1202, July 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/22165.

Fabes:1999:ICP

[FKS99] Eugene Fabes, Hyeonbae Kang, and Jin Keun Seo. Inverse conductivity problem with one measurement: Error estimates and approximate identification for perturbed disks. SIAM Journal on Mathematical Analysis, 30(4):699–720, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32495.

Friedman:1987:BTS

[FL87] Avner Friedman and Andrew A. Lacey. The blow-up time for solutions of nonlinear heat equations with small diffusion. SIAM Journal on Mathematical Analysis, 18 (3):711–721, May 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fabrie:1992:MAM

[FL92] Pierre Fabrie and Michel Langlais. Mathematical analysis of miscible displacement in porous medium. SIAM Journal on Mathematical Analysis, 23(6): 1375–1392, November 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fonseca:1998:BCE

[FL98] Irene Fonseca and Giovanni Leoni. Bulk and contact energies: Nucleation and relaxation. SIAM Journal on Mathematical Analysis, 30(1):190–219, 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/32556.

Farr:1989:DHB

[FLLL89] W. W. Farr, Chengzhi Li, I. S. Labouriau, and W. F. Langford. Degenerate Hopf bifurcation formulas and Hilbert’s 16th problem. SIAM Journal on Mathematical Analysis, 20(1):13–30, January 1989. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Flores:1991:SAS

[Flo91] Gilberto Flores. Stability analysis for the slow traveling pulse of the FitzHugh–Nagumo system. SIAM Journal on Mathematical Analysis, 22(2):392–399, March 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fonseca:1992:QCI

[FM92] Irene Fonseca and Stefan Müller.

Quasi-convex integrands and lower semicontinuity in L^{1}. SIAM Journal on Mathematical Analysis, 23(5):1081–1098, September 1992. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Flucher:1998:RSD

[FM98] M. Flucher and S. Müller. Radial symmetry and decay rate of variational ground states in the zero mass case. SIAM Journal on Mathematical Analysis, 29 (3):712–719, May 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/31402.

Fonseca:1999:QLS

[FM99] Irene Fonseca and Stefan Müller.

A-quasiconvexity, lower semicontinuity, and Young measures. SIAM Journal on Mathematical Analysis, 30(6):1355– 1390, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/33988.

Fasano:1986:PPI

[FMP86] A. Fasano, G. H. Meyer, and M. Primicerio. On a problem in the polymer industry: theoretical and numerical investigation of swelling. SIAM Journal on Mathematical Analysis, 17(4): 945–960, July 1986. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fonseca:1998:ACO

[FMP98] Irene Fonseca, Stefan Müller, and Pablo Pedregal. Analysis of concentration and oscillation effects generated by gradients. SIAM Journal on Mathematical Analysis, 29(3): 736–756, May 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/30653.

FinziVita:1996:QLR

[FMT96] Stefano Finzi Vita, François Murat, and Nicoletta A. Tchou. Quasi-linear relaxed Dirichlet problems. SIAM Journal on Mathematical Analysis, 27(4): 977–996, July 1996. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

                                        http://epubs.siam.org/sam-
                                        bin/dbq/article/26615.

Forte:1998:CGI

[FMV98] B. Forte, F. Mendivil, and E. R. Vrscay. “Chaos games” for iterated function systems with grey level maps. SIAM Journal on Mathematical Analysis, 29 (4):878–890, July 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30691.

Fonda:1993:SSS

[FMZ93] Alessandro Fonda, Raul Manasevich, and Fabio Zanolin. Subharmonic solutions for some secondorder differential equations with singularities. SIAM Journal on Mathematical Analysis, 24(5): 1294–1311, September 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Feinerman:1975:DTS

[FN75] Robert P. Feinerman and Donald J. Newman. Dual trigonometric series — a related problem. SIAM Journal on Mathematical Analysis, 6(6):989–997, November 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Forrester:1990:TFG

[For90] P. J. Forrester. Theta function generalizations of some constant term identities in the theory of random matrices. SIAM Journal on Mathematical Analysis, 21 (1):270–280, January 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fox:1972:LBE

[Fox72] David W. Fox. Lower bounds for eigenvalues with displacement of essential spectra. SIAM Journal on Mathematical Analysis, 3(4):

617–624, November 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fasano:1986:PHF

[FP86] Antonio Fasano and Mario Primicerio. A parabolic-hyperbolic free boundary problem. SIAM Journal on Mathematical Analysis, 17(1):67–73, January 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Feng:1984:SPO

[FQ84] Y. Y. Feng and D. X. Qi. A sequence of piecewise orthogonal polynomials. SIAM Journal on Mathematical Analysis, 15(4): 834–844, July 1984. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Forrester:1986:EZC

[FR86a] P. J. Forrester and J. B. Rogers. Electrostatics and the zeros of the classical polynomials. SIAM Journal on Mathematical Analysis, 17(2):461–468, March 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Furkotter:1986:PSF

[FR86b] M. Fürkotter and H. M. Rodrigues. Periodic solutions of forced nonlinear second order equations: symmetry and bifurcations. SIAM Journal on Mathematical Analysis, 17(6):1319– 1331, November 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Furkotter:1991:SBP

[FR91] M. Fürkotter and H. M. Rodrigues. Symmetry and bifurcation to 2π-periodic solutions of nonlinear second-order equations with 2π/m-periodic forcings. SIAM Journal on Mathematical Analysis, 22(1):169–180, January 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Frankel:1994:FBP

[FR94] Michael L. Frankel and Victor Roytburd. A free boundary problem modeling thermal instabilities: Well-posedness. SIAM Journal on Mathematical Analysis, 25(5):1357–1374, September 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/23322.

Franklin:1978:MPI

[Fra78] Joel N. Franklin. Minimum principles for ill-posed problems. SIAM Journal on Mathematical Analysis, 9(4):638–650, August 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Francfort:1983:HLT

[Fra83] Gilles A. Francfort. Homogenization and linear thermoelasticity. SIAM Journal on Mathematical Analysis, 14(4):696–708, July 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Frank:1990:SPE

[Fra90] L. S. Frank. On a singularly perturbed eigenvalue problem in the theory of elastic rods. SIAM Journal on Mathematical Analysis, 21(5):1245–1263, September 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Frenzen:1987:EBA

[Fre87a] C. L. Frenzen. Error bounds for asymptotic expansions of the ratio of two gamma functions. SIAM Journal on Mathematical Analysis, 18(3):890–896, May 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Frenzen:1987:AEM

[Fre87b] C. L. Frenzen. On the asymptotic expansion of Mellin transforms. SIAM Journal on Mathematical Analysis, 18(1):273–282, January 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Frenzen:1990:EBU

[Fre90] C. L. Frenzen. Error bounds for a uniform asymptotic expansion of the Legendre function Q_{n}^{−m}(cosh z). SIAM Journal on Mathematical Analysis, 21 (2):523–535, March 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Frenzen:1992:EBA

[Fre92] C. L. Frenzen. Error bounds for the asymptotic expansion of the ratio of two gamma functions with complex argument. SIAM Journal on Mathematical Analysis, 23(2):505–511, March 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Friedland:1981:GKM

[Fri81a] Shmuel Friedland. A generalization of the Kreiss matrix theorem. SIAM Journal on Mathematical Analysis, 12(6):826–832, November 1981. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Friedman:1981:VIS

[Fri81b] Avner Friedman. Variational inequalities in sequential analysis. SIAM Journal on Mathematical Analysis, 12(3):385–397, May 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Friedman:1981:MSN

[Fri81c] Mark J. Friedman. Mathematical study of the nonlinear singular integral magnetic field equation. III. SIAM Journal on Mathematical Analysis, 12(4):536–540, July 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fridman:1995:URG

[Fri95] B. L. Fridman. A uniqueness result for a generalized Radon transform. SIAM Journal on Mathematical Analysis, 26(6): 1467–1472, November 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/25138.

Fryant:1978:IAG

[Fry78] Allan J. Fryant. Interpolation and approximation of generalized axisymmetric potentials. SIAM Journal on Mathematical Analysis, 9(5):906–914, October 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fryant:1991:IGS

[Fry91] Allan Fryant. Inductively generating the spherical harmonics. SIAM Journal on Mathematical Analysis, 22(1):268–271, January 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Friedman:1995:SPR

[FRZ95] Avner Friedman, David S. Ross, and Jianhua Zhang. A Stefan problem for a reaction-diffusion system. SIAM Journal on Mathematical Analysis, 26(5):1089– 1112, September 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/25585.

Fox:1977:EEM

[FS77] David W. Fox and James T. Stadter. An eigenvalue estimation method of weinberger and Weinstein’s intermediate problems. SIAM Journal on Mathematical Analysis, 8(3):491–503, May 1977. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Franke:1979:CMV

[FS79] John E. Franke and James F.

Selgrade. A computer method for verification of asymptotically stable periodic orbits. SIAM Journal on Mathematical Analysis, 10(3):614–628, May 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Finch:1983:CRD

[FS83] David V. Finch and Donald C.

Solmon. A characterization of the range of the divergent beam x-ray transform. SIAM Journal on Mathematical Analysis, 14(4): 767–771, July 1983. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Freedman:1989:PDS

[FS89] H. I. Freedman and Joseph W.- H. So. Persistence in discrete semidynamical systems. SIAM Journal on Mathematical Analysis, 20(4):930–938, July 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Freistuhler:1995:EBV

[FS95] H. Freistühler and P. Szmolyan.

Existence and bifurcation of viscous profiles for all intermediate magnetohydrodynamic shock waves. SIAM Journal on Mathematical Analysis, 26(1):112–128, January 1995. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

http://epubs.siam.org/sam-
bin/dbq/article/24736.

Fraile:1996:PPW

[FS96] José M. Fraile and José Sabina de Lis. Periodic and positive wave front solutions of semilinear diffusion equations. SIAM Journal on Mathematical Analysis, 27 (1):170–203, January 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/25784.

Friedlander:1999:RIT

[FSV99] Susan Friedlander, Walter Strauss, and Misha Vishik. Robustness of instability for the twodimensional Euler equations. SIAM Journal on Mathematical Analysis, 30(6):1343–1354, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/33927.

Fournier:1996:SEE

[FSW96] Gilles Fournier, Andrzej Szulkin, and Michel Willem. Semilinear elliptic equations in R^{N} with almost periodic or unbounded forcing term. SIAM Journal on Mathematical Analysis, 27(6): 1653–1660, November 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/27869.

Friedman:1988:CPM

[FT88] Avner Friedman and Athanassios E. Tzavaras. Combustion in a porous medium. SIAM Journal on Mathematical Analysis, 19(3): 509–519, May 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Foias:1994:AAA

[FT94] C. Foias and R. Temam. Approximation of attractors by algebraic or analytic sets. SIAM Journal on Mathematical Analysis, 25(5): 1269–1302, September 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/22455.

Fasano:1999:CMM

[FT99] A. Fasano and F. Talamucci.

A comprehensive mathematical model for a multispecies flow through ground coffee. SIAM Journal on Mathematical Analysis, 31(2):251–273, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/33669.

Foias:1988:SEF

[FTZ88] Ciprian Foiaş, Allen Tannenbaum, and George Zames. Some explicit formulae for the singular values of certain Hankel operators with factorizable symbol. SIAM Journal on Mathematical Analysis, 19(5):1081– 1089, September 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fuhrmann:1980:OMS

[Fuh80] Paul A. Fuhrmann. Operator measures, selfadjoint operators and dynamical systems. SIAM Journal on Mathematical Analysis, 11(4):737–757, July 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fujita:1992:EII

[Fuj92] Yasuhiro Fujita. Energy inequalities for integro-partial differential equations with Riemann– Liouville integrals. SIAM Journal on Mathematical Analysis, 23 (5):1182–1188, September 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fulling:1982:LGA

[Ful82] S. A. Fulling. The local geometric asymptotics of continuum eigenfunction expansions. I. Overview. SIAM Journal on Mathematical Analysis, 13(6):891–912, November 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Fulling:1983:LGA

[Ful83] S. A. Fulling. The local geometric asymptotics of continuum eigenfunction expansions. II. Onedimensional systems. SIAM Journal on Mathematical Analysis, 14(4):780–795, July 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Frenzen:1988:UAE

[FW88] C. L. Frenzen and R. Wong.

Uniform asymptotic expansions of Laguerre polynomials. SIAM Journal on Mathematical Analysis, 19(5):1232–1248, September 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Freedman:1992:PSS

[FW92] H. I. Freedman and Jian Hong Wu. Periodic solutions of singlespecies models with periodic delay. SIAM Journal on Mathematical Analysis, 23(3):689–701, May 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gruyer:1998:HE

[GA98] E. Le Gruyer and J. C. Archer.

Harmonious extensions. SIAM Journal on Mathematical Analysis, 29(1):279–292, January 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/29406.

Gabutti:1984:ETB

[Gab84] Bruno Gabutti. An existence theorem for a boundary value problem related to that of Falkner and Skan. SIAM Journal on Mathematical Analysis, 15(5): 943–956, September 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Galaktionov:1991:BPU

[Gal91] Victor A. Galaktionov. Best possible upper bound for blowup solutions of the quasilinear heat conduction equation with source. SIAM Journal on Mathematical Analysis, 22(5):1293–1302, September 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Galaktionov:1995:ASS

[Gal95] Victor A. Galaktionov. On asymptotic self-similar behaviour for a quasilinear heat equation: Single point blow-up. SIAM Journal on Mathematical Analysis, 26(3):675–693, May 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/22341.

Gao:1993:TBP

[Gao93] Wei-Zheng Gao. Threshold behavior and propagation for a differential- difference system. SIAM Journal on Mathematical Analysis, 24(1):89–115, January 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gardner:1981:CST

[Gar81] Robert A. Gardner. Comparison and stability theorems for reaction-diffusion systems. SIAM Journal on Mathematical Analysis, 12(4):603–616, July 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gardner:1987:SNC

[Gar87] Robert Gardner. Solutions of a nonlocal conservation law arising in combustion theory. SIAM Journal on Mathematical Analysis, 18(1):172–183, January 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Garvan:1988:BIA

[Gar88] F. G. Garvan. A beta integral associated with the root system G_{2}. SIAM Journal on Mathematical Analysis, 19(6):1462–1474, November 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Garvan:1990:PMM

[Gar90] F. G. Garvan. A proof of the MacDonald–Morris root system conjecture for F_{4}. SIAM Journal on Mathematical Analysis, 21(3): 803–821, May 1990. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gardner:1992:SHB

[Gar92] Robert Gardner. Stability and Hopf bifurcation of steady state solutions of a singularly perturbed reaction-diffusion system. SIAM Journal on Mathematical Analysis, 23(1):99–149, January 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gasper:1973:TCA

[Gas73] George Gasper. On two conjectures of Askey concerning normalized Hankel determinants for the classical polynomials. SIAM Journal on Mathematical Analysis, 4(3):508–513, May 1973. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gasper:1975:PIB

[Gas75] George Gasper. Positive integrals of Bessel functions. SIAM Journal on Mathematical Analysis, 6 (5):868–881, October 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gasper:1977:PSC

[Gas77] George Gasper. Positive sums of the classical orthogonal polynomials. SIAM Journal on Mathematical Analysis, 8(3):423–447, May 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gasper:1981:SFB

[Gas81] George Gasper. Summation formulas for basic hypergeometric series. SIAM Journal on Mathematical Analysis, 12(2):196–200, March 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gasper:1985:RLF

[Gas85] George Gasper. Rogers’ linearization formula for the continuous q-ultraspherical polynomials and quadratic transformation formulas. SIAM Journal on Mathematical Analysis, 16(5): 1061–1071, September 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gasper:1989:ECF

[Gas89] George Gasper. q-extensions of Clausen’s formula and of the inequalities used by de Branges in his proof of the Bieberbach, Robertson, and Milin conjectures. SIAM Journal on Mathematical Analysis, 20(4):1019– 1034, July 1989. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gatteschi:1987:NIZ

[Gat87] Luigi Gatteschi. New inequalities for the zeros of Jacobi polynomials. SIAM Journal on Mathematical Analysis, 18(6):1549–1562, November 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gautschi:1974:HMI

[Gau74a] Walter Gautschi. A harmonic mean inequality for the gamma function. SIAM Journal on Mathematical Analysis, 5(2): 278–281, April 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gautschi:1974:SMV

[Gau74b] Walter Gautschi. Some mean value inequalities for the gamma function. SIAM Journal on Mathematical Analysis, 5(2): 282–292, April 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gautschi:1989:ZPO

[Gau89] Walter Gautschi. On the zeros of polynomials orthogonal on the semicircle. SIAM Journal on Mathematical Analysis, 20(3): 738–743, May 1989. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gaudenzi:1990:SPT

[Gau90] Marcellino Gaudenzi. On the Sturm–Picone theorem for N th order differential equations. SIAM Journal on Mathematical Analysis, 21(4):980–994, July 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Granoff:1972:ASOa

[GB72] B. Granoff and N. Bleistein. Asymptotic solutions of a 6 th order differential equation with two turning points. Part I: Derivation by method of steepest descent. SIAM Journal on Mathematical Analysis, 3(1):45–57, February 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gomes:1995:SRC

[GC95] Sônia M. Gomes and Elsa Cortina. Some results on the convergence of sampling series based on convolution integrals. SIAM Journal on Mathematical Analysis, 26(5):1386–1402, September 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/25509.

Georgiou:1987:TAN

[Geo87] Tryphon T. Georgiou. A topological approach to Nevanlinna– Pick interpolation. SIAM Journal on Mathematical Analysis, 18 (5):1248–1260, September 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Geronimo:1994:STO

[Ger94] Jeffrey S. Geronimo. Scattering theory, orthogonal polynomials, and q-series. SIAM Journal on Mathematical Analysis, 25 (2):392–419, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23899.

Gideon:1977:SAE

[GG77] Rudy A. Gideon and John Gurland. Some alternative expansions for the distribution function of a noncentral chi-square random variable. SIAM Journal on Mathematical Analysis, 8 (1):100–110, February 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gamboa:1996:SSM

[GG96] F. Gamboa and E. Gassiat.

Sets of superresolution and the maximum entropy method on the mean. SIAM Journal on Mathematical Analysis, 27(4): 1129–1152, July 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23320.

Giga:1992:GEW

[GGI92] Yoshikazu Giga, Shun’ichi Goto, and Hitoshi Ishii. Global existence of weak solutions for interface equations coupled with diffusion equations. SIAM Journal on Mathematical Analysis, 23(4): 821–835, July 1992. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Goldstein:1975:SPP

[GH75] A. J. Goldstein and A. D. Hall. Solutions to a problem in power series reversion. SIAM Journal on Mathematical Analysis, 6(1): 192–198, February 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Goldberg:1979:HET

[GH79] Wallace Goldberg and Harry Hochstadt. On a Hill’s equation with two gaps in its spectrum. SIAM Journal on Mathematical Analysis, 10(5):1069– 1076, September 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Greenspan:1984:RRH

[GH84] Bernie Greenspan and Philip Holmes. Repeated resonance and homoclinic bifurcation in a periodically forced family of oscillators. SIAM Journal on Mathematical Analysis, 15(1):69–97, January 1984. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gingold:1986:GSS

[GH86] Harry Gingold and Po Fang Hsieh. Global simplification of a singularly perturbed almost diagonal system. SIAM Journal on Mathematical Analysis, 17(1):7– 18, January 1986. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gingold:1987:GAP

[GH87a] Harry Gingold and Po-Fang Hsieh. Global approximation of perturbed Hamiltonian differential equations with several turning points. SIAM Journal on Mathematical Analysis, 18(5): 1275–1293, September 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gyllenberg:1987:ADD

[GH87b] M. Gyllenberg and H. J. A. M.

Heijmans. An abstract delaydifferential equation modelling size dependent cell growth and division. SIAM Journal on Mathematical Analysis, 18(1):74–88, January 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gingold:1988:ASH

[GH88] Harry Gingold and Po-Fang Hsieh. Asymptotic solutions of a Hamiltonian system in intervals with several turning points. SIAM Journal on Mathematical Analysis, 19(5):1142–1150, September 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ghidaglia:1986:FDA

[Ghi86] J.-M. Ghidaglia. On the fractal dimension of attractors for viscous incompressible fluid flows. SIAM Journal on Mathematical Analysis, 17(5):1139–1157, September 1986. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ganser:1996:STD

[GHL96] Gary Ganser, Xiaoping Hu, and Dening Li. Solutions for a twodimensional hyperbolic-elliptic coupled system. SIAM Journal on Mathematical Analysis, 27 (4):1024–1037, July 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/26465.

Glassey:1991:IDT

[GHLS91] Robert Glassey, Ernst Horst, Andrew Lenard, and Jack Schaeffer. On the inverse of the discrete two-dimensional wave operator at CFL. SIAM Journal on Mathematical Analysis, 22(6): 1763–1775, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gilbert:1970:CSB

[Gil70] R. P. Gilbert. The construction of solutions for boundary value problems by function theoretic methods. SIAM Journal on Mathematical Analysis, 1(1):96– 114, February 1970. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gillis:1975:IPL

[Gil75] J. Gillis. Integrals of products of Laguerre polynomials. SIAM Journal on Mathematical Analysis, 6(2):318–339, April 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gilding:1979:SFT

[Gil79] B. H. Gilding. Stabilization of flows through porous media. SIAM Journal on Mathematical Analysis, 10(2):237–246, March 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gimse:1993:CLD

[Gim93] Tore Gimse. Conservation laws with discontinuous flux functions. SIAM Journal on Mathematical Analysis, 24(2):279–289, March 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gingold:1978:MGB

[Gin78] H. Gingold. A method of global blockdiagonalization for matrixvalued functions. SIAM Journal on Mathematical Analysis, 9(6): 1076–1082, December 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gingold:1979:CTM

[Gin79] H. Gingold. On continuous triangularization of matrix functions. SIAM Journal on Mathematical Analysis, 10(4):709–720, July 1979. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Gingold:1985:ADM

[Gin85] H. Gingold. An asymptotic decomposition method applied to multi-turning point problems. SIAM Journal on Mathematical Analysis, 16(1):7–27, January 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gillis:1990:APD

[GIO90a] J. Gillis, Mourad E. H. Ismail, and T. Offer. An asymptotic problem in derangement theory. SIAM Journal on Mathematical Analysis, 21(1):262–269, January 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Giovangigli:1990:NPL

[Gio90b] Vincent Giovangigli. Nonadiabatic plane laminar flames and their singular limits. SIAM Journal on Mathematical Analysis, 21 (5):1305–1325, September 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Giovangigli:1993:ETF

[Gio93] Vincent Giovangigli. An existence theorem for a free boundary problem of hypersonic flow theory. SIAM Journal on Mathematical Analysis, 24(3):571–582, May 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Golomb:1982:ECF

[GJ82] Michael Golomb and Joseph Jerome. Equilibria of the curvature functional and manifolds of nonlinear interpolating spline curves. SIAM Journal on Mathematical Analysis, 13(3):421–458, May 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gillis:1988:CII

[GJZ88] J. Gillis, J. Jedwab, and D. Zeilberger. A combinatorial interpretation of the integral of the product of Legendre polynomials. SIAM Journal on Mathematical Analysis, 19(6):1455–1461, November 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gonzalez:1977:TDN

[GK77] R. González and R. Kress. On the treatment of a Dirichlet– Neumann mixed boundary value problem for harmonic functions by an integral equation method. SIAM Journal on Mathematical Analysis, 8(3):504–517, May 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Golubitsky:1980:QSS

[GK80] M. Golubitsky and B. L. Keyfitz.

A qualitative study of the steadystate solutions for a continuous flow stirred tank chemical reactor. SIAM Journal on Mathematical Analysis, 11(2):316–339, March 1980. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Grimmer:1984:SER

[GK84] R. C. Grimmer and F. Kappel. Series expansions for resolvents of Volterra integrodifferential equations in Banach space. SIAM Journal on Mathematical Analysis, 15(3):595–604, May 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gesztesy:1987:TBS

[GK87] F. Gesztesy and G. Karner. On three-body scattering near thresholds. SIAM Journal on Mathematical Analysis, 18(4): 1064–1086, July 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gilbert:1971:ASE

[GL71] Robert P. Gilbert and Chi Yeung Lo. On the approximation of solutions of elliptic partial differential equations in two and three dimensions. SIAM Journal on Mathematical Analysis, 2(1):17– 30, February 1971. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Goldstein:1974:BSN

[GL74] Jerome A. Goldstein and Arthur Lubin. On bounded solutions of nonlinear differential equations in Hilbert space. SIAM Journal on Mathematical Analysis, 5 (5):837–840, October 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grace:1984:OTD

[GL84] S. R. Grace and B. S. Lalli. Oscillation theorems for damped differential equations of even order with deviating arguments. SIAM Journal on Mathematical Analysis, 15(2):308–316, March 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gasull:1990:LCC

[GL90] A. Gasull and J. Libre. Limit cycles for a class of Abel equations. SIAM Journal on Mathematical Analysis, 21(5):1235– 1244, September 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Glasser:1971:SS

[Gla71] M. Lawrence Glasser. The summation of series. SIAM Journal on Mathematical Analysis, 2(4): 595–600, November 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Glasser:1976:TAG

[Gla76] M. L. Glasser. Time to attain a given temperature at the center of a sphere due to radial heat flow. SIAM Journal on Mathematical Analysis, 7(3):384–385, May 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Glasser:1984:SUP

[Gla84] M. L. Glasser. Some useful properties of the Hilbert transform. SIAM Journal on Mathematical Analysis, 15(6):1228– 1230, November 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gyori:1991:COD

[GLP91] I. Györi, G. Ladas, and L. Pakula. Conditions for oscillation of difference equations with applications to equations with piecewise constant arguments. SIAM Journal on Mathematical Analysis, 22(3):769–773, May 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gohberg:1979:PTD

[GLR79] I. Gohberg, P. Lancaster, and L. Rodman. Perturbation theory for divisors of operator polynomials. SIAM Journal on Mathematical Analysis, 10(6):1161–1183, November 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grace:1984:OTN

[GLY84] S. R. Grace, B. S. Lalli, and C. C. Yeh. Oscillation theorems for nonlinear second order differential equations with a nonlinear damping term. SIAM Journal on Mathematical Analysis, 15(6): 1082–1093, November 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grace:1988:AOT

[GLY88] S. R. Grace, B. S. Lalli, and C. C. Yeh. Addendum: “Oscillation theorems for nonlinear second-order differential equations with a nonlinear damping term” [SIAM J. Math. Anal. 15 (1984), no. 6, 1082–1093; MR 85j:34071]. SIAM Journal on Mathematical Analysis, 19(5): 1252–1253, September 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Goldstein:1976:TIS

[GM76] Jerome A. Goldstein and Charles J.

Monlezun. Temporally inhomogeneous scattering theory. II: Approximation theory and second order equations. SIAM Journal on Mathematical Analysis, 7(2):276–290, April 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gutmann:1979:SIS

[GM79] R. L. Gutmann and A. N. Michel.

Stability of interconnected systems described by stochastic nonlinear Volterra integral equations. SIAM Journal on Mathematical Analysis, 10(2):217–236, March 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Golubitsky:1983:MLI

[GM83] Martin Golubitsky and Jerrold Marsden. The Morse lemma in infinite dimensions via singularity theory. SIAM Journal on Mathematical Analysis, 14(6): 1037–1044, November 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See remark [CGR88, BMS83].

Grossmann:1984:DHF

[GM84] A. Grossmann and J. Morlet. Decomposition of Hardy functions into square integrable wavelets of constant shape. SIAM Journal on Mathematical Analysis, 15(4): 723–736, July 1984. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gustafsson:1989:IIS

[GM89] B. Gustafsson and J. Mossino. Isoperimetric inequalities for the Stefan problem. SIAM Journal on Mathematical Analysis, 20(5): 1095–1108, September 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ghidaglia:1991:LBS

[GM91] J. M. Ghidaglia and A. Marzocchi. Longtime behaviour of strongly damped wave equations, global attractors and their dimension. SIAM Journal on Mathematical Analysis, 22(4): 879–895, July 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Goodman:1992:RED

[GM92] T. N. T. Goodman and Charles A. Micchelli. On refinement equations determined by Polya frequency sequences. SIAM Journal on Mathematical Analysis, 23(3): 766–784, May 1992. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gianni:1993:SET

[GM93] Roberto Gianni and Paola Mannucci. Some existence theorems for an N-dimensional parabolic equation with a discontinuous source term. SIAM Journal on Mathematical Analysis, 24(3): 618–633, May 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gupta:1994:WBA

[GM94] Dharma P. Gupta and David R.

Masson. Watson’s basic analogue of Ramanujan’s entry 40 and its generalization. SIAM Journal on Mathematical Analysis, 25 (2):429–440, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/22512.

Giga:1996:EPS

[GM96] Yoshikazu Giga and Noriko Mizoguchi. Existence of periodic solutions for equations of evolving curves. SIAM Journal on Mathematical Analysis, 27 (1):5–39, January 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/26583.

Geronimo:1983:NSC

[GN83] Jeffrey S. Geronimo and Paul G.

Nevai. Necessary and sufficient conditions relating the coefficients in the recurrence formula to the spectral function for orthogonal polynomials. SIAM Journal on Mathematical Analysis, 14(3):622–637, May 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gervois:1989:III

[GN89] A. Gervois and H. Navelet. Infinite integrals involving three spherical Bessel functions. SIAM Journal on Mathematical Analysis, 20(4):1006–1018, July 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Greenberg:1993:ASM

[GN93] J. M. Greenberg and Anne Nouri. Antiplane shearing motions of a visco-plastic solid. SIAM Journal on Mathematical Analysis, 24(4): 943–967, July 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gao:1995:CAR

[GNP95] Bo Gao, Donald J. Newman, and V. A. Popov. Convex approximation by rational functions. SIAM Journal on Mathematical Analysis, 26(2):488–499, March 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/23285.

Gautesen:1971:SIE

[GO71] A. K. Gautesen and W. E. Olmstead. On the solution of the integral equation for the potential of two strips. SIAM Journal on Mathematical Analysis, 2 (3):293–306, August 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gomes:1986:SNE

[Gom86] Sônia M. Gomes. On a singular nonlinear elliptic problem. SIAM Journal on Mathematical Analysis, 17(6):1359–1369, November 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gordon:1975:MAE

[Gor75a] Noam Gordon. Matched asymptotic expansion solutions of nonlinear partial differential equations with a small parameter. SIAM Journal on Mathematical Analysis, 6(6):1007–1016, November 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gordon:1975:SMO

[Gor75b] Paul Gordon. Summability methods obtained by substitution of power series. SIAM Journal on Mathematical Analysis, 6 (5):860–867, October 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gordon:1995:PSI

[Gor95] Michael K. Gordon. Perturbed scale-invariant initial value problems in one-dimensional dynamic elastoplasticity. SIAM Journal on Mathematical Analysis, 26(6): 1564–1587, November 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/25642.

Goubet:1992:CAI

[Gou92] Olivier Goubet. Construction of approximate inertial manifolds using wavelets. SIAM Journal on Mathematical Analysis, 23(6): 1455–1481, November 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Granovsky:1984:CSM

[GP84] B. L. Granovsky and Eli Passow. Chebyshev systems of minimal degree. SIAM Journal on Mathematical Analysis, 15(1):166–169, January 1984. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Greco:1993:CSS

[GP93] Antonio Greco and Giovanni Porru. Convexity of solutions to some elliptic partial differential equations. SIAM Journal on Mathematical Analysis, 24(4): 833–839, July 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Giorgi:1999:BSD

[GP99] T. Giorgi and D. Phillips. The breakdown of superconductivity due to strong fields for the Ginzburg–Landau model. SIAM Journal on Mathematical Analysis, 30(2):341–359, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/32316.

Gurtin:1991:HTE

[GPG91] Morton E. Gurtin and Paolo Podio-Guidugli. A hyperbolic theory for the evolution of plane curves. SIAM Journal on Mathematical Analysis, 22(3):575–586, May 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grassmann:1975:ECS

[GR75] E. G. Grassmann and J. Rokne.

An explicit calculation of some sets of minimal capacity. SIAM Journal on Mathematical Analysis, 6(2):242–249, April 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gingold:1976:DEM

[GR76] H. Gingold and S. Rosenblat.

Differential equations with moving singularities. SIAM Journal on Mathematical Analysis, 7(6): 942–957, November 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grassmann:1978:CSE

[GR78] E. Grassmann and J. Rokne. Calculation of some extremal conformal mappings. SIAM Journal on Mathematical Analysis, 9(1):87– 105, February 1978. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grassmann:1979:CEP

[GR79] Eckhard Grassmann and Jon Rokne. Calculation of extremum problems for univalent functions. SIAM Journal on Mathematical Analysis, 10(4): 850–874, July 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gasper:1983:PPK

[GR83] George Gasper and Mizan Rahman. Positivity of the Poisson kernel for the continuous q-ultraspherical polynomials. SIAM Journal on Mathematical Analysis, 14(2):409–420, March 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gasper:1984:PFW

[GR84] George Gasper and Mizan Rahman. Product formulas of Watson, Bailey and Bateman types and positivity of the Poisson kernel for q-Racah polynomials. SIAM Journal on Mathematical Analysis, 15(4):768–789, July 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gupta:1985:HFS

[GR85] Rameshwar D. Gupta and Donald St. P. Richards. Hypergeometric functions of scalar matrix argument are expressible in terms of classical hypergeometric functions. SIAM Journal on Mathematical Analysis, 16(4): 852–858, July 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gasper:1986:PPK

[GR86] George Gasper and Mizan Rahman. Positivity of the Poisson kernel for the continuous q-Jacobi polynomials and some quadratic transformation formulas for basic hypergeometric series. SIAM Journal on Mathematical Analysis, 17(4):970–999, July 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gasper:1989:NCF

[GR89] George Gasper and Mizan Rahman. A nonterminating q-Clausen formula and some related product formulas. SIAM Journal on Mathematical Analysis, 20(5):1270–1282, September 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gimse:1992:SCP

[GR92] Tore Gimse and Nils Henrik Risebro. Solution of the Cauchy problem for a conservation law with a discontinuous flux function. SIAM Journal on Mathematical Analysis, 23(3):635–648, May 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Granoff:1972:ASOb

[Gra72] B. Granoff. Asymptotic solutions of a 6 th order differential equation with two turning points. Part II: Derivation by reduction to a first order system. SIAM Journal on Mathematical Analysis, 3(1):93–104, February 1972. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Granata:1980:CFD

[Gra80] Antonio Granata. Canonical factorizations of disconjugate differential operators. SIAM Journal on Mathematical Analysis, 11 (1):160–172, January 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Granata:1988:CFD

[Gra88] Antonio Granata. Canonical factorization of disconjugate differential operators — Part II. SIAM Journal on Mathematical Analysis, 19(5):1162–1173, September 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grant:1995:SMO

[Gra95] Christopher P. Grant. Slow motion in one-dimensional Cahn– Morral systems. SIAM Journal on Mathematical Analysis, 26 (1):21–34, January 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/22605.

Grant:1998:IBC

[Gra98] Christopher P. Grant. Interior blowup in a convection-diffusion equation. SIAM Journal on Mathematical Analysis, 29(6): 1447–1458, November 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/32745.

Greville:1974:PFI

[Gre74] T. N. E. Greville. On a problem of E. L. De Forest in iterated smoothing. SIAM Journal on Mathematical Analysis, 5(3): 376–398, May 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gregory:1980:NMA

[Gre80] R. D. Gregory. A note on multiple asymptotic series. SIAM Journal on Mathematical Analysis, 11(1):115–118, January 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Greenberg:1989:SSN

[Gre89] Leon Greenberg. Sturm sequences for nonlinear eigenvalue problems. SIAM Journal on Mathematical Analysis, 20(1): 182–199, January 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Greenberg:1991:OMF

[Gre91] Leon Greenberg. An oscillation method for fourth-order, self-adjoint, two-point boundary value problems with nonlinear eigenvalues. SIAM Journal on Mathematical Analysis, 22 (4):1021–1042, July 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gripenberg:1978:ERN

[Gri78] Gustaf Gripenberg. An existence result for a nonlinear Volterra integral equation in a Hilbert space. SIAM Journal on Mathematical Analysis, 9(5):793–805, October 1978. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gripenberg:1979:FDC

[Gri79a] Gustaf Gripenberg. On a frequency domain condition used in the theory of Volterra equations. SIAM Journal on Mathematical Analysis, 10(4):839–843, July 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gripenberg:1979:AIE

[Gri79b] Gustaf Gripenberg. On an abstract integral equation. SIAM Journal on Mathematical Analysis, 10(5):1017–1021, September 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gripenberg:1980:NVE

[Gri80a] Gustaf Gripenberg. On nonlinear Volterra equations with nonintegrable kernels. SIAM Journal on Mathematical Analysis, 11(4): 668–682, July 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gripenberg:1980:ABR

[Gri80b] Gustaf Gripenberg. On the asymptotic behavior of resolvents of Volterra equations. SIAM Journal on Mathematical Analysis, 11(4):654–662, July 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gripenberg:1981:ASS

[Gri81a] Gustaf Gripenberg. Asymptotic solutions of some nonlinear Volterra integral equations. SIAM Journal on Mathematical Analysis, 12(4):595–602, July 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gripenberg:1981:IRS

[Gri81b] Gustaf Gripenberg. Integrability of resolvents of systems of Volterra equations. SIAM Journal on Mathematical Analysis, 12 (4):585–594, July 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gripenberg:1982:NSS

[Gri82] Gustaf Gripenberg. Nonexistence of smooth solutions for shearing flows in a nonlinear viscoelastic fluid. SIAM Journal on Mathematical Analysis, 13(6): 954–961, November 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gripenberg:1985:EPV

[Gri85] Gustaf Gripenberg. An eigenvalue problem for a Volterra integral operator with infinite delay. SIAM Journal on Mathematical Analysis, 16(3):541–547, May 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gripenberg:1987:PST

[Gri87] Gustaf Gripenberg. On periodic solutions of a thermostat equation. SIAM Journal on Mathematical Analysis, 18(3):694–702, May 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gripenberg:1989:SPS

[Gri89] Gustaf Gripenberg. Stability of periodic solutions in thermostat control. SIAM Journal on Mathematical Analysis, 20(6): 1459–1471, November 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gripenberg:1991:TSH

[Gri91] Gustaf Gripenberg. Thermostat systems having stable periodic solutions with short periods. SIAM Journal on Mathematical Analysis, 22(5):1332– 1348, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gripenberg:1993:UBW

[Gri93]  Gustaf Gripenberg.  Uncon-

ditional bases of wavelets for Sobolev spaces. SIAM Journal on Mathematical Analysis, 24 (4):1030–1042, July 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gripenberg:1996:SIS

[Gri96] Gustaf Gripenberg. On subdivision interpolation schemes. SIAM Journal on Mathematical Analysis, 27(2):578–593, March 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/25432.

Grossman:1976:PFA

[Gro76] Nathaniel Grossman. Polygamma functions of arbitrary order. SIAM Journal on Mathematical Analysis, 7(3):366–372, May 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Grossman:1977:EPF

[Gro77] Nathaniel Grossman. Errata:

Polygamma functions of arbitrary order. SIAM Journal on Mathematical Analysis, 8(5):922, October 1977. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gruendler:1985:EHO

[Gru85] Joseph Gruendler. The existence of homoclinic orbits and the method of Mel′nikov for systems in R^{n}. SIAM Journal on Mathematical Analysis, 16(5): 907–931, September 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gruendler:1992:HSA

[Gru92] Joseph Gruendler. Homoclinic solutions for autonomous dynamical systems in arbitrary dimension. SIAM Journal on Mathematical Analysis, 23(3):702–721, May 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gillis:1983:EMP

[GRZ83] J. Gillis, B. Reznick, and D. Zeilberger. On elementary methods in positivity theory. SIAM Journal on Mathematical Analysis, 14 (2):396–398, March 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gilbert:1977:DOC

[GS77] E. N. Gilbert and D. Slepian. Doubly orthogonal concentrated polynomials. SIAM Journal on Mathematical Analysis, 8(2): 290–319, April 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gudder:1980:GC

[GS80] S. Gudder and F. Schroeck. Generalized convexity. SIAM Journal on Mathematical Analysis, 11(6): 984–1001, November 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Glashoff:1981:AGT

[GS81] K. Glashoff and J. Sprekels. An application of Glicksberg’s theorem to set-valued integral equations arising in the theory of thermostats. SIAM Journal on Mathematical Analysis, 12(3): 477–486, May 1981. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gessel:1982:SEH

[GS82] Ira Gessel and Dennis Stanton. Strange evaluations of hypergeometric series. SIAM Journal on Mathematical Analysis, 13 (2):295–308, March 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Golubitsky:1986:SST

[GS86] Martin Golubitsky and Ian Stewart. Symmetry and stability in Taylor–Couette flow. SIAM Journal on Mathematical Analysis, 17(2):249–288, March 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Goldstein:1987:ADF

[GS87] Jerome A. Goldstein and James T.

Sandefur, Jr. An abstract d’Alembert formula. SIAM Journal on Mathematical Analysis, 18(3):842–856, May 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Goff:1989:BTO

[GS89] Stuart Goff and Donald F. St.

Mary. The Bohl transformation and oscillation of linear differential systems. SIAM Journal on Mathematical Analysis, 20 (1):215–221, January 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gasser:1993:GSP

[GS93] I. Gasser and P. Szmolyan. A geometric singular perturbation analysis of detonation and deflagration waves. SIAM Journal on Mathematical Analysis, 24(4): 968–986, July 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guo:1999:UOT

[GS99] Yan Guo and Walter A. Strauss. Unstable oscillatory-tail waves in collisionless plasmas. SIAM Journal on Mathematical Analysis, 30(5):1076–1114, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/33391.

Goodman:1994:RSV

[GSZ94] Jonathan Goodman, Anders Szepessy, and Kevin Zumbrun. A remark on the stability of viscous shock waves. SIAM Journal on Mathematical Analysis, 25(6): 1463–1467, November 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23964.

Ginibre:1989:USG

[GT89] J. Ginibre and Y. Tsutsumi. Uniqueness of solutions for the generalized Korteweg–de Vries equation. SIAM Journal on Mathematical Analysis, 20(6): 1388–1425, November 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gasper:1994:NMC

[GT94a] George Gasper and Walter Trebels. On necessary multiplier conditions for Laguerre expansions II. SIAM Journal on Mathematical Analysis, 25 (2):384–391, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/22668.

Giga:1994:GWS

[GT94b] Yoshikazu Giga and Shuji Takahashi. On global weak solutions of the nonstationary twophase Stokes flow. SIAM Journal on Mathematical Analysis, 25 (3):876–893, May 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23191.

Gingold:1998:SFS

[GT98] Harry Gingold and Alexander Tovbis. Solutions of finitely smooth nonlinear singular differential equations and problems of diagonalization and triangularization. SIAM Journal on Mathematical Analysis, 29 (3):757–778, May 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30771.

Goldberg:1977:NPR

[GU77] J. L. Goldberg and J. L. Ullman.

A note on positive real functions. SIAM Journal on Mathematical Analysis, 8(5):757–762, October 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gu:1992:EGC

[Gu92] Cai Xing Gu. Eliminating the genericity conditions in the skew Toeplitz operator algorithm for H^{∞} optimization. SIAM Journal on Mathematical Analysis, 23(6): 1623–1636, November 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guckenheimer:1984:MBP

[Guc84] John Guckenheimer. Multiple bifurcation problems of codimension two. SIAM Journal on Mathematical Analysis, 15(1):1– 49, January 1984. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Guggenheimer:1971:GTD

[Gug71] H. Guggenheimer. Geometric theory of differential equations. I: Second order linear equations. SIAM Journal on Mathematical Analysis, 2(2):233–241, May 1971. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gunderson:1973:RDI

[Gun73] Robert W. Gunderson. A result on differential inequalities and its application to higher order trajectory derivatives. SIAM Journal on Mathematical Analysis, 4(4):557–560, November 1973. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gustafsson:1985:AVI

[Gus85] Björn Gustafsson. Applications of variational inequalities to a moving boundary problem for Hele–Shaw flows. SIAM Journal on Mathematical Analysis, 16 (2):279–300, March 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gustafson:1987:MST

[Gus87a] R. A. Gustafson. Multilateral summation theorems for ordinary and basic hypergeometric series in U(n). SIAM Journal on Mathematical Analysis, 18(6): 1576–1596, November 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gustafson:1987:WTH

[Gus87b] R. A. Gustafson. A Whipple’s transformation for hypergeometric series in U(n) and multivariable hypergeometric orthogonal polynomials. SIAM Journal on Mathematical Analysis, 18 (2):495–530, March 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gustafson:1990:STH

[Gus90] Robert A. Gustafson. A summation theorem for hypergeometric series very-well-poised on G_{2}. SIAM Journal on Mathematical Analysis, 21(2):510–522, March 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gustafson:1992:SBM

[Gus92] Robert A. Gustafson. Some qbeta and Mellin–Barnes integrals with many parameters associated to the classical groups. SIAM Journal on Mathematical Analysis, 23(2):525–551, March 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gustafson:1994:SBI

[Gus94] Robert A. Gustafson. Some q-beta integrals on SU(n) and Sp(n) that generalize the Askey– Wilson and Nasrallah–Rahman integrals. SIAM Journal on Mathematical Analysis, 25(2): 441–449, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24861.

Gustafsson:1998:MBC

[Gus98] Björn Gustafsson. On mother bodies of convex polyhedra. SIAM Journal on Mathematical Analysis, 29(5):1106–1117, September 1998. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

http://epubs.siam.org/sam-
bin/dbq/article/31791.

Gutman:1982:CPA

[Gut82] S. Gutman. Compact perturbations of m-accretive operators in general Banach spaces. SIAM Journal on Mathematical Analysis, 13(5):789–800, September 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Greenberg:1985:AMR

[Gv85] W. Greenberg and C. V. M. van der Mee. An abstract model for radiative transfer in an atmosphere with reflection by the planetary surface. SIAM Journal on Mathematical Analysis, 16(4): 695–702, July 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Galaktionov:1993:RBS

[GV93] Victor A. Galaktionov and Juan L. Vázquez. Regional blow up in a semilinear heat equation with convergence to a Hamilton– Jacobi equation. SIAM Journal on Mathematical Analysis, 24(5): 1254–1276, September 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Goh:1994:ATC

[GW94] William M. Y. Goh and Jet Wimp. On the asymptotics of the Tricomi–Carlitz polynomials and their zero distribution (I). SIAM Journal on Mathematical Analysis, 25(2): 420–428, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/22926.

Guan:1996:EBS

[GW96] Zhicheng Guan and Xu-Jia Wang. Existence and blow-up of solutions to two-phase nonequilibrium problems. SIAM Journal on Mathematical Analysis, 27 (4):1038–1048, July 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/25100.

Gorenflo:1997:SFC

[GW97] Norbert Gorenflo and Matthias Werner. Solution of a finite convolution equation with a Hankel kernel by matrix factorization. SIAM Journal on Mathematical Analysis, 28(2): 434–451, March 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/28915.

Giga:1990:DFB

[GY90] Yoshikazu Giga and Zensho Yoshida. A dynamic freeboundary problem in plasma physics. SIAM Journal on Mathematical Analysis, 21(5):1118– 1138, September 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gill:1992:DIS

[GZ92] Tepper L. Gill and W. W. Zachary. Dimensionality of invariant sets for nonautonomous processes. SIAM Journal on Mathematical Analysis, 23(5): 1204–1229, September 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gotz:1995:TPS

[GZ95] Ivan G. Götz and Boris Zaltzman. Two-phase Stefan problem with supercooling. SIAM Journal on Mathematical Analysis, 26 (3):694–714, May 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23605.

Ha:1986:EDP

[Ha86] Chung Wei Ha. Eigenvalues of differentiable positive definite kernels. SIAM Journal on Mathematical Analysis, 17 (2):415–419, March 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Haberman:1976:NPO

[Hab76] Richard Haberman. Nonlinear perturbations of the Orr– Sommerfeld equation — asymptotic expansion of the logarithmic phase shift across the critical layer. SIAM Journal on Mathematical Analysis, 7(1):70– 81, February 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Habsieger:1988:ISA

[Hab88]  Laurent Habsieger.  Une q-

intégrale de Selberg et Askey. (French) [A q-integral of Selberg and Askey]. SIAM Journal on Mathematical Analysis, 19(6): 1475–1489, November 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Haddock:1974:ABS

[Had74] John R. Haddock. On the asymptotic behavior of solutions of x′(t) = −a(t)f(x(t − r(t))). SIAM Journal on Mathematical Analysis, 5(4):569–573, August 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hagan:1982:TWM

[Hag82]  Patrick S. Hagan.  Travel-

ing wave and multiple traveling wave solutions of parabolic equations. SIAM Journal on Mathematical Analysis, 13(5):717–738, September 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hahner:1998:UTI

[Häh98] Peter Hähner. A uniqueness theorem for an inverse scattering problem in an exterior domain. SIAM Journal on Mathematical Analysis, 29(5):1118– 1128, September 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/31861.

Haimo:1980:HGT

[Hai80] Deborah Tepper Haimo. Homogeneous generalized temperatures. SIAM Journal on Mathematical Analysis, 11(3):473–487, May 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hainzl:1982:GCS

[Hai82] J. Hainzl. On a general concept for separation of variables. SIAM Journal on Mathematical Analysis, 13(2):208–225, March 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hainzl:1992:MBP

[Hai92] J. Hainzl. Multiparameter bifurcation of a predator-prey system.

SIAM Journal on Mathematical Analysis, 23(1):150–180, January 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hall:1977:RSD

[Hal77] L. M. Hall. Regular singular differential equations whose conjugate equation has polynomial solutions. SIAM Journal on Mathematical Analysis, 8(5):778–784, October 1977. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hall:1984:RZT

[Hal84] Glen Richard Hall. Resonance zones in two-parameter families of circle homeomorphisms. SIAM Journal on Mathematical Analysis, 15(6):1075–1081, November 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Halpern:1991:ABC

[Hal91] Laurence Halpern. Artificial boundary conditions for incompletely parabolic perturbations of hyperbolic systems. SIAM Journal on Mathematical Analysis, 22(5):1256–1283, September 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hannsgen:1973:VEP

[Han73]  Kenneth B. Hannsgen.  A

Volterra equation with parameter. SIAM Journal on Mathematical Analysis, 4(1):22–30, February 1973. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Hannsgen:1974:LVE

[Han74a] Kenneth B. Hannsgen. A linear Volterra equation in Hilbert space. SIAM Journal on Mathematical Analysis, 5(6):927–940, November 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hannsgen:1974:VEH

[Han74b] Kenneth B. Hannsgen. A Volterra equation in Hilbert space. SIAM Journal on Mathematical Analysis, 5(3):412–416, May 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hannsgen:1975:UBC

[Han75] Kenneth B. Hannsgen. Uniform boundedness in a class of Volterra equations. SIAM Journal on Mathematical Analysis, 6 (4):689–697, August 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hannsgen:1976:CPD

[Han76a] Kenneth B. Hannsgen. Continuous parameter dependence in a class of Volterra equations. SIAM Journal on Mathematical Analysis, 7(1):45–58, February 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hannsgen:1976:RKI

[Han76b] Kenneth B. Hannsgen. The resolvent kernel of an integrodifferential equation in Hilbert space. SIAM Journal on Mathematical Analysis, 7(4):481–490, August 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hannsgen:1977:UBI

[Han77] Kenneth B. Hannsgen. Uniform L^{1} behavior for an integrodifferential equation with parameter. SIAM Journal on Mathematical Analysis, 8(4):626–639, August 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hankerson:1989:EUT

[Han89] Darrel Hankerson. An existence and uniqueness theorem for difference equations. SIAM Journal on Mathematical Analysis, 20(5): 1208–1217, September 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Handelman:1990:SPP

[Han90] David E. Handelman. Stable positivity of polynomials obtained from three-term difference equations. SIAM Journal on Mathematical Analysis, 21(4):1051– 1064, July 1990. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Han:1999:ACO

[Han99] Bin Han. Analysis and construction of optimal multivariate biorthogonal wavelets with compact support. SIAM Journal on Mathematical Analysis, 31(2):274–304, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33641.

Harrow:1970:FRB

[Har70a] Martin Harrow. Further results on the boundedness and the stability of solutions of some differential equations of the fourth order. SIAM Journal on Mathematical Analysis, 1(2):189–194, May 1970. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Harrow:1970:BSS

[Har70b] Martin Harrow. On the boundedness and the stability of solutions of some differential equations of the fourth order. SIAM Journal on Mathematical Analysis, 1 (1):27–32, February 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Harbertson:1972:PNS

[Har72a] Noal C. Harbertson. Perturbations in nonlinear systems. SIAM Journal on Mathematical Analysis, 3(4):647–653, November 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hartman:1972:ESS

[Har72b] Philip Hartman. On the existence of similar solutions of some boundary layer problems. SIAM Journal on Mathematical Analysis, 3(1):120–147, February 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hartman:1974:TPB

[Har74]  Philip Hartman.  On two-

point boundary value problems for nonlinear second order systems. SIAM Journal on Mathematical Analysis, 5(2):172–177, April 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hartman:1977:PSS

[Har77] Philip Hartman. On the products of solutions of second order disconjugate differential equations and the Whittaker differential equation. SIAM Journal on Mathematical Analysis, 8(3): 558–571, May 1977. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hartman:1978:MPI

[Har78] Philip Hartman. Monotony properties and inequalities for Green’s functions for multipoint boundary value problems. SIAM Journal on Mathematical Analysis, 9 (5):806–814, October 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hartman:1979:SSN

[Har79a] Philip Hartman. Some singular nonoscillatory boundary value problems. SIAM Journal on Mathematical Analysis, 10(6): 1199–1219, November 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hartung:1979:NRP

[Har79b] Joachim Hartung. A note on restricted pseudoinverses. SIAM Journal on Mathematical Analysis, 10(2):266–273, March 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hartfiel:1980:CEU

[Har80] D. J. Hartfiel. Concerning the eigenvalues of DA under variations of the entries of D and A. SIAM Journal on Mathematical Analysis, 11(4):646–653, July 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hartman:1983:AIO

[Har83] Philip Hartman. Asymptotic integration of ordinary differential equations. SIAM Journal on Mathematical Analysis, 14(4): 772–779, July 1983. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Haraoka:1994:CFD

[Har94a] Yoshishige Haraoka. Canonical forms of differential equations free from accessory parameters. SIAM Journal on Mathematical Analysis, 25(4): 1203–1226, July 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/23108.

Haraoka:1994:MRS

[Har94b] Yoshishige Haraoka. Monodromy representations of systems of differential equations free from accessory parameters. SIAM Journal on Mathematical Analysis, 25(6):1595–1621, November 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/24222.

Hastings:1988:MTW

[Has88] S. P. Hastings. Multiple traveling waves in a combustion model. SIAM Journal on Mathematical Analysis, 19(6):1365–1371, November 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hayashi:1991:ASK

[Hay91] Nakao Hayashi. Analyticity of solutions of the Korteweg–De Vries equation. SIAM Journal on Mathematical Analysis, 22(6): 1738–1743, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Handelsman:1973:AEI

[HB73] R. A. Handelsman and N. Bleistein. Asymptotic expansions of integral transforms with oscillatory kernels: A generalization of the method of stationary phase. SIAM Journal on Mathematical Analysis, 4(3):519–535, May 1973. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Holman:1976:HSW

[HBL76] W. J. Holman, III, L. C. Biedenharn, and J. D. Louck. On hypergeometric series well-poised in SU(n). SIAM Journal on Mathematical Analysis, 7(4):529–541, August 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Heard:1982:APV

[Hea82]  Melvin L. Heard.  An ab-

stract parabolic Volterra integrodifferential equation. SIAM Journal on Mathematical Analysis, 13(1):81–105, January 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Healey:1988:GBC

[Hea88] Timothy J. Healey. Global bifurcations and continuation in the presence of symmetry with an application to solid mechanics. SIAM Journal on Mathematical Analysis, 19(4):824–840, July 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hegner:1981:LDS

[Heg81]  Stephen J. Hegner.  Linear

decomposable systems in continuous time. SIAM Journal on Mathematical Analysis, 12 (2):243–273, March 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Heinig:1975:SEH

[Hei75] Hans P. Heinig. Some extensions of Hardy’s inequality. SIAM Journal on Mathematical Analysis, 6(4):698–713, August 1975.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Heimes:1978:GFL

[Hei78] K. A. Heimes. Green’s functions for linear second order systems. SIAM Journal on Mathematical Analysis, 9(1):207–214, February 1978. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Henderson:1987:BIL

[Hen87a] Johnny Henderson. Best interval lengths for boundary value problems for third order Lipschitz equations. SIAM Journal on Mathematical Analysis, 18 (2):293–305, March 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Henrici:1987:TPT

[Hen87b] Peter Henrici. A triple product theorem for hypergeometric series. SIAM Journal on Mathematical Analysis, 18(6):1513– 1518, November 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Henderson:1989:ETB

[Hen89] Johnny Henderson. Existence theorems for boundary value problems for n th-order nonlinear difference equations. SIAM Journal on Mathematical Analysis, 20 (2):468–478, March 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hermes:1990:CMI

[Her90] Henry Hermes. Contact maps and induced differential equations. SIAM Journal on Mathematical Analysis, 21(3):660–672, May 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Herron:1991:ESF

[Her91] Isom H. Herron. Exchange of stabilities for flow along a concave wall. SIAM Journal on Mathematical Analysis, 22(5):1349– 1353, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Herve:1995:CRF

[Her95] Loïc Hervé. Construction et regularité des fonctions d’échelle (French) [Construction and regularity of scaling functions]. SIAM Journal on Mathematical Analysis, 26(5):1361–1385, September 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/24002.

Hethcote:1970:EBA

[Het70] Herbert W. Hethcote. Error bounds for asymptotic approximations of zeros of transcendental functions. SIAM Journal on Mathematical Analysis, 1(2): 147–152, May 1970. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hounie:1999:BDV

[HFL99] J. Hounie, M. C. Lopes Filho, and H. J. Nussenzveig Lopes. Bounds on the dispersion of vorticity in 2D incompressible, inviscid flows with A priori unbounded velocity. SIAM Journal on Mathematical Analysis, 31(1):134–153, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/33750.

Hale:1999:ASR

[HG99] Jack K. Hale and José Domingo Salazar González. Attractors of some reaction diffusion problems. SIAM Journal on Mathematical Analysis, 30(5):963–984, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32764.

Heidel:1972:EOS

[HH72] J. W. Heidel and Don B. Hinton. The existence of oscillatory solutions for a nonlinear differential equation. SIAM Journal on Mathematical Analysis, 3(2): 344–351, May 1972. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hsu:1988:ALD

[HH88] Sze Bi Hsu and Shin Feng Hwang. Analysis of large deformation of a heavy cantilever. SIAM Journal on Mathematical Analysis, 19(4):854–866, July 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hills:1989:IOD

[HI89] Norman L. Hills and John M. Irwin. Infinite-order differential equations and the heat equation. SIAM Journal on Mathematical Analysis, 20(2):430–438, March 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hilhorst:1982:NEP

[Hil82] D. Hilhorst. A nonlinear evolution problem arising in the physics of ionized gases. SIAM Journal on Mathematical Analysis, 13(1):16–39, January 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hill:1997:EHK

[Hil97] Adrian T. Hill. Estimates on the heat kernel of parabolic equations with advection. SIAM Journal on Mathematical Analysis, 28(6):1309–1316, November 1997. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/30104.

Hill:1998:ESN

[Hil98] Adrian T. Hill. L_{∞} estimates on the solutions of nonselfadjoint elliptic and parabolic equations in bounded domains. SIAM Journal on Mathematical Analysis, 29 (3):720–735, May 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/31015.

Hinton:1981:EES

[Hin81] Don Hinton. Eigenfunction expansions for a singular eigenvalue problem with eigenparameter in the boundary condition. SIAM Journal on Mathematical Analysis, 12(4):572–584, July 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hairer:1985:RAE

[HIN85] Ernst Hairer, Arieh Iserles, and Syvert P. Nørsett. Rational approximations to the exponential function with two complex conjugate interpolation points. SIAM Journal on Mathematical Analysis, 16(4):814–821, July 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hirschman:1972:IEC

[Hir72] I. I. Hirschman, Jr. Integral equations on certain compact homogeneous spaces. SIAM Journal on Mathematical Analysis, 3 (2):314–343, May 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hirschman:1975:ECI

[Hir75] I. I. Hirschman, Jr. On the eigenvalues of certain integral operators. SIAM Journal on Mathematical Analysis, 6(6):1024– 1050, November 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hirsch:1982:SDE

[Hir82] Morris W. Hirsch. Systems of differential equations which are competitive or cooperative. I. Limit sets. SIAM Journal on Mathematical Analysis, 13 (2):167–179, March 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hirano:1983:LET

[Hir83] Norimichi Hirano. Local existence theorems for nonlinear differential equations. SIAM Journal on Mathematical Analysis, 14 (1):117–125, January 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hirsch:1985:SDE

[Hir85] Morris W. Hirsch. Systems of differential equations that are competitive or cooperative. II. Convergence almost everywhere. SIAM Journal on Mathematical Analysis, 16(3):423–439, May 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hirano:1986:ANV

[Hir86] Norimichi Hirano. Abstract nonlinear Volterra equations with positive kernels. SIAM Journal on Mathematical Analysis, 17 (2):403–414, March 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hirsch:1990:SDE

[Hir90] Morris W. Hirsch. Systems of differential equations that are competitive or cooperative. IV: Structural stability in threedimensional systems. SIAM Journal on Mathematical Analysis, 21(5):1225–1234, September 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hively:1978:CNI

[Hiv78] G. A. Hively. On a class of nonlinear integral equations arising in transport theory. SIAM Journal on Mathematical Analysis, 9 (5):787–792, October 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hsiao:1983:NTP

[HJ83] George C. Hsiao and Kirk E.

Jordan. A numerical treatment for parabolic equations with a small parameter. SIAM Journal on Mathematical Analysis, 14(3): 507–521, May 1983. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hassard:1992:UPD

[HJ92] Brian Hassard and Katie Jiang.

Unfolding a point of degenerate Hopf bifurcation in an enzymecatalyzed reaction model. SIAM Journal on Mathematical Analysis, 23(5):1291–1304, September 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Han:1998:MRE

[HJ98] Bin Han and Rong-Qing Jia. Multivariate refinement equations and convergence of subdivision schemes. SIAM Journal on Mathematical Analysis, 29(5): 1177–1199, September 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29403.

Hagstrom:1986:EBC

[HK86] Thomas Hagstrom and H. B. Keller. Exact boundary conditions at an artificial boundary for partial differential equations in cylinders. SIAM Journal on Mathematical Analysis, 17 (2):322–341, March 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Han:1990:DPS

[HK90] H. Han and R. B. Kellogg. Differentiability properties of solutions of the equation ε^{2}Δu+ru = f(x, y) in a square. SIAM Journal on Mathematical Analysis, 21 (2):394–408, March 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hooker:1987:OSO

[HKP87] John W. Hooker, Man Kam Kwong, and William T. Patula. Oscillatory second order linear difference equations and Riccati equations. SIAM Journal on Mathematical Analysis, 18 (1):54–63, January 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hinton:1991:HBS

[HKS91] D. B. Hinton, M. Klaus, and J. K. Shaw. Half-bound states and Levinson’s theorem for discrete systems. SIAM Journal on Mathematical Analysis, 22(3): 754–768, May 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Handelsman:1970:AEL

[HL70] Richard A. Handelsman and John S. Lew. Asymptotic expansion of Laplace transforms near the origin. SIAM Journal on Mathematical Analysis, 1(1): 118–130, February 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Handelsman:1974:AEL

[HL74] Richard A. Handelsman and John S. Lew. Asymptotic expansion of Laplace convolutions for large argument and tail densities for certain sums of random variables. SIAM Journal on Mathematical Analysis, 5(3): 425–451, May 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hangelbroek:1977:DHS

[HL77] R. J. Hangelbroek and C. G.

Lekkerkerker. Decompositions of a Hilbert space and factorization of a W–A determinant. SIAM Journal on Mathematical Analysis, 8(3):458–472, May 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hazard:1993:DSF

[HL93] C. Hazard and M. Lenoir. Determination of scattering frequencies for an elastic floating body. SIAM Journal on Mathematical Analysis, 24(6):1458–1514, November 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hattori:1994:STD

[HL94] Harumi Hattori and De Ning Li. Solutions for two-dimensional system for materials of Korteweg type. SIAM Journal on Mathematical Analysis, 25 (1):85–98, January 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23413.

Hazard:1996:STH

[HL96] Christophe Hazard and Marc Lenoir. On the solution of timeharmonic scattering problems for Maxwell’s equations. SIAM Journal on Mathematical Analysis, 27(6):1597–1630, November 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/27125.

Hagstrom:1998:ATE

[HL98] Thomas Hagstrom and Jens Lorenz. All-time existence of classical solutions for slightly compressible flows. SIAM Journal on Mathematical Analysis, 29(3):652–672, May 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/31531.

Hallam:1972:ABF

[HLL72] T. G. Hallam, G. Ladas, and V. Lakshmikantham. On the asymptotic behavior of functional differential equations. SIAM Journal on Mathematical Analysis, 3(1):58–64, February 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hart:1986:MSN

[HLM86] D. C. Hart, A. C. Lazer, and P. J. McKenna. Multiplicity of solutions of nonlinear boundary value problems. SIAM Journal on Mathematical Analysis, 17(6): 1332–1338, November 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hastings:1992:BVP

[HLM92] S. P. Hastings, C. Lu, and A. D.

MacGillivray. A boundary value problem with multiple solutions from the theory of laminar flow. SIAM Journal on Mathematical Analysis, 23(1):201–208, January 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Handa:1980:BCS

[HM80] B. R. Handa and S. G. Mohanty. On q-binomial coefficients and some statistical applications. SIAM Journal on Mathematical Analysis, 11(6):1027– 1035, November 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hurley:1984:NAC

[HM84] M. Hurley and C. Martin. Newton’s algorithm and chaotic dynamical systems. SIAM Journal on Mathematical Analysis, 15 (2):238–252, March 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Heijmans:1989:SPS

[HM89a] H. J. A. M. Heijmans and J. A. J. Metz. Small parameters in structured population models and the Trotter–Kato theorem. SIAM Journal on Mathematical Analysis, 20(4):870–885, July 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hosono:1989:LCS

[HM89b] Yuzo Hosono and Masayasu Mimura. Localized cluster solutions of nonlinear degenerate diffusion equations arising in population dynamics. SIAM Journal on Mathematical Analysis, 20(4): 845–869, July 1989. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hollis:1987:GEB

[HMP87] Selwyn L. Hollis, Robert H. Martin, Jr., and Michel Pierre. Global existence and boundedness in reaction-diffusion systems. SIAM Journal on Mathematical Analysis, 18(3):744–761, May 1987. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Hayashi:1998:STH

[HNO98] Nakao Hayashi, Pavel I. Naumkin, and Tohru Ozawa. Scattering theory for the Hartree equation. SIAM Journal on Mathematical Analysis, 29(5):1256– 1267, September 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/31222.

Hanouzet:1998:CJL

[HNT98] Bernard Hanouzet, Roberto Natalini, and Alberto Tesei. On the Chapman–Jouguet limit for a combustion model. SIAM Journal on Mathematical Analysis, 29 (3):619–636, May 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29935.

Ho:1974:UTC

[Ho74] Chen-Kuang Ho. A uniqueness theorem for convolution equations in L^{p}(R^{n}) spaces. SIAM Journal on Mathematical Analysis, 5(3):463–468, May 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hayashi:1994:FES

[HO94] Nakao Hayashi and Tohru Ozawa. Finite energy solutions of nonlinear Schrödinger equations of derivative type. SIAM Journal on Mathematical Analysis, 25(6):

1488–1503, November 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/24612.

Hochstadt:1984:WII

[Hoc84] Harry Hochstadt. On the width of the instability intervals of the Mathieu equation. SIAM Journal on Mathematical Analysis, 15 (1):105–107, January 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hoff:1996:CDI

[Hof96] David Hoff. Continuous dependence on initial data for discontinuous solutions of the Navier–Stokes equations for onedimensional, compressible flow. SIAM Journal on Mathematical Analysis, 27(5):1193–1211, September 1996. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hofmann:1999:DTA

[Hof99] Bernd Hofmann. A denseness theorem with an application to a two-dimensional inverse potential refraction problem. SIAM Journal on Mathematical Analysis, 30(4):896–911, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33610.

Hogan:1998:NMR

[Hog98] Thomas A. Hogan. A note on matrix refinement equations.

SIAM Journal on Mathematical Analysis, 29(4):849–854, July 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/30135.

Holland:1977:SPF

[Hol77] Charles J. Holland. Singular perturbations in the first boundary value problem for parabolic equations. SIAM Journal on Mathematical Analysis, 8(2):368–374, April 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Holman:1980:STH

[Hol80] Wayne J. Holman, III. Summation theorems for hypergeometric series in U(n). SIAM Journal on Mathematical Analysis, 11(3): 523–532, May 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Holmaaker:1987:SMA

[Hol87] Kjell Holmåker. Some mathematical aspects on a problem of the optimal design of a vibrating beam. SIAM Journal on Mathematical Analysis, 18(5): 1367–1377, September 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Holmaaker:1993:GAS

[Hol93] Kjell Holmåker. Global asymptotic stability for a stationary solution of a system of integrodifferential equations describing the formation of liver zones.

SIAM Journal on Mathematical Analysis, 24(1):116–128, January 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Homentcovschi:1979:CMD

[Hom79] Dorel Homentcovschi. Conformal mapping of the domain exterior to a thin region. SIAM Journal on Mathematical Analysis, 10(6): 1246–1257, November 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Horton:1975:JPI

[Hor75] Robert Horton. Jacobi polynomials. IV: A family of variation diminishing kernel. SIAM Journal on Mathematical Analysis, 6(3): 544–550, May 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Howard:1974:PLP

[How74] F. T. Howard. Properties of Liron’s polynomials. SIAM Journal on Mathematical Analysis, 5(6):963–973, November 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Howes:1975:SPN

[How75] F. A. Howes. Singularly perturbed nonlinear boundary value problems with turning points. SIAM Journal on Mathematical Analysis, 6(4):644–660, August 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Howes:1978:ASC

[How78a] F. A. Howes. The asymptotic solution of a class of singularly perturbed nonlinear boundary value problems via differential inequalities. SIAM Journal on Mathematical Analysis, 9(2):215–249, April 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Howes:1978:SPN

[How78b] F. A. Howes. Singularly perturbed nonlinear boundary value problems with turning points. II. SIAM Journal on Mathematical Analysis, 9(2):250–271, April 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Howland:1981:QTT

[How81] Robert A. Howland, Jr. A quadratic transformation technique for the solution of nonlinear systems. SIAM Journal on Mathematical Analysis, 12 (1):90–103, January 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Howes:1982:DIH

[How82] F. A. Howes. Differential inequalities of higher order and the asymptotic solution of nonlinear boundary value problems. SIAM Journal on Mathematical Analysis, 13(1):61–80, January 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Howell:1985:DDA

[How85] Kenneth B. Howell. Directionally dependent asymptotic behavior of biharmonic functions with applications to elasticity. SIAM Journal on Mathematical Analysis, 16(4):822–847, July 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Howell:1986:ABP

[How86] Kenneth B. Howell. Asymptotic behavior of periodic strain states. SIAM Journal on Mathematical Analysis, 17(1):197–219, January 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Howes:1987:SLP

[How87] F. A. Howes. Shock layers in perturbed systems related to steady conservation laws. SIAM Journal on Mathematical Analysis, 18 (1):127–136, January 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Harrington:1973:NTM

[HP73] W. J. Harrington and K. A. Patel. Naylor transforms of Mellin type. SIAM Journal on Mathematical Analysis, 4(1):1– 7, February 1973. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hastings:1983:NPA

[HP83] S. P. Hastings and A. B. Poore. A nonlinear problem arising from combustion theory: Li nán’s problem. SIAM Journal on Mathematical Analysis, 14(3): 425–430, May 1983. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hastings:1985:LNP

[HP85] S. P. Hastings and A. B. Poore.

Liñán’s problem from combustion theory, Part II. SIAM Journal on Mathematical Analysis, 16 (2):331–340, March 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Horgan:1989:DEC

[HP89] C. O. Horgan and L. E. Payne.

Decay estimates for a class of nonlinear boundary value problems in two dimensions. SIAM Journal on Mathematical Analysis, 20(4):782–788, July 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Huang:1992:SEP

[HP92] Zhi Da Huang and Mario Primicerio. The surface evaporation problem with Signorini boundary condition. SIAM Journal on Mathematical Analysis, 23 (2):334–345, March 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hess:1993:BPP

[HP93] Peter Hess and Peter Poláčik.

Boundedness of prime periods of stable cycles and convergence to fixed points in discrete monotone dynamical systems. SIAM Journal on Mathematical Analysis, 24(5):1312–1330, September 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Henrot:1998:SOB

[HP98] Antoine Henrot and Gérard A. Philippin. Some overdetermined boundary value problems with elliptical free boundaries. SIAM Journal on Mathematical Analysis, 29(2):309–320, March 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/30721.

He:1997:AZR

[HPR97] Matthew He, K. Pan, and Paolo E. Ricci. Asymptotics of the zeros of relativistic Hermite polynomials. SIAM Journal on Mathematical Analysis, 28(5): 1248–1257, September 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29642.

Hall:1974:PIF

[HQW74] N. S. Hall, D. W. Quinn, and R. J. Weinacht. Poisson integral formulas in generalized biaxially symmetric potential theory. SIAM Journal on Mathematical Analysis, 5(1):111–118, February 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hrusa:1988:MEV

[HR88] William J. Hrusa and Michael Renardy. A model equation for viscoelasticity with a strongly singular kernel. SIAM Journal on Mathematical Analysis, 19 (2):257–269, March 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Heywood:1994:ST

[HR94] P. Heywood and P. G. Rooney.

On the Struve transformation. SIAM Journal on Mathematical Analysis, 25(2):450–461, March 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/22838.

Holden:1995:MMT

[HR95] Helge Holden and Nils Henrik Risebro. A mathematical model of traffic flow on a network of unidirectional roads. SIAM Journal on Mathematical Analysis, 26 (4):999–1017, July 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24328.

Holden:1999:RPK

[HR99] Helge Holden and Nils Henrik Risebro. Riemann problems with a kink. SIAM Journal on Mathematical Analysis, 30(3):497–515, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32703.

Hrabowski:1985:MHF

[Hra85] Jan Hrabowski. Multiple hypergeometric functions and simple Lie groups sl and sp. SIAM Journal on Mathematical Analysis, 16 (4):876–886, July 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hrusa:1985:GEA

[Hru85] William J. Hrusa. Global existence and asymptotic stability for a semilinear hyperbolic Volterra equation with large initial data. SIAM Journal on Mathematical Analysis, 16(1): 110–134, January 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

He:1998:GSC

[HRX98] Xue-Zhong He, Shigui Ruan, and Huaxing Xia. Global stability in chemostat-type equations with distributed delays. SIAM Journal on Mathematical Analysis, 29 (3):681–696, May 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/31110.

Hargrave:1977:LPL

[HS77] B. A. Hargrave and B. D. Sleeman. Lamé polynomials of large order. SIAM Journal on Mathematical Analysis, 8(5):800–842, October 1977. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hartwig:1979:DDI

[HS79] Robert E. Hartwig and Jim Shoaf. On the derivative of the Drazin inverse of a complex matrix. SIAM Journal on Mathematical Analysis, 10(1):207–216, January 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hinton:1984:BVP

[HS84a] D. B. Hinton and J. K. Shaw.

On boundary value problems for Hamiltonian systems with two singular points. SIAM Journal on Mathematical Analysis, 15 (2):272–286, March 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Holder:1984:BCM

[HS84b] E. J. Holder and D. Schaeffer.

Boundary conditions and mode jumping in the von Kármán equations. SIAM Journal on Mathematical Analysis, 15(3): 446–458, May 1984. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hinton:1986:ACS

[HS86a] D. B. Hinton and J. K. Shaw.

Absolutely continuous spectra of second order differential operators with short and long range potentials. SIAM Journal on Mathematical Analysis, 17(1): 182–196, January 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hoffmann:1986:IPG

[HS86b] K.-H. Hoffmann and J. Sprekels. On the identification of parameters in general variational inequalities by asymptotic regularization. SIAM Journal on Mathematical Analysis, 17(5):1198– 1217, September 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Halpern:1989:ABC

[HS89] Laurence Halpern and Michelle Schatzman. Artificial boundary conditions for incompressible viscous flows. SIAM Journal on Mathematical Analysis, 20 (2):308–353, March 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Haario:1994:RDG

[HS94] Heikki Haario and Thomas I. Seidman. Reaction and diffusion at a gas/liquid interface. II. SIAM Journal on Mathematical Analysis, 25(4): 1069–1084, July 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23471.

Hornung:1995:EPE

[HS95] Ulrich Hornung and R. E. Showalter. Elliptic-parabolic equations with hysteresis boundary conditions. SIAM Journal on Mathematical Analysis, 26 (4):775–790, July 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/22829.

Hsiao:1996:GES

[HS96] L. Hsiao and D. Serre. Global existence of solutions for the system of compressible adiabatic flow through porous media. SIAM Journal on Mathematical Analysis, 27(1):70–77, January 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/26707.

Habets:1997:TPP

[HS97] P. Habets and L. Sanchez. A twopoint problem with nonlinearity depending only on the derivative. SIAM Journal on Mathematical Analysis, 28(5):1205– 1211, September 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29570.

Heibig:1998:MCS

[HS98a] Arnaud Heibig and Amina Sahel. A method of characteristics for some systems of conservation laws. SIAM Journal on Mathematical Analysis, 29(6): 1467–1480, November 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/31035.

Huang:1998:EMS

[HS98b] Chaocheng Huang and Thomas Svobodny. Evolution of mixedstate regions in Type-II superconductors. SIAM Journal on Mathematical Analysis, 29 (4):1002–1021, July 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/31504.

Hsiao:1973:SPN

[Hsi73] George C. Hsiao. Singular perturbations for a nonlinear differential equation with a small parameter. SIAM Journal on Mathematical Analysis, 4(2):283–301, May 1973. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hsiao:1978:SPE

[Hsi78] George C. Hsiao. Singular perturbation of an exterior Dirichlet problem. SIAM Journal on Mathematical Analysis, 9(1): 160–184, February 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hsu:1983:SSS

[Hsu83] Sze Bi Hsu. Steady states of a system of partial differential equations modeling microbial ecology. SIAM Journal on Mathematical Analysis, 14(6): 1130–1138, November 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hsu:1989:STA

[Hsu89] Pei Hsu. Short-time asymptotics of the heat kernel on a concave boundary. SIAM Journal on Mathematical Analysis, 20(5):

1109–1127, September 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hsiao:1987:ABS

[HSX87] L. Hsiao, Y. Su, and Z. P. Xin.

On the asymptotic behavior of solutions of a reacting–diffusing system: a two predators–one prey model. SIAM Journal on Mathematical Analysis, 18(3): 647–669, May 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hunt:1971:AEC

[HT71] R. A. Hunt and M. H. Taibleson. Almost everywhere convergence of Fourier series on the ring of integers of a local field. SIAM Journal on Mathematical Analysis, 2(4):607–625, November 1971. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hastings:1987:OSF

[HT87] S. P. Hastings and W. C.

Troy. Oscillating solutions of the Falkner–Skan equation for negative β. SIAM Journal on Mathematical Analysis, 18(2):422–429, March 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hastings:1989:SCT

[HT89] S. P. Hastings and W. C. Troy.

On some conjectures of Turcotte, Spence, Bau, and Holmes. SIAM Journal on Mathematical Analysis, 20(3):634–642, May 1989.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hastings:1998:TAM

[HT98] Stuart P. Hastings and William C. Troy. There are asymmetric minimizers for the one-dimensional Ginzburg–Landau model of superconductivity. SIAM Journal on Mathematical Analysis, 30(1): 1–18, 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/30195.

Hu:1990:QVI

[Hu90] Bei Hu. A quasi-variational inequality arising in elastohydrodynamics. SIAM Journal on Mathematical Analysis, 21 (1):18–36, January 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hu:1991:DPP

[Hu91a] Bei Hu. Diffusion of penetrant in a polymer: A free boundary problem. SIAM Journal on Mathematical Analysis, 22(4): 934–956, July 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hu:1991:CPA

[Hu91b] Yingkang Hu. Convexity preserving approximation by free knot splines. SIAM Journal on Mathematical Analysis, 22(4):1183– 1191, July 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Huang:1996:NOR

[Hua96] Ying Huang. A nonlinear operator related to scaling functions and wavelets. SIAM Journal on Mathematical Analysis, 27(6): 1770–1790, November 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/25719.

Hughes:1975:SSC

[Hug75] O. F. Hughes. A simplification of the Schwarz–Christoffel formula for symmetric quadrilateral transformation. SIAM Journal on Mathematical Analysis, 6 (2):258–261, April 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hulshof:1997:SSS

[Hul97] Josephus Hulshof. Self-similar solutions of Barenblatt’s model for turbulence. SIAM Journal on Mathematical Analysis, 28 (1):33–48, January 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29003.

Hundhausen:1973:GOS

[Hun73] Joan Rohrer Hundhausen. A generating operator for solutions of certain partial difference and differential equations. SIAM Journal on Mathematical Analysis, 4 (1):15–21, February 1973. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hutson:1986:SRD

[Hut86]  V. Hutson.  Stability in a

reaction-diffusion model of mutualism. SIAM Journal on Mathematical Analysis, 17(1):58–66, January 1986. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Herrero:1987:ODN

[HV87] Miguel A. Herrero and Juan L. Vázquez. The one-dimensional nonlinear heat equation with absorption: regularity of solutions and interfaces. SIAM Journal on Mathematical Analysis, 18(1): 149–167, January 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Helton:1997:AFO

[HV97a] J. William Helton and Andrei E. Vityaev. Analytic functions optimizing competing constraints. SIAM Journal on Mathematical Analysis, 28(3): 749–767, May 1997. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29308.

Herrero:1997:MIB

[HV97b] Miguel A. Herrero and Juan J. L. Velázquez. On the melting of ice balls. SIAM Journal on Mathematical Analysis, 28 (1):1–32, January 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/28215.

Henry:1981:ATM

[HW81] M. S. Henry and K. Wiggins. Approximation theory methods for linear and nonlinear differential equations with deviating arguments. SIAM Journal on Mathematical Analysis, 12(3):342–354, May 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hannsgen:1982:CMR

[HW82] Kenneth B. Hannsgen and Robert L. Wheeler. Complete monotonicity and resolvents of Volterra integro-differential equations. SIAM Journal on Mathematical Analysis, 13(6): 962–969, November 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hsiao:1983:SPS

[HW83] George C. Hsiao and Richard J.

Weinacht. Singular perturbations for a semilinear hyperbolic equation. SIAM Journal on Mathematical Analysis, 14(6): 1168–1179, November 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hannsgen:1984:UBC

[HW84] Kenneth B. Hannsgen and Robert L. Wheeler. Uniform L^{1} behavior in classes of integrodifferential equations with completely monotonic kernels. SIAM Journal on Mathematical Analysis, 15(3):579–594, May 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hale:1989:PID

[HW89] Jack K. Hale and Paul Waltman. Persistence in infinitedimensional systems. SIAM Journal on Mathematical Analysis, 20(2):388–395, March 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Hu:1992:FBP

[HW92] Bei Hu and Lihe Wang. A free boundary problem arising in electrophotography: Solutions with connected toner region. SIAM Journal on Mathematical Analysis, 23(6):1439–1454, November 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Huang:1995:CPH

[HY95] Chao Cheng Huang and Jiong Min Yong. Coupled parabolic and hyperbolic equations modeling agedependent epidemic dynamics with nonlinear diffusion. SIAM Journal on Mathematical Analysis, 26(6):1588–1610, November 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/24445.

Hassard:1994:EHO

[HZ94] Brian Hassard and Jianhe Zhang. Existence of a homoclinic orbit of the Lorenz system by precise shooting. SIAM Journal on Mathematical Analysis, 25(1): 179–196, January 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23482.

IldefonsoDiaz:1994:NPP

[Id94] Jesus Ildefonso Diaz and François de Thélin. On a nonlinear parabolic problem arising in some models related to turbulent flows. SIAM Journal on Mathematical Analysis, 25(4): 1085–1111, July 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/21773.

Iffland:1987:PSP

[Iff87] Georges Iffland. Positive solution of a problem of Emden– Fowler type with a free boundary. SIAM Journal on Mathematical Analysis, 18(2):283–292, March 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ismail:1976:BPS

[IK76] Mourad E. H. Ismail and Douglas H. Kelker. The Bessel polynomials and the student t distribution. SIAM Journal on Mathematical Analysis, 7(1):82– 91, February 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ismail:1979:SFS

[IK79] Mourad E. H. Ismail and Douglas H. Kelker. Special functions, Stieltjes transforms and infinite divisibility. SIAM Journal on Mathematical Analysis, 10(5):

884–901, September 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Iwaniec:1996:PFN

[IL96] Tadeusz Iwaniec and Adam Lutoborski. Polyconvex functionals for nearly conformal deformations. SIAM Journal on Mathematical Analysis, 27(3):609–619, May 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ismail:1989:QBD

[ILV89] Mourad E. H. Ismail, Jean Letessier, and Galliano Valent. Quadratic birth and death processes and associated continuous dual Hahn polynomials. SIAM Journal on Mathematical Analysis, 20(3):727–737, May 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ismail:1978:MZC

[IM78] Mourad E. H. Ismail and Martin E. Muldoon. Monotonicity of the zeros of a cross-product of Bessel functions. SIAM Journal on Mathematical Analysis, 9 (4):759–767, August 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ismail:1987:GCP

[IM87] Mourad E. H. Ismail and Fuad S. Mulla. On the generalized Chebyshev polynomials. SIAM Journal on Mathematical Analysis, 18(1):243–258, January 1987.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ikeda:1991:SAS

[IM91] Hideo Ikeda and Masayasu Mimura. Stability analysis of stationary solutions of bistable reaction-variable diffusion systems. SIAM Journal on Mathematical Analysis, 22(6):1651– 1678, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ingerman:1998:CKD

[IM98] David Ingerman and James A.

Morrow. On a characterization of the kernel of the Dirichletto-Neumann map for a planar region. SIAM Journal on Mathematical Analysis, 29(1): 106–115, January 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30048.

Immink:1990:NRB

[Imm90] G. K. Immink. A note on the relationship between Stokes multipliers and formal solutions of analytic differential equations. SIAM Journal on Mathematical Analysis, 21(3):782–792, May 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Immink:1991:ABC

[Imm91a] G. K. Immink. On the asymptotic behavior of the coefficients of asymptotic power series and its relevance to Stokes phenomena. SIAM Journal on Mathematical Analysis, 22(2):524–542, March 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Immink:1991:RCF

[Imm91b] G. K. Immink. Reduction to canonical forms and the Stokes phenomenon in the theory of linear difference equations. SIAM Journal on Mathematical Analysis, 22(1):238–259, January 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Isaacson:1990:TWC

[IMP90] Eli L. Isaacson, Dan Marchesin, and Bradley J. Plohr. Transitional waves for conservation laws. SIAM Journal on Mathematical Analysis, 21(4):837–866, July 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Iannelli:1992:ANR

[IMP92] M. Iannelli, F. A. Milner, and A. Pugliese. Analytical and numerical results for the agestructured S–I–S epidemic model with mixed inter-intracohort transmission. SIAM Journal on Mathematical Analysis, 23(3): 662–688, May 1992. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Iserles:1990:ZTP

[IN90] A. Iserles and S. P. Nørsett. Zeros of transformed polynomials. SIAM Journal on Mathematical Analysis, 21(2):483–509, March 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Iooss:1989:HBP

[IR89] G. Iooss and M. Rossi. Hopf bifurcation in the presence of spherical symmetry: analytical results. SIAM Journal on Mathematical Analysis, 20(3):511–532, May 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Iraniparast:1989:GRF

[Ira89]  Nezam Iraniparast.  Green–

Riemann functions for a class of hyperbolic focal point problems. SIAM Journal on Mathematical Analysis, 20(2):408–414, March 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ishii:1998:RCC

[IS98] Katsuyuki Ishii and Halil Mete Soner. Regularity and convergence of crystalline motion. SIAM Journal on Mathematical Analysis, 30(1):19–37, 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/31734.

Iserles:1979:APA

[Ise79]  Arieh Iserles.  On the A-

acceptability of Padé approximations. SIAM Journal on Mathematical Analysis, 10(5):1002– 1007, September 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Iserles:1985:OSAa

[Ise85a] A. Iserles. Order stars, approximations and finite differences. I. The general theory of order stars. SIAM Journal on Mathematical Analysis, 16(3):559–576, May 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Iserles:1985:OSAb

[Ise85b] A. Iserles. Order stars, approximations and finite differences. II. Theorems in approximation theory. SIAM Journal on Mathematical Analysis, 16(4):785–802, July 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Iserles:1985:OSAc

[Ise85c] A. Iserles. Order stars, approximations and finite differences. III finite differences for u_{t} = ωu_{xx}. SIAM Journal on Mathematical Analysis, 16(5): 1020–1033, September 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ishige:1996:GTP

[Ish96a] Kazuhiro Ishige. The gradient theory of the phase transitions in Cahn–Hilliard fluids with Dirichlet boundary conditions. SIAM Journal on Mathematical Analysis, 27(3):620–637, May 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ishige:1996:ESC

[Ish96b] Kazuhiro Ishige. On the existence of solutions of the Cauchy problem for a doubly nonlinear parabolic equation. SIAM Journal on Mathematical Analysis, 27(5):1235–1260, September 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/27037.

Ismail:1974:OGF

[Ism74] Mourad El-Houssieny Ismail. On obtaining generating functions of Boas and Buck type for orthogonal polynomials. SIAM Journal on Mathematical Analysis, 5 (2):202–208, April 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ismail:1981:BBF

[Ism81] Mourad E. H. Ismail. The basic Bessel functions and polynomials. SIAM Journal on Mathematical Analysis, 12(3):454–468, May 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ismail:1985:SOP

[Ism85]  Mourad E. H. Ismail.  On

sieved orthogonal polynomials. I. Symmetric Pollaczek analogues. SIAM Journal on Mathematical Analysis, 16(5):1093–1113, September 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ismail:1986:AAW

[Ism86] Mourad E. H. Ismail. Asymptotics of the Askey–Wilson and q-Jacobi polynomials. SIAM Journal on Mathematical Analysis, 17 (6):1475–1482, November 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ismail:1994:APP

[Ism94] Mourad E. H. Ismail. Asymptotics of Pollaczek polynomials and their zeros. SIAM Journal on Mathematical Analysis, 25 (2):462–473, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/22029.

Ismail:1979:CAS

[IT79] Mourad E. H. Ismail and M. V. Tamhankar. A combinatorial approach to some positivity problems. SIAM Journal on Mathematical Analysis, 10(3):478–485, May 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ito:1994:GES

[Ito94] Kazuo Ito. On a globally existing solution to the inviscid Burgers equation with a nonlocal term. SIAM Journal on Mathematical Analysis, 25(4): 1112–1121, July 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23487.

Iaia:1998:EVS

[IW98] Joseph Iaia and Henry Warchall. Encapsulated-vortex solutions to equivariant wave equations: Existence. SIAM Journal on Mathematical Analysis, 30(1):118–139, 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/31692.

Iwasaki:1988:RFH

[Iwa88] Katsunori Iwasaki. Riemann function of harmonic equation and Appell’s F_{4}. SIAM Journal on Mathematical Analysis, 19(4): 902–917, July 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jaffard:1997:MFFa

[Jaf97a] S. Jaffard. Multifractal formalism for functions. Part I: Results valid for all functions. SIAM Journal on Mathematical Analysis, 28(4):944–970, July 1997. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/28299.

Jaffard:1997:MFFb

[Jaf97b] S. Jaffard. Multifractal formalism for functions. Part II: Selfsimilar functions. SIAM Journal on Mathematical Analysis, 28 (4):971–998, July 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/28300.

Jain:1981:STB

[Jai81] V. K. Jain. Some transformations of basic hypergeometric functions. Part II. SIAM Journal on Mathematical Analysis, 12(6): 957–961, November 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jodar:1988:RMR

[JAK88] L. Jódar and H. Abou-Kandil. A resolution method for Riccati differential systems coupled in their quadratic terms. SIAM Journal on Mathematical Analysis, 19(6): 1425–1430, November 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

James:1998:CRS

[Jam98] François James. Convergence results for some conservation laws with a reflux boundary condition and a relaxation term arising in chemical engineering. SIAM Journal on Mathematical Analysis, 29(5):1200– 1223, September 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/30793.

Janssen:1981:PWW

[Jan81] A. J. E. M. Janssen. Positivity of weighted Wigner distributions. SIAM Journal on Mathematical Analysis, 12(5):752–758, September 1981. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Janssen:1984:NHT

[Jan84] A. J. E. M. Janssen. A note on Hudson’s theorem about functions with nonnegative Wigner distributions. SIAM Journal on Mathematical Analysis, 15(1): 170–176, January 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Janssen:1986:EPU

[Jan86] Rainer Janßen. Elliptic problems on unbounded domains. SIAM Journal on Mathematical Analysis, 17(6):1370–1389, November 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Janfalk:1996:BLM

[Jan96] Ulf Janfalk. Behaviour in the limit, as p → ∞, of minimizers of functionals involving p-Dirichlet integrals. SIAM Journal on Mathematical Analysis, 27 (2):341–360, March 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/25261.

Jayne:1988:RGP

[Jay88] John W. Jayne. Recursively generated polynomials and Geronimus’ version of orthogonality on a contour. SIAM Journal on Mathematical Analysis, 19(3): 676–686, May 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Chen:1999:NSC

[jCcW99] Kuan ju Chen and Hwai chiuan Wang. A necessary and sufficient condition for Palais– Smale conditions. SIAM Journal on Mathematical Analysis, 31(1):154–165, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33801.

Jeffries:1991:SCS

[Jef91] John S. Jeffries. A strongly coupled singularly perturbed quasilinear second-order system. SIAM Journal on Mathematical Analysis, 22(4):896–904, July 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jensen:1991:IR

[Jen91] S. Jensen. An H_{0}^{m} interpolation result. SIAM Journal on Mathematical Analysis, 22(3):785–791, May 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jerome:1978:EAW

[Jer78] Joseph W. Jerome. Existence and approximation of weak solutions of nonlinear Dirichlet problems with discontinuous coefficients. SIAM Journal on Mathematical Analysis, 9(4):730–742, August 1978. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jerrard:1999:LBG

[Jer99] Robert L. Jerrard. Lower bounds for generalized Ginzburg–Landau functionals. SIAM Journal on Mathematical Analysis, 30(4): 721–746, 1999. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

http://epubs.siam.org/sam-
bin/dbq/article/30058.

Jiang:1988:SFL

[Jia88] Jyh-Ming Jiang. Starlike functions and linear functions of a Dirichlet distributed vector. SIAM Journal on Mathematical Analysis, 19(2):390–397, March 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jiang:1996:PMS

[Jia96] Ji-Fa Jiang. Periodic monotone systems with an invariant function. SIAM Journal on Mathematical Analysis, 27(6):1738– 1744, November 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/26063.

Jiang:1998:RMR

[Jia98] Qingtang Jiang. On the regularity of matrix refinable functions. SIAM Journal on Mathematical Analysis, 29(5):1157– 1176, September 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30817.

Jerri:1975:SED

[JK75] Abdul J. Jerri and Dale W. Kreisler. Sampling expansions with derivatives for finite Hankel and other transforms. SIAM Journal on Mathematical Analysis, 6(2):262–267, April 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jones:1986:IMS

[JK86] C. Jones and T. Küpper. On the infinitely many solutions of a semilinear elliptic equation. SIAM Journal on Mathematical Analysis, 17(4):803–835, July 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jones:1996:TIM

[JKK96] C. K. R. T. Jones, Tasso J. Kaper, and Nancy Kopell. Tracking invariant manifolds up to exponentially small errors. SIAM Journal on Mathematical Analysis, 27(2):558–577, March 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/25966.

Jordan:1998:VFF

[JKO98] Richard Jordan, David Kinderlehrer, and Felix Otto. The variational formulation of the Fokker– Planck equation. SIAM Journal on Mathematical Analysis, 29 (1):1–17, January 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30335.

Joly:1999:TSS

[JKV99] Patrick Joly, Alexander Komech, and Olivier Vacus. On transitions to stationary states in a Maxwell–Landau–Lifschitz– Gilbert system. SIAM Journal on Mathematical Analysis, 31 (2):346–374, 1999. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

http://epubs.siam.org/sam-
bin/dbq/article/32994.

Jackson:1972:IPB

[JM72] A. D. Jackson and L. C. Maximon. Integrals of products of Bessel functions. SIAM Journal on Mathematical Analysis, 3 (3):446–460, August 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jimbo:1996:GLE

[JM96] Shuichi Jimbo and Yoshihisa Morita. Ginzburg–Landau equations and stable solutions in a rotational domain. SIAM Journal on Mathematical Analysis, 27(5): 1360–1385, September 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Joly:1996:GSA

[JMR96] J. L. Joly, G. Metivier, and J. Rauch. Global solvability of the anharmonic oscillator model from nonlinear optics. SIAM Journal on Mathematical Analysis, 27(4):905–913, July 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/27367.

Jones:1994:ASP

[JNW94] William B. Jones, Olav Njåstad, and Haakon Waadeland. Application of Szegő polynomials to frequency analysis. SIAM Journal on Mathematical Analysis, 25 (2):491–512, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/22928.

Johnson:1975:IEF

[Joh75] H. L. Johnson. An integral equation formulation of a mixed boundary value problem on a sphere. SIAM Journal on Mathematical Analysis, 6(2):417–426, April 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Johnson:1977:MBV

[Joh77] H. L. Johnson. On a mixed boundary value problem for the reduced wave equation on a sphere. SIAM Journal on Mathematical Analysis, 8(1):1– 15, February 1977. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Johnson:1980:MBV

[Joh80] H. L. Johnson. A mixed boundary value problem viewed as a type V problem. SIAM Journal on Mathematical Analysis, 11(3):

565–576, May 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Johnston:1987:FER

[Joh87] Elgin H. Johnston. Faber expansions of rational and entire functions. SIAM Journal on Mathematical Analysis, 18(5):1235– 1247, September 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Joly:1992:STT

[Jol92] Patrick Joly. Some trace theorems in anisotropic Sobolev spaces. SIAM Journal on Mathematical Analysis, 23(3):799–819, May 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jones:1976:OCT

[Jon76] Gary D. Jones. Oscillation criteria for third order differential equations. SIAM Journal on Mathematical Analysis, 7(1):13– 15, February 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jones:1981:OOT

[Jon81] Gary D. Jones. An ordering of oscillation types for y^{(n)} + py = 0.. SIAM Journal on Mathematical Analysis, 12(1):72–77, January 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jones:1994:AR

[Jon94] D. S. Jones. Asymptotic remainders. SIAM Journal on Mathematical Analysis, 25(2): 474–490, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/22868.

Joshi:1999:RAC

[Jos99] M. S. Joshi. Recovering asymptotics of Coulomb-like potentials from fixed energy scattering data. SIAM Journal on Mathematical Analysis, 30 (3):516–526, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32763.

Joye:1997:EAS

[Joy97] Alain Joye. Exponential asymptotics in a singular limit for nlevel scattering systems. SIAM Journal on Mathematical Analysis, 28(3):669–703, May 1997. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/28884.

Joye:1995:SAB

[JP95] Alain Joye and Charles-Edouard Pfister. Semiclassical asymptotics beyond all orders for simple scattering systems. SIAM Journal on Mathematical Analysis, 26(4):944–977, July 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/25085.

Johnson:1987:EPL

[JPS87] Russell A. Johnson, Kenneth J.

Palmer, and George R. Sell. Ergodic properties of linear dynamical systems. SIAM Journal on Mathematical Analysis, 18(1):1– 33, January 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jorba:1997:ERQ

[JRRV97] Angel Jorba, Rafael Ramirez-Ros, and Jordi Villanueva. Effective reducibility of quasi-periodic linear equations close to constant coefficients. SIAM Journal on Mathematical Analysis, 28 (1):178–188, January 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/28096.

Jetter:1994:HIL

[JRS94] Kurt Jetter, Sherman D. Riemenschneider, and Zuowei Shen. Hermite interpolation on the lattice Z^{d}. SIAM Journal on Mathematical Analysis, 25(3):962–975, May 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/22691.

Johnson:1987:SIF

[JS87] G. W. Johnson and D. L. Skoug.

A stochastic integration formula for two-parameter Wiener × twoparameter Wiener space. SIAM Journal on Mathematical Analysis, 18(4):919–932, July 1987.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jepson:1989:RPN

[JS89] A. D. Jepson and A. Spence. On a reduction process for nonlinear equations. SIAM Journal on Mathematical Analysis, 20 (1):39–56, January 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jorba:1996:QPP

[JS96] Àngel Jorba and Carles Simó. On quasi-periodic perturbations of elliptic equilibrium points. SIAM Journal on Mathematical Analysis, 27(6):1704–1737, November 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/27691.

Jordan:1987:COF

[JSW87a] G. S. Jordan, Olof J. Staffans, and Robert L. Wheeler. Convolution operators in a fading memory space: the critical case. SIAM Journal on Mathematical Analysis, 18(2):366–386, March 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jordan:1987:SSU

[JSW87b] G. S. Jordan, Olof J. Staffans, and Robert L. Wheeler. Subspaces of stable and unstable solutions of a functional-differential equation in a fading memory space: the critical case. SIAM Journal on Mathematical Analysis, 18(5):1323–1340, September 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jones:1975:CPA

[JT75] William B. Jones and W. J.

Thron. On convergence of padé approximants. SIAM Journal on Mathematical Analysis, 6(1):9– 16, February 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jones:1979:SMF

[JT79] William B. Jones and W. J.

Thron. Sequences of meromorphic functions corresponding to a formal Laurent series. SIAM Journal on Mathematical Analysis, 10(1):1–17, January 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jones:1994:RQS

[JT94] D. A. Jones and E. S. Titi.

A remark on quasi-stationary approximate inertial manifolds for the Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 25(3):894–914, May 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/23042.

Juang:1995:CIL

[Jua95] Jonq Juang. On coupled integral H-like equations of Chandrasekhar. SIAM Journal on Mathematical Analysis, 26(4): 869–879, July 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/25047.

Justice:1974:SRR

[Jus74] James H. Justice. The Szegő recursion relation and inverses of positive definite Toeplitz matrices. SIAM Journal on Mathematical Analysis, 5(3):503–508, May 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jordan:1974:ABP

[JW74] G. S. Jordan and Robert L. Wheeler. On the asymptotic behavior of perturbed Volterra integral equations. SIAM Journal on Mathematical Analysis, 5 (2):273–277, April 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jordan:1980:SRV

[JW80a] G. S. Jordan and Robert L. Wheeler. Structure of resolvents of Volterra integral and integrodifferential systems. SIAM Journal on Mathematical Analysis, 11 (1):119–132, January 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Jordan:1980:WRT

[JW80b] G. S. Jordan and Robert L. Wheeler. Weighted L^{1}-remainder theorems for resolvents of Volterra equations. SIAM Journal on Mathematical Analysis, 11(5):885–900, September 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Johansen:1988:SRP

[JW88] Thormod Johansen and Ragnar Winther. The solution of the Riemann problem for a hyperbolic system of conservation laws modeling polymer flooding. SIAM Journal on Mathematical Analysis, 19(3):541–566, May 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Johansen:1989:RPM

[JW89] Thormod Johansen and Ragnar Winther. The Riemann problem for multicomponent polymer flooding. SIAM Journal on Mathematical Analysis, 20(4): 908–929, July 1989. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

ka:1976:SIPa

[ka76a] Ivo Babu ka. Solution of interface problems by homogenization. I. SIAM Journal on Mathematical Analysis, 7(5):603–634, October 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

ka:1976:SIPb

[ka76b] Ivo Babu ka. Solution of interface problems by homogenization. II. SIAM Journal on Mathematical Analysis, 7(5):635–645, October 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

ka:1977:SIP

[ka77] Ivo Babu ka. Solution of interface problems by homogenization. III. SIAM Journal on Mathematical Analysis, 8(6):923–937, November 1977. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kuttler:1988:QEE

[KA88] Kenneth Kuttler and Elias Aifantis. Quasilinear evolution equations in nonclassical diffusion. SIAM Journal on Mathematical Analysis, 19(1):110–120, January 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kadane:1972:CPH

[Kad72] Joseph B. Kadane. A concave property of the hypergeometric function with respect to a parameter. SIAM Journal on Mathematical Analysis, 3(4):642–643, November 1972. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kadell:1987:PPR

[Kad87] Kevin W. J. Kadell. A probabilistic proof of Ramanujan’s _{1}ψ_{1} sum. SIAM Journal on Mathematical Analysis, 18(6):1539– 1548, November 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kadell:1988:PAC

[Kad88a] Kevin W. J. Kadell. A proof of Askey’s conjectured q-analogue of Selberg’s integral and a conjecture of Morris. SIAM Journal on Mathematical Analysis, 19(4): 969–986, July 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kadell:1988:PSA

[Kad88b] Kevin W. J. Kadell. A proof of some q-analogues of Selberg’s integral for k = 1. SIAM Journal on Mathematical Analysis, 19(4): 944–968, July 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kaizu:1991:PEN

[Kai91] Satoshi Kaizu. The Poisson equation with nonautonomous semilinear boundary conditions in domains with many time holes. SIAM Journal on Mathematical Analysis, 22(5):1222–1245, September 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kaiser:1992:ATW

[Kai92] Gerald Kaiser. An algebraic theory of wavelets. I. operational calculus and complex structure. SIAM Journal on Mathematical Analysis, 23(1):222–243, January 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kalnins:1975:SVL

[Kal75] E. Kalnins. On the separation of variables for the Laplace equation Δψ + K^{2}ψ = 0 in two- and three-dimensional Minkowski space. SIAM Journal on Mathematical Analysis, 6 (2):340–374, April 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kamin:1979:ESP

[Kam79] S. Kamin. On elliptic singular perturbation problems with turning points. SIAM Journal on Mathematical Analysis, 10(3): 447–455, May 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kaminski:1989:AEP

[Kam89] David Kaminski. Asymptotic expansion of the Pearcey integral near the caustic. SIAM Journal on Mathematical Analysis, 20 (4):987–1005, July 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kaminski:1992:ASI

[Kam92] D. Kaminski. Asymptotics of the swallowtail integral near the cusp of the caustic. SIAM Journal on Mathematical Analysis, 23 (1):262–285, January 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kaneko:1993:SIH

[Kan93] Jyoichi Kaneko. Selberg integrals and hypergeometric functions associated with Jack polynomials. SIAM Journal on Mathematical Analysis, 24(4):1086– 1110, July 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kaplan:1972:ABV

[Kap72] James L. Kaplan. On the asymptotic behavior of Volterra integral equations. SIAM Journal on Mathematical Analysis, 3(1): 148–156, February 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kaper:1979:BVP

[Kap79a] Hans G. Kaper. Boundary value problems of mixed type arising in the kinetic theory of gases. SIAM Journal on Mathematical Analysis, 10(1):161–178, January 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kaper:1979:SRU

[Kap79b] Hans G. Kaper. Spectral representation of an unbounded linear transformation arising in the kinetic theory of gases. SIAM Journal on Mathematical Analysis, 10(1):179–191, January 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kapitula:1998:EFG

[Kap98] Todd Kapitula. The Evans function and generalized Mel′nikov integrals. SIAM Journal on Mathematical Analysis, 30(2): 273–297, 1998. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

http://epubs.siam.org/sam-
bin/dbq/article/32796.

Karaa:1998:SEE

[Kar98] Samir Karaa. Sharp estimates for the eigenvalues of some differential equations. SIAM Journal on Mathematical Analysis, 29(5): 1279–1300, September 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/30784.

Katsevich:1997:RRT

[Kat97] Alexander I. Katsevich. Range of the Radon transform on functions which do not decay fast at infinity. SIAM Journal on Mathematical Analysis, 28 (4):852–866, July 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/28951.

Kawamura:1979:ILC

[Kaw79] Yoshiaki Kawamura. Invertibility of Lipschitz continuous mappings and its application to electrical network equations. SIAM Journal on Mathematical Analysis, 10(2):253–265, March 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kazmin:1984:PBS

[Kaz84] Yu. A. Kaz′min. On a polynomial basis in some spaces of analytic functions. SIAM Journal on Mathematical Analysis, 15(6): 1169–1176, November 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

King:1973:GAT

[KB73] B. J. King and A. L. Buren.

A general addition theorem for spheroidal wave functions. SIAM Journal on Mathematical Analysis, 4(1):149–160, February 1973. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Keener:1972:OSC

[Kee72] M. S. Keener. On oscillatory solutions of certain fourth order linear differential equations. SIAM Journal on Mathematical Analysis, 3(4):599–605, November 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Keener:1977:VTT

[Kee77] James P. Keener. On the validity of the two-timing method for large times. SIAM Journal on Mathematical Analysis, 8(6): 1067–1091, November 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kegley:1983:GEU

[Keg83] J. Colby Kegley. A global existence and uniqueness theorem for a Riccati equation. SIAM Journal on Mathematical Analysis, 14 (1):47–59, January 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Keller:1972:ETM

[Kel72] Herbert B. Keller. Existence theory for multiple solutions of a singular perturbation problem.

SIAM Journal on Mathematical Analysis, 3(1):86–92, February 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Keller:1973:STM

[Kel73] Herbert B. Keller. Stability theory for multiple equilibrium states of a nonlinear diffusion process: A singularly perturbed eigenvalue problem. SIAM Journal on Mathematical Analysis, 4 (1):134–140, February 1973. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kelley:1979:SEI

[Kel79a] C. T. Kelley. Solution of Hequations by iteration. SIAM Journal on Mathematical Analysis, 10(4):844–849, July 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kelley:1979:NSP

[Kel79b] Walter G. Kelley. A nonlinear singular perturbation problem for second-order systems. SIAM Journal on Mathematical Analysis, 10(1):32–37, January 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kelley:1984:BIL

[Kel84]  Walter G. Kelley.  Bound-

ary and interior layer phenomena for singularly perturbed systems. SIAM Journal on Mathematical Analysis, 15(4):635–641, July 1984. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Kellogg:1988:DSL

[Kel88] R. Bruce Kellogg. Discontinuous solutions of the linearized, steady state, compressible, viscous, Navier–Stokes equations. SIAM Journal on Mathematical Analysis, 19(3):567–579, May 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kelley:1989:UAS

[Kel89] Walter G. Kelley. Uniform approximation of singular perturbation problems having singular regular expansions. SIAM Journal on Mathematical Analysis, 20 (2):479–493, March 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kemnitz:1982:PES

[Kem82a] Hans Kemnitz. Polynomial expansions for solutions of D_{x}^{r}u(x, t) = D_{t}u(x, t), r = 2, 3, 4, · · ·. SIAM Journal on Mathematical Analysis, 13(4): 640–650, July 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kemperman:1982:HMT

[Kem82b] J. H. B. Kemperman. A Hurwitz matrix is totally positive. SIAM Journal on Mathematical Analysis, 13(2):331–341, March 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kemnitz:1985:PES

[Kem85] Hans Kemnitz. Polynomial expansions for solutions of the system D_{X_{1}}^{k}U(X_{1}, ⋯, X_{r}) = D_{X_{k}}U(X_{1}, ⋯, X_{r})^{∗}. SIAM Journal on Mathematical Analysis, 16(2):379–391, March 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Keyfitz:1991:ACS

[Key91] Barbara Lee Keyfitz. Admissibility conditions for shocks in conservation laws that change type. SIAM Journal on Mathematical Analysis, 22(5):1284– 1292, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kelman:1974:DOS

[KF74] Robert B. Kelman and Robert P. Feinerman. Dual orthogonal series. SIAM Journal on Mathematical Analysis, 5(3):489–502, May 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kuang:1990:MOS

[KF90] Yang Kuang and Alan Feldstein. Monotonic and oscillatory solutions of a linear neutral delay equ ation with infinite lag. SIAM Journal on Mathematical Analysis, 21(6):1633–1641, November 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Korevaar:1971:FDE

[KG71] J. Korevaar and T. Geveci.

Fields due to electrons on an analytic curve. SIAM Journal on Mathematical Analysis, 2(3): 445–453, August 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kawashima:1995:ESP

[KH95] Shuichi Kawashima and Harumi Hattori. Existence of shock profiles for viscoelastic materials with memory. SIAM Journal on Mathematical Analysis, 26(5): 1130–1142, September 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24929.

Khalsa:1987:ATT

[Kha87] Sat Nam S. Khalsa. Application of topological techniques to the analysis of asymptotic behavior of numerical solutions of a reaction-diffusion equation. SIAM Journal on Mathematical Analysis, 18(1):184–190, January 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Khosrovshahi:1977:NNS

[Kho77] G. B. Khosrovshahi. Nonexistence of nontrivial solutions of Schrödinger type systems. SIAM Journal on Mathematical Analysis, 8(6):998–1013, November 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

ki:1972:RSI

[ki72] Eduard Prugove ki. Riemann– Stieltjes integration of operatorvalued functions with respect to vector-valued functions in Banach spaces. SIAM Journal on Mathematical Analysis, 3(2): 183–205, May 1972. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Klibanov:1995:USC

[KI95] Michael V. Klibanov and Olga V. Ioussoupova. Uniform strict convexity of a cost functional for three-dimensional inverse scattering problem. SIAM Journal on Mathematical Analysis, 26(1): 147–179, January 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/24403.

Kiffe:1977:VEN

[Kif77] T. R. Kiffe. A Volterra equation with a nonconvolution kernel. SIAM Journal on Mathematical Analysis, 8(6):938–949, November 1977. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kiffe:1980:PAV

[Kif80] T. Kiffe. A perturbation of an abstract Volterra equation. SIAM Journal on Mathematical Analysis, 11(6):1036–1046, November 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kim:1983:SEO

[Kim83] Jong Uhn Kim. Solutions to the equations of one-dimensional viscoelasticity in BV. SIAM Journal on Mathematical Analysis, 14 (4):684–695, July 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kim:1986:DCT

[Kim86] W. J. Kim. Disconjugacy and comparison theorems for secondorder linear systems. SIAM Journal on Mathematical Analysis, 17 (5):1104–1112, September 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kim:1987:WSI

[Kim87] Jong Uhn Kim. Weak solutions of an initial-boundary value problem for an incompressible viscous fluid with nonnegative density. SIAM Journal on Mathematical Analysis, 18(1):89–96, January 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kim:1990:SMT

[Kim90]  Jong Uhn Kim.  Semidis-

cretization method for threedimensional motion of a Bingham fluid. SIAM Journal on Mathematical Analysis, 21(1): 53–75, January 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kim:1992:EDL

[Kim92] Jong Uhn Kim. On the energy decay of a linear thermoelastic bar and plate. SIAM Journal on Mathematical Analysis, 23(4): 889–899, July 1992. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kim:1995:LRO

[Kim95] Jong Uhn Kim. Local regularity of the one-dimensional motion of a viscoelastic medium. SIAM Journal on Mathematical Analysis, 26(3):738–749, May 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/22759.

Kim:1998:WEB

[Kim98] Jong Uhn Kim. On a wave equation with a boundary condition associated with capillary waves. SIAM Journal on Mathematical Analysis, 30(1):53–71, 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32821.

Kinderlehrer:1986:RBS

[Kin86] David Kinderlehrer. A relation between semi-inverse and Saint– Venant solutions for prisms. SIAM Journal on Mathematical Analysis, 17(3):626–640, May 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kohler-Jobin:1982:SED

[KJ82]  Marie-Thérèse  Kohler-Jobin.

Symmetrization with equal Dirichlet integrals. SIAM Journal on Mathematical Analysis, 13 (1):153–161, January 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kaper:1992:TCC

[KK92] Hans G. Kaper and Man Kam Kwong. On two conjectures concerning the multiplicity of solutions of a Dirichlet problem. SIAM Journal on Mathematical Analysis, 23(3):571–578, May 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kweon:1997:CNS

[KK97] Jae Ryong Kweon and R. Bruce Kellogg. Compressible Navier– Stokes equations in a bounded domain with inflow boundary condition. SIAM Journal on Mathematical Analysis, 28(1): 94–108, January 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/28425.

Kreiss:1999:SCL

[KKL99] Gunilla Kreiss, Heinz-Otto Kreiss, and Jens Lorenz. On stability of conservation laws. SIAM Journal on Mathematical Analysis, 30(2):401–430, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32247.

Kreith:1984:OPN

[KKY84] Kurt Kreith, Takaŝi Kusano, and Norio Yoshida. Oscillation properties of nonlinear hyperbolic equations. SIAM Journal on Mathematical Analysis, 15(3): 570–578, May 1984. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kaper:1984:RTC

[KKZ84] Hans G. Kaper, Man Kam Kwong, and Anton Zettl. Regularizing transformations for certain singular Sturm–Liouville boundary value problems. SIAM Journal on Mathematical Analysis, 15(5):957–963, September 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kaper:1986:CFE

[KKZ86] Hans G. Kaper, Man Kam Kwong, and Anton Zettl. Characterizations of the Friedrichs extensions of singular Sturm– Liouville expressions. SIAM Journal on Mathematical Analysis, 17(4):772–777, July 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kilpelainen:1996:DBV

[KL96] T. Kilpeläinen and P. Lindqvist. On the Dirichlet boundary value problem for a degenerate parabolic equation. SIAM Journal on Mathematical Analysis, 27(3):661–683, May 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Krylov:1998:SST

[KL98] N. V. Krylov and S. V. Lototsky. A Sobolev space theory of SPDE with constant coefficients on a half line. SIAM Journal on Mathematical Analysis, 30(2):298–325, 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32690.

Krylov:1999:SST

[KL99] N. V. Krylov and S. V. Lototsky. A Sobolev space theory of SPDEs with constant coefficients in a half space. SIAM Journal on Mathematical Analysis, 31(1):19–33, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/33884.

Klamkin:1971:EDM

[Kla71] Murray S. Klamkin. Extensions of Dirichlet’s multiple integral. SIAM Journal on Mathematical Analysis, 2(3):467–469, August 1971. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Klamkin:1977:CP

[Kla77] M. S. Klamkin. On Chaplygin’s problem. SIAM Journal on Mathematical Analysis, 8 (2):288–289, April 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Klibanov:1997:GCT

[Kli97] Michael V. Klibanov. Global convexity in a three-dimensional inverse acoustic problem. SIAM Journal on Mathematical Analysis, 28(6):1371–1388, November 1997. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/29736.

Kloeden:1985:GEC

[Klo85] P. E. Kloeden. Global existence of classical solutions in the dissipative shallow water equations. SIAM Journal on Mathematical Analysis, 16(2):301–315, March 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kawohl:1993:BEC

[KLV93] Bernhard Kawohl, Howard A. Levine, and Waldemar Velte. Buckling eigenvalues for a clamped plate embedded in an elastic medium and related questions. SIAM Journal on Mathematical Analysis, 24(2):327–340, March 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kwon:1994:COP

[KLY94] K. H. Kwon, L. L. Littlejohn, and B. H. Yoo. Characterizations of orthogonal polynomials satisfying differential equations. SIAM Journal on Mathematical Analysis, 25(3):976–990, May 1994. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/23643.

Kalnins:1978:LTW

[KM78] E. G. Kalnins and Willard Miller, Jr. Lie theory and the wave equation in space-time. 2. the group SO(4, C). SIAM Journal on Mathematical Analysis, 9 (1):12–33, February 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kalnins:1980:KTV

[KM80] E. G. Kalnins and Willard Miller, Jr. Killing tensors and variable separation for Hamilton–Jacobi and Helmholtz equations. SIAM Journal on Mathematical Analysis, 11(6):1011–1026, November 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kalnins:1981:KTN

[KM81] E. G. Kalnins and Willard Miller, Jr. Killing tensors and nonorthogonal variable separation for Hamilton–Jacobi equations. SIAM Journal on Mathematical Analysis, 12(4):617–629, July 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kalnins:1983:CKT

[KM83] E. G. Kalnins and Willard Miller, Jr. Conformal Killing tensors and variable separation for Hamilton–Jacobi equations.

SIAM Journal on Mathematical Analysis, 14(1):126–137, January 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kalnins:1985:REE

[KM85] E. G. Kalnins and Willard Miller, Jr. Related evolution equations and Lie symmetries. SIAM Journal on Mathematical Analysis, 16 (2):221–232, March 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Klaasen:1986:SWS

[KM86] Gene A. Klaasen and Enzo Mitidieri. Standing wave solutions for a system derived from the FitzHugh–Nagumo equations for nerve conduction. SIAM Journal on Mathematical Analysis, 17 (1):74–83, January 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kalnins:1988:SOP

[KM88] E. G. Kalnins and Willard Miller, Jr. q-series and orthogonal polynomials associated with Barnes’ first lemma. SIAM Journal on Mathematical Analysis, 19(5): 1216–1231, September 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kalnins:1991:AHE

[KM91a] E. G. Kalnins and W. Miller, Jr. Addendum: Hypergeometric expansions of Heun polynomials. SIAM Journal on Mathematical Analysis, 22(6):1803, September 1991. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Kalnins:1991:HEH

[KM91b] E. G. Kalnins and W. Miller, Jr. Hypergeometric expansions of Heun polynomials. SIAM Journal on Mathematical Analysis, 22(5):1450–1459, September 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kostova:1991:NAD

[KM91c] Tanya Kostova and Fabio Milner.

Nonlinear age-dependent population dynamics with constant size. SIAM Journal on Mathematical Analysis, 22(1):129–137, January 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Koekoek:1993:GLP

[KM93] R. Koekoek and H. G. Meijer. A generalization of Laguerre polynomials. SIAM Journal on Mathematical Analysis, 24 (3):768–782, May 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kupper:1995:BSN

[KM95] Tassilo Küpper and Thomas Mrziglod. On the bifurcation structure of nonlinear perturbations of Hill’s equations at boundary points of the continuous spectrum. SIAM Journal on Mathematical Analysis, 26(5): 1284–1305, September 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/25087.

Kalnins:1994:MAR

[KMM94] E. G. Kalnins, W. Miller, Jr., and S. Mukherjee. Models of Q-algebra representations: The group of plane motions. SIAM Journal on Mathematical Analysis, 25(2):513–527, March 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/22461.

Kalnins:1991:FOB

[KMT91] E. G. Kalnins, Willard Miller, Jr., and M. V. Tratnik. Families of orthogonal and biorthogonal polynomials on the n-sphere. SIAM Journal on Mathematical Analysis, 22(1):272–294, January 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kolbe:1972:STC

[KN72] W. Kolbe and R. J. Nessel. Saturation theory in connection with Mellin transform methods. SIAM Journal on Mathematical Analysis, 3(2):246–262, May 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kangro:1997:ABS

[KN97] Urve Kangro and Roy Nicolaides. Asymptotic behavior of solutions of two-dimensional periodic scattering problems in electromagnetics. SIAM Journal on Mathematical Analysis, 28(6):1406– 1419, November 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29808.

Kawashima:1998:SWM

[KN98] Shuichi Kawashima and Shinya Nishibata. Shock waves for a model system of the radiating gas. SIAM Journal on Mathematical Analysis, 30 (1):95–117, 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32216.

Knabner:1986:FBP

[Kna86] Peter Knabner. A free boundary problem arising from the leaching of saline soils. SIAM Journal on Mathematical Analysis, 17(3): 610–625, May 1986. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Knightly:1972:CPN

[Kni72] George H. Knightly. A Cauchy problem for the Navier–Stokes equations in R^{n}. SIAM Journal on Mathematical Analysis, 3 (3):506–511, August 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kamran:1989:EDO

[KO89] Niky Kamran and Peter J. Olver.

Equivalence of differential operators. SIAM Journal on Mathematical Analysis, 20(5):1172– 1185, September 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kichenassamy:1992:ENS

[KO92] Satyanad Kichenassamy and Peter J. Olver. Existence and nonexistence of solitary wave solutions to higher-order model evolution equations. SIAM Journal on Mathematical Analysis, 23 (5):1141–1166, September 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kan-on:1995:PDP

[Ko95a] Yukio Kan-on. Parameter dependence of propagation speed of travelling waves for competitiondiffusion equations. SIAM Journal on Mathematical Analysis, 26 (2):340–363, March 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24455.

Korman:1995:MRT

[KO95b] Philip Korman and Tiancheng Ouyang. Multiplicity results for two classes of boundaryvalue problems. SIAM Journal on Mathematical Analysis, 26(1): 180–189, January 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23187.

Kalachev:1996:RLD

[KO96] Leonid V. Kalachev and Robert E. O’Malley, Jr. The regularization of linear differential-algebraic equations. SIAM Journal on Mathematical Analysis, 27(1): 258–273, January 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/21867.

Kan-on:1998:BSS

[Ko98]  Yukio Kan-on.  Bifurcation

structure of stationary solutions of a Lotka–Volterra competition model with diffusion. SIAM Journal on Mathematical Analysis, 29(2):424–436, March 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/30578.

Koelink:1994:AFC

[Koe94]  H. T. Koelink.  The addi-

tion formula for continuous q-Legendre polynomials and associated spherical elements on the SU(2) quantum group related to Askey–Wilson polynomials. SIAM Journal on Mathematical Analysis, 25(1): 197–217, January 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/18611.

Koh:1970:HTN

[Koh70a] E. L. Koh. The Hankel transformation of negative order for distributions of rapid growth. SIAM Journal on Mathematical Analysis, 1(3):322–327, August 1970. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Koh:1970:RHT

[Koh70b] E. L. Koh. A representation of Hankel transformable generalized functions. SIAM Journal on Mathematical Analysis, 1(1):33– 36, February 1970. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Koliha:1974:SLE

[Kol74a] J. J. Koliha. The solution of linear equations in normed spaces by averaging iteration. SIAM Journal on Mathematical Analysis, 5(2):178–186, April 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kollett:1974:TTM

[Kol74b] F. W. Kollett. Two-timing methods valid on expanding intervals. SIAM Journal on Mathematical Analysis, 5(4):613–625, August 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kolata:1979:WPF

[Kol79] William Kolata. Weakly perturbed forms and the spectral properties of associated eigenvalue problems. SIAM Journal on Mathematical Analysis, 10 (2):297–305, March 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kolbig:1986:NGP

[Köl86] K. S. Kölbig. Nielsen’s generalized polylogarithms. SIAM Journal on Mathematical Analysis, 17(5):1232–1258, September 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kan-on:1998:SPA

[KoM98] Yukio Kan-on and Masayasu Mimura. Singular perturbation approach to a 3-component reaction-diffusion system arising in population dynamics. SIAM Journal on Mathematical Analysis, 29(6):1519–1536, November 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/31832.

Koornwinder:1974:JPI

[Koo74] Tom Koornwinder. Jacobi polynomials, II. an analytic proof of the product formula. SIAM Journal on Mathematical Analysis, 5(1):125–137, February 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Koornwinder:1975:JPI

[Koo75] Tom Koornwinder. Jacobi polynomials, III. an analytic proof of the addition formula. SIAM Journal on Mathematical Analysis, 6(3):533–543, May 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Koornwinder:1977:AFL

[Koo77] Tom Koornwinder. The addition formula for Laguerre polynomials. SIAM Journal on Mathematical Analysis, 8(3):535–540, May 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Koornwinder:1982:KPU

[Koo82] Tom H. Koornwinder. Krawtchouk polynomials, a unification of two different group theoretic interpretations. SIAM Journal on Mathematical Analysis, 13(6): 1011–1023, November 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Koornwinder:1985:MEI

[Koo85] Tom H. Koornwinder. Matrix elements of irreducible representations of SU(2) × SU(2) and vector-valued orthogonal polynomials. SIAM Journal on Mathematical Analysis, 16(3):602–613, May 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Koornwinder:1991:AFL

[Koo91] Tom H. Koornwinder. The addition formula for little q-legendre polynomials and the su(2) quantum group. SIAM Journal on Mathematical Analysis, 22(1): 295–301, January 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Koornwinder:1993:AWP

[Koo93] Tom H. Koornwinder. Askey– Wilson polynomials as zonal spherical functions on the su(2) quantum group. SIAM Journal on Mathematical Analysis, 24(3): 795–813, May 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Korman:1988:ESS

[Kor88] Philip Korman. On existence of solutions for several classes of free boundary problems. SIAM Journal on Mathematical Analysis, 19(4):814–823, July 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kotorynski:1994:DPS

[Kot94]  W. P. Kotorynski.  Disper-

sion in pipes with slowly varying cross-sections. SIAM Journal on Mathematical Analysis, 25 (3):915–940, May 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/18373.

Kovacic:1995:SPT

[Kov95] Gregor Kovačič. Singular perturbation theory for homoclinic orbits in a class of near- integrable dissipative systems. SIAM Journal on Mathematical Analysis, 26(6):1611–1643, November 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/24542.

Kowalski:1982:ORF

[Kow82a] M. A. Kowalski. Orthogonality and recursion formulas for polynomials in n variables. SIAM Journal on Mathematical Analysis, 13(2):316–323, March 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kowalski:1982:RFO

[Kow82b] M. A. Kowalski. The recursion formulas for orthogonal polynomials in n variables. SIAM Journal on Mathematical Analysis, 13 (2):309–315, March 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kreiss:1974:RSP

[KP74] Heinz Otto Kreiss and Seymour V. Parter. Remarks on singular perturbations with turning points. SIAM Journal on Mathematical Analysis, 5(2):230–251, April 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kinderlehrer:1992:WCI

[KP92] David Kinderlehrer and Pablo Pedregal. Weak convergence of integrands and the Young measure representation. SIAM Journal on Mathematical Analysis, 23 (1):1–19, January 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Klamkin:1976:NDE

[KR76] Murray S. Klamkin and James L. Reid. Nonlinear differential equations equivalent to solvable nonlinear equations. SIAM Journal on Mathematical Analysis, 7(3): 305–310, May 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kieffer:1981:MCA

[KR81] John C. Kieffer and Maurice Rahe. Markov channels are asymptotically mean stationary. SIAM Journal on Mathematical Analysis, 12(3):293–305, May 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kranzler:1972:SCO

[Kra72] Stan K. Kranzler. Sums of convolution operators. SIAM Journal on Mathematical Analysis, 3(2): 380–388, May 1972. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kraft:1974:UEI

[Kra74] Richard Kraft. Uniqueness and existence for the integral equation of interreflections. SIAM Journal on Mathematical Analysis, 5(2):293–302, April 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Krabbe:1975:PPR

[Kra75] Gregers L. Krabbe. The periodic part of the response to a periodic excitation. SIAM Journal on Mathematical Analysis, 6 (2):375–403, April 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Krall:1978:OPT

[Kra78] Allan M. Krall. Orthogonal polynomials through moment generating functionals. SIAM Journal on Mathematical Analysis, 9 (4):600–603, August 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Krall:1989:TSHa

[Kra89a] Allan M. Krall. M(λ) theory for singular Hamiltonian systems with one singular point. SIAM Journal on Mathematical Analysis, 20(3):664–700, May 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Krall:1989:TSHb

[Kra89b] Allan M. Krall. M(λ) theory for singular Hamiltonian systems with two singular points. SIAM Journal on Mathematical Analysis, 20(3):701–715, May 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kratz:1991:RSF

[Kra91] Werner Kratz. On the representation of Stokes flows. SIAM Journal on Mathematical Analysis, 22(2):414–423, March 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kreyszig:1994:NSI

[Kre94] Erwin Kreyszig. A new standard isometry of developable surfaces in CAD/CAM. SIAM Journal on Mathematical Analysis, 25 (1):174–178, January 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24609.

Kristensson:1980:UTH

[Kri80] Gerhard Kristensson. A uniqueness theorem for Helmholtz’ equation: penetrable media with an infinite interface. SIAM Journal on Mathematical Analysis, 11 (6):1104–1117, November 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Krupa:1990:BRE

[Kru90] Martin Krupa. Bifurcations of relative equilibria. SIAM Journal on Mathematical Analysis, 21(6): 1453–1486, November 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Krylov:1996:TSP

[Kry96] N. V. Krylov. On L_{p}-theory of stochastic partial differential equations in the whole space. SIAM Journal on Mathematical Analysis, 27(2):313–340, March 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/26331.

Knightly:1975:NAB

[KS75] George H. Knightly and D. Sather.

Nonlinear axisymmetric buckled states of shallow spherical caps. SIAM Journal on Mathematical Analysis, 6(6):913–924, November 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kuttler:1978:BEE

[KS78] J. R. Kuttler and V. G. Sigillito. Bounding eigenvalues of elliptic operators. SIAM Journal on Mathematical Analysis, 9 (4):768–773, August 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kiffe:1979:SVI

[KS79a] T. Kiffe and M. Stecher. L^{2} solutions of Volterra integral equations. SIAM Journal on Mathematical Analysis, 10(2):274–280, March 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Knightly:1979:NBS

[KS79b] George H. Knightly and D. Sather. Nonlinear buckling and stability of cylindrical panels. SIAM Journal on Mathematical Analysis, 10 (2):389–403, March 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kojima:1979:SHS

[KS79c] M. Kojima and R. Saigal. A study of PC^{1} homeomorphisms on subdivided polyhedrons. SIAM Journal on Mathematical Analysis, 10(6):1299– 1312, November 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kiffe:1980:PAR

[KS80a] T. Kiffe and M. Stecher. Properties and applications of the resolvent operator to a Volterra integral equation in Hilbert space. SIAM Journal on Mathematical Analysis, 11(1):82–91, January 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kuiper:1980:SSI

[KS80b] Hendrik J. Kuiper and Steven M.

Shew. Strong solutions for infinite-dimensional Riccati equations arising in transport theory. SIAM Journal on Mathematical Analysis, 11(2):211–222, March 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kusano:1985:ABS

[KS85a] Takaŝi Kusano and Bhagat Singh. Asymptotic behavior of solutions of a class of higher order ordinary differential equations. SIAM Journal on Mathematical Analysis, 16(4):770–784, July 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kusano:1985:GPS

[KS85b] Takaŝi Kusano and Charles A.

Swanson. Growth properties of stationary Klein–Gordon equations. SIAM Journal on Mathematical Analysis, 16(3):440–446, May 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kurth:1986:GRS

[KS86] Theo Kurth and Dieter Schmidt. On the global representation of the solutions of second-order linear differential equations having an irregular singularity of rank one in ∞ by series in terms of confluent hypergeometric functions. SIAM Journal on Mathematical Analysis, 17(5):1086– 1103, September 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kusano:1991:ESR

[KS91] Takaşi Kusano and Charles A. Swanson. Entire solutions of real and complex Monge–Ampère equations. SIAM Journal on Mathematical Analysis, 22(1): 157–168, January 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kannappan:1992:RIS

[KS92] Pl. Kannappan and P. K. Sahoo. Rotation invariant separable functions are Gaussian. SIAM Journal on Mathematical Analysis, 23(5):1342–1351, September 1992. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kim:1996:UDC

[KS96] Hyunseok Kim and Jin Keun Seo. Unique determination of a collection of a finite number of cracks from two boundary measurements. SIAM Journal on Mathematical Analysis, 27(5): 1336–1340, September 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/27548.

Koornwinder:1978:GPS

[KSK78] Tom Koornwinder and Ida Sprinkhuizen-Kuyper. Generalized power series expansions for a class of orthogonal polynomials in two variables. SIAM Journal on Mathematical Analysis, 9(3): 457–483, June 1978. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kang:1997:ICP

[KSS97] Hyeonbae Kang, Jin Keun Seo, and Dongwoo Sheen. The inverse conductivity problem with one measurement: Stability and estimation of size. SIAM Journal on Mathematical Analysis, 28(6): 1389–1405, November 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29937.

Kartsatos:1979:OAB

[KT79] A. G. Kartsatos and J. Toro. Oscillation and asymptotic behavior of forced nonlinear equations. SIAM Journal on Mathematical Analysis, 10(1):86–95, January 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kurganov:1997:SSH

[KT97] A. Kurganov and E. Tadmor.

Stiff systems of hyperbolic conservation laws: Convergence and error estimates. SIAM Journal on Mathematical Analysis, 28(6): 1446–1456, November 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/30148.

Kusaka:1999:CSS

[KT99] Yoshiaki Kusaka and Atusi Tani. On the classical solvability of the Stefan problem in a viscous incompressible fluid flow. SIAM Journal on Mathematical Analysis, 30(3):584–602, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33493.

Kolmanovskii:1994:SST

[KTV94] V. B. Kolmanovskiĭ, L. Torelli, and R. Vermiglio. Stability of some test equations with delay. SIAM Journal on Mathematical Analysis, 25(3):948–961, May 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/23802.

Kuga:1974:BFP

[Kug74] Kiyoshi Kuga. Brouwer’s fixedpoint theorem: An alternative proof. SIAM Journal on Mathematical Analysis, 5(6):893–897, November 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kuiper:1976:SNB

[Kui76] Hendrik J. Kuiper. Some nonlinear boundary value problems.

SIAM Journal on Mathematical Analysis, 7(4):551–564, August 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kuiken:1979:HEH

[Kui79] Kathryn Kuiken. Heun’s equation and the hypergeometric equation. SIAM Journal on Mathematical Analysis, 10(3): 655–657, May 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kuiper:1980:ISN

[Kui80] Hendrik J. Kuiper. Invariant sets for nonlinear elliptic and parabolic systems. SIAM Journal on Mathematical Analysis, 11 (6):1075–1103, November 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kuiper:1985:GOR

[Kui85]  Hendrik J. Kuiper.  Gener-

alized operator Riccati equations. SIAM Journal on Mathematical Analysis, 16(4):675–694, July 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kurtz:1979:VFG

[Kur79] Thomas G. Kurtz. A variational formula for the growth rate of a positive operator semigroup. SIAM Journal on Mathematical Analysis, 10(1):112–117, January 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kushner:1988:LPT

[Kus88] Howard B. Kushner. The linearization of the product of two zonal polynomials. SIAM Journal on Mathematical Analysis, 19 (3):687–717, May 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kuttler:1984:NLT

[Kut84] J. R. Kuttler. A nodal line theorem for the sloshing problem. SIAM Journal on Mathematical Analysis, 15(6):1234– 1237, November 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kuwamura:1996:SRS

[Kuw96] Masataka Kuwamura. The stability of roll solutions of the twodimensional Swift–Hohenberg equation and the phase–diffusion equation. SIAM Journal on Mathematical Analysis, 27(5): 1311–1335, September 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/27039.

Kamin:1991:ABS

[KV91] Shoshana Kamin and Juan Luis Vázquez. Asymptotic behaviour of solutions of the porous medium equation with changing sign. SIAM Journal on Mathematical Analysis, 22(1):34–45, January 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Koelink:1998:COP

[KV98] H. T. Koelink and J. Van Der Jeugt. Convolutions for orthogonal polynomials from Lie and quantum algebra representations. SIAM Journal on Mathematical Analysis, 29(3): 794–822, May 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30673.

Kitaev:1999:ASM

[KV99] A. V. Kitaev and A. H. Vartanian. Asymptotics of solutions to the modified nonlinear Schrödinger equation: Solitons on a nonvanishing continuous background. SIAM Journal on Mathematical Analysis, 30(4):787–832, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/33201.

Kwong:1983:OTB

[KW83] Man Kam Kwong and James S. W. Wong. On an oscillation theorem of Belohorec. SIAM Journal on Mathematical Analysis, 14(3):474–476, May 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kaplan:1975:SPS

[KY75] James L. Kaplan and James A.

Yorke. On the stability of a periodic solution of a differential delay equation. SIAM Journal on Mathematical Analysis, 6 (2):268–282, April 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Krall:1988:SSA

[KZ88] Allan M. Krall and Anton Zettl. Singular selfadjoint Sturm–Liouville problems. II. Interior singular points. SIAM Journal on Mathematical Analysis, 19(5):1135–1141, September 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Kong:1995:IOC

[KZ95] Q. Kong and A. Zettl. Interval oscillation conditions for difference equations. SIAM Journal on Mathematical Analysis, 26 (4):1047–1060, July 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/25128.

Kopell:1990:MCC

[KZE90] N. Kopell, W. Zhang, and G. B. Ermentrout. Multiple coupling in chains of oscillators. SIAM Journal on Mathematical Analysis, 21 (4):935–953, July 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

LanzadeCristoforis:1991:LDN

[LA91] Massimo Lanza de Cristoforis and Stuart S. Antman. The large deformation of nonlinearly elastic tubes in two-dimensional flows. SIAM Journal on Mathematical Analysis, 22(5):1193– 1221, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

LeGruyer:1996:SCE

[LA96] E. Le Gruyer and J. C. Archer.

Stability and convergence of extension schemes to continuous functions in general metric spaces. SIAM Journal on Mathematical Analysis, 27(1): 274–285, January 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/25565.

Labouriau:1985:DHB

[Lab85] Isabel Salgado Labouriau. Degenerate Hopf bifurcation and nerve impulse. SIAM Journal on Mathematical Analysis, 16(6): 1121–1133, November 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Labouriau:1989:DHB

[Lab89] Isabel Salgado Labouriau. Degenerate Hopf bifurcation and nerve impulse. Part II. SIAM Journal on Mathematical Analysis, 20(1):1–12, January 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Labreuche:1999:GSR

[Lab99] Christophe Labreuche. Generalization of the Schwarz reflection principle in scattering theory for dissipative systems: Application to purely imaginary resonant frequencies. SIAM Journal on Mathematical Analysis, 30(4):848–878, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32921.

Lada:1994:ISN

[Ład94] A. Łada. Investigations of solutions of nonlinear hyperbolic equations with a small nonlinearity and applications. SIAM Journal on Mathematical Analysis, 25(5):1303–1319, September 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/24251.

Laetsch:1975:ABPa

[Lae75a] Theodore Laetsch. Asymptotic branch points and multiple positive solutions of nonlinear integral equations. SIAM Journal on Mathematical Analysis, 6(1): 168–177, February 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Laetsch:1975:ABPb

[Lae75b] Theodore Laetsch. Asymptotic branch points and multiple positive solutions of nonlinear integral equations. SIAM Journal on Mathematical Analysis, 6(1): 178–191, February 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lagnese:1972:ESS

[Lag72a] John Lagnese. Exponential stability of solutions of differential equations of Sobolev type. SIAM Journal on Mathematical Analysis, 3(4):625–636, November 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lagnese:1972:GBV

[Lag72b] John E. Lagnese. General boundary value problems for differential equations of Sobolev type. SIAM Journal on Mathematical Analysis, 3(1):105–119, February 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lagnese:1973:SDE

[Lag73] John Lagnese. Singular differential equations in Hilbert space. SIAM Journal on Mathematical Analysis, 4(4):623–637, November 1973. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lagnese:1975:PCN

[Lag75] John Lagnese. Perturbations in a class of nonlinear abstract equations. SIAM Journal on Mathematical Analysis, 6(4):616–627, August 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lahtinen:1993:PHI

[Lah93] Aatos Lahtinen. Positive Hermite interpolation by quadratic splines. SIAM Journal on Mathematical Analysis, 24(1):223–233, January 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Laine:1980:PFC

[Lai80] Thomas P. Laine. The product formula and convolution structure for the generalized Chebyshev polynomials. SIAM Journal on Mathematical Analysis, 11 (1):133–146, January 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Laine:1982:PFN

[Lai82] Thomas P. Laine. Projection formulas and a new proof of the addition formula for the Jacobi polynomials. SIAM Journal on Mathematical Analysis, 13 (2):324–330, March 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lakos:1990:ESS

[Lak90]  Nela Lakoš.  Existence of

steady-state solutions for a one-predator–two-prey system. SIAM Journal on Mathematical Analysis, 21(3):647–659, May 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lampone:1978:PPE

[Lam78] Leo W. Lampone. A positive polynomial expansion problem in minimum variance smoothing. SIAM Journal on Mathematical Analysis, 9(3):500–514, June 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lambert:1980:SII

[Lam80] Alan Lambert. Statistical independence and invariant subspaces for point transformation. SIAM Journal on Mathematical Analysis, 11(3):468–472, May 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lange:1974:AAN

[Lan74] Charles G. Lange. Asymptotic analysis of nonlinear diffusion and related multidimensional integrals. SIAM Journal on Mathematical Analysis, 5(3):517–550, May 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Langford:1978:GIU

[Lan78] W. F. Langford. The generalized inverse of an unbounded linear operator with unbounded constraints. SIAM Journal on Mathematical Analysis, 9(6):1083– 1095, December 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Langebartel:1980:FER

[Lan80] R. G. Langebartel. Fourier expansions of rational fractions of elliptic integrals and Jacobian elliptic functions. SIAM Journal on Mathematical Analysis, 11(3): 506–513, May 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Landau:1983:IPV

[Lan83a] H. J. Landau. The inverse problem for the vocal tract and the moment problem. SIAM Journal on Mathematical Analysis, 14(5):

1019–1035, September 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lange:1983:FEA

[Lan83b] L. J. Lange. δ-fraction expansions of analytic functions. SIAM Journal on Mathematical Analysis, 14(2):323–368, March 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Langlais:1985:NPA

[Lan85] Michel Langlais. A nonlinear problem in age-dependent population diffusion. SIAM Journal on Mathematical Analysis, 16(3): 510–529, May 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lang:1996:OWC

[Lan96] W. Christopher Lang. Orthogonal wavelets on the Cantor dyadic group. SIAM Journal on Mathematical Analysis, 27(1): 305–312, January 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/24804.

Larsen:1981:BMP

[Lar81] Edward W. Larsen. Bounds and maximum principles for the solution of the linear transport equation. SIAM Journal on Mathematical Analysis, 12(2):282–292, March 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Larrouturou:1988:EOD

[Lar88] B. Larrouturou. The equations of one-dimensional unsteady flame propagation: existence and uniqueness. SIAM Journal on Mathematical Analysis, 19(1):32–59, January 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Larsen:1998:QOJ

[Lar98] Christopher J. Larsen. Quasiconvexification in W^{1,1} and optimal jump microstructure in bv relaxation. SIAM Journal on Mathematical Analysis, 29 (4):823–848, July 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29599.

Lasser:1983:LPA

[Las83] Rupert Lasser. Linearization of the product of associated Legendre polynomials. SIAM Journal on Mathematical Analysis, 14 (2):403–408, March 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lauwerier:1975:NVW

[Lau75] H. A. Lauwerier. A note on the van Wijngaarden Transformation. SIAM Journal on Mathematical Analysis, 6(1):96–104, February 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Laginestra:1974:CES

[LB74] Anthony V. Laginestra and William E. Boyce. Convergence and evaluation of sums of reciprocal powers of eigenvalues of boundary value problems nonlinear in the eigenvalue parameter. SIAM Journal on Mathematical Analysis, 5(1):64–90, February 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:1993:HOR

[LB93] W. B. Liu and John W. Barrett. Higher-order regularity for the solutions of some degenerate quasilinear elliptic equations in the plane. SIAM Journal on Mathematical Analysis, 24(6): 1522–1536, November 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lohe:1994:BHF

[LB94] M. A. Lohe and L. C. Biedenharn. Basic hypergeometric functions and the Borel–Weil construction for U_{q}(3). SIAM Journal on Mathematical Analysis, 25 (1):218–241, January 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/22639.

Li:1996:ASW

[LB96] Yi A. Li and Jerry L. Bona. Analyticity of solitary-wave solutions of model equations for long waves. SIAM Journal on Mathematical Analysis, 27(3):725–737, May 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Leung:1984:PST

[LC84] Anthony W. Leung and Gen Shun Chen. Positive solutions for temperature-dependent twogroup neutron flux equations: equilibria and stabilities. SIAM Journal on Mathematical Analysis, 15(3):490–499, May 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liou:1988:ULC

[LC88] Lii Perng Liou and Kuo Shung Cheng. On the uniqueness of a limit cycle for a predatorprey system. SIAM Journal on Mathematical Analysis, 19(4): 867–878, July 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lin:1997:GLV

[LD97] Fang-Hua Lin and Qiang Du.

Ginzburg–Landau vortices: Dynamics, pinning, and hysteresis. SIAM Journal on Mathematical Analysis, 28(6):1265– 1293, November 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29806.

Lee:1974:STT

[Lee74] Will Y. K. Lee. On spaces of type H_{μ} and their Hankel transformations. SIAM Journal on Mathematical Analysis, 5(2):336–348, April 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lee:1975:SHT

[Lee75] Will Y. Lee. On Schwartz’s Hankel transformation of certain spaces of distributions. SIAM Journal on Mathematical Analysis, 6(2):427–432, April 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lefevere:1982:NBV

[Lef82] Anne-Marie Lefevere. A nonlinear boundary value problem suggested by the Laplace equation for an elastic and axisymmetric membrane. SIAM Journal on Mathematical Analysis, 13 (1):48–60, January 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Leggett:1976:NAE

[Leg76] Richard W. Leggett. A new approach to the H-equation of Chandrasekhar. SIAM Journal on Mathematical Analysis, 7(4): 542–550, August 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Leitman:1981:IDE

[Lei81] M. J. Leitman. An integrodifferential equation for plane waves propagating into a random fluid: asymptotic behavior. SIAM Journal on Mathematical Analysis, 12(4):560–571, July 1981. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Leitman:1985:ABS

[Lei85] M. J. Leitman. On the asymptotic behavior of solutions to kinetic equations. SIAM Journal on Mathematical Analysis, 16(4): 663–674, July 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Leonard:1982:OPD

[Leo82] Douglas A. Leonard. Orthogonal polynomials, duality and association schemes. SIAM Journal on Mathematical Analysis, 13(4): 656–663, July 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lerner:1990:WWF

[Ler90] Nicolas Lerner. Wick–Wigner functions and tomographic methods. SIAM Journal on Mathematical Analysis, 21(4):1083– 1092, July 1990. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Letessier:1994:SRC

[Let94] Jean Letessier. Some results on co-recursive associated Laguerre and Jacobi polynomials. SIAM Journal on Mathematical Analysis, 25(2):528–548, March 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/22855.

Leung:1973:CFA

[Leu73] Anthony Wing-Kwok Leung. Connection formulas for asymptotic solutions of second order turning points in unbounded domains. SIAM Journal on Mathematical Analysis, 4(1):89–103, February 1973. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Leung:1974:DAS

[Leu74] Anthony Leung. Doubly asymptotic series for n th order differential equations in unbounded domains. SIAM Journal on Mathematical Analysis, 5(2):187–201, April 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Leung:1975:ECF

[Leu75] Anthony Leung. Errata: Connection formulas for asymptotic solutions of second order turning points in unbounded domains. SIAM Journal on Mathematical Analysis, 6(3):600, May 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Levine:1973:UBS

[Lev73] Howard A. Levine. On the uniquenes of bounded solutions to u′(t) = A(t)u(t) and u′′(t) = A(t)u(t) in Hilbert space. SIAM Journal on Mathematical Analysis, 4(2):250–259, May 1973. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Levine:1974:NNT

[Lev74a] Howard A. Levine. A note on a nonexistence theorem for nonlinear wave equations. SIAM Journal on Mathematical Analysis, 5 (4):644–648, August 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Levine:1974:SAR

[Lev74b] Howard A. Levine. Some additional remarks on the nonexistence of global solutions to nonlinear wave equations. SIAM Journal on Mathematical Analysis, 5(1):138–146, February 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Levin:1976:SPB

[Lev76] J. J. Levin. Some a priori bounds for nonlinear Volterra equations. SIAM Journal on Mathematical Analysis, 7(6):872–897, November 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Levine:1983:QSL

[Lev83]  Howard A. Levine.  The

quenching of solutions of linear parabolic and hyperbolic equations with nonlinear boundary conditions. SIAM Journal on Mathematical Analysis, 14(6): 1139–1153, November 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Levine:1988:SIS

[Lev88] Howard A. Levine. Stability and instability for solutions of Burgers’ equation with a semilinear boundary condition. SIAM Journal on Mathematical Analysis, 19 (2):312–336, March 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lew:1974:FWF

[Lew74] John S. Lew. Functions whose Fourier transforms decay at infinity: Qualitative criteria for an additional case. SIAM Journal on Mathematical Analysis, 5(3): 469–472, May 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lewis:1979:ACT

[Lew79] John L. Lewis. Applications of a convolution theorem to Jacobi polynomials. SIAM Journal on Mathematical Analysis, 10(6): 1110–1120, November 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lewanowicz:1986:RRC

[Lew86] Stanisław Lewanowicz. Recurrence relations for the coefficients in Jacobi series solutions of linear differential equations. SIAM Journal on Mathematical Analysis, 17(5):1037–1052, September 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lyashenko:1998:NIP

[LF98] Andrei A. Lyashenko and Susan J. Friedlander. Nonlinear instability of a precessing body with a cavity filled by an ideal fluid. SIAM Journal on Mathematical Analysis, 29 (3):600–618, May 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30216.

Li:1985:TUB

[Li85] Rong Hua Li. A theorem for the uniform boundedness of a family of compact operators. SIAM Journal on Mathematical Analysis, 16(6):1306–1308, November 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:1993:RPC

[Li93] Tong Li. On the Riemann problem for a combustion model. SIAM Journal on Mathematical Analysis, 24(1):59–75, January 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:1997:SSM

[Li97] Ming-Chia Li. Structural stability of Morse–Smale gradientlike flows under discretizations. SIAM Journal on Mathematical Analysis, 28(2):381–388, March 1997. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/29238.

Li:1999:SSE

[Li99] Ming-Chia Li. Structural stability for the Euler method. SIAM Journal on Mathematical Analysis, 30(4):747–755, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32255.

Liang:1994:NID

[Lia94] Jin Liang. On a nonlinear integrodifferential drift-diffusion semiconductor model. SIAM Journal on Mathematical Analysis, 25(5):1375–1392, September 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/23826.

Lim:1978:ABS

[Lim78]  Eng-Bin Lim.  Asymptotic

bounds of solutions of the functional differential equation x′(t) = ax(λt) + bx(t) + f(t), 0 < λ < 1. SIAM Journal on Mathematical Analysis, 9 (5):915–920, October 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ling:1974:SSH

[Lin74] Chih-Bing Ling. On summation of series of hyperbolic functions. SIAM Journal on Mathematical Analysis, 5(4):551–562, August 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ling:1975:CWE

[Lin75a] Chih-Bing Ling. On a class of Weierstrass elliptic functions at half and quarter periods. SIAM Journal on Mathematical Analysis, 6(1):117–128, February 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ling:1975:SSH

[Lin75b] Chih-Bing Ling. On summation of series of hyperbolic functions. II. SIAM Journal on Mathematical Analysis, 6(1):129–139, February 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ling:1978:GCS

[Lin78] Chih-Bing Ling. Generalization of certain summations due to Ramanujan. SIAM Journal on Mathematical Analysis, 9(1):34– 48, February 1978. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lindquist:1986:SRP

[Lin86] W. B. Lindquist. The scalar Riemann problem in two spatial dimensions: piecewise smoothness of solutions and its breakdown. SIAM Journal on Mathematical Analysis, 17(5):1178– 1197, September 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lin:1991:IPC

[Lin91] Yanping Lin. An inverse problem for a class of quasilinear parabolic equations. SIAM Journal on Mathematical Analysis, 22 (1):146–156, January 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lin:1994:ANS

[Lin94] Yan Ping Lin. Analytical and numerical solutions for a class of nonlocal nonlinear parabolic differential equations. SIAM Journal on Mathematical Analysis, 25(6):1577–1594, November 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/24306.

Lin:1995:SIE

[Lin95] Hwei Ting Lin. On a system of integrodifference equations modelling the propagation of genes. SIAM Journal on Mathematical Analysis, 26 (1):35–76, January 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/23708.

Lin:1997:SGS

[Lin97]  Song-Sun Lin.  Stability of

gaseous stars in spherically symmetric motions. SIAM Journal on Mathematical Analysis, 28 (3):539–569, May 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/29288.

Lippman:1972:SMC

[Lip72] G. E. Lippman. Spherical means of conjugate Fourier integrals in E_{2} at the critical index. SIAM Journal on Mathematical Analysis, 3(3):389–400, August 1972.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lippman:1973:SSC

[Lip73]  G. E. Lippman.  Spherical

summability of conjugate multiple Fourier series and integrals at the critical index. SIAM Journal on Mathematical Analysis, 4(4): 681–695, November 1973. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lipton:1998:SSE

[Lip98] Robert Lipton. The second Stekloff eigenvalue and energy dissipation inequalities for functionals with surface energy. SIAM Journal on Mathematical Analysis, 29(3):673–680, May 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/31014.

Liron:1971:RCR

[Lir71a] N. Liron. A recurrence concerning Rayleigh functions. SIAM Journal on Mathematical Analysis, 2(4):496–499, November 1971. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liron:1971:SIS

[Lir71b] N. Liron. Some infinite sums.

SIAM Journal on Mathematical Analysis, 2(1):105–112, February 1971. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:1991:BIE

[Liu91] Y. W. Liu. A boundary integral equation for the two-dimensional floating-body problem. SIAM Journal on Mathematical Analysis, 22(4):973–981, July 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:1995:EPS

[Liu95a] Yong Liu. The equilibrium plasma subject to skin effect. SIAM Journal on Mathematical Analysis, 26(5):1157–1183, September 1995. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

                                        http://epubs.siam.org/sam-
                                        bin/dbq/article/24931.

Liu:1995:IBS

[Liu95b] Yue Liu. Instability and blowup of solutions to a generalized Boussinesq equation. SIAM Journal on Mathematical Analysis, 26(6):1527–1546, November 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/25809.

Liu:1996:DST

[Liu96] Youming Liu. A distributional sampling theorem. SIAM Journal on Mathematical Analysis, 27 (4):1153–1157, July 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/26693.

Liu:1997:GES

[Liu97] Changmei Liu. Global estimates for solutions of partial differential equations. SIAM Journal on Mathematical Analysis, 28 (4):924–943, July 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29062.

Lei:1991:APE

[LJ91] Junjiang Lei and Rong Qing Jia. Approximation by piecewise exponentials. SIAM Journal on Mathematical Analysis, 22(6): 1776–1789, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lei:1997:ASI

[LJC97] Junjiang Lei, Rong-Qing Jia, and E. W. Cheney. Approximation from shift-invariant spaces by integral operators. SIAM Journal on Mathematical Analysis, 28 (2):481–498, March 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/27986.

Lu:1989:EST

[LK89] Chun Qing Lu and Nicholas D.

Kazarinoff. On the existence of solutions of a two-point boundary value problem arising from flows in a cylindrical floating zone. SIAM Journal on Mathematical Analysis, 20(2):494–503, March 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lu:1988:ESS

[LKMT88] Chun Qing Lu, Nicholas D. Kazarinoff, J. Bryce McLeod, and William C. Troy. Existence of solutions of the similarity equations for floating rectangular cavities and disks. SIAM Journal on Mathematical Analysis, 19(5):1119–1126, September 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ladeveze:1979:LBF

[LL79] J. Ladevèze and P. Ladevèze. Lower bounds for the fundamental free membrane eigenvalue. SIAM Journal on Mathematical Analysis, 10(6):1280–1292, November 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Levin:1990:MBI

[LL90] A. L. Levin and D. S. Lubinsky. L_{∞} Markov and Bernstein inequalities for Freud weight s. SIAM Journal on Mathematical Analysis, 21(4):1065–1082, July 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:1993:SNE

[LL93] Lige Li and Yaping Liu. Spectral and nonlinear effects in certain elliptic systems of three variables. SIAM Journal on Mathematical Analysis, 24(2):480–498, March 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lloyd:1980:HAS

[Llo80] Stuart P. Lloyd. Hamming association schemes and codes on spheres. SIAM Journal on Mathematical Analysis, 11(3): 488–505, May 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lawton:1997:SOM

[LLS97] W. Lawton, S. L. Lee, and Zuowei Shen. Stability and orthonormality of multivariate refinable functions. SIAM Journal on Mathematical Analysis, 28 (4):999–1014, July 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/28815.

Levine:1975:ABL

[LM75] Howard A. Levine and Amy C.

Murray. Asymptotic behavior and lower bounds for semilinear wave equations in Hilbert space with applications. SIAM Journal on Mathematical Analysis, 6 (5):846–859, October 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lewis:1977:MCP

[LM77] J. T. Lewis and M. E. Muldoon. Monotonicity and convexity properties of zeros of Bessel functions. SIAM Journal on Mathematical Analysis, 8(1): 171–178, February 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Levine:1980:QSS

[LM80] Howard A. Levine and John T. Montgomery. The quenching of solutions of some nonlinear parabolic equations. SIAM Journal on Mathematical Analysis, 11(5):842–847, September 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Leviatan:1982:RMS

[LM82] D. Leviatan and H. N. Mhaskar. The rate of monotone spline approximation in the L_{p}-norm. SIAM Journal on Mathematical Analysis, 13(5):866–874, September 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Laforgia:1983:IAZ

[LM83] Andrea Laforgia and Martin E. Muldoon. Inequalities and approximations for zeros of Bessel functions of small order. SIAM Journal on Mathematical Analysis, 14(2):383–388, March 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lubinsky:1995:NSC

[LM95] D. S. Lubinsky and D. M. Matjila. Necessary and sufficient conditions for mean convergence of Lagrange interpolation for Freud weights. SIAM Journal on Mathematical Analysis, 26 (1):238–262, January 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23705.

Li:1996:GAG

[LM96] Michael Y. Li and James S.

Muldowney. A geometric approach to global-stability problems. SIAM Journal on Mathematical Analysis, 27(4):1070– 1083, July 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/26644.

Lubinsky:1987:QSI

[LMN87] D. S. Lubinsky, A. Máté, and P. Nevai. Quadrature sums involving p th powers of polynomials. SIAM Journal on Mathematical Analysis, 18(2):531–544, March 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:1998:BSB

[LMN98] Tai-Ping Liu, Akitaka Matsumura, and Kenji Nishihara. Behaviors of solutions for the Burgers equation with boundary corresponding to rarefaction waves. SIAM Journal on Mathematical Analysis, 29(2):293–308, March 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/30600.

Leaute:1990:CPS

[LMPS90] B. Leaute, G. Marcilhacy, R. Pons, and J. Skinazi. On the connection problem for some Schrödinger equations in relation to the biconfluent Heun differential equation. SIAM Journal on Mathematical Analysis, 21(3): 793–798, May 1990. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lorch:1973:SMP

[LMS73] Lee Lorch, M. E. Muldoon, and Peter Szego. Some monotonicity properties of Bessel functions. SIAM Journal on Mathematical Analysis, 4(2):385–392, May 1973. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lau:1996:SSR

[LMW96] Ka-Sing Lau, Mang-Fai Ma, and Jianrong Wang. On some sharp regularity estimations of L_{2}- scaling functions. SIAM Journal on Mathematical Analysis, 27(3): 835–864, May 1996. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

LopesFilho:1998:EMB

[LN98] M. C. Lopes Filho and H. J. Nussenzveig Lopes. An extension of Marchioro’s bound on the growth of a vortex patch to flows with L^{p} vorticity. SIAM Journal on Mathematical Analysis, 29 (3):596–599, May 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/31091.

Lo:1972:SES

[Lo72] Chi Yeung Lo. Series expansions of solutions of U_{xx}+U_{yy}+ε^{2}U_{tt} = U_{t}. SIAM Journal on Mathematical Analysis, 3(3):461–473, August 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Logan:1979:IMN

[Log79a] B. F. Logan. Inequalities and minimum norm kernels for the Hardy class H_{p}. SIAM Journal on Mathematical Analysis, 10(4): 741–751, July 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Logan:1979:LCT

[Log79b] B. F. Logan. Limits in L_{p} of convolution transforms with kernels aK(at), a → 0. SIAM Journal on Mathematical Analysis, 10(4): 733–740, July 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Logan:1979:NCW

[Log79c] B. F. Logan. Note on construction of weight functions. SIAM Journal on Mathematical Analysis, 10(4):752–756, July 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Logan:1983:EHT

[Log83a] B. F. Logan. Erratum: “Hilbert transform of a function having a bounded integral and a bounded derivative”. SIAM Journal on Mathematical Analysis, 14(4):845, July 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [Log83e].

Logan:1983:EPPa

[Log83b] B. F. Logan. Extremal problems for positive-definite bandlimited functions. I. Eventually positive functions with zero integral. SIAM Journal on Mathematical Analysis, 14(2):249–252, March 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Logan:1983:EPPb

[Log83c] B. F. Logan. Extremal problems for positive-definite bandlimited functions. II. Eventually negative functions. SIAM Journal on Mathematical Analysis, 14 (2):253–257, March 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Logan:1983:EPPc

[Log83d] B. F. Logan. Extremal problems for positive-definite bandlimited functions. III. The maximum number of zeros in an interval [0, T]. SIAM Journal on Mathematical Analysis, 14 (2):258–268, March 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Logan:1983:HTF

[Log83e] B. F. Logan. Hilbert transform of a function having a bounded integral and a bounded derivative. SIAM Journal on Mathematical Analysis, 14(2):247–248, March 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [Log83a].

Logan:1983:IEC

[Log83f] B. F. Logan. An integral equation connected with the Jacobi polynomials. SIAM Journal on Mathematical Analysis, 14 (2):269–322, March 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Logan:1984:EIE

[Log84a] B. F. Logan. Errata: “An integral equation connected with the Jacobi polynomials” [SIAM J. Math. Anal. 14 (1983), no. 2, 269–322; MR 84f:45006]. SIAM Journal on Mathematical Analysis, 15(1):213, January 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Logan:1984:IHP

[Log84b] B. F. Logan. Integrals of highpass functions. SIAM Journal on Mathematical Analysis, 15 (2):389–405, March 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Logan:1984:ECI

[Log84c] B. F. Logan. On the eigenvalues of a certain integral equation. SIAM Journal on Mathematical Analysis, 15(4):712–717, July 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Logan:1985:NCC

[Log85] B. F. Logan. The norm of certain convolution transforms on L_{p} spaces of entire functions of exponential type. SIAM Journal on Mathematical Analysis, 16 (1):167–179, January 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Logan:1988:BTS

[Log88] B. F. Logan. Bounds for the tails of sharp-cutoff filter kernels. SIAM Journal on Mathematical Analysis, 19(2):372–376, March 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Logan:1990:ROP

[Log90] B. F. Logan. The recovery of orthogonal polynomials from a sum of squares. SIAM Journal on Mathematical Analysis, 21 (4):1031–1050, July 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Londen:1971:ABS

[Lon71] Stig-Olof Londen. On the asymptotic behavior of the solution of a nonlinear integrodifferential equation. SIAM Journal on Mathematical Analysis, 2(3): 356–367, August 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Londen:1973:SNT

[Lon73]  Stig-Olof Londen.  Some

nonoscillation theorems for a second order nonlinear differential equation. SIAM Journal on Mathematical Analysis, 4(3): 460–465, May 1973. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Londen:1974:ABB

[Lon74a] Stig-Olof Londen. On the asymptotic behavior of the bounded solutions of a nonlinear Volterra equation. SIAM Journal on Mathematical Analysis, 5(6): 849–875, November 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Longman:1974:BRF

[Lon74b] I. M. Longman. Best rational function approximation for Laplace transform inversion. SIAM Journal on Mathematical Analysis, 5(4):574–580, August 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Londen:1977:IEH

[Lon77] Stig-Olof Londen. On an integral equation in a Hilbert space. SIAM Journal on Mathematical Analysis, 8(6):950–970, November 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Londen:1983:SIE

[Lon83] Stig-Olof Londen. On some integral equations with locally finite measures and L^{∞}-perturbations. SIAM Journal on Mathematical Analysis, 14(6):1187–1203, November 1983. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Londen:1987:VES

[Lon87] Stig-Olof Londen. A Volterra equation with L^{2}-solutions. SIAM Journal on Mathematical Analysis, 18(1):168–171, January 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Longinetti:1988:SII

[Lon88] Marco Longinetti. Some isoperimetric inequalities for the level curves of capacity and Green’s functions on convex plane domains. SIAM Journal on Mathematical Analysis, 19(2):377–389, March 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lopes:1989:DSM

[Lop89] Artur O. Lopes. The dimension spectrum of the maximal measure. SIAM Journal on Mathematical Analysis, 20(5):1243– 1254, September 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lorch:1980:TWZ

[Lor80] Lee Lorch. Turánians and Wronskians for the zeros of Bessel functions. SIAM Journal on Mathematical Analysis, 11(2): 223–227, March 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lorch:1993:SIF

[Lor93] Lee Lorch. Some inequalities for the first positive zeros of Bessel functions. SIAM Journal on Mathematical Analysis, 24(3): 814–823, May 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Loud:1980:BAG

[Lou80] W. S. Loud. A bifurcation application of the generalized inverse of a linear differential operator. SIAM Journal on Mathematical Analysis, 11(3):545–558, May 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Louis:1984:OFS

[Lou84] Alfred K. Louis. Orthogonal function series expansions and the null space of the Radon transform. SIAM Journal on Mathematical Analysis, 15(3): 621–633, May 1984. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Louisell:1994:VFA

[Lou94] James Louisell. A velocity functional for an analysis of stability in delay-differential equations. SIAM Journal on Mathematical Analysis, 25(4):1181– 1194, July 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23319.

Levine:1976:NES

[LP76] Howard A. Levine and Lawrence E. Payne. On the nonexistence of entire solutions to nonlinear second order elliptic equations. SIAM Journal on Mathematical Analysis, 7(3):337–343, May 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Luning:1981:PSN

[LP81] C. D. Luning and W. L. Perry. Positive solutions of negative exponent generalized Emden– Fowler boundary value problems. SIAM Journal on Mathematical Analysis, 12(6):874–879, November 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lin:1991:EMP

[LP91] Song-Sun Lin and Feng Ming Pai. Existence and multiplicity of positive radial solutions for semilinear elliptic equations in annular domains. SIAM Journal on Mathematical Analysis, 22(6): 1500–1515, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lin:1994:IDD

[LP94] Chang Hao Lin and L. E. Payne. The influence of domain and diffusivity perturbations on the decay of end effects in heat conduction. SIAM Journal on Mathematical Analysis, 25(5):1241– 1258, September 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23355.

Lowe:1992:RPF

[LPR92] Bruce D. Lowe, Michael Pilant, and William Rundell. The recovery of potentials from finite spectral data. SIAM Journal on Mathematical Analysis, 23 (2):482–504, March 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Levine:1989:ACR

[LPSS89] H. A. Levine, L. E. Payne, P. E.

Sacks, and B. Straughan. Analysis of a convective reactiondiffusion equation. II. SIAM Journal on Mathematical Analysis, 20(1):133–147, January 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Levinson:1972:STH

[LR72] Norman Levinson and Ray Redheffer. Schur’s theorem for Hurwitz polynomials. SIAM Journal on Mathematical Analysis, 3 (1):17–19, February 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lancaster:1977:SNO

[LR77] Peter Lancaster and Jon G.

Rokne. Solutions of nonlinear operator equations. SIAM Journal on Mathematical Analysis, 8(3): 448–457, May 1977. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Little:1984:EAK

[LR84] G. Little and J. B. Reade. Eigenvalues of analytic kernels. SIAM Journal on Mathematical Analysis, 15(1):133–136, January 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lalli:1971:BSS

[LS71] B. S. Lalli and W. A. Skrapek. On the boundedness and the stability of some differential equations of the fourth order. SIAM Journal on Mathematical Analysis, 2(2):221–225, May 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lundin:1979:CTA

[LS79] L. Lundin and F. Stenger. Cardinal-type approximations of a function and its derivatives. SIAM Journal on Mathematical Analysis, 10(1):139–160, January 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lorenz:1986:SON

[LS86] Jens Lorenz and Richard Sanders. Second order nonlinear singular perturbation problems with boundary conditions of mixed type. SIAM Journal on Mathematical Analysis, 17(3):580–594, May 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lorenz:1987:RCV

[LS87] Jens Lorenz and Richard Sanders. On the rate of convergence of viscosity solutions for boundary value problems. SIAM Journal on Mathematical Analysis, 18 (2):306–320, March 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lasiecka:1988:RSC

[LS88a] Irena Lasiecka and Jan Sokołowski.

Regularity and strong convergence of a variational approximation to a nonhomogeneous Dirichlet hyperbolic boundary value problem. SIAM Journal on Mathematical Analysis, 19(3): 528–540, May 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Levin:1988:DAR

[LS88b] A. L. Levin and E. B. Saff. Degree of approximation of real functions by reciprocals of real and complex polynomials. SIAM Journal on Mathematical Analysis, 19(1):233–245, January 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lorch:1988:ZDB

[LS88c] Lee Lorch and Peter Szego.

On the zeros of derivatives of Bessel functions. SIAM Journal on Mathematical Analysis, 19(6): 1450–1454, November 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lasiecka:1989:SAH

[LS89] I. Lasiecka and J. Sokołowski.

Semidiscrete approximations of hyperbolic boundary value problems with nonhomogeneous Dirichlet boundary conditions. SIAM Journal on Mathematical Analysis, 20(6):1366–1387, November 1989. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lubinsky:1993:MBN

[LS93] D. S. Lubinsky and E. B. Saff. Markov–Bernstein and Nikol′skiĭ inequalities, and Christoffel functions for exponential weights on (−1, 1). SIAM Journal on Mathematical Analysis, 24(2): 528–556, March 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lorch:1994:BMZ

[LS94] Lee Lorch and Peter Szego. Bounds and monotonicities for the zeros of derivatives of ultraspherical Bessel functions. SIAM Journal on Mathematical Analysis, 25(2):549–554, March 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/23145.

Legner:1996:TPQ

[LS96] Mary M. Legner and Victor L. Shapiro. Time-periodic quasilinear reaction-diffusion equations. SIAM Journal on Mathematical Analysis, 27(1):135–169, January 1996. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:1992:GRP

[LSZ92] Ta-Tsien Li, D. Serre, and Hao Zhang. The generalized Riemann problem for the motion of elastic strings. SIAM Journal on Mathematical Analysis, 23(5): 1189–1203, September 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lavoie:1981:JMP

[LT81] J. L. Lavoie and R. Tremblay. The Jabotinsky matrix of a power series. SIAM Journal on Mathematical Analysis, 12(6): 819–825, November 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lasiecka:1992:RDE

[LT92] I. Lasiecka and R. Triggiani. Riccati differential equations with unbounded coefficients and nonsmooth terminal condition — the case of analytic semigroups. SIAM Journal on Mathematical Analysis, 23(2):449–481, March 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lubinsky:1994:BWP

[LT94] Doron S. Lubinsky and Vilmos Totik. Best weighted polynomial approximation via Jacobi expansions. SIAM Journal on Mathematical Analysis, 25 (2):555–570, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23290.

LeFloch:1999:RWL

[LT99] Philippe G. LeFloch and Athanasios E. Tzavaras. Representation of weak limits and definition of nonconservative products. SIAM Journal on Mathematical Analysis, 30(6):1309–1342, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/34179.

Lu:1987:MSS

[Lu87] Chun Qing Lu. Multiple steady states in a biochemical system. SIAM Journal on Mathematical Analysis, 18(6):1771–1783, November 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lu:1990:EBL

[Lu90a] Chun Qing Lu. Existence, bifurcation, and limit of solutions of the similarity equations for floating rectangular cavities and disks. SIAM Journal on Mathematical Analysis, 21(3):721–728, May 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lu:1990:GBS

[Lu90b] Chun Qing Lu. Global bifurcation of steady-state solutions on a biochemical system. SIAM Journal on Mathematical Analysis, 21 (1):76–84, January 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lu:1997:ASL

[Lu97] Chunqing Lu. On the asymptotic solution of laminar channel flow with large suction. SIAM Journal on Mathematical Analysis, 28(5):1113–1134, September 1997. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/29770.

Lu:1999:SDS

[Lu99] Xuguang Lu. Spatial decay solutions of the Boltzmann equation: Converse properties of long time limiting behavior. SIAM Journal on Mathematical Analysis, 30(5):1151–1174, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/33498.

Lubich:1986:DFC

[Lub86] Ch. Lubich. Discretized fractional calculus. SIAM Journal on Mathematical Analysis, 17(3): 704–719, May 1986. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lucier:1988:RTA

[Luc88] Bradley J. Lucier. Regularity through approximation for scalar conservation laws. SIAM Journal on Mathematical Analysis, 19(4): 763–773, July 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lui:1982:NIOa

[Lui82a] Roger Lui. A nonlinear integral operator arising from a model in population genetics. I. Monotone initial data. SIAM Journal on Mathematical Analysis, 13(6): 913–937, November 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lui:1982:NIOb

[Lui82b] Roger Lui. A nonlinear integral operator arising from a model in population genetics. II. Initial data with compact support. SIAM Journal on Mathematical Analysis, 13(6):938–953, November 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lui:1985:NIO

[Lui85] Roger Lui. A nonlinear integral operator arising from a model in population genetics. III. Heterozygote inferior case. SIAM Journal on Mathematical Analysis, 16(6):1180–1206, November 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lui:1986:NIO

[Lui86] Roger Lui. A nonlinear integral operator arising from a model in population genetics. IV. Clines. SIAM Journal on Mathematical Analysis, 17(1):152–168, January 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Luke:1970:EGF

[Luk70] Yudell L. Luke. Evaluation of the gamma function by means of padé approximations. SIAM Journal on Mathematical Analysis, 1(2):266–281, May 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Luke:1975:EPA

[Luk75] Yudell L. Luke. On the error in the padé approximants for a form of the incomplete gamma function including the exponential function. SIAM Journal on Mathematical Analysis, 6(5): 829–839, October 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Luning:1978:ITO

[Lun78] C. D. Luning. An iterative technique for obtaining solutions of a Thomas–Fermi equation. SIAM Journal on Mathematical Analysis, 9(3):515–523, June 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lunardi:1987:LSR

[Lun87] Alessandra Lunardi. Local stability results for the elastic beam equation. SIAM Journal on Mathematical Analysis, 18(5): 1341–1366, September 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lunardi:1990:LHE

[Lun90] Alessandra Lunardi. On the linear heat equation with fading memory. SIAM Journal on Mathematical Analysis, 21(5): 1213–1224, September 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Luther:1978:TRT

[Lut78] Wolfram Luther. A Tauberian remainder theorem for the Hankel transform. SIAM Journal on Mathematical Analysis, 9(4): 595–599, August 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lenoir:1992:VFD

[LVLH92] M. Lenoir, M. Vullierme-Ledard, and C. Hazard. Variational formulations for the determination of resonant states in scattering problems. SIAM Journal on Mathematical Analysis, 23(3): 579–608, May 1992. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lee:1977:SOS

[LW77] T. C. Lee and D. Willett. Second order singular boundary value problems. SIAM Journal on Mathematical Analysis, 8 (4):741–755, August 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lowengrub:1979:SGA

[LW79] M. Lowengrub and J. Walton. Systems of generalized Abel equations. SIAM Journal on Mathematical Analysis, 10(4): 794–807, July 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Li:1994:EBA

[LW94]  X. Li and R. Wong.  Er-

ror bounds for asymptotic expansions of Laplace convolutions. SIAM Journal on Mathematical Analysis, 25(6):1537– 1553, November 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24585.

Lau:1995:CST

[LW95] Ka-Sing Lau and Jianrong Wang.

Characterization of L^{p}-solutions for the two-scale dilation equations. SIAM Journal on Mathematical Analysis, 26(4):1018– 1046, July 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23877.

Liu:1998:SRM

[LWY98] Hailiang Liu, Jinghua Wang, and Tong Yang. Stability of a relaxation model with a nonconvex flux. SIAM Journal on Mathematical Analysis, 29 (1):18–29, January 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29903.

Levesley:1996:COR

[LXLC96] Jeremy Levesley, Yuan Xu, Will Light, and Ward Cheney. Convolution operators for radial basis approximation. SIAM Journal on Mathematical Analysis, 27 (1):286–304, January 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/25523.

Logan:1970:NCI

[LY70] N. A. Logan and K. S. Yee. Note on contour integral representations for products of Airy functions. SIAM Journal on Mathematical Analysis, 1(1):115–117, February 1970. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Liu:1995:NSS

[LY95] Tai-Ping Liu and Long An Ying. Nonlinear stability of strong detonations for a viscous combustion model. SIAM Journal on Mathematical Analysis, 26 (3):519–528, May 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/25906.

Lasota:1996:WLT

[LY96] Andrzej Lasota and James A. Yorke. When the long-time behavior is independent of the initial density. SIAM Journal on Mathematical Analysis, 27(1): 221–240, January 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Lin:1995:SDS

[LYC95] Charles S. C. Lin, Bin Yang, and Fudong Chen. On stability of a dynamical system. SIAM Journal on Mathematical Analysis, 26 (2):428–435, March 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23832.

Levi:1995:BSQ

[LZ95] M. Levi and E. Zehnder. Boundedness of solutions for quasiperiodic potentials. SIAM Journal on Mathematical Analysis, 26(5): 1233–1256, September 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24907.

Mikelic:1988:HSF

[MA88] A. Mikelić and I. Aganović. Homogenization of stationary flow of miscible fluids in a domain with a grained boundary. SIAM Journal on Mathematical Analysis, 19(2):287–294, March 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Maass:1996:FOT

[Maa96] Peter Maass. Families of orthogonal two-dimensional wavelets. SIAM Journal on Mathematical Analysis, 27(5):1454–1481, September 1996. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

http://epubs.siam.org/sam-
bin/dbq/article/24649.

Macdonald:1982:SCR

[Mac82] I. G. Macdonald. Some conjectures for root systems. SIAM Journal on Mathematical Analysis, 13(6):988–1007, November 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

MacGillivray:1990:JMT

[Mac90] A. D. MacGillivray. Justification of matching with the transition expansion of van der Pol’s equation. SIAM Journal on Mathematical Analysis, 21(1):221–240, January 1990. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Maddocks:1985:RQF

[Mad85a] John H. Maddocks. Restricted quadratic forms and their application to bifurcation and stability in constrained variational principles. SIAM Journal on Mathematical Analysis, 16(1): 47–68, January 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See errata [Mad88].

Madych:1985:LDC

[Mad85b] W. R. Madych. Limits of dilated convolution transforms. SIAM Journal on Mathematical Analysis, 16(3):551–558, May 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Maddocks:1988:ERQ

[Mad88] John H. Maddocks. Errata: “Restricted quadratic forms and their application to bifurcation and stability in constrained variational principles” [SIAM J.

Math. Anal. 16 (1985), no. 1, 47–68; MR 86f:49111]. SIAM Journal on Mathematical Analysis, 19(5):1256–1257, September 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [Mad85a].

Magalhaes:1987:PSH

[Mag87] Luis T. Magalhães. Presistence and smoothness of hyperbolic invariant manifolds for functional differential equations. SIAM Journal on Mathematical Analysis, 18(3):670–693, May 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mair:1994:TRF

[Mai94] B. A. Mair. Tikhonov regularization for finitely and infinitely smoothing operators. SIAM Journal on Mathematical Analysis, 25(1):135–147, January 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/23806.

Maki:1976:BDP

[Mak76] Daniel P. Maki. On birthdeath processes with rational growth rates. SIAM Journal on Mathematical Analysis, 7(1):29– 36, February 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Manfredi:1991:ISH

[Man91] Juan J. Manfredi. Isolated singularities of p-harmonic functions in the plane. SIAM Journal on Mathematical Analysis, 22 (2):424–439, March 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mao:1997:RTT

[Mao97] Xuerong Mao. Razumikhintype theorems on exponential stability of neutral stochastic functional differential equations. SIAM Journal on Mathematical Analysis, 28(2):389–401, March 1997. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/29083.

Martin:1970:ABA

[Mar70] R. H. Martin, Jr. On the asymptotic behavior of autonomous differential equations. SIAM Journal on Mathematical Analysis, 1(4):507–514, November 1970. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Martin:1978:ABS

[Mar78] R. H. Martin, Jr. Asymptotic behavior for semilinear differential equations in Banach spaces. SIAM Journal on Mathematical Analysis, 9(6):1105–1119, December 1978. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Martin:1980:CPQ

[Mar80] D. H. Martin. Conditional positivity of quadratic forms in Hilbert space. SIAM Journal on Mathematical Analysis, 11(6): 1047–1057, November 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Marcati:1981:ABA

[Mar81] Pierangelo Marcati. Asymptotic behavior in age-dependent population dynamics with hereditary renewal law. SIAM Journal on Mathematical Analysis, 12(6): 904–916, November 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Markowich:1982:TAS

[Mar82] Peter A. Markowich. A theory for the approximation of solutions of boundary value problems on infinite intervals. SIAM Journal on Mathematical Analysis, 13(3): 484–513, May 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Markett:1983:CTI

[Mar83a] C. Markett. Cohen type inequalities for Jacobi, Laguerre and Hermite expansions. SIAM Journal on Mathematical Analysis, 14(4): 819–833, July 1983. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Markowich:1983:ABV

[Mar83b] Peter A. Markowich. Analysis of boundary value problems on infinite intervals. SIAM Journal on Mathematical Analysis, 14 (1):11–37, January 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Markoe:1984:FIA

[Mar84] Andrew Markoe. Fourier inversion of the attenuated xray transform. SIAM Journal on Mathematical Analysis, 15(4): 718–722, July 1984. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Markowich:1985:NEP

[Mar85a] Peter A. Markowich. A nonlinear eigenvalue problem modelling the avalanche effect in semiconductor diodes. SIAM Journal on Mathematical Analysis, 16(6): 1268–1283, November 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Marti:1985:LCF

[Mar85b] J. T. Marti. The least constant in Friedrichs’ inequality in one dimension. SIAM Journal on Mathematical Analysis, 16(1): 148–150, January 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Markett:1986:NPW

[Mar86] C. Markett. A new proof of Watson’s product formula for Laguerre polynomials via a Cauchy problem associated with a singular differential operator. SIAM Journal on Mathematical Analysis, 17(4):1010–1032, July 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Marion:1989:FDA

[Mar89a] Martine Marion. Finitedimensional attractors associated with partly dissipative reaction-diffusion systems. SIAM Journal on Mathematical Analysis, 20(4):816–844, July 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Marshall:1989:RSN

[Mar89b] Bernard Marshall. The regularity of solutions of nonlinear wave equations. SIAM Journal on Mathematical Analysis, 20 (2):415–429, March 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Masani:1978:DPH

[Mas78] P. Masani. Dilations as propagators of Hilbertian varieties. SIAM Journal on Mathematical Analysis, 9(3):414–456, June 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mathsen:1970:ISS

[Mat70] Ronald M. Mathsen. Interpolation in the solution sets of ordinary differential equations. SIAM Journal on Mathematical Analysis, 1(1):41–51, February 1970. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Matsumoto:1978:IFT

[Mat78] T. Matsumoto. An implicit function theorem and its applications to nonlinear electrical networks. SIAM Journal on Mathematical Analysis, 9(3):492–499, June 1978. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

May:1970:PFN

[May70] L. E. May. Perturbations in fully nonlinear systems. SIAM Journal on Mathematical Analysis, 1 (3):376–391, August 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

May:1979:STM

[May79] C. P. May. A saturation theorem for modified Bernstein polynomials in L_{p}-spaces. SIAM Journal on Mathematical Analysis, 10 (2):321–330, March 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Morgan:1991:ACFa

[MB91a] R. C. Morgan and I. Babuška. An approach for constructing families of homogenized equations for periodic media. I. An integral representation and its consequences. SIAM Journal on Mathematical Analysis, 22(1):1– 15, January 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Morgan:1991:ACFb

[MB91b] R. C. Morgan and I. Babuška. An approach for constructing families of homogenized equations for periodic media. II. Properties of the kernel. SIAM Journal on Mathematical Analysis, 22 (1):16–33, January 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mas-Colell:1979:HCC

[MC79] Andreu Mas-Colell. Homeomorphisms of compact, convex sets and the Jacobian matrix. SIAM Journal on Mathematical Analysis, 10(6):1105–1109, November 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

McBride:1975:TFI

[McB75] Adam C. McBride. A theory of fractional integration for generalized functions. SIAM Journal on Mathematical Analysis, 6(3): 583–599, May 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

McCann:1977:IZB

[McC77] Roger C. McCann. Inequalities for the zeros of Bessel functions. SIAM Journal on Mathematical Analysis, 8(1):166–170, February 1977. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

McCalla:1978:ZSF

[McC78] Clement McCalla. Zeros of the solutions of first-order functional-differential equations. SIAM Journal on Mathematical Analysis, 9(5):843–847, October 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

McCracken:1981:RPS

[McC81] Marjorie McCracken. The resolvent problem for the Stokes equations on halfspace in L_{p}. SIAM Journal on Mathematical Analysis, 12(2):201–228, March 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

McCall:1983:MGS

[McC83] James D. McCall. Mean growth, H_{p} spaces and subharmonic functions in the upper halfplane. SIAM Journal on Mathematical Analysis, 14(5):982–986, September 1983. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

McCoy:1997:CTS

[McC97] Peter A. McCoy. A classical theorem on the singularities of Legendre series in C^{3} and associated system of hyperbolic partial differential equations. SIAM Journal on Mathematical Analysis, 28 (3):704–714, May 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/27349.

McLaughlin:1976:IEP

[McL76] Joyce R. McLaughlin. An inverse eigenvalue problem of order four. SIAM Journal on Mathematical Analysis, 7(5):646–661, October 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

McLaughlin:1978:IEP

[McL78] Joyce R. McLaughlin. An inverse eigenvalue problem of order four — an infinite case. SIAM Journal on Mathematical Analysis, 9(3):

395–413, June 1978. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

McLean:1990:IEM

[McL90] William McLean. An integral equation method for a problem with mixed boundary conditions. SIAM Journal on Mathematical Analysis, 21(4):917–934, July 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Meidan:1972:PTD

[Mei72] Reuven Meidan. A proof of a theorem by dolezal. SIAM Journal on Mathematical Analysis, 3 (3):426–427, August 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Meidan:1976:CTV

[Mei76a] Reuvan Meidan. Causality of time-varying nonlinear operators and the extendibility of their domains. SIAM Journal on Mathematical Analysis, 7(4):473–480, August 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Meidan:1976:DTB

[Mei76b] Reuven Meidan. Distribution theory as the basis of a postulational foundation of linear optical systems. SIAM Journal on Mathematical Analysis, 7(6):837–842, November 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Meidan:1977:LCO

[Mei77] Reuven Meidan. Linear continuous operators in L^{p} and generalized random processes: A kernel representation. SIAM Journal on Mathematical Analysis, 8 (2):337–345, April 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Meidan:1979:RKH

[Mei79] Reuven Meidan. Reproducingkernel Hilbert spaces of distributions and generalized stochastic processes. SIAM Journal on Mathematical Analysis, 10 (1):62–70, January 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Melkman:1985:WOI

[Mel85]  Avraham A. Melkman.  n-

widths and optimal interpolation of time- and band-limited functions. II. SIAM Journal on Mathematical Analysis, 16(4): 803–813, July 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Memory:1989:BAB

[Mem89] Margaret C. Memory. Bifurcation and asymptotic behavior of solutions of a delay-differential equation with diffusion. SIAM Journal on Mathematical Analysis, 20(3):533–546, May 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mercer:1975:CCF

[Mer75a] A. McD. Mercer. The characterization of the cylinder functions by a functional equation. SIAM Journal on Mathematical Analysis, 6(3):488–495, May 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mercer:1975:GIB

[Mer75b] A. McD. Mercer. Grünbaum’s inequality for Bessel functions and its extensions. SIAM Journal on Mathematical Analysis, 6(6): 1021–1023, November 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mercer:1976:CSF

[Mer76] A. McD. Mercer. The characterization of solutions of ∇^{2}ϕ + λ^{2}ϕ = 0 by a functional equation. SIAM Journal on Mathematical Analysis, 7(2):271–275, April 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mercer:1977:IRI

[Mer77] A. McD. Mercer. Integral representations and inequalities for Bessel functions. SIAM Journal on Mathematical Analysis, 8(3): 486–490, May 1977. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mercier:1987:AAO

[Mer87] B. Mercier. Application of accretive operators theory to the radiative transfer equations. SIAM Journal on Mathematical Analysis, 18(2):393–408, March 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Meyer:1974:RDS

[Mey74] H. D. Meyer. A representation for a distributional solution of the heat equation. SIAM Journal on Mathematical Analysis, 5 (4):708–722, August 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Meyer:1976:HPR

[Mey76] H. D. Meyer. Half-plane representations and harmonic continuation. SIAM Journal on Mathematical Analysis, 7(5):713–722, October 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Meyer:1977:RDS

[Mey77] Harold D. Meyer. A representation for distributional solutions of parabolic problems. SIAM Journal on Mathematical Analysis, 8(2):356–367, April 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Meyer:1984:SHS

[Mey84] K. R. Meyer. Scaling Hamiltonian systems. SIAM Journal on Mathematical Analysis, 15(5): 877–889, September 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Meyer:1986:RCF

[Mey86] Burnett Meyer. Real continued fractions and asymptotic expansions. SIAM Journal on Mathematical Analysis, 17(5):1218– 1221, September 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miller:1971:SSV

[MF71] Richard K. Miller and Alan Feldstein. Smoothness of solutions of Volterra integral equations with weakly singular kernels. SIAM Journal on Mathematical Analysis, 2(2):242–258, May 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Moran:1970:RDE

[MG70] M. J. Moran and R. A. Gaggioli. On the reduction of differential equations to algebraic equations. SIAM Journal on Mathematical Analysis, 1(1):37– 40, February 1970. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mischaikow:1993:TWM

[MH93] Konstantin Mischaikow and Vivian Hutson. Travelling waves for mutualist species. SIAM Journal on Mathematical Analysis, 24 (4):987–1008, July 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Morgan:1995:EPS

[MH95] J. J. Morgan and S. L. Hollis.

The existence of periodic solutions to reaction-diffusion systems with periodic data. SIAM Journal on Mathematical Analysis, 26(5):1225–1232, September 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/25717.

Mhaskar:1988:QDJ

[Mha88] H. N. Mhaskar. A quantitative Dirichlet–Jordan type theorem for orthogonal polynomial expansions. SIAM Journal on Mathematical Analysis, 19(2):484–492, March 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miamee:1987:DPN

[Mia87]  A. G. Miamee.  On deter-

mining the predictor of nonfullrank multivariate stationary random processes. SIAM Journal on Mathematical Analysis, 18(4): 909–918, July 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miamee:1993:EFB

[Mia93] A. G. Miamee. Explicit formula for the best linear predictor of periodically correlated sequences. SIAM Journal on Mathematical Analysis, 24(3): 703–711, May 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Michelson:1992:BFS

[Mic92]  Daniel Michelson.  Bunsen

flames as steady solutions of the Kuramoto–Sivashinsky equation. SIAM Journal on Mathematical Analysis, 23(2):364–386, March 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Michelson:1996:SBF

[Mic96] Daniel Michelson. Stability of the Bunsen flame profiles in the Kuramoto–Sivashinsky equation. SIAM Journal on Mathematical Analysis, 27(3):765–781, May 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mierczynski:1987:SCS

[Mie87] Janusz Mierczyński. Strictly cooperative systems with a first integral. SIAM Journal on Mathematical Analysis, 18(3):642–646, May 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miklavcic:1988:LSL

[Mik88] Milan Miklavčič. On limit states of a linearized Boltzmann equation. SIAM Journal on Mathematical Analysis, 19(1):150–152, January 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Millar:1970:LSTa

[Mil70a] R. F. Millar. The location of singularities of two-dimensional harmonic functions. I: Theory. SIAM Journal on Mathematical Analysis, 1(3):333–344, August 1970. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Millar:1970:LSTb

[Mil70b] R. F. Millar. The location of singularities of two-dimensional harmonic functions. II: Applications. SIAM Journal on Mathematical Analysis, 1(3):345–353, August 1970. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miller:1970:LSM

[Mil70c] Keith Miller. Least squares methods for ill-posed problems with a prescribed bound. SIAM Journal on Mathematical Analysis, 1(1):52–74, February 1970. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miller:1970:LTD

[Mil70d] Willard Miller, Jr. Lie theory and q-difference equations. SIAM Journal on Mathematical Analysis, 1(2):171–188, May 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miller:1970:LTSb

[Mil70e] Willard Miller, Jr. Lie theory and some special solutions of the hypergeometric equations. SIAM Journal on Mathematical Analysis, 1(4):405–425, November 1970. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miller:1970:LTSa

[Mil70f] Willard Miller, Jr. Lie theory and special functions satisfying second order nonhomogeneous differential equations. SIAM Journal on Mathematical Analysis, 1 (2):246–265, May 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miller:1971:AAP

[Mil71a] R. K. Miller. Asymptotically almost periodic solutions of a nonlinear Volterra system. SIAM Journal on Mathematical Analysis, 2(3):435–444, August 1971. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miller:1971:LAD

[Mil71b] Willard Miller, Jr. On Lie algebras of difference operators and the special functions of mathematical physics. SIAM Journal on Mathematical Analysis, 2 (3):307–327, August 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miller:1972:LTG

[Mil72] Willard Miller, Jr. Lie theory and generalized hypergeometric functions. SIAM Journal on Mathematical Analysis, 3(1):31– 44, February 1972. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miller:1973:LTA

[Mil73a] Willard Miller, Jr. Lie theory and the Appell functions F_{1}. SIAM Journal on Mathematical Analysis, 4(4):638–655, November 1973. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miller:1973:SDE

[Mil73b] Willard Miller, Jr. Symmetries of differential equations. the hypergeometric and Euler– Darboux equations. SIAM Journal on Mathematical Analysis, 4 (2):314–328, May 1973. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miller:1974:LTM

[Mil74a] Willard Miller, Jr. Lie theory and Meijers’s G-function. SIAM Journal on Mathematical Analysis, 5(2):309–318, April 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miller:1974:LTSa

[Mil74b] Willard Miller, Jr. Lie theory and separation of variables. I: Parabolic cylinder coordinates. SIAM Journal on Mathematical Analysis, 5(4):626–643, August 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miller:1974:LTSb

[Mil74c] Willard Miller, Jr. Lie theory and separation of variables. II: Parabolic coordinates. SIAM Journal on Mathematical Analysis, 5(5):822–836, October 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miller:1975:STS

[Mil75] William Brunner Miller. Separation theorems for self-adjoint linear differential equations of the fourth order. SIAM Journal on Mathematical Analysis, 6 (4):742–759, August 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Millar:1976:SSE

[Mil76]  R. F. Millar.  Singularities

of solutions to exterior analytic boundary value problems for the Helmholtz equation in three independent variables. I: The plane boundary. SIAM Journal on Mathematical Analysis, 7(1):131–156, February 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Millar:1979:SSE

[Mil79] R. F. Millar. Singularities of solutions to exterior analytic boundary value problems for the Helmholtz equation in three independent variables. II. The axisymmetric boundary. SIAM Journal on Mathematical Analysis, 10(4):682–694, July 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Millar:1984:ACP

[Mil84] Robert F. Millar. The analytic Cauchy problem for fourth order elliptic equations in two independent variables. SIAM Journal on Mathematical Analysis, 15(5): 964–978, September 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miller:1987:NWP

[Mil87] Willard Miller, Jr. A note on Wilson polynomials. SIAM Journal on Mathematical Analysis, 18 (5):1221–1226, September 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Milne:1994:RSI

[Mil94] Stephen C. Milne. The C_{ℓ} Rogers–Selberg identity. SIAM Journal on Mathematical Analysis, 25(2):571–595, March 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/23748.

Mimachi:1988:PRI

[Mim88] Katsuhisa Mimachi. A proof of Ramanujan’s identity by use of loop integrals. SIAM Journal on Mathematical Analysis, 19(6): 1490–1493, November 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mingarelli:1986:ECP

[Min86] Angelo B. Mingarelli. On the existence of conjugate points for a second order ordinary differential equation. SIAM Journal on Mathematical Analysis, 17(1):1– 6, January 1986. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mirsky:1983:IMB

[Mir83] L. Mirsky. An inequality of the Markov–Bernstein type for polynomials. SIAM Journal on Mathematical Analysis, 14(5):1004– 1008, September 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mirollo:1994:SPO

[Mir94] Renato E. Mirollo. Splay-phase orbits for equivariant flows on tori. SIAM Journal on Mathematical Analysis, 25(4):1176– 1180, July 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/21706.

Mischaikow:1995:GAD

[Mis95] Konstantin Mischaikow. Global asymptotic dynamics of gradientlike bistable equations. SIAM Journal on Mathematical Analysis, 26(5):1199–1224, September 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/25082.

Mitchell:1975:SAD

[Mit75] George E. Mitchell. Self-adjoint differential expressions in two variables. SIAM Journal on Mathematical Analysis, 6(6): 1017–1020, November 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mizony:1983:TLJ

[Miz83] Michel Mizony. Une transformation de Laplace–Jacobi. (French) [A Laplace–Jacobi transformation]. SIAM Journal on Mathematical Analysis, 14(5):987– 1003, September 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Morton:1978:DWF

[MK78] Robert D. Morton and Allan M. Krall. Distributional weight functions for orthogonal polynomials. SIAM Journal on Mathematical Analysis, 9(4):604–626, August 1978. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mingarelli:1983:LTP

[MK83] Angelo B. Mingarelli and Allan M. Krall. Legendre type polynomials under an indefinite inner product. SIAM Journal on Mathematical Analysis, 14 (2):399–402, March 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

McKenna:1984:SPR

[ML84] P. J. McKenna and I. Lasiecka. Some perturbation results and their applications to stabilization of hyperbolic systems. SIAM Journal on Mathematical Analysis, 15(6):1140–1152, November 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

McComb:1993:SEC

[ML93] I-Heng McComb and Chjan C. Limt. Stability of equilibria for a class of time-reversible, D_{n} × O(2)-symmetric homogeneous vector fields. SIAM Journal on Mathematical Analysis, 24 (4):1009–1029, July 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

McLaughlin:1997:EUR

[MLM97a] D. J. McLaughlin, W. Lamb, and A. C. McBride. An existence and uniqueness result for a coagulation and multiple-fragmentation equation. SIAM Journal on Mathematical Analysis, 28(5): 1173–1190, September 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29171.

McLaughlin:1997:SAF

[MLM97b] D. J. McLaughlin, W. Lamb, and A. C. McBride. A semigroup approach to fragmentation models. SIAM Journal on Mathematical Analysis, 28(5): 1158–1172, September 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29170.

Medeiros:1977:EUP

[MM77a] L. A. Medeiros and G. Perla Menzala. Existence and uniqueness for periodic solutions of the Benjamin–Bona–Mahony equation. SIAM Journal on Mathematical Analysis, 8(5):792–799, October 1977. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miller:1977:SIL

[MM77b] R. K. Miller and A. N. Michel.

L_{2}-stability and -instability of large-scale systems described by integrodifferential equations. SIAM Journal on Mathematical Analysis, 8(3):547–557, May 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miller:1979:SLV

[MM79] R. K. Miller and A. N. Michel. Stability of linear Volterra integro-differential equations of order n. SIAM Journal on Mathematical Analysis, 10(5):1089– 1091, September 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Manselli:1980:DRM

[MM80] Paolo Manselli and Keith Miller. Dimensionality reduction methods for efficient numerical solution, backward in time, of parabolic equations with variable coefficients. SIAM Journal on Mathematical Analysis, 11 (1):147–159, January 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miersemann:1992:SOP

[MM92] E. Miersemann and H. D. Mittelmann. Stability in obstacle problems for the von Karman plate. SIAM Journal on Mathematical Analysis, 23(5):1099– 1116, September 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Majorana:1997:SHS

[MM97] A. Majorana and S. A. Marano. Space homogeneous solutions to the Cauchy problem for semiconductor Boltzmann equations. SIAM Journal on Mathematical Analysis, 28(6):1294–1308, November 1997. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

http://epubs.siam.org/sam-
bin/dbq/article/29139.

Makagon:1997:LTA

[MMM97] A. Makagon, C. A. Mensah, and A. G. Miamee. A Levinson-type algorithm for discrete stationary random fields. SIAM Journal on Mathematical Analysis, 28 (4):897–902, July 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/28938.

Malek-Madani:1985:FSC

[MMN85] Reza Malek-Madani and John A.

Nohel. Formation of singularities for a conservation law with memory. SIAM Journal on Mathematical Analysis, 16(3):530–540, May 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miller:1975:SMT

[MN75] R. K. Miller and J. A. Nohel. A stable manifold theorem for a system of Volterra integrodifferential equations. SIAM Journal on Mathematical Analysis, 6(3):506–522, May 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Madych:1986:RRR

[MN86] W. R. Madych and S. A. Nelson. Reconstruction from restricted Radon transform data: resolution and ill-conditionedness. SIAM Journal on Mathematical Analysis, 17(6):1447–1453, November 1986. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

MaTe:1986:AGZ

[MNT86] Attila MáTé, Paul Nevai, and Vilmos Totik. Asymptotics for the greatest zeros of orthogonal polynomials. SIAM Journal on Mathematical Analysis, 17(3): 745–751, May 1986. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Morita:1994:NPB

[MNY94] Yoshihisa Morita, Hirokazu Ninomiya, and Eiji Yanagida. Nonlinear perturbation of boundary values for reaction-diffusion systems: Inertial manifolds and their applications. SIAM Journal on Mathematical Analysis, 25(5): 1320–1356, September 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/23147.

McLeod:1983:CBP

[MO83] J. B. McLeod and P. J. Olver. The connection between partial differential equations soluble by inverse scattering and ordinary differential equations of Painlevé type. SIAM Journal on Mathematical Analysis, 14(3):488–506, May 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mock:1974:IVP

[Moc74] M. S. Mock. An initial value problem from semiconductor device theory. SIAM Journal on Mathematical Analysis, 5(4): 597–612, August 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Moet:1979:NAB

[Moe79] H. J. K. Moet. A note on the asymptotic behavior of solutions of the KPP equation. SIAM Journal on Mathematical Analysis, 10(4):728–732, July 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Moeckel:1984:HPI

[Moe84] R. Moeckel. Heteroclinic phenomena in the isosceles threebody problem. SIAM Journal on Mathematical Analysis, 15(5): 857–876, September 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Monegato:1980:SNI

[Mon80]  Giovanni Monegato.  Some

new inequalities related to certain ultra-spherical polynomials. SIAM Journal on Mathematical Analysis, 11(4):663–667, July 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Moore:1973:IPI

[Moo73] Michael H. Moore. An inner product inequality. SIAM Journal on Mathematical Analysis, 4 (3):514–518, May 1973. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

MoralMedina:1985:LOR

[Mor85] José M. Moral Medina. Linear operators related to timeinvariant discrete filters. SIAM Journal on Mathematical Analysis, 16(4):703–712, July 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Morita:1987:PWS

[Mor87] Yoshihisa Morita. A periodic wave and its stability to a circular chain of weakly coupled oscillators. SIAM Journal on Mathematical Analysis, 18(6):1681– 1698, November 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Morgan:1989:GES

[Mor89] Jeff Morgan. Global existence for semilinear parabolic systems. SIAM Journal on Mathematical Analysis, 20(5):1128–1144, September 1989. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Morgan:1990:BDR

[Mor90]  Jeff Morgan.  Boundedness

and decay results for reactiondiffusion systems. SIAM Journal on Mathematical Analysis, 21(5):

1172–1189, September 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Moseley:1983:ASD

[Mos83a] J. L. Moseley. Asymptotic solutions for a Dirichlet problem with an exponential nonlinearity. SIAM Journal on Mathematical Analysis, 14(4):719–735, July 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Moseley:1983:TDD

[Mos83b] J. L. Moseley. A two-dimensional Dirichlet problem with an exponential nonlinearity. SIAM Journal on Mathematical Analysis, 14(5):934–946, September 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Micchelli:1977:MTW

[MP77] C. A. Micchelli and Allan Pinkus.

Moment theory for weak Chebyshev systems with applications to monosplines, quadrature formulae and best one-sided L^{1}- approximation by spline functions with fixed knots. SIAM Journal on Mathematical Analysis, 8(2):206–230, April 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Meyer:1983:CWM

[MP83] R. E. Meyer and J. F. Painter.

Connection for wave modulation. SIAM Journal on Mathematical Analysis, 14(3):450–462, May 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mischler:1997:BEI

[MP97] S. Mischler and B. Perthame. Boltzmann equation with infinite energy: Renormalized solutions and distributional solutions for small initial data and initial data close to a Maxwellian. SIAM Journal on Mathematical Analysis, 28(5): 1015–1027, September 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29810.

Monasse:1998:OWB

[MP98] Pascal Monasse and Valérie Perrier. Orthonormal wavelet bases adapted for partial differential equations with boundary conditions. SIAM Journal on Mathematical Analysis, 29(4): 1040–1065, July 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29512.

Mallet-Paret:1989:DDE

[MPN89] John Mallet-Paret and Roger D. Nussbaum. A differential-delay equation arising in optics and physiology. SIAM Journal on Mathematical Analysis, 20(2): 249–292, March 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Markoe:1985:EPL

[MQ85] Andrew Markoe and Eric Todd Quinto. An elementary proof of local invertibility for generalized and attenuated Radon transforms. SIAM Journal on Mathematical Analysis, 16(5):1114– 1119, September 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mahfoud:1979:SPS

[MR79a] W. E. Mahfoud and S. M.

Rankin. Some properties of solutions of (r(t)ψ(x)x′)′ + a(t)f(x) = 0. SIAM Journal on Mathematical Analysis, 10 (1):49–54, January 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mullis:1979:WEL

[MR79b] C. T. Mullis and R. A. Roberts.

On weak equivalence of linear systems and finite state systems. SIAM Journal on Mathematical Analysis, 10(3):498–511, May 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Murdock:1980:NAE

[MR80] James Murdock and Clark Robinson. A note on the asymptotic expansion of eigenvalues. SIAM Journal on Mathematical Analysis, 11(3):458–459, May 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Markowich:1983:NVI

[MR83a] P. Markowich and M. Renardy. A nonlinear Volterra integrodifferential equation describing the stretching of polymeric liquids. SIAM Journal on Mathematical Analysis, 14(1):66–97, January 1983. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Markowich:1983:SPP

[MR83b] Peter A. Markowich and C. A. Ringhofer. Singular perturbation problems with a singularity of the second kind. SIAM Journal on Mathematical Analysis, 14(5): 897–914, September 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Markowich:1983:SSP

[MR83c] Peter A. Markowich and C. A. Ringhofer. Systems of singular perturbation problems with a first order turning point. SIAM Journal on Mathematical Analysis, 14(5):875–896, September 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Masson:1991:STJ

[MR91] D. R. Masson and J. Repka. Spectral theory of Jacobi matrices in l^{2}(Z) and the su(1, 1) Lie algebra. SIAM Journal on Mathematical Analysis, 22(4):1131– 1146, July 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mischaikow:1993:TWP

[MR93] Konstantin Mischaikow and James F. Reineck. Travelling waves in predator-prey systems. SIAM Journal on Mathematical Analysis, 24(5):1179–1214, September 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Misiewicz:1994:PIB

[MR94] Jolanta K. Misiewicz and Donald St. P. Richards. Positivity of integrals of Bessel functions. SIAM Journal on Mathematical Analysis, 25(2): 596–601, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/22693.

Mallordy:1995:PEK

[MR95] Jean-François Mallordy and Jean-Michel Roquejoffre. A parabolic equation of the KPP type in higher dimensions. SIAM Journal on Mathematical Analysis, 26(1):1–20, January 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/24610.

Mikelic:1998:EDR

[MR98a] Andro Mikelic and Raoul Robert.

On the equations describing a relaxation toward a statistical equilibrium state in the twodimensional perfect fluid dynamics. SIAM Journal on Mathematical Analysis, 29(5):1238– 1255, September 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/30650.

Mikulevicius:1998:LPS

[MR98b] R. Mikulevicius and B. Rozovskii. Linear parabolic stochastic PDEs and Wiener chaos. SIAM Journal on Mathematical Analysis, 29(2):452–480, March 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/29906.

McEliece:1981:TIA

[MRS81] R. J. McEliece, B. Reznick, and J. B. Shearer. A Turán inequality arising in information theory. SIAM Journal on Mathematical Analysis, 12(6):931–934, November 1981. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Marsden:1987:CSI

[MrS87] Jerrold Marsden and Jürgen rgen Scheurle. The construction and smoothness of invariant manifolds by the deformation method. SIAM Journal on Mathematical Analysis, 18(5):1261–1274, September 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Miamee:1979:BPE

[MS79] A. G. Miamee and H. Salehi. On the bilateral prediction error matrix of a multivariate stationary stochastic process. SIAM Journal on Mathematical Analysis, 10 (2):247–252, March 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Muldoon:1984:SRZ

[MS84] M. E. Muldoon and R. Spigler.

Some remarks on zeros of cylinder functions. SIAM Journal on Mathematical Analysis, 15(6): 1231–1233, November 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

McTaggart:1986:CSR

[MS86] Carol L. McTaggart and Brian Straughan. Chemical surface reactions and nonlinear stability by the method of energy. SIAM Journal on Mathematical Analysis, 17(2):342–351, March 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Marquez:1987:EPO

[MS87] Viviana Marquez and Meir Shillor. The electropainting problem with overpotentials. SIAM Journal on Mathematical Analysis, 18(3):788–811, May 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Meyer:1988:BRE

[MS88] Kenneth R. Meyer and Dieter S.

Schmidt. Bifurcations of relative equilibria in the N-body and Kirchhoff problems. SIAM Journal on Mathematical Analysis, 19 (6):1295–1313, November 1988.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Menyuk:1992:TSR

[MS92] Curtis R. Menyuk and Thomas I. Seidman. Transient stimulated Raman scattering. SIAM Journal on Mathematical Analysis, 23 (2):346–363, March 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Moeckel:1995:BSC

[MS95] Richard Moeckel and Carles Simó. Bifurcation of spatial central configurations from planar ones. SIAM Journal on Mathematical Analysis, 26(4):978–998, July 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/24841.

Melenk:1999:ARS

[MS99] J. M. Melenk and C. Schwab. Analytic regularity for a singularly perturbed problem. SIAM Journal on Mathematical Analysis, 30(2):379–400, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/31754.

Muller:1996:ITP

[MST96] Stefan Müller, Scott J. Spector, and Qi Tang. Invertibility and a topological property of Sobolev maps. SIAM Journal on Mathematical Analysis, 27 (4):959–976, July 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/26376.

Muschietti:1995:PAL

[MT95] M. A. Muschietti and B. Torrésani.

Pyramidal algorithms for Littlewood– Paley decompositions. SIAM Journal on Mathematical Analysis, 26(4):925–943, July 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/25222.

Mimura:1980:MST

[MTH80] Masayasu Mimura, Masahisa Tabata, and Yuzo Hosono. Multiple solutions of two-point boundary value problems of Neumann type with a small parameter. SIAM Journal on Mathematical Analysis, 11(4):613–631, July 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Muckenhoupt:1970:EAE

[Muc70] Benjamin Muckenhoupt. Equiconvergence and almost everywhere convergence of Hermite and Laguerre series. SIAM Journal on Mathematical Analysis, 1(3): 295–321, August 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mugler:1975:CCP

[Mug75] Dale H. Mugler. Completely convex and positive harmonic functions. SIAM Journal on Mathematical Analysis, 6(4):681–688, August 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mugler:1979:CCF

[Mug79] Dale H. Mugler. Completely convex functions and convergence. SIAM Journal on Mathematical Analysis, 10(2):292–296, March 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Muki:1972:UAE

[Muk72] R. Muki. Uniform asymptotic expansions of integrals of the Lipschitz–Hankel type. SIAM Journal on Mathematical Analysis, 3(2):285–290, May 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Muller:1975:SRF

[Mul75]  George M. Muller.  A set

of rational functions related to the Euler–Fobenius polynomials. SIAM Journal on Mathematical Analysis, 6(6):948–959, November 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Muldowney:1978:CTL

[Mul78a] James S. Muldowney. Comparison theorems for linear boundary value problems. SIAM Journal on Mathematical Analysis, 9 (5):943–955, October 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Muldowney:1978:LDI

[Mul78b] James S. Muldowney. Linear differential inequalities. SIAM Journal on Mathematical Analysis, 9(1):106–120, February 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mullin:1978:EAK

[Mul78c] Melvin Mullin. The energy approach to the karman–Föppl equations. SIAM Journal on Mathematical Analysis, 9(1): 151–156, February 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Muldowney:1981:ILO

[Mul81] James S. Muldowney. On invertibility of linear ordinary differential boundary value problems. SIAM Journal on Mathematical Analysis, 12(3):368–384, May 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Muldowney:1987:CPM

[Mul87] James S. Muldowney. The converse of Pólya’s mean value theorem. SIAM Journal on Mathematical Analysis, 18(5):1317– 1322, September 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Murdock:1983:SAE

[Mur83] James A. Murdock. Some asymptotic estimates for higher order averaging and a comparison with iterated averaging. SIAM Journal on Mathematical Analysis, 14 (3):421–424, May 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Muldowney:1974:EPE

[MW74] J. S. Muldowney and D. Willett. An elementary proof of the existence of solutions to second order nonlinear boundary value problems. SIAM Journal on Mathematical Analysis, 5 (4):701–707, August 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Macki:1975:BVP

[MW75] Jack W. Macki and Paul Waltman. A boundary value problem for a two-dimensional system with a parameter. SIAM Journal on Mathematical Analysis, 6 (1):91–95, February 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Marcus:1978:ASF

[MW78] Steven I. Marcus and Alan S. Willsky. Algebraic structure and finite dimensional nonlinear estimation. SIAM Journal on Mathematical Analysis, 9(2):312–327, April 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Markowich:1983:NEP

[MW83] Peter A. Markowich and Richard Weiss. Nonlinear eigenvalue problems on infinite intervals. SIAM Journal on Mathematical Analysis, 14(3):431–449, May 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

McClure:1987:AEM

[MW87] J. P. McClure and R. Wong.

Asymptotic expansion of a multiple integral. SIAM Journal on Mathematical Analysis, 18(6): 1630–1637, November 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

McClure:1991:TDS

[MW91] J. P. McClure and R. Wong.

Two-dimensional stationary phase approximation: Stationary point at a corner. SIAM Journal on Mathematical Analysis, 22 (2):500–523, March 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Marz:1993:SBV

[MW93] Roswitha März and Ewa B.

Weinmüller. Solvability of boundary value problems for systems of singular differentialalgebraic equations. SIAM Journal on Mathematical Analysis, 24 (1):200–215, January 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Mhaskar:1991:MCE

[MX91] H. N. Mhaskar and Y. Xu.

Mean convergence of expansions in Freud-type orthogonal polynomials. SIAM Journal on Mathematical Analysis, 22(3):847–855, May 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Murray:1998:ETW

[MX98] Regan Murray and J. X. Xin. Existence of traveling waves in a biodegradation model for organic contaminants. SIAM Journal on Mathematical Analysis, 30(1):72–94, 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/31339.

Mizoguchi:1998:BLS

[MY98] Noriko Mizoguchi and Eiji Yanagida. Blowup and life span of solutions for a semilinear parabolic equation. SIAM Journal on Mathematical Analysis, 29(6):1434–1446, November 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/32493.

Montaldi:1992:SFR

[MZ92] Emilio Montaldi and Giuseppe Zucchelli. Some formulas of Ramanujan, revisited. SIAM Journal on Mathematical Analysis, 23 (2):562–569, March 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Micu:1998:ASF

[MZ98] Sorin Micu and Enrique Zuazua. Asymptotics for the spectrum of a fluid/structure hybrid system arising in the control of noise. SIAM Journal on Mathematical Analysis, 29(4):967–1001, July 1998. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/31234.

Nagle:1978:NBV

[Nag78] R. Kent Nagle. Nonlinear boundary value problems for ordinary differential equations with a small parameter. SIAM Journal on Mathematical Analysis, 9 (4):719–729, August 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nakajima:1978:SCD

[Nak78] F. Nakajima. A stability criterion of diagonal dominance type. SIAM Journal on Mathematical Analysis, 9(5):815–824, October 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nakane:1988:FSS

[Nak88] Shizuo Nakane. Formation of shocks for a single conservation law. SIAM Journal on Mathematical Analysis, 19(6):1391– 1408, November 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nakamura:1991:GRF

[Nak91] Yoshimasa Nakamura. Geometry of rational functions and nonlinear integrable systems. SIAM Journal on Mathematical Analysis, 22(6):1744–1754, September 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nambu:1978:ABS

[Nam78] Takao Nambu. Asymptotic behavior of solutions of a class of nonlinear differential equations in Banach space. SIAM Journal on Mathematical Analysis, 9 (4):687–718, August 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Narita:1993:ABS

[Nar93] K. Narita. Asymptotic behavior of solutions of SDE for relaxation oscillations. SIAM Journal on Mathematical Analysis, 24 (1):172–199, January 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Naasell:1978:RBR

[Nås78] Ingemar Nåsell. Rational bounds for ratios of modified Bessel functions. SIAM Journal on Mathematical Analysis, 9(1):1– 11, February 1978. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nasr:1984:SOD

[Nas84] A. H. Nasr. A study of onedimensional Schrödinger equation with nonlocal potential. SIAM Journal on Mathematical Analysis, 15(3):459–467, May 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Naundorf:1976:CPS

[Nau76] Friedrich Naundorf. A connection problem for second order linear differential equations with two irregular singular points. SIAM Journal on Mathematical Analysis, 7(2):157–175, April 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Naumann:1979:CFO

[Nau79]  J. Naumann.  On a class

of first-order evolution inequalities arising in heat conduction with memory. SIAM Journal on Mathematical Analysis, 10(6): 1144–1160, November 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Navarro:1999:SDW

[Nav99] Jaime Navarro. Singularities of distributions via the wavelet transform. SIAM Journal on Mathematical Analysis, 30(2): 454–467, 1999. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

http://epubs.siam.org/sam-
bin/dbq/article/29249.

Naylor:1977:ITO

[Nay77] D. Naylor. On an integral transform occurring in the theory of diffraction. SIAM Journal on Mathematical Analysis, 8(3): 402–411, May 1977. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Naylor:1978:NAE

[Nay78] D. Naylor. On a nonself adjoint eigenfunction expansion. SIAM Journal on Mathematical Analysis, 9(5):967–978, October 1978.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Naylor:1984:SAF

[Nay84] D. Naylor. On simplified asymptotic formulas for a class of Mathieu functions. SIAM Journal on Mathematical Analysis, 15(6): 1205–1213, November 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Naylor:1987:SAF

[Nay87]  D. Naylor.  On a simpli-

fied asymptotic formula for the Mathieu function of the third kind. SIAM Journal on Mathematical Analysis, 18(6):1616– 1629, November 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Naylor:1989:ITI

[Nay89] D. Naylor. On an integral transform involving a class of Mathieu functions. SIAM Journal on Mathematical Analysis, 20(6): 1500–1513, November 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Naylor:1975:EPO

[NB75] D. Naylor and D. W. Barclay. On an expansion problem occurring in the theory of diffraction. SIAM Journal on Mathematical Analysis, 6(1):199–207, February 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Naylor:1982:FI

[NC82] D. Naylor and P. H. Chang. On a formula of inversion. SIAM Journal on Mathematical Analysis, 13 (6):1053–1071, November 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nevai:1979:AAP

[ND79] Paul G. Nevai and Jesús S. Dehesa. On asymptotic average properties of zeros of orthogonal polynomials. SIAM Journal on Mathematical Analysis, 10(6): 1184–1192, November 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Norris:1981:UTO

[ND81] M. J. Norris and R. D. Driver. A uniqueness theorem for ordinary differential equations. SIAM Journal on Mathematical Analysis, 12(2):141–144, March 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nevai:1994:GJW

[NEM94] Paul Nevai, Tamás Erdélyi, and Alphonse P. Magnus. Generalized Jacobi weights, Christoffel functions, and Jacobi polynomials. SIAM Journal on Mathematical Analysis, 25(2):602–614, March 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/23686. See erratum

[EMN94].

Neuman:1988:CSF

[Neu88] Edward Neuman. On complete symmetric functions. SIAM Journal on Mathematical Analysis, 19(3):736–750, May 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nevai:1979:AFD

[Nev79] Paul G. Nevai. An asymptotic formula for the derivatives of orthogonal polynomials. SIAM Journal on Mathematical Analysis, 10(3):472–477, May 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nevai:1984:AOP

[Nev84] Paul Nevai. Asymptotics for orthogonal polynomials associated with exp(−x^{4}). SIAM Journal on Mathematical Analysis, 15(6): 1177–1187, November 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nishiura:1987:SSP

[NF87] Yasumasa Nishiura and Hiroshi Fujii. Stability of singularly perturbed solutions to systems of reaction-diffusion equations. SIAM Journal on Mathematical Analysis, 18(6):1726–1770, November 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nguetseng:1989:GCR

[Ngu89] Gabriel Nguetseng. A general convergence result for a functional related to the theory of homogenization. SIAM Journal on Mathematical Analysis, 20(3): 608–623, May 1989. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nguetseng:1990:AAS

[Ngu90] Gabriel Nguetseng. Asymptotic analysis for a stiff variational problem arising in mechanics. SIAM Journal on Mathematical Analysis, 21(6):1394–1414, November 1990. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Niijima:1978:BSS

[Nii78] Koichi Niijima. On the behavior of solutions of a singularly perturbed boundary value problem with a turning point. SIAM Journal on Mathematical Analysis, 9(2):298–311, April 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nii:1997:ESI

[Nii97a] Shunsaku Nii. An extension of the stability index for travelingwave solutions and its application to bifurcations. SIAM Journal on Mathematical Analysis, 28 (2):402–433, March 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/27878.

Nii:1997:STM

[Nii97b] Shunsaku Nii. Stability of travelling multiple-front (multipleback) wave solutions of the FitzHugh–Nagumo equations. SIAM Journal on Mathematical Analysis, 28(5):1094–1112, September 1997. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

                                        http://epubs.siam.org/sam-
                                        bin/dbq/article/28829.

Nipp:1986:BSSa

[Nip86a] K. Nipp. Breakdown of stability in singularly perturbed autonomous systems. I. Orbit equations. SIAM Journal on Mathematical Analysis, 17(3):512–532, May 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nipp:1986:BSSb

[Nip86b] K. Nipp. Breakdown of stability in singularly perturbed autonomous systems. II. Estimates for the solutions and application. SIAM Journal on Mathematical Analysis, 17(5):1068– 1085, September 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nishiura:1982:GSB

[Nis82] Y. Nishiura. Global structure of bifurcating solutions of some reaction-diffusion systems. SIAM Journal on Mathematical Analysis, 13(4):555–593, July 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nishiura:1990:SLA

[NMIF90] Y. Nishiura, M. Mimura, H. Ikeda, and H. Fujii. Singular limit analysis of stability of traveling wave solutions in bistable reactiondiffusion systems. SIAM Journal on Mathematical Analysis, 21 (1):85–122, January 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nitsche:1988:EUR

[NNB88] Ludwig C. Nitsche, Johannes M.

Nitsche, and Howard Brenner. Existence, uniqueness and regularity of a time-periodic probability density distribution arising in a sedimentation-diffusion problem. SIAM Journal on Mathematical Analysis, 19(1):153–166, January 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nohel:1974:AQE

[Noh74] J. A. Nohel. Addendum: Quantitative estimates for a nonlinear system of integrodifferential equations arising in reactor dynamics. SIAM Journal on Mathematical Analysis, 5(2): 349, April 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Noren:1988:SLP

[Nor88a] Richard Noren. A singular limit problem for a Volterra equation. SIAM Journal on Mathematical Analysis, 19(5):1103– 1107, September 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Noren:1988:UBS

[Nor88b] Richard Noren. Uniform L^{1} behavior for the solution of a Volterra equation with a parameter. SIAM Journal on Mathematical Analysis, 19(2):270–286, March 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Novak:1988:CII

[Nov88] Erich Novak. On the convergence of interpolated iteration methods. SIAM Journal on Mathematical Analysis, 19(5):1174– 1182, September 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nohel:1993:NSA

[NP93] John A. Nohel and Robert L. Pego. Nonlinear stability and asymptotic behavior of shearing motions of a non-Newtonian fluid. SIAM Journal on Mathematical Analysis, 24(4):911–942, July 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nouri:1995:SSB

[NP95] A. Nouri and F. Poupaud. Stationary solutions of boundary value problems for a Maxwell– Boltzmann system modelling degenerate semiconductors. SIAM Journal on Mathematical Analysis, 26(5):1143–1156, September 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/24560.

Nassrallah:1985:PFR

[NR85] B. Nassrallah and Mizan Rahman. Projection formulas, a reproducing kernel and a generating function for q-Wilson polynomials. SIAM Journal on Mathematical Analysis, 16(1):186–197, January 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nouri:1995:GEU

[NR95] A. Nouri and M. Rascle. A global existence and uniqueness theorem for a model problem in dynamic elasto-plasticity with isotropic strain-hardening. SIAM Journal on Mathematical Analysis, 26(4):850–868, July 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/19960.

Ni:1985:NPP

[NS85] Wei Ming Ni and Paul Sacks.

The number of peaks of positive solutions of semilinear parabolic equations. SIAM Journal on Mathematical Analysis, 16(3): 460–471, May 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Noussair:1989:DSS

[NS89] Ezzat S. Noussair and Charles A.

Swanson. Decaying solutions of semilinear elliptic equations in R^{N}. SIAM Journal on Mathematical Analysis, 20(6):1336– 1343, November 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nedelec:1991:IEM

[NS91] J.-C. Nédélec and F. Starling. Integral equation methods in a quasi-periodic diffraction problem for the time-harmonic Maxwell’s equations. SIAM Journal on Mathematical Analysis, 22(6):1679–1701, September 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nishiura:1998:NHD

[NS98] Yasumasa Nishiura and Hiromasa Suzuki. Nonexistence of higher dimensional stable Turing patterns in the singular limit. SIAM Journal on Mathematical Analysis, 29(5):1087– 1105, September 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/31323.

Nurnberger:1984:GCS

[NSSS84] G. Nürnberger, L. L. Schumaker, M. Sommer, and H. Strauss. Generalized Chebyshevian splines. SIAM Journal on Mathematical Analysis, 15(4):790–804, July 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nagata:1986:BDDa

[NT86a] Wayne Nagata and James W. Thomas. Bifurcation in doublydiffusive systems. I. Equilibrium solutions. SIAM Journal on Mathematical Analysis, 17(1): 91–113, January 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nagata:1986:BDDb

[NT86b] Wayne Nagata and James W.

Thomas. Bifurcation in doublydiffusive systems. II. Time periodic solutions. SIAM Journal on Mathematical Analysis, 17 (1):114–127, January 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nagata:1986:BDDc

[NT86c] Wayne Nagata and James W.

Thomas. Bifurcations in doublydiffusive systems. III. Interaction of equilibrium and time periodic solutions. SIAM Journal on Mathematical Analysis, 17 (2):289–311, March 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nishiyama:1996:IIB

[NT96] Takahiro Nishiyama and Atusi Tani. Initial and initial-boundary value problems for a vortex filament with or without axial flow. SIAM Journal on Mathematical Analysis, 27(4): 1015–1023, July 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/27379.

Nakamura:1995:IPB

[NU95] Gen Nakamura and Gunther Uhlmann. Inverse problems at the boundary for an elastic medium. SIAM Journal on Mathematical Analysis, 26 (2):263–279, March 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24749.

Nurnberger:1980:NCS

[Nür80] Günther Nürnberger. Nonexistence of continuous selections of the metric projection and weak Chebyshev systems. SIAM Journal on Mathematical Analysis, 11 (3):460–467, May 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nussbaum:1978:PTT

[Nus78] Roger Nussbaum. A periodicity threshold theorem for some nonlinear integral equations. SIAM Journal on Mathematical Analysis, 9(2):356–376, April 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nussbaum:1990:SNW

[Nus90] Roger D. Nussbaum. Some nonlinear weak ergodic theorems. SIAM Journal on Mathematical Analysis, 21(2):436–460, March 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nochetto:1995:CDO

[NV95] Ricardo H. Nochetto and Claudio Verdi. Convergence of double obstacle problems to the generalized geometric motion of fronts. SIAM Journal on Mathematical Analysis, 26(6):1514– 1526, November 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/25542.

Nashed:1974:GIR

[NW74] M. Z. Nashed and Grace Wahba.

Generalized inverses in reproducing kernel spaces: An approach to regularization of linear operator equations. SIAM Journal on Mathematical Analysis, 5(6): 974–987, November 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Nishibata:1997:ABH

[NY97] Shinya Nishibata and Shih-Hsien Yu. The asymptotic behavior of the hyperbolic conservation laws with relaxation on the quarter-plane. SIAM Journal on Mathematical Analysis, 28 (2):304–321, March 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/27650.

Obi:1978:NRP

[Obi78] E. C. Obi. A note on the Rayleigh polynomials. SIAM Journal on Mathematical Analysis, 9(5):825–834, October 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See corrigendum [Obi81].

Obi:1981:CNR

[Obi81] E. C. Obi. Corrigendum: “A note on the Rayleigh polynomials” [SIAM J. Math. Anal. 9 (1978), no. 5, 825–834; MR 80j:33010]. SIAM Journal on Mathematical Analysis, 12(1): 139, January 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [Obi78].

Obi:1986:SSR

[Obi86] E. C. Obi. The summability significance of roots of Bessel functions. SIAM Journal on Mathematical Analysis, 17(6):1489– 1494, November 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See corrigenda [Obi87].

Obi:1987:CSS

[Obi87] E. C. Obi. Corrigenda: “The summability significance of roots of Bessel functions”. SIAM Journal on Mathematical Analysis, 18(4):1219, July 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [Obi86].

ODonnell:1984:BCL

[O’D84] Mark A. O’Donnell. Boundary and corner layer behavior in singularly perturbed semilinear systems of boundary value problems. SIAM Journal on Mathematical Analysis, 15(2):317–332, March 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ohtsuki:1977:ETS

[OFK77] T. Ohtsuki, T. Fujisawa, and S. Kumagai. Existence theorems and a solution algorithm for piecewise-linear resistor networks. SIAM Journal on Mathematical Analysis, 8(1):69–99, February 1977. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ohkohchi:1991:UST

[Ohk91] Shigemi Ohkohchi. Uniform simplification at a transition point and the problem of resonance. SIAM Journal on Mathematical Analysis, 22(1):213–237, January 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ohta:1979:NAM

[Oht79] Yuzo Ohta. Nonlinear accretive mappings in Banach spaces: the solvability and a solution algorithm. SIAM Journal on Mathematical Analysis, 10(2):337–353, March 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Oja:1978:OPO

[Oja78] Erkki Oja. S-orthogonal projection operators as asymptotic solutions of a class of matrix differential equations. SIAM Journal on Mathematical Analysis, 9 (5):848–854, October 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

OMalley:1994:RND

[OK94] Robert E. O’Malley, Jr. and Leonid V. Kalachev. Regularization of nonlinear differentialalgebraic equations. SIAM Journal on Mathematical Analysis, 25 (2):615–629, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/22640.

Okrasinski:1991:NSN

[Okr91] W. Okrasiński. Nontrivial solutions to nonlinear Volterra integral equations. SIAM Journal on Mathematical Analysis, 22 (4):1007–1015, July 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Oldham:1970:ECS

[Old70] Keith B. Oldham. On the evaluation of certain sums involving the natural numbers raised to an arbitrary power. SIAM Journal on Mathematical Analysis, 1(4): 538–546, November 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [Old72].

Oldham:1972:EEC

[Old72] Keith B. Oldham. Errata: On the evaluation of certain sums involving the natural numbers raised to an arbitrary power. SIAM Journal on Mathematical Analysis, 3(1):71, February 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [Old70].

Oldham:1983:RIS

[Old83] Keith B. Oldham. The reformulation of an infinite sum via semi-integration. SIAM Journal on Mathematical Analysis, 14(5): 974–981, September 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olmstead:1977:ELB

[Olm77] W. E. Olmstead. Extent of the left branching solution to certain bifurcation problems. SIAM Journal on Mathematical Analysis, 8(3):392–401, May 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olsder:1974:EAS

[Ols74] G. J. Olsder. On the existence of analytic solutions of systems of equations if the Jacobian is zero. SIAM Journal on Mathematical Analysis, 5(1):30– 42, February 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olver:1970:PA

[Olv70] F. W. J. Olver. A paradox in asymptotics. SIAM Journal on Mathematical Analysis, 1(4): 533–534, November 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olver:1974:EBS

[Olv74] F. W. J. Olver. Error bounds for stationary phase approximations. SIAM Journal on Mathematical Analysis, 5(1):19–29, February 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olver:1977:CFSb

[Olv77a] F. W. J. Olver. Connection formulas for second-order differential equations having an arbitrary number of turning points of arbitrary multiplicities. SIAM Journal on Mathematical Analysis, 8(4):673–700, August 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olver:1977:CFSa

[Olv77b] F. W. J. Olver. Connection formulas for second-order differential equations with multiple turning points. SIAM Journal on Mathematical Analysis, 8(1): 127–154, February 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olver:1978:SCA

[Olv78] F. W. J. Olver. Sufficient conditions for ackerberg–O’malley resonance. SIAM Journal on Mathematical Analysis, 9(2):328–355, April 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olver:1991:UEIb

[Olv91a] F. W. J. Olver. Uniform, exponentially improved, asymptotic expansions for the confluent hypergeometric function and other integral transforms. SIAM Journal on Mathematical Analysis, 22 (5):1475–1489, September 1991.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olver:1991:UEIa

[Olv91b] F. W. J. Olver. Uniform, exponentially improved, asymptotic expansions for the generalized exponential integral. SIAM Journal on Mathematical Analysis, 22 (5):1460–1474, September 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olver:1992:NDO

[Olv92] Peter J. Olver. A nonlinear differential operator series that commutes with any function. SIAM Journal on Mathematical Analysis, 23(1):209–221, January 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olver:1993:EIAb

[Olv93] F. W. J. Olver. Exponentiallyimproved asymptotic solutions of ordinary differential equations. I. The confluent hypergeometric function. SIAM Journal on Mathematical Analysis, 24(3): 756–767, May 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

OMalley:1970:BVP

[O’M70] R. E. O’Malley, Jr. On boundary value problems for a singularly perturbed differential equation with a turning point. SIAM Journal on Mathematical Analysis, 1(4):479–490, November 1970. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

OMalley:1971:ASI

[O’M71] R. E. O’Malley, Jr. On the asymptotic solution of initial value problems for differential equations with small delay. SIAM Journal on Mathematical Analysis, 2(2):259–268, May 1971. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

OMalley:1979:SSP

[O’M79] R. E. O’Malley, Jr. A singular singularly-perturbed linear boundary value problem. SIAM Journal on Mathematical Analysis, 10(4):695–708, July 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ortiz:1987:LRS

[OP87] E. L. Ortiz and A. Pham Ngoc Dinh. Linear recursive schemes associated with some nonlinear partial differential equations in one dimension and the tau method. SIAM Journal on Mathematical Analysis, 18(2): 452–464, March 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

ORegan:1987:TTA

[O’R87] Daniel J. O’Regan. Topological transversality: applications to third order boundary value problems. SIAM Journal on Mathematical Analysis, 18(3):630–641, May 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

ORegan:1993:SGE

[O’R93] Donal O’Regan. Some general existence principles and results for (ϕ(y′)) = qf(t, y, y′), 0 < t < 1. SIAM Journal on Mathematical Analysis, 24(3):648–668, May 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Orton:1973:HTP

[Ort73] Marion Orton. Hilbert transforms, plemelj relations, and Fourier transforms of distributions. SIAM Journal on Mathematical Analysis, 4(4):656–670, November 1973. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Orton:1980:GAE

[Ort80] Marion Orton. The generalized Abel equations for Schwartz distributions. SIAM Journal on Mathematical Analysis, 11(3): 596–611, May 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ortega:1994:SEN

[Ort94] Rafael Ortega. The stability of the equilibrium of a nonlinear Hill’s equation. SIAM Journal on Mathematical Analysis, 25(5): 1393–1401, September 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/23920.

Osher:1993:LSF

[Osh93] Stanley Osher. A level set formulation for the solution of the Dirichlet problem for Hamilton– Jacobi equations. SIAM Journal on Mathematical Analysis, 24(5): 1145–1152, September 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Osler:1970:FDC

[Osl70] Thomas J. Osler. The fractional derivative of a composite function. SIAM Journal on Mathematical Analysis, 1(2):288–293, May 1970. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Osler:1971:TSG

[Osl71] Thomas J. Osler. Taylor’s series generalized for fractional derivatives and applications. SIAM Journal on Mathematical Analysis, 2(1):37–48, February 1971. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Osler:1972:FEL

[Osl72] Thomas J. Osler. A further extension of the Leibniz rule to fractional derivatives and its relation to Parseval’s formula. SIAM Journal on Mathematical Analysis, 3(1):1–16, February 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Osler:1973:CLR

[Osl73] Thomas J. Osler. A correction to Leibniz rule for fractional derivatives. SIAM Journal on Mathematical Analysis, 4(3):456–459, May 1973. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ostrov:1996:ABT

[Ost96]  Daniel N. Ostrov.  Asymp-

totic behavior of two interreacting chemicals in a chromatography reactor. SIAM Journal on Mathematical Analysis, 27(6): 1559–1596, November 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/27570.

OldeDaalhuis:1994:UAT

[OT94] A. B. Olde Daalhuis and N. M.

Temme. Uniform Airy-type expansions of integrals. SIAM Journal on Mathematical Analysis, 25(2):304–321, March 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/18768.

Pagani:1986:SSD

[Pag86] Carlo Domenico Pagani. Stability of a surface determined from measures of potential. SIAM Journal on Mathematical Analysis, 17(1):169–181, January 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pal:1977:CDA

[Pal77] Jenö Pal. On a concept of a derivative among functions defined on the dyadic field. SIAM Journal on Mathematical Analysis, 8(3):375–391, May 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pandey:1971:RTC

[Pan71] J. N. Pandey. A representation theorem for a class of convolution transformable generalized functions. SIAM Journal on Mathematical Analysis, 2(2): 286–289, May 1971. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pandey:1972:DPL

[Pan72] J. N. Pandey. The dual Poisson– Laguerre transform of a class of generalized functions. SIAM Journal on Mathematical Analysis, 3(2):211–229, May 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pan:1996:RWL

[Pan96] K. Pan. A refined Wiener– Levinson method in frequency analysis. SIAM Journal on Mathematical Analysis, 27(5): 1448–1453, September 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/27387.

Pao:1972:SAS

[Pao72] C. V. Pao. Semigroups and asymptotic stability of nonlinear differential equations. SIAM Journal on Mathematical Analysis, 3(2):371–379, May 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pao:1974:SAS

[Pao74] C. V. Pao. Successive approximations of some nonlinear initialboundary value problems. SIAM Journal on Mathematical Analysis, 5(1):91–102, February 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pao:1980:NGS

[Pao80] C. V. Pao. Nonexistence of global solutions for an integrodifferential system in reactor dynamics. SIAM Journal on Mathematical Analysis, 11(3): 559–564, May 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pao:1983:CSS

[Pao83] C. V. Pao. Comparison and stability of solutions for a neutron transport problem with temperature feedback. SIAM Journal on Mathematical Analysis, 14(1): 167–184, January 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pao:1987:CRD

[Pao87] C. V. Pao. On a coupled reaction diffusion system with time delays. SIAM Journal on Mathematical Analysis, 18(4):1026– 1039, July 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Papanicolaou:1995:TFB

[Pap95] Vassilis G. Papanicolaou. Trace formulas and the behaviour of large eigenvalues. SIAM Journal on Mathematical Analysis, 26 (1):218–237, January 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/22460.

Parter:1970:NEPa

[Par70a] Seymour V. Parter. Nonlinear eigenvalue problems for some fourth order equations. I: Maximal solutions. SIAM Journal on Mathematical Analysis, 1(4): 437–457, November 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Parter:1970:NEPb

[Par70b] Seymour V. Parter. Nonlinear eigenvalue problems for some fourth order equations. II: Fixedpoint methods. SIAM Journal on Mathematical Analysis, 1(4): 458–478, November 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Parter:1972:RSP

[Par72a] Seymour V. Parter. Remarks on singular perturbation of certain nonlinear two-point boundary value problems. SIAM Journal on Mathematical Analysis, 3 (2):295–299, May 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Parter:1972:RET

[Par72b] Seymour V. Parter. Remarks on the existence theory for multiple solutions of a singular perturbation problem. SIAM Journal on Mathematical Analysis, 3 (3):496–505, August 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Park:1975:EPE

[Par75] Chull Park. An extension of Parseval’s equation. SIAM Journal on Mathematical Analysis, 6 (4):714–723, August 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Paris:1984:IBF

[Par84] R. B. Paris. An inequality for the Bessel function J_{ν}(νx). SIAM Journal on Mathematical Analysis, 15(1):203–205, January 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Parrott:1992:LSI

[Par92] Mary E. Parrott. Linearized stability and irreducibility for a functional differential equation. SIAM Journal on Mathematical Analysis, 23(3):649–661, May 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Paris:1994:GPI

[Par94] R. B. Paris. A generalization of Pearcey’s integral. SIAM Journal on Mathematical Analysis, 25 (2):630–645, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23054.

Passow:1974:ECS

[Pas74] Eli Passow. Extended Chebyshev systems on (−∞, ∞). SIAM Journal on Mathematical Analysis, 5(5):762–763, October 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pasa:1989:BEC

[Paş89] Gelu I. Paşa. Bounds for effective coefficients of periodic fiberreinforced materials. SIAM Journal on Mathematical Analysis, 20 (1):106–115, January 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Patrick:1971:SIC

[Pat71] Merrell L. Patrick. Some inequalities concerning Jacobi polynomials. SIAM Journal on Mathematical Analysis, 2(2):213–220, May 1971. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Patula:1979:GOP

[Pat79a] William T. Patula. Growth and oscillation properties of second order linear difference equations. SIAM Journal on Mathematical Analysis, 10(1):55–61, January 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Patula:1979:GOC

[Pat79b] William T. Patula. Growth, oscillation and comparison theorems for second-order linear difference equations. SIAM Journal on Mathematical Analysis, 10(6): 1272–1279, November 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Payne:1970:PBC

[Pay70a] L. E. Payne. On a priori bounds in the Cauchy problem for elliptic equations. SIAM Journal on Mathematical Analysis, 1(1):82– 89, February 1970. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Payne:1970:SII

[Pay70b] L. E. Payne. Some isoperimetric inequalities for harmonic functions. SIAM Journal on Mathematical Analysis, 1(3):354–359, August 1970. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [Pay74].

Payne:1974:ESI

[Pay74] L. E. Payne. Erratum: Some isoperimetric inequalities for harmonic functions. SIAM Journal on Mathematical Analysis, 5(5):847, October 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [Pay70b].

Pazy:1972:ALT

[Paz72] A. Pazy. On the applicability of Lyapunov’s theorem in Hilbert space. SIAM Journal on Mathematical Analysis, 3(2):291–294, May 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Percell:1985:FDB

[PB85] Peter B. Percell and Peter N. Brown. Finite determination of bifurcation problems. SIAM Journal on Mathematical Analysis, 16(1):28–46, January 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Poracka-Divi:1971:ETC

[PD71] Zita Poracká-Divi. Existence theorem and convergence of minimizing sequences in extremum problems. SIAM Journal on Mathematical Analysis, 2(4): 505–510, November 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peherstorfer:1981:CPQ

[Peh81] Franz Peherstorfer. Characterization of positive quadrature formulas. SIAM Journal on Mathematical Analysis, 12(6):935–942, November 1981. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peherstorfer:1984:CQF

[Peh84a] Franz Peherstorfer. Characterization of quadrature formula. II. SIAM Journal on Mathematical Analysis, 15(5):1021–1030, September 1984. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peherstorfer:1984:SCB

[Peh84b] Franz Peherstorfer. On a sufficient condition for best L^{1}- approximation. SIAM Journal on Mathematical Analysis, 15(6): 1188–1195, November 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peherstorfer:1990:BSO

[Peh90] Franz Peherstorfer. On Bernstein– Szegő orthogonal polynomials on several intervals. SIAM Journal on Mathematical Analysis, 21 (2):461–482, March 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Penzel:1992:SSM

[Pen92] F. Penzel. Sobolev space methods for dual integral equations in axialsymmetric screen problems. SIAM Journal on Mathematical Analysis, 23(5):1167– 1181, September 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Perry:1973:FMC

[Per73] William L. Perry. Finite Mellin convolution equations. SIAM Journal on Mathematical Analysis, 4(2):238–244, May 1973. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Perry:1974:IGI

[Per74] William L. Perry. Inversion of general integral transforms. SIAM Journal on Mathematical Analysis, 5(5):764–773, October 1974. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Persek:1981:HIA

[Per81] Stephen C. Persek. Hierarchies of iterated averages for systems of ordinary differential equations with a small parameter. SIAM Journal on Mathematical Analysis, 12(3):413–420, May 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Persek:1982:ASM

[Per82] Stephen C. Persek. Asymptotic stability and the method of iterated averages for almost periodic systems. SIAM Journal on Mathematical Analysis, 13(5): 828–841, September 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Perlstadt:1984:POP

[Per84] Marci Perlstadt. A property of orthogonal polynomial families with polynomial duals. SIAM Journal on Mathematical Analysis, 15(5):1043–1054, September 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Perlstadt:1986:PAP

[Per86] Marci Perlstadt. Polynomial analogues of prolate spheroidal wave functions and uncertainty. SIAM Journal on Mathematical Analysis, 17(1):240–248, January 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Perlstadt:1988:OPS

[Per88] Marci A. Perlstadt. Oscillation properties for some polynomial analogues of the prolate spheroidal wave functions. SIAM Journal on Mathematical Analysis, 19(3):751–761, May 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peroueme:1993:BPO

[Pér93] Marie-Christine Pérouème. Bifurcation from a periodic orbit for a strongly resonant reversible autonomous vector field. SIAM Journal on Mathematical Analysis, 24(6):1577–1596, November 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peszek:1995:OST

[Pes95] Robert Peszek. Oscillations of solutions to the two-dimensional Broadwell model, an H-measure approach. SIAM Journal on Mathematical Analysis, 26(3): 750–760, May 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24870.

Peterson:1981:EUF

[Pet81] Allan C. Peterson. Existenceuniqueness for focal-point boundary value problems. SIAM Journal on Mathematical Analysis, 12 (2):173–185, March 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peterson:1984:BVP

[Pet84] Allan C. Peterson. Boundary value problems for an n th order linear difference equation. SIAM Journal on Mathematical Analysis, 15(1):124–132, January 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Petrushev:1998:ARF

[Pet98]  Pencho P. Petrushev.  Ap-

proximation by ridge functions and neural networks. SIAM Journal on Mathematical Analysis, 30(1):155–189, 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/32295.

Peyser:1971:HEM

[Pey71] Gideon Peyser. Hyperbolic equations with multiple characteristics and time dependent coefficients. SIAM Journal on Mathematical Analysis, 2(3):402–412, August 1971. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peyser:1975:SSH

[Pey75] Gideon Peyser. Stability of symmetric hyperbolic systems with nonlinear feedback. SIAM Journal on Mathematical Analysis, 6(6):925–936, November 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pfaltzgraff:1986:UCS

[Pfa86] John A. Pfaltzgraff. Univalence constraints on the Schwarz– Christoffel parameters. SIAM Journal on Mathematical Analysis, 17(1):231–235, January 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peckham:1995:BBS

[PFK95] Bruce B. Peckham, Christos E. Frouzakis, and Ioannis G. Kevrekidis. Bananas and banana splits: A parametric degeneracy in the Hopf bifurcation for maps. SIAM Journal on Mathematical Analysis, 26(1):190–217, January 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/24307.

Prasad:1993:NGK

[PG93] V. C. Prasad and P. B. L.

Gaur. A note on the generalised Katzenelson’s method for piecewise linear resistive networks. SIAM Journal on Mathematical Analysis, 24(6):1662–1665, November 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Podio-Guidugli:1989:EES

[PGRSV89] P. Podio-Guidugli, M. Rosati, A. Schiaffino, and V. Valente. Equilibrium of an elastic spherical cap pulled at the rim. SIAM Journal on Mathematical Analysis, 20(3):643–663, May 1989.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peng:1975:CNC

[PH75] S. T. Peng and A. Hessel. Convergence of noncommutative continued fractions. SIAM Journal on Mathematical Analysis, 6 (4):724–727, August 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Piele:1973:ANF

[Pie73] D. T. Piele. Asymptotically neutral families in E^{3}. SIAM Journal on Mathematical Analysis, 4 (2):260–268, May 1973. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Piele:1974:AHF

[Pie74] D. T. Piele. An approximation of harmonic functions in E^{3} by potentials of unit charges. SIAM Journal on Mathematical Analysis, 5(4):563–568, August 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pierre:1983:PCS

[Pie83] Michel Pierre. Parabolic capacity and Sobolev spaces. SIAM Journal on Mathematical Analysis, 14 (3):522–533, May 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pilipovic:1986:GZS

[Pil86] Stevan Pilipović. Generalization of Zemanian spaces of generalized functions which have orthonormal series expansions. SIAM Journal on Mathematical Analysis, 17(2):477–484, March 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pinsky:1980:EET

[Pin80] Mark A. Pinsky. The eigenvalues of an equilateral triangle. SIAM Journal on Mathematical Analysis, 11(5):819–827, September 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pinsky:1985:CEE

[Pin85] Mark A. Pinsky. Completeness of the eigenfunctions of the equilateral triangle. SIAM Journal on Mathematical Analysis, 16(4): 848–851, July 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pinkus:1988:SI

[Pin88a] A. Pinkus. On smoothest interpolants. SIAM Journal on Mathematical Analysis, 19(6):1431– 1441, November 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pinsky:1988:GDP

[Pin88b] Ross G. Pinsky. A generalized Dirichlet principle for second order nonselfadjoint elliptic operators. SIAM Journal on Mathematical Analysis, 19(1):204–213, January 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pippenger:1993:EAS

[Pip93] Nicholas Pippenger. An elementary approach to some analytic asymptotics. SIAM Journal on Mathematical Analysis, 24(5): 1361–1377, September 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Piquette:1986:AEC

[Piq86] Jean C. Piquette. An analytical expression for coefficients arising when implementing a technique for indefinite integration of products of special functions. SIAM Journal on Mathematical Analysis, 17(4):1033–1035, July 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Piquette:1989:ATE

[Piq89] Jean C. Piquette. Applications of a technique for evaluating indefinite integrals containing products of the special functions of physics. SIAM Journal on Mathematical Analysis, 20(5): 1260–1269, September 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pitkaranta:1980:EDS

[Pit80] Juhani Pitkäranta. Estimates for the derivatives of solutions to weakly singular Fredholm integral equations. SIAM Journal on Mathematical Analysis, 11(6): 952–968, November 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Petty:1973:PSE

[PJ73] C. M. Petty and W. E. Johnson. Properties of solutions of u′′ + c(t)f(u)h(u′) = 0 with explicit initial conditions. SIAM Journal on Mathematical Analysis, 4(2):269–282, May 1973. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peckham:1991:PDH

[PK91] Bruce B. Peckham and Ioannis G. Kevrekidis. Period doubling with higher-order degeneracies. SIAM Journal on Mathematical Analysis, 22(6):1552– 1574, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Painter:1982:TPC

[PM82] J. F. Painter and R. E. Meyer.

Turning-point connection at close quarters. SIAM Journal on Mathematical Analysis, 13(4): 541–554, July 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Poggiolini:1999:AES

[Pog99] Laura Poggiolini. Almost everywhere solutions of partial differential equations and systems of any order. SIAM Journal on Mathematical Analysis, 30(5):1136–1150, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/33612.

Poreda:1976:GCF

[Por76] S. J. Poreda. A generalization of the Carathéodory–Fejér problem. SIAM Journal on Mathematical Analysis, 7(2):200–212, April 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Porter:1980:SPS

[Por80] William A. Porter. Synthesis of polynomic systems. SIAM Journal on Mathematical Analysis, 11 (2):308–315, March 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pospiech:1990:DPE

[Pos90] Christoph Pospiech. A Dirichlet problem exhibiting global bifurcation with symmetry breaking. SIAM Journal on Mathematical Analysis, 21(6):1426– 1452, November 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pourciau:1983:GPP

[Pou83] B. H. Pourciau. Global properties of proper Lipschitzian maps. SIAM Journal on Mathematical Analysis, 14(4):796–799, July 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Poulsen:1989:NCH

[Pou89] Ebbe Thue Poulsen. Nonmonotone clines in homogeneous space cannot be stable. SIAM Journal on Mathematical Analysis, 20 (1):148–159, January 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Payne:1979:SMP

[PP79] L. E. Payne and G. A. Philippin.

On some maximum principles involving harmonic functions and their derivatives. SIAM Journal on Mathematical Analysis, 10 (1):96–104, January 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See remark [Ban79].

Payne:1983:IIT

[PP83] L. E. Payne and G. A. Philippin.

Isoperimetric inequalities in the torsion and clamped membrane problems for convex plane domains. SIAM Journal on Mathematical Analysis, 14(6):1154– 1162, November 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Piila:1993:EER

[PP93] Jyrki Piila and Juhani Pitkäranta.

Energy estimates relating different linear elastic models of a thin cylindrical shell I. the membranedominated case. SIAM Journal on Mathematical Analysis, 24(1): 1–22, January 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Piila:1995:EER

[PP95] Jyrki Piila and Juhani Pitkäranta.

Energy estimates relating different linear elastic models of a thin cylindrical shell II: The case of free boundary. SIAM Journal on Mathematical Analysis, 26(4):820–849, July 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/19803.

Pandey:1971:MLE

[PR71] J. N. Pandey and M. Rahman. The maximum likelihood estimate of the noncentrality parameter of a noncentral F variate. SIAM Journal on Mathematical Analysis, 2(2):269–276, May 1971. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pilant:1987:IPN

[PR87] Michael Pilant and William Rundell. An inverse problem for a nonlinear elliptic differential equation. SIAM Journal on Mathematical Analysis, 18(6): 1801–1809, November 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peterson:1991:ASO

[PR91] Allan C. Peterson and Jerry Ridenhour. Atkinson’s superlinear oscillation theorem for matrix difference equations. SIAM Journal on Mathematical Analysis, 22 (3):774–784, May 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Prabhakar:1972:HIE

[Pra72] Tilak Raj Prabhakar. Hypergeometric integral equations of a general kind and fractional integration. SIAM Journal on Mathematical Analysis, 3(3):422–425, August 1972. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Price:1975:LTD

[Pri75]  Douglas B. Price.  On the

Laplace transform for distributions. SIAM Journal on Mathematical Analysis, 6(1):49–80, February 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pruss:1988:BSV

[Prü88] Jan Prüss. Bounded solutions of Volterra equations. SIAM Journal on Mathematical Analysis, 19 (1):133–149, January 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pethe:1972:MAE

[PS72] S. P. Pethe and A. Sharma. Modified Abel expansion and a subclass of completely convex functions. SIAM Journal on Mathematical Analysis, 3(3):546–558, August 1972. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Payne:1976:SPI

[PS76] L. E. Payne and R. P. Sperb.

Some a priori and isoperimetric inequalities associated with a class of boundary value problems for doubly connected domains. SIAM Journal on Mathematical Analysis, 7(4):451–460, August 1976. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Pence:1982:APB

[PS82] D. D. Pence and P. W. Smith. Asymptotic properties of best L_{p}[0, 1] approximation by splines. SIAM Journal on Mathematical Analysis, 13(3):409–420, May 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Perthame:1988:NPN

[PS88] Benoît Perthame and Richard Sanders. The Neumann problem for nonlinear second order singular perturbation problems. SIAM Journal on Mathematical Analysis, 19(2):295–311, March 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Paivarinta:1991:IDS

[PS91a] Lassi Päivärinta and Erkki Somersalo. Inversion of discontinuities for the Schrödinger equation in three dimensions. SIAM Journal on Mathematical Analysis, 22 (2):480–499, March 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pittaluga:1991:IZA

[PS91b] Giovanna Pittaluga and Laura Sacripante. Inequalities for the zeros of the Airy functions. SIAM Journal on Mathematical Analysis, 22(1):260–267, January 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pucci:1994:ASI

[PS94] Patrizia Pucci and James Serrin. Asymptotic stability for intermittently controlled nonlinear oscillators. SIAM Journal on Mathematical Analysis, 25 (3):815–835, May 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24067.

Pierre:1997:BRD

[PS97] Michel Pierre and Didier Schmitt.

Blowup in reaction-diffusion systems with dissipation of mass. SIAM Journal on Mathematical Analysis, 28(2):259–269, March 1997. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/29543.

Paivarinta:1998:RSM

[PS98a] Lassi Päivärinta and Valeri Serov. Recovery of singularities of a multidimensional scattering potential. SIAM Journal on Mathematical Analysis, 29(3):697–711, May 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30579.

Plonka:1998:CMF

[PS98b] G. Plonka and V. Strela. Construction of multiscaling functions with approximation and symmetry. SIAM Journal on Mathematical Analysis, 29(2):

481–510, March 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/29718.

Patil:1978:MGF

[PT78] K. R. Patil and N. K. Thakare. Multilinear generating function for the Konhauser biorthogonal polynomial sets. SIAM Journal on Mathematical Analysis, 9 (5):921–923, October 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Peletier:1997:SPD

[PT97] L. A. Peletier and W. C. Troy. Spatial patterns described by the extended Fisher–Kolmogorov equation: Periodic solutions. SIAM Journal on Mathematical Analysis, 28(6):1317–1353, November 1997. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

                                        http://epubs.siam.org/sam-
                                        bin/dbq/article/28095.

Pohozaev:1999:ENS

[PT99] Stanislav I. Pohozaev and Alberto Tesei. Existence and nonexistence of solutions of nonlinear Neumann problems. SIAM Journal on Mathematical Analysis, 31(1):119–133, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/33494.

Piquette:1984:TEI

[PV84] Jean C. Piquette and A. L. Van Buren. Technique for evaluating indefinite integrals involving products of certain special functions. SIAM Journal on Mathematical Analysis, 15(4):845–855, July 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Pollard:1970:ICT

[PW70] Harry Pollard and D. V. Widder. Inversion of a convolution transform related to heat conduction. SIAM Journal on Mathematical Analysis, 1(4):527–532, November 1970. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Parker:1982:CS

[PW82] G. Edgar Parker and Terry J.

Walters. Constructing solutions to ∫_{0}^{1} f(j)A(j, x)/f(x) dj = ∫_{0}^{1} f(x)A(x, j)/f(j) dj. SIAM Journal on Mathematical Analysis, 13(5):856–865, September 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Qi:1996:DFD

[Qi96] Yuan-Wei Qi. The degeneracy of a fast-diffusion equation and stability. SIAM Journal on Mathematical Analysis, 27 (2):476–485, March 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/26002.

Quak:1993:LSA

[QSW93] E. Quak, N. Sivakumar, and J. D. Ward. Least squares approximation by radial functions. SIAM Journal on Mathematical Analysis, 24(4):1043– 1066, July 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Queen:1971:GWT

[Que71] W. C. Queen. A generalized Weierstrass transformation for the case of several independent variables. SIAM Journal on Mathematical Analysis, 2(1):52– 63, February 1971. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Quinto:1993:SXR

[Qui93] Eric Todd Quinto. Singularities of the X-ray transform and limited data tomography in R^{2} and R^{3}. SIAM Journal on Mathematical Analysis, 24(5):1215– 1225, September 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Quintanilla:1996:SDE

[Qui96] R. Quintanilla. A spatial decay estimate for the hyperbolic heat equation. SIAM Journal on Mathematical Analysis, 27 (1):78–91, January 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/25693.

Qiu:1996:SEC

[QV96] S.-L. Qiu and M. K. Vamanamurthy. Sharp estimates for complete elliptic integrals. SIAM Journal on Mathematical Analysis, 27(3):823–834, May 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Qiu:1999:IPN

[QV99] S.-L. Qiu and M. Vuorinen.

Infinite products and normalized quotients of hypergeometric functions. SIAM Journal on Mathematical Analysis, 30(5):1057–1075, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32680.

Qiu:1998:SIG

[QVV98] S. L. Qiu, M. K. Vamanamurthy, and M. Vuorinen. Some inequalities for the growth of elliptic integrals. SIAM Journal on Mathematical Analysis, 29(5): 1224–1237, September 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/31049.

Rathore:1979:UPM

[RA79] R. K. S. Rathore and P. N.

Agrawal. The unsettled problem of M. G. Kreĭn on nonnegative polynomials in T -systems. SIAM Journal on Mathematical Analysis, 10(5):1092–1094, September 1979. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Rabinowitz:1982:PSH

[Rab82] Paul H. Rabinowitz. Periodic solutions of Hamiltonian systems: a survey. SIAM Journal on Mathematical Analysis, 13(3): 343–352, May 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rabier:1984:NSC

[Rab84] Patrick Rabier. A necessary and sufficient condition for the coerciveness of a class of functionals and its applications. SIAM Journal on Mathematical Analysis, 15 (2):367–388, March 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rabier:1987:SSO

[Rab87] Patrick Rabier. A synthetic study of one-parameter nonlinear problems. SIAM Journal on Mathematical Analysis, 18 (4):1116–1141, July 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rahman:1976:CFP

[Rah76a] M. Rahman. Construction of a family of positive kernels from Jacobi polynomials. SIAM Journal on Mathematical Analysis, 7 (1):92–116, February 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rahman:1976:FPF

[Rah76b] M. Rahman. A five-parameter family of positive kernels from Jacobi polynomials. SIAM Journal on Mathematical Analysis, 7 (3):386–413, May 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rahman:1976:SPK

[Rah76c] M. Rahman. Some positive kernels and bilinear sums for Hahn polynomials. SIAM Journal on Mathematical Analysis, 7 (3):414–435, May 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rahman:1977:GPK

[Rah77] Mizan Rahman. On a generalization of the Poisson kernel for Jacobi polynomials. SIAM Journal on Mathematical Analysis, 8(6): 1014–1031, November 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rahman:1978:PKH

[Rah78] Mizan Rahman. A positive kernel for Hahn–Eberlein polynomials. SIAM Journal on Mathematical Analysis, 9(5):891–905, October 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rahman:1979:EPD

[Rah79] Mizan Rahman. An elementary proof of Dunkl’s addition theorem for Krawtchouk polynomials. SIAM Journal on Mathematical Analysis, 10(2):438–445, March 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rahman:1981:FBR

[Rah81a] Mizan Rahman. Families of biorthogonal rational functions in a discrete variable. SIAM Journal on Mathematical Analysis, 12(3):355–367, May 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rahman:1981:SMB

[Rah81b] Mizan Rahman. A stochastic matrix and bilinear sums for Racah– Wilson polynomials. SIAM Journal on Mathematical Analysis, 12 (2):145–160, March 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rahman:1986:ACS

[Rah86a] Mizan Rahman. Another conjectured q-Selberg integral. SIAM Journal on Mathematical Analysis, 17(5):1267–1279, September 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rahman:1986:WFS

[Rah86b] Mizan Rahman. q-Wilson functions of the second kind. SIAM Journal on Mathematical Analysis, 17(5):1280–1286, September 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rahman:1991:BSR

[Rah91] Mizan Rahman. Biorthogonality of a system of rational functions with respect to a positive measure on [−1, 1]. SIAM Journal on Mathematical Analysis, 22(5): 1430–1441, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rall:1982:IIF

[Ral82] L. B. Rall. Integration of interval functions. II. The finite case. SIAM Journal on Mathematical Analysis, 13(4):690–697, July 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rall:1983:MVT

[Ral83] L. B. Rall. Mean value and Taylor forms in interval analysis. SIAM Journal on Mathematical Analysis, 14(2):223–238, March 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rao:1975:BPF

[Rao75] V. R. Gopala Rao. On the behavior of principal fundamental solutions of elliptic equations. SIAM Journal on Mathematical Analysis, 6(6):978–988, November 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Raphael:1982:SSM

[Rap82] Louise A. Raphael. The Stieltjes summability method and summing Sturm–Liouville expansions. SIAM Journal on Mathematical Analysis, 13(4):676–689, July 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rathore:1974:SLT

[Rat74] R. K. S. Rathore. On a sequence of linear trigonometric polynomial operators. SIAM Journal on Mathematical Analysis, 5(6): 908–917, November 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Raymon:1981:PMI

[Ray81] L. Raymon. Piecewise monotone interpolation in polynomial time. SIAM Journal on Mathematical Analysis, 12(1):110–114, January 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Roberts:1971:SEO

[RB71] Charles E. Roberts and Geneva G. Belford. Stability and entering of the origin for real, nonlinear, autonomous differential equations of third order. SIAM Journal on Mathematical Analysis, 2(1): 133–148, February 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rodriguez-Bernal:1998:LSH

[RB98] Aníbal Rodríguez-Bernal. Localized spatial homogenization and large diffusion. SIAM Journal on Mathematical Analysis, 29(6): 1361–1380, November 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/31864.

Rognlie:1973:AIT

[RC73] D. M. Rognlie and B. C. Carlson. Averaged integral transforms. SIAM Journal on Mathematical Analysis, 4(2):393–407, May 1973. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ronveaux:1977:BNP

[RDLH77] A. Ronveaux, M-C. Dumont-Lepage, and J. Habay. Bounds on the number of poles of the solution of a general Riccati equation. SIAM Journal on Mathematical Analysis, 8(6): 1038–1056, November 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ren:1998:MPS

[RE98] Liwei Ren and G. Bard Ermentrout. Monotonicity of phaselocked solutions in chains and arrays of nearest-neighbor coupled oscillators. SIAM Journal on Mathematical Analysis, 29(1): 208–234, January 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29883.

Reade:1983:EPD

[Rea83] J. B. Reade. Eigenvalues of positive definite kernels. SIAM Journal on Mathematical Analysis, 14 (1):152–157, January 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Reade:1984:EPD

[Rea84] J. B. Reade. Eigenvalues of positive definite kernels. II. SIAM Journal on Mathematical Analysis, 15(1):137–142, January 1984.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Reade:1985:SWE

[Rea85] J. B. Reade. On the sharpness of Weyl’s estimate for eigenvalues of smooth kernels. SIAM Journal on Mathematical Analysis, 16(3): 548–550, May 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Reade:1986:OMF

[Rea86a] J. B. Reade. On the order of magnitude of Fourier coefficients. SIAM Journal on Mathematical Analysis, 17(2):469–476, March 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Reade:1986:PDK

[Rea86b] J. B. Reade. Positive definite C^{p} kernels. SIAM Journal on Mathematical Analysis, 17 (2):420–421, March 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Reach:1987:RRS

[Rea87] Michael Reach. Recursion relations for solutions to the Schrödinger equation. SIAM Journal on Mathematical Analysis, 18(2):465–472, March 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Reade:1988:SWE

[Rea88] J. B. Reade. On the sharpness of Weyl’s estimates for eigenvalues of smooth kernels. II. SIAM Journal on Mathematical Analysis, 19(3):627–631, May 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Redinbo:1971:NCG

[Red71] G. Robert Redinbo. A note on the construction of generalized Walsh functions. SIAM Journal on Mathematical Analysis, 2(2): 166–167, May 1971. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Redheffer:1984:CFS

[Red84] Ray Redheffer. Convergence of Fourier series at a discontinuity. SIAM Journal on Mathematical Analysis, 15(5):1007–1009, September 1984. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Redlinger:1985:VPE

[Red85]  Reinhard Redlinger.  On

Volterra’s population equation with diffusion. SIAM Journal on Mathematical Analysis, 16(1): 135–142, January 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Reid:1970:MPS

[Rei70] William T. Reid. Monotoneity properties of solutions of Hermitian Riccati matrix differential equations. SIAM Journal on Mathematical Analysis, 1(2): 195–213, May 1970. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Reid:1973:BPS

[Rei73] William T. Reid. Boundary problems of Sturmian type on an infinite interval. SIAM Journal on Mathematical Analysis, 4(1): 185–197, February 1973. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Reid:1975:VMQ

[Rei75] William T. Reid. Variational methods and quadratic functional inequalities. SIAM Journal on Mathematical Analysis, 6 (2):404–416, April 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Reid:1979:SAO

[Rei79] William T. Reid. Some aspects of oscillation and boundary problem theory for Hamiltonian systems in a B^{∗}-algebra. SIAM Journal on Mathematical Analysis, 10(4):667–681, July 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Reid:1986:SHS

[Rei86] Gregory John Reid. R-separation for heat and Schrödinger equations. I. SIAM Journal on Mathematical Analysis, 17(3): 646–687, May 1986. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Reitich:1991:RSP

[Rei91]  Fernando Reitich.  Rapidly

stretching plastic jets: The linearized problem. SIAM Journal on Mathematical Analysis, 22 (1):107–128, January 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Reissel:1993:TBV

[Rei93] Martin Reissel. On a transmission boundary-value problem for the time-harmonic Maxwell equations without displacement currents. SIAM Journal on Mathematical Analysis, 24(6): 1440–1457, November 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Reif:1997:OCB

[Rei97] Ulrich Reif. Orthogonality of cardinal B-splines in weighted Sobolev spaces. SIAM Journal on Mathematical Analysis, 28(5): 1258–1263, September 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29685.

Renardy:1982:QPE

[Ren82]  M. Renardy.  A quasilinear

parabolic equation describing the elongation of thin filaments of polymeric liquids. SIAM Journal on Mathematical Analysis, 13 (2):226–238, March 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Renardy:1984:SPH

[Ren84] Michael Renardy. Singularly perturbed hyperbolic evolution problems with infinite delay and an application to polymer rheology. SIAM Journal on Mathematical Analysis, 15(2):333–349, March 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Renardy:1990:LES

[Ren90] Michael Renardy. Local existence of solutions of the Dirichlet initial-boundary value problem for incompressible hypoelastic materials. SIAM Journal on Mathematical Analysis, 21(6): 1369–1385, November 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Renardy:1991:ETM

[Ren91] Michael Renardy. An existence theorem for model equations resulting from kinetic theories of polymer solutions. SIAM Journal on Mathematical Analysis, 22 (2):313–327, March 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Renardy:1996:IVP

[Ren96] Michael Renardy. Initial-value problems with inflow boundaries for Maxwell fluids. SIAM Journal on Mathematical Analysis, 27 (4):914–931, July 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/26591.

Renardy:1997:DPH

[Ren97] Michael Renardy. A degenerate parabolic-hyperbolic system modeling the spreading of surfactants. SIAM Journal on Mathematical Analysis, 28(5):1048– 1063, September 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29912.

Richards:1989:CRP

[Ric89] Donald St. P. Richards. Constant regression polynomials and the Wishart distribution. SIAM Journal on Mathematical Analysis, 20(2):504–510, March 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ricci:1990:TWS

[Ric90] Riccardo Ricci. Traveling wave solutions of the Stefan and the ablation problems. SIAM Journal on Mathematical Analysis, 21 (6):1386–1393, November 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rioul:1992:SRC

[Rio92] Olivier Rioul. Simple regularity criteria for subdivision schemes. SIAM Journal on Mathematical Analysis, 23(6):1544–1576, November 1992. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rivera:1993:DDV

[Riv93] Jaime E. Muñoz Rivera. Decomposition of the displacement vector field and decay rates in linear thermoelasticity. SIAM Journal on Mathematical Analysis, 24 (2):390–406, March 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Renardy:1986:HBT

[RJ86] M. Renardy and D. D. Joseph. Hopf bifurcation in two-component flow. SIAM Journal on Mathematical Analysis, 17(4):894–910, July 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rawlins:1978:MPA

[RM78] A. D. Rawlins and J. D. Morgan. Monotonic properties of analytic functions. SIAM Journal on Mathematical Analysis, 9(6): 1172–1178, December 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Robinson:1983:SRN

[Rob83]  Clark Robinson.  Sustained

resonance for a nonlinear system with slowly varying coefficients. SIAM Journal on Mathematical Analysis, 14(5):847–860, September 1983. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Roberts:1985:ITR

[Rob85] A. J. Roberts. An introduction to the technique of reconstitution. SIAM Journal on Mathematical Analysis, 16(6):1243– 1257, November 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Roberts:1989:UIM

[Rob89] A. J. Roberts. The utility of an invariant manifold description of the evolution of a dynamical system. SIAM Journal on Mathematical Analysis, 20(6): 1447–1458, November 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Robinson:1992:HBT

[Rob92] Clark Robinson. Homoclinic bifurcation to a transitive attractor of Lorenz type, II. SIAM Journal on Mathematical Analysis, 23(5): 1255–1268, September 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rogers:1988:NVP

[Rog88] Robert C. Rogers. Nonlocal variational problems in nonlinear electromagneto-elastostatics. SIAM Journal on Mathematical Analysis, 19(6):1329–1347, November 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ron:1997:SRF

[Ron97] Amos Ron. Smooth refinable functions provide good approximation orders. SIAM Journal on Mathematical Analysis, 28 (3):731–748, May 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/28248.

Rosenzweig:1971:WSM

[Ros71] M. B. Rosenzweig. On weak solutions of a mildly nonlinear Dirichlet problem. SIAM Journal on Mathematical Analysis, 2(4): 483–495, November 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rose:1978:LEG

[Ros78] Nicholas J. Rose. The Laurent expansion of a generalized resolvent with some applications. SIAM Journal on Mathematical Analysis, 9(4):751–758, August 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rosengren:1998:MOP

[Ros98] Hjalmar Rosengren. Multivariable orthogonal polynomials and coupling coefficients for discrete series representations. SIAM Journal on Mathematical Analysis, 30(2):233–272, 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/32568.

Rothe:1993:RPP

[Rot93a] Franz Rothe. Remarks on periods of planar Hamiltonian systems. SIAM Journal on Mathematical Analysis, 24(1):129–154, January 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rother:1993:NEN

[Rot93b] W. Rother. Negative eigenvalues for a nonlinear differential equation on R. SIAM Journal on Mathematical Analysis, 24(3): 669–680, May 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Redheffer:1984:QTD

[RR84] Ray Redheffer and Reinhard Redlinger. Quenching in timedelay systems: a summary and a counterexample. SIAM Journal on Mathematical Analysis, 15(6): 1114–1124, November 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rivera:1995:SPD

[RR95] Jaime E. Muñoz Rivera and Reinhard Racke. Smoothing properties, decay, and global existence of solutions to nonlinear coupled systems of thermoelastic type. SIAM Journal on Mathematical Analysis, 26(6): 1547–1563, November 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/25505.

Redheffer:1988:TSL

[RRW88] Ray Redheffer, Reinhard Redlinger, and Wolfgang Walter. A theorem of La Salle–Lyapunov type for parabolic systems. SIAM Journal on Mathematical Analysis, 19 (1):121–132, January 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ridenhour:1974:PRT

[RS74] J. R. Ridenhour and R. P. Soni.

Parseval relation and Tauberian theorems for the Hankel transform. SIAM Journal on Mathematical Analysis, 5(5):809–821, October 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rundell:1976:MAP

[RS76] William Rundell and Michael Stecher. A method of ascent for parabolic and pseudoparabolic partial differential equations. SIAM Journal on Mathematical Analysis, 7(6):898–912, November 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rundell:1978:RAU

[RS78] William Rundell and Michael Stecher. A Runge approximation and unique continuation theorem for pseudoparabolic equations. SIAM Journal on Mathematical Analysis, 9(6):1120–1125, December 1978. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rahman:1985:ISP

[RS85] Mizan Rahman and Mihr J. Shah. An infinite series with products of Jacobi polynomials and Jacobi functions of the second kind. SIAM Journal on Mathematical Analysis, 16(4): 859–875, July 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ruf:1986:CSS

[RS86] Bernhard Ruf and Sergio Solimini. On a class of superlinear Sturm–Liouville problems with arbitrarily many solutions. SIAM Journal on Mathematical Analysis, 17(4):761–771, July 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ringland:1991:GBS

[RS91a] J. Ringland and M. Schell. Genealogy and bifurcation skeleton for cycles of the iterated two-extremum map of the interval. SIAM Journal on Mathematical Analysis, 22(5):1354–1371, September 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rynne:1991:ITP

[RS91b] B. P. Rynne and B. D. Sleeman.

The interior transmission problem and inverse scattering from inhomogeneous media. SIAM Journal on Mathematical Analysis, 22(6):1755–1762, September 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rahman:1994:BRT

[RS94a] Mizan Rahman and Sergei K.

Suslov. Barnes and Ramanujantype integrals on the q-linear lattice. SIAM Journal on Mathematical Analysis, 25(3):1002– 1022, May 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23367.

Rahman:1994:PEB

[RS94b] Mizan Rahman and Sergei K.

Suslov. The Pearson equation and the beta integrals. SIAM Journal on Mathematical Analysis, 25(2):646–693, March 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/22874.

Rubinstein:1995:SMV

[RS95] Jacob Rubinstein and Peter Sternberg. On the slow motion of vortices in the Ginzburg– Landau heat flow. SIAM Journal on Mathematical Analysis, 26(6): 1452–1466, November 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/25940.

Ramanathan:1993:TFL

[RT93] Jayakumar Ramanathan and Pankaj Topiwala. Timefrequency localization via the Weyl correspondence. SIAM Journal on Mathematical Analysis, 24(5):1378–1393, September 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rota:1994:CUC

[RT94] G.-. C. Rota and B. D. Taylor. The classical umbral calculus. SIAM Journal on Mathematical Analysis, 25(2):694–711, March 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/24561.

Roquejoffre:1996:GST

[RTV96] Jean-Michel Roquejoffre, David Terman, and Vitaly A. Volpert. Global stability of traveling fronts and convergence towards stacked families of waves in monotone parabolic systems. SIAM Journal on Mathematical Analysis, 27(5):1261–1269, September 1996. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

http://epubs.siam.org/sam-
bin/dbq/article/26752.

Ruan:1998:LAL

[Rua98] W. H. Ruan. Local average Liapunov functions and persistence in population dynamics. SIAM Journal on Mathematical Analysis, 29(1):134–154, January 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/29737.

Ruf:1995:HSF

[Ruf95] Bernhard Ruf. Higher singularities and forced secondary bifurcation. SIAM Journal on Mathematical Analysis, 26(5):1342– 1360, September 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24384.

Rulla:1987:WSS

[Rul87]  Jim Rulla.  Weak solutions

to Stefan problems with prescribed convection. SIAM Journal on Mathematical Analysis, 18 (6):1784–1800, November 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Russell:1977:ARS

[Rus77] A. M. Russell. Applications of Riemann–Stieltjes integrals of order k in functional analysis. SIAM Journal on Mathematical Analysis, 8(5):879–890, October 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ruscheweyh:1978:KET

[Rus78] Stephan Ruscheweyh. On the Kakeya–Eneström theorem and Gegenbauer polynomial sums. SIAM Journal on Mathematical Analysis, 9(4):682–686, August 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ruscheweyh:1988:EST

[Rus88] Stephan Ruscheweyh. Extension of Szegő’s theorem on the sections of univalent functions. SIAM Journal on Mathematical Analysis, 19(6):1442–1449, November 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rahman:1986:PAF

[RV86] Mizan Rahman and Arun Verma. Product and addition formulas for the continuous q-ultraspherical polynomials. SIAM Journal on Mathematical Analysis, 17(6):1461–1474, November 1986. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Redheffer:1987:EIW

[RV87] Ray Redheffer and Alexander Voigt. An elementary inequality for which equality holds in an infinite-dimensional set. SIAM Journal on Mathematical Analysis, 18(2):486–489, March 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Reddien:1978:NAN

[RW78] G. W. Reddien and G. F.

Webb. Numerical approximation of nonlinear functional differential equations with L^{2} initial functions. SIAM Journal on Mathematical Analysis, 9(6): 1151–1171, December 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Rybka:1999:CVM

[Ryb99] Piotr Rybka. The crystalline version of the modified Stefan problem in the plane and its properties. SIAM Journal on Mathematical Analysis, 30(4):756–786, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32543.

Roach:1991:TPS

[RZ91] G. F. Roach and Bo Zhang.

On transmission problems for the Schrödinger equation. SIAM Journal on Mathematical Analysis, 22(4):991–1006, July 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Saari:1974:ETT

[Saa74] Donald G. Saari. An elementary Tauberian theorem for absolutely continuous functions and for series. SIAM Journal on Mathematical Analysis, 5(4):649–662, August 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sack:1974:GFS

[Sac74] R. A. Sack. Generating functions for spherical harmonics Part I: Three-dimensional harmonics. SIAM Journal on Mathematical Analysis, 5(5):774–796, October 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sachs:1983:CDS

[Sac83] Robert L. Sachs. Completeness of derivatives of squared Schrödinger eigenfunctions and explicit solutions of the linearized KdV equation. SIAM Journal on Mathematical Analysis, 14(4): 674–683, July 1983. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sachs:1984:JKA

[Sac84] Robert L. Sachs. A justification of the KdV approximation to first order in the case of N-soliton water waves in a canal. SIAM Journal on Mathematical Analysis, 15(3):468–489, May 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sacks:1985:GBC

[Sac85] Paul E. Sacks. Global behavior for a class of nonlinear evolution equations. SIAM Journal on Mathematical Analysis, 16 (2):233–250, March 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Safonov:1999:NSO

[Saf99] Mikhail V. Safonov. Nonuniqueness for second-order elliptic equations with measurable coefficients. SIAM Journal on Mathematical Analysis, 30(4):879– 895, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/30904.

Sainsaulieu:1996:TWS

[Sai96] Lionel Sainsaulieu. Travelingwave solutions of convectiondiffusion systems in nonconservation form. SIAM Journal on Mathematical Analysis, 27(5): 1286–1310, September 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sakai:1993:RBQ

[Sak93] Makoto Sakai. Regularity of boundaries of quadrature domains in two dimensions. SIAM Journal on Mathematical Analysis, 24(2):341–364, March 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sala:1989:TJA

[Sal89] Kenneth L. Sala. Transformations of the Jacobian amplitude function and its calculation via the arithmetic-geometric mean. SIAM Journal on Mathematical Analysis, 20(6):1514–1528, November 1989. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sanders:1979:PTR

[San79] Jan Sanders. On the passage through resonance. SIAM Journal on Mathematical Analysis, 10 (6):1220–1243, November 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sanders:1980:AAE

[San80] Jan A. Sanders. Asymptotic approximations and extension of time-scales. SIAM Journal on Mathematical Analysis, 11(4): 758–770, July 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sandefur:1983:EUS

[San83a] James T. Sandefur, Jr. Existence and uniqueness of solutions of second order nonlinear differential equations. SIAM Journal on Mathematical Analysis, 14(3): 477–487, May 1983. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sanders:1983:FTA

[San83b] J. A. Sanders. On the fundamental theorem of averaging. SIAM Journal on Mathematical Analysis, 14(1):1–10, January 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sandstede:1998:SFB

[San98] Björn Sandstede. Stability of Nfronts bifurcating from a twisted heteroclinic loop and an application to the FitzHugh–Nagumo equation. SIAM Journal on Mathematical Analysis, 29(1): 183–207, January 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29738.

Saperstone:1971:DEE

[Sap71] Stephen H. Saperstone. On a differential equation for the eigenvectors of a real symmetric matrix. SIAM Journal on Mathematical Analysis, 2(3):413–419, August 1971. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sasaki:1991:CRA

[Sas91] Takeshi Sasaki. Contiguity relations of Aomoto–Gel′fand hypergeometric functions and applications to Appell’s system F_{3} and Goursat’s system _{3}F_{2}. SIAM Journal on Mathematical Analysis, 22(3):821–846, May 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sattinger:1977:GRT

[Sat77] D. H. Sattinger. Group representation theory and branch points of nonlinear functional equations. SIAM Journal on Mathematical Analysis, 8(2):179–201, April 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sattinger:1990:FCS

[Sat90] D. H. Sattinger. Flat connections and scattering theory on the line. SIAM Journal on Mathematical Analysis, 21(3):729–756, May 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Solomon:1987:IVE

[SAW87] A. Solomon, V. Alexiades, and D. G. Wilson. The initial velocity of the emerging free boundary in a two-phase Stefan problem with imposed flux. SIAM Journal on Mathematical Analysis, 18(5): 1438–1452, September 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Saylor:1971:TRA

[Say71] Richard Saylor. Time reversal in abstract Cauchy problems. SIAM Journal on Mathematical Analysis, 2(3):454–457, August 1971. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sanders:1986:LCJ

[SC86] Jan A. Sanders and Richard Cushman. Limit cycles in the Josephson equation. SIAM Journal on Mathematical Analysis, 17 (3):495–511, May 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schaefer:1973:IES

[Sch73a] Philip W. Schaefer. Improvable estimates in some non-well-posed problems for a system of elliptic equations. SIAM Journal on Mathematical Analysis, 4(3): 447–455, May 1973. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schindler:1973:EIT

[Sch73b] Susan Schindler. Explicit integral transform proofs of some transplantation theorems for the Hankel transform. SIAM Journal on Mathematical Analysis, 4(2): 367–384, May 1973. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schleusner:1974:NBP

[Sch74] John W. Schleusner. A note on biorthogonal polynomials in two variables. SIAM Journal on Mathematical Analysis, 5(1):11– 18, February 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schach:1976:NILa

[Sch76a] S. R. Schach. New identities for Legendre associated functions of integral order and degree. I. SIAM Journal on Mathematical Analysis, 7(1):59–69, February 1976. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Schach:1976:NILb

[Sch76b] S. R. Schach. New identities for Legendre associated functions of integral order and degree. II: Extension to other polynomials. SIAM Journal on Mathematical Analysis, 7(5):754–766, October 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schechter:1976:STH

[Sch76c] Murray Schechter. A solvability theorem for homogeneous functions. SIAM Journal on Mathematical Analysis, 7(5):696–701, October 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schafke:1980:CPT

[Sch80] Reinhard Schäfke. The connection problem for two neighboring regular singular points of general linear complex ordinary differential equations. SIAM Journal on Mathematical Analysis, 11(5): 863–875, September 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schreiber:1981:RST

[Sch81] Robert Schreiber. Regularity of singular two-point boundary value problems. SIAM Journal on Mathematical Analysis, 12(1): 104–109, January 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schafke:1984:CPR

[Sch84a] Reinhard Schäfke. A connection problem for a regular and an irregular singular point of complex ordinary differential equations. SIAM Journal on Mathematical Analysis, 15(2):253–271, March 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schonbek:1984:ESS

[Sch84b] Maria Elena Schonbek. Existence of solutions to singular conservation laws. SIAM Journal on Mathematical Analysis, 15(6): 1125–1139, November 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schaefer:1987:SGL

[Sch87a] Philip W. Schaefer. Solution, gradient, and Laplacian bounds in some nonlinear fourth order elliptic equations. SIAM Journal on Mathematical Analysis, 18 (2):430–434, March 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schecter:1987:MMS

[Sch87b] Stephen Schecter. Mel′nikov’s method at a saddle-node and the dynamics of the forced Josephson junction. SIAM Journal on Mathematical Analysis, 18(6): 1699–1715, November 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schecter:1987:SNS

[Sch87c] Stephen Schecter. The saddlenode separatrix-loop bifurcation.

SIAM Journal on Mathematical Analysis, 18(4):1142–1156, July 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schechter:1988:EEP

[Sch88] Martin Schechter. Exact estimates for potentials. SIAM Journal on Mathematical Analysis, 19 (6):1324–1328, November 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schmeiser:1990:SPA

[Sch90] Christian Schmeiser. A singular perturbation analysis of reverse biased pn-junctions. SIAM Journal on Mathematical Analysis, 21 (2):313–326, March 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schweizer:1997:FBF

[Sch97] Ben Schweizer. Free boundary fluid systems in a semigroup approach and oscillatory behavior. SIAM Journal on Mathematical Analysis, 28(5):1135– 1157, September 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29989.

Scheel:1998:BSW

[Sch98] Arnd Scheel. Bifurcation to spiral waves in reaction-diffusion systems. SIAM Journal on Mathematical Analysis, 29(6):1399– 1418, November 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/31894.

Sethi:1984:GSS

[SDL84] S. P. Sethi, N. A. Derzko, and J. Lehoczky. General solution of the stochastic price-dividend integral equation: a theory of financial valuation. SIAM Journal on Mathematical Analysis, 15(6): 1100–1113, November 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Secchi:1988:MVF

[Sec88] Paolo Secchi. On the motion of viscous fluids in the presence of diffusion. SIAM Journal on Mathematical Analysis, 19 (1):22–31, January 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Seifert:1987:NEE

[Sei87] George Seifert. Nonlinear evolution equations with almost periodic time dependence. SIAM Journal on Mathematical Analysis, 18(2):387–392, March 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Seip:1991:RFD

[Sei91] Kristian Seip. Reproducing formulas and double orthogonality in Bargmann and Bergman spaces. SIAM Journal on Mathematical Analysis, 22(3):856–876, May 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Selgrade:1981:ABS

[Sel81] James F. Selgrade. Asymptotic behavior of solutions to multiple loop positive feedback systems. SIAM Journal on Mathematical Analysis, 12(5):669–678, September 1981. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Senderovizh:1978:CPO

[Sen78] Rivka Senderovizh. On convexity preserving operators. SIAM Journal on Mathematical Analysis, 9(1):157–159, February 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sentis:1985:SCE

[Sen85] Rémi Sentis. Study of the corrector of the eigenvalue of a transport operator. SIAM Journal on Mathematical Analysis, 16(1): 151–166, January 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Senouf:1996:ANA

[Sen96]  David Senouf.  Asymptotic

and numerical approximations of the zeros of Fourier integrals. SIAM Journal on Mathematical Analysis, 27(4):1102–1128, July 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/27221.

Senouf:1997:DCCa

[Sen97a] David Senouf. Dynamics and condensation of complex singularities for Burgers’ equation. I. SIAM Journal on Mathematical Analysis, 28(6):1457– 1489, November 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/28937.

Senouf:1997:DCCb

[Sen97b] David Senouf. Dynamics and condensation of complex singularities for Burgers’ equation. II. SIAM Journal on Mathematical Analysis, 28(6):1490– 1513, November 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/28970.

Serge:1975:AMS

[Ser75]  Vasilach Serge.  Algebraic

method for solving linear differential equations whose coefficients are functions of one variable. SIAM Journal on Mathematical Analysis, 6(2):295–311, April 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sung:1991:IPM

[SF91] Li-Yeng Y. Sung and A. S. Fokas.

Inverse problems in multidimensions. SIAM Journal on Mathematical Analysis, 22(5):1303– 1331, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sox:1971:CCO

[SH71] J. L. Sox, Jr. and W. J. Harrington. A class of complete orthogonal sequences of periodic functions. SIAM Journal on Mathematical Analysis, 2(3):458–466, August 1971. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shapiro:1973:CNN

[Sha73]  Victor L. Shapiro.  Capac-

ity and the nonlinear Navier– Stokes equations. SIAM Journal on Mathematical Analysis, 4(2): 329–343, May 1973. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shapiro:1976:ISS

[Sha76a] Victor L. Shapiro. Isolated singularities in steady state fluid flow. SIAM Journal on Mathematical Analysis, 7(4):577–601, August 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sharma:1976:NCF

[Sha76b] B. L. Sharma. A note on Carlitz’s formula. SIAM Journal on Mathematical Analysis, 7(6):820–823, November 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shaw:1976:SEL

[Sha76c] J. K. Shaw. Series expansions and linear differential operators. SIAM Journal on Mathematical Analysis, 7(3):311–331, May 1976. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Shaw:1977:APG

[Sha77] J. K. Shaw. Analytic properties of generalized completely convex functions. SIAM Journal on Mathematical Analysis, 8 (2):271–279, April 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shail:1980:IRL

[Sha80a] R. Shail. On integral representations for Lamé and other special functions. SIAM Journal on Mathematical Analysis, 11(4): 702–723, July 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shapiro:1980:SPR

[Sha80b] Victor L. Shapiro. Saddle points and Ritz–Galerkin approximations. SIAM Journal on Mathematical Analysis, 11(3):577–595, May 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shankar:1987:FSS

[Sha87] Shiva Shankar. Filippov solutions to singular differential equations. SIAM Journal on Mathematical Analysis, 18(5): 1463–1466, September 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shahgholian:1991:NPH

[Sha91a] Henrik Shahgholian. On the Newtonian potential of a heterogeneous ellipsoid. SIAM Journal on Mathematical Analysis, 22(5): 1246–1255, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shapiro:1991:SFS

[Sha91b] Victor L. Shapiro. Special functions and singular quasilinear partial differential equations. SIAM Journal on Mathematical Analysis, 22(5):1411–1429, September 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shere:1972:MAEa

[She72a] Kenneth D. Shere. Multiple asymptotic expansions and singular problems. SIAM Journal on Mathematical Analysis, 3(2): 263–271, May 1972. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shere:1972:MAEb

[She72b] Kenneth D. Shere. On multiple asymptotic expansions. SIAM Journal on Mathematical Analysis, 3(2):272–284, May 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shearer:1980:SBN

[She80a]  M. Shearer.  Secondary bi-

furcation near a double eigenvalue. SIAM Journal on Mathematical Analysis, 11(2):365–389, March 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sheffer:1980:CPS

[She80b] I. M. Sheffer. On a class W [k, s; C_{ij}] of polynomial sets: Part I. SIAM Journal on Mathematical Analysis, 11(2):228–241, March 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shepherd:1983:CFB

[She83] J. J. Shepherd. Concurrent flow in a bubble-column reactor at high Péclet numbers. SIAM Journal on Mathematical Analysis, 14(1):38–46, January 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shearer:1985:EWS

[She85] Michael Shearer. Elementary wave solutions of the equations describing the motion of an elastic string. SIAM Journal on Mathematical Analysis, 16(3): 447–459, May 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sheu:1986:ABI

[She86] Shuenn Jyi Sheu. Asymptotic behavior of the invariant density of a diffusion Markov process with small diffusion. SIAM Journal on Mathematical Analysis, 17 (2):451–460, March 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shen:1988:NSE

[She88a] Chao Liang Shen. On the nodal sets of the eigenfunctions of the string equation. SIAM Journal on Mathematical Analysis, 19(6): 1419–1424, November 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shen:1988:RSE

[She88b] Chao Liang Shen. Remarks on the second eigenvalue of a symmetric simply connected plane region. SIAM Journal on Mathematical Analysis, 19(1):167–171, January 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shen:1993:NSE

[She93] Chao Liang Shen. On the nodal sets of the eigenfunctions of certain homogeneous and nonhomogeneous membranes. SIAM Journal on Mathematical Analysis, 24 (5):1277–1282, September 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shen:1995:NPW

[She95] Zuowei Shen. Nontensor product wavelet packets in L_{2}(R^{s}). SIAM Journal on Mathematical Analysis, 26(4):1061–1074, July 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/24364.

Shen:1998:RFV

[She98] Zuowei Shen. Refinable function vectors. SIAM Journal on Mathematical Analysis, 29(1):

235–250, January 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30268.

Shinbrot:1974:EEN

[Shi74] Marvin Shinbrot. The energy equation for the Navier–Stokes system. SIAM Journal on Mathematical Analysis, 5(6):948–954, November 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shi:1993:WSE

[Shi93] Peter Shi. Weak solution to an evolution problem with a nonlocal constraint. SIAM Journal on Mathematical Analysis, 24 (1):46–58, January 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Speck:1989:SDPb

[SHM89] F.-O. Speck, R. A. Hurd, and E. Meister. Sommerfeld diffraction problems with third kind boundary conditions II. SIAM Journal on Mathematical Analysis, 20(3):589–607, May 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Showalter:1970:WPP

[Sho70] R. E. Showalter. Well-posed problems for a partial differential equation of order 2m + 1. SIAM Journal on Mathematical Analysis, 1(2):214–231, May 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Showalter:1972:ERT

[Sho72] R. E. Showalter. Existence and representation theorems for a semilinear Sobolev equation in Banach space. SIAM Journal on Mathematical Analysis, 3(3): 527–543, August 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Showalter:1975:NDE

[Sho75] R. E. Showalter. Nonlinear degenerate evolution equations and partial differential equations of mixed type. SIAM Journal on Mathematical Analysis, 6(1):25– 42, February 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Showalter:1976:RAS

[Sho76] R. E. Showalter. Regularization and approximation of second order evolution equations. SIAM Journal on Mathematical Analysis, 7(4):461–472, August 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shultz:1978:OCF

[Shu78] Harris S. Shultz. Operational calculus for functions of two variables. SIAM Journal on Mathematical Analysis, 9(4):660–666, August 1978. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Si:1978:RSF

[Si78] Do Tan Si. Representation of special functions by differintegral and hyperdifferential operators. SIAM Journal on Mathematical Analysis, 9(6):1068–1075, December 1978. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sibuya:1981:TCU

[Sib81] Yasutaka Sibuya. A theorem concerning uniform simplification at a transition point and the problem of resonance. SIAM Journal on Mathematical Analysis, 12(5): 653–668, September 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sidi:1985:AEM

[Sid85] Avram Sidi. Asymptotic expansions of Mellin transforms and analogues of Watson’s lemma. SIAM Journal on Mathematical Analysis, 16(4):896–906, July 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sidi:1986:BSC

[Sid86] Avram Sidi. Borel summability and converging factors for some everywhere divergent series. SIAM Journal on Mathematical Analysis, 17(5):1222– 1231, September 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sigillito:1974:EDF

[Sig74] V. G. Sigillito. Exponential decay of functionals of solutions of a pseudoparabolic equation. SIAM Journal on Mathematical Analysis, 5(4):581–585, August 1974.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sigillito:1976:PID

[Sig76] Vincent G. Sigillito. A priori inequalities and the direchlet problem for a pseudo-parabolic equation. SIAM Journal on Mathematical Analysis, 7(2):222–229, April 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Silverstein:1981:DBE

[Sil81] Jack W. Silverstein. Describing the behavior of eigenvectors of random matrices using sequences of measures on orthogonal groups. SIAM Journal on Mathematical Analysis, 12 (2):274–281, March 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Silverstein:1985:LED

[Sil85] Jack W. Silverstein. The limiting eigenvalue distribution of a multivariate F matrix. SIAM Journal on Mathematical Analysis, 16 (3):641–646, May 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Simar:1977:NSM

[Sim77] Leopold Simar. A note on Stieltjes moment sequences. SIAM Journal on Mathematical Analysis, 8(3):533–534, May 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Simons:1980:UPE

[Sim80] S. Simons. The use of polynomials in evaluating certain probabilistic integrals. SIAM Journal on Mathematical Analysis, 11(5): 771–786, September 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Simsa:1984:AIP

[Šim84] Jaromír Šimša. Asymptotic integration of perturbed linear differential equations under conditions involving ordinary integral convergence. SIAM Journal on Mathematical Analysis, 15 (1):116–123, January 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Simsa:1985:COI

[Šim85] Jaromír Šimša. The condition of ordinary integral convergence in the asymptotic theory of linear differential equations with almost constant coefficients. SIAM Journal on Mathematical Analysis, 16(4):757–769, July 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Simsa:1988:ETP

[Šim88] Jaromír Šimša. An extension of a theorem of Perron. SIAM Journal on Mathematical Analysis, 19 (2):460–472, March 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Simon:1990:NVI

[Sim90] Jacques Simon. Nonhomogeneous viscous incompressible fluids: Existence of velocity, density, and pressure. SIAM Journal on Mathematical Analysis, 21(5): 1093–1117, September 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Singh:1975:ANN

[Sin75] Bhagat Singh. Asymptotic nature of nonoscillatory solutions of n th order retarded differential equations. SIAM Journal on Mathematical Analysis, 6(5): 784–795, October 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Singh:1976:AVO

[Sin76] Bhagat Singh. Asymptotically vanishing oscillatory trajectories in second order retarded equations. SIAM Journal on Mathematical Analysis, 7(1):37–44, February 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Singh:1979:NSC

[Sin79] Bhagat Singh. Necessary and sufficient condition for maintaining oscillations and nonoscillations in general functional equations and their asymptotic properties. SIAM Journal on Mathematical Analysis, 10(1):18–31, January 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Singal:1980:SND

[Sin80] R. P. Singal. A short note on a differential recursion formula for Appell’s hypergeometric function F_{3}. SIAM Journal on Mathematical Analysis, 11(2):390–391, March 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sinestrari:1997:RPI

[Sin97]  Carlo Sinestrari.  The Rie-

mann problem for an inhomogeneous conservation law without convexity. SIAM Journal on Mathematical Analysis, 28(1): 109–135, January 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/27446.

Sivazlian:1971:CMI

[Siv71] B. D. Sivazlian. A class of multiple integrals. SIAM Journal on Mathematical Analysis, 2(1):72– 75, February 1971. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shamma:1973:AES

[SK73] S. E. Shamma and S. N. Karp.

Asymptotic eigenfunctions of a singular integro-differential equation. SIAM Journal on Mathematical Analysis, 4(1):116–127, February 1973. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sprinkhuizen-Kuyper:1976:OPT

[SK76] I. G. Sprinkhuizen-Kuyper. Orthogonal polynomials in two variables. A further analysis of the polynomials orthogonal over a region bounded by two lines and a parabola. SIAM Journal on Mathematical Analysis, 7 (4):501–518, August 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shih:1987:AAS

[SK87] Shagi-Di Shih and R. Bruce Kellogg. Asymptotic analysis of a singular perturbation problem. SIAM Journal on Mathematical Analysis, 18(5):1467– 1511, September 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Skinner:1980:NAB

[Ski80a] L. A. Skinner. Note on the asymptotic behavior of multidimensional Laplace integrals. SIAM Journal on Mathematical Analysis, 11(5):911–917, September 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Skinner:1980:UVE

[Ski80b] L. A. Skinner. Uniformly valid expansions for Laplace integrals. SIAM Journal on Mathematical Analysis, 11(6):1058–1067, November 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [Ski81].

Skinner:1981:EUV

[Ski81] L. A. Skinner. Erratum: “Uniformly valid expansions for Laplace integrals”. SIAM Journal on Mathematical Analysis, 12(3):487, May 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [Ski80b].

Skinner:1988:UVC

[Ski88] Lindsay A. Skinner. Uniformly valid composite expansions for Laplace integrals. SIAM Journal on Mathematical Analysis, 19(4): 918–925, July 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Skinner:1994:MES

[Ski94] L. A. Skinner. Matched expansion solutions of the first-order turning point problem. SIAM Journal on Mathematical Analysis, 25(5):1402–1411, September 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/23996.

Slepian:1978:NEH

[SL78] David Slepian and Henry J. Landau. A note on the eigenvalues of Hermitian matrices. SIAM Journal on Mathematical Analysis, 9 (2):291–297, April 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Slepian:1972:SKP

[Sle72] David Slepian. On the symmetrized Kronecker power of a matrix and extensions of Mehler’s formula for Hermite polynomials. SIAM Journal on Mathematical Analysis, 3(4): 606–616, November 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Slemrod:1995:CTV

[Sle95] M. Slemrod. A comparison of two viscous regularizations of the Riemann problem for Burgers’s equation. SIAM Journal on Mathematical Analysis, 26(6): 1415–1424, November 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/25368.

Sloss:1975:NIE

[Slo75] James M. Sloss. A new integral equation for certain plane Dirichlet problems. SIAM Journal on Mathematical Analysis, 6(6): 998–1006, November 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shen:1981:PTF

[SLS81] M. C. Shen, K. C. Lin, and S. M. Shih. Peristaltic transport of a fluid-particle mixture. SIAM Journal on Mathematical Analysis, 12(1):49–59, January 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Scheurle:1984:BQP

[SM84] Jürgen Scheurle and Jerrold Marsden. Bifurcation to quasiperiodic tori in the interaction of steady state and Hopf bifurcations. SIAM Journal on Mathematical Analysis, 15(6):

1055–1074, November 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Soderlind:1985:SAE

[SM85] Gustaf Söderlind and Robert M. M. Mattheij. Stability and asymptotic estimates in nonautonomous linear differential systems. SIAM Journal on Mathematical Analysis, 16(1):69–92, January 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Smith:1971:AAC

[Smi71] Donald R. Smith. An asymptotic analysis of a certain hyperbolic Cauchy problem. SIAM Journal on Mathematical Analysis, 2 (3):375–392, August 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Smith:1975:NBV

[Smi75] Donald R. Smith. A nonlinear bboundary value problem on an unbounded interval. SIAM Journal on Mathematical Analysis, 6 (4):601–615, August 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Smillie:1984:CCT

[Smi84] John Smillie. Competitive and cooperative tridiagonal systems of differential equations. SIAM Journal on Mathematical Analysis, 15(3):530–534, May 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Smith:1986:ICM

[Smi86a] Hal L. Smith. Invariant curves for mappings. SIAM Journal on Mathematical Analysis, 17(5): 1053–1067, September 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Smith:1986:PSP

[Smi86b] Hal L. Smith. Periodic solutions of periodic competitive and cooperative systems. SIAM Journal on Mathematical Analysis, 17(6): 1289–1318, November 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Smith:1989:HQP

[Smi89] Richard A. Smith. On a hyperbolic quenching problem in several dimensions. SIAM Journal on Mathematical Analysis, 20(5): 1081–1094, September 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Smith:1991:PTC

[Smi91a] Hal L. Smith. Periodic tridiagonal competitive and cooperative systems of differential equations. SIAM Journal on Mathematical Analysis, 22(4):1102– 1109, July 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Smith:1991:AMW

[Smi91b] William V. Smith. The analysis of a model for wave motion in a liquid semiconductor: Boundary interaction and variable conductivity. SIAM Journal on Mathematical Analysis, 22 (2):352–378, March 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Smyth:1979:CLS

[Smy79] W. F. Smyth. Constant limit of a sequence of iterates. SIAM Journal on Mathematical Analysis, 10 (3):463–471, May 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Saffren:1971:RAS

[SN71] M. M. Saffren and E. W. Ng. Recursive algorithms for the summation of certain series. SIAM Journal on Mathematical Analysis, 2(1):31–36, February 1971. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Specovius-Neugebauer:1999:ASD

[SN99] Maria Specovius-Neugebauer.

Approximation of the Stokes Dirichlet problem in domains with cylindrical outlets. SIAM Journal on Mathematical Analysis, 30(3):645–677, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32508.

Sommer:1980:CCS

[Som80] Manfred Sommer. Characterization of continuous selections for the metric projection for generalized splines. SIAM Journal on Mathematical Analysis, 11 (1):23–40, January 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sommer:1982:CCS

[Som82] Manfred Sommer. Characterization of continuous selections of the metric projection for a class of weak Chebyshev spaces. SIAM Journal on Mathematical Analysis, 13(2):280–294, March 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Soni:1970:UTR

[Son70] Kusum Soni. A unitary transform related to some integral equations. SIAM Journal on Mathematical Analysis, 1(4): 426–436, November 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Soni:1980:EET

[Son80] Kusum Soni. Exact error terms in the asymptotic expansion of a class of integral transforms. I. Oscillatory kernels. SIAM Journal on Mathematical Analysis, 11(5):828–841, September 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Soni:1983:NUA

[Son83] K. Soni. A note on uniform asymptotic expansion of incomplete Laplace integrals. SIAM Journal on Mathematical Analysis, 14(5):1015–1018, September 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Souplet:1998:BNR

[Sou98] Philippe Souplet. Blow-up in nonlocal reaction-diffusion equations. SIAM Journal on Mathematical Analysis, 29(6):1301– 1334, November 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/31890.

Siewert:1979:STE

[SP79] C. E. Siewert and J. S. Phelps, III. On solutions of a transcendental equation basic to the theory of vibrating plates. SIAM Journal on Mathematical Analysis, 10(1):105–111, January 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sperb:1979:CAM

[Spe79] René P. Sperb. Corrigendum and addendum: “Maximum principles and bounds in some inhomogeneous elliptic boundary problems” [SIAM J. Math. Anal. 8 (1977), no. 5, 871–878; MR 58 #11889] by P. W. Schaefer and R. P. Sperb. SIAM Journal on Mathematical Analysis, 10 (6):1326, November 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [SS77].

Speck:1989:SDPa

[Spe89] F.-O. Speck. Sommerfeld diffraction problems with first and second kind boundary conditions. SIAM Journal on Mathematical Analysis, 20(2):396–407, March 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Spikes:1979:CLC

[Spi79] Paul W. Spikes. Criteria of limit circle type for nonlinear differential equations. SIAM Journal on Mathematical Analysis, 10(3): 456–462, May 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schulz-Rinne:1993:CRP

[SR93] Carsten W. Schulz-Rinne. Classification of the Riemann problem for two-dimensional gas dynamics. SIAM Journal on Mathematical Analysis, 24(1):76–88, January 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Srivastava:1970:CRI

[Sri70] H. M. Srivastava. Certain results involving generalized hypergeometric functions. SIAM Journal on Mathematical Analysis, 1 (1):75–81, February 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Srivastav:1972:DIE

[Sri72] R. P. Srivastav. Dual integral equations with trigonometric kernels and tempered distributions. SIAM Journal on Mathematical Analysis, 3(3):413–421, August 1972. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Srivastava:1983:NMG

[Sri83] H. M. Srivastava. A note on a multilinear generating function for the Konhauser biorthogonal polynomials. SIAM Journal on Mathematical Analysis, 14 (2):369–371, March 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

SriRanga:1990:RFA

[Sri90] A. Sri Ranga. On a recurrence formula associated with strong distributions. SIAM Journal on Mathematical Analysis, 21(5): 1335–1348, September 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Soni:1973:ABC

[SS73a] K. Soni and R. P. Soni. Asymptotic behavior of a class of integral transforms. SIAM Journal on Mathematical Analysis, 4(3): 466–481, May 1973. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Soni:1973:LBC

[SS73b] K. Soni and R. P. Soni. Lipschitz behavior and characteristic functions. SIAM Journal on Mathematical Analysis, 4(2): 302–308, May 1973. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shih:1975:UAA

[SS75] S. M. Shih and M. C. Shen. Uniform asymptotic approximation for viscous fluid flow down an inclined plane. SIAM Journal on Mathematical Analysis, 6(3): 560–582, May 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schaefer:1977:MPB

[SS77] P. W. Schaefer and R. P. Sperb. Maximum principles and bounds in some inhomogeneous elliptic boundary value problems. SIAM Journal on Mathematical Analysis, 8(5):871–878, October 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [Spe79].

Sficas:1978:OAB

[SS78] Y. G. Sficas and I. P. Stavroulakis. On the oscillatory and asymptotic behavior of a class of differential equations with deviating arguments. SIAM Journal on Mathematical Analysis, 9(5): 956–966, October 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schafke:1980:CPG

[SS80] Reinhard Schäfke and Dieter Schmidt. The connection problem for general linear ordinary differential equations at two regular singular points with applications in the theory of special functions. SIAM Journal on Mathematical Analysis, 11(5): 848–862, September 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Saxe:1984:HFT

[SS84] Michael B. Saxe and Thomas O. Sherman. Hilbert and Fourier transforms on a sphere. SIAM Journal on Mathematical Analysis, 15(3):605–620, May 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Soni:1985:NSA

[SS85] K. Soni and R. P. Soni. A note on summability and asymptotics. SIAM Journal on Mathematical Analysis, 16(2):392–404, March 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Scherer:1989:EBP

[SS89] K. Scherer and P. W. Smith. Existence of best parametric interpolation by curves. SIAM Journal on Mathematical Analysis, 20 (1):160–168, January 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shen:1992:SPF

[SS92] Chao Liang Shen and Chung Tsun Shieh. Some properties of the first eigenvalue of the Laplace operator on the spherical bands in S^{2}. SIAM Journal on Mathematical Analysis, 23(5):1305–1308, September 1992. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shearer:1993:IVP

[SS93a] Michael Shearer and David G.

Schaeffer. The initial value problem for a system modelling unidirectional longitudinal elasticplastic waves. SIAM Journal on Mathematical Analysis, 24(5):

1111–1144, September 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sun:1993:SWT

[SS93b] S. M. Sun and M. C. Shen. Solitary waves in a two-layer fluid with surface tension. SIAM Journal on Mathematical Analysis, 24 (4):866–891, July 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Simion:1994:SGL

[SS94] R. Simion and D. Stanton. Specializations of generalized Laguerre polynomials. SIAM Journal on Mathematical Analysis, 25 (2):712–719, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/22854.

Stojanovic:1995:FBP

[SS95] Srdjan Stojanovic and Thomas Svobodny. Free boundary problems for potential and Stokes flows via nonsmooth analysis. SIAM Journal on Mathematical Analysis, 26(3):633–658, May 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/24652.

Soravia:1996:PFT

[SS96] Pierpaolo Soravia and Panagiotis E. Souganidis. Phase-field theory for Fitzhugh–Nagumotype systems. SIAM Journal on Mathematical Analysis, 27(5):

1341–1359, September 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/26131.

Showalter:1970:PPD

[ST70] R. E. Showalter and T. W. Ting.

Pseudoparabolic partial differential equations. SIAM Journal on Mathematical Analysis, 1(1):1– 26, February 1970. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sharma:1976:CCT

[ST76] A. Sharma and J. Tzimbalario.

A class of cardinal trigonometric splines. SIAM Journal on Mathematical Analysis, 7(6):809–819, November 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

StMary:1984:CTH

[St.84] D. F. St. Mary. On a comparison theorem of Hille for selfadjoint second order linear differential systems. SIAM Journal on Mathematical Analysis, 15(6): 1094–1099, November 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

StVincent:1988:ELC

[St.88] Michael St. Vincent. Entrainment of a limit-cycle oscillator with shear by large amplitude forcing. SIAM Journal on Mathematical Analysis, 19(3): 648–666, May 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Saff:1990:ZCP

[ST90a] E. B. Saff and V. Totik. Zeros of Chebyshev polynomials associated with a compact set in the plane. SIAM Journal on Mathematical Analysis, 21(3):799–802, May 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Smith:1990:QCS

[ST90b] Hal L. Smith and Horst R. Thieme. Quasi convergence and stability for strongly orderpreserving semiflows. SIAM Journal on Mathematical Analysis, 21(3):673–692, May 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Smith:1991:CSO

[ST91] Hal L. Smith and Horst R. Thieme. Convergence for strongly order-preserving semiflows. SIAM Journal on Mathematical Analysis, 22(4):1081– 1101, July 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

StVincent:1996:BFP

[St.96] M. St. Vincent. Bifurcation of fixed points in coupled Josephson junctions. SIAM Journal on Mathematical Analysis, 27 (4):1084–1094, July 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/23834.

Stallmann:1971:EUT

[Sta71] Friedmann W. Stallmann. An existence and uniqueness theorem for boundary value problems of nonlinear ordinary differential equations. SIAM Journal on Mathematical Analysis, 2(2): 277–285, May 1971. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Staffans:1978:APP

[Sta78a] Olof J. Staffans. An asymptotic problem for a positive definite operator-valued Volterra kernel. SIAM Journal on Mathematical Analysis, 9(5):855–866, October 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Staffans:1978:ABF

[Sta78b] Olof J. Staffans. On the asymptotic behavior of finite energy solutions of an abstract integral equation. SIAM Journal on Mathematical Analysis, 9(5): 867–875, October 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Staffans:1979:ASC

[Sta79] Olof J. Staffans. On the asymptotic spectrum of a convolution. SIAM Journal on Mathematical Analysis, 10(6):1138–1143, November 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Staffans:1980:NHV

[Sta80a] Olof J. Staffans. On a nonlinear hyperbolic Volterra equation. SIAM Journal on Mathematical Analysis, 11(5):793–812, September 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stanton:1980:PFH

[Sta80b] Dennis Stanton. Product formulas for q-Hahn polynomials. SIAM Journal on Mathematical Analysis, 11(1):100–107, January 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Staffans:1985:EIF

[Sta85] Olof J. Staffans. Extended initial and forcing function semigroups generated by a functional equation. SIAM Journal on Mathematical Analysis, 16(5):1034– 1048, September 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Staffans:1986:SGN

[Sta86a] Olof J. Staffans. Semigroups generated by a neutral functionaldifferential equation. SIAM Journal on Mathematical Analysis, 17(1):46–57, January 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stanton:1986:SVM

[Sta86b] Dennis Stanton. Sign variations of the Macdonald identities. SIAM Journal on Mathematical Analysis, 17(6):1454– 1460, November 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Staffans:1988:DLA

[Sta88]  Olof J. Staffans.  A direct

Lyapunov approach to Volterra integro-differential equations. SIAM Journal on Mathematical Analysis, 19(4):879–901, July 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Steinig:1970:RZS

[Ste70a] J. Steinig. The real zeros of Struve’s function. SIAM Journal on Mathematical Analysis, 1 (3):365–375, August 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stenger:1970:AAC

[Ste70b] Frank Stenger. The asymptotic approximation of certain integrals. SIAM Journal on Mathematical Analysis, 1(3):392–404, August 1970. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stenger:1972:TMO

[Ste72] Frank Stenger. Transform methods for obtaining asymptotic expansions of definite integrals. SIAM Journal on Mathematical Analysis, 3(1):20–30, February 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stenger:1973:ASC

[Ste73a] Frank Stenger. The approximate solution of convolution-type integral equations. SIAM Journal on Mathematical Analysis, 4(3):

536–555, May 1973. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stephen:1973:WHO

[Ste73b] G. Gregory Stephen. On Wiener– Hopf operators associated with the Hankel transform. SIAM Journal on Mathematical Analysis, 4(2):221–237, May 1973. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stecher:1975:ION

[Ste75] Michael Stecher. Integral operators and the noncharacteristic Cauchy problem for parabolic equations. SIAM Journal on Mathematical Analysis, 6(5): 796–811, October 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stengle:1977:AEA

[Ste77] Gilbert Stengle. Asymptotic estimates for the adiabatic invariance of a simple oscillator. SIAM Journal on Mathematical Analysis, 8(4):640–654, August 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stech:1980:PSN

[Ste80a]  Harlan W. Stech.  Peri-

odic solutions to a nonlinear Volterra integro-differential equation. SIAM Journal on Mathematical Analysis, 11(3): 533–544, May 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stevens:1980:PS

[Ste80b] Robert R. Stevens. Periodic solutions of x′′ + g(t, x) = 0. SIAM Journal on Mathematical Analysis, 11(2):400–403, March 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Steinberg:1984:FPE

[Ste84] Stanly Steinberg. Factored product expansions of solutions of nonlinear differential equations. SIAM Journal on Mathematical Analysis, 15(1):108–115, January 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stech:1985:NHB

[Ste85a] Harlan W. Stech. Nongeneric Hopf bifurcations in functionaldifferential equations. SIAM Journal on Mathematical Analysis, 16(6):1134–1151, November 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stengle:1985:TEA

[Ste85b] Gilbert Stengle. Transcendental estimates for the adiabatic variation of linear Hamiltonian systems. SIAM Journal on Mathematical Analysis, 16(5):932–940, September 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Steinruck:1991:ODW

[Ste91]  H. Steinrück.  The one-

dimensional Wigner–Poisson problem and its relation to the Schrödinger–Poisson problem. SIAM Journal on Mathematical Analysis, 22(4):957–972, July 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stolarsky:1980:SLI

[Sto80] Kenneth B. Stolarsky. A stronger logarithmic inequality suggested by the entropy inequality. SIAM Journal on Mathematical Analysis, 11(2):242–247, March 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stokman:1997:MBL

[Sto97] Jasper V. Stokman. Multivariate big and little q-Jacobi polynomials. SIAM Journal on Mathematical Analysis, 28 (2):452–480, March 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/28719.

Straughan:1976:GUT

[Str76] Brian Straughan. Growth and uniqueness theorems for an abstract nonstandard wave equation. SIAM Journal on Mathematical Analysis, 7(4):519–528, August 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Struble:1977:LTO

[Str77a] Raimond A. Struble. Linear transformations in the operational calculus. SIAM Journal on Mathematical Analysis, 8 (2):258–270, April 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Struble:1977:NVF

[Str77b]  Raimond A. Struble.  The

numerical-valued Fourier transform in the two-sided operational calculus. SIAM Journal on Mathematical Analysis, 8 (2):243–257, April 1977. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Struble:1979:OR

[Str79] Raimond A. Struble. Operational rules. SIAM Journal on Mathematical Analysis, 10(3): 629–642, May 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Straughan:1980:CAL

[Str80] Brian Straughan. Consequences of analyticity in linear elastostatics and related systems. SIAM Journal on Mathematical Analysis, 11(1):181–189, January 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stuart:1989:BHO

[Stu89] C. A. Stuart. Bifurcation of homoclinic orbits and bifurcation from the essential spectrum. SIAM Journal on Mathematical Analysis, 20(5):1145– 1171, September 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Stuart:1996:MPS

[Stu96] D. Stuart. Minimal periods for solutions of some classical field equations. SIAM Journal on Mathematical Analysis, 27 (4):1095–1101, July 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/27418.

Shub:1988:NGC

[STW88] Michael Shub, David Tischler, and Robert F. Williams. The Newtonian graph of a complex polynomial. SIAM Journal on Mathematical Analysis, 19(1): 246–256, January 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shen:1999:ZRL

[STW99] Wen Shen, Aslak Tveito, and Ragnar Winther. On the zero relaxation limit for a system modeling the motions of a viscoelastic solid. SIAM Journal on Mathematical Analysis, 30(5):1115–1135, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/31984.

Stefanov:1997:IBA

[SU97] Plamen Stefanov and Gunther Uhlmann. Inverse backscattering for the acoustic equation. SIAM Journal on Mathematical Analysis, 28(5):1191– 1204, September 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30185.

Sun:1993:TDS

[Sun93] Tien Yu Sun. Three-dimensional steady water waves generated by partially localized pressure disturbances. SIAM Journal on Mathematical Analysis, 24(5): 1153–1178, September 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Suzuki:1995:BSQ

[Suz95] Ryuichi Suzuki. Blow-up solutions of quasi-linear degenerate parabolic equations with convection. SIAM Journal on Mathematical Analysis, 26(1): 77–97, January 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23642.

Saff:1976:ZFP

[SV76] E. B. Saff and R. S. Varga.

Zero-free parabolic regions for sequences of polynomials. SIAM Journal on Mathematical Analysis, 7(3):344–357, May 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sharma:1983:MCI

[SV83] A. Sharma and P. Vértesi. Mean convergence and interpolation in roots of unity. SIAM Journal on Mathematical Analysis, 14(4): 800–806, July 1983. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Spigler:1994:DCL

[SV94] Renato Spigler and Marco Vianello. Discrete and continuous Liouville–Green–Olver approximations: A unified treatment via Volterra–Stieltjes integral equations. SIAM Journal on Mathematical Analysis, 25 (2):720–732, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23121.

Schumitzky:1973:ICL

[SW73] Alan Schumitzky and Tom Wenska. Imbedding a class of linear integral equations through the first critical point. SIAM Journal on Mathematical Analysis, 4(4): 592–608, November 1973. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schumitzky:1975:ORT

[SW75] Alan Schumitzky and Tom Wenska. An operator residue theorem with applications to branching processes and renewal type integral equations. SIAM Journal on Mathematical Analysis, 6 (2):229–235, April 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schmidt:1979:MGI

[SW79a] Dieter Schmidt and Gerhard Wolf. A method of generating integral relations by the simultaneous separability of generalized Schrödinger equations. SIAM Journal on Mathematical Analysis, 10(4):823–838, July 1979.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shivakumar:1979:AEM

[SW79b] P. N. Shivakumar and R. Wong.

Asymptotic expansion of multiple Fourier transforms. SIAM Journal on Mathematical Analysis, 10(6):1095–1104, November 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Schmeiser:1986:AAS

[SW86] Christian Schmeiser and Richard Weiss. Asymptotic analysis of singular singularly perturbed boundary value problems. SIAM Journal on Mathematical Analysis, 17(3):560–579, May 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Showalter:1991:DFF

[SW91] R. E. Showalter and N. J. Walkington. Diffusion of fluid in a fissured medium with microstructure. SIAM Journal on Mathematical Analysis, 22(6):1702– 1722, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Swartz:1972:CLF

[Swa72] Charles Swartz. Continuous linear functionals on certain K{M_{p}} spaces. SIAM Journal on Mathematical Analysis, 3(4): 595–598, November 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Swanson:1978:WI

[Swa78] C. A. Swanson. Wirtinger’s inequality. SIAM Journal on Mathematical Analysis, 9(3):484–491, June 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Swanson:1982:BPS

[Swa82]  C. A. Swanson.  Bounded

positive solutions of semilinear Schrödinger equations. SIAM Journal on Mathematical Analysis, 13(1):40–47, January 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sweers:1989:SMP

[Swe89] Guido Sweers. A strong maximum principle for a noncooperative elliptic system. SIAM Journal on Mathematical Analysis, 20 (2):367–371, March 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sweldens:1998:LSC

[Swe98] Wim Sweldens. The lifting scheme: A construction of second generation wavelets. SIAM Journal on Mathematical Analysis, 29 (2):511–546, March 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/28905.

Swick:1976:MSS

[Swi76] K. E. Swick. A model of single species population growth. SIAM Journal on Mathematical Analysis, 7(4):565–576, August 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Swick:1980:PSN

[Swi80] K. E. Swick. Periodic solutions of a nonlinear age-dependent model of single species population dynamics. SIAM Journal on Mathematical Analysis, 11(5): 901–910, September 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sakurai:1989:FSA

[SY89] Kouichi Sakurai and Masaaki Yoshida. Fuchsian systems associated with the P^{2}(F_{2})- arrangement. SIAM Journal on Mathematical Analysis, 20(6): 1490–1499, November 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Symes:1981:SSI

[Sym81] W. W. Symes. Stable solution of the inverse reflection problem for a smoothly stratified elastic medium. SIAM Journal on Mathematical Analysis, 12(3): 421–453, May 1981. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Symes:1986:LSI

[Sym86a] W. W. Symes. Linearization stability for an inverse problem in several-dimensional wave propagation. SIAM Journal on Mathematical Analysis, 17(1):132–151, January 1986. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Symes:1986:RBC

[Sym86b] William W. Symes. On the relation between coefficient and boundary values for solutions of Webster’s horn equation. SIAM Journal on Mathematical Analysis, 17(6):1400–1420, November 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Symes:1991:LVI

[Sym91] William W. Symes. Layered velocity inversion: A model problem from reflection seismology. SIAM Journal on Mathematical Analysis, 22(3):680–716, May 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sokolowski:1987:SSA

[SZ87] Jan Sokołowski and Jean-Paul Zolésio. Shape sensitivity analysis of unilateral problems. SIAM Journal on Mathematical Analysis, 18(5):1416–1437, September 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Shi:1997:NTP

[SZL97] Yongdong Shi, Qinde Zhou, and Yong Li. A note on a two-point boundary value problem arising from a liquid metal flow. SIAM Journal on Mathematical Analysis, 28(5):1086–1093, September 1997. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/29025.

Szwarc:1992:OPDa

[Szw92a] Ryszard Szwarc. Orthogonal polynomials and a discrete boundary value problem I. SIAM Journal on Mathematical Analysis, 23(4):959–964, July 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Szwarc:1992:OPDb

[Szw92b] Ryszard Szwarc. Orthogonal polynomials and a discrete boundary value problem II. SIAM Journal on Mathematical Analysis, 23(4):965–969, July 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Sprekels:1998:ABS

[SZZ98] Jürgen Sprekels, Songmu Zheng, and Peicheng Zhu. Asymptotic behavior of the solutions to a Landau–Ginzburg system with viscosity for Martensitic phase transitions in shape memory alloys. SIAM Journal on Mathematical Analysis, 29(1): 69–84, January 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29769.

Tadmor:1986:WPK

[Tad86]  Eitan Tadmor.  The well-

posedness of the Kuramoto– Sivashinsky equation. SIAM Journal on Mathematical Analysis, 17(4):884–893, July 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tahraoui:1990:CFN

[Tah90] Rabah Tahraoui. Sur une classe de fonctionnelles non convexes et applications. (French) [On a class of nonconvex functionals, and applications]. SIAM Journal on Mathematical Analysis, 21 (1):37–52, January 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [Tah91].

Tahraoui:1991:ECF

[Tah91]  Rabah Tahraoui.  Erratum:

“Sur une classe de fonctionelles non convexes et applications” (French) [On a class of nonconvex functionals and applications] [SIAM J. Math. Anal. 21 (1990), no. 1, 37–52]. SIAM Journal on Mathematical Analysis, 22(3):877, May 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [Tah90].

Takac:1992:LTB

[Tak92] Peter Takáč. Large-time behavior of a time-periodic cooperative system of reaction-diffusion equations depending on parameters. SIAM Journal on Mathematical Analysis, 23(2):387–411, March 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Taliaferro:1980:AVS

[Tal80] Steven D. Taliaferro. On the angular variation of solutions of second order linear systems. SIAM Journal on Mathematical Analysis, 11(1):173–180, January 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Taliaferro:1981:ABS

[Tal81] Steven D. Taliaferro. Asymptotic behavior of solutions of y′′ = φ(t)f(y). SIAM Journal on Mathematical Analysis, 12(6): 853–865, November 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Takac:1996:AEB

[TBD^{+}96] P. Takáč, P. Bollerman, A. Doelman, A. van Harten, and E. S. Titi. Analyticity of essentially bounded solutions to semilinear parabolic systems and validity of the Ginzburg–Landau equation. SIAM Journal on Mathematical Analysis, 27(2): 424–448, March 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/26251.

Tchamitchian:1996:ICO

[Tch96] Philippe Tchamitchian. Inversion de certains opérateurs elliptiques à coefficients variables. (French) [Inversion of certain elliptic operators with variable coefficients]. SIAM Journal on Mathematical Analysis, 27(6):

1680–1703, November 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/26806.

Tefteller:1978:NSO

[Tef78] S. C. Tefteller. Nonhomogeneous second order differential systems with integral boundary conditions. SIAM Journal on Mathematical Analysis, 9(4):583–590, August 1978. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Temme:1974:AMS

[Tem74] N. M. Temme. Analytical methods for a singular perturbation problem in a sector. SIAM Journal on Mathematical Analysis, 5(6):876–887, November 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Temme:1976:RPE

[Tem76] N. M. Temme. Remarks on a paper of A. erdélyi. SIAM Journal on Mathematical Analysis, 7 (5):767–770, October 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Temme:1979:AEI

[Tem79] N. M. Temme. The asymptotic expansion of the incomplete gamma functions. SIAM Journal on Mathematical Analysis, 10(4): 757–766, July 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Temme:1982:UAE

[Tem82] N. M. Temme. The uniform asymptotic expansion of a class of integrals related to cumulative distribution functions. SIAM Journal on Mathematical Analysis, 13(2):239–253, March 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Temme:1987:ILI

[Tem87]  N. M. Temme.  Incomplete

Laplace integrals: uniform asymptotic expansion with application to the incomplete beta function. SIAM Journal on Mathematical Analysis, 18(6): 1638–1663, November 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Temam:1990:IMM

[Tem90a] R. Temam. Inertial manifolds and multigrid methods. SIAM Journal on Mathematical Analysis, 21(1):154–178, January 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Temme:1990:UAE

[Tem90b] N. M. Temme. Uniform asymptotic expansions of a class of integrals in terms of modified Bessel functions with application to confluent hypergeometric functions. SIAM Journal on Mathematical Analysis, 21(1): 241–261, January 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tepper:1974:FBP

[Tep74] David E. Tepper. Free boundary problem. SIAM Journal on Mathematical Analysis, 5(5): 841–846, October 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tepper:1975:FBP

[Tep75] David E. Tepper. On a free boundary problem, the starlike case. SIAM Journal on Mathematical Analysis, 6(3):503–505, May 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Terman:1983:FBP

[Ter83] David Terman. A free boundary problem arising from a bistable reaction-diffusion equation. SIAM Journal on Mathematical Analysis, 14(6):1107– 1129, November 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Terras:1985:SFS

[Ter85] Audrey Terras. Special functions for the symmetric space of positive matrices. SIAM Journal on Mathematical Analysis, 16(3): 620–640, May 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Terman:1988:TWS

[Ter88] David Terman. Traveling wave solutions arising from a two-step combustion model. SIAM Journal on Mathematical Analysis, 19 (5):1057–1080, September 1988.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Terman:1990:SPW

[Ter90] David Terman. Stability of planar wave solutions to a combustion model. SIAM Journal on Mathematical Analysis, 21(5): 1139–1171, September 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Terracina:1999:FBP

[Ter99] Andrea Terracina. A free boundary problem for scalar conservation laws. SIAM Journal on Mathematical Analysis, 30 (5):985–1009, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32530.

Tang:1975:AES

[TF75] Min Ming Tang and Paul C.

Fife. Asymptotic expansions of solutions of quasi-linear elliptic equations with small parameter. SIAM Journal on Mathematical Analysis, 6(3):523–532, May 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tremblay:1984:EOR

[TF84] R. Tremblay and B. J. Fugère.

Expansions of operators related to xD and the fractional derivative. SIAM Journal on Mathematical Analysis, 15(6):1214– 1219, November 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Trefethen:1985:CDR

[TG85] Lloyd N. Trefethen and Martin H. Gutknecht. On convergence and degeneracy in rational Padé and Chebyshev approximation. SIAM Journal on Mathematical Analysis, 16(1):198–210, January 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Thach:1993:GOC

[Tha93] Phan Thien Thach. Global optimality criterion and a duality with a zero gap in nonconvex optimization. SIAM Journal on Mathematical Analysis, 24(6): 1537–1556, November 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Thieme:1993:PUR

[Thi93] Horst R. Thieme. Persistence under relaxed point-dissipativity (with application to an endemic model). SIAM Journal on Mathematical Analysis, 24 (2):407–435, March 1993. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Thompson:1977:IPS

[Tho77] Russell C. Thompson. An invariance property of solutions to second order differential inequalities in ordered Banach spaces. SIAM Journal on Mathematical Analysis, 8(4):592–603, August 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Thron:1994:TEL

[Thr94] W. J. Thron. Truncation error for limit periodic Schur algorithms. SIAM Journal on Mathematical Analysis, 25(2): 733–748, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/22201.

Tippett:1974:EUT

[Tip74] James Tippett. An existenceuniqueness theorem for two-point boundary value problems. SIAM Journal on Mathematical Analysis, 5(1):153–157, February 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tang:1993:SNC

[TKS93] Bao Rong Tang, Yang Kuang, and Hal Smith. Strictly nonautonomous cooperative system with a first integral. SIAM Journal on Mathematical Analysis, 24 (5):1331–1339, September 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Taniguchi:1994:IPI

[TN94] Masaharu Taniguchi and Yasumasa Nishiura. Instability of planar interfaces in reactiondiffusion systems. SIAM Journal on Mathematical Analysis, 25 (1):99–134, January 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/23350.

Tolmasky:1998:EGS

[Tol98] Carlos F. Tolmasky. Exponentially growing solutions for nonsmooth first-order perturbations of the laplacian. SIAM Journal on Mathematical Analysis, 29 (1):116–133, January 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/30103.

Tomastik:1981:CFP

[Tom81] E. C. Tomastik. Conjugate and focal points of second-order differential systems. SIAM Journal on Mathematical Analysis, 12(3): 314–320, May 1981. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tomastik:1983:CTS

[Tom83] E. C. Tomastik. Comparison theorems for second order nonselfadjoint differential systems. SIAM Journal on Mathematical Analysis, 14(1):60–65, January 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tomastik:1986:CTC

[Tom86] E. C. Tomastik. Comparison theorems for certain differential systems of arbitrary order. SIAM Journal on Mathematical Analysis, 17(1):30–37, January 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ton:1981:EUC

[Ton81] Bui An Ton. Existence and uniqueness of a classical solution of an initial-boundary value problem of the theory of shallow waters. SIAM Journal on Mathematical Analysis, 12(2):229–241, March 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ton:1982:EUL

[Ton82] Bui An Ton. On the existence and uniqueness of a local classical solution of an initial-boundary value problem for incompressible nonhomogeneous viscous fluids. SIAM Journal on Mathematical Analysis, 13(5):699–716, September 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Torres:1996:TPS

[Tor96] Rodolfo H. Torres. A transmission problem in the scattering of electromagnetic waves by a penetrable object. SIAM Journal on Mathematical Analysis, 27(5): 1406–1423, September 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/26738.

Trench:1973:NAE

[Tre73] William F. Trench. Nonnegative and alternating expansions of one set of orthogonal polynomials in terms of another. SIAM Journal on Mathematical Analysis, 4 (1):111–115, February 1973. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Trench:1976:AIL

[Tre76a] William F. Trench. Asymptotic integration of linear differential equations subject to integral smallness conditions involving ordinary convergence. SIAM Journal on Mathematical Analysis, 7(2):213–221, April 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Trench:1976:OPE

[Tre76b] William F. Trench. Orthogonal polynomial expansions with nonnegative coefficients. SIAM Journal on Mathematical Analysis, 7(6):824–833, November 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tremblay:1979:SOF

[Tre79] Richard Tremblay. Some operational formulas involving the operators xD, xΔ and fractional derivatives. SIAM Journal on Mathematical Analysis, 10(5): 933–943, September 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Trench:1984:AIL

[Tre84] William F. Trench. Asymptotic integration of linear differential equations subject to mild integral conditions. SIAM Journal on Mathematical Analysis, 15(5): 932–942, September 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Trench:1985:FPS

[Tre85] William F. Trench. Functional perturbations of second order differential equations. SIAM Journal on Mathematical Analysis, 16 (4):741–756, July 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Trench:1988:AIP

[Tre88] William F. Trench. Asymptotic integration of a perturbed constant coefficient differential equation under mild integral smallness conditions. SIAM Journal on Mathematical Analysis, 19 (2):444–459, March 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tricot:1997:FNF

[Tri97] Claude Tricot. Function norms and fractal dimension. SIAM Journal on Mathematical Analysis, 28(1):189–212, January 1997. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/27879.

Troch:1981:IDL

[Tro81] I. Troch. On the interval of disconjugacy of linear autonomous differential equations. SIAM Journal on Mathematical Analysis, 12(1):78–89, January 1981.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Troy:1987:EBS

[Tro87] William C. Troy. The existence of bounded solutions of a semilinear heat equation. SIAM Journal on Mathematical Analysis, 18 (2):332–336, March 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Troy:1990:BSS

[Tro90] William C. Troy. Bounded solutions of Δu+|u|^{p−1}u−|u|^{q−1}u = 0 in the supercritical case. SIAM Journal on Mathematical Analysis, 21(5):1326–1334, September 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Troy:1993:STO

[Tro93] William C. Troy. Solutions of third-order differential equations relevant to draining and coating flows. SIAM Journal on Mathematical Analysis, 24(1):155–171, January 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tselnik:1997:TSE

[Tse97] D. S. Tselnik. On trigonometric series expansions of twelve Jacobian elliptic functions. SIAM Journal on Mathematical Analysis, 28(3):715–730, May 1997. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/29125.

Tsutsumi:1984:NGS

[Tsu84] Masayoshi Tsutsumi. Nonexistence of global solutions to the Cauchy problem for the damped nonlinear Schrödinger equations. SIAM Journal on Mathematical Analysis, 15(2): 357–366, March 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tsutsumi:1989:CPK

[Tsu89] Yoshio Tsutsumi. The Cauchy problem for the Korteweg–de Vries equation with measures as initial data. SIAM Journal on Mathematical Analysis, 20(3): 582–588, May 1989. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Twombly:1983:BIF

[TT83] Evan Eugene Twombly and J. W.

Thomas. Bifurcating instability of the free surface of a ferrofluid. SIAM Journal on Mathematical Analysis, 14(4):736–766, July 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tudor:1987:SFD

[Tud87] Constantin Tudor. On stochastic functional-differential equations with unbounded delay. SIAM Journal on Mathematical Analysis, 18(6):1716–1725, November 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Turyn:1984:PPB

[Tur84]  Lawrence Turyn.  Perturba-

tion of periodic boundary conditions. SIAM Journal on Mathematical Analysis, 15(4):648–655, July 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Turkington:1989:VRS

[Tur89] Bruce Turkington. Vortex rings with swirl: axisymmetric solutions of the Euler equations with nonzero helicity. SIAM Journal on Mathematical Analysis, 20 (1):57–73, January 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Turyn:1996:CUM

[Tur96] Lawrence Turyn. A centerunstable manifold theorem for parametrically excited surface waves. SIAM Journal on Mathematical Analysis, 27(1):241–257, January 1996. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Travis:1979:ASO

[TW79] C. C. Travis and G. F. Webb. An abstract second-order semilinear Volterra integro-differential equation. SIAM Journal on Mathematical Analysis, 10(2): 412–424, March 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Toland:1989:DFN

[TW89] J. F. Toland and J. R. Willis. Duality for families of natural variational principles in nonlinear electrostatics. SIAM Journal on Mathematical Analysis, 20(6): 1283–1292, November 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tveito:1991:EUC

[TW91] Aslak Tveito and Ragnar Winther.

Existence, uniqueness, and continuous dependence for a system of hyperbolic conservation laws modeling polymer flooding. SIAM Journal on Mathematical Analysis, 22(4):905–933, July 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Tveito:1997:RCE

[TW97] Aslak Tveito and Ragnar Winther.

On the rate of convergence to equilibrium for a system of conservation laws with a relaxation term. SIAM Journal on Mathematical Analysis, 28(1): 136–161, January 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/26375.

Trimble:1989:SFS

[TWW89] S. Y. Trimble, Jim Wells, and F. T. Wright. Superadditive functions and a statistical application. SIAM Journal on Mathematical Analysis, 20(5):1255– 1259, September 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Travis:1977:USS

[TY77] C. C. Travis and E. C. Young. Uniqueness of solutions to singular boundary value problems. SIAM Journal on Mathematical Analysis, 8(1):111–117, February 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ubhaya:1976:MMA

[Ubh76] Vasant A. Ubhaya. Moduli of monotonicity with applications to monotone polynomial approximation. SIAM Journal on Mathematical Analysis, 7(1):117–130, February 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ungar:1974:DT

[Ung74a] Abraham Ungar. A differential transform. SIAM Journal on Mathematical Analysis, 5(6): 920–926, November 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ungar:1974:ORI

[Ung74b] Abraham Ungar. An operator related to the inverse Laplace transform. SIAM Journal on Mathematical Analysis, 5(3): 367–375, May 1974. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Urakawa:1984:NCS

[Ura84]  Hajime Urakawa.  Numeri-

cal computations of the spectra of the Laplacian on 7dimensional homogeneous manifolds SU(3)/T (k, l). SIAM Journal on Mathematical Analysis, 15(5):979–987, September 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Urbach:1996:ADI

[Urb96] H. P. Urbach. Analysis of the domain integral operator for anisotropic dielectric waveguides. SIAM Journal on Mathematical Analysis, 27(1):204–220, January 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/24130.

Utz:1970:NST

[Utz70] W. R. Utz. The nonoscillation of a solution of a third order equation. SIAM Journal on Mathematical Analysis, 1(4):535–537, November 1970. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Valent:1986:ITI

[Val86] G. Valent. An integral transform involving Heun functions and a related eigenvalue problem. SIAM Journal on Mathematical Analysis, 17(3):688–703, May 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Valent:1994:AAS

[Val94] Galliano Valent. Asymptotic analysis of some associated orthogonal polynomials connected with elliptic functions. SIAM Journal on Mathematical Analysis, 25(2):749–775, March 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/22786.

vanDuyn:1979:DIF

[van79] C. J. van Duyn. On the diffusion of immiscible fluids in porous media. SIAM Journal on Mathematical Analysis, 10(3):486–497, May 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

vanDoorn:1984:OPI

[van84] Erik A. van Doorn. On oscillation properties and the interval of orthogonality of orthogonal polynomials. SIAM Journal on Mathematical Analysis, 15(5): 1031–1042, September 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

VanAssche:1985:WZD

[Van85] Walter Van Assche. Weighted zero distribution for polynomials orthogonal on an infinite interval. SIAM Journal on Mathematical Analysis, 16(6):1317–1334, November 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [Van01].

vanHarten:1986:SPC

[van86] A. van Harten. On the spectral properties of a class of elliptic functional-differential operators arising in feedback control theory for diffusion processes. SIAM Journal on Mathematical Analysis, 17(2):352–383, March 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

VanTuyl:1989:EPT

[Van89] Andrew H. Van Tuyl. Electrostatic problems for two conducting spheres. SIAM Journal on Mathematical Analysis, 20(6): 1293–1320, November 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vanderbauwhede:1990:SBR

[Van90] A. Vanderbauwhede. Subharmonic branching in reversible systems. SIAM Journal on Mathematical Analysis, 21(4): 954–979, July 1990. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

VanTuyl:1996:AEE

[Van96] Andrew H. Van Tuyl. Asymptotic expansions with error bounds for the coefficients of capacity and induction of two spheres. SIAM Journal on Mathematical Analysis, 27(3):782– 804, May 1996. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vancostenoble:1998:WAS

[Van98] Judith Vancostenoble. Weak asymptotic stability of secondorder evolution equations by nonlinear and nonmonotone feedbacks. SIAM Journal on Mathematical Analysis, 30(1):

140–154, 1998. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

                                        http://epubs.siam.org/sam-
                                        bin/dbq/article/33237.

VanAssche:2001:EWZ

[Van01] Walter Van Assche. Erratum to “Weighted Zero Distribution For Polynomials Orthogonal on an Infinite Interval”. SIAM Journal on Mathematical Analysis, 32(5): 1169–1170, 2001. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

                                        http://epubs.siam.org/sam-
                                        bin/dbq/article/35987. See

[Van85].

Vardi:1988:DLM

[Var88] Ilan Vardi. Determinants of Laplacians and multiple gamma functions. SIAM Journal on Mathematical Analysis, 19(2): 493–507, March 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vasilach:1979:AMSb

[Vas79a]  Serge Vasilach.  Algebraic

method for solving linear partial differential equations with variable coefficients. II. Fundamental solutions. SIAM Journal on Mathematical Analysis, 10(5): 1077–1088, September 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vasilach:1979:AMSa

[Vas79b]  Serge Vasilach.  Algebraic

method for solving linear partial differential equations with variable coefficients. Part I: Basic theory. SIAM Journal on Mathematical Analysis, 10(3):586–602, May 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vasilach:1981:AMS

[Vas81]  Serge Vasilach.  Algebraic

method for solving systems of linear differential equations with variable coefficients. SIAM Journal on Mathematical Analysis, 12 (1):23–39, January 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

vanEijndhoven:1986:ASS

[vd86] S. J. L. van Eijndhoven and J. de Graaf. Analyticity spaces of selfadjoint operators subjected to perturbations with applications to Hankel invariant distribution spaces. SIAM Journal on Mathematical Analysis, 17(2):485–494, March 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

vanDuyn:1987:LPC

[vd87] C. J. van Duyn and J. M.

de Graaf. Limiting profiles in contaminant transport through porous media. SIAM Journal on Mathematical Analysis, 18(3): 728–743, May 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

vanEijndhoven:1983:CSM

[vdP83] S. J. L. van Eijndhoven, J. de Graaf, and R. S. Pathak. A characterization of the spaces S_{∞/∥+∞}^{∥/∥+∞} by means of holomorphic semigroups. SIAM Journal on Mathematical Analysis, 14(6): 1180–1186, November 1983. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

vanderSluis:1976:ESS

[vdS76] A. van der Sluis. Estimating the solutions of slowly varying recursions. SIAM Journal on Mathematical Analysis, 7(5):662–695, October 1976. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

vandeWetering:1975:VDF

[vdW75] R. L. van de Wetering. Variation diminishing Fourier–Jacobi transforms. SIAM Journal on Mathematical Analysis, 6(5): 774–783, October 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vega:1988:IRG

[Veg88] José M. Vega. Invariant regions and global asymptotic stability in an isothermal catalyst. SIAM Journal on Mathematical Analysis, 19(4):774–796, July 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vega:1993:AEA

[Veg93] José M. Vega. On the amplitude equations arising at the onset of the oscillatory instability in pattern formation. SIAM Journal on Mathematical Analysis, 24(3):

603–617, May 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vemuri:1995:SPF

[Vem95] M. K. Vemuri. A simple proof of Fryant’s theorem. SIAM Journal on Mathematical Analysis, 26(6): 1644–1646, November 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/25818.

Venkatesh:1978:GVC

[Ven78] Y. V. Venkatesh. Global variation criteria for the L_{2}-stability of nonlinear time varying systems. SIAM Journal on Mathematical Analysis, 9(3):568–581, June 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Verma:1972:CTZ

[Ver72] Arun Verma. Characterizations of σ-type zero polynomial sets. SIAM Journal on Mathematical Analysis, 3(4):637–641, November 1972. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Verma:1980:QTB

[Ver80] Arun Verma. A quadratic transformation of a basic hypergeometric series. SIAM Journal on Mathematical Analysis, 11(3): 425–427, May 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vessella:1997:SEI

[Ves97] Sergio Vessella. Stability estimates in an inverse problem for a three-dimensional heat equation. SIAM Journal on Mathematical Analysis, 28(6):1354– 1370, November 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29426.

VanAssche:1987:ABC

[VFO87] W. Van Assche, G. Fano, and F. Ortolani. Asymptotic behaviour of the coefficients of some sequences of polynomials. SIAM Journal on Mathematical Analysis, 18(6):1597–1615, November 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Victory:1978:SRI

[Vic78] H. D. Victory, Jr. On the spectral radius of integral operators defined by a class of difference kernels. SIAM Journal on Mathematical Analysis, 9(3):558–567, June 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Villemoes:1992:EMT

[Vil92] Lars F. Villemoes. Energy moments in time and frequency for two-scale difference equation solutions and wavelets. SIAM Journal on Mathematical Analysis, 23(6):1519–1543, November 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Villemoes:1994:WAR

[Vil94] Lars F. Villemoes. Wavelet analysis of refinement equations. SIAM Journal on Mathematical Analysis, 25(5):1433–1460, September 1994. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL

http://epubs.siam.org/sam-
bin/dbq/article/22817.

Visintin:1986:EPH

[Vis86] A. Visintin. Evolution problems with hysteresis in the source term. SIAM Journal on Mathematical Analysis, 17(5):1113– 1138, September 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Visintin:1990:NFR

[Vis90] Augusto Visintin. Nonconvex functionals related to multiphase systems. SIAM Journal on Mathematical Analysis, 21(5):1281– 1304, September 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vitins:1979:EBM

[Vit79] M. Vitins. Error bounds in the method of averaging based on properties of average motion. SIAM Journal on Mathematical Analysis, 10(2):331–336, March 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Verma:1981:STB

[VJ81] A. Verma and V. K. Jain.

Some transformations of basic hypergeometric functions. Part I. SIAM Journal on Mathematical Analysis, 12(6):943–956, November 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Verma:1985:SSF

[VJ85] A. Verma and V. K. Jain. Some summation formulae for nonterminating basic hypergeometric series. SIAM Journal on Mathematical Analysis, 16(3):647–655, May 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

VanAssche:1991:ABW

[VK91] Walter Van Assche and Tom H. Koornwinder. Asymptotic behaviour for Wall polynomials and the addition formula for little q-Legendre polynomials. SIAM Journal on Mathematical Analysis, 22(1):302–311, January 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vogel:1985:FBP

[Vog85]  Thomas I. Vogel.  A free-

boundary problem arising from a galvanizing process. SIAM Journal on Mathematical Analysis, 16(5):970–979, September 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Volkmer:1982:IRL

[Vol82] Hans Volkmer. Integral relations for Lamé functions. SIAM Journal on Mathematical Analysis, 13(6):978–987, November 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Volkmer:1984:IRP

[Vol84] Hans Volkmer. Integral representations for products of Lamé functions by use of fundamental solutions. SIAM Journal on Mathematical Analysis, 15(3): 559–569, May 1984. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Volkmer:1987:ETF

[Vol87] Hans Volkmer. On an expansion theorem of F. V. Atkinson and P. Binding. SIAM Journal on Mathematical Analysis, 18(6): 1669–1680, November 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Volkmer:1990:EHF

[Vol90] Hans Volkmer. The expansion of a holomorphic function in a Laplace series. SIAM Journal on Mathematical Analysis, 21(3): 771–781, May 1990. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Volkmer:1995:ARC

[Vol95] Hans Volkmer. Asymptotic regularity of compactly supported wavelets. SIAM Journal on Mathematical Analysis, 26(4): 1075–1087, July 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24896.

vanderMee:1986:KER

[vP86] C. V. M. van der Mee and V. Protopopescu. Kinetic equations with reflecting boundary conditions. SIAM Journal on Mathematical Analysis, 17(4):933–944, July 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vretare:1984:FES

[Vre84] Lars Vretare. Formulas for elementary spherical functions and generalized Jacobi polynomials. SIAM Journal on Mathematical Analysis, 15(4):805–833, July 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Vanden-Broeck:1979:OPF

[VS79] Jean-Marc Vanden-Broeck and Leonard W. Schwartz. A oneparameter family of sequence transformations. SIAM Journal on Mathematical Analysis, 10(3): 658–666, May 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

vanderAa:1984:AIP

[vV84] Els van der Aa and Ferdinand Verhulst. Asymptotic integrability and periodic solutions of a Hamiltonian system in 1:2:2resonance. SIAM Journal on Mathematical Analysis, 15(5): 890–911, September 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

vanHarten:1985:QLS

[vV85] A. van Harten and R. R. Van Hassel. A quasi-linear, singular perturbation problem of hyperbolic type. SIAM Journal on Mathematical Analysis, 16(6): 1258–1267, November 1985. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wade:1980:UWS

[Wad80] William R. Wade. Uniqueness of Walsh series which satisfy an averaged growth condition. SIAM Journal on Mathematical Analysis, 11(6):933–937, November 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wagner:1983:RPT

[Wag83] David H. Wagner. The Riemann problem in two space dimensions for a single conservation law. SIAM Journal on Mathematical Analysis, 14(3):534–559, May 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wagner:1989:EBV

[Wag89] David H. Wagner. The existence and behavior of viscous structure for plane detonation waves. SIAM Journal on Mathematical Analysis, 20(5):1035– 1054, September 1989. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Waid:1976:SON

[Wai76] Margaret C. Waid. A second order nonlinear degenerate parabolic equation with nonlinear boundary conditions. SIAM Journal on Mathematical Analysis, 7(3):373–383, May 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Walker:1971:ACN

[Wal71a] Philip W. Walker. Asymptotics for a class of nonanalytic second order differential equations. SIAM Journal on Mathematical Analysis, 2(3):328–339, August 1971. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Walter:1971:SPS

[Wal71b] Gilbert G. Walter. Singular points of Sturm–Liouville series. SIAM Journal on Mathematical Analysis, 2(3):393–401, August 1971. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Walter:1975:ENL

[Wal75a] Wolfgang Walter. On existence and nonexistence in the large of solutions of parabolic differential equations with a nonlinear boundary condition. SIAM Journal on Mathematical Analysis, 6 (1):85–90, February 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Walton:1975:DAD

[Wal75b] J. R. Walton. A distributional approach to dual integral equations of Titchmarsh type. SIAM Journal on Mathematical Analysis, 6(4):628–643, August 1975.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wallgren:1976:OSD

[Wal76] Thomas Wallgren. Oscillation of solutions of the differential equation y′′ + p(x)y = f(x). SIAM Journal on Mathematical Analysis, 7(6):848–857, November 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Walton:1979:SGA

[Wal79] J. R. Walton. Systems of generalized Abel integral equations with applications to simultaneous dual relations. SIAM Journal on Mathematical Analysis, 10(4): 808–822, July 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Walter:1988:SBF

[Wal88] Gilbert G. Walter. Sampling bandlimited functions of polynomial growth. SIAM Journal on Mathematical Analysis, 19(5): 1198–1203, September 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Walter:1992:DW

[Wal92a] Gilbert G. Walter. Discrete wavelets. SIAM Journal on Mathematical Analysis, 23(4): 1004–1014, July 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Walter:1992:NSB

[Wal92b] Gilbert G. Walter. Nonuniform sampling of bandlimited functions of polynomial growth. SIAM Journal on Mathematical Analysis, 23(4):995–1003, July 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Waldron:1997:MFH

[Wal97] Shayne Waldron. A multivariate form of Hardy’s inequality and L_{p}-error bounds for multivariate Lagrange interpolation schemes. SIAM Journal on Mathematical Analysis, 28(1): 233–258, January 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/27550.

Wang:1978:ABS

[Wan78] Frank J. S. Wang. Asymptotic behavior of some deterministic epidemic models. SIAM Journal on Mathematical Analysis, 9(3): 529–534, June 1978. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:1979:PMN

[Wan79] Frank J. S. Wang. Perturbation and maximum number of infectives of some sir epidemic models. SIAM Journal on Mathematical Analysis, 10(4):721–727, July 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:1980:SAD

[Wan80] Frank J. S. Wang. Stability of an age-dependent population. SIAM Journal on Mathematical Analysis, 11(4):683–689, July 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:1998:SLI

[Wan98a] Jianzhong Wang. Stability and linear independence associated with scaling vectors. SIAM Journal on Mathematical Analysis, 29(5):1140–1156, September 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/30330.

Wang:1998:CRV

[Wan98b] Wei-Cheng Wang. On L_{1} convergence rate of viscous and numerical approximate solutions of genuinely nonlinear scalar conservation laws. SIAM Journal on Mathematical Analysis, 30(1):38–52, 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/31640.

Wasow:1970:STP

[Was70]  Wolfgang Wasow.  Simple

turning-point problems in unbounded domains. SIAM Journal on Mathematical Analysis, 1(2): 153–170, May 1970. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wasow:1973:AIS

[Was73] Wolfgang Wasow. Adiabatic invariance of a simple oscillator. SIAM Journal on Mathematical Analysis, 4(1):78–88, February 1973. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wasow:1974:CAI

[Was74] Wolfgang Wasow. Calculation of an adiabatic invariant by turning point theory. SIAM Journal on Mathematical Analysis, 5 (4):673–700, August 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wimp:1995:FTP

[WB95] Jet Wimp and Bernhard Beckermann. Families of two-point Padé approximants and some 4F_{3}(1) identities. SIAM Journal on Mathematical Analysis, 26 (3):761–773, May 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/22349.

Webb:1979:BPN

[Web79] G. F. Webb. A bifurcation problem for a nonlinear hyperbolic partial differential equation. SIAM Journal on Mathematical Analysis, 10(5):922–932, September 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Weinberg:1971:ADE

[Wei71] Laurence Weinberg. The asymptotic distribution of eigenvalues for the boundary value problem y′′(x) − λ^{2}p(x)y(x) = 0, y ∈ L_{2}(−∞, +∞). SIAM Journal on Mathematical Analysis, 2(4): 546–566, November 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Weinacht:1974:SPG

[Wei74] R. J. Weinacht. Some properties of generalized biaxially symmetric Helmholtz potentials. SIAM Journal on Mathematical Analysis, 5(1):147–152, February 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Weinberger:1982:LTB

[Wei82] H. F. Weinberger. Long-time behavior of a class of biological models. SIAM Journal on Mathematical Analysis, 13(3): 353–396, May 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Weinmuller:1984:BVP

[Wei84] Ewa Weinmüller. On the boundary value problem for systems of ordinary second-order differential equations with a singularity of the first kind. SIAM Journal on Mathematical Analysis, 15 (2):287–307, March 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Weinstein:1985:MSG

[Wei85] Michael I. Weinstein. Modulational stability of ground states of nonlinear Schrödinger equations. SIAM Journal on Mathematical Analysis, 16(3):472–491, May 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Weinitschke:1988:UAD

[Wei88] Hubertus J. Weinitschke. On uniqueness of axisymmetric deformations of elastic plates and shells. SIAM Journal on Mathematical Analysis, 19(3):580–592, May 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Weiss:1999:SSB

[Wei99] G. S. Weiss. Self-similar blowup and Hausdorff dimension estimates for a class of parabolic free boundary problems. SIAM Journal on Mathematical Analysis, 30(3):623–644, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/32740.

Wells:1991:MEF

[Wel91] Benjamin B. Wells. Minimal extrapolations of filters. SIAM Journal on Mathematical Analysis, 22(3):796–809, May 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Weston:1978:ASP

[Wes78] V. H. Weston. On the asymptotic solution of a partial differential equation with an exponential nonlinearity. SIAM Journal on Mathematical Analysis, 9(6):

1030–1053, December 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wexler:1979:FDS

[Wex79] D. Wexler. Frequency domain stability for a class of equations arising in reactor dynamics. SIAM Journal on Mathematical Analysis, 10(1):118–138, January 1979. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wexler:1980:FDS

[Wex80] D. Wexler. On frequency domain stability for evolution equations in Hilbert spaces via the algebraic Riccati equation. SIAM Journal on Mathematical Analysis, 11(6):969–983, November 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Webb:1987:AEG

[WG87] G. F. Webb and A. Grabosch. Asynchronous exponential growth in transition probability models of the cell cycle. SIAM Journal on Mathematical Analysis, 18(4):897–908, July 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wiggins:1987:HOS

[WH87a] Stephen Wiggins and Philip Holmes. Homoclinic orbits in slowly varying oscillators. SIAM Journal on Mathematical Analysis, 18(3):612–629, May 1987.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See errata [WH88].

Wiggins:1987:POS

[WH87b] Stephen Wiggins and Philip Holmes. Periodic orbits in slowly varying oscillators. SIAM Journal on Mathematical Analysis, 18(3):592–611, May 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wiggins:1988:EHO

[WH88] Stephen Wiggins and Philip Holmes. Errata: “Homoclinic orbits in slowly varying oscillators” [SIAM J. Math. Anal. 18 (1987), no. 3, 612–629; MR 88i:34087b]. SIAM Journal on Mathematical Analysis, 19(5): 1254–1255, September 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [WH87a].

Waldstatter:1988:LMM

[WHG88] R. Waldstätter, K. P. Hadeler, and G. Greiner. A Lotka– McKendrick model for a population structured by the level of parasitic infection. SIAM Journal on Mathematical Analysis, 19 (5):1108–1118, September 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

White:1979:WS

[Whi79] R. E. White. Weak solutions of (p(x)u′(x))′+g(x)u′(x)+qu(x) = f with q, f ∈ H_{−1}[a, b], 0 < p(x) ∈ L_{∞}[a, b], g(x) ∈ L_{∞}[a, b] and u ∈ H_{1}[a, b]. SIAM Journal on Mathematical Analysis, 10(6): 1313–1325, November 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

White:1980:VSA

[Whi80] R. E. White. Value set at x ∈ Ω for an arbitrary distribution with applications to local extrema of f ∈ C(Ω) and a maximum principle for ordinary differential equations. SIAM Journal on Mathematical Analysis, 11 (1):61–72, January 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wiebers:1987:ENS

[Wie87]  Henning Wiebers.  The ex-

act number of solutions of finite difference approximations of some nonlinear elliptic boundary value problems. SIAM Journal on Mathematical Analysis, 18(4): 954–965, July 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wilson:1970:HP

[Wil70] M. Wayne Wilson. On the Hahn polynomials. SIAM Journal on Mathematical Analysis, 1 (1):131–139, February 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Willett:1971:DTS

[Wil71] D. Willett. Disconjugacy tests for singular linear differential equations. SIAM Journal on Mathematical Analysis, 2(4):

536–545, November 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Willett:1972:EDT

[Wil72a] D. Willett. Errata: Disconjugacy tests for singular linear differential equations. SIAM Journal on Mathematical Analysis, 3(3): 559, August 1972. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wilson:1972:NDA

[Wil72b] M. Wayne Wilson. On a new discrete analogue of the Legendre polynomials. SIAM Journal on Mathematical Analysis, 3(1): 157–169, February 1972. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wilkins:1973:EBD

[Wil73] J. Ernest Wilkins, Jr. The eigenvalues of the Bethe differential system. SIAM Journal on Mathematical Analysis, 4(1):62– 77, February 1973. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wilson:1979:IIF

[Wil79] F. J. Wilson. Ideal inversion formulae for the Fourier transform. SIAM Journal on Mathematical Analysis, 10(1):80–85, January 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wilkes:1980:CDI

[Wil80a] N. S. Wilkes. Continuous dependence and instability in linear thermoelasticity. SIAM Journal on Mathematical Analysis, 11 (2):292–299, March 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Willson:1980:SSC

[Wil80b] Stephen J. Willson. A semigroup on the space of compact convex bodies. SIAM Journal on Mathematical Analysis, 11(3): 448–457, May 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wilson:1980:SHO

[Wil80c] James A. Wilson. Some hypergeometric orthogonal polynomials. SIAM Journal on Mathematical Analysis, 11(4):690–701, July 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Willson:1985:CGC

[Wil85] Stephen J. Willson. On coherent growth of configurations. SIAM Journal on Mathematical Analysis, 16(2):316–330, March 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wilson:1986:RFL

[Wil86] F. J. Wilson. The representation of functions as Laplace and Laplace–Stieltjes transforms. SIAM Journal on Mathematical Analysis, 17(3):720–733, May 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wilson:1991:OFG

[Wil91] James A. Wilson. Orthogonal functions from Gram determinants. SIAM Journal on Mathematical Analysis, 22(4):1147– 1155, July 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wimp:1974:ETS

[Wim74] Jet Wimp. Equiconvergence theorems for series whose terms satisfy a difference equation. SIAM Journal on Mathematical Analysis, 5(3):509–516, May 1974. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wimp:1985:SEP

[Wim85] Jet Wimp. Some explicit Padé approximants for the function Φ′/Φ and a related quadrature formula involving Bessel functions. SIAM Journal on Mathematical Analysis, 16(4):887–895, July 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Winston:1974:ASO

[Win74] Elliot Winston. Asymptotic stability for ordinary differential equations with delayed perturbations. SIAM Journal on Mathematical Analysis, 5(2):303–308, April 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:1992:LCC

[WK92] Xian Wang and Robert E. Kooij. Limit cycles in a cubic system with a cusp. SIAM Journal on Mathematical Analysis, 23(6): 1609–1622, November 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Williams:1982:NSN

[WL82a] Lynn R. Williams and Richard W.

Leggett. Nonzero solutions of nonlinear integral equations modeling infectious disease. SIAM Journal on Mathematical Analysis, 13(1):112–121, January 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Williams:1982:UMS

[WL82b] Lynn R. Williams and Richard W.

Leggett. Unique and multiple solutions of a family of differential equations modeling chemical reactions. SIAM Journal on Mathematical Analysis, 13(1): 122–133, January 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Willner:1982:TDE

[WM82] B. E. Willner and T. J. Mahar. The two-dimensional eigenvalue range and extremal eigenvalue problems. SIAM Journal on Mathematical Analysis, 13(4): 621–631, July 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wolfe:1987:CSS

[WM87] Robert Wolfe and Hedley C.

Morris. Chaotic solutions of systems of first order partial differential equations. SIAM Journal on Mathematical Analysis, 18 (4):1040–1063, July 1987. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Walter:1981:SOP

[WN81] Gilbert G. Walter and Paul G. Nevai. Series of orthogonal polynomials as boundary values. SIAM Journal on Mathematical Analysis, 12(4):502–513, July 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Watanabe:1991:PIT

[WO91] Masaji Watanabe and Hans G. Othmer. Persistence of invariant tori in systems of coupled oscillators. II. degenerate problems. SIAM Journal on Mathematical Analysis, 22(6):1584– 1630, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wolfe:1971:DDF

[Wol71] Peter Wolfe. Distribution derivatives of functions harmonic in the unit disc. SIAM Journal on Mathematical Analysis, 2(3): 431–434, August 1971. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wolanski:1993:GBP

[Wol93] Noemí Wolanski. Global behavior of positive solutions to nonlinear-diffusion problems with nonlinear absorption through the boundary. SIAM Journal on Mathematical Analysis, 24(2):317–326, March 1993.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wolfe:1996:GBE

[Wol96] Peter Wolfe. Global bifurcation of an elastic conducting rod in a magnetic field. SIAM Journal on Mathematical Analysis, 27 (2):528–542, March 1996. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/26305.

Wong:1970:LII

[Won70] R. Wong. On a Laplace integral involving logarithms. SIAM Journal on Mathematical Analysis, 1(3):360–364, August 1970. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wong:1976:EBA

[Won76] R. Wong. Error bounds for asymptotic expansions of Hankel transforms. SIAM Journal on Mathematical Analysis, 7(6): 799–808, November 1976. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wong:1978:AEF

[Won78] R. Wong. Asymptotic expansions of fractional integrals involving logarithms. SIAM Journal on Mathematical Analysis, 9 (5):835–842, October 1978. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wong:1980:AEH

[Won80] R. Wong. Asymptotic expansion of the Hilbert transform. SIAM Journal on Mathematical Analysis, 11(1):92–99, January 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wong:1988:SON

[Won88] James S. W. Wong. Second order nonlinear forced oscillations. SIAM Journal on Mathematical Analysis, 19(3):667–675, May 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wolf:1980:HAB

[WS80] Kurt Bernardo Wolf and Thomas H. Seligman. Harmonic analysis on bilateral classes. SIAM Journal on Mathematical Analysis, 11(6): 1068–1074, November 1980. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wu:1990:LCM

[Wu90] Li Wu. Lipschitz continuous metric selections in C_{0}(T). SIAM Journal on Mathematical Analysis, 21(1):205–220, January 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wasilkowski:1988:OAA

[WW88a] G. W. Wasilkowski and H. Woźniakowski. On optimal algorithms in an asymptotic model with Gaussian measure. SIAM Journal on Mathematical Analysis, 19(3):

632–647, May 1988. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wilde:1988:DSF

[WW88b] Peter Wilde and Achim Willers.

Dense sets and far field patterns for the vector Helmholtz equation under transmission boundary conditions. SIAM Journal on Mathematical Analysis, 19 (2):348–357, March 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wong:1992:AJF

[WW92] R. Wong and Q.-Q. Wang. On the asymptotics of the Jacobi function and its zeros. SIAM Journal on Mathematical Analysis, 23(6):1637–1649, November 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:1993:GEB

[WW93] Ming Xin Wang and Yong Hui Wu. Global existence and blow-up problems for quasilinear parabolic equations with nonlinear boundary conditions. SIAM Journal on Mathematical Analysis, 24(6):1515–1521, November 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wei:1998:SCH

[WW98] Juncheng Wei and Matthias Winter. On the stationary Cahn– Hilliard equation: Bubble solutions. SIAM Journal on Mathematical Analysis, 29(6):1492– 1518, November 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/32066.

Wei:1999:TDG

[WW99] Juncheng Wei and Matthias Winter. On the two-dimensional Gierer–Meinhardt system with strong coupling. SIAM Journal on Mathematical Analysis, 30(6):1241–1263, 1999. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/34723.

Weiss:1975:ISM

[WY75] J. R. Weiss and E. Y. Yu. An investigation of stability of motion under constantly acting disturbances. SIAM Journal on Mathematical Analysis, 6(1):43– 48, February 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wang:1992:AVC

[WY92] Sheng Wang and Yi Song Yang. Abrikosov’s vortices in the critical coupling. SIAM Journal on Mathematical Analysis, 23(5): 1125–1140, September 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wyrwich:1977:ESC

[Wyr77] Herbert Wyrwich. An explicit solution of the central connection problem for an n th order linear ordinary differential equation with polynomial coefficients. SIAM Journal on Mathematical Analysis, 8(3):412–422, May 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wysocki:1993:SET

[Wys93] Krzysztof Wysocki. Some ergodic theorems for solutions of homogeneous differential equations. SIAM Journal on Mathematical Analysis, 24(3):681–702, May 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Walter:1987:RSS

[WZ87] Gilbert G. Walter and Ahmed J.

Zayed. Real singularities of singular Sturm–Liouville expansions. SIAM Journal on Mathematical Analysis, 18(1):219–227, January 1987. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Wong:1994:REJ

[WZ94] R. Wong and J.-M. Zhang.

On the relative extrema of the Jacobi polynomials P_{n}^{(0,−1)}(x). SIAM Journal on Mathematical Analysis, 25(2):776–811, March 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/22389.

Wang:1997:CPB

[WZ97] Xuefeng Wang and Bin Zeng. On concentration of positive bound states of nonlinear Schrödinger equations with competing potential functions. SIAM Journal on Mathematical Analysis, 28 (3):633–655, May 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/29024.

Xie:1991:SCN

[XA91] Hong Xie and W. Allegretto. C^{α}(\overline{Ω}) solutions of a class of nonlinear degenerate elliptic systems arising in the thermistor problem. SIAM Journal on Mathematical Analysis, 22(6):1491– 1499, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Xie:1990:SPK

[Xie90] Wei Qing Xie. The Stefan problem with a kinetic condition at the free boundary. SIAM Journal on Mathematical Analysis, 21 (2):362–373, March 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Xu:1992:DSL

[Xu92] Xiang Sheng Xu. A degenerate Stefan-like problem with Joule’s heating. SIAM Journal on Mathematical Analysis, 23(6): 1417–1438, November 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Xu:1993:MOP

[Xu93] Yuan Xu. On multivariate orthogonal polynomials. SIAM Journal on Mathematical Analysis, 24(3):783–794, May 1993.

CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Xu:1994:ZMQ

[Xu94] Yuan Xu. On zeros of multivariate quasi-orthogonal polynomials and Gaussian cubature formulae. SIAM Journal on Mathematical Analysis, 25(3): 991–1001, May 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/23720.

Xu:1998:OPC

[Xu98] Yuan Xu. Orthogonal polynomials and cubature formulae on spheres and on balls. SIAM Journal on Mathematical Analysis, 29 (3):779–793, May 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/30735.

Yagasaki:1992:CDQ

[Yag92] Kazuyuki Yagasaki. Chaotic dynamics of quasi-periodically forced oscillators detected by Mel′nikov’s method. SIAM Journal on Mathematical Analysis, 23 (5):1230–1254, September 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yamamoto:1971:NSM

[Yam71] Tetsuro Yamamoto. A note on the spectral mapping theorem. SIAM Journal on Mathematical Analysis, 2(1):49–51, February 1971. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yamada:1990:SSS

[Yam90] Yoshio Yamada. Stability of steady states for prey-predator diffusion equations with homogeneous Dirichlet conditions. SIAM Journal on Mathematical Analysis, 21(2):327–345, March 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yang:1992:QAO

[Yan92] Kuang Yang. Qualitative analysis of one- or two-species neutral delay population models. SIAM Journal on Mathematical Analysis, 23(1):181–200, January 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yanagida:1996:SRS

[Yan96]  Eiji Yanagida.  Structure

of radial solutions to Δu + K(|x|)|u|^{p−1}u = 0 in R^{n}. SIAM Journal on Mathematical Analysis, 27(4):997–1014, July 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/26132.

Yao:1995:ILT

[Yao95] Peng Fei Yao. On the inversion of the Laplace transform of C_{0} semigroups and its applications. SIAM Journal on Mathematical Analysis, 26(5):1331– 1341, September 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24003.

Yarmish:1975:NMT

[Yar75]  Joshua Yarmish.  Newton’s

method techniques for singular perturbations. SIAM Journal on Mathematical Analysis, 6(4): 661–680, August 1975. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Young:1982:ZCJ

[YH82] Andrew Young and Hassan Hamideh. The zeros of certain Jacobi polynomials. SIAM Journal on Mathematical Analysis, 13(5):875–878, September 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yin:1991:PVE

[Yin91] Hong Ming Yin. On parabolic Volterra equations in several space dimensions. SIAM Journal on Mathematical Analysis, 22(6): 1723–1737, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yin:1998:MEE

[Yin98] Hong-Ming Yin. On Maxwell’s equations in an electromagnetic field with the temperature effect. SIAM Journal on Mathematical Analysis, 29(3):637–651, May 1998. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/31615.

Yuan:1994:EUS

[YL94] Guang Wei Yuan and Zu Han Liu. Existence and uniqueness of the C^{α}-solution for the thermistor problem with mixed boundary value. SIAM Journal on Mathematical Analysis, 25 (4):1157–1166, July 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/23789.

Yokoyama:1997:MGS

[Yok97] Toshiaki Yokoyama. Monodromy groups of systems of total differential equations of two variables. SIAM Journal on Mathematical Analysis, 28(5):1227– 1247, September 1997. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/29687.

Yomdin:1983:FRS

[Yom83] Y. Yomdin. On functions representable as a supremum of a family of smooth functions. SIAM Journal on Mathematical Analysis, 14(2):239–246, March 1983. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yomdin:1986:CVR

[Yom86a] Y. Yomdin. Critical values and representation of functions by means of compositions. SIAM Journal on Mathematical Analysis, 17(1):236–239, January 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yomdin:1986:FRS

[Yom86b] Y. Yomdin. On functions representable as a supremum of a family of smooth functions. II. SIAM Journal on Mathematical Analysis, 17(4):961–969, July 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yoshida:1983:OPS

[Yos83] Norio Yoshida. Oscillation properties of solutions of second order elliptic equations. SIAM Journal on Mathematical Analysis, 14(4): 709–718, July 1983. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yoshikawa:1984:BSL

[Yos84] Atsushi Yoshikawa. On the behavior of the solutions to the Lamm equation of the ultracentrifuge. SIAM Journal on Mathematical Analysis, 15(4):686–711, July 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yoshida:1985:FOE

[Yos85] Norio Yoshida. Forced oscillations of extensible beams. SIAM Journal on Mathematical Analysis, 16(2):211–220, March 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yoshikawa:1986:BSL

[Yos86] Atsushi Yoshikawa. On the behavior of the solutions to the Lamm equation of the ultracentrifuge. II. SIAM Journal on Mathematical Analysis, 17(3): 641–645, May 1986. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Young:1975:MPE

[You75] Eutiquio C. Young. A mixed problem for the Euler–Poisson– Darboux equation in two space variables. SIAM Journal on Mathematical Analysis, 6(1):17– 24, February 1975. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Young:1999:ESD

[You99]  Robin Young.  Exact solu-

tions to degenerate conservation laws. SIAM Journal on Mathematical Analysis, 30(3):537– 558, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/32723.

Yad-Shalom:1993:SSD

[YS93]  I. Yad-Shalom.  Subdivision

schemes determined by coefficients of a Hurwitz polynomial. SIAM Journal on Mathematical Analysis, 24(1):234–245, January 1993. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yu:1974:IRA

[Yu74] Ching-Ling Yu. Integral representation, analytic continuation and the reflection principle under the complementing boundary condition for higher order elliptic equations in the plane. SIAM Journal on Mathematical Analysis, 5(2):209–223, April 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yu:1977:CPS

[Yu77] Chung-Ling Yu. Cauchy’s problem for systems of first order analytic elliptic equations in the plan. SIAM Journal on Mathematical Analysis, 8(4):719–740, August 1977. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yuen:1973:AWT

[Yue73] C. K. Yuen. Analysis of Walsh transforms using integration by parts. SIAM Journal on Mathematical Analysis, 4(4):574–584, November 1973. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Yu:1994:MSA

[YZ94] X. M. Yu and S. P. Zhou.

On monotone spline approximation. SIAM Journal on Mathematical Analysis, 25(4):1227– 1239, July 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/22825.

Zachmann:1974:BSE

[Zac74] David W. Zachmann. Branching solutions of equations containing several parameters. SIAM Journal on Mathematical Analysis, 5(6):898–907, November 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zagatti:1998:DPV

[Zag98] Sandro Zagatti. On the Dirichlet problem for vectorial Hamilton– Jacobi equations. SIAM Journal on Mathematical Analysis, 29(6): 1481–1491, November 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/32127.

Zajaczkowski:1994:NMC

[Zaj94] W. M. Zajaczkowski. On nonstationary motion of a compressible barotropic viscous capillary fluid bounded by a free surface. SIAM Journal on Mathematical Analysis, 25(1):1– 84, January 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/17320.

Zayed:1982:LSB

[Zay82] Ahmed I. Zayed. Laguerre series as boundary values. SIAM Journal on Mathematical Analysis, 13 (2):263–279, March 1982. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ziolkowski:1984:APL

[ZD84] Richard W. Ziolkowski and Georges A. Deschamps. The asymptotic Poincaré lemma and its applications. SIAM Journal on Mathematical Analysis, 15(3): 535–558, May 1984. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zeilberger:1980:ALP

[Zei80] Doron Zeilberger. The algebra of linear partial difference operators and its applications. SIAM Journal on Mathematical Analysis, 11(6):919–932, November 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zeilberger:1987:PCM

[Zei87] Doron Zeilberger. A proof of the G_{2} case of Macdonald’s root system-Dyson conjecture. SIAM Journal on Mathematical Analysis, 18(3):880–883, May 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zeilberger:1988:UAM

[Zei88] Doron Zeilberger. A unified approach to Macdonald’s rootsystem conjectures. SIAM Journal on Mathematical Analysis, 19 (4):987–1013, July 1988. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zeilberger:1994:TEM

[Zei94] Doron Zeilberger. Towards a W Z evolution of the Mehta integral. SIAM Journal on Mathematical Analysis, 25(2): 812–814, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/22875.

Zemanian:1979:CIN

[Zem79a] A. H. Zemanian. Countably infinite nonlinear time-varying active electrical networks. SIAM Journal on Mathematical Analysis, 10(5):944–960, September 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zemanian:1979:CIT

[Zem79b] A. H. Zemanian. Countably infinite time-varying electrical networks. SIAM Journal on Mathematical Analysis, 10(6):1193– 1198, November 1979. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zemanian:1981:CRM

[Zem81] A. H. Zemanian. The characteristicresistance method for grounded semi-infinite grids. SIAM Journal on Mathematical Analysis, 12 (1):115–138, January 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zemanian:1982:NSI

[Zem82] A. H. Zemanian. Nonuniform semi-infinite grounded grids. SIAM Journal on Mathematical Analysis, 13(5):770–788, September 1982. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zeng:1990:LPP

[Zen90]  Jiang Zeng.  Linéarisation

de produits de polynômes de Meixner, Krawtchouk, et Charlier. (French) [Linearization of products of Meixner, Krawtchouk and Charlier polynomials]. SIAM Journal on Mathematical Analysis, 21(5): 1349–1368, September 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhang:1986:ASO

[Zha86] Gong Li Zhang. p-ary sequency and orderings of the Chrestenson functions. SIAM Journal on Mathematical Analysis, 17(5): 1259–1266, September 1986. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhang:1992:UCK

[Zha92a] Bing-Yu Zhang. Unique continuation for the Korteweg–De Vries equation. SIAM Journal on Mathematical Analysis, 23 (1):55–71, January 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhao:1992:GLI

[Zha92b] Kang Zhao. Global linear independence and finitely supported dual basis. SIAM Journal on Mathematical Analysis, 23(5): 1352–1355, September 1992. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhang:1994:LET

[Zha94] Bo Zhang. On a local existence theorem for a simplified one-dimensional hydrodynamic model for semiconductor devices. SIAM Journal on Mathematical Analysis, 25 (3):941–947, May 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/22459.

Zhang:1995:ASG

[Zha95a] Bing Yu Zhang. Analyticity of solutions of the generalized Korteweg–de Vries equation with respect to their initial values. SIAM Journal on Mathematical Analysis, 26(6):1488– 1513, November 1995. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/24260.

Zhang:1995:PSS

[Zha95b] Yong Zhang. Positive solutions of singular sublinear Dirichlet boundary value problems. SIAM Journal on Mathematical Analysis, 26(2):329–339, March 1995. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/24608.

Zhai:1999:EBS

[Zha99a] Jian Zhai. Existence and behavior of solutions to the Landau– Lifshitz equation. SIAM Journal on Mathematical Analysis, 30(4):833–847, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32795.

Zhang:1999:GES

[Zha99b] Yongqian Zhang. Global existence of steady supersonic potential flow past a curved wedge with a piecewise smooth boundary. SIAM Journal on Mathematical Analysis, 31(1):166– 183, 1999. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/33105.

Zhou:1989:RHP

[Zho89]  Xin Zhou.  The Riemann–

Hilbert problem and inverse scattering. SIAM Journal on Mathematical Analysis, 20(4):966–986, July 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhou:1996:SRF

[Zho96] Ding-Xuan Zhou. Stability of refinable functions, multiresolution analysis, and Haar bases. SIAM Journal on Mathematical Analysis, 27(3):891–904, May 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zinner:1991:STW

[Zin91] B. Zinner. Stability of traveling wavefronts for the discrete Nagumo equation. SIAM Journal on Mathematical Analysis, 22 (4):1016–1020, July 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ziolkowski:1985:SPC

[Zio85] Richard W. Ziolkowski. n-series problems and the coupling of electromagnetic waves to apertures: a Riemann–Hilbert approach. SIAM Journal on Mathematical Analysis, 16(2):358–378, March 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zucker:1979:SSH

[Zuc79] I. J. Zucker. The summation of series of hyperbolic functions. SIAM Journal on Mathematical Analysis, 10(1):192–206, January 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zucker:1984:SIS

[Zuc84] I. J. Zucker. Some infinite series of exponential and hyperbolic functions. SIAM Journal on Mathematical Analysis, 15 (2):406–413, March 1984. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zayed:1982:SOP

[ZW82] Ahmed I. Zayed and Gilbert G. Walter. Series of orthogonal polynomials as hyperfunctions. SIAM Journal on Mathematical Analysis, 13(4):664–675, July 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zayed:1985:SSS

[ZW85] Ahmed I. Zayed and Gilbert G.

Walter. On the singularities of singular Sturm–Liouville expansions and an associated class of elliptic PDEs. SIAM Journal on Mathematical Analysis, 16(4): 725–740, July 1985. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zhang:1990:CSS

[ZZ90] Tong Zhang and Yu Xi Zheng.

Conjecture on the structure of solutions of the Riemann problem for two-dimensional gas dynamics systems. SIAM Journal on Mathematical Analysis, 21(3): 593–630, May 1990. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).